A Complete Bibliography of Publications in the Journal of the American Society for Information Science and Technology

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

31 January 2018 Version 1.05

Title word cross-reference

(∞, n − 1) [QCGB^{+}08]. 1\frac{1}{2} [TD02]. 25 [Wal11b]. 26 [Sch08]. 7 [Coo02a]. ^{2} [RJ08]. _{m} [Sch09a]. A [Sch08]. c_{f} = c/c_{0} [BD09]. E [Bor11, Zha10a]. f [Käk06]. g [ADV10, ADV13, De 11b, De 13, Egg08a, Egg08b, Egg12, Sch08, Sch09b, Sch10, Sch13a, Woe09, Zha10a]. H [Pra12a, Tre12, ADV10, ADV13, AL08, Bar12, BB13, Bor11, Bur09b, Bur13, CM06, De 12, Egg07b, ER08b, Egg08a, Egg08b, ER08a, ELR09, Egg10d, Egg13b, EF10, FR13, GMM09, GG09, GR09, HvdW09, Hov13, Jar11, KLP09, MR08, Opp07, Pil13, Pra11a, Pra12b, RJ08, RY08, Sch08, Van07, WvE12a, Ye11b, Zha10a, ZTY12]. n [AN09, Egg06, ER08b, Sta11]. n + 1 [Lak03]. P [YDS11]. R [Sch08, AJR04a, Whi03a]. w [Wu10].

-core [Sch10]. -gram [AN09]. -grams [ER08b, Sta11]. -index [Bar12, BB13, Bor11, Bur13, CM06, De 11b, De 13, Egg07b, ER08b, Egg08a, Egg08b, ELR09, Egg10d, Egg12, Egg13b, EF10, FR13, GMM09, GG09, GR09, HvdW09, Hov13, Jar11, KLP09, MR08, Opp07, Pil13, Pra11a, Sch08, Sch09b, Sch10, Sch13a, Van07, WvE12a, Woe09, Wu10, Ye11b, Zha10a, Bor11, Pra12a, RJ08, Sch08, Tre12]. -indexes [ADV10, ADV13]. -indices [ER08a, ZTY12]. -overlap [Egg06]. -Rank [YDS11]. -Type [Bur09b, De 12, Pra12b, RJ08, RY08]. -year [Wal11b]. -year-old [Coo02a].

//www.hotelodisej.hr [Ano03d]. //www.hr/iuc [Ano03d]. //www.pedos.hr/lida [Ano03d].

1 [EM07, Haw01, Kel06a, Mar07a, rP08a, RU07, SB07b, SWF^{+}02a, Whi07a]. 10.1002 [Ano12, Ano13b, Ano13c, Ano14]. 10.1002/asi.20981 [Ano14]. 10.1002/asi.21688 [Ano13b]. 10.1002/asi.22620 [Ano12]. 10.1002/asi.22710 [Ano13c]. 101 [Zin09]. 107 [Ano13d]. 113 [Ano06b]. 1183 [Ano08b]. 12th [Ano01d]. 1344 [Ano13a]. 1440 [JP08b]. 15th [Ano03f]. 1677 [Tan07a]. 1840 [WCC10b]. 1852 [Ano08c]. 1929 [Bla13]. 1945 [Cer10]. 1976 [Smi04]. 1986 [Ano10].

2 [AHH06b, HLC03, Kel06b, Mar07b, rP08b, SB07c, UR07, Whi07b, WFE^{+}02]. 2.0-ness [CHH09]. 2003 [Ano03c, Ano03h]. 2011/2012 [BMD13, WCMK^{+}12]. 20th [CSL03, CSL04]. 21st [Gar01b, Lar10b]. 245 [Ano08a]. 2558 [Ano11]. 2D [JG06]. 2nd [Fou01a, Har06a, Hem07, Koe11].

3 [CBL^{+}13]. 346 [AFR07]. 3rd [Jul12].

4 [CBL^{+}13]. 4th [McL10b].

51 [Ano08a]. 55 [Mac04b]. 57 [Ano06b, AFR07, Tan07a]. 59 [Ano08b, Ano08c, JP08b, WCC10b]. 5th [Vil10a].

60 [Ano14]. 61 [Ano10, Ano11, Rob10b]. 63 [Ano12, Ano13b, Ano13c]. 64 [Ano13a]. 658 [Rob10b].

7 [AKW07, BZC07, EM07, Fat07, Jör07, LKK07, Mar07c, RJ07, SB07a].

A. [AFR07, Egg12]. Abbott [Wil06c]. abbreviation [HKKP10]. ABC [YBC^{+}06]. ABC-Clio [YBC^{+}06]. ABCs [Hol04, Sma12]. ability [GL12]. able [BSMD11]. aboutness [Hjø01b]. absence [HC07, War10a]. abstract [HB08]. abstracting [Sag06]. Abstraction [WH10]. Abstracts [Haw01, YBC^{+}06, HB07, HB09, HLC03, LC13, OKG07, ŠKL08, NMD09]. Abstracts/America [YBC^{+}06]. ACA [LV06]. Academic [Pil13, SNRB11, Tie10, Ahn11, BF08, CGY09, Chr03, CC13, FFMPM13, FN12, GRSFV^{+}12b, HLLS07, HLVA07, Hyl03, JCDR^{+}06, JHW^{+}08, KTA12, Lak08, Lim07, LJ13, Luz09, Mat13, Mat09, MB13, NHL^{+}10, NH12, PT05, TW03a, TK08, TS11, VY12, VŽ08, WT08, Cro04, Fou01a]. Academics [Sch03, San08, DWA13]. Academy [YH04]. acceptance [Bro03a, Bro03b, HLC05, Kim06, LCC07, SF08, SZ08]. accepted [Ano08c, BD08]. Access [Gau09, Abb05, Ano01h, Ano01f, Ano04c, Ano13d, BHDK13, Bjö12, BSMD11, CLYL06, DRW08, Dav09b, ESL^{+}10, Enn07b, GZS13, HVP06, HS01, KH13, Kra08, KJ12, Kwo05, LT12, LB13a, MC12, MCRdMA11, Moe07, Muk09, MH04, Nel06, NOR08, PLZA04, Pal05, PR07, Pet13, Pet02, RHWB11, SB12a, SB12b, SH07b, SH07c, Spe01, TrP13, Wal07, WL11a, WJR05, Whi06, WG10, Xia10, Mar07c, Yeo11]. Accessibility [Par10, YH04, DLM09, LDM12, PZ05, YCE10]. accessible [Enn01a]. Accessing [San02]. accommodate [CMD05]. according [FK06, GRSFV^{+}12b, GRSFV12a, GS11, HNB^{+}09, LIL09, VŽ08]. Accounting [Ley12, Sch09a]. accreditation [Cro11a]. accrued [SSPA10]. Accuracy [GP10, ADV13, BSK13, FF04, LSJR11, LS11d, LATH08, WM06]. accurate [KB11, The08a]. accurately [BK10]. achievement [LGM^{+}09, OJ04]. acknowledgment [Ano14, SMAB09]. acknowledgments [CvL12]. ACM [Ano03c]. acquaintance [MPR^{+}03]. acquaintanceship [Pep11]. acquisition [KS10, Vei09]. Across [Hil11, AWBB09, ASQR09, BH11, BGY^{+}13, CKR13, Gen12, Har06b, KP03, LZB13, MBBS02, Mel04, PAR13, TS01, WBJM12]. Act [JBM04, LHLL08, YWC08, Max04]. Acting [RRG11]. action [Hal04a, KY08, MBBS02]. actions [HB11, Hem07]. Active [KLY08, LS05, WP05]. Activist [Kim11b]. activists [Sav07]. activities [ADC10, BP06b, Gor01, HJ04, SJ10, ZATH09, ZS08a]. activity [AGOP06, AKS11, BGY^{+}13, CK08, CF09, JdMAA12, KTL^{+}13, LAL09, MCK08, Sun12, Xu07a, XL07]. actor [SEDC11]. actual [KO12]. ad [JLS13]. adaptable [HM09]. Adapting [BKRN03, KRK^{+}10]. Adaptive [San10, SBT^{+}07, CMFZ01, CL12, LHA07]. added [KdHWA04, SEDC11]. addicts [TL05]. adding [Egg10d, Ley09, LFC01, JG03]. addition [BMD09, Bor11]. Additional [uHG10, Lun09]. address [KRK^{+}10]. Addressing [JW09, RRW05, KH07]. adjusted [SBF13]. administrators [Zac05]. Adolescent [FDY^{+}13, Koh13, Ahn11, Bow10, CL12]. Adoption [YK05, BDS12, DWA13, DLM09, Kim10b, MQH13, NY09, Sun12, Vis05, YSD^{+}08, vRD09]. ads [HTT07]. adult [Xie11a]. adults [Xie11b]. advanced [Hol01, WmcsW09, ESL^{+}10]. Advancement [RKBS11]. Advances [Ahn11, Bor12, Bor13a, Cho12, CC13, Cro10a, Efr11, Fur10, Fur11, Gaz11, IS12, LSZC13, Sha13b, Shi12, Sma12, SML12, TrP13, VW12, Win11, PGL13]. Advancing [Des08a]. Advantage [BdMAM13, NOR08]. advantages [War07d, vR06a]. adverse [CAC10]. advertising [JLS13]. advocacy [Enn04]. aerospace [JJ10]. aesthetic [Reu07]. aesthetics [LBTN09]. affect [Lai09, NH12, RH12, SZ08]. affecting [BDS12, KS08, KO05, jL05, PZ12, YWM10]. affective [DMR06, Yoo10]. affects [LB11a]. affiliates [PBN02]. Affiliation [JZW^{+}11, CM08a, FC01, PBN02]. Affinity [LBS11]. Affinity-driven [LBS11]. affordable [Kwo05]. AffRank [LBS11]. Africa [Cap08, KMD06]. African [MKMM09]. after [HHHR13, WV13, War06, Tol01b]. Afya [MBBS02]. Against [RGLCTSC13, Cui10, GKL09]. Age [Fry10, Van10, Ame12, Ano02e, Ano03d, AFR06, AFR07, BF08, BAB07, BMK03, CvLB10, Egg01, Enn06b, FMP01, Hal01, Him03, HDB07, LLG12, LATH08, RJ08, Vau03, Wag03, Bla10, Sch01, Kaz11, Urq09, Van10, Ago05, Enn03a, Hea05, Sch02, Yak05]. age-dependent [RJ08]. aged [YR08]. agencies [LC05b]. Agency [HHfHC10, HW02]. agenda [Cho12, GRJS01, Lan04, Ste03a, TKB05b]. agent [MPR^{+}03, MPR^{+}05, WCLZ01, YC02]. agents [HCH^{+}05, LPSS06]. Ages [Raw10]. Agewandte [Ano08c]. Aggregate [FS11]. aggregated [Che08, LB05, Ley06, LHS11, LRC13]. Aggregation [OM13, Hen12]. aggregations [PMBdS10]. aggressiveness [GV13]. aging [LAG08]. agree [Hal04a]. Agreements [Col07b]. agricultural [PR10, TWKH13]. ahead [LHB^{+}08, Ros02]. aid [Whi02a]. AIDS [NHCW13, Vei09]. al [BD09, Bur04b]. alcohol [TL05]. Alex [Pon01]. algorithm [CF07c, DML02, JBLOHM10, LSZ04, MG09b, QCGB^{+}08, WP05]. Algorithmic [FS09, PG02, GPI03, RL09]. algorithms [ASABS^{+}10, ASR^{+}10, BCT02, CLMB10, LHHVH09, LPGBdMA03, WZ08, DLH^{+}08]. Alien [Tol01b]. aligning [CLL^{+}07, GZS13]. alignment [OYM09, RMC12]. allocation [HKLG05, SLR^{+}11]. alone [Cro04, Ekb12]. also [Hjø02d]. AltaVista [JSP05]. alterative [Luy11]. Alternative [Bur12, Mer09, FM05, HvdW09, Lak03, LB11b, Kim11b]. Amanda [Jul12]. ambiguities [Ano02e, Sch13d]. Ambiguity [Bur01, MG09b, Boo01]. Ambivalence [For09]. America [CRBPdMAM12, Shi08, YBC^{+}06]. American [Ano06b, Ano08b, Ano08c, Ano10, Ano11, Ano12, Ano13b, Ano13c, Ano14, AFR07, JP08b, Mis09, NMD09, Rob10b, SLR^{+}11, Tan07a, Tie10, WCC10b, Bro07, CR03, GTHZ05, MKMM09, NB06, PM11, Wal11b, Kap09]. among [CH13, CK08, CLHHVH11, GL12, Hei06, HC09, LP10, LS11b, LB11a, LHS11, MS02, MG10, Sav09, Vin11a, YH04, YRXS11, ZC04]. Analogies [War07a]. analyses [Åst07, Bro07, TWF08]. Analysis [Ale03, BdMAM13, CFS05, Cho13b, Cop11, FR01, Lee10, Nic10, NNY10, PR10, Whi02b, WGW12, ZWW09, AJR04a, ASAK04, ACH^{+}04, Ano08c, Ard13, Arg03, Arg06, ASW05, AR13, BL12, BMJARB06, BR12a, BT08, BT05, BNUA04, BJC^{+}07, Ben06, Ben04b, Ben08, BD08, Bor13b, BK10, BBB^{+}08, Bud07, CHE10a, Cam10a, CSCC07, lC01, CPO01, CISH10, CFW^{+}12, CR03, CKF07, CB13, CLHHVH11, CLHHVH12, CvL12, CD06, Cri11, DJWR02, DS03, Dav04a, Dav09b, DS05a, Din07, DO01, DMR06, Egg01, EHD05, FPZAGBdMA05, FCjM12, FS09, FARR09, FMM05a, Fra09, FAC08, Glu01, Gre01b, HM13, HH13, Ham11, HSH10, HWT^{+}13, Hel10, Hen13a, Hjø01c, Hjø02d, Hjø11b, HTdV11, HNJT06, HHvdHdR10, Jac01, JBM04, JHR01, Jär07, JLK09, JSI^{+}04, JLB08, KR08, Kim10b, KBP10, Kno06, KT07b, KT08, KT09]. analysis [KTA12, KW02, KP03, KHC09, KL11, Kup06, LSJR11, Ley05, LZ08a, LBMO11, LO13, LY05, LY06, LKL10, LK10, LLA08, LS11c, LL12, LZG13, LR13, LAL08, Ma02b, MM10, MHL05, MC12, McD07, MBBK07, Moe07, MLO02, NDD^{+}09, OM13, OM08, PPH06, POC09, rPLM09, PZ12, Pet13, PAR13, Pom05b, Rob10a, Rob10b, RFV12, STS12, SS04, SXCC07, SW12, SF07, Sha06a, SLS06, SYL^{+}13, SKM07, SM03, SS06, SZ12, SPH08, SK13, The06, TVF06, TK08, The08c, TB13, TSRGJC^{+}13, TLdR10, TC04, VV12, VGM^{+}11, VBMM11, Wal11b, WCCS13, WG05, WY06, Whi02a, Whi03a, Whi03b, WT08, Win08b, WWZ09, YD09, YBC^{+}06, YC10, YKT03b, ZW01, ZWW^{+}08, ZJS09, ZZL^{+}09, ZATH09, ZQI^{+}10, ZZL12, ZDM13, Zha13, ZS07, ZS08a, ZS08b, ZS11a, vEW08, vEW09, Bro03a, Bro03b, DO01]. Analysis [Fat07, GACCNM^{+}08, MSYD11, NSE^{+}06, Sto11, Zuc06a, The09a]. analytic [CHMS13, KO12, RBLH13, War07d]. analytical [BD05, ESK06, KWR^{+}07]. analytics [DSG^{+}12]. analyze [Bor13b, CvLvR10, DZC^{+}09, DO01]. analyzer [CCR09]. analyzers [KC09]. Analyzing [Alb10a, HH13, KKF^{+}12, LS10a, SP11, ZKSE13, Hub01, LLL09, MMCC04, Whi02c, YC11, Ley04, SLR^{+}11]. anaphora [WL08]. anchor [HBSZ13, WHTS12]. anchor-based [HBSZ13]. ancient [Wei10]. and/or [HKBY^{+}10]. Andy [Enn01b]. Angeles [Buc11]. Angewandte [BD08, NMD09]. animation [HTT07]. animations [LS08]. Ann [Sti01a]. Anna [BM12]. Anne [Gal12b]. annotation [CBS10, Cui12, HSJB12, MMMM^{+}01, OYM09, PCR13]. Annotations [Win08a, CCCJ10, FL01]. Announcement [Ano01c, Ano02a, Ano06a]. Annual [Ano03d, Ano03f, Sch05, Vit08]. anomalies [Eas02, TW03a]. Anomalous [CLB^{+}05]. anonymity [Tsi13]. anonymous [ADD^{+}06]. Answer [Hjø10b, CB13, Him04, OYM09]. answerers [Oh12]. Answering [LJ13, CB13, KWR^{+}07, LWY09, LK06, RFQ^{+}05, WKK^{+}07, WTP12, Yag06]. answers [FF04, Hua13, KO09, RLF02, Sti05]. antecedents [Cho13a, KH09]. Anti [Lak08, HKLG05]. Anti-plagiarism [Lak08]. anti-terrorism [HKLG05]. anticipated [SV05]. Antonio [Day02b]. anxiety [Kra02a, KW02, OJ04, OJ04]. Anxiety-Expectation [OJ04]. any [Muk09]. anytime [Sti05, EM07]. anywhere [Sti05]. appearing [Mac04b]. apples [SK12]. appliances [Han01]. applicable [Gre01b]. applicant [ACMP11]. Application [DWA13, Kos04, Vin11a, DGKS04, Ehi02, Fat07, FMM13, HBSC13, Jör07, LGCA10, LCW06, rP08b, PR10, TR05, TLdR10, Fan10]. Applications [Ano01d, Cho10b, EG12, BZC07, CHE10a, JLCL04, JP02, LV06, LKK07, OCO08, Soa06, TS08, Wil06b, ZZ10, Fou02]. applied [Dai06, KASGAM09, LPGBdMA03, Meg01, Tij10]. Applying [AJBAFHCE08, CM08a, Din11a, Hua13, KS10, Rok12, WLLH09, WS06, YD09, ZZ03, Hov13]. appreciation [Cro10b]. Approach [Ago13, Rok12, AMFiA11, BR12a, BT05, BBMN03, BBMN08, BEK^{+}01, BL11, BK10, Cha09, Che06c, CLH13, CL06, CL09, CL01, CMH10, CdSG^{+}09, CvLB10, DGA11, DZC^{+}09, Dav02, DPV11, DLVÁ12, DK07, ESK06, FPCB04, FDH04, Fra11, FMM13, FCM04, FAC08, GMV13, GLLL11, Has04, HLP04, Hen12, HV01, HASAT^{+}10, Hol08, HRC12b, Iva01, JG05, KTL^{+}13, KHAN08, KHC12, LM01, LBC02, LLL09, LMRB08, LKL10, LL12, LMV^{+}12, MCRdMA11, MAA08, OL13, Pat11, PN09, RS10, Sav07, SRDZ06, SC11, SK13, TYS13, VG06, WGD09, War08, fBWLBC06, WLC09, XR07, Yan01, YL05, YLZ10, Ye11a, Yi10, YR08, ZIB01, NJ07a, Ben06, ZL07d]. approaches [AN09, ACHVH10, Ano04c, Bur05, Dow04, FMM03, FMM05b, HT11, HVP06, Jär07, LB08, LLC12, LW12, O’B07, POMVULMC13, SC06, SC12, Zin07c]. Approaching [CvL12]. Appropriate [vEW08]. Appropriation [Kap09]. appropriations [SHO11]. Approximate [GMA07, Mus05, NBYA03]. April [Ano01e]. Arabic [Abd04, AN09, ASABS^{+}10, ASAK04, AFG02, ABJ^{+}06, AMFiA11, BAT08b, BEES10, CC09, Duw06, DARK09, FR08, HASAT^{+}10, KASGAM09, MAR04, Mus05, MAA08, RSMVULPO11, SR09]. Aramaic [HKKP10]. arbitrary [FR05, KZ01]. architect [MW10]. architects [BIB07]. architectural [JG06]. Architecture [Ano01g, Bur11, CJJ^{+}02, Dil02, HW02, Hav02, HDG^{+}07, LBC02, Lat02, Ros02, Tom02, UZ04]. architectures [DHW13, Sri07, SML12]. Archival [Cox08, YK05, Sin12, Sti03]. archive [Hjø09b, HHvdHdR10]. archived [KPEO10]. Archives [Cox09, McL10a, PSO06, TD02, Tar08]. archiving [Col07b, Kim10a]. area [NMPBA11]. areas [GRSFV11, GACCNM^{+}08, MBG03, WL11a]. aren’t [HW02]. argue [Hal04a]. argument [Sav12, WS11]. argumentation [PS03]. Arguments [Hjø03, Egg07c]. Arizona [DZC^{+}09]. ARL [CGY09]. Arms [Hjø01a]. ARR [Hen12]. array [CJA^{+}04]. art [Ago04b, CS12, lC01, Nay05b, Nay07b, TrP13, Rob01, LS10b]. Arthur [Hjø02a]. Article [FR05, HNJT06, BSP10, CF06, ESF^{+}08, Gel09, KEA^{+}05b, Mac04b, MSYD11, ST08, SB12b, YD10]. articles [Åst07, ARC11, Ano12, Bjö12, BD07b, Cam10b, CKT06, CHMS13, Fig06, FW09, GLAA07, GACCNM^{+}08, Hen13b, JD12, Ler13, LS11a, LT10, NOR08, OIM^{+}11, SM13, SKM07, Wal11b, Zha12]. artifacts [Ekb09, Hen13a, PMBdS10]. Artificial [Rok12, YKCL12, Gar01b, RM13, RKBS11]. artistic [Pal08]. Artists [Rin09]. Arts [AŞT06, Ben01a, Kar06, SHS09, Zac05, LS10b, LHS11]. arts-based [SHS09]. ArXiv [Moe07, uHG09, uHG10]. asi.20981 [Ano14]. asi.21688 [Ano13b]. asi.22620 [Ano12]. asi.22710 [Ano13c]. Asia [Bol07, AH05b, Ano02b, Ano03b, JLCL04]. ASIS [Mac06]. ASIS&T [Mac06, SL10]. ASIST [Buc11, VH05]. Ask [JHR01, LY11, Tan07a, XTY06, Bil12]. Ask-an-expert [JHR01]. aspect [Liu13a]. aspects [DMR06]. Asserting [Enn04]. assess [BKRN03, Oak09, SPH08, TRJ05, Wu10]. assessed [Bor13a, KTR10, SV05]. Assessing [ADV13, Col07a, KT08, KTR11, McC08, McC12, MS09b, Reu07, ST09, CC04, SEDC11]. Assessment [DeN09, ZC01, AD09, ADR11, BdRB^{+}13, BLA07, CvLB10, GPG^{+}11, HLC11, Jar11, KKF^{+}12, LZB13, Liu10, LJLlW08, NSLD13, Smi01, SGTS07, Tre12, VS05, WBY^{+}09, Whi02c, WL06b]. assessments [BL13, DGA11, KJ02, LKJ08, VvdHH^{+}13, WCC10a]. assigned [HRC12a, LS12]. assigning [Los13]. assignment [CMH10, DDO11]. assignments [MM01]. assist [CLH13, HH01]. assistance [JM05]. assistant [LPSS06, WH10]. assistants [KAB05]. assisted [BB03, DDO11, LJLlW08]. associated [NHCW13]. Associating [MHL05]. Association [Ano03f, AMFiA11, CL01, Egg10b, NSQI03, RCT05, SWT04, SY07, YHHL12, RKBS11, Wal11b]. associations [NJ07a, NJ07b, VG07a, VC08, Yoo12]. associative [vdEvMK^{+}04]. assumptions [Ma12a]. assurance [LMV^{+}12]. astronomers [Bro01b, TKB^{+}05a]. Astrophysics [KEA^{+}05a]. asymmetric [CL09]. asymmetries [Ano02e]. ATDS [ESL^{+}10]. Atherton [Pet02]. Atlas [Nea11]. atmospheric [BSMD11]. attachedness [ZQI^{+}10]. attachedness-based [ZQI^{+}10]. attachment [Cho13a, Mil10a, RRG11]. attitude [TWW^{+}06, YR02]. attitudes [FMM05b, Jan02, Lus07, MHR13, Xia10, ZS09]. attribute [Has04]. attributed [Hen13b]. attributes [Cho13b, NMPBA11, RR12, YKL07]. attribution [FT12, KSA09, PPN07, SBF13, Sta09, SK02b]. audience [OCD08, WvE10, XZW13, ZS08c]. audiences [BL12]. audio [BNUA04, DMR06, TC04]. audiovisual [HHvdHdR10, Spe01]. auditory [JG03]. augmented [Glu01, RPYZ07]. Augmenting [CHE10a]. authentication [LF03]. Author [AJR04a, Dav09b, KR01, SZ12, WJD^{+}13, Whi03a, Zuc06a, AK13, Aks08, Ben04b, Bir06, CM08a, DGA11, Dan11, Din11a, Din11b, Fei11, LKBJ12, Ley05, Ley08b, LW12, McC08, NSJ13, OIM^{+}11, OM08, RZ04, RR12, SGJ13, TWSS05, yT04, Whi03b, YKT03b, ZS08a, ZS08b, vEW08, MM01]. author-based [SZ12]. Author-choice [Dav09b]. Author-level [WJD^{+}13]. Authority [Spo08, FC01, He09, HWJ10, PK11a, Rie02, TPLC05]. Authors [Ano13b, Whi01, AŞT06, CH13, Cro09e, DYFC09, ER08a, FMP01, GLAA07, Hu09, Mar09a, MG07, OIM^{+}11, Pal08, SL10, TVY12, WJD^{+}13, ZS11a, GSD12, Min10]. authorship [Egg08b, HS02, Hu09, IS06, KSA09, SBF13, Sta09, ZLCH06, ASR^{+}10]. Automated [Bre05, NKS^{+}06, SXCC07, YEV08, BBB^{+}08, JM05, LCR^{+}05, Nic03a, SPR09, TR05, VGM^{+}11, Wan09]. Automatic [FMCL07, Gre01c, JP02, LY05, LKK07, LVdA^{+}04, OKG07, SCA12, Tse01, Tse02, TK06, Wan06, YL03a, YL03b, AF12, Ano13a, CMH10, CH07, FCjM12, HNB^{+}09, Kar03, KK04, Kos04, KT08, LWY01, MW08, MO07, OCO08, WY06, WH10, Wil10b, WL13, Mel04, Lin02]. automatically [OTP13, RNLdL01, WCLZ01]. Automating [GS03]. autopoietic [Riv13]. available [LB13a]. average [Sch10]. avoid [Sch12, Sch13d]. avoidance [BDBA13]. Award [RKBS11]. aware [Fei11, SC08, Ma07]. Awareness [Ano10, SM10, Sha13a, SV03, SM05b, YCE10].

back [WK13]. background [BIZGMS10, JBCJ06, Su03a, Kel06a, Mos04]. Backlink [CSCC07]. Backstage [Ano14, SMAB09]. baker [WS11]. balance [CH13, CM08b]. balanced [LMRB08]. Balsamo [Gal12b]. bandwagon [Ano12, FS11, JD12]. Bang [TSV12]. Banking [CS02]. banks [DLM09]. banner [HTT07]. baring [FMMF13]. Barriers [KJ12, Kim10a]. base [Che06c, Ley04]. based [Ack02a, AXP11, Ago02, ALW10, ACHVH10, Ano06b, Ben04a, Bis11, BLA07, BCT06, BR12b, BV08, BP10b, BL11, BMD13, Bro07, CL11, CCG^{+}06, dAP13, CLMB10, CC01a, CC02, CF07c, CL08, CL09, Chu09, CL01, CMH10, Coo01, CFN^{+}10, DML02, Din11b, DGKS04, Efr05, ER02, Egg09a, FFPW04, FDH04, HKBY^{+}10, Har08b, HMdR11, HVPLHP06, Hjø11b, HBSZ13, HW13, HRC12b, HCO03, HFC10, HMFW12, Hua13, HRK^{+}06, HNB^{+}09, HC07, Iye01, JP08a, JP08b, JT10, JG06, KdHWA04, Käk06, KWS07, KW01, Kim05, Kim06, Kim11a, Kle08, KT07a, Kos05, KAB05, KSS03, KJ12, KO05, LWW01, LCR^{+}05, LF03, LP03, LB08, LWH11, LSJR11, LR02, LCB^{+}07, LC05a, LMRB08, LKBJ12, LR09, Ley09, LO10, LS13, LLA05, LZLC08, LC05b, Liu10, LS11c, LH11, LLC12, LPGBdMA03, LS05, LW12, LAL08, Ma02a, MSWK11, Mar09a]. based [MMBF09, MvL12, NSQI03, NMD09, NDD^{+}09, Nic03a, NS07, NRD11, OLLL10, ÖAU^{+}04, PN09, PRNZ^{+}11, PS03, Pom05a, QCGB^{+}08, RL09, RMC12, RM13, Ros08, SXCC07, SRDZ06, SYL^{+}13, SYNC08, Sma12, SHS09, SJTB13, SPR09, SZ12, SL06, SBNB11, SST06, TJLJ07, TS09, The02, TH03, TB13, TMPP13, TK06, VTD10, WS13, WA03, WKdJvdBV01, WTF08, Wil06a, WL08, WL09, XR07, XJ10, YL03b, YL05, YL09, YLZ10, YWC08, Yi08, ZIB01, ZI01, Zha01, ZWW^{+}08, ZATH09, ZQI^{+}10, vRC11, vdEvMK^{+}04, PK11b, Che10b, Ruc10]. bases [Has03a]. Basic [Rou12, BI04, Hei06, RI08, Szo11, Tij10, RJ08]. basics [Mar09b]. basis [Ano08a, LdMAGB10, LS10b, LHS11, LRC13, Los01, Wil00]. Bates [Bat08a, Bat11, Hjø09c]. Bawden [Ave13]. Bayes [LMRB08]. Bayesian [LM01, LC06, SBF13, dCFLH03]. Bazillion [Fou01a]. be [Bir06, BL11, Bor13a, Buc12, Bur02b, Dan11, Eas02, Hen13b, HB11, Hjø02d, KTR10, Lak08, Ley06, Ley07b, LZB13, WvE07]. bearing [KW01]. beautiful [Ost08]. become [LTO11]. behavior [All11, AKS11, AE10, BD10, Bar08a, BT05, BR13, Ben04a, BR08, BBC01, Bur03b, CCC10, Coo02a, Cot02, DC01b, DS03, Dav04a, EAPL05, Fos04, FN12, GKE11, HHWD05, HLR09, HLVA07, HHNK12, HSCL07, HHvdHdR10, Joh07, KRS13, KA02, Kim09a, Koh13, KC12, Lan06, LSB13, LB10, LCW06, MMC05, MHL05, MT03, MA12c, MEH07, NHL^{+}10, NH12, POC09, RB08, Ros06, RU07, RLvR03, STS12, SKM01, SR06, SC06, Ste08, SL13, TDC13, TD02, Tra08, UR07, VW12, WRB^{+}09, Whi02d, WG10, Xu07a, YBC^{+}06, Zac05, ZZZR04, vT08, Sar07b, Ago13, Jul12, Ago07a, Wan07]. Behavioral [LX13, YD13]. Behaviors [CAC10, BMK03, Bil01, Bil02, BP06a, KKGB11, MBC08, MW10, Rob10a, Rob10b, Xia10]. Behaviour [Efr12, Ing12, Jul12]. beholder [OA12]. belief [KH09]. Believe [WB02]. benchmark [ADV10]. Benchmarks [Har05]. benefit [Cam10a]. benefits [BOGdMA13, Enn01b]. Benford [EG12]. Benjamin [Sch01, Ham11]. Bergman [Han01]. Besselaar [Ley03b]. Best [Bea06, Dav06, Gre06, Hen06, JBCJ06, Kal06, KSM06, Moc06, RV06, Ros06, Rot06, Sac06, VR06c, Wil06a, Ano06b, HRK^{+}06, VR06c]. better [Ben07, BMD08, Dav08, Eas02, HW02, KB06, LCB^{+}07, LC05c, LMV^{+}12, ZS07]. between [ADC10, AG02, BR11a, Bur05, Cas08, Egg01, Egg05b, Egg09a, ELR09, EL09, Egg10b, EF10, FFMPM13, Fig06, GKL09, GMM09, GW12, Has03a, LG10b, LTO11, Li09b, LB10, LC05b, LC13, LLG12, Moe05, NHCW13, Pep11, RFV12, Ror07, RCD08, RR12, SB07b, SB07c, SKS11, The02, TH03, WvE10, War07a, YS11, YKL07, YRXS11, ZL07a]. Betweenness [Ley07a]. Beyond [Har02, Koe04a, Koe04b, Col07a, Han01, HBSZ13, Tom01a, XC06, KC07, Mil10a, Sma12, Wal06, Him03]. Bhavani [Cul01]. Bias [Kva10, LSZC13, BIKLY09, BR11a, CH11, FR13, HWJ10, SRV07]. BIBE [HML02]. Bibliographic [VS03, WW03, WL11a, BL09, BK10, CdSG^{+}09, CFN^{+}10, HW03, Mor05, NSLD13, Pet13, RKBS11, Tan05, Tan06, Tan07c, YD12, ZS08a, Sha06a]. bibliographic-coupling [ZS08a]. bibliographies [Egg02b]. Bibliometric [LRR12, VBMM11, vR08a, Ano12, ACGL09, AR13, BBB04, BMD09, CH12, CvLB10, FARR09, Fra09, FMM13, FN08, Iva01, JLK09, JCDR^{+}06, JD12, Kup06, KEA^{+}05b, LSC12, Ley09, LLA08, Mar09a, Muk09, NKLC09, SYL^{+}13, Sin11, TSRGJC^{+}13, WS13, vEWDvdB10, vR06a, vR06b, AŞT06, NGJ03, SL12]. bibliometricians [CM11]. Bibliometrics [Hov13, Bor13b, DGA11, De 12, FMM13, KdRH^{+}01, LG10a, LYJ^{+}10, Mar11, Pra12b, Whi07a, Whi07b, WOE^{+}13, Nic10]. Bibliomining [Nic03a]. bidirectional [AFG02]. bids [CP13]. Big [Pra10, TSV12, WBB10a, WBB10b]. bigrams [HASAT^{+}10]. bilingual [CJJ^{+}02, LWY09, POMVULMC13, TK06, YWC08]. Bing [Bil12]. binning [Mil10b]. Biobibliometric [SC12]. Bioinformatics [RW05, BT05, BC05, Fen05, Hem05, HLB05, HAT12, JLK09, MD05, SF08, TR05, Ano02f]. biological [ASW05, Dav09b, HCH^{+}05, Nah07, RCT05, Sha06a, HML02]. biologists [SF08]. biology [Dai06, MD05, McC09]. Biomed [KSF04]. biomedical [HSSV07, HNB^{+}09, KR08, LC05a, LZLC08, LH11, Pet13]. biomedicine [BD07b, BMD08]. biosystematics [CBS10]. biotechnology [LH01, RW05]. bioterrorism [DZC^{+}09]. biplot [TSRGJC^{+}13]. Birger [Bat11]. Bish [Hol01]. bit [KMK03]. bits [Bla11, Joh01]. black [VV12]. blends [LGCA10]. Blind [ESF^{+}08]. Blobgects [SBBF09]. Block [WM07, TCW11]. blog [KCC12, YHWL13]. Blogging [CPP09, CK08, FMMF13, YZLC13, CPP09, Edw07]. blogs [FMMF13]. Blown [Joh01]. Blueprint [DeN09]. boards [Cab12]. body [NHCW13]. bonding [CLD^{+}13]. Book [Ack02a, Ack02b, Ack03, Afi03, Ago01, Ago04b, Ago04a, Ago05, Ago06a, Ago06b, Ago07a, Ago07b, Ago08, Ago13, Ahm07, And02, Ano03a, Ave13, Awa06, Bad03, BI13, Bar01a, Bar01b, Baw10, Ben01a, Ben06, Ben11, Bia10a, Bia10b, Bla13, Bla10, Bol07, Bol08, Bor09, Bra09, Bra11, Bri10, Bro02, Buc10, Bur09a, Cal08, Cas03, Cer10, Cha03, Cha13, Che06a, Che03a, Che04, Che09, Che10b, Cho09b, Cho10b, Coo07, Cop12, Cox09, Cox10, Cul01, Cul03, Cul04, Cul06, Dav09a, Dav11, Day02a, De 11a, DeN09, Des07, Des08a, Des08b, Det11, Dia02, Dim09, Din12, Dot13, Dum01b, Dum01a, Edw07, Efr12, Egg02a, Egg07a, Ekb12, Els09, Enn01a, Enn01b, Enn02b, Enn02a, Enn03a, Enn03c, Enn03b, Enn03d, Enn04, Enn06b, Enn06a, Enn07a, Enn07b, Enn08a, Enn08b, Esc10b, Esc10a, Ess10, Exn02]. Book [Fah02, Fan10, Fer04, Fle06a, For09, For13, Fos01, Fou01b, Fou01a, Fou02, Fou03, Fou12, Fry10, Gaf07, Gal06, Gal12a, Gal12b, Gaz07, Gaz09, Glu08, Gon08, Gra05, Gru07, Gua11, Hai12, Hal01, Hal10, Han01, Har09, Har04, Har05, Har06a, Har07b, Har08a, Has03b, Has09, Hay07, Hea05, Hem07, Hil11, Him03, Him04, Hjø01a, Hjø02a, Ho08, Hol01, Hol04, Hut08a, Hut08b, Ing12, Ivo03, Jam10, Jap07, Jap08, Joh01, Jon02, Jul12, Kal01, Kal05, Kan04, Kap09, Kaz11, Kim11b, Koc06, Koe03, Koe04a, Koe04b, Koe11, Kos09b, Kva10, Lap07, Lar10a, Lar10c, Lar10b, Law02, Les11, Lin02, Lin03, Lip13, Luk09, Ma07, McI10, McL10a, McL10b, Mea10, Mer09, Mil12, Mis09, Mit08, Mon04, MC03b, Mos09, Mue10, Nay04, Nay05b, Nay05a, Nay06, Nay07b, Nay07a]. Book [Nea08, Nea11, Nel06, Nic03b, Nic10, O’B07, O’B12, Opp09, Opp10, Orm01, Orm07, Ost08, Par10, Pee12, Pet12, Pet02, Pet03, Pet07a, Pet07b, Pon01, Pre09, Raw10, Ren04, Reu05, Rho04, Row09, Ruc10, Rya09, Sag06, San10, San02, Sat08, Sav03, Saw10, Sch03, Sch04a, Sch04b, Sch01, Sch02, Sch05, Sha08, Sha06b, She10, Sim07, Smi04, Soa06, Son10, Spe01, Sta10, Sti01b, Sti01a, Sti02, Sti03, Sti05, Su06, Sun06, Swa11, Tad06, Tar08, The09a, Tho09, Tie10, Tol01b, Tom01a, Tom05a, Tom05b, Tom01b, Tom04, Tom06, Tra09, Urq09, Van03, Van10, Vei06a, VG07b, VG07a, VT13, Vil10b, Vil10a, Vit08, Wad01, Wal04a, Wal02, Wan07, War02, War04, War05, War06, War07b, War09a, War10b, War09b, War10d, Wes07a, Whe03, Whi06, Whi02c]. Book [Wil05, Wil10a, Wil06b, Win09, Wor09, Wri10, Yak05, Yan10, Yeo11, ZL07d, ZL07c, ZL07b, Zin09, ZvL11, vD10, Day03, Har06b, KT09, Reu07, WBY^{+}09, Bia10b, GPG13, De 11a, Him04]. bookmarking [AdG12]. bookmarks [GD01, GD03]. Books [Bia10b, DWA13, KTR11, MQH13, VE13, KTR11, Rya09]. bookselling [BBC01]. bookstore [Ago06a]. Boolean [LHHVH09]. boost [DLA^{+}10]. boosting [YHL11]. bootstrapping [LKBJ12]. borrowing [LR08]. both [Rou12]. bottom [Tom04]. boundaries [CKR13, Luy12]. boundary [Huv11]. Bounded [Ago02]. Bowling [Cro04]. box [TD02, VV12]. boxes [ST13]. brains [Ago08]. Brand [JZS09]. Braun [Fou01a]. Braunstein [Buc13]. breakdown [KR01]. Breast [MMBF09, Has04, Hjø11a]. Brian [Cro10b, Tom01a]. bridge [FN11]. Bridging [Ma07, CLD^{+}13, Sav06]. Brief [Aks08, Ard13, AJBAFHCE08, AD04, BBC09, Bat08a, BMK03, Ben07, BD07a, BMD09, BBB^{+}08, Bur01, Bur06a, Bur07, Bur09b, Bur12, CH13, Cam10b, Cam10a, CS02, Cro04, CM06, CM08a, Cro08, DC01a, Dav08, De 11b, DA05, DS10a, DG12, Egg07b, ER08b, Egg10a, Egg10b, EG12, Egg12, FS07, GMM09, GACCNM^{+}08, Har08b, Hjø01b, HC02, HW05, Hu09, HWD03, Jam08a, Jör04, KH13, LGA09, LG10a, Ley05, LZ08a, LP09, LO10, LO13, LSZ04, LC05c, NY08, NY09, OW04, Opp07, Pra11b, Pra11a, PT05, Rok12, RM13, Rou06, Rou07, RY08, Rou12, SD01, San08, Sch13a, Sch13b, Sch13c, Sch13d, Sto06, TS09, TWKH13, TWF08, Tol09, Van12, Van07, VE13, VBMM11, WvE07, WSJS01, Xia10, Ye11a, Ye11b, Zha10a, ZZL12, ZTG13, ZS08c, vdB03a]. British [Ben01b, Opp07]. Broad [Hei06, TPF06]. broader [MHM12]. broadness [ER08b]. Brownson [Jay12]. browser [PLZA04, Sre02]. browsers [Lak07, MMMM^{+}01]. Browsing [CC09, Gor01, HML02, NSQI03, Tan07b, Bod06, CZH^{+}04, GRS^{+}07, HDB07, JTDG13, San02, Hjø11c]. Browsing-based [NSQI03]. Buckley [Ano11]. Buckminsterfullerenes [HHHR13]. build [SPR09]. Building [Ago06a, Cho12, Har08a, KHAN08, LK06, QH06, Cal08, CS06b, WCLZ01, BD10, Orm01]. Bulica [Ano03d]. Burrell [YKT03a]. Burton [SK02b]. Business [CZH^{+}04, CCR09, Det11, AHK08, Bar01a, BDS12, CSCC07, DJW^{+}11, HvdW09, MMCC04, Pon01, QZCC06, SDJB12, VTD10, VY12, WL06b, Mis09]. business-to-business [WL06b]. Butcher [WS11]. BVideoQA [LWY09]. bypassing [BDBA13].

C [Ano06b, Ano10, Ano13b, Sti01a, Tan07a]. caching [OAU11]. Cai [Ano11]. CAL [LJLlW08]. calculated [FR05]. calculation [Sch13b, WS13]. Call [Ano01a, Ano01b, Ano01g, Ano01h, Ano01d, Ano01e, Ano01f, Ano02c, Ano02b, Ano02e, Ano02d, Ano03d, Ano03c, Ano03e, Ano03h, Ano03f, Ano03b, Ano03g, Ano04a, Ano04c, Ano04b, Ano05, Arg03, Hau05, Kra04]. Cambodia [Luy13]. cameras [SHO11]. campaigns [JLS13, PB01]. campus [BGY^{+}13, HZ12]. campuses [BGY^{+}13]. Can [Ago04a, CL12, Dan11, DHKW10, KTR10, Ley06, Lus07, RMC12, ZS07, BL11, Bor13a, Buc12, DP06, Hen13b, HB11, LZB13]. Cancer [MMBF09, CK08, Has04, Hjø11a]. candlestick [WS11]. Canonicity [Cro13a]. capabilities [KKF^{+}12]. capability [HLP13, PPH06]. capacity [TWW^{+}06]. capital [BPT07, CS02, Day02b, Gal06, Joh07, Kwo05, Pon01]. capitalism [Day02b, Day02a]. Capitalizing [CP13]. Capturing [GS11]. Card [Cer10]. Cards [Bla13]. care [KMD06, NS07, OMS^{+}08, PGR^{+}13, SSM^{+}08a, SSM^{+}08b]. Career [Mer09]. carers [Hep04]. Caribbean [CRBPdMAM12]. Carlo [Ano03h]. Carr [Bar01a]. carries [FN11]. Case [Bia10b, Bol07, Enn01a, Jul12, AB09, AJBAFHCE08, BR12a, BI06, BIP09, BAB07, BPT07, BGY^{+}13, BLA07, Bud07, CCMM02, CCQ^{+}08, Coo02a, DLM09, Egg08b, EGR11, Enn07b, FPCB04, GP10, GRSFV^{+}12b, Glu01, Has03a, Has04, HH07, Hjø11a, HBT^{+}09, IS06, KMD06, KK08, Kim09b, LM11, LH01, LP09, LR10, LDM12, Luy13, MBG^{+}07, Mar02, McC11, Mil10b, MLO02, MK13, NHCW13, NMD09, NCRJ13, NY08, OM13, PN06, PPH06, PV03, PBS^{+}04, RM13, RFV12, RS10, RPH^{+}12, RJ08, SBT^{+}07, SG11, Sch09a, Shi08, TPLC05, Ten12, TPF06, TVF06, VC09, WT10, WHP10, ZQI^{+}10, BH11, JJ10, The08a]. cases [APSA12, Hal10]. cast [CSL03]. catalog [CC01b, Coo01, KC12, RH12]. Cataloger [Šau04]. catalogers [Tan07c]. Cataloging [rPLM09, Spe01, Tan04]. catalogs [SBBF09, Whi06, WG10, Bla13]. cataloguing [Tse01]. catastrophic [Wee07]. categorical [Gre01b]. categorically [CS04]. categories [ASR^{+}10, BW05, Coo02b, GRSFV12a, LR09, MG10, NKLC09]. Categorization [FJRNV^{+}06, CFCZ01, CMH10, CHdS^{+}11, DRM^{+}04, Duw06, DARK09, GS03, GBK05, HNB^{+}09, PCY02, QCL^{+}12, RKO10, RI08, WLHO11, vRC11]. Categorization-driven [FJRNV^{+}06]. categorized [BJC^{+}07]. categorizing [RNLdL01, SSB01]. category [LB13a, LWH11, RI08, YLW10, YLW10]. Category-tree [YLW10]. causality [SZ08]. causes [KBG12]. cautionary [Bur04b]. Caveats [Ley08a, McC12]. CBizPort [CZH^{+}04]. CCA [ZDM13]. CD [SD03]. CD-ROM [SD03]. cell [ZS11a]. Celtic [ÖL06]. census [PGFMP13]. Center [HPGG^{+}07, RW05]. Centered [Kos09b, Erc06, HML02, Pat11, PBS^{+}04, SR06, TWF08, WEA^{+}12, ZL08, ZL07d]. centers [SS10]. central [KSF04, BLA07]. centrality [Ley07a, YD09, ZZ10]. centre [Ano03d]. centric [KR08]. Centuries [Rya09]. Century [Lar10b, CSL03, CSL04, NB06, Gar01b, Har04, Ago08]. ceremony [Cro13e]. certification [Lak08]. chain [FN13]. chains [NLC06]. challenge [Day02b, Shi12]. challenges [Lip03, LHB^{+}08, PNHG07, PAR13, War07b]. Change [Bla10, HHNK12, Koe02, NY08, NY09, NS12, PN06, Pal08, For09, Tom01b]. Changes [Åst07, KGL12, HLC11, LS08, LX13, PK11a, WT10]. Changing [Cro09b, JBCJ06, Vau03, Bro03a, CJA^{+}04, hChC05, Cro09e, Mar07d, MHR13, Ste03b, TS09]. channels [Xie11a]. chaos [Sha07]. Chapter [Buc11]. character [Cui10, KDL02, LZ08a]. Characteristics [Ano12, hHwC08, JSI^{+}04, KR08, KC13, Pet13, AŞT06, ACMP11, BR11a, BV08, CBvLvR09, Cus13, CTS04, Oh12, RLvR02, The05, yT04, VŽ08, ZC01, ZYH13, vR08b, JD12]. Characterization [Vin04, HAT12, JG06]. characterize [CCS11, HS01]. Characterizing [CHH09, HSCL07, HJ13, KWS07, TP07, NBG11]. CharaParser [Cui12]. charges [SB12b]. Charles [For13, Fou12]. charting [GBMH^{+}09, GD07]. charts [ZLB11]. chat [KO12, Pom05a, Rad06, Mon04]. chat-based [Pom05a]. chats [EZ08]. checking [CW04, OW04]. Chemical [Bro07, NMD09, Bro03b, NMD09]. Chemie [Ano08c, BD08, NMD09]. chemistry [BD09, BSMD11, BD05, Chr03, CSL04, DS03, OM08, Vau03, Men03]. chemists [Bro07, Dav04a]. Chen [Ano13c]. child [God07, PSWL13]. Children [BIB07, Lar10a, Abb05, Bil12, Coo02a, Coo02b, HDB07, LBR02, LBC02, Reu05, Bil01, Bil02, BW05, FDB^{+}12, JBM04, LBTN09, Reu07]. China [GKL09, JLR04, LKZC11, SLS06, TH13, ZL07a, Mos09]. Chinese [Cha11, CFY07, CKH^{+}03, Che06c, CZH^{+}04, GTHZ05, KDL02, LWW01, LWY09, LB05, LZ08a, LWY01, QZCC06, Tse02, WY07, WL08, YC02, YL03a, YL03b, YL05, YLZ10, ZZL^{+}09, ZSL10]. Chinese-language [Cha11]. Ching [Ano13c]. Ching-Yi [Ano13c]. choice [Ano13d, BBI12, CF09, Dav09b, SB12a]. choices [MM01]. chosen [SL10].

Christopher [Gal12a]. chronicling [LSC12]. CIKM [Ano03c]. CIO [Hut08b]. CIPA [JBM04]. circuits [Day02a]. circumvention [EHD05]. Citation [Aks06, Ame12, Ano13c, BBRdMA12, BL09, GC07, GPG13, Ham11, HAT12, KdRH^{+}01, LKBJ12, LT09b, LB05, LdMAGB10, LS10b, LHS11, LO13, MR08, Nic10, SF07, Wei10, WHC12, ZDM13, ZL07a, ALW10, APSA12, Ano08c, Ard13, BF08, Ben07, Ben08, BSP10, BD07b, BD08, BD09, Bor13b, BMD13, BB03, BK10, BSK13, BHC06, Bro03b, Bur03b, CKT06, CPO01, Che08, Che12, CBvLvR09, DC01b, Dav08, Dav09b, DT13, DYFC09, Din11a, Egg01, ESF^{+}08, Fra12, FMM13, FN11, GP10, GRZ12, HM13, uHG09, uHG10, Hel10, HRC12b, Hyl03, KGL12, Kno06, KTR11, Ley06, Ley07c, Ley08a, Ley08b, Ley09, LS11b, LB11a, LBMO11, LC13, LZG13, MM10, MBC10, McC09, McD07, MvL12, MY07, MS09b, Moe07, Moe11, NMD09, NOR08, PK11a, Rou06, SG11, Sch12, SKM07]. citation [SKTM09, SKS11, SKS12, Sin11, SH07b, SZ12, TS11, VS05, Wal11b, WvEvR12, WWN04, Whi02b, WT08, YD10, YS11, YD12, ZS07, ZS11a, ZS08c, ZLB11, vEW08, vR08c, vR08b, Bro07, FN13, KT09, SS04]. Citation-based [LKBJ12, ALW10, BMD13, Ley09]. citation-word [ZLB11]. Citations [ZL07a, ARC11, ACMP11, Ben07, Cam10b, Cam10a, Cam11, Che03a, Col05, CdSG^{+}09, CFN^{+}10, EF10, FR13, FR01, GMM09, KAB09, KT07b, KT07c, LGA09, Ler13, LO10, LS11b, LB11a, LHS11, LRC13, LR10, LLG12, LT10, Moe05, OA12, PZ12, Riv13, RZ04, Sch10, Sel05, Sma03, SSPA10, TSW12, VS03, Vin11a, GC07]. cite [CM11]. cited [ARPS11, Bor13b, Bur02b, CA07, KT07c, KTA12, LT08, MM10, Sch13c, Wu10]. citedness [HW05]. citee [Har08b]. Citer [Har08b, ALW10, RM13]. citer-based [RM13]. citers [Whi01]. CiteSpace [Che06b]. CiteULike [LS12]. cities [BL11, LP10, Sto11]. citing [FN12, LC13, SR07, TK08, WdV04]. Citizenship [vD10, Gon08, Wri10]. civil [Nay06]. claims [Nic09]. clarification [Alm13]. clarity [PBS^{+}04]. Class [BdMAM13, CN07, Mos09]. classes [BSMD11, Bor13b, PGSH06]. Classic [Cul01]. Classification [CC11, Che08, LB06, Zin07a, ADC12, AF12, AWC^{+}07, AMFiA11, BBB^{+}08, CCG^{+}06, dAP13, CLL^{+}07, DG12, FS10, GD01, HFC10, HC07, KASGAM09, Kar07, KT07c, Liu09, Liu10, LS11c, Liu13a, MH11, MSWK11, ME03, MS09a, MG09a, MG10, MPR^{+}03, POMVULMC13, QCL^{+}12, RKO10, RZ04, SLN03, Ten12, TH03, Tij10, WvE12b, YEV08, YLL06, ZLCH06, ZS11b, HKBY^{+}10, Wan09, Hjø02a]. classifications [FP04, Ley08c, RL09]. classificatory [CvLB10, Ley02, SS13]. classified [Lar09, Ley06]. classifier [GCR09, LMRB08, SBF13]. classify [FK06]. Classifying [OCJ12, CCR09, JJBC01, PCBL04, Su06]. classrooms [PGSH06]. cleansing [LZC^{+}07]. Clearing [Day05]. CLEF [PB01]. clickthrough [SBA12, ZJS09]. Client [FG03, LF03]. Client-side [FG03]. clinical [DLVÁ12, GPG^{+}11, Hua13, McK04]. Clio [YBC^{+}06]. closing [Jam08a]. closure [TTT12]. clothing [Gup09]. clumping [BKRN03]. Cluster [LC05c, Fra09, HMdR11, KR08, LS05, SRDZ06, Tol01a, WWZ09]. cluster-based [LS05, SRDZ06]. clustering [BYHM07, BC09, BSK13, CL11, CDLY11, CMRS09, CJB^{+}07, CC11, CC01a, DLH^{+}08, Kis10, KB05, LYJ^{+}10, Men04, Mor05, PT05, RPG^{+}08, SPR09, TLM06, ZWW09, ZA13]. clustering-based [SPR09]. clusters [CISH10, LC05c, Ste08, VC09, VC06, vRC11]. CMC [GLLL11]. Co [BK10, LV06, LZ08a, Mar02, AR13, BSK13, DYFC09, HJ13, LC05a, Ley07b, Ley08b, RZ04, TSW12, Vau06, WvE07, YRXS11, ZWW^{+}08, vEW08]. Co-citation [BK10, BSK13, DYFC09, Ley08b, vEW08]. co-citations [RZ04]. Co-evolution [Mar02]. co-inlink [TSW12]. co-link [Vau06]. co-located [YRXS11]. Co-occurrence [LV06, LC05a, Ley07b, WvE07]. co-occurrences [ZWW^{+}08]. co-occurring [HJ13]. Co-word [LZ08a, AR13]. coauthor [Egg08c]. coauthors [ADC12, Sch09a]. coauthorship [LS09, MG07, Pep11, Sin11, YD09, YD12, CSL03]. Cochrane [Pet02]. cocitation [Åst07, AJR03, AJR04a, Ben04b, CHE10a, CISH10, Din11b, Ley05, LW12, MM01, Mor05, OM08, PCBL04, SPH08, WGD09, Whi03a, Whi03b, YD12, ZS08b, ZY13, Zuc06a]. Code [Sha06b]. Coding [Ano01e, GS03]. coefficient [AJR03, EL09, Mor05]. coefficients [WBB10a]. cognition [Cal08, Col05, Cro04, RI08, TTS13]. Cognitive [CM02, KA02, New01, WRB^{+}09, CMD05, Coo02b, DS11, For04b, FMCL07, FC01, Gru07, Gwi10, HC02, Hjø02b, HWJ10, LB08, LLL09, MSYD11, Rie02, SKWH07, YS11, Bil01, Bil02, FWF^{+}02, FMM05b]. coherence [FAC08]. cohesion [ER03]. cohort [LHMW11]. cohorts [HSSV07]. Cole [For13]. Colink [Zuc06a]. CoLIS4 [Ano01b]. collaborate [Bir07]. Collaboration [Dim09, Fra11, Ano13b, Bir06, Bir07, BGY^{+}13, BOGdMA13, CRBPdMAM12, CSL03, CSL04, GSD12, HSKS03, He09, HDN11, HTC11, LT09b, Mil10a, MG07, Pel11, Pep11, SZ11, SSM^{+}08a, SS10, SSM^{+}08b, TH13, ZQI^{+}10, Son10]. collaborations [BBRdMA12, Har08a, Hay06, LP07]. Collaborative [BP06a, BP06b, Sha06b, Zha02, Ano08a, Ano10, BPT07, Cal08, FPCB04, KRS13, LC12, RRW05, SM10, Sha13a, SKM01, SV03, SXZ11, Wil00, WLC09, Xie11b, ZS11a, Win08a]. collected [LY06]. collecting [TS11, MMCC04]. Collection [Cas08, Afi03, AW05, CC04, DW13, JTLG12, KPEO10, KLP09, LK06, Nic03a, Ror10, SH07c, Ste03b, SJ10, WCMK^{+}12, WT08, CS06b]. collection-evaluation [KLP09]. Collections [Gal12a, Ano01f, Chr03, CC04, DSG^{+}12, Fei11, Gor01, HJ13, Kis10, MH04, Sau09, Sin12, ZL08, dMFRN^{+}10, vHVD^{+}13, AKW07]. Collective [SS09, CCCJ10, HZ12, SK01, SK02a, Yi12, WK13]. collectors [LT09a]. College [Che09, GKL09, GL12, Lim09, Zuc06b, Cox09]. collide [Pel11]. collocative [Ten12]. colonial [War02]. combine [WLL10]. combined [FMM05a]. Combining [MO04, RLvR02, Whi07a, Whi07b, DRM^{+}04, ESK06, LYJ^{+}10, POMVULMC13]. come [Dil02]. Comment [Szo10, BL13]. commentary [Ben04b, Ben07]. Commenting [TSV12]. Comments [YKT03a, Bur03a, Bur03c, Bur06b, Bur09b, WvE07]. commerce [Fah02, HCC04, Pee12]. commercial [JLCL04]. commission [Hen13b]. committee [Hal04a]. commodity [ME03]. Common [HB09, Tie10, Gor01, Šau04, SGJ13]. common-sense [Gor01].

Commonality [HC09]. Commons [Opp09, Des08b]. commonsense [FMM10]. Communicate [Wil10a]. communicating [Gaz07, Tom04]. Communication [Ard13, AD04, BMK03, BD07a, Bri10, Bur01, Bur06a, Bur07, Bur12, CH13, Cam10a, CM06, DC01a, DA05, DS10a, DG12, Egg07b, ER08b, Egg10a, Egg10b, EG12, Egg12, Gup09, Hjø01b, HW05, Hu09, Jam08a, Jam10, KMK03, KH13, LGA09, Ley05, LO10, LO13, LSZ04, LC05c, Luy13, Mos09, OW04, Opp07, PT05, Rok12, RM13, Rou06, Rou07, Rou12, SD01, San08, Sch13a, Sch13b, Sch13c, Sch13d, Sta10, Sto06, TWKH13, Van12, Van07, Van10, VE13, WSJS01, Ye11a, Ye11b, ZZL12, ZTG13, Ack02a, AGOP06, Bar06a, Bro01b, Bro03b, Bro07, Cha09, CPP09, CvL12, Cro01, DHKW10, Dum01a, Ehi02, Gaz07, Har08a, HS13a, Jac01, KO12, KTA12, KO05, LR02, Ler10, LP09, Ley10, MDM06, rP08a, rP08b, Per03, Rad06, RL09, RB08, ST13, TWU10, VL13]. communication [VC08, WBJM12, WOE^{+}13, YZLC13, ZSL10, Ess10, Swa11, San02]. communication-based [LR02]. Communications [Aks08, AJBAFHCE08, BBC09, Bat08a, Ben07, BMD09, BBB^{+}08, Bur09b, Cam10b, CS02, Cro04, CM08a, Cro08, Dav08, De 11b, FS07, GMM09, GACCNM^{+}08, Har08b, HC02, HWD03, Jör04, LG10a, LZ08a, LP09, NY08, NY09, Pra11b, Pra11a, RY08, TS09, TWF08, Tol09, VBMM11, WvE07, Xia10, Zha10a, ZS08c, vdB03a, Ano08c, BD08, Has03a, SK02a]. communities [BBC01, CLD^{+}13, CLL13, Dum01a, Har07a, Hol01, LFLP13, NSJ13, O’C13, Riv13, RS10, RPH^{+}12, SKS11, SBBF09, Tsi13, WK13, Cox10]. Community [Kwo05, Cho13a, CB13, HSH10, Har08a, KKGB11, KH09, Kom11, Kra13, KY04, KO05, LT09a, LDS^{+}11, LP07, NNY10, Orm01, PM11, POC09, PDU02, Sri07, SL13, VW12, VL13, WJD^{+}13, XFCR13, Kwo05, LDS^{+}11, Dum01a]. community-based [KO05]. community-driven [KKGB11, KH09]. Compaine [Sch01]. companies [Lai09]. companions [CK08]. Comparative [FPZAGBdMA05, GTHZ05, Kno06, SKTM09, CL01, CD06, DBP^{+}07, Gor12, LHHVH09, MPR^{+}05, Ror10, STS12, Shi08, SS06, WBY^{+}09, YKT03b, ZSL10, MLO02]. compared [LB11b]. Comparing [ACGL09, BSM^{+}05, HNB^{+}09, LBMO11, SK12, Chu07, LAGVG06, NV09, SH07a]. Comparison [HSSV07, ALW10, ASABS^{+}10, BMD08, CTS04, Egg09b, FS10, GKL09, HS02, HH01, JSP05, KASGAM09, LSB13, LS12, LdMAGB10, LRB^{+}13, LW12, MS02, MBC10, MR08, MKMM09, SB07b, SB07c, Sch08, SD03, TS11, TSJ09, VS05, VY12, VC08, WT12, WW13, WWZ09, WRM^{+}02, ZS11a, ZL07a, vEWDvdB10]. comparisons [The08b]. competence [Min10]. competencies [FFMPM13]. Competency [Cui10, Hjø02c]. competent [LY11]. competing [CCMM02, CL08, LCW06]. competition [DC09]. Compiler [Fei11]. complementarity [Egg10c, TS01]. Complementary [HML02, SSB01]. complements [Hon12]. complete [Sti05, The08a, Whi02c]. completeness [GP10, MMMM12]. Complex [GBF^{+}12, Szo11, LM09, MTS^{+}13, TWW^{+}06, Swa11]. Complexity [Bur11, AM11, Bys02, CF07c, JTDG13, ZS09]. component [Gel09]. Components [ŠKL08, Din07, ST08, YEV08]. Composition [SHS^{+}11, WS12]. comprehension [CTZ05, Lus07]. comprehensive [ASAK04, Bha05, BBJ^{+}06, Su03a, Su03b, ZS11a]. compression [ANdlF07, CW01, WM07]. Compromised [Nic09]. computation [DGKS04, GP10]. Computational [EAPL05, KSA09, Arg06, HLR09, War10a, WHP10, Arg03]. computed [DLA^{+}10]. Computer [Ano01e, Bia10a, Cal08, DDO11, Ess10, Vil10a, BEES10, Coo07, Fra11, Gu02, Lai09, LJLlW08, MR08, rP08a, rP08b, RKO10, WV13, YH04, Sim07, Ho08]. Computer-assisted [DDO11]. computer-mediated [rP08a]. Computerization [Has09]. Computers [Plo10]. Computing [Ano01e, Ivo03, CTZ05, HVPLHP06, Kos04, Tal05, Che10b]. concentration [Bur05, Bur06a, Egg02b, Egg05c, LGA09, YKT03b]. Concept [GOR06, Hjø09a, MSWK11, Bor03, CS08a, CBL^{+}13, Hjø09c, Hjø10a, Hjø10b, HMFW12, HD06, LZC09, Szo10, TS08, vdEvMK^{+}04]. Concept-based [MSWK11, HMFW12]. Concept-matching [GOR06]. conception [CMH10, IS12]. conception-based [CMH10]. Conceptions [Zin07b, Ano01b]. Concepts [Bia10b, Sto10, Szo11, WKdJvdBV01, ZL07c, Din12]. Conceptual [BH05, LY06, Yoo12, Zin07c, APLG13, BW05, CC04, DW13, HM13, JJBC01, KK10b, KT07a, Nah07, NSQI03, OT08, PV03, Pom05a, Tan04, ZB07]. conceptualization [SW12]. Conceptualizing [HLP13, SH07a, The02, ZBK09]. conceptually [GD03]. concern [BPJR07, TPF06]. concerns [YRW07]. conclusive [BIP05]. condensed [Moe07]. conditional [SZC11, TCW11]. Conduct [HJ04]. Conference [Ano01d, Ano01e, Ano03h, Ano03f, LLA08, Ano01b, Ano03d, Ano04a, MO08, Sel05, CKT06]. conferences [CP13]. confidence [SRS^{+}13, SCC^{+}12]. conflation [AN09, DML02]. conflicts [AYN13, TPLC05]. conformity [Tsi13]. Confronting [Him03]. confuse [Cro11a]. confusion [Jam08b]. Congress [FP04, GGR10, YC10, Yi10]. Congressional [CFW^{+}12]. connected [Bar01a, Tom05a]. Connecting [BJ09, For13, MC03b, Hay07]. connection [TH03]. Connections [Pre09, SCC^{+}12, SST06, Har09]. connects [Col11]. Connie [Fou01a]. Conrad [McC08, McC09]. consensus [Gor02, KB09, Zha02]. Consequences [Ago04a, FN08, Ma12a, TTS13, Wes09, vDH12]. consider [XC06]. consideration [MB13, YKT03b]. considerations [Ma12b, SH07c]. Considering [GOR06]. consilience [De 12, Pra12b]. consistency [WOB09]. consistent [BM11]. consortium [KK08]. constrained [All11]. constraint [For04a]. constraints [Kle08, Slo07]. construct [CJ04, DW04]. Constructing [Gen12, JTLG12, VTD10, vdEvMK^{+}04, JZW^{+}11, KLY08, WvE12b, YLL06]. Construction [Kap09, Bur07, HZ03a, NJN08, Rou07, Sto11, Tal05, YL03a, ER02]. constructionism [TTS03]. constructs [JR10]. consulting [SC11]. Consumer [MMBF09, RHS10, CL06, DJWR02, Gen06, Han01, MK13, SMY09, VC09, Zha13, vRC11]. consumer-generated [Zha13]. Consumers [Row09, LMV^{+}12]. consumption [DAW06]. contemporary [Cul01, Him08]. Content [RL09, ZIB01, dV01, Ahm07, AYX^{+}13, BP10b, Bro07, CH11, CFW^{+}12, EHD05, GKE11, GRJS01, Hjø01b, HFC10, KW01, KTA12, Lak07, LP03, LGCA10, Lus07, NDD^{+}09, rPLM09, RPH^{+}12, SW12, SRDZ06, SYNC08, TL01, TYS13, VGM^{+}11, XZW13, YKCL12, ZI01, ZDM13, KW02]. Content-based [RL09, BP10b, LP03, NDD^{+}09, SRDZ06, SYNC08]. content-bearing [KW01]. content-sharing [XZW13]. Context [LCR^{+}05, Liu09, Liu10, AXP11, AF12, Ano14, Aus01, Bai04, BD10, BBM02, BJ09, CLMB10, CCC10, DLA^{+}10, FDH04, GRS^{+}07, Gre06, HM13, Käk06, KK11, Kel06a, Kel06b, MMC05, MHM12, MA12c, PK05, RPYZ07, SMAB09, SB03, XC09, Zha13, AKS11, Pre09]. Context-based [LCR^{+}05, Liu10, AXP11, CLMB10, FDH04]. context-dependent [DLA^{+}10]. context/permutation [Wil10b]. contexts [DH13, FFMPM13, Gaz07, GLLL11, LKK03, LC13, Wes09]. Contextual [CHE10a, HBBdR10, WTP12, Cho10a, HDN11, HCO03, KDL02, RPG^{+}08, SCZZ13]. contextualization [HH13]. Contextualizing [BDH12]. Contingency [Atk01]. continuance [CLL13, HBT^{+}09]. continued [ZS09]. continuities [VW12]. continuity [BIP04]. continuous [Bur06a, Egg05b, Egg05c]. continuum [Wor07]. contradiction [Atk01]. contrarian [SC08]. contrast [Ror07]. Contrasting [MM01]. contravention [WG07]. contribute [Cam10a]. Contributing [LFLP13, YD13, XW02]. contribution [Aks06, Bir06, SV05]. contributions [AŞT06, ASR^{+}10]. Control [Wil10a, BN08, Ben04a, CAC10, CC04, For04a, Him03, Miz03, MC03b, PK11a, SW12, War06]. Controlled [JLB08, Yeo11, Abb05, LS12]. controversies [Ahn11]. controversy [Ben01b, Ben04b, Hjø09c]. conundrum [Bor12, Pra10, WBB10b]. conventions [Ami01, Bod09]. Convergence [For09, For04a]. conversation [KO12]. Conveying [PK05, TTT12, LB08]. cooccurrence [vEW09]. cooperative [CLHHVH11]. coordinate [Kil04]. coordination [Bod09, DS11, Sha13a]. CoPE [SXZ11]. CopeOpi [KHC09]. COPLINK [HLC05]. Copyright [Enn08b, Max04, BNUA04, GKL09, Lip03, Col07b, Lip13]. core [Sch10, SKM07]. corpora [POMVULMC13, TK06, YL03a, ZBK09]. corporate [hChC05, Huv10, Xie06]. corporation [EAPL05]. Corporations [Mea10]. Corpus [KW01, TJLJ07, CL01, LY05, LY06, RSMVULPO11, WTLC06, YL03b, ZBK09]. Corpus-based [KW01, TJLJ07, CL01]. corpus-related [ZBK09]. correct [Põd10]. Correction [Ano11, Rob10b, WCC10b, Whi04, Ano10]. correlates [BLH^{+}10, WFE^{+}02]. Correlation [Ler13, AJR03, EL09, FCjM12, Pra10, WBB10a, WBB10b, WM06]. correlations [vR06b]. correspondence [DO01, SLS06]. corresponding [Hu09]. cosine [EL09, Ley08b]. Cost [TS06]. costs [Wal07]. costume [Sau09]. coterminal [KAB09]. counted [LRB^{+}13]. COUNTER [DP06]. Counting [HLC11, ZS11a, Egg08b, LO10, LS11b, LB11a, MR08, Sch09b, The02]. countries [Ano13b, CF07a, DO01, GSD12, HC09, LDM12, MDM06, MKMM09, WJD^{+}13, Win08b]. country [HLC11, YYL13, ZRG11]. counts [BD07b, BB03, Che12, Dav08, LZB13, MY07, TH03]. coupled [LF03]. coupling [BK10, Mor05, NSJ13, YD12, ZS08a]. course [Ano03d, CLB^{+}05, CLL^{+}07, GBMH^{+}09, LCB^{+}07]. courses [BM06, LGM^{+}09]. court [Har07a, HLR09]. Coverage [HW11, Egg13a, FN08, LM11, WW03, WL11a]. covered [ZS11b]. Covert [Hay07]. coword [YD12]. craft [McQ09]. crawler [FAC10]. crawling [Cot04, TS06]. CRCTOL [JT10]. create [KB11]. Creating [Bro02, Enn02a, Koh13, Ago06a, Enn01a]. creation [FR08, HL05, SW12, Awa06]. Creative [Fau09, Gup09, Hal09, Lar09, LMH09, Lun09, MR09, Mar09b, McQ09, Rin09, Sau09, SHS09, SHO11]. Creativity [For04a, War13, CM07, Van03, War10a, Sim07]. credibility [Hon06, LMNS13, LT10, Met07, RHS10, Sav11, WB02, Hyl03]. credit [Bir06]. crime [KL11, LY05, Wal06]. crime-psychology [Wal06]. crimes [KL11]. criminal [MCK08, SXCC07]. crises [HZ12]. crisis [Enn02a, HPGG^{+}07]. criteria [BR11b, CG06, KWR^{+}07, KO09, LBR02, MS02, MHM12, TS01]. criterion [KAD06]. Critical [Enn03c, Day01, Day07, Jap08, Wij12, vT08, Hil11]. critique [Bat08a]. Croatia [Ano03d]. Croatian [IS06]. Cross [BGY^{+}13, HSH10, OCO08, SK06, BAT08b, BLA07, Che06c, CS06a, FJRNV^{+}06, LCR^{+}05, PBS^{+}04, PLBS06, SKM11, TJLJ07, VC08, WOE^{+}13, YHHL12, PB01]. Cross-campus [BGY^{+}13]. Cross-cultural [HSH10, BLA07, VC08]. cross-disciplinary [WOE^{+}13]. Cross-Evaluation [SK06, SKM11]. cross-language [Che06c, CS06a, FJRNV^{+}06, PBS^{+}04, PLBS06, TJLJ07, YHHL12, PB01]. cross-lingual [LCR^{+}05]. Cross-validation [OCO08]. crossborder [BLA07]. crosslingual [LY05]. crosswalking [RMC12]. Crossword [Efr08]. CS [WV13]. cube [NJN08]. cues [BJ09, FS11, LMV^{+}12, Sav09, SKWH07]. Cuisine [HKBY^{+}10]. Cultural [CH11, Esc10b, O’B12, Pet12, BLA07, GSB08, HSH10, Jap08, Pat11, Sri07, SPR09, Vau06, VC08, vHVD^{+}13, Cha11, JLR04]. Culture [Gal12b, Swa11, Wes07a, CBDH08, LDM12, vDH12, Enn03c, Enn08b]. cultures [CF06, TVFW07, Ben01a]. cumulated [JP08a, JP08b]. Cumulative [vR06a]. curation [RPH^{+}12]. curiosity [Bow10]. current [DJWR02, OV01, TrP13, Ahn11, Pow03, Cas03]. curriculum [Lat02]. curves [Bur07, Rou07, ERR07]. cusp [BMK03]. customer [Ago06a]. customer-driven [Ago06a]. customers [OCJ12]. customizable [BP10b]. Cyber [Day02a]. Cyber-Marx [Day02a]. cyberlearning [LJ13, Liu13b, MC12]. cybermetrics [AGOP06, Nic10]. cyberporn [LC12]. Cyberspace [Yeo11, Fle06a, Him03, Nay06]. cycle [Wal11b]. Cycles [Day02a, PMBdS10].

D [Ano06b, Ano11, Sti01b]. daily [FFMPM13, YC11]. Dancing [Det11]. Dark [Les11, CCQ^{+}08, HKC09, FAC10, Enn03c]. Daryl [Cul01]. Data [Ben02, BdMAM13, DHW13, KEA^{+}05a, LIL09, LZC^{+}07, TWU10, WEF^{+}07, BWW03, Bor12, BMD08, BD09, Bor13b, CS04, Cro13g, Dai06, DJF^{+}10, EG12, Glu01, Ham07, HS01, HSJB12, HWD03, JSI^{+}04, KWR^{+}07, Ley07b, Ley08b, LdMAGB10, LB12b, MH10, MS05, MY07, NJN08, Nic03a, NS07, PD08, RTM04, RKBS11, SF07, Shi12, SBA12, SH07c, Spo08, Ste03b, SP11, Su06, TS11, Tse01, VS05, Vau06, VY12, WvE07, WY07, WGW12, WP05, WM06, YNF09, ZKSE13, Zin07c, vEW09, vdHR11, CTZ05, GZS13, HH13, NHJ03, PGFMP13, Pel11, WCMK^{+}12, Sch02, Che04]. Database [Ano01d, Mac04b, Bro03a, FN08, GLAA07, Gor12, Kim06, LYJ^{+}10, Ma02a, Mac04a, RPWM06, VRL^{+}09, LB05, ZL07a]. databases [DGA11, Dai06, Egg13a, Enn01a, FMM13, Gel09, HZ03a, HW01, HW03, HW05, MS09b, Rob01, RCT05, Sha06a, Whi06, dV01]. date [HNJT06, CvLvR10]. David [Ave13, Pon01]. Davis [Fou12]. dawn [Atk01]. DCI [JP08b, AJ10, JP08a]. DCI-index [JP08b]. dCollection [KK08]. de-identification [NS07]. Dead [Kaz11]. Death [Day11]. debate [CFW^{+}12]. debates [DC09]. Deborah [Fou12]. decade [GD07, ZS08b]. decay [GN07, HNJT06]. deceive [Lyn01]. Deceiving [Kra13]. decentalization [War06]. deception [WBJM12]. Decimal [Wan09]. Decision [BR11b, AD09, Ago02, All11, Beg05, Ben04a, CL12, ESK06, FG01, Haa03, KBG12, LC06, Nic03b, RRW05, RBLH13]. decision-making [Beg05, KBG12, Nic03b, RRW05]. decisions [RR12, Fou01a]. decline [LGA09]. decomposing [AK13]. decomposition [Bar12, BB13, Egg13b, NB06]. decreases [Cam11]. DeCSS [EHD05]. deep [HNJT06]. defenders [Har07a]. defense [AJR04b]. defined [Abb05, Tol09]. Defining [JSBK07, Bat08a, OT08, Zin07c, Hep04]. Definition [BL09, NJ07a, CCCJ10, Haw01]. Definitions [RZ04]. Deflating [Sun06]. degrades [VH12]. degree [CHH09, YKL07]. déjà [CS06b]. Delayed [LB13a, CKF07, Van12]. delays [TDC13]. DELETE [Bla10]. deleting [Egg10d]. Delicious [LLD^{+}12, Yi12]. delimiter [YLZ10]. delimiter-based [YLZ10]. delineation [LP09]. delivered [CL12]. delivery [TÜ05]. Delphi [BAB02]. Delusion [Les11]. demand [LR08]. Demner [Ano06b]. Demner-Fushman [Ano06b]. Democracy [vD10, Ren04]. Democratic [Buc08, Bus07, Wri10]. Democratizing [Cul06]. demographic [FMM05b]. demonstrated [YS11]. Demonstrating [DW13]. DeNardis [Cha13]. denial [TS06]. dentists [Lan06]. Department [LY05]. departments [BM11]. departures [Kaz10]. dependence [Bur04a, Los01, MS09a]. dependencies [PR07, SML12]. Dependency [YKT03a, Bur03a, KCC12, YKL07, YKT03b]. dependent [DLA^{+}10, LAL08, RJ08]. deployment [Awa06]. depository [Glu08]. Depth [JLB08, Har02, KSK01]. Deriving [KLAS13, LDvM12]. Derrick [Ano12]. describe [LDL^{+}08]. Describing [Kim09a]. Description [War07c, Jör07, RJ07, SB07a, Sti03, TrP13, Whi06, YK05]. descriptions [BZC07, CH07, Cui12, rPLM09, TL01]. Descriptive [Ley03b, vdB03b, FG01, Tan05]. Descriptor [Ano06b, HRK^{+}06, WIY01]. descriptors [GLAA07, JJBC01, SL10]. Design [ALT09, AHK08, Fau09, Gup09, Hal09, LBR02, Lar09, LMH09, Lun09, MR09, Mar09b, McQ09, MO07, O’B12, Rin09, Sau09, SHS09, Tan04, Ano01f, BDH12, BBMN08, Bha05, BW05, BJD10, BSM^{+}05, CMD05, Chu05, CTS04, EZ08, Fle06b, Fou01a, GRS^{+}07, Han01, Hav02, HML02, HS01, JHW^{+}08, KS10, Koc06, LB11b, Li09a, LKG^{+}10, Ma02b, MW10, MTS^{+}13, Mos04, Nay05b, PBS^{+}04, RHS10, SWF^{+}02a, Tom06, Whi06, Wij12, Wil06a, Xie06, YR02, ZL07d, HDB07, OKG07, PLBS06, Bol07, Ost08, Sti02]. designers [FDB^{+}12]. Designing [AH05a, Bur02a, Dav06, Fou01b, LBNB04, Lar10a, Wal02, Fei11, LBC02, PBS^{+}04, Bro02, Gal12b, Vil10a]. designs [Cha11, Jär07]. desktop [MC03a, RHWB11]. detect [APSA12, CC01a, ZS07, ZKSE13]. Detecting [MG10, SKTM09, WGD09, Che06b]. Detection [ESL^{+}10, ACK12, ACH^{+}04, Ano11, CKB^{+}10, DLVÁ12, MG09a, PN09, Sta11, SL13, TBP^{+}10, TBP12]. Determinants [DT13, HBT^{+}09, LR08, LDM12]. Determining [HLLS07, LKZC11, LHMW11, RPYZ07, Tom04, BIKLY09]. deterrents [WG07]. develop [GLDM02]. Developing [BT05, CCS11, KLP09, LBTN09, Zha10c, CF07a, Ma02b, MDM06, Reu05, WEA^{+}12, Abb05]. Development [BPJR07, KK08, TPLC05, WLHO11, Afi03, Ahn11, ABJ^{+}06, CWZ^{+}10, CF09, CC04, HDG^{+}07, JLR04, KT12, MDM06, MBBK07, NGJ03, Nic03a, OT10, PPH06, Ste03b, SLR^{+}11, TS08, TTS03, Wan06, Win08a, XFCR13, vRFBvW10, CM02, CPP09, LL12, Che09, Jam10]. developmental [McC09, Per03]. developments [JLCL04, LS08]. Dewey [Wan09]. DEXA [Ano01d]. diachronic [FT12]. diagrams [TSW12]. dialog [YWC08]. dialogs [TW10]. Dialogue [LHLL08]. Diane [Ing12, Efr12]. Diaries [TD02]. Diasporic [SP07]. dichotomy [Tij10]. Dick [Wad01]. dictionary [CW04, LKJ08, NSQI03, VGM^{+}11, YHHL12]. did [Dil02]. didactics [God07]. Didegah [Ano13b]. didn’t [Jap07]. difference [DS10a, Lus07, SZ12, VS03]. Differences [AWBB09, Ben08, Cha11, FMM01, FMM05b, FMM05a, FN08, LS11b, LB11a, LC13, MKMM09, MC03a, Mor10, PAR13, TWU10, Vau06, VL13, WY06, Whi02d, vR06a, BSP10]. different [Aks08, ACHVH10, BMD08, Bur05, CMD05, CS02, ER08a, HW01, KH13, MBC10, Meg01, NKLC09, PNHG07, SKTM09, SM05a, Sun12, The02, WBJM12, YB10b, YZLC13, ZS09, vRD09, KH13]. differently [CM11]. difficult [Bha05]. Difficulty [Efr08, CTC05, Jam08a]. Diffusion [Sha06a, Gen06, LR11, LR10, LKZC11, ST09, VG06, YK05, Has09]. Digital [ACMS05, Ano08a, Bla10, Bor09, Det11, Ekb09, Esc10b, Fry10, Gal12a, Gua11, Jam08b, Jan02, Kva10, MWG06, MHM12, Swa11, Tho09, Urq09, War09b, War10d, Wes07b, Wil00, vD10, Abb05, Alb10b, Ano01c, Ano03d, Ano04a, Awa06, Bar01a, BAB07, BR11b, Ben01a, BH11, Bol07, BSM^{+}05, Cop11, Cus13, Dav06, Duf11, Enn02a, Enn06b, Enn08b, Erc06, FMP01, Fle06a, FCM04, FMCL07, GLDM02, Hea05, HBSC13, HNJT06, HDB07, Jam08a, Jap08, JZW^{+}11, JP02, KTL^{+}13, KFW08, Koe04a, Koe04b, Koh13, KEA^{+}05a, Lan04, LS10b, Lip13, LB12c, LLG12, MBG^{+}07, MBBS02, Mel04, Nay05b, NC13, Nic03a, NY08, NY09, NV09, Pal05, PGSH06, rPLM09, PBV08, Reu07, SHO11, Sha06b, Sin12, Sre02, SC11, TVFW07, Tom01a, TP11, TMPP13, TTS03, Van03, Vau03, Vei06a, Whi06]. digital [Win08a, XW02, Xie06, XC09, Yak05, Yan01, YC10, Zha10c, Ano02a, Max04, SBBF09, Enn03a, Gon08, Hjø01a, Kaz11, Lip13, Pon01, Row09, Tom06, Van03]. Digitalization [For09]. Digitized [Sau09]. dimension [GdMA01]. dimensional [KHC12, Lak03, MCDA03, RPWM06]. Dimensions [Zha10b, Bri08, HSJB12, LS09, MBBK07]. direct [BK10]. Directions [Jul12, Gru07, Mar07b, Wan07]. directories [BIB07, BW05, CMD05]. directory [ZATH09, CC09]. Dirichlet [SLR^{+}11]. Disadvantage [BdMAM13]. disadvantaged [LKG^{+}10]. disaggregation [ST08]. disagreements [AYN13]. disambiguation [Ano06b, CFN^{+}10, DGA11, HKKP10, HKKP13, HRK^{+}06, JZW^{+}11, LKBJ12, LSD13, MPS12, RHCP12, SZ12, TWSS05]. disappearance [BIP04]. disasters [WZL09]. Discarding [Tij10, JCDR^{+}06]. disciplinarity [CY08]. Disciplinary [DeN09, Whi02d, ADC12, Buc11, HM13, KT08, SH07b, SLR^{+}11, WOE^{+}13]. discipline [HO09, KT07b, VŽ08]. disciplines [ALW10, DO01, Har06b, KH13, LS11b, Men03, MBG03, NSJ13, VS05, Hil11]. disconnect [Cro11c]. discount [DLA^{+}10]. Discounted [JP08b, JP08a]. discounting [FR13]. Discourse [Exn02, Bro03a, HLR09, Jac01, ME03, Jap08]. discover [Nic03a, SKM07]. discoveries [CA07, RN09, WKdJvdBV01]. Discovering [BR01, Che04, Cri11, Rya09, ASQR09, CH03, NJ07a, NJ07b]. discovery [Ben02, FFPW04, HHHR13, KT07a, KHC09, LT12, RW05, RKO10, Sma12, Sre02, Ste03b, SSB01, SST06, TR05, WKdJvdBV01, vdEvMK^{+}04, Har09, Ruc10]. discrete [Egg05b]. discretion [CAC10]. discretization [Ben02]. discrimination [DGKS04, ZW01]. discursive [Tal05]. discussed [Wei10]. discussion [Kel06b, Meh06, Sav11]. Disease [SZC11, Liu13a, NHCW13]. Disengaging [Kaz10]. disk [Whi02c]. Disorder [Tho09]. Disorderly [Hal09]. display [JG03, KK10b, Tan07b, Whi06, Wil10a]. displays [Erc06, MLO02, SRS^{+}13]. Dispositional [NS12]. dispositions [CAC10]. disputes [Him08]. disruptive [Enn01b]. dissect [McC09]. dissemination [Mar02, ST13]. dissenting [Bur01]. dissertation [SNRB11]. dissertations [SLR^{+}11]. distance [BM06, Har05, LP07]. distance-learning [BM06]. distillation [WHTS12]. distinctive [ACMP11]. distribute [RK01]. Distributed [MPR^{+}05, Son10, vT08, Cro04, Kaz10, PD08, SYNC08, Tho12, YRXS11]. Distribution [Cam10b, Gwi10, AFR06, AFR07, BF08, Egg09b, HS02, Mor05, RS08, SH07b, SSPA10, Vin11a, ZY13, LP10]. Distributional [Ben08]. distributions [BD09, Bor13b, Bur05, ER02, Iva01, Mil10b, Sto06, WvEvR12, Win02, WZ02, vR06b]. Distributive [ZTY12]. diverse [Ano04a, SC06, SBBF09]. diversification [HMdR11]. diversity [WS12].

Divide [Kva10, Duf11, Enn02a, Har07b, Jam08a, Jam08b, MBBS02, Koe04a, Koe04b]. DMCA [Lip03]. Do [APLG13, BMD09, Bor11, BdMAM13, CM11, Cro12a, Dra03, Fle06b, Lim07, Muk09, Sch13a, TH04, Bir07, BD07a, ESF^{+}08, GPI03, GL12, HTT07, HBSC13, Kal05, Lim09, XC06, CvLB12, SM05b]. Do-it-yourself [Fle06b]. DocCube [MCDA03]. Doctor [HWJ10]. doctoral [LSB13]. Document [BBB^{+}08, LLC12, Mos04, RKO10, ZRG11, ZA13, BEK^{+}01, BJD10, CMFZ01, CZT06, Fei11, GCR09, GdMA01, Hoe03, hHyW04, HMFW12, Huv11, JP02, KCC12, KK04, Kis10, KB11, LWW01, LWH11, LB03, LC05c, Los13, LAL09, MSWK11, MCDA03, MH04, NDD^{+}09, NKS^{+}06, NRD11, RPG^{+}08, SU11, SCZZ13, TS01, TW03b, TÜ05, Wal12, WCC10a, WZ02, WZ08, YLW10, ZZL^{+}09, ZBK09, ZSR09, Ano01f]. document-category [LWH11]. document-level [KB11]. document-management [BJD10]. Document-type [ZRG11]. documentation [Day01, The02, HS02, Sun06]. Documents [BW06, Urq09, ANdlF07, AK13, BIP04, CCG^{+}06, CLH13, CSA^{+}02, CS06b, CvLvR10, FK06, Glu08, GC07, Gor02, GO03, HZ03b, HW01, hHyW04, HNB^{+}09, HC07, LCR^{+}05, LBMO11, LFC01, LLD^{+}02, Lyn01, Moe05, NJN08, RNLdL01, SM02, TJLJ07, Tse02, fBWLBC06, YW08, ZZL^{+}09]. Does [GV13, HZ07, Hjø09b, Ju07, WWN04, Bir06, De 11b, De 13, Egg12, He09, SZ12, ZTG13, BD07b, Dav08]. Dogpile.com [JSK07]. DOI [Ano12, Ano13b, Ano13c, Ano14]. Domain [BWD02, MW08, Opp09, Ano13c, CKR13, CPO01, CLL^{+}05, Cui10, Dra03, Hjø01b, JT10, Ju07, LVdA^{+}04, Max05, MK13, Pat11, RM13, RRA^{+}13, SXCC07, SW12, SK13, TP11, VV12, VRL^{+}09, WEA^{+}12, Wil04, WTP12, WHC12, Xie06, YBC^{+}06, BBJ^{+}06, ZL07c]. Domain-independent [MW08]. domain-specific [YBC^{+}06]. domains [HW01, TYS13]. domestic [DRW08, Wes07b]. Dominant [Luy11]. Don [Ano03d, Pon01]. Donald [Hol01, Jul12]. Done [Det11]. Don’t [Cro11a, WBB10a]. double [Fon07]. down [LY11]. downloads [Moe05]. drawing [Luy12]. Dreams [GBF^{+}12]. driven [Ago06a, FJRNV^{+}06, HLP04, KKGB11, KTL^{+}13, KH09, LLL09, LBS11, LZ08b, LPSS06, Sri07]. dropout [SRV07]. drug [BBB04]. dual [All11, ER02, PPH06]. dual-level [PPH06]. dual-processing [All11]. Duality [ER02]. Dubrovnik [Ano03d]. dunnit [DC01a]. duplicate [CS06b, Fra03, Kno06]. durability [CvLvR10]. during [Bow10, CG06, Efr05, HZ12, IA12, JR06, JBS09, PNHG07, RB08, Slo02, XJ10, ZS08b]. Dutch [VVLB12]. Dutch-speaking [VVLB12]. DVD [EHD05]. Dyer [Day02a]. Dyer-Witheford [Day02a]. Dynamic [Bur09b, Egg07b, HPAS04, Lak03, Ley02, LS08, Sac06, Bro02, LLD^{+}12, LGM^{+}09, McI02, MPR^{+}03, RY08, Sun12, WLC09, YYL13, vT08]. dynamics [CISH10, DW13, HHWD05, MAL^{+}05, POC09, RL09, Vei09, Wee07, WIY01, Xu07a, XL07, ZQI^{+}10]. dyslexia [Per03].

E-Biomed [KSF04]. e-books [MQH13, DWA13]. E-Collaboration [Son10].

E-commerce [HCC04]. E-government [DRW08, KT12, HLP13, Enn06a]. e-health [Xie11a, Xie11b]. e-learning [LZ08b, LGM^{+}09, SBT^{+}07]. E-mail [Meh06, TLM^{+}12, Ano03d, CKR13, KPEO10, PSO06, YZLC13, PN09, ZKSE13]. e-mailing [Mat09]. E-memory [Bla10]. e-research [JdMAA12]. e-Science [ZZ10]. E-volution [Bro01b]. Each [Ekb12, YD12]. Earlier [BHC06]. Early [Cer10, LMCRRPA09, LGM^{+}09, Cho09a, Jay12]. ease [NY08]. East [JLCL04]. easy [HW02]. EBizPort [MMCC04]. eBusiness [Wil06b]. Ecological [Ago13]. ecologies [Nay07b]. ecology [Ack02b]. ecommerce [JR06]. Economic [DLM09, Ehi02, LB12c, ME03]. economics [HvdW09, Joh01, Tom01a, War05]. economists [Tol09]. economizing [Dai06]. economy [LS13, WRB^{+}09, Saw10, Che03a, Des08a, Sha06b]. ed [Hem07, Jul12, Koe11, McL10b]. Eddan [Pet12]. edit [LATH08]. Edited [Cha13, Dum01a, Fou12, Gal12a, Ing12, Jul12, VT13, Bar01a, Cul01, Enn01b, Hal01, Hjø01a, Hjø02a, Hol01, Pet02, Tom01a, Efr12]. Editing [War10d]. Edition [Fou01a, Har06a, Sag06, Vil10a, Gaz09, Ano08c, BD08]. Editor [AJR04a, AJR04b, Bar12, Ben04b, BB13, Bor11, Bor13c, BL13, Buc08, Bur03a, Bur04b, Bur04a, Bur06b, Che03d, Chr04, Cot01, Dav04b, Day03, Day10, De 12, De 13, Egg13b, EF10, Fid13, Fig06, Fra03, GV13, Hal04a, Hal04b, Hau08, Hjø01c, Hjø01d, Hjø02c, Hjø03, Hjø09c, Hjø09b, Hjø10b, Jar11, Kos03, Kos12, Kra13, Lak03, Lak07, Lak08, Ley03a, Ley03b, Ley07b, LO11, Ley12, LB12a, Mac04b, Moe11, Pil13, Põd10, Pra10, Pra12a, Pra12b, RGLCTSC13, Sma05, Sre02, Szo10, Tol01a, Tre12, Vin11b, Vin13, Wal04b, WBB10b, Whi02a, Whi04, Win02, YKT03a, vdB03b, Han01]. Editorial [Ano02f, Boy01a, Boy01b, Boy01c, Boy01d, Boy01e, Boy01f, Boy02a, Boy02b, Boy02c, Boy02d, Boy02e, Boy02f, Boy02g, Boy02h, Boy02i, Boy03b, Boy03c, Boy03d, Boy03e, Boy03f, Boy03g, Boy03h, Boy03i, Boy03j, Boy03k, Boy04a, Boy04b, Boy04c, Boy04d, Boy04e, Boy04f, Boy04g, Boy04h, Boy04i, Boy04j, Boy05a, Boy05b, Boy05c, Che03b, Cro09b, Cro09e, Cro09c, Cro09d, Cro09a, Cro09f, Cro10a, Cro11a, Cro11c, Cro11b, Cro12a, Cro12b, Cro12c, Cro13a, Cro13f, Cro13b, Cro13e, Cro13d, Cro13c, Cro13g, Dil02, Fur10, Fur11, Kra08, KE08, MD02, Mos04, SL10, Spi02, ZZL12, Cab12, GC07, Jar11, RR12, Tre12, Sti01a]. Editorials [Ano01c, Ano01g, Ano01h, Ano01d, Ano01e, Ano01f, Ano01i, Boy01g, Boy01h, Boy01i, Boy01j, Boy01k, Boy01l, Boy01m, Boy02j, Boy02k, Boy02l, Boy03l, Bro01a, Iye01, Kra02b, Kra03a, Kra03b]. editors [CH13, Hjø02d, Zha04, Sch01, Sti01b]. education [AR13, BLA07, CWZ^{+}10, HH13, Har05, LJ13, SHS09, Buc11]. Educational [Che10b, CL12, DDO11]. educators [FDB^{+}12]. effect [Ahn11, BL09, BWS^{+}10, Cam11, DC01b, FR13, FN13, FS11, JZS09, JLS13, KKGB11, KSK01, KGL12, LKZC11, Los13, Moe07, Nic09, PK11a, PK11b, RR06, SC08, SKM01, Tol09, TSA11, WG10, WM06, XW08, YKT03b, vRD09, Hjø11a, Kra02a, LG10a, SF08]. Effective [CS12, KZ01, LLA05, Vil10a, Wil02, BJC^{+}04, HC11, Mel04, UZ04, WJR05, YLL06]. effectively [TTT12, Cho10b]. effectiveness [Cho09a, DLH^{+}08, DS10b, HB07, JM05, MM04, Smi01, VS04]. Effects [Ago04a, BD10, Cho10a, FN12, FL01, Sha13a, VRL^{+}09, ADR11, Ano13c, ACMP11, BIZGMS10, BHDK13, BV08, CP13, Che12, CvLB10, DLM09, FFPW04, uHG09, uHG10, HB07, HGGL05, HTT07, JBM04, KT12, RL09, Sar07b, SCZZ13, Spo08, TDC13, Tsi13, Wil04, WHC12, YR02, ZC01, KS10, Lus07, SBIS11]. efficacy [Lai09]. Efficiency [Ler10, WG05]. Efficient [HZ03a, HC11, CW01, dCCRdM^{+}12, CK02, SRDZ06]. effort [HB08, PMK11, VH12]. Egghe [Bur06a, Bur07, Rou07, Woe09]. ego [HRC12b]. Eigenfactor [Dav08, Pra10, WvE10, WBB10a, WBB10b, WJD^{+}13]. Eigenvalue [Efr05]. Eigenvalue-based [Efr05]. eight [Zin07a]. eighteenth [NB06]. eighteenth-century [NB06]. eJournal [DP06]. elaboration [CL12, LMV^{+}12]. Election [DC09]. Electronic [Bia10b, Bra09, Mac04b, McL10b, BMK03, Bro03b, Chr03, Chr04, CC04, DS03, Dav04b, Fah02, GD01, GD03, HHfHC10, JZSC09, KTW05, LFC01, LKK07, Mac04a, MBC08, Men03, NS12, Rob01, RKO10, RU07, Spe01, TKB^{+}05a, UR07, Vak08, Vau03, Wag03, WL06b, YRXS11, Zha01, Afi03, Fos01, She10]. electronics [VC09, vRC11]. element [PK11b]. elementary [HDB07, LBNB04]. elementary-age [HDB07]. Elements [HLR09, Gre01b, Zha10b]. elephants [ESF^{+}08]. elicit [CJ04]. elicitation [WL03, WL11b]. eliminating [OIM^{+}11, WM06]. elite [JLK09, SK01, SK02a]. Ellis [MBC08, MT03]. elsewhere [Ano08c, BD08]. elusive [LT12]. Email [AF12]. Emanuel [Hai12]. embedded [Koh13]. Embedding [LB03]. emendation [Bus07]. Emergence [Bod09, LR11, ME03]. emergency [SSM^{+}08a, SS10, SSM^{+}08b]. emergent [TP06]. Emerging [Cho10b, Dra02, Che06b, HSKS03, MSC13, Mor05, NBG11, NHCW13, SBT^{+}07, Shi12, Ste03a, Ano01b]. Emotion [TTS13, dAP13, TWU10]. emotion-based [dAP13]. emotional [SS09]. emotions [Kal06, SS09]. Emphasizing [SBIS11]. Empires [Les11]. Empirical [Ley03a, Sch13b, vdB03a, CAC10, Cho13a, CTS04, ESK06, FMM01, HFW09, HH01, HD04, HLP13, JJBC01, KTW05, Lai09, LCC07, LFC01, Ma02a, OJ04, Rob10a, Rob10b, Sch08, SL12, SBNB11, Tan07a, TWW^{+}06, Vis05, WL09, XFCR13, XTY06, XL07, YSD^{+}08, ZZL12, AHH06b]. empirically [SM03]. employed [Kil01]. employees [CKR13, YRXS11]. Employing [KCC12, PJ01, HNJT06]. empowerment [MMBF09]. enabled [HBSZ13]. Enabling [BZC07, Dum01a, SXZ11]. Enclosing [Opp09]. encoded [Gre01a, YK05, Sti03]. encoding [LSZ04]. encounter [CA07]. Encountering [Rad06]. Encyclopedia [RPH^{+}12, Wor09]. End [Cro09c, Kaz11, CTZ05, Mar07a, Mar07b, WKK^{+}07]. end-to-end [WKK^{+}07]. end-user [CTZ05, Mar07a, Mar07b]. ends [Jam08b]. energies [Nay07b]. Energy [GBMH^{+}09]. enforcement [HW02]. engagement [God07, OT08, OT10, SBIS11]. engine [BR12a, BI02, BI06, BIKYL07, Bil01, Bil02, CFS05, CFY07, CWZ^{+}10, CJ04, HM09, JLCL04, JSK07, JZS09, KSJ06, LDvM12, NSE^{+}06, SY07, Sre02, Ste08, The08b, TSA11, Wal11a, WCCS13, WA03, WEF^{+}07]. engineer [Rob10a, Rob10b]. Engineering [Ano03h, Laz07, Sre02, Vil10b, FM05, JJ10, JTLG12, LAGVG06, MM01, MO08, MV08, Wor08]. engineers [ALT09, BD10, YH04]. Engines [Yan10, BYHM07, JSBK07, LS11a, OAU11, Su03a, Su03b, The08a, WCC08, WCC10b]. English [Ano14, AFG02, BAT08b, Cha11, Che06c, FR08, LWY09, QZCC06, SMAB09, The05, TK06, WT12, YL03a, YL03b]. English-Arabic [FR08]. English-language [Cha11]. English-written [Ano14, SMAB09]. English/ [LWY09, YL03a, YL03b]. enhance [DHKW10, HC11, LZG13]. Enhanced [KK04, LC05a, LFC01, SR06, WRM^{+}02]. enhancement [LH11]. Enhancing [BYHM07, CL11, KC07, LYM07, LHA07, LZC^{+}07, PR07, PS03, SY07, WY07, Ano05, OMS^{+}08]. enlist [Dra03]. enough [HW02, Zac05]. enriched [HDN11, ZS08b]. enrichment [CS08a]. Enron [KPEO10]. ensemble [WLL10]. enterprise [Bar01a, FM05]. entities [BL09, CLYL06, SJTB13]. entitled [Ley03b]. Entity [RHCP12, SR09, PRNZ^{+}11, SZC11]. entropy [DGKS04, LHLL08, WL08]. entropy-based [DGKS04, WL08]. Environment [HML02, BLA07, BJ09, hChC05, Haa03, HNJT06, KH09, LB08, LV06, LFC01, Mat13, Moc06, rPLM09, Rie04, SR06, Sun12, Wal07, YC10, ZZZR04, vT08, Gaz09, Esc10b]. environmental [Hen13a, Sav07]. environmentally [Cho12]. environments [Ano13c, FPZAGBdMA05, HH07, LKK03, Ley07c, Oh12, SP07, WWZ09, WHC12, Bor09]. ePaper [SSTM09]. epidemic [KRK^{+}10, NHCW13]. episodes [jL05]. epistemic [Cro08, He09, Xu07b]. Epistemology [Hjø02b, Day05, Fal08, Hjø03]. Eponymy [McC11]. equal [Hu09]. equality [SOH08, SM05a]. equations [Ye11a]. Equilibrium [McC11]. equivalence [Fig06]. equivalents [Vau03]. EquiX [CKK^{+}02]. era [Cro09c, Gal12a, Glu08]. Eric [Han01]. Erik [Kal01]. ERP [PNHG07]. errata [Pow03]. Erratum [Ano02g, Ano06b, Ano08a, Ano08b, Ano08c, Ano10, Ano11, Ano12, Ano13b, Ano13c, Ano13d, Ano13a, Ano14, AFR07, CF07b, EGRL12, JP08b, Rob10b, Sma05, Tan07a, WCC10b, Yoo08]. errors [Fid13, Hen13b, RPWM06]. Érudit [LM11]. ESI [LR10]. ESI-fields [LR10]. essay [CLB^{+}05, CLL^{+}07, LCB^{+}07, Wil06c]. essay-type [LCB^{+}07]. essays [LCB^{+}07]. Essential [Ago05]. established [WmcsW09]. establishing [ZC04]. establishment [LF03]. estimate [SCC^{+}12]. estimates [Dav08, VY12]. Estimating [OCD08, WZL09, FMM13, Jam08a]. eTax [HBT^{+}09]. Ethical [Beg05, Bri08, LB12c]. Ethics [Ano08b, Bri08, Bro08, Cap08, CS08b, Cox08, Him08, RS08, Wor09, Enn07a, Fro08, Hau05, TS06, WG07, Win09]. Ethiopia [KMD06]. ethnographic [Nat12, VL13]. Ethnomethodological [Sri07]. ethological [Dav02]. eugenics [Ten12]. Europe [BLA07]. European [BT08, Day01, DLM09, KLFG04, LH01, LDM12, PB01, RFV10, vR08a, Bur09a]. Europeana [NCRJ13]. Eurovision [CS06a]. evaluate [DWA13, LS11b, MBC08, Oya08, SKM11, VC06, WT08]. Evaluating [CWZ^{+}10, Cly04, FC01, Hen12, HVPLHP06, JM05, MMMM^{+}01, SC08, SLS06, SBA12, WmcsW09, YB10a, vHVD^{+}13, KO09, Met07, Pra11b, Zha02, RHWB11, RKBS11, WJD^{+}13]. Evaluation [AN09, Hjø11a, KLFG04, PMK11, SK06, AJBAFHCE08, Ben11, BLH12, BMD08, BB03, dCCRdM^{+}12, CTZ05, CLL13, CL01, CJ04, Cui10, DBP^{+}07, ESK06, ECP08, FS09, FR13, FDH04, FG01, GV13, hHwC08, HDB07, JP08a, JP08b, KSK01, KJ02, KWR^{+}07, KS13, KK08, KLP09, KAB05, KHC12, LWW01, LCB^{+}07, LMRB08, Ley12, LFC01, LMNS13, Ma02a, MBC08, MS05, MO04, MO07, NKLC09, NJ07b, OT10, OTP13, OKG07, PZ05, PB01, QCL^{+}12, RN09, RU07, SR06, Su03a, Su03b, SBNB11, SKM11, TS01, The02, TP11, TMPP13, UR07, VG07b, WKK^{+}07, WG05, WTF08, WLHO11, Yan01, Yi08, ZL08, Zha10c, Bil12, JP02, Tom06]. evaluations [CMRS09, FCjM12, JLB08, Ley08a, Vin11a, ZvL11, Hut08a]. Evans [Joh01]. event [CKB^{+}10]. events [CAC10, JLK09, TBP11, Wee07]. ever [Bur02b]. Every [Mea10]. Everyday [SHO11, AHH06a, AHH06b, DPV11, Nay04, Dav09a]. Everything [Bla10, Tho09]. Evidence [Hjø11b, PŠD08, Wil06a, APSA12, dCCRdM^{+}12, Cro01, Ley03a, Nic09, RMC12, Sch13b, SL12, SYL^{+}13, Tan05, Tan06, Tan07c, WWN04, vdB03a, BBMN08, TH13]. Evidence-based [Hjø11b, Wil06a, SYL^{+}13]. Evolution [BIP04, ZS08a, BIP09, CH12, HKC09, Mar02, SCLC06, ST09, SH07c, TYS13, TP06, YYL13]. evolutionary [CLMB10, Ley02, LHHVH09, McC09]. evolutions [LWH11]. Evolving [CC13, Cro13f, DW13, Gen12, ZQI^{+}10]. Exactly [War10b]. examination [JR06, Luy12, MW12, Nic09, Oak09, SNRB11, SC12, WBJM12]. examine [Cha09]. examiner [ACMP11]. examining [Coo02b, HLP13, SSS13, DHKW10, HW02, Joh07, PBN02]. example [Dai06, GD03, Ham11, HC02, Hjø11d, Hjø11c, LL12, Van12]. examples [Bur03c, CBS10, ER08b, FS10, Whi07a, ÖL06, Vil10b]. excellent [BL11]. exception [IS12]. exchange [ZKSE13]. exclusion [MA12c]. exclusively [Sre02]. exemplars [Mor05]. exemplified [BD05]. exercise [FW09, HHNK12, Ley07c]. exercises [AD09]. exergy [Pra11b]. exhaustivity [WZ02]. exists [War06]. expanded [HHHR13]. expansion [Ale03, Gre01a, Gre01c, KK04, Nie03, SR06]. Expect [Ekb12]. Expectation [OJ04]. expectations [Fos01, Ley10, VH05]. expected [BL11]. experience [FN12, HTT07, KP03, MEH07, OL13, Tad06, VVLB12, YSD^{+}08, Har05]. experiences [Jan02, KW02, KP03, Mat13, Oh12]. Experiment [VG07b, Ano06b, Bjö12, CCCJ10, CZH^{+}04, CCR09, CC09, HRK^{+}06, Kil04, Kra02a, KW02, LG10a, QZCC06, SF08]. experimental [DLH^{+}08, DBM02, FS10, GKE11, Kar03, MM04, RNLdL01, TS01, VRL^{+}09, Xie11b]. Experimenting [Xie11a]. Experiments [LKJ08, CS06a, Din07]. Expert [Ano01d, CP13, JHR01, Luy12, Ma02a, ZSR09]. expertise [BdRB^{+}13, HGGL05, Ju07, LS11d, YCE10]. experts [Dra03, GLDM02, HBBdR10, SK12]. explain [BLT07]. explaining [Egg10a]. explanation [KO05, Pra11a, RK01]. explicating [HS13b]. Explicit [KK11, AdG12, DAW06, KT07a]. Exploiting [CJB^{+}07, GW12, OAU11, WTLC06, ZBK09, YLW10, SCZZ13]. exploration [CK03, CGY09, Cho13a, Hei06, HY09, LR02, LR11, LB10, MC03a, MCDA03, SZ08, TLM06, TMPP13]. explorative [Sav09]. exploratory [BIB07, BLH^{+}10, CBDH08, Glu01, Gu02, KT07b, KTR10, MKMM09, STS02, SM03]. Exploring [BHDK13, Gen06, GZS13, GG09, Hil11, Li09b, RFV12, RCD08, SSM^{+}08a, SSM^{+}08b, YD13, BN08, PCY02, Cus13, God07, KPEO10, YS11, O’B07]. Explosion [Cer10]. exponential [LAG08]. export [EAPL05]. Expressions [Esc10b, Suz09]. expressive [CK02]. extended [MO08]. Extending [LV06, RRA^{+}13, TS08, vRC11, SEDC11]. eXtensible [Hol04]. extension [BC09, VC08]. extensive [CvLB12, Wan09]. extent [LTO11]. External [LZC09]. extracted [HB11]. Extracting [NLC06, Suz09, The01, The08a, dBvSW10, CdSG^{+}09]. extraction [HASAT^{+}10, JP02, KK11, KK10b, LWY01, LZLC08, PBV08, VL13, VT13, Wan06, WP05, YLZ10]. extractor [Liu13a]. Extramural [BC05]. extraordinaire [Gar01a]. extreme [McQ09]. extrinsic [SKM01]. Eye [LHB^{+}08, BR12a, KMD06, OA12]. eye-tracking [BR12a].

F [Ano13b, KAB09]. face [Bro03a, rP08a, rP08b]. face-work [rP08a, rP08b]. Facebook [Nea08]. facet [MBBK07]. faceted [KC12, Tan07b]. facets [RH12]. facilitate [WJR05]. Facilitating [PDU02]. facilitation [WOE^{+}13]. facilitators [HKC09]. Facing [Enn02a, vdHR11]. factoid [OYM09]. Factor [KB05, AWBB09, Ben07, Ben08, BV08, Cam10a, Col07a, DO01, ELR09, Egg09b, ER12, Fig06, GC07, Kra13, KC09, LG10a, LO10, LB11a, LLG12, Vin04, Wal06, WvE10, WCC08, WCC10b, ZS08c, NMD09]. Factors [BDS12, DAW06, KS08, YWM10, AMS10, Ano13d, BLT07, Cam11, Cho10a, DLM09, Ehi02, FFMPM13, FMM05b, GL09, HSKS03, HLLS07, HBBdR10, KO05, LB11b, LRB^{+}13, LZB13, jL05, LHMW11, Lyn01, NH12, PZ12, PR10, RPWM06, SB12a, SM05a, SZ08, XW02, ZJS09, FMMF13, WB02]. Factual [Fid13, LS11d]. Faculty [Kim10a, BMK03, BSM^{+}05, LSB13, MS05, MY07, WS12, WT08]. failed [Bjö12]. Fair [LS11a]. Fall [Les11]. familiarity [Alb10b, LMNS13, WG10]. family [Cas08, TP06]. famous [CH11]. Farsi [MH11]. fashions [Ami01]. fast [CB13, LSZ04]. favor [Cro12a]. FCC [CFW^{+}12]. Fear [Cop12]. fears [TP07]. Feature [DARK09, KAD06, CMFZ01, HKBY^{+}10, JG06, LP03, MH11, MG09b, RCD08, WIZ01]. feature-based [JG06]. feature-selection [MG09b]. Features [CMFZ01, AWC^{+}07, FT12, GRJS01, Hon06, Lee10, LLA05, LLD^{+}12, LZC^{+}07, Mar11, NSLD13, ÖL06, Ror07, SBIS11, TVY12, ZLCH06, ZSR09]. Federal [LC05b]. Federated [AYSC06]. FedLemur [AYSC06]. FedStats [CS04]. feedback [CF07c, Fu10, HGGL05, LKJ08, LPGBdMA03, RLvR03, YHL11, dCFLH03]. feeds [TPF06, TP07]. feeling [Kal06]. fees [Ano13d, KH13, SB12a]. Feist [War10a, War10c, War13]. Fellowship [RKBS11]. female [ARPS11]. feminist [BBC01]. Fiber [Enn03c]. Fidel [Ago13]. Field [LRB^{+}13, MS09a, Ame12, Ano12, AMFiA11, lC01, CBvLvR09, FMM13, Hjø01b, JD12, KWS07, MTS^{+}13, SBT^{+}07, SF08, SPH08, VL13, WW03, Wal06, ZYH13, ZS11a, Zin07a, vR08b]. field-association [AMFiA11]. Field-normalized [LRB^{+}13]. field-specific [CBvLvR09, vR08b]. fields [AWBB09, KB11, LB11a, LZB13, LR10, Min10, PAR13, SM05a, SZC11, SPH08, TKV^{+}10, TCW11, ZB07, ZLB11, EGR11, Hen13b]. Fight [GALC11]. Fighting [God07, Kaz11]. figures [AYX^{+}13]. file [BWS^{+}10, GD03, LWW01, YNF09]. file-sharing [YNF09]. files [RK01, Ste08, TK08]. filtering [DRM^{+}04, Käk06, KB05, KSS03, LC12, MPR^{+}05, SDWG01]. filters [ESK06]. Final [KE08]. finance [Huv10]. financial [DLM09, GD03, YC02]. find [Bha05, BBJ^{+}06]. Finding [CS04, ST08, SS13, fBWLBC06, YHL11, ASABS^{+}10, HBBdR10, PSB04, PG02, SKM01, ZSR09]. findings [GZS13, Mar07a, Har08b, IA12, Kel06b, PDU02]. Fine [MLZ10, Cui12]. Fine-grained [MLZ10, Cui12]. finite [iCG09, GMA07]. finite-state [GMA07]. Finland [NV09]. Firm [HHWD05]. Firm-like [HHWD05]. firms [LDM12, SC11]. First [Har04, Ago08, GRZ12, GL12, Hu09, KLFG04, Kil01, LSC12, MM04, TVY12, ZS08b, ZS11a, Whi07a]. first-citation [GRZ12]. first-year [GL12]. fish [WKdJvdBV01]. fit [DW04, DWA13]. fitness [FFPW04, LPGBdMA03]. Fitting [Bur04b, CPO01]. Five [RPWM06, Rya09, Mar07a, Mar07b]. Five-dimensional [RPWM06]. flatter [SSS13]. flattery [FN11]. flexible [BP10b, CMD05, CdSG^{+}09, NBYA03, PR07]. Flickr [CCS11, LLD^{+}12, Ror10, SJ10]. Flood [Dav11]. flooding [MCK08]. Flow [Cop12, McI10, Pet12, CJA^{+}04, HM13, LSJR11]. flows [BBRdMA12, HM03, LLC12]. focused [FAC10, PK05, SP07]. focusing [Wu10]. fog [Hen13a]. Folder [BGBIBM13, BWS^{+}10]. folders [GD03]. folksonomies [PLN09]. folksonomy [Kim11a]. folktales [LT12]. follow [Per03]. follow-up [Per03]. Following [GLDM02, VW12]. Food [Cro13d]. fool [WBB10a]. footprint [TH13]. Foraging [San10]. Forced [SRV07]. Forced-response [SRV07]. Fordism [Day01]. foreign [HS02]. forests [Suz09]. Forgetting [Bla10]. Form [LT02, Bar08a, LR02]. formal [AJR04b, CDLvR03, HJ04, NJ07a, RN09]. formalism [PAL12]. Formalizing [Wil06b]. Formally [WdV04]. formation [LR11, For13]. formed [KOS09a]. forms [Bat06, CS02]. Formulae [Bur13]. formulas [LMRB08]. formulation [Wil04]. forum [FAC08]. forums [FAC10, Sav11, KMK03]. Fostering [Cox10]. Found [Els09]. Foundation [Par10, Hjø10a, RW05]. foundational [Ma12a]. Foundations [Egg07a, Egg02a, Har06a, Ho08]. Four [PGR^{+}13, Koe02, PGSH06, VS05]. four-year [Koe02]. Fourth [Ano01b]. fractals [Egg05a]. fractional [ER02, Egg08b, HRC12a, LO10, LB11a, LZB13, Sch13b, ZS08c, LS11b]. Fractionalized [Pra11b, Sch09b]. fractionally [LRB^{+}13]. frame [KK10b, OYM09]. frame-alignment [OYM09]. frames [Max04]. framework [BN08, BI04, BLH12, BH05, CMH10, CC04, DG12, Gaz07, GLLL11, Hep04, HS13b, JP01, JJBC01, KK10b, Nah07, OT08, PCR13, rP08a, rP08b, PRNZ^{+}11, Pom05a, RRW05, Sar07a, Sar07b, SEDC11, SWF^{+}02a, SGTS07, SLN03, The06, Tom02, TLdR10, WLL10, ZB07, ZDM13, ZLCH06, ZZ03]. frameworks [Ano01b, CL11, Jär07]. francaise [IS12]. Frank [Pee12]. FRBR [McD11]. free [Rig02, SM13]. freedom [Li09a, Les11, Tie10]. French [IS12]. Frequency [LP07, ER02, Egg08c, iCG09, Liu10, LB03, Orm07, PT05, RTM04, Win02, Käk06]. Frequency-based [Käk06]. frequent [Aks08]. friending [The08c]. friendliness [VŽ08]. friends [AE10, Wal12]. front [Åst07, BK10, LAL09]. frontier [Bro01a]. frontiers [Sch04a]. fronts [ASQR09, MYWA03, SKTM09, ZS07]. Full [LZG13, BSK13, HSSV07]. Full-text [LZG13, HSSV07]. fully [Bil02]. function [Egg05b, Egg07b, Egg08c, ER12, Hua13, LT02, Luz09, MSWK11, Riv13]. function-based [Hua13]. functional [AWC^{+}07, PR07, Suz09, VG06, ZL08]. functionality [MBC08]. functions [BBC09, BR11a, FFPW04, Hu09, LPGBdMA03, Zha12]. Fundamental [Bat06, ELR08]. Fundamentals [Gaz09]. funding [Ano13d, BC05, BB03, CvL12, SL12, SB12a]. further [Bur03c, OV01]. Fushman [Ano06b]. Fusion [BJC^{+}04, dCCRdM^{+}12, LIL09, OYM09, Spo08, WM06]. Future [Bea06, Dav06, Gre06, Hen06, JBCJ06, Kal06, KSM06, Moc06, RV06, Ros06, Rot06, Sac06, VR06c, Wil06a, BAB02, Bur03b, CS08b, GBMH^{+}09, Hjø02a, JP01, LJLlW08, Met07, Sau09, SKM07, SS10, Mar07b, Bia10b]. futures [Enn03a]. fuzzified [GdMA01]. Fuzzy [ADD^{+}06, BC09, CS08a, CH03, GOR06, HV01, HVPLHP06, ZL07b].

G [Ano10, Ano11, Ano12, Bar01a, Pee12]. game [Bod09, Chu05, GALC11]. gameplay [LGCA10]. games [Ago08, Win11]. gap [KY08, Sav06]. gap-bridging [Sav06]. gaps [JW09]. garden [Hal01]. Garfield [Ben01b, Vin04]. Gatekeeping [BD05, BN08]. gateways [Fou01a]. gathering [JSI^{+}04]. Gaussian [Iva01]. gaze [KC12]. Gazni [Ano13b]. Gender [HD06, HB11, The08c, XZW13, TWU10, Wes07a]. genealogy [Cas08, SNRB11]. General [CW01, BBC09, CLH13, CvLvR10, HJ13, PS03, Ror10, YLZ10, ZIB01, CvL12]. General-purpose [CW01]. Generalizations [Woe09]. generalized [Egg07a, EG12, YKL07]. generated [Bil02, LY05, OTP13, SS13, WW13, Zha13]. generates [Hub02]. Generating [Liu13b, OCJ12, Sri04]. Generation [Det11, CS12, Has04, Lak07, LVdA^{+}04, Mel04, MO07, Shi08, Tse02, TK06, YL03b, YL09]. Generative [Luk09, BC09]. generic [AK13, Hua13, LCR^{+}05, PRNZ^{+}11]. genes [HSSV07]. Genescene [LC05a]. genetic [BCT02, FFPW04, LPGBdMA03]. genetics [CJA^{+}04]. Genome [Dai06, HSJB12]. genomic [Bro03a, Sha06a]. genomics [UGF08]. genre [FK06]. genres [GALC11, PN06, Pal08, Ros08]. geo [OTP13]. geo-referenced [OTP13]. Geographic [HPGG^{+}07, APLG13, DH13, JG03, VG07a, YS11]. geographical [Hen12]. geography [LP10]. geomorphology [FMM10]. geophysics [FR01]. Georeferencing [VG07a]. GeoSearcher [WA03]. geospatial [Glu01]. geotemporal [DH13]. German [CTS04, SS04]. German-language [SS04]. get [LT04]. Getting [RLF02, SSS13]. Ghosts [Gor02]. Giddens [Ley10]. given [GLAA07]. giving [YKT03b]. Glänzel [Pra11a]. Global [Ano08b, Bra09, Bri08, Bro08, Cap08, CS08b, Cox08, Him08, Pet12, RS08, SG11, Yeo11, Beg05, Cha13, DZC^{+}09, KB11, LR09, LR11, MS09a, Nay07a, ST09, Sre02, SSPA10, Su12, Ano04a]. globalization [YYL13]. Globenet [WWN04]. Glut [Raw10]. go [NCRJ13]. goals [Hon12, Slo02, SLSK09, WGW12]. God [Wal12]. going [RKBS11]. Goldberg [Hai12, Hal01]. Good [Egg10c, Tie10, Bri08, CB13, MDM06, YHL11]. Google [Les11, BR12a, BI06, Bil12, BL11, GP10, HvdW09, KT07b, KT07c, KT09, KTR11, LB12b, LY11, Meg01, MY07, Muk09, OA12, Wal11a, Kva10, Sat08]. got [GL12]. Governance [Mue10, SW12, KT12, MDM06, MC03b, Yeo11]. governing [DAW06, Bra09]. Government [Mea10, DRW08, HLP13, KP11, KT12, LS09, YYL13, LC05b, Glu08, Enn06a]. governmental [CS04]. graded [KJ02, LKJ08]. graduate [AH05a, PR10]. grained [Cui12, MLZ10]. gram [AN09]. grammar [OW04]. grams [ER08b, MAR04, Sta11]. grants [BB03]. granularity [BD10]. Graph [ASW05, TW03a, HCC04, Meg01]. Graphical [Wil10a, MLO02]. Graphs [TWKH13, GMA07]. Grasping [Zha12]. gratifications [CK08, LGCA10]. greater [He09]. Greek [GSB08, Laz07]. Greeklish [Kar03]. green [Cho12]. grid [Ben01a]. Grønbaek [Fou01b]. ground [Šau04]. grounded [Wor08]. grounds [FDH04]. Group [GMV13, Cha09, DHKW10, HBSC13, Kaz10, LZC09, LLC12, WWN04, Zha02, ZKSE13, vR06b, Cal08]. group-based [LLC12]. groups [CCS11, ER08a, JLK09, SM05b, WG05, vR06a, vR08b]. groupthink [Tsi13]. growth [CCMM02, GD07, GR09, HHWD05, HHHR13, Hu09, LAG08, MAL^{+}05, Mor05, MG07, yT04, WZL09, ZQI^{+}10]. guard [Cro09b]. Guatemalan [TSV12]. guide [Ago08, Ben01a, CL03, Enn08a, Fou01a, Glu08, Hol04, Koc06, LKK07, Mit08, Mon04, Reu05, Sti05, Spe01, Afi03, Edw07, Jon02, Wad01]. guided [Sac06]. Guideline [Pee12]. Guidelines [NS07]. guiding [SW12]. Gunilla [BI13]. Gurstein [Dum01a].

H [Ano06b, Fou01b, Fou12, O’B12, Sch01, BD07a, BMD08, BMD09, GP10, San08, Sch09a, AJBAFHCE08, Gar01a]. HAADS [HKKP10]. Hal [McC08, McC09, Tom01a]. Hale [Dot13]. half [SK02b]. half-life [SK02b]. hall [LY11]. Handbook [Bri10, Lap07, Orm07, Sch01, Whi02c, Ess10, The09a, Ago06b]. hands [Jon02]. hands-on [Jon02]. Hans [Wil10b]. happens [WV13]. haptic [JG03]. hardness [DS05a]. Harnessing [Enn01b, Pon01, Yi12]. Harold [Wil06c]. harvesting [SYNC08]. hashtags [MSC13]. Have-Less [Mos09]. headings [FP04, SST06, Yi10, LS12, LRR12, LO13, YC10]. health [BP06a, Cas08, CL12, CG06, Gen12, GRS^{+}07, Hon12, LMRB08, LR13, Lus07, LKG^{+}10, MK13, NS07, Oh12, OMS^{+}08, PGR^{+}13, RHS10, RRA^{+}13, SML12, SMY09, WP04, WTP12, Xie11a, Xie11b, YR08, ZWW^{+}08, ZATH09, Zha13]. health-subject [ZWW^{+}08]. healthcare [BP10a, SS10]. Heap [Egg07c]. hearings [CFW^{+}12]. Hearst [Efr12]. heat [Ye11a]. Hebrew [HKKP10, Gor12]. Hector [Lip13]. hedonic [Xu07b]. Hegelian [Wij12]. Hegemonic [Huv11]. Heinström [Jul12]. Helen [Jay12]. helix [KP11, SK13, LS09, YYL13]. help [AE10, BBJ^{+}06, Bla02, BMTV02, HZ07, SK12, XC09]. HelpfulMed [CLZC03]. Herbert [Wil10b]. Herdan [Egg07c]. here [KJ03]. heritage [Jap08, LB12c, Pat11, vHVD^{+}13, ZYH13]. hermeneutic [FM05]. heterogeneous [NJN08, SCZZ13, SJTB13, YDS11]. heuristic [CFN^{+}10, DGA11, Mus05, The02, YL05]. heuristic-based [CFN^{+}10, The02]. heuristics [FS11, SKWH07]. Hewlett [HD04]. Hierarchical [YW08, ADR11, CJB^{+}07, CFN^{+}10, FS10, GCR09, Liu09, SLN03, YLW10, YLL06]. hierarchically [FL01, vR06a]. hierarchies [RBLH13, ZL07a]. hierarchy [YL09]. High [CN07, Kis10, Ano08c, BD08, BSMD11, Day02a, Fou01a, LB13a, Lai09, LSJR11, LFLP13, OCJ12, WM06, GBMH^{+}09, Hol01]. High-Energy [GBMH^{+}09]. high-impact [Ano08c, BD08, BSMD11, LB13a]. High-speed [Kis10]. high-tech [Fou01a, Lai09]. high-technology [Day02a]. higher [BS09, HH13, KH07, SM13, TH04]. highly [ANdlF07, CA07, Kno06, LT08, Sch13c, TJLJ07, ZS11a]. Hip [NBG11]. HIPAA [NS07]. Hirsch [AL08, BS09, Egg07b, Egg09b, ER12, LL12, Sch09a]. Hirsch-index [Egg09b]. HistCite^{T M} [LAL08]. HistCite^{T M}-based [LAL08]. historic [Sau09]. Historical [YBC^{+}06, Luy11, MQH13, Mar11, Sin12, SJ10, dBvSW10]. histories [Cas08, KSM06]. historiograms [LAL08]. historiographic [McC08]. historiography [GPI03, Luy11]. History [Dav11, Luy13, War09a, YBC^{+}06, Bla11, lC01, CR03, CLB^{+}05, CLL^{+}07, CBL^{+}13, Day01, Dra02, Exn02, HH13, LCB^{+}07, LT10, NRD11, Plo10, Sha13b, SYL^{+}13, Sma03, Smi04, BW06, Ste03a, Fer04]. hits [Eas02]. HIV [NHCW13, Vei09]. HIV/ [NHCW13, Vei09]. Hjørland [Bat08a, Bat11, Szo10]. hobbyist [LT09a]. holistic [Zha10c]. Holmberg [BI13]. holy [Wal12]. home [FDB^{+}12, Rie04]. homebuyers [Sav09]. homeland [RRW05]. homeless [WH09]. homepage [SL06, WHTS12]. homepages [BLT07]. homonymous [OIM^{+}11]. homonyms [Aks08]. Homophily [The09b]. Hong [CFY07, LY05]. honor [Pet02]. Honoring [Kra03a]. hope [Bla02]. Horacio [VT13]. horizons [Sav07]. hormone [Gen06]. hospital [NS12, Plo10]. host [TH03]. hotel [Ano03d]. HotMap [HY09]. hours [TD02]. HPSG [MAA08]. Hsieh [Spe01]. Hsieh-Yee [Spe01]. Hsuan [Ano13c]. http [Ano03d]. Huang [Ano13c]. hubs [BBJ^{+}06]. Human [Bia10a, Ivo03, Kos09b, Lar10b, SC06, Vil10a, WCC10a, CFW^{+}12, FARR09, FPCB04, Fid13, FMM05b, FMM05a, Gaz07, MR08, RPWM06, SEDC11, SC11, Tra08, Wan07, Xie06, Dai06, FARR09, Ago13, Hem07, Ho08, Nay07a, Wes07a]. Human-Centered [Kos09b]. Human-Computer [Vil10a, MR08, Ho08]. human-information [FPCB04]. Humanists [BAB07]. Humanities [LS10b, AŞT06, Ard13, Hel10, HW11, hHwC08, KT09, LAGVG06, LM11, WBY^{+}09, ZvL11, LHS11]. Humanizing [Tom05b]. humans [APLG13]. humble [Jay12]. humming [DBP^{+}07]. Humphrey [Ano06b]. hundred [LSC12]. Hurricane [CKF07]. hurricanes [Chu07]. Hybrid [LS11c, ZLB11, Bjö12, FAC08, LYJ^{+}10, LLC12, TP06]. Hybrid-patent [LS11c]. hybridization [Fle06b]. hydrogeology [SF07]. Hype [Bla02]. hyper [Mel04]. hyper-textbooks [Mel04]. Hyperauthorship [Cro01, DC01a]. Hyperlink [PBN02, KOS09a, KBP10, RFV12]. hyperlinks [PS03]. hypermedia [CM02, Fou01b, LB08]. hypermedia-based [LB08]. Hypertext [War09b, Ami01, Che04, KS10]. hyperwriting [Luz09]. hypotheses [Sri04]. hypothesis [CLB^{+}05, FN11].

I-Commerce [Pee12]. I3 [LB12a, Rou12]. IA [LBC02]. Ian [Ing12, Efr12]. icon [MLO02, Wal12]. ICR [SK01, SK02a]. ICT [AAN12, PR10, ZS09]. Ict4d [Jam10]. ideas [TP06]. identical [Aks08]. Identification [ZJS09, BL09, Bre05, CLL^{+}05, CW04, CS06b, FMCL07, HPAS04, KTR10, NS07, Orm07, OCO08, PBV08, Ros08, SCA12, YKCL12, YWC08, ZLCH06]. identified [CG06]. identify [BSMD11, KWR^{+}07, KDL02, KL11, MCK08, MBBK07, OCJ12]. Identifying [ADC12, HKC09, KB06, PCR13, Ste08, TP07, WWZ09, HZ03b, Hol08, NSJ13, TP06, Whi02c]. identity [ACMS05, FC01]. ideology [Bud07]. IEEE [Ano01e]. IFs [LRB^{+}13, LRB^{+}13]. II [Bil01, BP06b, Che06b, NJ07b, Sar07a, Su03b, SK02a, Tan07c, XL07]. II. [KW02, SSM^{+}08a]. III [Bil02, Sar07b]. IKE’03 [Ano03h]. illicit [YKCL12]. Illuminations [Ham11]. illustrated [Hjø11a]. illustration [MPT09]. I’m [Kal06]. Image [AKW07, AYX^{+}13, Gel09, JJ05, Mat13, MPT09, TLM06, BIZGMS10, BP10b, lC01, CWZ^{+}01, CR03, Cho10a, Cho13b, Chu01, CS06a, GALC11, GKE11, GRJS01, Gor01, Gre01b, HTdV11, KTR10, LSB13, LP03, LLA05, McC09, MMMM^{+}01, NDD^{+}09, PBV08, RI08, Ror10, SRDZ06, Smi01, SBA12, SBNB11, WEA^{+}12, WLHO11, WIZ01, Yoo10, YC11, ZIB01, ZI01, ZL08, GRS^{+}07]. image-applicable [Gre01b]. Image-browsing [GRS^{+}07]. image-identification [PBV08]. Images [Wal12, CR03, KTR10, Nay05a, OTP13, Ror07, SS09, MPT09]. imaginaire [Gaf07]. Imagination [Gal12b]. immediate [HT11]. immigrant [SP07]. immigrant-focused [SP07]. immigrants [FDH04]. immune [YKCL12]. Impact [BF08, CF09, MY07, SS04, Sin12, TVFW07, Vis05, AJ10, AWBB09, Ano04a, Ano08c, BLT07, BBI12, Ben07, Ben08, BV08, BD08, BSMD11, Bor13a, Bor13b, BHC06, Cam10a, Cam11, Cho13a, Col07a, CBvLvR09, CvLB10, DJWR02, DAB^{+}07, De 10, DT13, DHKW10, ELR09, Egg09b, ER12, FR13, Fig06, FR05, GC07, Ham07, HSKS03, HvdW09, HHWD05, HLC11, HAT12, JP08a, JP08b, KAB09, KGL12, KT08, KTR10, KTR11, Kra13, KEA^{+}05a, LB13a, LG10a, LG10b, LC06, Ley07c, LO10, LB11a, LB11b, LB12a, LRB^{+}13, LZB13, Lip03, LHMW11, LLG12, LT02, MBC10, MS09b, Moe07, Moe11, Muk09, NMD09, NSLD13, OYM09, Pet07b, PGL13, Sch12, SM13, Sel05, Sin11, Slo07, SM05a, SPH08, TH04, TK08, TJSL11, VVLB12, VS05, Vin04, Vin13, Wal06, WY06, Wu10]. impact [WHP10, Xie11a, YD09, YCE10, ZTY12, ZS08c, vR08c, vR08b, AL08]. impact-based [JP08a, JP08b]. impact-reinforcing [vR08c]. impacts [LS11b, MHM12, ST08, Vau03]. impaired [STS12]. imperative [DA05]. impermanent [Sel05]. implementation [CG08, Chu05, MO04, MO07, PNHG07, Enn06a]. implementations [BBMN08]. Implementing [dCFLH03, WL09, Kim05]. Implications [Boo01, BP10a, BW06, Hel10, JBM04, LB11b, SML12, Ste03b, TTS03, Whi07b, Bha05, BSM^{+}05, DP06, Day05, FDB^{+}12, GD01, MS09a, MTS^{+}13, Mor10, OMS^{+}08, Xie06]. implicit [AdG12, CBL^{+}13, Day05, Fu10, KT07a, SBA12]. importance [Hjø11d, Hjø11c, MCK08, NMPBA11, TCW11, VŽ08]. important [SB07b]. improve [GV13, GZS13, Gre06, JG05, LT12, dMFRN^{+}10]. improved [HS13b, LSZ04, SY07]. improvement [Dav08, Enn08a]. improves [VH12]. Improving [BYHM07, Bea06, BB13, BSK13, DRM^{+}04, Haa03, LM11, LATH08, LT10, MH11, POMVULMC13, WM06, DH13, MMC05, Nel06]. IMS [ZL07c]. In-Depth [JLB08, Har02]. in-service [JJ10]. incentive [Rab08]. incidents [IA12]. included [GC07]. Inclusive [GRS^{+}07]. inclusivity [Luy12]. income [Hol01]. incomplete [Egg02b]. Inconsistencies [Sch12, Sch13c]. inconsistency [Pra12a, WvE12a]. Incorporating [RLvR03, NSQI03]. Incorporation [McC11, McC12]. increase [CL12, SRV07]. increases [Cam11]. incremental [WS06]. Indagator [LGCA10]. independence [dV01]. independent [LZC^{+}07, MW08]. Index [Bur09b, Nic10, AJ10, Bar12, BB13, BS09, BD07a, BMD08, BMD09, Bor11, Bur13, CM06, CM08a, De 10, De 11b, De 13, Egg07b, ER08b, Egg08a, Egg08b, ELR09, Egg09b, Egg10d, Egg12, ER12, Egg13b, EF10, FR13, GMM09, GG09, GR09, HvdW09, HZ03a, Hov13, Jar11, JP08a, JP08b, Käk06, KAB09, KLP09, Ley08b, LL12, MR08, Opp07, Pil13, Pra11a, Pra12a, Ror10, RJ08, RY08, Sch08, Sch09a, Sch09b, Sch10, Sch13a, SK02b, Tre12, Van07, WW03, WL11a, WvE12a, WCLZ01, Woe09, Wu10, Ye11b, Zha10a, ZLXW11, GPG13, LS10b, LHS11, ZL07a, GC07]. index-number [SK02b]. indexed [Win08b]. indexer [Bod09, MM01, RL09]. indexes [ADV10, ADV13, Whi06]. Indexing [Ano06b, DS09, Ano13a, Chu01, CMH10, Efr05, GOR06, Gel09, LWW01, LS12, LK10, Los13, LLD^{+}02, MW08, Oya08, PK11a, RPYZ07, Rob01, SS09, yT04, WEA^{+}12, Wil10b, WL13, WZ02, WZ08, Din05, Hjø11d, HRK^{+}06, Sag06]. India [Gau09, LKZC11]. Indicator [BB03, BD09, Cro13f, JdMAA12, Ley07a, LO11, LB12a, LAL09, Moe11, Rou12, Sch13c, SNRB11, Vin13, YD10, YDS11, vR06b]. Indicator-assisted [BB03]. Indicators [FF04, AGOP06, BD05, Fra09, HRC12b, KK08, Ley08a, Ley08c, Ley09, LB11b, Li09a, Lia05, Sch12, TSRGJC^{+}13, WCMK^{+}12, WS13, vR06b, LS08, Ben11]. indices [AJBAFHCE08, BMD08, De 12, ER08a, GP10, Pra12b, ZTY12]. Indirect [SK01, SK02a, SST06]. Indirect-collective [SK01, SK02a]. Individual [MC03a, Mor10, CBvLvR09, CvLB10, CvLB12, FMM01, FMM05b, FMM05a, Hut08a, KO05, Moe05, SH05, TKV^{+}10, VŽ08]. individuality [BR13]. individuals [GD03]. induced [CH07]. inducers [VT03]. Inducing [MK13]. induction [WS06]. inductive [Hav02]. industrial [ME03]. industry [BH11, GD03, IA12, JJ10, KP11, Ley04, LS09, YYL13, YH04]. Inequality [HTC11, Koe04a, Koe04b]. INEX’02 [KLFG04]. infectious [NHCW13]. inference [Nie03, SB03]. inferences [DS03]. inferential [Ley03b, vdB03b]. inflation [HLC11, KR01]. Influence [Egg10d, KC12, BL12, BBRdMA12, BD07b, BB03, DP06, ER08b, Egg08a, Hon06, hHyW04, McC08, MA12c, Slo02, VS04, Vak08, WvE10, WJD^{+}13, WZ02, WZ08, ZLXW11, Ame12, CBvLvR09, vR08b]. Influences [Alb10b, YSD^{+}08, ASQR09, HJ04, JLR04, KA02, MM10, MDM06, Per03, ZS08a]. influencing [AMS10, Ano13d, FMMF13, PR10, SB12a, WB02]. influential [CM06, GRSFV12a, Opp07]. influenza [KRK^{+}10]. Info [Mer09]. Infocomm [TPLC05]. inform [HS01, MW10]. Informal [Cox10, Ano11, Bar06a, Hep04, TBP^{+}10, Tra08]. Informatics [Ano03g, Ano04b, Che10b, Che05, Day07, Dum01a, Gaz07, God07, GMV13, LKG^{+}10, VW12, HLC05]. Information [Ago13, Ahn11, All11, Ano01g, Ano01e, Ano03h, Ano03f, Ano06b, Ano08b, Ano08c, Ano11, Ano12, Ano13b, Ano13c, Ano14, AFR07, BYHM07, BI13, Bia10b, Bol08, Bor12, Bor09, Bor13a, Bri08, Bri10, Bro08, Bud07, Bur09a, Bys02, CF07a, CKB^{+}08, Cap08, CS08b, Cer10, CKT06, Che10b, Cho13a, CBDH08, Cho09a, Cho10b, Cho12, Cox08, CC13, CJJ^{+}02, Dav09a, Dav04a, Dav11, DeN09, Dil02, Din12, DS05b, Efr11, Efr12, EWF^{+}02, Els09, Fau09, FDH04, FWF^{+}02, For13, Fou12, Fry10, Gaz09, Gaz11, GBMH^{+}09, Gup09, Hal09, Har07a, Hav02, Haw01, HLC03, HLVA07, Hil11, Him08, HHNK12, Hjø07, HW13, Huv10, IS12, IA12, Ing12, Jac01, Jam10, JP08b, Jul12, Kar06, KH07, Kaz11, KC07, LYM07, LBC02, Lar09, Lar10c, Lar10b, Lat02, LB08, LSZC13, Les11, LH01]. Information [LHA07, LZC^{+}07, LMH09, Lun09, Luy13, MMC05, MD05, Mac04b, MW10, ME03, MR09, Mar09b, McQ09, Mos09, NCRJ13, OJ04, PPN07, PR07, Pet12, Plo10, Raw10, RS08, Rin09, Ros02, Row09, SR11, Sau09, Sav06, Sav07, Saw10, Sch03, Sha13b, SY07, Shi12, Sma12, SHS09, SWF^{+}02a, SWF^{+}02b, Sta10, SML12, Ste03a, STSG08, SSB01, TWF08, Tom02, TrP13, Tra09, Urq09, Van10, VW12, WY07, Wes09, WvdHH04, WFE^{+}02, Win11, Wor09, WCC10b, Yan10, Yi08, ZI01, ZS08b, AXP11, Ago05, Ago06b, AHH06a, AHH06b, Ago07a, Ago07b, AJ10, ADR11, AMS10, ALT09, AKS11, APSA12, Ano01b, Ano01h, Ano03e, Ano04c, Ano05, Ano10, AK11, Ave13, AHK08, Bai04, BIZGMS10, Bar08a, BT05, BAB02, BAB07, BN08, Bat06, Bat08a, BR13, BJC^{+}04, BMK03]. information [BAT08b, Ben02, Ben06, Ben05, BBMN03, BBMN08, BGBIBM13, BPT07, Bha05, BBJ^{+}06, BP10a, Bil12, BBM02, BP06a, BP06b, BSM^{+}05, BCT02, Bow10, BMTV02, Bro07, Buc11, Buc08, Buc12, BR08, Bur11, Bus07, BH05, Cab12, CS08a, CMRS09, Cas03, CJA^{+}04, CH12, CC01a, CC02, CLZC03, Che06c, CZT06, CLYL06, CAC10, CL08, CK02, CY08, CL03, CLB^{+}05, CLL^{+}07, Col11, CBL^{+}13, Col07b, Coo07, Coo02a, Coo02b, Cop12, CvLB12, Cot02, CDLvR03, CM06, CM07, CM08a, CM08b, Cro10a, CJ04, CG06, DW04, Dai06, Dan11, DLA^{+}10, DRW08, Day05, DJW^{+}11, DK07, Dot13, DHW13, Dow04, Dra02, DBT11, Dum01a, Efr10, Egg02a, Egg07a, Ehi02, ESK06, ERJ07, Enn03a, Exn02, FMM10, FMP01, FFMPM13, FPCB04, Fid13, Fle06b, FM05, Fon07, For04a, For04b, Fos04, Fou02, FL01]. information [FJRNV^{+}06, Fro08, Gal06, GOR06, Gar01b, Gaz07, Gen06, Gen12, GRS^{+}07, GN07, GACCNM^{+}08, GLDM02, GPG^{+}11, Gru07, Gu02, GL09, Haa03, Hal04b, Har06a, Har07b, HPGG^{+}07, HT11, HW02, HLP04, Hay07, HDN11, HML02, HL05, Hei06, HH01, Hep04, HV01, HVPLHP06, HVP06, HDG^{+}07, HM03, HD04, Hjø01c, Hjø02b, Hjø02d, Hjø02c, Hjø03, Hjø09c, Hjø11a, Hjø11d, Ho08, Hol01, Hol08, HO09, HCH^{+}05, HCO03, HSCL07, HBSC13, HM09, HWJ10, HD06, Huv13, JJ10, JR10, Jär07, Jay12, JG03, Joh01, Joh07, JBCJ06, Ju07, JLB08, KdHWA04, Kal05, Kal06, Kar07, Kar09, KRS13, KS08, KWS07, Kel06a, Kel06b, KS13, KFW08, KRK^{+}10, KDL02, Kim09a, Kim09b, Koh13, Kos05, Kos04, KT12, KSS03, KJ12, KEA^{+}05b, KO05, LT12, LKK03, Lan06, Lap07, LSC12, LM09]. information [LLL09, LT09a, Lee10, LCB^{+}07, LMRB08, LT09b, Ley05, LV06, LWY01, Li09b, LB10, Lim07, jLB05, LB13b, Lip03, LLA08, LC05b, Liu13b, Liu13a, LB03, LC05c, LS05, LY11, LS11d, LMNS13, Lyn01, Ma07, Ma12a, Ma12b, Mac04a, Mai13, MBG^{+}07, MBC08, MMCC04, Mar07d, Mat09, Max04, MPT09, MS09a, MMBF09, McK04, MT03, MS05, MS09b, MBBS02, MO04, MA12c, MDM06, Met07, Mil10b, MB13, MEH07, MC03b, Mor10, MLO02, Muk09, MPR^{+}03, MPR^{+}05, Nah07, NSQI03, Nat12, Nay07a, New01, NHL^{+}10, NH12, O’B07, O’C13, Oak09, Oh12, OV01, Opp07, OLLL10, ÖAU^{+}04, Pal08, PM11, PBS^{+}04, PLBS06, Pet07a, PM01, PDU02, PGR^{+}13, RR06, Rab08, Rab09, RHWB11, Rie02, RB08, RHS10, Rob10a, Rob10b, Ros06, RU07, RLvR02, STS12, SHO11, San02, Sar07a, Sar07b, Sav09, Sav11]. information [SH07a, SS09, SEDC11, Sch01, Sch02, Sch05, SDJB12, SV05, SM10, Sha13a, SKM01, SBIS11, SD03, Shi08, SYNC08, Smi04, SB03, SOO02, SC06, SV03, SP07, SM03, Sti01b, SS06, Sto10, SGTS07, SMY09, SC11, SK06, Sun06, SK13, TJLJ07, Tan07a, TTT12, Tan07b, TDC13, TTS13, The01, Tho12, TKB05b, TRJ05, Tom01a, Tom05b, Tom01b, TD02, Tra08, UZ04, UR07, VS04, Vak08, VH05, Vei09, VG07b, VT13, Vit08, Wal11a, WCCS13, Wan07, War02, WP04, War07a, War07b, War07c, War07d, War07e, War08, WRB^{+}09, Wes07b, Whi02a, Whi03b, WJR05, Whi06, WR06, Whi07a, Whi07b, Whi02d, Wij12, Wil02, WmcsW09, WH09, Wor08, WL03, WP05, WLC09, XR07, Xie11a, XC06, XTY06, Xu07a, XL07, Xu07b, XY08, YL06, YHHL12, YBC^{+}06, YH04, YR08, YR02, YB10a, YB10b, YZLC13, Zac05]. information [ZIB01, ZC01, Zha04, ZZZR04, ZB07, Zha08, Zha10b, Zha12, ZYH13, Zha13, ZS08a, ZZ03, Zin07c, Zin07b, Zin07d, dMFRN^{+}10, vT08, Ano02d, Ano03f, Bro01a, CKT06, CPO01, CN07, CLL^{+}05, FM05, Fur10, Fur11, GL12, LB13b, LMH09, MSYD11, RW05, Rie04, SSM^{+}08b, Sri07, SNRB11, SLR^{+}11, SKWH07, Zin07a, Gar01a, Ago05, And02, Bra11, Buc10, Des07, Enn07a, For13, Gra05, Han01, Hea05, Kos09b, Lar10b, Mil12, Pee12, San10, War05, War09a, Whi06, Zin09, Ack02b, Awa06, Dia02, Sha08, VG07a]. Information-based [HW13]. Information-centered [TWF08]. Information-conveying [LB08]. information-creating [Koh13]. information-related [FFMPM13]. information-rich [Zha10b]. information-searching [Cot02]. Information-seeking [Dav04a, SWF^{+}02a, Coo02a, Fos04, KWS07, Kim09a, MT03, NH12, PGR^{+}13, Ros06, STS12, Whi02d, WmcsW09, YR02, YB10a, YB10b, Zac05, ZZZR04]. information-use [BMK03]. Informational [LKK03, Sto11, Sav09, TS09, Ano02e]. Informed [Haa03]. informetric [Bur05, Egg05a, Egg07c, EG12, HWD03, Ye11a]. Informetrics [YKT03a, BIP04, BIP09, Bur03a, Bur03c, Bur04a, Bur13, Egg03, HM13, NHJ03]. Informing [Bur09a]. Infrastructure [Fry10, Mue10, JdMAA12, KT12, Shi08, SML12]. infrastructures [Huv13]. infringement [GKL09]. Ingrid [Spe01]. initial [MH10, ZS09]. Initialism [HKKP13]. Initiating [Son10]. INitiative [KLFG04]. inlink [TSW12]. Innovation [Saw10, CFW^{+}12, LRR12, LS13, Pet02, WL06b, Dot13, Cul06, War05]. Innovations [Ano03f, Che03a, TWW^{+}06]. innovativeness [vRD09, vRC11]. innovators [CM07]. inquiring [Wij12, WL03, WL11b]. Inquiry [Chu09]. inside-out [RHWB11]. insider [WG07]. Insights [GMM09, EF10, BHDK13, PPN07, TTT12]. inspection [WLC09]. inspiration [MW10]. Inspiring [Sau09]. instability [KAD06]. instances [NLC06, NZY13]. institute [FR01]. institutes [LP10]. institution [BGY^{+}13]. Institutional [BdMAM13, Wal07, YS11, AJBAFHCE08, KK08, Lip03]. Institutions [BdMAM13, Ano13b, GRSFV^{+}12b, GSD12, KHC12, WJD^{+}13, YS11, Mue10, Hut08a]. instruction [DJW^{+}11, SDJB12]. Instruments [Col05, Kim09b]. insults [SCA12]. Integer [Win02]. IntegraL [HBSZ13]. integrate [PNHG07]. Integrated [LB11b, TR05, AHK08, BT05, CL11, DZC^{+}09, LB12a, LFC01, LL12, Nah07, RW05, TVF06, YB10a]. Integrating [AKW07, Tad06, ZSR09, DJF^{+}10, MLZ10, DW04, DS11, DSG^{+}12, KdRH^{+}01, McD11, SZ11, SHS09, SC06]. integration [AT02, BPT07, CL03, HBSZ13, HCH^{+}05, LC05a, YLW10, vdHR11]. integrative [CCC10, NY09, WL09]. integrity [Ten12, MC03b]. Intellectual [Esc10b, FARR09, Opp10, SS06, WS12, CM07, CM08b, Him08, Ley08c, LAL09, Tom01a, ZYH13, ZS08a, Zuc06a]. Intelligence [CZH^{+}04, RKBS11, Rok12, CSCC07, Cro11c, EAPL05, Gar01b, Hay07, MMCC04, QZCC06, RM13, Yi12, Ano03g, Ano04b, Che05, HLC05, Tol01b]. intelligent [Ano01f, LZ08b, CLZC03, Fou02]. Intended [vDH12]. intensifiers [dAP13]. intensity [dAP13]. intent [TSA11, VvdHH^{+}13]. intention [CLL13, Cho13a, HBT^{+}09, KY08]. Intentions [vRC11]. intents [LC12, LDvM12]. inter [Ano03d, GCR09]. inter-university [Ano03d]. Interaction [Ago13, BM06, Han01, KMK03, Lar10a, Nay05b, O’B12, San10, Alb10a, BIZGMS10, Bea06, BCT06, EWF^{+}02, FPCB04, Fid13, Ho08, JSK07, Kos05, LFC01, MR08, MBBK07, POC09, PLBS06, Slo02, Tan07b, Tom02, Xie06, vRFBvW10, DS05b, Vil10a, Bia10a, Cop12]. interactional [FAC08]. Interactions [RR12, Cha09, Has03a, JM05, Ju07, KO12, OL13, YS11]. Interactive [LRC13, Vei09, Alb10b, AM11, BSMD11, CL11, CvL12, DBM02, DHKW10, HS01, HCO03, KWR^{+}07, KS13, LB10, SKM11, VS04, WR06, Win08a, Xu07a, XL07, XY08, Bor09, Efr12, Ing12]. interactivity [Lus07]. Intercultural [CS08b]. Interdisciplinarity [MBG03, ADC12, CH12, GBF^{+}12, LG10b, Ley07a, LRC13, SNRB11]. interdisciplinary [AH05a, Hay06, LTO11, LS08, LP09, WWN04]. interest [Bow10, Bud07, LM01, LT13, NSQI03, SL13, WIY01]. interested [AE10]. interesting [MCK08]. interests [Jac01, LfBW08, PCY02]. Interface [Vil10a, Ano01f, CMD05, DP06, Dav06, Fle06b, GRS^{+}07, HW02, JG05, Kar03, KC12, Li09a, Mos04, RHWB11, WR06, Wil06a]. Interfaces [Bea06, Dav06, Gre06, Hen06, JBCJ06, Kal06, KSM06, Moc06, RV06, Ros06, Rot06, Sac06, VR06c, Wil06a, BMTV02, FCM04, JLB08, Mar02, RH12, WmcsW09, vdHR11, Efr12]. intergenerational [LBNB04]. interjournal [ZL07a]. intermediarie [WL11b]. intermediary [EWF^{+}02, RMC12, WL03]. intermediary-based [RMC12]. intermittent [iCG09]. internal [Hav02, MC03b, OM13]. International [Ano01d, Ano01e, Ano02a, Ano03h, Ano03f, Ano08c, BD08, Bra09, CRBPdMAM12, He09, Sin11, Ano01b, BOGdMA13, BD05, DW04, HTC11, HDB07, LS09, MHR13, SS04, WT12, Ano01c, CKT06]. Internet [Enn03c, Gaf07, Gon08, Kap09, Les11, Mis09, Nay04, Tol01b, Wes07a, Wil06b, Wri10, Yeo11, vD10, Bar06a, BF08, Bea06, BLA07, BPJR07, BR08, Bur02a, CJA^{+}04, CLZC03, CL06, CL09, CZH^{+}04, DJWR02, DAB^{+}07, Eas02, FS07, FMM01, FMM05b, FF04, FC01, Hon12, Hov13, JBM04, Kup06, Kwo05, PZ12, PDU02, PSWL13, RB08, Sav11, SD03, Slo07, Vei06b, Wij12, Zha01, Hal01, Ago04a, Dim09, Fah02, Fry10, Mue10, Sti03, Tom01a, Van03]. Internet-based [BLA07, CL09, Zha01]. Internetworking [jL05]. interoperability [ZC04, Cha13]. interoperable [MV08, QH06]. interorganizational [TPLC05]. interpersonal [KK10a, Rad06, Tan07a, XTY06, YRXS11]. interplay [YRXS11]. interpret [YCKW09]. interpretation [DGKS04, Egg05a, Iva01, Mor10, WCMK^{+}12]. Interpreting [NC13, Sav09, The06]. interpretive [Yan01]. interrater [Oak09]. Interreg [BLA07]. interrelations [GOR06]. intersection [Bir06]. Intertextual [MR09]. Intertopic [OLLL10]. interventions [BDH12, Huv11, JLR04, RPWM06]. interview [Har08b, KWR^{+}07]. interview-based [Har08b]. interviewing [Meh06]. intimate [HW13, Wes09]. Intradisciplinary [FN08]. intranet [Ste08]. intrinsic [YD13]. introduce [HWJ10]. Introducing [LS05, ZS08a, Pet07a]. Introduction [AH05b, Arg06, Bar05a, Bar05b, Bar05c, Bar05d, Bar06j, Bar06k, Bar06l, Bar06m, Bar06n, Bar06o, Boy05d, Boy05e, Boy05f, Boy05g, Boy05h, Che03c, Che05, Dow04, Fen05, Hem05, Jör01, LYM07, MD02, Men03, Spi02, SC04a, SC04b, TV04, YL06, Che03b, CDLvR03, HVP06, Mos04, Iye01, Ave13, Fou12, Lar10c, Yak05]. invariant [LP03, LLA05]. Inventing [O’B12]. invention [Exn02, HHHR13]. inventor [LKZC11]. inverse [LB03]. investigate [Suz09]. Investigating [CLB^{+}05, GD03, HDG^{+}07, SRS^{+}13, WCCS13, YB10b, GALC11, LGCA10, Van10]. investigation [AXP11, Gen12, HFW09, KBG12, Lai09, Nat12, Rad06, RPG^{+}08, Sch08, SH07b, TH03, WZ02, Wor08, YSD^{+}08]. investigations [CLL^{+}07]. investment [Cam10a]. investment-benefit-yield [Cam10a]. Investor [O’C13]. invisibility [Gor12]. Invisible [Che09, CSL04, Zuc06b]. invited [JLK09]. IP [ZL07c, DS03]. IP-level [DS03]. iPod [Vei06b]. IQp [AL08]. IR [BR11a, Bor03, GOR06, KJ02, Kle08, PJ06, San08, VŽ08]. IR-rational [PJ06]. Iranian [OM08]. iRANK [WLL10]. Iraq [RB08]. Ireland [War07b]. IRMA [Ano03f]. Irrational [PJ06]. Irving [Gar01a]. iSchools [WS12]. ISI [Ano04b, BL09, Col07b, LR09, LdMAGB10, SM05a]. ISI-2004 [Ano04b]. ISI-ranked [Col07b]. island [Ano03d]. ISP [Kra02a]. iSpace [Cro13f]. Israel [BIA12, BAB07, YH04]. Issue [Ano01g, Arg03, Bar06b, Bar06c, Bar06d, Bar06e, Bar06f, Bar06g, Bar06h, Bar06i, Bar07a, Bar07b, Bar07c, Bar07d, Bar07e, Bar07f, Bar07g, Bar07h, Bar07i, Bar07j, Bar07k, Bar07l, Bar07m, Bar07n, Bar08b, Bar08c, Bar08d, Bar08e, Bar08f, Bar08g, Bar08h, Bar08i, Bar08j, Bar08k, Bar08l, Bar08m, Bar08n, Bar08o, Boy03a, Boy03m, Mac04b, Ano01a, Ano01h, Ano01f, Ano01i, Ano02c, Ano02b, Ano02e, Ano02d, Ano02f, Ano03e, Ano03g, Ano04c, Ano05, Bar05a, Bar05b, Bar05c, Bar05d, Bar06j, Bar06k, Bar06l, Bar06m, Bar06n, Bar06o, Boy01g, Boy01h, Boy01i, Boy01a, Boy01j, Boy01b, Boy01c, Boy01d, Boy01k, Boy01l, Boy01f, Boy01m, Boy02a, Boy02b, Boy02c, Boy02d, Boy02e, Boy02j, Boy02f, Boy02g, Boy02k, Boy02l, Boy02h, Boy02i, Boy03c, Boy03d, Boy03e, Boy03f, Boy03l, Boy03g, Boy03h, Boy03i, Boy03j, Boy03k, Boy04a, Boy04b, Boy04c, Boy04d]. issue [Boy04e, Boy04f, Boy04g, Boy04h, Boy04i, Boy04j, Boy05a, Boy05b, Boy05c, Boy05d, Boy05e, Boy05f, Boy05g, Boy05h, Che03c, Che05, CDLvR03, Hem05, KBP10, MD02, Spi02, TV04, TPF06, TVF06, TB13, Bar08d, Bar08n, Bar08o, Boy01e, Boy03b]. issue-related [TB13]. Issues [CH13, Kos09b, McC12, LB12c, PLBS06, SB07b, ZC04]. ITCC [Ano01e]. item [Kil01]. items [ACHVH10, Egg10d]. iuc [Ano03d]. Ivana [Ano03d].

J [Ano06b, Fou01a, Hol01, Pet02]. Jaccard [Ley08b]. Jakub [VT13]. James [Tol01b]. Janet [O’B12]. Jannica [Jul12]. Japan [LS09]. Japanese [AFR07, AFR06, CTS04, KOS09a, TK06]. JASIST [Ano08a, Mac04b, Ano01a, Ano01g, Ano01h, Ano01f, Ano01i, Ano02e, Ano02d, Ano02f, Ano03e, Ano03g, Ano04c, Ano05, Arg03, CH13, Che03c, Cro09e, Dil02, HS02, Hjø09b, Kra02b, Kra03a, Kra03b, Kra04, Kra08, KE08, MD02, SL10]. Javascript [Bro02]. JCDL [Ano04a]. Jeremiad [Wil05]. jew [BI06]. Jewish [BAB07, Wei10]. jigsaw [CPO01]. Jihad [CCQ^{+}08]. job [rPLM09]. Joint [Ano04a]. joke [ST09]. Jones [RT08]. Jonkers [Ano12]. Jottings [Wil05]. Journal [Ano06b, Ano08b, Ano10, Ano11, Ano12, Ano13b, Ano13c, Ano14, AFR07, BL09, JP08b, LdMAGB10, NSLD13, Rob10b, Tan07a, WCC10b, Åst07, Ano08c, Ano13d, Ben08, BD08, Bro03b, Cam10a, Cam11, Che08, Col07a, CSL03, CF06, Egg01, FARR09, Fra09, Fra12, FW09, GRSFV^{+}12b, GACCNM^{+}08, HM13, HvdW09, HW11, IS06, Jar11, JCDR^{+}06, LM11, Ley02, LB05, Ley06, Ley08a, LS08, LP09, Ley09, LO10, LZB13, LRC13, LYJ^{+}10, McD07, MvL12, Moe05, Moe11, MO08, Mor05, MG07, RR12, ST08, SB12a, SSPA10, SK01, SK02a, Tre12, yT04, Vau03, VT03, Wal07, Xia10, ZLXW11, ZZL12, Zha12, ZL07a, ZS08c, Che08, HRK^{+}06, IS06, Ley06, LdMAGB10, NMD09, HS02, LS10b, NMD09]. journal-article [ST08]. Journals [Rok12, BL09, BSP10, BD05, Cab12, Cam10b, Col07b, CM08a, DS03, Fig06, Fos01, GRSFV12a, GN07, GC07, HvdW09, HHWD05, HRC12a, Hol08, HAT12, KH13, Ley06, Ley07a, Ley07c, LS10b, LHS11, LRC13, LHMW11, LMCRRPA09, LJLlW08, MM01, Men03, Min10, MO08, Muk09, Pow03, PG02, RL09, SS04, SG11, Sha06a, Sin11, SB12b, SH07b, SH07c, TWKH13, TKB^{+}05a, TK08, Tij10, TÜ05, Vau03, Wal06, WL11a, Wal11b, ZSL10, Bro07]. JSTOR [Fer04]. judged [hHyW04]. Judging [Sav11, FG01]. Judgment [Rie02, BD10, Bil12, NKS^{+}06, SM03, XC06, Xu07b, XW08, Zha02]. judgments [BR12a, BJ09, EW10, HB07, hHyW04, HWJ10, MS02, SBA12, SG01]. Julie [Enn01a]. June [Ano03h]. just [MBG^{+}07, Wal11a, Dil02]. Justice [Kaz11, Bri08]. justices [HLR09]. justification [Him08].

K. [Ano12]. Kahin [Tom01a]. Kaj [Fou01b]. Kanti [Sti01b]. Kappas [Ano11]. Karen [RT08]. Karenina [BM12]. Kathleen [Dot13, Orm01]. Katrina [CKF07, Chu07]. Katz [Pet12]. Kebler [SK02b]. Keep [LB13b]. Keeping [Els09]. Kelly [Efr12, Ing12]. Ken [Hal01]. Kernel [LZLC08, XR07]. Kernel-based [LZLC08]. Kevin [Wad01]. Key [FFMPM13, ASQR09, KK10b, LF03, VT03]. key-frame [KK10b]. keyframes [IL07]. keyphrase [MW08]. keyphrases [JP02]. keys [PJ01]. keyword [JLS13, Käk06, NSQI03, VTD10, Wil10b, OA12]. keyword-in-context [Käk06, Wil10b]. keyword-in-context/permutation [Wil10b]. Keywords [GLAA07, BBI12, OTP13, ZWW09]. Kids [Bil12]. Kilicoglu [Ano06b]. Kim [BI13]. kind [Buc12]. kindred [FT12]. Kling [Day07]. KMS [AT02]. know [BD07a, Nea11]. Knowledge [AT02, Ano03h, Baw10, Bla02, BDBA13, Cox10, For13, GD07, Hal10, HH07, KT07a, LR10, McI02, McI10, MMBF09, SCZZ13, STS02, Yag06, Zin07d, Ack02a, AH05a, AB09, And02, Ano02b, Ano03b, Ano13c, AD04, BDH12, Beg05, BPT07, Bis11, CG08, Che03a, Che04, Che06c, CLL13, CCC10, Chu05, CKF07, Chu07, Col11, CBL^{+}13, CH07, Dav02, Day01, Day05, DA05, DAW06, Des08a, Des08b, GTHZ05, Hai12, Has03a, Hay06, HFW09, HL05, HDG^{+}07, Hjø11d, Hjø11c, HJ04, HBSC13, Jap07, JW09, Ju07, KK10a, KKGB11, KTW05, KS10, KH09, KY08, LT12, Lai09, LR02, Ley04, LS13, LKZC11, LLC12, LFLP13, NC13, Oh12, OV01, OMS^{+}08, PNHG07, PDJCRBBM12, RBLH13, Šau04, Sch04a, SXCC07, Sre02, Ste03b, Sto11, SC11, TPLC05, TH13, TS08, Tij10, VRL^{+}09, WG05, WEA^{+}12, Wil04]. knowledge [WL08, WL09, WHC12, YD13, YLW10, YRXS11, YZLC13, ZC04, Zin07c, ZS11b, AH05b, CLB^{+}05, Day05, GRSFV12a, LO13, MD02, Cul01, Koe11, Kva10, Pre09, Sti01b, Cal08]. knowledge-based [Ack02a, LS13]. knowledge-sharing [CCC10]. Knowledge-system [GD07]. Known [Kil01, vEW09]. Known-item [Kil01]. knows [Huv13]. Koenig [Sti01b]. Kohonen [GdMA01]. Kong [LY05, CFY07]. Korea [KP11, PBN02, Shi08]. Korean [KC13]. Krajewski [Bla13]. Kuhlthau [CBL^{+}13, Kra02a]. KWIC [Käk06].

L [Fou01a, Tan07a, WCC10b, Jay12]. lab [SHS^{+}11]. label [Nic09, VE13]. labeled [GC07, HC07]. labels [RI08]. labor [War07c, War07e, War08]. Labs [HD04]. lack [Bur13]. lacunae [LJLlW08]. landscape [Cab12, CK03, FMM10, GBMH^{+}09]. Language [Cro12b, FS07, Hol04, TP06, XR07, BAT08b, BWW03, Cha11, Che06c, CK02, CS06a, CKK^{+}02, CSA^{+}02, CTC05, Cro09f, DDO11, DS09, FJRNV^{+}06, HPAS04, KRK^{+}10, Lee10, LLXS13, LR13, NDD^{+}09, NJ07a, NJ07b, PB01, PBS^{+}04, PLBS06, RLF02, SD01, SS04, SJTB13, TJLJ07, TL01, The05, TVY12, VVLB12, WY06, YHHL12, ZZ03, NHJ03]. language-based [SJTB13]. Language-modeling [XR07]. languages [KGL12]. Laptop [Ivo03]. Large [Cam11, ARPS11, ACHVH10, Ano01f, BJC^{+}07, CHdS^{+}11, DGA11, DHKW10, Fra12, HHfHC10, Kis10, KHC12, LKBJ12, LYJ^{+}10, MCDA03, MH04, OIM^{+}11, SM13, SS10, TWF08, TD02, XBM13, YW08]. large-group [DHKW10]. large-scale [ARPS11, Ano01f, CHdS^{+}11, DGA11, Fra12, HHfHC10, LKBJ12, LYJ^{+}10, TWF08]. largest [vR08a]. Last [uHG10, Kil01, ZS11a]. Latent [BdMAM13, Cri11, Efr05, LK10, LC13, RPYZ07, SLR^{+}11, Din05]. later [BHC06]. Latin [CRBPdMAM12, MKMM09]. Laughing [Hut08b]. Laura [Cha13]. laureates [Hen13b, EGR11]. law [Aus01, ELR09, EG12, HW02, Mil10b, Tar08, ZY13, Bur04b, Egg07c, Egg10a, Hub02, KR01, MG07, NGJ03]. laws [Ack02b, ER02, Egg05a, Mil10a, TW03a, YL03b, Ben01b]. layered [FL01, Zha13, vR06a]. LCSH [JTLG12]. LDA [LDS^{+}11]. lead [Jar11, Tre12]. leadership [TLM^{+}12, Hut08b]. leading [Cab12, GN07, McL10a]. leads [KR01]. leakage [ZKSE13]. lean [DHKW10]. learn [Har07a, LHHVH09, WLL10, dCCRdM^{+}12]. learned [LHB^{+}08, TD02, VH05]. Learning [Che10b, Cox10, DJW^{+}11, FK06, Hay06, Hol04, HMFW12, LLXS13, WIY01, WLC09, BM06, CMFZ01, CM02, Che03c, Chu09, Coo07, Dav02, DLVÁ12, DHKW10, Duw06, GBK05, GLLL11, HGGL05, JT10, KLY08, LB08, LZLC08, LKK07, LZ08b, LJLlW08, LGM^{+}09, MBBK07, PŠD08, PGSH06, POMVULMC13, QH06, SBT^{+}07, SDJB12, SHS09, SZC11, TWW^{+}06, WW13, WP05, fBWLBC06, Xie11a, Xie11b, Kal01, Lar10a]. least [ADR11, uHG10]. lecture [LZ08b]. Lectures [Bia10b]. Lee [Gal12a]. Legal [Pet12, CLH13, Him03, Him04, JBM04, Lip03, RKO10]. Legitimate [Soa06]. Leiden [BMD13, WCMK^{+}12]. Lempel [ANdlF07]. lengths [AG02]. lens [CKF07]. Leonardo [Ivo03, LS10b]. LePrEF [dCCRdM^{+}12]. Less [Ekb12, Mos09, ARPS11, CSL04]. lessons [Men03, NS07, AD04, LHB^{+}08, RU07, TD02, UR07]. let [WBB10a, Põd10]. Letter [Bar12, BB13, Bor13c, BL13, Buc08, Bur04a, Bur06b, Cot01, Day10, De 12, De 13, Egg13b, EF10, Fig06, GV13, Hau08, Hjø01d, Hjø02d, Hjø02c, Hjø09c, Hjø09b, Jar11, KLFG04, Kos03, Kos12, Kra13, Lak07, Lak08, Ley07b, Ley12, LB12a, Mac04b, Põd10, Pra10, Pra12a, Pra12b, RGLCTSC13, Sma05, Sre02, Tol01a, Tre12, Vin13, WBB10b, Whi02a, Ley03b, PB01]. Letters [AJR04a, AJR04b, Ben04b, Bor11, Bur03a, Bur04b, Che03d, Chr04, Dav04b, Day03, Fid13, Fra03, Hal04a, Hal04b, Hjø01c, Hjø03, Hjø10b, Lak03, Ley03a, Ley03b, LO11, Moe11, Pil13, Szo10, Vin11b, Wal04b, Whi04, Win02, YKT03a, Zha04, vdB03b]. level [Ame12, BD09, BBB^{+}08, CL11, CvLB10, GG09, Ham11, KB11, Moe05, OCD08, PPH06, RI08, SCC^{+}12, SK01, SK02a, WvE12b, WLLH09, WJD^{+}13, WHP10, YYL13, ZZL^{+}09, DS03]. levels [LT09b, OYM09, PGR^{+}13, SG01, VS04, vRD09]. leveraging [LT12]. Lexical [Men04, AWC^{+}07, Che06c, Gre01c, WL08]. lexicons [KLY08]. Leydesdorff [Ack02a]. Libcitations [WBY^{+}09]. librarian [Ago08, Chu09, Edw07, Enn04, Hol04, Jan02, Ma02a, PBV08, Ste03b]. Librarians [Ago04a, Mer09, hChC05, Sre02, Lap07, Orm07]. librarianship [Ben05, Cly04, Glu08, Opp07, BL09]. librarie [NV09]. libraries [Alb10b, Ano04a, Bol07, BSM^{+}05, CGY09, Chr03, DP06, Enn01a, Erc06, FG01, Fou01a, Har08a, Hol04, JBM04, JZW^{+}11, JCDR^{+}06, JP02, Kap09, KOS09a, KP03, Lim07, LR08, MBG^{+}07, Mel04, NY08, NY09, Sre02, Tar08, TTS03, Whi06, XC09, Yan01, Ano03d, Hjø01a, Sat08, Enn08a, Koc06, Pet07b, Vei06a]. Library [GACCNM^{+}08, Hil11, Whi02a, Abb05, Ago06a, Ano01b, Ano01c, Ano08a, Atk01, BAB02, Ben05, BDS12, Buc08, Bus07, Cas03, CH12, CH05, CGY09, Col07b, CC01b, Coo01, Coo02b, Cop11, CC04, CC13, CM08a, DJWR02, DAB^{+}07, Dav06, Enn02b, Fou02, FCM04, FMCL07, GLDM02, GZS13, Har06a, Har07b, HLLS07, Him04, HBSZ13, HDB07, Jon02, Kim10b, KO12, KP03, KC12, KEA^{+}05a, LSC12, LT09b, Ma12b, MBG^{+}07, MS05, Mit08, Muk09, Nel06, NC13, Nic03a, Nic03b, NS07, OCD08, OJ04, PGSH06, PLN09, Pet02, PDU02, RPL10, RH12, Reu05, Reu07, SLS06, TVFW07, TTS13, Tom06, TP11, TMPP13, Wil00, WT08, Win08a, XW02, Xie06, Zha10c, ZS11b, Ano02a, BC05, FP04, MWG06, MHM12, MSYD11, SNRB11, SLR^{+}11, YC10, Yi10, Mit08, War09b, Win09, Tom04]. LibraryFind [JHW^{+}08]. license [CF09]. Lida [Ano03d]. lies [ACMP11, WBJM12]. Life [YBC^{+}06, AHH06a, AHH06b, BT08, CH13, CLB^{+}05, CF06, Dia02, FFMPM13, KH07, PMBdS10, SK02b, Wal11b, YC11, RPH^{+}12, Nay04]. Lifting [Hen13a]. Lightweight [CHdS^{+}11]. like [HHWD05]. likelihood [BBC09, BEK^{+}01, BWW03, LMV^{+}12]. limitations [GPG13, MKMM09, NMD09]. limits [Him03]. Linde [Pee12]. line [KJ03, MYWA03, Tom04, WM07]. Linear [Efr10, LKL10, Hu09]. lines [WV13]. Ling [Ano13c]. Ling-Ling [Ano13c]. lingual [LCR^{+}05]. Linguistic [rP08a, rP08b, FCjM12, HV01, LC05a, Pom05b, Vau06]. Linguistics [War07d, ŠKL08, War07a, WHP10]. Link [CCG^{+}06, GN07, SKS12, TSW12, ACH^{+}04, Cas08, FL01, LNK07, NLC06, SXCC07, SCC^{+}12, The06, Vau06, YL09, ZY13, Ben06]. Link-based [CCG^{+}06]. link-prediction [LNK07]. Linked [PGFMP13, GZS13]. Linking [Pow03, GRJS01, RFV10, SLSK09]. links [BI06, CS04, Col05, JR06, Luz09, Mel04, Men04, The01, The02, TH03, VT03]. LIS [Har05, Whi02a, Åst07, Ma12a, MY07, San08, SS04, Sin11, SPH08]. LIS/ [SPH08]. lisa [GACCNM^{+}08]. list [Laz07]. listed [LDM12]. lists [DS10a, HW11, Kup06, Luy12, Mat09, MPR^{+}03]. literacies [Mar09b, Tra08]. literacy [Ago05, DJW^{+}11, GL12, Mat13, Oak09, SDJB12, Xie11a, Xie11b, DeN09, Coo07]. literary [PM11, Pal08]. Literature [DZC^{+}09, Sma12, yT04, AŞT06, Ano12, Bor13a, Bro03b, Che06b, CHMS13, Chu01, CvLvR10, CSL03, Cui10, CBS10, Dav09b, Gau09, Gen06, Haw01, HLC03, HSSV07, JD12, Kos12, LB13a, LAG08, LKZC11, LJLlW08, Mor05, MG07, NHCW13, Pet13, RPWM06, Sar07a, Sar07b, ŠKL08, SST06, SK01, SK02a, SK02b, TÜ05, TMPP13, UGF08, VBMM11, WKdJvdBV01, Win11, vdEvMK^{+}04, JLB08, Ruc10]. Literature-based [Sma12, SST06, WKdJvdBV01, vdEvMK^{+}04, Ruc10]. literatures [SH07a, SSB01]. little [AE10, PN06]. live [Mon04, The08a, Tom05a]. lived [Ten12]. Lives [Les11]. living [Mil12]. load [Gwi10]. Loads [Hu09]. Local [LR11, WT08, Bis11, CH05, Enn06b, LR08]. Locality [ZQI^{+}10]. Localization [Abd04, CTS04]. located [Sel05, YRXS11]. Location [Bis11, WA03]. Location-based [Bis11, WA03]. locations [Hen12]. Loet [Ack02a]. Log [The09a, BJC^{+}07, Dav04a, HS01, HTdV11, HPAS04, HNJT06, HHvdHdR10, MHL05, Ste08, WCCS13]. logarithmic [LB06, Mil10b]. logic [Ano01h, Kar09, ZL07b]. logical [CDLvR03, LB03, Nie03, ŠKL08]. logs [CFS05, Har02, HCO03, SY07, WEF^{+}07, YBC^{+}06]. Long [JLR04, LAG08, LB12c, LP07, McD11, Ten12]. long-lived [Ten12]. Long-term [JLR04, LAG08, LB12c, McD11]. LongHouse [Lar09]. longitudinal [BIP04, Cot02, Koe02, Mar02, PV03, Per03, RU07, UR07, Wal11b, WBY03, WRB^{+}09, Xia10, KP11, WWN04]. Looking [Ros02, Jul12]. Lorenz [Bur07, Rou07]. losers [DC09]. losing [Har07a]. Lotka [Bur04b, ER02, Egg05b, Egg10a, ER12, Hub02, KR01, MG07, NGJ03]. Lotkaian [Bur13, Egg05a, Egg05c]. low [Hol01, PT05]. low-income [Hol01]. Lowy [Pon01]. LSA [JBLOHM10]. LSI [ZA13]. lucky [Kal06]. Luhmann [Ley10]. Luhn [Los01, Wil10b]. lunch [SM13]. Lunenfeld [Ben01a]. Lyn [Ave13]. Lyndon [Tom01b].

M [Ano06b, Ano11, Enn01a, Rob10b, Sch01, Wil10b]. Machine [Duw06, Che03c, DLVÁ12, DS10b, FCjM12, HKKP13, Hem07, POMVULMC13, fBWLBC06, Hai12]. machine-learning [DLVÁ12]. Machines [Bla13, DRM^{+}04]. macro [PAR13]. macrolevel [LT08]. macromolecules [ASW05]. macroscopic [The01]. Macs [Pra10, WBB10a, WBB10b]. made [ARC11, Enn08a]. magazine [WLHO11]. magazines [PN06]. magnesium [WKdJvdBV01]. mail [Ano03d, CKR13, KPEO10, Meh06, PN09, PSO06, TLM^{+}12, YZLC13, ZKSE13]. mailing [Kup06, Mat09]. Main [LAL08, BD05, LL12, SS06]. Main-path [LAL08]. mainstream [LMCRRPA09]. Maintaining [DA05]. Major [Mea10, IA12]. Make [McI10, DS10a, Lus07, SZ12]. maker [WS11]. Making [Bri08, Chr03, KPEO10, KTL^{+}13, Mel04, Met07, SM05a, Ago02, All11, Beg05, CL12, FG01, Gen12, Haa03, HZ12, HBSZ13, Huv11, KBG12, LC06, Nic03b, Plo10, RRW05, RBLH13, Sav06, DS03, MHM12, Mil10b]. Malaysia [EAPL05]. Malaysian [Gu02]. malpractice [Kos03]. Maltby [Hjø02a]. Man [HKKP13]. Manage [Ago04a]. Management [Ano03f, Hal10, Koe11, Pre09, Sch03, AH05a, AH05b, AB09, Ano02b, Ano03b, ADD^{+}06, AD04, BDH12, BBMN03, BBMN08, BGBIBM13, Bla02, BJD10, BWD02, BDBA13, CG08, CPP09, Chu05, CKF07, Chu07, Cul01, Dai06, Dav02, Day01, DA05, ERJ07, EZ08, FARR09, GTHZ05, HFW09, Ho08, HJ04, HHfHC10, KRK^{+}10, Lai09, LR02, McI02, MB13, Nic03b, OMS^{+}08, PNHG07, Pet07a, RPL10, RHWB11, SCZZ13, STS02, SXZ11, Ste03b, Sti01b, TS08, TVF06, Tom01b, WL09, YLW10, Day05, FARR09, MD02, Baw10, Bol08, Els09, Jap07, Nel06]. manager [Wad01]. Managing [CS06b, KC09, McL10a, Son10, TPLC05, Wag03, Lap07, McI10, McL10b]. mandate [Lak08]. Manichean [Bat11]. Manifestation [Mor05, MG07]. manifestations [Sar07a]. manifested [CSL03]. manifesto [Jör04, Miz12]. manipulation [Kra13]. Mann [HWD03, Wal04b]. manner [BM11]. Mantel [SB07c]. manual [RKO10, Nea08]. manually [HC07]. manuals [WS06]. Manuscript [She10]. manuscripts [Ano08c, BD08, Cro12c, RGLCTSC13]. many [HW01]. Map [BIA12, DP07, Enn02b, GSB08, Haw01, KB09, LLL09, LR09, NC13, WIZ01, Zin07a, Zin07d]. Mapper [DZC^{+}09]. Mapping [Ano13b, GRSFV^{+}12b, GSD12, LB05, LP10, LB12b, Max04, ZS11b, AR13, BL11, BC09, CLHHVH11, CLHHVH12, Has03a, KB06, KB11, KC09, Ley07c, LHS11, McC08, Min10, SG01, ZLB11, vEWDvdB10, Ju07, KLP09, Sch04a]. Maps [LS10b, GSB08, GBK05, KB11, LS08, LdMAGB10, MvL12, OA12, PDJCRBBM12, RPL10, Sma03, Yi08, BL11, LB12b]. maps-based [Yi08]. Marcella [Hjø02a]. March [RB08]. Marchionini [Ano10]. Marcia [Bat11]. Mark [Cul01]. markers [KAD06]. market [DJWR02, Rab08]. marketing [Enn04]. marketplaces [WL06b]. Markets [Pee12]. Markov [LHA07]. Markup [Hol04, CH07, DPV11, KdHWA04]. Markus [Bla13]. marrow [dMAVQCR^{+}07]. Marti [Efr12]. Martin [Tol01b]. Marx [Day02a]. massively [Win11]. Master [Les11]. Mastering [McI10, Raw10]. matching [GMA07, GOR06, Mus05, NBYA03, PLN09]. Matchsimile [NBYA03]. material [Bla11, GC07]. Materialist [Nay07b]. materiality [Ekb09]. materials [Col05, TL01]. Mathematical [CDLvR03, Egg07c, Egg08b, SR07, Egg02a]. mathematics [BSP10, Ler13]. matrices [LV06]. Matrix [SB07b, SB07c, KB05, LBC02]. matter [Cox08, Ju07, Moe07]. Matters [Rya09, ACMS05, BP10a, Cro12b, Dot13, PZ12, YCKW09]. Matthew [LG10a, Tol09]. maximum [BEK^{+}01, BWW03, LHLL08]. May [Ano03d, Eas02]. Maybury [Cul01]. McCook [Orm01]. me [Cro12a, Cus13]. Mean [BSP10, Bir06]. meaning [De 12, Ley10, Pra12b]. meaningfully [Bor13b]. Meanings [Ma12a, Yoo10]. means [Jam08b, LPSS06, Tom05a]. measurability [AK11]. measure [AJR03, ALW10, BR11b, DLH^{+}08, Day02b, De 10, De 11b, De 13, DK07, ER03, EL09, Egg12, HvdW09, HMFW12, KB06, MG09b, MBBK07, NSJ13, OT10, WL06a, WBY^{+}09, Wu10, ZLXW11, CPP09]. measured [AGOP06, San08]. measurement [Boo01, Bur01, LRC13, PAR13, SLN03, Tol01a]. measurements [FPZAGBdMA05]. Measures [YKT03a, Ame12, BL09, BKRN03, BMD09, BPJR07, Bur03a, Bur04a, CCG^{+}06, CL01, Egg09a, Egg10b, Egg10c, Har02, LR02, Ley05, RM13, SB07b, SB07c, YD09, YKT03b, ZI01, vEW08, vEW09]. Measuring [AJ10, BBC09, Bur05, JdMAA12, Kel06a, Kel06b, KP11, LR13, LW12, MM04, SKS11, Wal06, WOB09, ADV13, BI02, BIP09, HRC12b, MMMM12, MHR13, OL13, SB07b, WW13, YDS11, ZTY12, BL12, Ham07, LAL09, SB07c, SG01]. mechanical [War10c]. mechanism [Lak03, vR08c]. mechanisms [SW12, WR06]. medalists [Hen13b, EGR11]. Media [For09, Ben01a, CLD^{+}13, Ehi02, GLLL11, Hov13, HHvdHdR10, KJ03, Kom11, RFV12, SM03, SM05b, TTT12, Van03, WCLZ01, YKCL12, KJ03, Koh13, Gua11, Kim11b, Nay07b, Ren04]. Mediated [Ess10, LM09, EWF^{+}02, FWF^{+}02, MH04, rP08a, rP08b, SWF^{+}02a, SWF^{+}02b, WFE^{+}02]. mediating [ADR11, YD13]. mediation [Ekb09, OJ04, SZ08]. mediator [Lip03]. Medical [CC09, Plo10, Ano14, CLZC03, Dav09b, FJRNV^{+}06, Gen06, IS06, LRR12, LO13, NDD^{+}09, NS12, RNLdL01, SMAB09, SSM^{+}08a, SS10, SSM^{+}08b, SST06, WEA^{+}12, WS06, ZZZR04, CRBPdMAM12, LS12]. medical-image [WEA^{+}12]. medicine [BP06a, KJ12, SYL^{+}13, BC05]. Medium [O’B12, DHKW10, Enn08a, KBP10]. mediums [Lar09]. MEDLINE [HNB^{+}09, Sri04, TWSS05]. Meet [Hal04a]. meeting [Hal04a]. Meets [Che10b, Efr08]. melody [MO04, PSB04]. Member [SJ10, LSB13, SBIS11, SC11, The08c, XFCR13]. members [HBSC13, PM11, TP06]. membership [VH05]. memetics [ST09]. Memex [Vei06b]. memorial [Wil06c]. Memoriam [Buc13, Cro10b, RT08, Wil06c, Boo13, VG12]. memory [Bla10, DHW13, ERJ07, HBSC13, WTF08, Esc10a]. men [ESF^{+}08]. menopausal [Gen06]. mental [CL03, CLL^{+}07, Coo07, LB08, MBG^{+}07, Slo02, ZC01, Zha08, Zha10b]. mentions [TSW12]. menu [YR02]. merely [Kwo05]. merging [WM07]. merit [MHM12]. Mertonian [HS13a]. MeSH [LO13]. meso [HDN11]. message [Hon06]. messages [CL12, ZLCH06]. messaging [LCC07]. Meta [Hjø02d, Hjø01c, JLB08, Liu13b, Whi02a, CFCZ01]. Meta-analysis [Hjø02d, JLB08, Whi02a]. meta-search [Liu13b]. Meta-searching [CFCZ01]. metabrowsing [WvdHH04]. Metadata [AKW07, BZC07, EM07, Fat07, Jör07, LKK07, Mar07c, ÖAU^{+}04, RJ07, SB07a, SYNC08, CdSG^{+}09, Gre01b, HZ07, HSSV07, HDG^{+}07, HJ13, Liu13b, Los13, MMMM12, MMBF09, Rin09, SS13, Tan05, Wan06, WGW12, ZL08, Pet07b]. Metadata-based [ÖAU^{+}04, MMBF09]. metaphor [FMM10]. metasearch [JSK07, JHW^{+}08]. MetaSpider [CFCZ01]. Metatags [DC01a]. method [AK13, BIP09, BC09, CC11, CH03, CFN^{+}10, Gor02, HTdV11, Hub01, Kel06a, KLP09, LZG13, MvL12, OIM^{+}11, OM13, Oya08, PCBL04, RBLH13, SSS13, SRDZ06, Su06, YL05, YKT03b, Zha02]. methodological [GMV13, Meh06]. methodologies [GZS13, HML02, MLO02]. Methodology [CG08, Coo02b, BIP05, CvLvR10, HVPLHP06, MO07, Oak09, Sav06, WvE12b, MO04, SP11]. Methods [Ahm07, BI02, Bia10a, HZ03b, Sch03, AJR04b, Ano01b, Arg03, ASW05, BL09, CH12, CHdS^{+}11, CDLvR03, GLDM02, GZS13, Gre01a, HLC11, HNJT06, KSA09, KH13, LLXS13, Ma12b, MS05, MV08, OL13, OYM09, PGR^{+}13, Sha08, SKTM09, Sta09, TS11, Wil02, WG10, WT08, Win08a, Xie11a, Opp10]. Metric [PZ05, HvdW09, TWSS05, ZLXW11]. metrical [Kle08]. Metrics [AW05, Cro13b, BV08, CCS11, SG11, TW03b, WJD^{+}13, ZTG13]. Mexican [LMCRRPA09]. mHealth [ADR11]. Michael [Dum01a, Sti01b]. micro [Ame12, BD09, Ham11, YZLC13]. micro-blogging [YZLC13]. microbibliometrics [FW09]. microblogging [HZ12]. microblogs [Efr11]. microdata [PGFMP13, RK01]. microlevel [Dav02]. microscopy [HHHR13]. middle [Abb05, CBL^{+}13, YR08]. middle-aged [YR08]. midwifery [McK04]. migraine [WKdJvdBV01]. migration [Kom11]. Millennium [Max04, Atk01, Bro01a, Enn02b]. mind [Koc06, Opp09]. minds [WL03, WL11b]. minimum [CC11]. Mining [Ano01a, BYHM07, HDN11, KC07, LYM07, LC12, LHA07, LZC^{+}07, PR07, RCT05, SY07, WBY03, WY07, YHHL12, CWZ^{+}01, Che03c, CLH13, FG03, Glu01, Iye01, KdRH^{+}01, KHC09, LZC09, LYJ^{+}10, MC12, MLZ10, NZY13, Nic03a, OLLL10, RMC12, RKBS11, Sri04, SL06, SL13, TWU10, WEF^{+}07, ZC08, Ano05, Che04]. MIR [JLCL04]. mirrors [Nay05b]. MIS [SPH08]. Miscellaneous [Tho09]. misconstruction [Bat11]. misfits [vRFBvW10]. missing [Nea08]. mistake [Põd10]. Mitchell [Buc13, Hol01]. Mitigating [FR13]. mitigation [HW13]. Mixed [OL13, Ma12b, PGR^{+}13, WG10]. Mixed-methods [OL13]. mixture [LHA07]. mixtures [KC09]. Mljet [Ano03d]. Mobile [CMRS09, DHKW10, LGCA10, MA12c, NCRJ13, Rot06, SSTM09, ZL07c, Lar10a]. mobilization [CF06]. Möbius [MWG06]. Modal [ÖL06, JLCL04]. modality [dAP13]. mode [Cro13b]. Model [Efr08, HL05, Mac04b, AHH06a, AHH06b, ADR11, BSP10, BBM02, Bir07, Bjö12, BP06a, BCT06, BR04, BP10b, dAP13, CM02, CCC10, Chr03, CL03, DW04, DWA13, Din05, DS11, EAPL05, Efr05, Egg08c, ELR09, Fos04, Fu10, GPG^{+}11, HLR09, HCC04, Hub02, HD06, JCDR^{+}06, JG05, KRS13, Kaz10, Kra02a, KC09, LC06, LDS^{+}11, jLB05, LB03, Ma02b, Mac04a, MBC08, MPT09, McD11, Mor05, MG07, NY09, OJ04, OMS^{+}08, PS03, Pra11a, RTM04, Ror07, SDJB12, SEMK05, SY07, SJTB13, SSPA10, SMY09, Su03a, Su03b, SK06, TWSS05, Vis05, WKK^{+}07, WLHO11, Win08b, WCC08, WCC10b, Zha10c, Zha13, ZA13, dCFLH03, LWW01, NHJ03, SEDC11, Ago06a]. Modeling [AN06, DH13, For04b, HV01, KO05, LM01, LDS^{+}11, MT03, NHCW13, Pat11, Zuc06b, Ben04a, CTZ05, CC02, DP07, HDG^{+}07, LB08, LLXS13, LT13, MH04, ÖAU^{+}04, Tan04, WOB09, XR07, GO03, PMBdS10, Zac05]. models [APLG13, BBC09, BWW03, BR12b, CL08, CLL^{+}07, CTC05, Coo07, CJ04, Efr10, GD07, HPAS04, HWJ10, Lak03, LHA07, Los01, MBG^{+}07, QCL^{+}12, Sim07, Slo02, SH07b, The02, UGF08, WZL09, WmcsW09, Win08a, Ye11a, Ye11b, ZC01, Zha08, Zha10b, ER08b, Met07, Nic03b]. Moderating [KT12, MDM06, ADR11, Ano13c, LY11, WHC12]. Modern [Son10, Exn02, Mil10a, Sta09, Abd04]. Modernism [Bur09a]. modernity [Day01]. Modes [Mil10a, KBG12, WBJM12, vRFBvW10]. modification [Cho13b]. modified [Sch09a]. Modifying [ZS08c]. modules [PGSH06]. Moed [LO11]. molecular [KJ12, MD05]. moment [BS09, RB08]. momentum [SC08]. Mongolia [Joh07]. monitor [DZC^{+}09]. monitoring [FG03, RU07, UR07]. Monte [Ano03h]. mood [TB13]. Mooers [Aus01]. moral [FMP01]. Morey [Cul01]. morphological [ASAK04, Cui12, FS09]. mortems [PPN07]. most [BK10, PZ12, Tol09]. Motivation [WCC08, WCC10b, SKM01, YD13, SB07b]. motivations [Oh12, Kim10a]. Motives [vRFBvW10]. mouse [Ben07]. mouth [JZSC09]. mouths [Abb05]. Move [Mea10]. movement [Lip13]. Movements [Has09]. moving [ZL08, AKW07]. MPEG [AKW07, BZC07, EM07, Fat07, Jör07, LKK07, Mar07c, RJ07, SB07a]. MPEG-7 [AKW07, BZC07, EM07, Fat07, Jör07, LKK07, Mar07c, RJ07, SB07a]. MST [QCGB^{+}08]. MST-based [QCGB^{+}08]. Mu [Ano13c]. Mu-Hsuan [Ano13c]. much [HO09, NSE^{+}06]. Multi [MPR^{+}03, AK13, CY08, HLR09, JLCL04, KT07b, LHHVH09, MPR^{+}05, MCDA03, VT13]. Multi-agent [MPR^{+}03, MPR^{+}05]. multi-author [AK13]. Multi-dimensional [MCDA03]. multi-disciplinarity [CY08]. multi-discipline [KT07b]. multi-modal [JLCL04]. multi-objective [LHHVH09]. multi-party [HLR09]. Multi-source [VT13]. multiclass [GS03, MH11]. Multidimensional [Goo01, NHJ03, FPCB04, GRSFV12a, NNY10, NS12, SC11, VC06, vEWDvdB10]. multidisciplinary [CL06, LT08, WW03]. multidocument [OKG07, PAL12, WLLH09]. multifaceted [SC12, WLHO11]. multigroup [Kim10b]. multilayered [dMCMNN09]. Multilevel [BMD13, HBSC13, KK10a, PZ12, YCE10, ZTY12]. Multilevel-statistical [BMD13]. Multilingual [QZCC06, CLYL06, CZH^{+}04, VT13, WTLC06, WY06, YL06, YHHL12, Ano03e]. Multimedia [AKW07, BZC07, EM07, Fat07, Glu01, Jör07, LKK07, Mar07c, RJ07, SB07a, LCC07, Mat13, TSJ09, ZL07c, dV01, Jör07, Ahm07]. Multimodal [JG03, BCT02, HM09]. multinational [EAPL05]. Multinationals [Cha11]. Multiobjective [CLMB10]. multipanel [AYX^{+}13]. Multiperspective [TTS03]. multiplayer [Win11]. Multiple [BD07b, Erc06, FCM04, IS06, ALW10, CKH^{+}03, CISH10, Egg13a, HW05, jL05, jLB05, MLZ10, Sch09a, TYS13, WT08, WOB09, YB10a, ZSR09]. multiple-perspective [CISH10]. Multiple-version [Erc06]. Multiplicative [HRC12a]. multipliers [MBC10]. multisession [LX13, MW12]. Multitasking [SOO02, DS11]. multitheoretic [LR02]. multivariate [TSRGJC^{+}13]. Mundane [Dav02]. Murray [O’B12]. MUSART [DBP^{+}07]. museum [Huv13, Jör04, SBBF09]. museums [Enn01a, Mar07d]. Music [Bai04, TC04, Dov04, Dow04, DMR06, HLP04, Kos04, Lee10, LDL^{+}08, PLZA04, UZ04, Win08a, AE10, GZS13]. musical [Sim07, Win08a]. musicians [Win08a]. my [AE10]. MySpace [AE10, The08c, The09b, TWU10, WT10]. Myth [Fle06a, vD10, Enn02a, Lip03, Nay05b]. myths [Wes09].

N [MAR04, PD08, SRS^{+}13]. N-grams [MAR04]. naïve [LMRB08]. naïveté [Bro08]. Name [PSB04, Aks08, ABJ^{+}06, CFN^{+}10, DGA11, GMA07, HKBY^{+}10, PK11a, SZ12, TWSS05]. name-based [HKBY^{+}10]. name-matching [GMA07]. named [CLYL06, SZC11, SR09]. names [Har08b, KC13, NBYA03]. nanoscience [DT13]. nanotechnology [DT13, LKZC11, TH13, ZQI^{+}10]. Naranan [Egg10a]. NASA [KEA^{+}05a]. Nash [McC11]. nation [War02]. National [LS09, NHL^{+}10, AD09, HLC11, Li09a, Mar02, SL12, SHS^{+}11, TW03a, BC05, MHM12, RW05]. nations [SM13]. Natives [Det11]. Natural [NDD^{+}09, SD01, DDO11, FMM10, KRK^{+}10, LAGVG06, LG10a, Lee10, RLF02, TL01, ZZ03]. natural-language [TL01]. naturalistic [TS01, Tan07b]. nature [dMCMNN09, Iva01, Luy11, Mar07d, McI02, RR06, Sar07a]. navigation [AFMM02, BWS^{+}10, CM02, HW02, LPSS06, MHL05, RTM04]. navigational [OAU11]. navigator [Sre02]. necessary [Alm13]. Need [For13, Kar07, BMD09, Bor11, CL03, CLL^{+}05, Col11, FDH04, GPI03, Kal05, LY11, Luy13, Nic09, PNHG07, Sch13a]. need-based [FDH04]. needed [HW01]. Needs [Ivo03, Jul12, AHH06a, AHH06b, Hep04, HWJ10, Lee10, McK04, VH05, Whi02c, WLC09, YC11]. negation [dAP13, DLVÁ12]. negative [BL12, FS10]. neglected [Hjø01c]. negotiation [McK04]. Negri [Day02b]. neo [Glu08]. neo-depository [Glu08]. NERA [SR09]. ness [CHH09]. Net [Les11, CFW^{+}12, Ago04a]. NetLens [KPEO10]. Network [Mos09, Per03, dCFLH03, AAN12, Ahn11, Ano02e, BMJARB06, BN08, BGY^{+}13, BDBA13, CS08a, Cha09, CC11, Fra11, GdMA01, HWT^{+}13, HDG^{+}07, HM03, HKC09, KOS09a, KBP10, KT07a, LR11, LS11c, MvL12, Nay06, OCO08, PŠD08, PBN02, PDU02, PDJCRBBM12, SCZZ13, SKS11, SLSK09, SC11, SK13, TYS13, TWU10, TW10, TSW12, Tom05a, WJD^{+}13, WT10, YD09, YD12, YCE10, ZJS09, ZQI^{+}10, POC09, Fan10]. networked [FC01]. networking [HS13a]. Networks [Bra09, Dot13, McI10, Yeo11, Che08, DYFC09, Din11a, Din11b, DC09, ER03, Fra12, Hay06, Hen12, HRC12b, KKGB11, KKF^{+}12, KS10, KS01, KY04, Kwo05, KO05, LP10, LBS11, LNK07, LZG13, LGM^{+}09, MCK08, Pep11, QCGB^{+}08, RHCP12, SGJ13, SZ11, SKM07, SKS12, SJTB13, SXZ11, SNRB11, TLM^{+}12, The08c, Vei09, VL13, WGD09, Whi03b, XZW13, YS11, YDS11, YD12, ZQI^{+}10, ZY13, vR06a, KMK03, Sta10, Soa06]. neural [LGM^{+}09, OCO08, ZJS09]. neutrality [CFW^{+}12, Lip03]. Nevada [Ano01e, Ano03h]. Newby [Bur04b]. newly [MSC13]. News [SKWH07, CKH^{+}03, Fat07, OL13, PCR13, SC08, SSS13, SDWG01, SCA12, WZL09, YC02]. news-aware [SC08]. newspaper [NB06, SSTM09]. Next [Shi08, Lak07, RKBS11]. Nicholas [Bar01a]. Nick [Day02a]. Nigerian [Ehi02]. nine [BMD08, Win08b]. NLPIR [ZZ03]. No [BR13, LS12]. Nobel [EGR11, FN13, Hen13b]. Nodes [GO03, BYN02]. nominal [MAA08]. Non [CW04, HTT07, Iva01]. non-banner-ads [HTT07]. non-Gaussian [Iva01]. Non-word [CW04]. nondomain [Dra03]. nonidentical [CS06b]. noninformetric [Egg01]. nonlinear [CL09, Fos04]. nonpoetic [FT12]. nonprofit [BIA12]. nonprofits [Enn01a]. nonsponsored [JR06]. normal [JLR04]. normality [TDC13]. normalization [HAT12, Ley08b, LS11b, LB11a]. normalize [vEW09]. normalized [Ley07b, LO10, LRB^{+}13, LZB13, Moe11, WvE07]. Normative [BBC01]. norms [Ami01, Egg09a, LB13b, PSWL13]. North [SLR^{+}11]. Norwegian [ARPS11, LR08]. not-relevant [MS02]. Note [Egg13b]. Notes [Lat02]. notion [FM05, HS13b, Sar07a, Sar07b]. notions [GO03]. noun [FR08]. nouns [BAT08b]. novel [AFMM02, Ben02, CH03, FMM13, HASAT^{+}10]. Novelty [ACK12, XY08]. novice [CTZ05, CLL^{+}05]. NSF [MHM12]. nuclear [KKF^{+}12]. nuggets [fBWLBC06]. number [BD07b, BB03, hHyW04, KHC12, OIM^{+}11, Sch10, SSPA10, SK02b]. numbers [AG02, ACHVH10]. numerical [Fig06, KH12]. nursing [VVLB12]. NWICO [Luy13].

O [Jul12]. OAIS [McD11]. Object [CS08a]. Object-fuzzy [CS08a]. objective [LHHVH09, Hjø07]. objectivity [For04a]. objects [APLG13, BH11, Day11, Ekb09, Huv11, Lar09, QH06, WCLZ01]. Obliteration [McC11, McC12, Wei10]. obscurity [KJ03]. observational [GLDM02]. observations [Bur04b, VL13]. Observing [PBS^{+}04, ALT09]. obsolescence [SK02b, TÜ05]. obtain [QCGB^{+}08]. obtained [ACGL09]. OCA [RSMVULPO11]. occupations [WS11]. occurrence [BR01, LC05a, LV06, Ley07b, Mus05, WvE07]. occurrences [ZWW^{+}08]. occurring [HJ13]. OCR [BIP05, Tse01]. odious [Cro13c]. Odisej [Ano03d]. oeuvre [McC08]. off [ADC10, WM07]. off-line [WM07]. office [BJ09]. officials [CH05]. offs [SL12]. OhioLINK [HNJT06]. Ohlman [Wil10b]. oil [IA12, WKdJvdBV01]. old [Coo02a]. Older [Xie11b, Xie11a]. omission [Hen13b]. omitted [FMM13]. omitted-citation [FMM13]. OMRAS [Dov04]. on-line [KJ03]. One [PGSH06, Cam11, LBMO11, LC05c, TD02]. Online [Cha09, Kva10, MMBF09, TK08, BDS12, BPJR07, CAC10, CLL13, CLL^{+}05, CLL^{+}07, Col05, EZ08, FS11, HH07, HBT^{+}09, HSCL07, HFC10, HWJ10, Kel06a, Kel06b, Kil01, KT08, KTR10, KTA12, LKK03, Ley07c, LBS11, LHB^{+}08, LFLP13, LMV^{+}12, LMNS13, Lus07, Met07, MBBK07, NNY10, OL13, Oh12, POC09, Pow03, Rab08, RS10, Smi04, SXZ11, SRV07, SMY09, TH04, Tsi13, WG10, Win11, XZW13, YRW07, ZLCH06, Dov04, LWY09, Har08a]. only [Cam11]. onscreen [BLH^{+}10]. ontogeny [Ten12]. ontological [KTL^{+}13, KHAN08, TP11, TMPP13]. Ontologies [AB09, Alm13, CS08a, Cui10, FM05, Fon07, SPR09, XBM13, ZBK09]. ontologization [SR11]. Ontology [WTF08, YWC08, ZC08, Ano13a, DJF^{+}10, GMV13, HJ04, JT10, KdHWA04, Kim05, LWH11, LC05a, NZY13, Pat11, WL13, Yi08]. Ontology-based [WTF08, YWC08, KdHWA04, Kim05, LWH11]. ontology-enhanced [LC05a]. Ontology-supported [ZC08]. ONTOWEB [Kim05]. Open [AD04, MCRdMA11, NGJ03, Pel11, Ano13d, BHDK13, Bjö12, BSMD11, CF07a, Dav09b, Enn07b, Fou01b, GRJS01, KH13, Kra08, LB13a, Moe07, Muk09, NOR08, Pet13, Pom05a, SBF13, SW12, SB12a, SB12b, SH07b, SH07c, Wal07, WL11a, Xia10, CF09, MWG06, PGFMP13, Cho09b]. open-access [Dav09b, Muk09, NOR08, WL11a, Xia10]. open-set [SBF13]. Opening [VV12, Cha13]. openly [LB13a]. Operationalization [GPG^{+}11]. operator [LT02]. operators [Hoe03]. Opinion [BR13, HS13a, KHC09, Miz12, Tho12, GW12, MLZ10, RSMVULPO11]. Opinionated [SU11]. opinions [KC07]. Opportunities [GPG13, BC05, VW12]. opportunity [MQH13]. Optimal [Gre01a]. optimization [BCT02, LLL09]. optimization-driven [LLL09]. Optimizing [HKLG05]. Options [Mer09, Whi06]. Oram [Enn01b]. oranges [SK12]. orchestration [BD05]. Order [LPGBdMA03, Luy13, Sha07, XW08, CP13, Cri11, Egg09b, HD04, hHyW04, Suz09]. Order-based [LPGBdMA03]. ordering [Los13]. ordinal [HV01]. ordination [SB07b, SB07c]. organism [Cui12]. organisms [Win08b]. organization [Ack02a, BIA12, Bar08a, Beg05, LT12, LH01, Ley03a, Ley08c, Ley10, Lip03, LAL09, RBLH13, Sha13b, STSG08, TS08, Whi06, Yi08, ZC04, vdB03a]. Organizational [LR02, McI10, TWW^{+}06, AF12, BDH12, BDBA13, Dav02, DHW13, HFW09, HL05, Mar02, PŠD08, WTF08, XW02, YWM10, LKK03]. Organizations [Son10, AD09, Bre05, CBDH08, KRS13, RK01, STS02, TS11, YZLC13]. organized [CS04, FL01]. organizing [DP07, GBK05, GS11, WIZ01, Spe01]. orientation [HD06, POMVULMC13]. oriented [DBM02, Mus05, WLLH09]. origin [OV01]. original [War10c]. origins [Gun13, Wil10b]. Orleans [Ano01e]. other [AAN12, CA07, CM11, Egg10b, Enn01a, Lim07, Men03, YD12, ZTG13, Ekb12]. Otlet [vdHR11]. Our [Les11, Him03]. outcome [FDH04, PV03, VH12]. outcomes [ALW10, CK08, DJW^{+}11, PGR^{+}13, SDJB12]. output [GdMA01, HHWD05, hHwC08, LM11, MAL^{+}05, VH12]. outsourcing [SK13]. overcome [MKMM09]. Overcoming [PNHG07]. Overlap [HW03, Egg06, HW11]. overlapping [DPV11]. overlay [RPL10]. overlays [LB12b, LRC13]. overload [KRK^{+}10]. overlooked [LB13a]. overt [Hay07]. Overview [Dov04, Dow04, Jör01, Men03]. overviews [ASQR09]. own [GLDM02, LSB13]. ownership [Sha06b, VC09]. Oxford [Bri10]. oxymoron [Cro13c].

P [WCC10b]. P2P [YNF09]. pace [JBCJ06]. Pacific [Bol07]. Packaging [McD11]. Packard [HD04]. page [CS04, Koe02, Meg01]. PageRank [DYFC09, Din11a, Din11b, DS05b, MPS12]. pages [GOR06, KOS09a, Luy12, PCBL04, TH03, TRJ05, Bro02]. paid [NSE^{+}06]. pairs [Egg08c, FR08]. Paltoglou [Ano11]. Paper [HS13a, Miz12, AG02, Bur02b, Bur04b, De 13, DSG^{+}12, Egg12, LO10, LLG12, Men03, Moe11, ZY13, Bla13, Urq09]. Papers [Ano01a, Ano01b, Ano02c, Ano02b, Ano02e, Ano02d, Ano03d, Ano03c, Ano03e, Ano03h, Ano03f, Ano03b, Ano03g, Ano04a, Ano04c, Ano04b, Ano05, Arg03, Hau05, Kra04, Pet02, ZL07a, Ano01g, Ano01h, Ano01d, Ano01e, Ano01f, Ano14, ASQR09, BL11, Bor13b, BB03, CA07, FW09, LS12, LC13, MBC10, Mar11, MO08, Mor05, MG07, SMAB09, SSPA10, Vin04, Wu10]. Papyrus [War09b]. parable [Hut08b]. paradigm [AFMM02, LS05]. paradigmatic [Whi03b]. Paradigms [Sma03, Ano01i, CCMM02, Che03b, DBM02]. paradoxical [HS13a]. parallel [LY06, POMVULMC13, YL03a, YL03b]. parallelism [CJB^{+}07]. paratext [Ano14, CF06, SMAB09]. parents [FDB^{+}12]. Parsimonious [RM13]. Parsing [Max05]. part [AHH06a, AHH06b, EWF^{+}02, FWF^{+}02, Haw01, HLC03, Kel06a, Kel06b, Kra02a, KW02, Mar07a, Mar07b, NJ07a, NJ07b, rP08a, rP08b, RU07, Sar07a, Sar07b, SB07b, SB07c, SSM^{+}08a, SSM^{+}08b, SWF^{+}02a, SWF^{+}02b, Tan07c, UR07, Whi07a, Whi07b, WFE^{+}02, Xu07a, XL07]. partial [ADR11, YWC08]. partially [MS02]. participant [WW13]. participant-generated [WW13]. Participating [Orm01, AD09]. participation [Cho13a, NNY10, vD10, Gon08]. Participatory [LKG^{+}10, Shi12, MBBS02, Mit08]. particle [LLL09]. parties [RFV12]. partition [CH03]. partitioned [LWW01]. partitioner [ACHVH10]. partitioning [ACHVH10]. partitions [Win02]. partner [HW13, Wes09]. partnership [Chu09, XFCR13]. party [HLR09]. Pask [For04b]. pass [HZ03a]. Passage [MG09a, Liu13a, UGF08]. passages [KZ01]. Past [Bia10b, CS08b, Sau09, Bur09a]. Patent [Ley08c, SLSK09, ACMP11, CS12, GG09, HRC12b, LB12b, LS11c, SL12]. patent-network [LS11c]. patents [HRC12b, Ley04, Che03a]. path [LL12, LAL08]. path-dependent [LAL08]. Pathfinder [Whi03b, QCGB^{+}08]. Patient [KBG12, MMC05]. patients [CK08, Has04]. patron [EZ08, Koc06]. pattern [Sha06a, YWC08]. Patterns [Ack02b, Har09, JBS09, POC09, Ami01, Ano13c, CKT06, Cha09, CC01a, CC02, CH03, Che06b, Cho13b, Coo01, CvLB12, Hei06, HHHR13, HS01, HO09, JM05, LP10, LR11, LMCRRPA09, NGJ03, ÖL06, RFV10, Sav12, Slo02, SM03, TKB^{+}05a, Vau03, VTD10, WBY03, WT08, WWZ09, WHC12, CSL04]. Paul [vdHR11]. Pauline [Pet02]. pay [BD07b]. Pearson [AJR03, AJR04a, Ben04b, EL09, Whi03a]. pedos.hr [Ano03d]. Peer [Cox10, Cro11b, MHR13, Ano08c, ADD^{+}06, BD08, CP13, CvL12, KY04, LSZC13, SL10, Soa06, Sti01a, VE13, Enn01b]. peer-reviewed [CP13, VE13]. Peer-to-Peer [Cox10, ADD^{+}06, KY04, Soa06, Enn01b]. Pena [Orm01]. People [LfBW08, Ago02, Har02, HWT^{+}13, Ma07, PSWL13, RRG11, SH07a, Sun12, VC08, WH09, Zha10b]. people-tagging [RRG11]. Perceived [Vak08, VC09, Huv10, Lai09, LGCA10, NMPBA11, NY08, Ror07, Tsi13]. percentage [Hu09]. Percentile [LB12a, BSMD11, Bor13b, Bor13c, Ley12, LZB13, Rou12, Sch13b, WS13]. percentile-based [WS13]. percentiles [Sch13d]. perception [JZS09]. Perceptions [VŽ08, Aks06, JR06, Kra02a, KW02, RHS10, SH05]. perceptual [FCjM12, RCD08]. Performance [LC05c, SLN03, vR06a, Ale03, AN06, BI02, BIP05, CTZ05, CvLB10, DRM^{+}04, FL01, Hen12, HTC11, JZS09, JLS13, Jar11, KHC12, NKLC09, OJ04, OM13, Pra11b, SSM^{+}08a, SH07b, TLM^{+}12, Tre12, Wor07, YR02, Zha02]. Performance-related [vR06a]. performing [Win08a]. periods [FR05, MBC10, dBvSW10]. Peripheral [SV03]. Perish [De 11a, vDH12]. permutation [Wil10b]. Perseus [Yan01]. persistence [Bar08a, Koe02]. Personal [Els09, Gal12a, SML12, VC08, AYN13, BLT07, Bar08a, BBMN03, BBMN08, BWS^{+}10, BGBIBM13, BJD10, CKR13, ERJ07, GMA07, KS08, KC13, KSS03, MB13, PCR13, Shi12, Sma03, SCA12, YC02, Bol08]. personal-based [KSS03]. Personality [Cop12, NY08, Vis05]. personalization [FMCL07]. Personalized [LHA07, SZ11, DP07, LKK07, LLC12, SSTM09, WCLZ01]. persons [Aks08, CH11]. perspective [All11, BDH12, Bri08, Cab12, Che03c, CISH10, Chu07, HDN11, Hjø02b, hHwC08, LB12c, NNY10, NS12, OV01, PŠD08, PPN07, SEDC11, Sun12, Xu07a, XL07, YCE10, Dav09a]. Perspectives [AKW07, Bea06, BZC07, Bra09, Bri08, Bro08, Cap08, CS08b, Cox08, Dav06, DJZ^{+}09, EM07, Fat07, Fau09, Gre06, Gup09, Hal09, Hen06, Him08, JBCJ06, Jör07, Kal06, KSM06, Lar09, LKK07, LMH09, Lun09, MR09, Mar09b, Mar07c, McQ09, Moc06, RS08, RV06, Rin09, RJ07, Ros06, Rot06, Sac06, SB07a, Sau09, SHS09, VR06c, Wil06a, Ano02b, Ano03b, HSKS03, LRR12, MS02, MKMM09, Wes07a, Ano08b, Cho09a, RL09, SS10, Pet12, Baw10, Kos09b, Mil12]. persuasiveness [LMV^{+}12]. Pervasive [Gua11]. perversion [Cro01]. Peter [Ben01a, Wil10b]. pharmacology [ŠKL08, SG11]. Phase [EM07]. Phenomena [Su06, MAL^{+}05]. Phenomenological [Dav09a, Sav07]. Philip [Joh01]. Philosophical [Mil12, Him08, Ma12b]. philosophy [CSL03, Hjø11b, Sha13b]. phishing [Wor08]. phones [Rot06]. photo [NNY10]. photo-sharing [NNY10]. photographs [SJ10]. photography [HT11]. phrase [PN09]. phrasing [BBI12]. physical [Bil01, Bil02, NMPBA11, Pow03]. physically [GD03]. Physician [GPG^{+}11, NS12]. physicists [Bro01b, Sch08]. physics [BSMD11, CK03, HTC11, LP07, PAL12, SK01, SK02a, GBMH^{+}09]. picture [Rob01]. Picturing [Wil10a]. pilot [Cly04, GTHZ05, PM11, PSB04]. PIM [CKR13]. Piskorski [VT13]. Place [Gua11, Atk01, Cro08, LAGVG06, Orm01, WH09, Cox10]. placement [NSE^{+}06]. Placing [SM13]. Plagiarism [Sta11, APSA12, Cro13c, Lak08, Wei10]. plagiarized [HZ03b]. plain [WRM^{+}02]. Plan [Mea10]. planned [LCW06]. planning [Nic03b]. Plans [Hem07, Lun09]. plant [Cui10]. platform [LM11]. players [BD05]. playing [Win11]. pleasurable [KH07]. Plex [Les11]. PLS [ADR11]. plus [Kil01]. podcast [TLdR10]. poetic [FT12]. poetics [Sha07]. poetry [FT12]. Poibeau [VT13]. polarity [dAP13, NZY13, POMVULMC13, ZC08]. Police [LY05]. Policies [Bra11, CF07a, Jar11, Tre12]. Policy [Bra09, TKV^{+}10, Dot13, HH13, JBM04, JCDR^{+}06, LB11b, LDM12, Max04, RPL10, Sch01, Su12, Cho09b]. policy-history [HH13]. Policy-relevant [TKV^{+}10]. polishing [HB08]. politeness [rP08a, rP08b]. political [Li09a, LB12c, RB08, RFV10, RFV12, Suz09]. political-economic [LB12c]. Politics [Yeo11, Cha13, Des07, Huv11, Max04, Tra09]. Polymath [Gar01a]. Polymath/ [Gar01a]. polyrepresentation [EW10, LIL09]. polysemy [JBLOHM10, KC09]. Pomena [Ano03d]. pool [KSK01]. Popper [For04b]. Popular [PLZA04, BL12, ST09]. popularity [FS11, MSC13, SCZZ13]. populi [WSJS01]. pornography [God07, PSWL13]. Portal [CZH^{+}04, AKW07, CL09, ZZZR04, MMBF09]. portals [BBJ^{+}06, LBR02, LBC02, LBNB04, SBIS11]. portfolio [SLSK09]. portfolios [KS08]. Portolan [ZLB11]. Positional [uHG09, uHG10]. Positioning [McK04]. Positive [Fau09, BL12, Hol01]. possessions [Cus13]. possible [Bar12, Egg13b, TTS13]. post [CLL13, Day01, PPN07, YSD^{+}08, YHWL13]. post-adoption [YSD^{+}08]. post-Fordism [Day01]. post-knowledge-sharing [CLL13]. post-mortems [PPN07]. Postigo [Lip13]. postindustrial [Duf11]. posting [EHD05]. postmodern [Cro01]. postprocessing [CS12]. posts [EHD05]. potential [AT02, KAD06, KO12, LDL^{+}08, SSM^{+}08a, SSM^{+}08b, SSB01]. Potentiality [YKT03a]. potentials [LB11a]. poverty [LB13b]. Power [Mil10b, Rok12, Tho09, Yeo11, ZY13, Ano02e, Egg05a, Egg05b, ELR09, FS07, Fle06a, Hol08, JZSC09, Lai09, Lim07, Mil10a, PJ01, Pon01, PT05, War07e, TW03a, Bra09, Enn07a, Exn02]. Power-law [ZY13, ELR09]. powerlaws [LB06]. PowerPoint [TK08]. practical [Afi03, Ago06b]. Practice [Cox10, Els09, Hal10, Pre09, All11, Bea06, Bur11, CLL13, Fat07, Har07a, Har07b, Hjø11b, Koh13, Kos03, MWG06, Nat12, RS10, ST08, Tom06, Van11, Wil06a, vT08, O’B12, Cox09, Has03b, Koe11, Yan10, Des08b, Gra05, Sag06, Su06, Tad06]. Practices [Bea06, Dav09a, Dav06, Esc10a, Fau09, Gre06, Gup09, Hal09, Hen06, JBCJ06, Kal06, KSM06, Lar09, LMH09, Lun09, MR09, Mar09b, McQ09, Moc06, RV06, Rin09, Ros06, Rot06, Sac06, Sau09, SHS09, VR06c, Wil06a, AAN12, CKR13, CGY09, Cop11, Cro01, KGL12, KY08, LAGVG06, Mat13, Pow03, PGL13, RK01, WZ08]. Practising [Van10]. practitioners [LKG^{+}10]. pragmatics [AF12, Sha07]. Praise [FN11, War10b]. PRAW [SEMK05]. pre [Ano13c, WHC12]. pre-web [Ano13c, WHC12]. precept [Buc11]. precise [Hei06]. precision [Bil12, Eas02]. precompute [dCCRdM^{+}12]. predicate [WS11]. predication [JBLOHM10]. predict [TL05]. predicted [Dan11]. Predicting [Bur03b, CTC05, CC01b, FP04, HM03, NMPBA11, TLdR10, XZW13, YRW07, Kim09a, MSC13, Yi10, ZJS09, WS11]. prediction [HC11, LNK07, LGM^{+}09, SKS12, TCW11]. Predictive [Che12, BSMD11, HWJ10]. predictors [BHC06, CK08]. predicts [LC06]. preembryo [VBMM11]. Preface [BYCMS02]. prefer [Sun12]. preference [BIKLY09, BGBIBM13, BLH^{+}10, LKK07, SM03, TLdR10, WL06a]. preferences [GD01, LWH11, Sav07, SM05b]. preferential [Mil10a, RRG11]. preferred [PK11b]. Preliminary [Ano06b, GZS13, KBG12, ZZZR04, Ano03c, CLL^{+}07, HRK^{+}06, JSI^{+}04]. premier [Hol08]. preparedness [Jam08b]. preparing [PV03, SV05]. preprints [Bro01b, Bro03b]. prescription [Hut08b]. presence [CvL12]. presences [BT08]. Present [Bia10b, Bur09a, GBMH^{+}09, War02]. Presentation [BIKLY09, hHyW04, OAU11, Rab09, TDC13, Xie11a]. presentations [TK08]. Presenting [KH12]. preservation [BR11b, Cop11, LB12c, McD11, Win11]. preserving [LWH11, PD08, Sau09, SRDZ06]. presidential [DC09]. pressures [YYL13]. prestige [FN12, GRSFV12a, TWKH13, YD10, YDS11]. pretext [Wor08]. prevalent [Ano13c, WHC12, vR08a]. prevalently [Hen13b]. primary [PGR^{+}13]. primer [Che06a]. primitives [NJ07a]. principal [Din07]. principle [BM12, Dav08, Enn07b, LIL09, RS08]. Principles [O’B12, Opp10, Tie10, Whi06, Ost08, SBT^{+}07, LBMO11, Hal10, Sti02]. Print [War10d, Chr03, Chr04, Dav04b, Ehi02, MQH13, Vau03]. prior [CS12]. Prioritization [HSJB12, IL07]. priority [SLSK09]. Privacy [Enn07a, PD08, VGM^{+}11, BPJR07, CPP09, LMV^{+}12, NS07, Sch02, SXZ11, SH05, TS06, YRW07, SEMK05]. Privacy-preserving [PD08]. Private [HO09]. Prize [FN13]. Probabilistic [Efr08, NB06, RFQ^{+}05, UGF08, Ben08, BR04, Din05, TWSS05]. probability [Ben05]. problem [Bor13c, Bro08, CLB^{+}05, KS10, LNK07, MTS^{+}13, RRW05, Tho12]. Problems [MM10, KO12, MD05]. procedural [AYN13]. procedure [Fra03, PG02, TP06, YLL06]. proceedings [LLA08]. Process [Vil10b, AHK08, BBI12, Bow10, DBM02, Fei11, iCG09, GL09, Hav02, HV01, Hon12, JG05, Ma07, PPH06, PV03, Tan04, War10a, WL09, XJ10, Ma07]. Process-aware [Ma07]. process-oriented [DBM02]. processes [CLL13, For04b, SOO02]. Processing [Kle08, Mac04b, All11, AYX^{+}13, DDO11, Gre01a, KRK^{+}10, Mac04a, MO04, NDD^{+}09, OYM09, OV01, SD01, SB12b, ZZ03]. Procrustes [SB07c]. produce [BL11, HD04]. producer [vRFBvW10]. product [CHdS^{+}11]. production [AHK08, LMCRRPA09, McQ09, Van11]. Productivity [Kup06, ADC10, ADV13, CvLB10, Egg01, Gu02, HLC11, Hub01, MS05, PAR13, SHS^{+}11, yT04, Vak08, YKT03b]. Products [Vil10b, LBS11, Han01]. profession [BAB02]. Professional [Swa11, Ago06b, HWT^{+}13, HTdV11, SM05b, Sti01b]. Professionals [Sch03, ALT09, HHvdHdR10, JJ05, rPLM09, SHS09, VH05]. Professions [Row09]. Professor [Hjø09c, Hjø10b, YKT03a]. proficiency [LR13]. Profile [BdMAM13, VH05, BMK03, BBB04, Cro09e, GRSFV^{+}12b]. profiles [BdRB^{+}13, NMD09, SDWG01, SM03, The08c, VTD10, ZRG11]. profiling [KdRH^{+}01, SC12, TVY12]. profound [KH07]. program [AH05a, BLA07, LKK07, Pel11]. programming [FFPW04, NGJ03]. programming-based [FFPW04]. Programs [McL10a, HLB05]. project [AYSC06, BNUA04, BPT07, Chu09, CBL^{+}13, Coo02b, Dov04, Kaz10, LLL09, MBBS02, PPN07, TPLC05, Dai06, DP06]. project-based [Chu09]. projects [CN07, MC12, XFCR13, Son10]. prolific [Tol09]. promoting [HFW09]. promotion [Hyl03, Pil13]. pronominal [WL08]. proof [Egg10a]. propagation [XBM13]. propensity [Bir07]. Proper [BAT08b, FR08, NBYA03]. proper-noun [FR08]. Properties [Ben04a, Egg06, ELR08, Egg10c, HHWD05, KEA^{+}05b, RJ08, Rou12, vR06a, vR06b, vR08a]. Properties-based [Ben04a]. Property [Esc10b, Opp10, BBC09, Enn07a, Him08, SS06, Tom01a]. Proposal [Bur09b, Lar09, MM04, Miz03, PV03, RY08, SV05]. proposed [LFC01, Sch12]. propositional [PS03]. Pros [Mer09]. prosocial [AE10]. prospective [Gar01b, Sav09]. Prospects [Hol01, War07b, Cho09a]. protect [NS07]. Protection [JBM04, BNUA04, BPJR07]. proteomic [Bro03a]. protocol [BT05, LF03, War06]. protoforms [Yag06]. prototype [Kos05, LBNB04]. Prototypes [CMRS09]. provenance [Lyn01]. provide [HBSC13]. providers [SS06]. providing [Sti05, Tom02]. proximal [BYN02, Wor07]. proximal-virtual [Wor07]. proximity [DH13, LH11, LPSS06, NSLD13, SB07b, SB07c, VC08, YHWL13]. proximity-based [LH11]. proximity-driven [LPSS06]. pseudo [LKJ08, YHL11]. pseudo-relevance [LKJ08, YHL11]. psycholinguistic [TVY12]. Psychological [Wal11b, BLH^{+}10]. psychology [CLB^{+}05, CLL^{+}07, CSL03, Day05, Wal06, Wes07a, GP10]. psychometric [Kim09b]. psychosis [Cha09]. PsycINFO [GP10, YBC^{+}06]. Public [Opp09, TW10, AD09, BP06a, Bud07, Cas08, CLH13, Cop11, DJWR02, DAB^{+}07, Dot13, Har07a, HO09, JBM04, KOS09a, LF03, LR08, Max05, MC12, NV09, PDU02, RK01, RRA^{+}13, SWJS01, SH05, TP07, WG10, WSJS01]. public-access [MC12]. public-key [LF03]. Publication [Ano14, KH13, Rok12, SB12a, Ame12, BGY^{+}13, Bjö12, BD07b, DO01, GP10, GRSFV^{+}12b, Gu02, Hau05, Hol08, HO09, HNJT06, Hyl03, KGL12, MO08, NMD09, PGFMP13, SHS^{+}11, Vak08, VL13, WvE12b, Ano13d, SMAB09]. publication-level [WvE12b]. publications [ADC12, APSA12, AYX^{+}13, Cly04, CvL12, DO01, FR01, Gor12, JSI^{+}04, Kno06, KTA12, LR10, Sch09b, Sch13c, Sel05, Vin11a, WBY^{+}09, Win08b, ZS07]. Publish [De 11a, SL12, PZ12, vDH12]. publish-or-perish [vDH12]. published [Ano08c, BD08, MBC10, SSPA10, WV13, Wal11b]. publisher [SM13, Sma05]. publishers [DP06]. publishing [Ano13d, BH11, BSMD11, Dav09b, HTC11, KSF04, Min10, Miz03, PAR13, SB12a, Tom01a, Wee07, WT08, Xia10]. PubMed [KSF04, LRR12, MTS^{+}13, Tan07b, VVLB12, VRL^{+}09, YCKW09]. Pugh [Tom01b]. Punched [Cer10]. Punched-Card [Cer10]. Punishment [WG07]. Purchase [Mea10]. purpose [CW01, TW10]. purposes [BMD08, PBV08, RK01]. Putting [CLL^{+}05, Nel06]. Puzzle [Efr08, LF03].

Q [LK10]. Q&A [Gaz11, KO09, LFLP13, Rab09, RS10, Sav12, SK12, WK13, YC11]. Q-analysis [LK10]. QA [SKM11]. QE [Gre01a]. Qualification [Ekb09]. qualitative [CD06, Gen12, Har08b, HH07, JG06, Mat13, Meh06, Rad06, Dai06, Ago06b]. qualities [Mai13, NKS^{+}06]. Quality [Hal04b, Lak07, Mat09, Miz03, ZLXW11, ADC10, AK11, CB13, Cly04, Dan11, Fig06, GV13, HVPLHP06, HBT^{+}09, HSJB12, HWJ10, KS08, LDL^{+}08, LS11a, LFLP13, Mai13, MMBF09, OMS^{+}08, Rie02, Sav11, SGTS07, STSG08, SMY09, SJ10, Tol01a, VY12, WL06a, ZZL12, Zha04]. Quality-structure [ZLXW11]. Quantifying [AL08, BOGdMA13, MMMM12, KAB09]. Quantitative [Smi01, The08b, Win08b, CD06, Gre01b, OM13, SC08, SP11, WKK^{+}07, Yoo10]. Quantity [Vin13, LFLP13]. quantization [GBK05, LSZ04]. quantum [PAL12]. quasi [Ekb09]. quasi-objects [Ekb09]. Queens [FDH04]. queries [Bil12, BW06, lC01, CR03, Cho13b, CD06, DS10b, EZ08, Kle08, KY04, Lee10, LS10a, LDvM12, LHHVH09, NJ07b, OCJ12, OAU11, SY07, SWJS01, TS09, WBY03, YCKW09, ZWW09]. Query [EW10, Nie03, SWT04, SR06, Ale03, ACHVH10, BYHM07, BI06, BR11a, BJC^{+}07, dCCRdM^{+}12, CS12, CTZ05, CFS05, CKK^{+}02, CSA^{+}02, DBP^{+}07, DGKS04, ER08b, Gre01a, Gre01c, HMdR11, HCO03, JBS09, KK04, KLAS13, LDvM12, LZC^{+}07, LR13, LT02, MPS12, NJ07a, NJ07b, OLLL10, PSB04, PCY02, SY07, VvdHH^{+}13, WL06a, WTLC06, WLLH09, YHL11, YBC^{+}06, ZWW^{+}08, ZWW09, Cho13b, NHJ03]. query-based [OLLL10]. query-by-humming [DBP^{+}07]. query-oriented [WLLH09]. query-related [YHL11]. query-specific [HMdR11]. query-term [WL06a]. Querying [SM02, JJ05]. quest [Sch04a]. Question [LZ08b, NV09, CDLY11, KWR^{+}07, LWY09, LK06, Pom05b, QCL^{+}12, RFQ^{+}05, Ten12, WKK^{+}07, WvE07, WTP12, Yag06, CB13, LC05c]. question-answering [KWR^{+}07, WKK^{+}07, Yag06, CB13]. Question-driven [LZ08b]. Questions [Har04, Bis11, CDLY11, FF04, Hua13, LJ13, Pom05a, RLF02, YC11, Zha13]. quick [QCGB^{+}08].

R [Ano13b, AFR07, Tom01a, Ben04b, EL09, Lia05]. R-sequences [Lia05]. RAA [BNUA04]. Radicchi [BD09]. Radio [Orm07]. Ramesh [Pet12]. Randall [Fou01b]. random [Ano13a, BW06, SZC11, Suz09, TCW11, WL13]. randomization [Hen12]. range [TW03b]. rank [BSMD11, Bor13b, Bor13c, CM06, Egg09b, HLC11, JLS13, LLXS13, Opp07, Rou12, Sch13b, WLL10, Win02, YDS11]. rank-frequency [Win02]. rank-learn-combine [WLL10]. rank-order [Egg09b]. Ranked [KS01, GRSFV12a, Col07b]. Ranker [LH11]. Ranking [Bil12, BM11, GRSFV11, LS11a, MS05, SST06, BI06, BMD13, CDLY11, DYFC09, EW10, FFPW04, HMdR11, KZ01, KAB09, Lak03, LBS11, LH11, LT13, Meg01, MR08, SM02, SJTB13, Spo08, WCMK^{+}12, WA03, WLLH09, WLL10, WJR05, BdMAM13]. Rankings [BdMAM13, BIKYL07, BS09, BMD13, Mar09a, MY07]. ranks [Cam10b, Ley12, LB12a, LZB13]. Rapid [DSG^{+}12]. rate [BSP10, FMM13, TL05]. rated [TH04]. rates [Aks06]. rating [BIP05, LBS11]. rational [PJ06]. rationale [Egg09b]. rationality [Ago02]. raw [Dav08, TPF06, WGD09]. Rawls [Duf11]. Rawls-Tawney [Duf11]. Ray [Kal01]. Raya [Ago13]. Raynaud [WKdJvdBV01]. RDA [TrP13]. Re [CDLY11, Jam08a, Ano08a, Hen12, JP08b, Ano08b, Ano08c]. Re-estimating [Jam08a]. Re-ranking [CDLY11]. re-sampling [Hen12]. reach [Ano04a]. reactance [SRV07]. reacting [RRG11]. Reaction [Day03]. reactions [FN13, Yoo10]. Read [Sch01]. readability [LMRB08]. reader [SS04, Pet07a]. readership [uHG09, uHG10, KEA^{+}05b]. Readersourcing [Miz12]. readiness [BBB^{+}08]. Reading [Har06b, BMK03, CTC05, JW09, WT08, Bia10b, TKB^{+}05a, Rya09]. Readings [Cox09, O’B07]. ready [FF04, Whi02c]. ready-to-use [Whi02c]. Real [CMFZ01, LT13, AYSC06, CLB^{+}05, KSF04, NSE^{+}06, Wil06b]. real-life [CLB^{+}05]. Real-time [CMFZ01, LT13]. reality [GBF^{+}12, Lak07, SM05b]. really [De 11b, De 13, Egg12]. reasonable [LS11a]. reasoning [Hal09]. reasons [GLLL11]. Recall [Bla10, Coo07]. received [ARC11]. recently [Sch12]. recipient [EZ08]. reciprocal [SPH08]. reciprocity [WK13]. Recognition [BNUA04, LHLL08, Liu09, SD01, SZC11, Van12, SR09]. Recognizing [ASR^{+}10]. Recollecting [Hay07]. Recollections [Gar01a, KP03]. recommendation [LHA07, PD08]. recommendations [LLC12, Met07, SRS^{+}13, BP06b]. recommender [HCC04]. recommending [Fur02, LJ13]. Reconceptualizing [Day11]. reconciliation [vHVD^{+}13]. Reconciling [Ros06]. Reconfigurable [HCH^{+}05]. reconfigurations [Hem07]. Reconfiguring [CC04]. record [Dan11, HHfHC10, Sha06a, SML12]. Recording [Tan05, PGL13, Tan06, Tan07c]. recordkeeping [Sha07]. Records [McL10a, She10, GP10, HW05, NS12, Tan05, Tan06, Tan07c, TWSS05, McL10b]. recovering [TL05]. recovery [Oya08, TL05]. recursion [RJ07]. redesign [JJ10]. Redips [CSCC07]. rediscovery [KBP10]. Reducing [JTDG13]. Reduction [GdMA01, DARK09, Has04, LAL09]. Refereeing [Kra03b]. referees [CP13, Kra03a, Kra04]. Reference [MBC10, Ago04b, AJR03, BCT06, Bor13c, dMCMNN09, EZ08, FF04, Hyl03, Jan02, KO12, Lan04, Luy12, Ma02a, McD11, Mon04, NV09, Pom05a, Rad06, SOH08, SK12, Sti05, Tad06, VRL^{+}09, Yak05, Jan02]. referenced [OTP13]. References [ARC11, AG02, Mor05, SV05, LT10]. Referencing [CvLB12, Hel10, LAGVG06, LSB13, SK01, SK02a]. referral [Dav04a]. Refining [Kaz10]. reflect [WWN04]. reflected [Chu01, TÜ05]. Reflections [AFR06, AFR07, OV01, PLBS06, RK01, O’B07]. reformulation [BMD13, JBS09]. Reframing [SP07]. Regional [Buc11, MKMM09, LS13]. Regions [SG01]. Regularities [BP10a]. regulation [Bow10]. regulative [GKL09]. reinforcement [WLLH09]. reinforcing [vR08c]. rejected [Ano08c, BD08, Cro12c, RGLCTSC13]. Rejecting [CA07]. Rejoinder [AJR04b, Chr04, Ley03b, Zha04, Hjø09c, Ley03a, Ley07b]. relate [RPWM06, YD12]. related [AYN13, CH05, ESL^{+}10, FFMPM13, Gen06, Kno06, KL11, PG02, SY07, TP07, TB13, VC09, YHL11, ZB07, ZWW09, ZBK09, vR06a]. relatedness [KB06, LW12, SKS11]. relation [ELR09, EL09, Egg10b, ER12, LZLC08, RU07, UR07, WvE10]. relational [KBP10]. Relations [Egg05b, Che08, Day11, Egg09a, LC05a, LB05, Ley06, LS09, LP10, NSQI03, Sto10, WS11, YYL13, ZL07a]. Relationship [Zha10a, AG02, BR11a, Ben02, Egg01, EF10, GMM09, HW05, KP11, KHC09, LG10b, Ley04, LLG12, NHCW13, NLC06, Pep11, RHS10, RCD08, Tra08]. Relationships [AAN12, Ror07, ASW05, GCR09, Gre01c, GW12, KHAN08, Li09b, LB10, MG10, Moe05, ÖL06, PSO06, RFV12, SL06, TH03, YS11, YLW10]. relative [BD09, DS05a, GRZ12, LS11b, Lia05]. Relevance [BWW03, Bod06, CG06, FG01, Gor02, HS13b, Sar07a, Sar07b, XC06, Xu07b, BD10, BR12a, Bil12, Bor03, BCT02, CF07c, CC01b, DLA^{+}10, EW10, Hjø01b, HC02, Hjø10a, hHyW04, Hua13, KJ02, KO09, KC07, LSJR11, LKJ08, LPGBdMA03, MS02, Reu07, RLvR03, SS04, Sch13b, SG01, Sto06, TS01, VS04, VV12, Whi07a, Whi07b, XBM13, XW08, YHL11, Zha02, dCFLH03, Luk09]. Relevant [HCO03, MS02, SJTB13, TJLJ07, TKV^{+}10]. Reliability [VvdHH^{+}13, BEES10, Bur06b, Cot04, Oak09, PMK11, TL05]. religious [WCCS13]. rely [CvLB12]. Relying [TKB^{+}05a]. remapping [Whi03b]. remarkable [Van12]. Remarks [De 13, Egg12]. removal [Fra03]. rendered [BLH^{+}10]. rendering [CF06]. repeated [Dav08]. replacing [MQH13]. Replies [Whi04]. reply [De 12, LO11]. report [GZS13, KLFG04]. reporting [MMBF09]. Reports [EZ08, Ben08, KL11, BL09, Che08, Ley06, LdMAGB10]. repositories [AFMM02, HBSC13, KTW05, KK08, MH10]. repository [OW04, She10]. represent [SM03, Yoo10]. representation [Beg05, CZT06, Day01, Day08, GRZ12, Hea05, Li09a, Luy11, TP11, TMPP13, WIY01]. representations [Gor01, LVdA^{+}04, ZLB11]. represents [BK10]. reproductive [Riv13]. Repurposing [BH11, SEDC11]. reputation [ADD^{+}06, Cro11a]. requests [EZ08, TÜ05]. require [BS09]. Requirements [AJR03, BJD10, JJ10, Hep04, HSJB12]. rescaling [LRB^{+}13]. Research [Ago07b, Aus01, BEK^{+}01, BR01, BdMAM13, BBC01, lC01, CPO01, CMFZ01, CWZ^{+}01, Chu01, CC01b, Coo01, Cro01, DC01b, Day01, Ess10, Glu01, Goo01, GD01, Gre01c, HH01, HV01, Hjø01d, HS01, JLCL04, Jul12, KZ01, KLFG04, KSK01, KW01, Kra02a, KW02, KR01, LWW01, Meg01, PJ01, Sch01, SG01, SWJS01, SSB01, SK01, TS01, The09a, WKdJvdBV01, WCLZ01, Whi01, WIY01, Wil00, WIZ01, ZIB01, ZI01, Zha01, ZW01, Åst07, AD09, ADC10, Ago06b, AGOP06, Ano14, ASQR09, Bai04, BT08, Ben11, Bil01, BR12b, BBB04, BSM^{+}05, Bor12, BD07b, Bor13a, BL13, BB03, BK10, Cab12, CS08b, Cho12, Cly04, CvLB10, DGA11, DZC^{+}09, DT13, ESF^{+}08, Enn07b, FARR09, Fon07, For04a, FN08, FR01, GV13, GMV13, GACCNM^{+}08, GRJS01, HM13, HHHR13, Hen13a, hHwC08]. research [HLC11, Hut08a, JBM04, JP01, JP08a, JP08b, Jay12, JdMAA12, JJBC01, Kaz10, Kim09b, KB11, KdRH^{+}01, Kos12, KT08, KTA12, Kra13, KHC12, Lan04, LT08, Ley08a, LB11b, LR11, LW12, LAL09, LJLlW08, Ma12a, Ma12b, MWG06, Mar07b, MQH13, MS05, Meh06, MBBS02, Met07, MEH07, MBG03, MYWA03, MG07, NKLC09, NC13, Nic03a, O’B07, OJ04, OM08, OKG07, PV03, PB01, Pet07a, PM01, Pom05a, Pra11b, RPL10, RK01, RN09, Rou06, SMAB09, ST08, SV05, SKTM09, Shi12, Sin12, Spi02, SWF^{+}02a, SP07, Ste03a, SP11, Su12, TVFW07, TYS13, TH03, The06, TWF08, TKB05b, Tij10, VW12, WV13, WG05, WJD^{+}13, WWN04, WT08, XFCR13, ZS07, ZS08a, ZS11a, ZLB11, ZvL11, vR06a, vR08a, vR08b, Ano02d, CRBPdMAM12, Mar07a, RN09, vR06b, BI13]. Research [Bia10a, Har04, Par10, Sch03, Sha08]. researcher [ADV13, FN12, LM11]. researchers [ARPS11, Bir07, CBvLvR09, CvLB12, FDB^{+}12, KR08, LMCRRPA09, MR08, MHR13, NHL^{+}10, Opp07, PJ06, RKBS11]. resemblance [SB07b, SB07c]. reshape [Bar01a]. Reshaping [Row09, Saw10]. resilience [Cro12c, RGLCTSC13]. Resistance [NY09, Tra08, CA07, NY08, NS12]. resolution [PRNZ^{+}11, PJ01, WL08]. resolved [Bur07]. Resort [Ano03h]. Resource [FARR09, HKLG05, MBG^{+}07, TVF06, VGM^{+}11, CF06, TrP13]. Resources [Ano03f, BYHM07, KC07, LYM07, LHA07, LZC^{+}07, PR07, SY07, WY07, Ano05, Ano08a, BDS12, DRW08, Erc06, FARR09, FC01, GBMH^{+}09, GZS13, HJ04, Kim10b, LJ13, Liu13b, Lun09, MBC08, Mat13, OCD08, ÖAU^{+}04, Pat11, RW05, Sch01, SD03, Spe01, TVFW07, Vak08, Wil00, YRXS11, Zha01]. respond [HTT07]. Response [De 13, WBB10b, CKF07, SRV07, Tre12]. responsiveness [ZZL12]. Result [HMdR11, Dav08, OAU11, Win02]. results [BD10, BIKYL07, BIKLY09, CDLY11, CMRS09, Gre06, HY09, Käk06, LT02, NSE^{+}06, SB07b, SB07c, The08a, The08b, TSA11, WA03, KAB05, Yi08]. Retaining [YLW10]. retracted [CHMS13, Pet13]. Retrieval [Ano01a, Ano01e, BYHM07, Bia10b, Bor09, Cho10b, DS10b, KC07, LYM07, Lar10b, LHA07, LZC^{+}07, PR07, SY07, WY07, Yan10, AN09, AMS10, Alb10a, AM11, Ano04c, Ano05, AYX^{+}13, BBC09, Bai04, BD10, BR11a, BJC^{+}04, BAT08b, BZC07, Ben02, Ben04a, Bil12, BBM02, BLH12, BEK^{+}01, BP10b, BCT02, BMTV02, CS08a, CKB^{+}08, CW01, CMRS09, CWZ^{+}01, Che03c, Che06c, CZT06, CLH13, CK02, CR03, Chu01, CS06a, CSA^{+}02, Col11, CDLvR03, Din07, DGKS04, DK07, Dov04, Dow04, Efr10, EW10, Efr11, Fat07, FPCB04, FJRNV^{+}06, GOR06, Gar01b, GW12, HLP04, Hea05, HH01, HV01, HVP06, Hjø11d, HM09, Iye01, JR10, Jär07, JLB08, KdHWA04, KLFG04, KS13, KK04, Kim05, Kos05, Kos04, KS01, LCR^{+}05, Laz07, LP03, LM09, LSJR11, LCB^{+}07, LDL^{+}08, LLA05, Liu13b, LB03, LC05c]. retrieval [LS05, Lyn01, Mac04a, MM04, MO04, MLO02, NSQI03, NDD^{+}09, PMK11, PBS^{+}04, PLBS06, PK11b, PS03, QZCC06, SS09, SU11, SRDZ06, SYNC08, Smi01, SBNB11, TJLJ07, TL01, Tho12, TC04, UZ04, VS04, War07c, War07e, War08, WS06, Whi06, WR06, Whi07a, Whi07b, WvdHH04, WG10, WIZ01, WM06, XR07, Xu07a, XL07, XY08, YEV08, YHHL12, YHWL13, YBC^{+}06, Yi08, YCE10, YB10b, ZIB01, ZI01, ZC01, Zha08, ZZ03, Din12, DS05b, Mac04b, dCFLH03, Ing12, Efr12, Lar10c, Lar10b, Egg02a, Gru07, VG07b]. retrieve [ACHVH10]. retrieved [OIM^{+}11, SV05]. retrospective [Gar01b, LHMW11, Tse01]. return [HS13a]. reusability [CH07, PMK11]. reusable [LK06, YEV08]. Reuters [DS05a]. Reuters-21578 [DS05a]. reveal [RI08]. revealed [BR12b, CF06]. revenue [OCJ12]. reversed [Fra03]. Review [Ack02a, Ack02b, Afi03, Ago04b, Ago04a, Ago05, Ago06a, Ago06b, Ago07a, Ago07b, Ago08, Ago13, Ahm07, And02, Ave13, Awa06, BI13, Bar01a, Baw10, Ben01a, Ben06, Ben11, Bia10a, Bla13, Bol07, Bol08, Bor09, Bra09, Bra11, Bri10, Bro02, Buc10, Bur09a, Cal08, Cas03, Cer10, Cha03, Cha13, Che06a, Che03a, Che04, Che09, Che10b, Cho09b, Cho10b, Coo07, Cox09, Cox10, Cul01, Cul04, Cul06, Dav09a, Dav11, Day02a, De 11a, DeN09, Des07, Des08a, Des08b, Det11, Dia02, Dim09, Dot13, Dum01a, Edw07, Egg02a, Egg07a, Els09, Enn01a, Enn01b, Enn02b, Enn02a, Enn03a, Enn03c, Enn03b, Enn04, Enn06b, Enn06a, Enn07a, Enn07b, Enn08a, Enn08b, Esc10b, Esc10a, Ess10, Exn02, Fah02, Fan10, Fer04, Fle06a, For09, For13, Fos01, Fou01b, Fou01a, Fou02, Fou12, Fry10, Gaf07, Gal06]. Review [Gal12a, Gaz07, Gaz09, Glu08, Gon08, Gra05, Gru07, Gua11, Hal01, Hal10, Han01, Har09, Har04, Har05, Har06a, Har07b, Har08a, Has03b, Has09, Hay07, Hea05, Hem07, Hil11, Him03, Him04, Hjø01a, Hjø02a, Ho08, Hol01, Hol04, Hut08a, Hut08b, Ing12, Ivo03, Jam10, Jap07, Jap08, Joh01, Jon02, JLB08, Kal01, Kal05, Kan04, Kap09, Kaz11, Kim11b, Koc06, Koe03, Koe04a, Koe04b, Koe11, Kos09b, Kva10, Lap07, Lar10a, Lar10c, Lar10b, Law02, Lin02, Lip13, Luk09, Ma07, McI10, McL10a, McL10b, Mea10, Mer09, Mis09, Mit08, Mon04, MC03b, Mos09, Mue10, Nay04, Nay05b, Nay05a, Nay06, Nay07b, Nay07a, Nea08, Nea11, Nel06, Nic03b, Nic10, O’B07, O’B12, Opp09, Opp10, Orm01, Orm07, Ost08, Par10, Pee12, Pet12, Pet02, Pet03, Pet07a, Pet07b, Pon01]. Review [Pre09, Raw10, Ren04, Reu05, Rho04, Row09, Ruc10, Rya09, Sag06, San10, San02, Sat08, Saw10, Sch03, Sch04a, Sch04b, Sch01, Sch02, Sch05, Sha08, Sha06b, She10, Sim07, Smi04, Soa06, Son10, Spe01, Sta10, Sti01b, Sti01a, Sti02, Sti03, Sti05, Su06, Sun06, Swa11, Tad06, Tar08, The09a, Tho09, Tie10, Tol01b, Tom01a, Tom05a, Tom05b, Tom01b, Tom04, Tom06, Tra09, Urq09, Van03, Van10, Vei06a, VG07b, VG07a, VT13, Vil10b, Vil10a, Vit08, Wad01, Wal04a, Wal02, Wan07, War02, War04, War05, War06, War07b, War09a, War10b, War09b, War10d, Wes07a, Whe03, Whi06, Whi02c, Wil05, Wil10a, Wil06b, Win09, Wor09, Wri10, Yak05, Yan10, Yeo11, ZL07d, ZL07c, ZL07b, Zin09, vD10, AMS10, Ano08c, BD08, Cro11b, Day03, Fid13, JP01, KS13]. review [LSZC13, MHM12, MLZ10, MHR13, RKO10, Sar07a, Sar07b, TKB05b, Win11, ZZL12, CLHHVH11, HLP13, Esc10b]. reviewed [CP13, VE13]. reviewers [SL10]. Reviews [Ack03, Ago01, Ano03a, Bad03, Bar01b, Bia10b, Bla10, Cop12, Cul03, Din12, Dum01b, Efr12, Ekb12, Enn03d, Fou03, Gal12b, Hai12, Jul12, Les11, Lin03, Mil12, Sav03, CP13, Har06b, ZvL11]. Revising [HB08]. revision [NRD11]. revisited [ER02, LS11a, MT03, TS06, WvEvR12]. Revisiting [Alm13, San08, Sch10, YC10]. Revolution [Bla10, Tra09, AD04, War07b, JLR04]. Revolutions [War09a, CK03]. reward [CvL12, Lai09]. Rhetoric [Swa11]. rhetorical [Ley03b, vdB03b]. rhetorically [Fei11]. rhythm [ELR08, Lia05]. rhythms [PSO06]. rich [RJ07, Zha10b]. Richard [Fou01a]. richness [TTT12]. Rights [Yeo11, FMP01, Lip03, Lip13, Nay07a]. Rigor [Cro09d]. Rindflesch [Ano06b]. ripeness [RJ07]. ripple [FN13, PK11a]. Rise [Les11, McC09]. rising [CK03]. risk [LLL09]. Rita [Hjø02a, Chu07]. Robinson [Ave13, Rob10b]. robot [Hal01]. robots [DG12]. robust [ADV10]. robustness [Bur13, Van07]. Rocchio [CF07c]. rock [TSV12]. Rogers [Ano06b]. Role [CH05, WEA^{+}12, AT02, Bar06a, Bro03b, Bro07, BDBA13, hChC05, CFW^{+}12, FMM10, FT12, Fon07, FMM01, Gen06, GL09, HFW09, Joh07, Kal06, KH09, KTR11, KTA12, LT09a, Lip13, LY11, LS11d, Nay06, SDWG01, Ste03b, SZ08, SST06, TB13, Win11, YD13, Fle06b]. role-playing [Win11]. Role-related [CH05]. roles [BPT07, CLD^{+}13, For04a, FDB^{+}12, FDY^{+}13, Huv13, Lan06, Wil10b, XFCR13]. ROM [SD03]. Roman [VT13]. Romania [Dra02]. root [ASABS^{+}10]. roots [HASAT^{+}10, HS13a]. Rotation [LP03]. Rough [Has04, Kis10]. Rousseau [Bur09b]. routine [LO13, War10c]. RSS [Edw07, TPF06, TP07]. rubrics [Oak09]. Rule [Yeo11, Has04, HNB^{+}09, SY07]. rule-based [HNB^{+}09]. rules [AAN12, CBvLvR09, KSS03, RCT05, Tan04, vR08a, vR08b]. run [Hal04a]. rural [LKG^{+}10]. Russia [LKZC11]. Russian [DS09, KGL12]. Ruthven [Ing12, Efr12].

S [Joh01, Atk01]. S. [Ano06b]. Saaty [RBLH13]. safe [LB13b]. safety [MMC05, vT08]. Saggion [VT13]. Saharan [KMD06, MKMM09]. Salton [EL09, Ley08b]. same [AG02, BJC^{+}04]. Sample [Bur04a, YKT03a, BV08, Bur03a, Dow04, LT10, NJ07b, YKT03b]. Sample-size [Bur04a]. samples [iCG09]. Sampling [Egg02b, Hen12]. sanction [GKL09]. sands [SLR^{+}11]. Sanyal [Hol01]. SAR [LY05]. SAT [RMC12]. satisfaction [AMS10, CL08, CL09, HHfHC10, WL11b]. satisficing [Ago02, WRB^{+}09]. Saving [Pet02]. scalable [CJB^{+}07]. scale [ARPS11, Ano01f, CHdS^{+}11, DGA11, Fra12, HHfHC10, LP03, LKBJ12, LYJ^{+}10, TWF08]. Scaling [CBvLvR09, MAL^{+}05, vR08b, BS09, Goo01, VC06, vEWDvdB10, vR08a, HHWD05]. scanning [HHHR13, TPF06]. Scatter [BP10a, TÜ05, Bha05, HW01]. Scenarios [MBBS02]. scent [SKWH07]. schemas [Gre01b]. scheme [Ten12]. schemes [Ley02, Zin07a]. Schneider [BL13]. scholar [BF08, Fra09, GP10, HvdW09, KT07b, KTR11, MY07, Muk09, OA12, Whi02b]. Scholarly [Luz09, Pal05, VT03, YD12, Zha01, Bar06a, Bjö12, Bro01b, Bro07, Cro01, CSL03, Jac01, JLK09, KTR10, LZG13, LP07, MS09b, Min10, Miz03, Muk09, Nic03a, Per03, ŠKL08, Sel05, TJSL11, Vak08, XMCM13, YDS11, AL08, KMK03]. Scholars [ST13, BAB07, CM11, Kil01, TH04, Xia10, Zin07a]. scholarship [SC12, Fry10, Enn07b]. Schön [Hol01]. school [Abb05, Ano01c, Chu09, CN07, Cly04, CBL^{+}13, LBNB04, SDJB12, Ano02a]. schools [DJW^{+}11, MS05]. Schubert [Pra11a]. Science [Ahn11, Ano06b, Ano08b, Ano08c, Ano10, Ano11, Ano12, Ano13b, Ano13c, Ano14, AFR07, Bor12, Bor13a, Buc10, Che09, Che10b, Cho12, CLHHVH11, CC13, Efr11, Fan10, Fos01, Fou12, Gaz11, Haw01, Hil11, IS12, JP08b, LSZC13, LB05, MHM12, Nea11, Nic10, OA12, RPL10, Rob10b, Sha13b, Shi12, Sma12, SML12, Tan07a, TrP13, Van10, VW12, Win11, WCC10b, Zin09, Ano01b, BDH12, BAB02, Ben06, Ben05, BW05, Bir06, BM12, BWD02, Bro01a, Buc11, Buc08, Buc12, Bud07, Bus07, CC11, CH12, CY08, CLHHVH12, Col07b, CBvLvR09, CM07, CM08a, Cro10a, Day05, Dra02, Fon07, For04a, Fra11, GN07, GACCNM^{+}08, Gor12, GBF^{+}12, Gu02, HLC03, Hjø02b, Hjø02d, Hjø03, Hjø11b, Jay12, JLR04, JTLG12, Kar06, Kar07, Kar09, KH07, Kim09b, KGL12, KB06]. science [KB09, KT07c, KT09, KC09, LAG08, LM11, LSC12, Ler10, LT09b, Ley04, LV06, LR09, LP10, LB11a, LZB13, Lia05, Ma12b, MS05, Mil10a, Mil10b, MBG03, Mor10, Muk09, Opp07, PM01, RMC12, Sar07a, Sar07b, SH07a, Sch05, Sma03, Ste03a, Sto10, SHS^{+}11, Su06, Su12, Sun06, SK01, SK02a, SK02b, TTS13, The06, TPF06, TP07, Tij10, VS05, Vit08, WV13, WvE12b, WJD^{+}13, Whi02a, Whi07b, Wij12, ZS08a, ZS08b, Zin07b, Zin07d, ZLB11, dMAVQCR^{+}07, vR08a, vR08c, vR08b, ACHVH10, ACGL09, Fur10, Fur11, GP10, MY07, MR08, MS09b, MSYD11, SNRB11, SLR^{+}11, ST13, Whi02b, ZL07a, ZZ10, Zin07a, Ave13, Ago04b, Cas03, Har06a, Har07b, Hay07]. science-related [TP07]. science/applied [Tij10]. Sciences [GC07, AG02, BT08, Ben08, CF06, HB09, HW11, hHwC08, LAGVG06, LTO11, Pow03, Shi12, WBY^{+}09, ZTG13, ZSL10, Esc10a]. Scientific [AGOP06, FR13, GRSFV12a, MH10, Riv13, WHP10, ZL07a, ADC12, Aks06, ARC11, APSA12, Ano01i, Ano13b, AYX^{+}13, BMJARB06, BBRdMA12, BSP10, BOGdMA13, Bro03a, Che03b, Che06b, Che08, CHMS13, CvLvR10, CvL12, Dan11, DSG^{+}12, GV13, Gau09, GSD12, GLAA07, HSKS03, Has03a, HDN11, HHHR13, Hen12, HS13a, Hub01, JSI^{+}04, KKF^{+}12, KGL12, LB13a, LAG08, LG10b, Ler10, Ley02, Ley06, Ley07a, Ley07c, LLA08, LMCRRPA09, MAL^{+}05, MCRdMA11, Mil12, Min10, NHCW13, PMBdS10, Pep11, PDJCRBBM12, PAR13, RL09, SCLC06, Sch04a, Sha07, SSPA10, SHS^{+}11, TYS13, TK08, TKV^{+}10, TSRGJC^{+}13, VL13, Vin04, VBMM11, WGD09, WGW12, Wu10, ZQI^{+}10, ZLXW11, Dim09]. scientist [BLT07, De 10, Gar01a, KAB09]. scientists [AJ10, ARPS11, AFR06, AFR07, BP06a, BM11, CA07, CvLB12, HLVA07, Hjø01c, Hut08a, MT03, NH12, Whi03b, YH04, HSKS03]. scientistss [CM06]. scientometric [Ben11, BGY^{+}13, Boo01, Bur01, Cab12, Hen13a, Iva01, SS04, Vin11a]. scientometricians [Vin11b]. Scientometrics [Vin11b, ZTG13]. Scifinder [Whi02b]. SCImago [BdMAM13]. SciMAT [CLHHVH12]. scope [AW05]. Scopus [ACGL09, GRSFV11, KTR11, LO10, LdMAGB10, LO11, MY07, MR08, MS09b, MCRdMA11]. Score [Mar09a, HLP04]. Score-based [Mar09a]. score-driven [HLP04]. scores [Bor13c, Pra10, Rou12, Sch13b, WBB10a, WBB10b, Win08a]. scoring [BBC09, Sch13b]. scratch [Koh13, WK13]. screen [NMPBA11]. screening [Hjø11a, KW01]. Search [Bea06, Dav06, For13, Gre06, Hen06, HHvdHdR10, JBCJ06, Kal06, KSM06, Moc06, RV06, Ros06, Rot06, Sac06, VH12, VR06c, Wil06a, Alb10b, AFG02, Ano01f, ABJ^{+}06, AYSC06, BYHM07, BD10, BR12a, BI02, BI06, BIKYL07, BIKLY09, BBI12, Bha05, Bil01, Bil02, Bow10, CDLY11, CMRS09, CLMB10, CS12, CFS05, CSCC07, CFY07, CWZ^{+}10, CMFZ01, Cho13b, CKK^{+}02, CC01b, CJ04, DBP^{+}07, DH13, DBM02, DP07, DS11, Efr11, FFPW04, FMM03, FMM05b, FMM05a, FDB^{+}12, FDY^{+}13, FL01, Fu10, GL09, Gwi10, Haa03, HZ07, HWT^{+}13, HGGL05, HY09, HTdV11, Hon12, HC11, HCO03, HM09, HWJ10, JSBK07, JZS09, Joh07, Käk06, KCC12, Kim11a, KSJ06, KT09, KAB05, KLAS13, KC12, Lak03, Lak07, LC12, LDL^{+}08, LS11a, LDvM12, LfBW08, Li09b, LB10, jLB05, Liu13b, LHB^{+}08, LT02, LT04, MHL05]. search [Mel04, MEH07, Mos04, NSE^{+}06, OIM^{+}11, OJ04, OCJ12, OAU11, PK05, PV03, PLN09, PJ01, PCY02, RH12, RZ04, RLvR03, SWT04, SZ11, SY07, SR06, Slo02, SBA12, Sre02, Ste08, Su03a, Su03b, TDC13, The08a, The08b, TSA11, UGF08, VRL^{+}09, Wal11a, WCCS13, War07c, WA03, WEF^{+}07, Wil04, WG10, WmcsW09, WWZ09, WRM^{+}02, WCC08, WCC10b, WHTS12, XJ10, YBC^{+}06, Zha02, dMFRN^{+}10, Efr12, Yan10]. searcher [Bod09, JR06, JSK07]. searchers [PJ06, STS12]. searches [CG06, Eas02, Kil01, Kil04, jL05, LX13, Sac06, VVLB12, Wal11a, Whi02b, Xu07b]. Searching [CR03, Cho10b, KY04, Lar10b, SWJS01, AN06, BCT06, Bod06, CFY07, CFCZ01, CLZC03, Cho10a, CZH^{+}04, Cot02, DBM02, DS09, Dov04, EWF^{+}02, FMM01, FWF^{+}02, HSSV07, HDB07, JP01, JM05, JSP05, JR06, JBS09, JR10, KA02, KSJ06, LC06, LM09, LLD^{+}02, Mar07a, Mar07b, MAR04, Mus05, NBYA03, Rie04, SWF^{+}02a, SWF^{+}02b, SOO02, Tan07b, TSJ09, Tse01, WCCS13, WFE^{+}02, WSJS01, XC09, ZJS09, Zha13, LfBW08]. seat [Cro09a]. Second [Ano04b, Gaz09, Cri11]. second-order [Cri11]. secondary [LHMW11]. secret [LB13b]. section [AH05b, Arg06, HVP06, LYM07, Moe07, Mos04, YL06]. sections [NMD09]. sector [ME03]. secure [CP13, LC05b]. Security [Ano03g, Che05, Lai09, MC03b, RRW05, SH05, WG07, Wor08, Ano04b, HLC05, Sta10, Wor09]. Seed [MvL12]. seeker [CJ04]. seekers [AXP11]. Seeking [CBL^{+}13, Efr12, HD04, HBSZ13, Ing12, Jul12, RK01, Ano10, BSM^{+}05, CJA^{+}04, CN07, CLB^{+}05, CLL^{+}07, Coo02a, DW04, Dav04a, EWF^{+}02, FWF^{+}02, For04b, Fos04, Gen12, HFW09, HML02, Hei06, HLVA07, HD06, Joh07, JBCJ06, Kal06, KTW05, KWS07, Kel06a, Kel06b, Kim09a, LT12, Lee10, jLB05, MW10, MT03, MBBS02, NHL^{+}10, NH12, PDU02, PGR^{+}13, Rie04, Ros06, STS12, SM10, Sha13a, SWF^{+}02a, SWF^{+}02b, SOO02, SC11, Tan07a, TRJ05, WP04, WRB^{+}09, Whi02d, Wil02, WFE^{+}02, WmcsW09, WL03, XC09, XTY06, YD13, YR08, YR02, YB10a, YB10b, YRXS11, Zac05, ZZZR04, Ano02d]. segmentation [LZ08b, SD01, WY07, YL05]. seldom [MM10]. seldom-cited [MM10]. selected [MKMM09, Muk09]. Selecting [BD08, FS10, Ano06b, HRK^{+}06, KAD06, RLvR02, Ano08c]. Selection [War07e, AAN12, dMCMNN09, Efr05, HGGL05, LY11, Ma02a, MH11, MG09b, Reu07, SC08, SM03]. Self [Cam10a, Col07b, Cro13c, GKL09, Hyl03, Rab09, vR08c, Ack02a, Bil02, Bow10, Cam11, DP07, Egg05a, EF10, FR13, GMM09, GBK05, GL12, HD06, Kim10a, Lai09, LH01, Ley03a, Ley10, VC08, WIZ01, vdB03a, AE10]. Self-archiving [Col07b, Kim10a]. Self-citation [Hyl03, vR08c]. Self-citations [Cam10a, Cam11, EF10, FR13, GMM09]. self-concept [HD06]. self-efficacy [Lai09]. self-generated [Bil02]. Self-interested [AE10]. self-organization [Ack02a, LH01, Ley03a, Ley10, vdB03a]. self-organizing [DP07, GBK05, WIZ01]. Self-plagiarism [Cro13c]. Self-presentation [Rab09]. self-reference [Hyl03]. self-regulation [Bow10]. self-similar [Egg05a]. self-views [GL12]. Seljuk [ÖL06]. ‘Sell [CvLvR10]. ‘Sell-By-Date’ [CvLvR10]. Semantic [Che06a, CBS10, Din05, DC09, HTdV11, NZY13, RJ07, Ahm07, Ano06b, BZC07, BJ09, CS08a, Cho13b, CH07, Cui12, Efr05, Gre01c, Gun13, GW12, HRK^{+}06, JT10, JZW^{+}11, KRK^{+}10, KS10, Kim11a, KS01, LDL^{+}08, LKL10, LK10, Men04, NJ07a, NJ07b, OYM09, PMBdS10, POMVULMC13, PBV08, RPYZ07, RHCP12, Sto10, TL01, UGF08, WCLZ01, Yi10, ZDM13, ZL07b, DPV11]. semantic-based [JT10]. semantically [Gre01a, PG02]. semantics [MR09]. semi [SPR09, WP05]. semi-automated [SPR09]. semi-supervised [WP05]. semistructured [Coo02a]. semisupervised [SZC11]. SemRep [RRA^{+}13]. Senate [KBP10]. senior [Zac05]. seniors [GRS^{+}07]. sense [Ano06b, Gen12, Gor01, HZ12, HBSZ13, HRK^{+}06, KPEO10, KTL^{+}13, LSD13, MHM12, Met07, PCR13, Sav06]. sense-making [HZ12, HBSZ13, Sav06]. sensemaking [WW13]. Sensitive [TKB05b, SB03]. Sentence [LSJR11, CL11, ZZL^{+}09]. Sentence-based [LSJR11]. sentence-level [CL11]. sentences [KH12, MAA08, WJR05]. Sentiment [Ano11, BL12, TBP^{+}10, TBP11, TBP12, ZZL^{+}09, TB13]. separate [AYX^{+}13]. sequences [ELR08, Lia05, Sha06a, TLM06]. sequential [CH03]. serendipity [MQH13]. serial [Chr03]. serials [Chr04, Dav04b, LAGVG06]. seriation [Glu01]. Series [Din12, Efr10, McC08]. servant [Jay12]. Service [SOH08, CF06, DRW08, Fou02, Hay07, Hep04, HBT^{+}09, HLP13, JJ10, NV09, Pom05a, TS06, Mit08]. services [ADR11, BR12b, Cho12, DJWR02, FDH04, GD03, HBT^{+}09, JHR01, KKGB11, LCC07, RU07, Smi04, UR07, ZL07c, Din12, Bia10b, Tom01b, Mon04]. Session [HWT^{+}13, Ano01e, HCO03, HPAS04, JSBK07, WWZ09]. sessions [jLB05]. set [Has04, LZ08a, LP09, SBF13]. sets [Bor13c, HKBY^{+}10, LBMO11, LAL09, MW08, MCDA03, Vin04]. setting [McK04, Nic03a]. seventeen [JR10, Win08b]. Seventh [Ano02a]. several [HRC12a]. Sex [FMMF13, SRV07, Ago08]. sex-specific [SRV07]. sexual [CL12]. Shah [Ano10]. Shall [LY11]. Shannon [Efr08]. shape [Enn08b, LLA05, LT04, Oya08, TDC13]. shape-based [LLA05]. shaped [War02]. shapers [YWM10]. Shapes [Les11, BP10b]. Shaping [Cab12, Yeo11, Pal05, Nay06]. share [PNHG07, ZTG13]. shared [GD01, Šau04, SPH08]. Sharing [Cox10, McI10, Bor12, BDBA13, CLL13, CCC10, HH07, HFC10, IA12, KY08, LGCA10, LfBW08, LC05b, NNY10, O’C13, Oh12, SSM^{+}08b, Vei09, XZW13, YNF09, YZLC13]. Shaw [Fou12]. Sher [Gar01a]. Sherry [Ekb12]. shift [CY08, CC04, Cro01, LM01]. shifted [ER12]. shifting [CM08b, OV01, SLR^{+}11]. shifts [DS11, Jar11, Tre12]. shootings [HZ12]. shopping [CL06, Moc06]. short [Ano11, Egg10a, TBP^{+}10]. Shortz [Efr08]. Should [Ley07b, Hjø02d, WvE07]. shut [Enn08b]. Side [Les11, FG03, KLP09]. sifting [TD02]. sighted [STS12]. signal [DMR06]. signals [TC04]. signature [LWW01, ZA13]. signatures [PBV08]. significance [BL11, Cro08]. significant [APSA12, BIKLY09, HKC09]. significantly [TH04]. silence [iCG09]. similar [Bre05, Egg05a, HBBdR10, LfBW08]. similarities [Cha11, YD12]. Similarity [Ley05, AJR03, BKRN03, Bur05, CCG^{+}06, CF07c, Cri11, Egg06, Egg09a, Egg10b, Egg10c, HMFW12, LB03, Mel04, NZY13, PLN09, PN09, Ror07, SRDZ06, TWSS05, WCC10a, Yi10, ZI01, vEW08, vEW09]. similarity-based [CF07c]. similarity-preserving [SRDZ06]. simple [CH03, Enn08a, TS08]. Simplifying [Fra03]. Simulating [GR09, WKdJvdBV01]. simulation [AHK08, BHDK13, Chu05]. Simultaneous [Has03a]. Singapore [TPLC05]. single [BGY^{+}13, BD07b, HZ03a, Moe05, RHWB11, Vin13]. single-pass [HZ03a]. site [Ano03d, CFS05, CL12, CTS04, ECP08, HLLS07, HS01, Hon06, HC09, KO09, Kim10b, Koc06, KLAS13, NLC06, Sav12, SCC^{+}12, WT10, YL09, YC11, vRFBvW10]. site-level [SCC^{+}12]. sites [Ahn11, CB13, CJJ^{+}02, ESL^{+}10, EHD05, HVPLHP06, HFC10, Jon02, Ma02b, PBN02, RTM04, SLS06, SCA12, The02, TW03b, TH04, The05, TW10, VT03, VTD10, Wal02]. Situated [AM11, Hem07, Yan01]. Situating [Kar09]. Situation [TDC13, CLB^{+}05, Coo02a]. situational [Hjø07]. six [Sin11, WL11a]. Sixty [Ard13]. Size [LDM12, YKT03a, Bur03a, Bur04a, Egg08c, SBF13, YKT03b]. size-adjusted [SBF13]. size-frequency [Egg08c]. SK [PDJCRBBM12]. skewed [Win02]. skill [GL12]. skills [Ago05, GL12, Har02, HSJB12, LMNS13, Tal05]. slave [Wij12]. sliced [Åst07]. slip [AYN13]. Slovakia [Ste03a]. Slow [Cro13d]. Small [BBC01, Bor13c, Enn08a, IS06, MW08, Põd10, SS10]. smoke [Hen13a]. Smoothing [ACMP11, LLXS13]. Snap [Ben01a]. SNIP [LO10, LO11]. soap [ST13]. Sociable [Gua11]. Social [CCCJ10, Dav09a, Day02b, Gaz11, Hil11, HS13a, Joh07, Kap09, Kaz11, Kom11, LS12, Nah07, Pet12, SK12, The08c, Tom06, WT10, XMCM13, AAN12, Ahn11, AdG12, BIZGMS10, Ben08, BPT07, Bri08, BDBA13, CLD^{+}13, Cha09, CC11, DRW08, Day05, Dia02, DJZ^{+}09, DJF^{+}10, Gaz07, God07, GKE11, HB09, HWT^{+}13, HW11, HC11, hHwC08, KKGB11, KO09, KT09, LAGVG06, LM11, LTO11, LBS11, LDS^{+}11, LNK07, MT03, RK01, Sav07, Sav12, SZ11, SBIS11, SCA12, SV03, SXZ11, SM03, SBNB11, SK13, SS13, Tal05, The06, TWU10, TW10, TBP12, TB13, WJD^{+}13, WWN04, WBY^{+}09, Wor08, WK13, XZW13, YS11, YKCL12, Yi10, Yi12, YC11, ZQI^{+}10, ZSL10, Day07, MA12c, BI13, Cho10b, Gal06, GC07, Har08a, Has03b]. societal [Bor13a]. Society [Ano06b, Ano08b, Ano08c, Ano10, Ano11, Ano12, Ano13b, Ano13c, Ano14, AFR07, Bro07, Bur09a, JP08b, Mos09, Rob10b, Tan07a, WCC10b, vD10, BR13, Bri08, Dra02, Duf11, Gon08, GD07, Mac06, Nay06, Sto11, Wes07a, Ano01e, LH01, Ack02a, NMD09, Nay07a, Tom05a]. Socio [Ehi02, HC02, Hjø02b, KMK03]. socio-cognitive [HC02, Hjø02b]. Socio-economic [Ehi02]. Socio-Technical [KMK03]. socioeconomic [MDM06]. sociological [Ack02a]. sociology [ŠKL08, VW12]. Sociospatial [MA12c]. sociotechnical [GMV13]. Soft [HVP06, Kos04, Ano02e, Ano04c]. Software [Has03b, CF07a, CLHHVH11, CLHHVH12, EHD05, Har08a, Ma07, MM01, MO08, NGJ03, YEV08, YH04, CF09]. solidarity [Ano14, SMAB09]. solution [LATH08, Tho12]. solving [KS10, MTS^{+}13]. Some [Bur06b, Bur09b, Ma12b, WvE07, YKT03a, ASABS^{+}10, TTS13, vEW09, Bur03a, Bur03c, Bur04b, OV01, RMC12, SH07c, Whi07b, YKT03a]. someone [WK13]. Sony [PLZA04]. sophisticated [KAB05, Moe11, PLN09]. soul [FMMF13]. Sound [RN09]. Source [Cho09b, CF09, AXP11, ACHVH10, AD04, CF07a, GRJS01, Hjø11a, LO10, LLA08, LY11, Moe11, NGJ03, RR06, Sav07, SM03, TK08, VT13]. source-based [ACHVH10]. Sources [Ano13d, Bys02, CJA^{+}04, CKH^{+}03, CvLB12, DBT11, Egg10d, Gen12, Huv10, MS05, MY07, MLZ10, SC11, TWF08, VY12, Wei10, YH04, SB12a, Wes09]. space [CL03, Cro13f, Fou01a, GdMA01, HT11, Has03a, Hoe03, Lak03, New01, SRDZ06, Tan07b, WZ02, WOB09, Zha10b, vdEvMK^{+}04]. Spaces [Gua11, FL01, GLDM02, KS01, RCD08, WZ08]. spacing [HB07]. Spain [OM13, RN09, RFV12]. spam [PN09]. SpamED [PN09]. span [CC11]. Spanish [Ano14, GRSFV^{+}12b, SMAB09]. Spanish- [Ano14, SMAB09]. Spanning [Har07b]. Spärck [RT08]. spatial [NMPBA11]. Spatialization [RTM04]. speak [LBR02]. speaker [JLK09, Suz09]. speaker-specific [Suz09]. speaking [VVLB12]. speaks [Cus13]. Special [Ano01a, Ano01g, Ano01h, Ano01f, Ano01i, Ano02c, Ano02e, Ano02d, Ano02f, Ano03e, Ano03g, Ano04c, Ano05, Arg03, Che03c, Mar11, AJR03, AH05b, Arg06, Che05, CDLvR03, Hem05, HVP06, Hjø02c, Hu09, LYM07, MD02, Mos04, Spi02, TV04, Tom04, YL06, Ano01e]. specialists [Hjø02c]. specialized [HJ13]. specialties [CC13, Mor05, WGD09]. specialty [HNB^{+}09, LP09]. Specific [BdMAM13, BIP04, CBvLvR09, De 10, HMdR11, HC02, SRV07, Suz09, YBC^{+}06, vR08b]. specificity [Hei06]. spectrum [iCG09]. speculation [DLVÁ12]. speech [LZ08b, YWC08]. speeches [Suz09]. speed [CB13, Kis10, Cro09d]. spell [CW04]. Spelling [OW04, WG10]. spent [TD02]. spillover [TH13]. Spink [Jul12]. spoken [CD06]. sponsored [Ano01e, JR06]. spontaneous [WBJM12]. sports [ZWW09]. sports-related [ZWW09]. spread [ST09, TP06]. spring [MLO02]. Spychips [Mea10]. squares [ADR11]. Srikantaiah [Sti01b]. SSCI [LJLlW08]. stability [RPG^{+}08, TSA11]. Stabilizing [WH09]. stage [KC12, LSD13, MH11, CBL^{+}13]. stages [HHNK12, TS01]. stakeholder [CCR09, MKMM09]. stakeholders [CCR09]. stakes [KSF04]. Standard [Abd04, BMD09, Jör07, RMC12]. standards [Ahm07, Cha13, DDO11, MBC10, RMC12]. Standing [Cro13e]. standpoint [MCRdMA11]. Stankus [Fos01]. State [XW02, GMA07, LAG08, TrP13, DRW08]. statements [PCR13]. States [KBP10, War02, Yeo11, Ben04a, Tsi13, CLB^{+}05, CTS04, HLB05, NHCW13, NHL^{+}10, Shi08]. Statistical [BL13, Din07, Moe05, SGJ13, SSPA10, YKT03a, vR06b, BT08, BL11, Bor13b, BMD13, Bur03a, Bur04a, CS04, CL01, CTC05, DMR06, Haa03, HDG^{+}07, HPAS04, HNB^{+}09, JBM04, KDL02, KW01, McD07, YL05, YKT03b, vR06a, vR08a]. statistics [ACGL09, BHC06, DS03, DP06, DSG^{+}12, Ley03b, Mar02, SB07c, vdB03b]. status [DGKS04, LKZC11, RR06]. Stay [AYN13]. Staying [Det11]. steady [LAG08]. steady-state [LAG08]. stem [ZS11a, AR13]. stemmer [MO07]. stemmers [FS09]. stemming [Kos03]. Step [Det11, CJA^{+}04, Koc06]. step-by-step [Koc06]. stereo [CZT06]. Stereotype [KSS03]. Stereotype-based [KSS03]. Stereotypical [HB11]. stigma [LB13b]. Still [Bro01a, Enn01a]. Stimulated [Coo07]. Stochastic [CC02]. stock [SC08, Pee12]. stopping [Zac05]. stopword [DS10a, Laz07, Sta11]. stories [BEES10, WZL09]. storyboard [IL07, KK10b]. strange [Ten12]. Strategic [BMTV02, Pee12, PDJCRBBM12, DA05, HL05, SLSK09]. Strategies [Enn03a, McL10a, Vil10a, BJC^{+}04, CS12, DS09, Dra03, FMM03, FMM05b, FMM05a, FMMF13, HRC12a, HWJ10, IL07, MKMM09, RZ04, SC08, SL12, YB10a, YB10b, Bra11, Fah02]. Strategy [BBJ^{+}06, Gra05, Ma02a, PLN09, WL09, Joh01]. strategy-based [WL09]. stratification [HH13]. strength [Ano11, BKRN03, Egg10b, TBP^{+}10, TBP12, ZY13]. strengths [Sti01a]. string [Mus05]. strip [MWG06]. Structural [DMR06, APSA12, ASW05, BZC07, Che12, Cro01, GLLL11, HH13, Hon06, HRC12b, LS08, SSS13, YC10, dMFRN^{+}10, BMJARB06]. structurally [NJN08]. structuration [HPGG^{+}07, Ley10, RS10]. Structure [GC07, SB07a, Ano13a, AR13, Ben08, BWS^{+}10, BR01, CISH10, CL08, Cri11, DWA13, FARR09, Fra12, GD03, Hua13, KOS09a, KS10, LHS11, LK10, LDS^{+}11, LP07, MSYD11, OM08, PBN02, PV03, Rab08, ŠKL08, Sav12, Su12, TW03a, VC06, WWN04, WL13, YL09, ZLXW11, Zha12, Zuc06a]. Structured [TL01, ANdlF07, Gre01a, HB07, HB08, Kim11a, MMMM^{+}01]. structures [BW05, DML02, Ley02, LVdA^{+}04, Meg01, QH06, VL13, YLW10, ZSL10, Huv13]. Structuring [LCB^{+}07]. struggle [Day02a]. Student [PV03, CN07, CBL^{+}13, God07, GLLL11, Kra02a, KW02, LSB13, LGM^{+}09, Mat13, RU07, SDJB12, UR07, Zha08]. Students [Sch03, BDS12, Bow10, GKL09, GL12, Hei06, LBNB04, Lim09, PR10, VVLB12]. Studies [Gaz09, Ard13, BAB07, BH11, Bol07, BH05, CCMM02, CM08b, DLH^{+}08, EGR11, Enn01a, JJ10, JP01, KS13, LP09, MS09b, PZ12, SBT^{+}07, Sun06, TS01, Enn06b]. studio [PGL13]. studios [SHS09]. Study [ER08a, Els09, FCjM12, APLG13, ARPS11, ASABS^{+}10, AF12, AB09, AJBAFHCE08, BIP04, BIB07, BIKLY09, BAB02, BHDK13, BPT07, BLH^{+}10, BGY^{+}13, BD07b, BSMD11, BEES10, Bud07, Cab12, CS04, CH12, CFY07, CCQ^{+}08, CHMS13, CAC10, CBDH08, Cly04, CLHHVH11, CLB^{+}05, Coo02a, CvLB10, Cot02, CM07, DBM02, FPCB04, FG01, FMM01, GP10, GRSFV^{+}12b, GTHZ05, GKE11, Gor12, Gu02, HT11, Har08b, HZ12, HH07, HD04, Hjø11a, HBSZ13, HTdV11, HLC05, HLP13, IS06, Jär07, KTW05, KWS07, KK08, Koe02, KO12, KTR10, KY08, LT12, LCC07, LT08, LR10, LKZC11, LHMW11, LHHVH09, MM10, MBG^{+}07, Mar02, Mat13, MT03, MDM06, MLO02, Muk09, MPR^{+}05, MHR13, MK13, NMD09, NCRJ13, NHL^{+}10, NH12, PPH06, PSB04, PV03, PBS^{+}04, PGR^{+}13, RNLdL01, RM13, Ror10, RJ08, SHO11, Sav09]. study [SS09, Sch09a, Sel05, SKTM09, Shi08, Sin11, SB12b, STS02, SWF^{+}02b, SM03, SK01, SK02a, SK02b, TPLC05, Tan07a, Tan07b, TWW^{+}06, TPF06, TVF06, The08a, TSJ09, Tsi13, VRL^{+}09, Wan09, WRB^{+}09, WR06, Wij12, WG10, WG07, Wor07, WL09, WHP10, XBM13, Xia10, XFCR13, Xie11b, XTY06, XL07, ZZZR04, ZQI^{+}10, ZSL10, vT08]. Studying [MTS^{+}13, FMM03, Gaz07, MBBS02, SC11, Wil02]. stuff [Cus13]. style [Arg03, Arg06, FMCL07, JG06, KAD06, ZLCH06]. styles [CM02, CMD05, FWF^{+}02, LB08, Suz09, WL03, WL11b]. Stylistic [AWC^{+}07, HKBY^{+}10, ÖL06]. sub [KMD06, MKMM09]. sub-Saharan [KMD06, MKMM09]. subauthorship [CSL03]. subcultural [PSWL13]. subfields [HRC12a]. Subject [BdMAM13, LS12, PCY02, YC10, Abb05, Ano13a, BL09, CMH10, CvLvR10, FP04, GRSFV11, GRSFV12a, Hjø01b, HW01, JTDG13, LR09, LRR12, LO13, NKLC09, Pet02, SM05a, SST06, SS13, Ten12, yT04, WL11a, WL13, Yi10, ZWW^{+}08, ZATH09]. subjective [BBMN03, BBMN08, GBK05, GL09, Hjø07, PCR13, RR06, SU11]. subjective/situational [Hjø07]. Subjectivity [Fro08, For04a]. subjects [Day11, LTO11]. sublime [Fle06a]. submissions [BSMD11]. submitting [SL10]. subpixel [BLH^{+}10]. subpixel-rendered [BLH^{+}10]. Subramanian [Pet12]. subscription [Kim06]. subsequent [HGGL05]. subsets [DS05a]. subsidiaries [EAPL05]. subspace [LKL10, ZA13]. substance [ZYH13]. substitute [LS12]. substitutes [YSD^{+}08]. substitution [KJ03]. subversive [Rin09]. subverting [Lip13]. success [BBI12, BM12, Ham07, Huv10, Lai09, vHVD^{+}13, McL10a]. successful [AN06]. successive [AJBAFHCE08, jL05, LX13, SWF^{+}02b]. such [BR13]. suggestion [HCO03]. suggestions [KLAS13]. Sugimoto [Ano13b]. suitable [TPF06]. summaries [ESF^{+}08, LCR^{+}05, WW13, WRM^{+}02]. summarization [CL11, CKH^{+}03, KK10b, OKG07, OLLL10, PAL12, WY06, WLLH09, YW08, VT13, Lin02]. Summary [Mac04b]. Summer [Ano02a, Ano01c]. sung [PSB04]. superstring [CK03]. supervised [WP05]. support [AD09, AB09, BP06b, Cal08, CJB^{+}07, Cha09, DRM^{+}04, Har07a, HDG^{+}07, KSM06, LZC09, LPSS06, SP11, VL13, Wes07b, WR06, WOE^{+}13, YB10b, YZLC13]. supported [ZC08]. Supporting [GBK05, HDB07, RPH^{+}12, ERJ07, Tan06, Tan07c, YB10a, HY09, JBCJ06]. supportive [Cha09]. Supposedly [Hen13b]. Supreme [HLR09]. surfactants [BMJARB06]. surname [Kil01]. surrogates [Goo01, SWT04]. Survey [HLB05, Jul12, ASAK04, Bor13a, GS03, Ham07, HHfHC10, KR08, Ler10, LLD^{+}02, MH10, OT10, PM11, SS04, Sta09, TrP13, VH05, vDH12]. Surveying [GBMH^{+}09]. surveys [Har02, SRV07]. survivors [DRW08]. suspected [Has04]. Sustainable [Nat12, Cho12]. Swanson [WKdJvdBV01]. swarm [LLL09]. Swedish [HH13, LS13]. Switch [Les11]. switching [YSD^{+}08]. syllabuses [KT08]. symbolic [CS02, VC08]. Symposium [Ano04b]. Synchronic [YKT03a]. synergies [LS13]. synonymy [KC09]. syntactic [GW12, ZDM13]. syntactical [YC10]. syntagmatic [Yoo12]. synthesis [Arg03, BP06a, BP06b, Whi07a, WGW12, Bia10b]. synthesized [TTT12]. System [ESL^{+}10, AD09, Alb10a, AHK08, BJC^{+}04, BLA07, BCT06, BJD10, BP10b, BDBA13, CC01a, CC02, CKH^{+}03, CS06a, CBvLvR09, DHW13, GD01, GD07, HKKP10, HH13, HV01, Him03, HNB^{+}09, JT10, JHW^{+}08, KSK01, Kim05, Kos05, KHC09, KL11, LS13, LKG^{+}10, Ma02a, MO04, MBBK07, OM13, PB01, PBS^{+}04, PK11b, RRG11, SKM01, SK06, Tan06, Tan07c, TS08, Tij10, VVLB12, WKK^{+}07, WvE12b, WY06, WH10, Wij12, WG10, YKCL12, YWC08, YLL06, YB10a, Zha08, vR08a, vR08c, vR08b, HML02, JHW^{+}08, KEA^{+}05a, Plo10]. systematic [KS13, Su03a, Su03b]. systemic [Kar07]. Systems [Ano01d, Cer10, Cho10b, Ano03e, ADD^{+}06, BD10, BDH12, Beg05, BBMN03, BBMN08, BPT07, BP06b, Cab12, CG08, CKT06, CL08, Cho13a, DA05, EW10, Egg05a, FM05, GOR06, GBF^{+}12, HFW09, Ho08, HBSZ13, Hol08, HO09, HHfHC10, HCC04, Iye01, JG03, Kal05, KWR^{+}07, Kle08, KSS03, Lai09, Lap07, LLL09, LDL^{+}08, LZC09, LZ08b, Ma07, MS09a, Mil12, MC03b, O’B07, PPN07, PDU02, PS03, Plo10, Riv13, SH07a, SEDC11, SCZZ13, Sri07, Ste03b, SKM11, Sun12, TC04, VŽ08, War05, Yag06, YL06, YNF09, YBC^{+}06, ZC04, ZC01, ZS11b, CKT06, Fou01b, Sch03]. Szostak [Hjø10b].

T [Ano06b, Kal01, Sti01b]. table [Cro09a, Orm01]. tables [KH12, LBMO11, Mus05]. tactic [Wil04]. tactics [XJ10]. Tag [SBNB11, BGBIBM13, DJF^{+}10]. Tag-based [SBNB11]. taggers [WOB09]. tagging [BIZGMS10, DJZ^{+}09, DJF^{+}10, GALC11, GKE11, HJ13, LS12, LDS^{+}11, RRG11, SL13, XMCM13, Yi12, Yoo12]. tags [AdG12, HC11, LS12, Ror10, SJ10, SS13, Yi10]. tailoring [CL12]. Taiwan [Cha09, hChC05]. Taken [Bur03c, Egg03]. takes [Van12]. taking [HLR09]. tale [Chu07, LT12, MS09b, RH12]. talk [Luy12]. tangled [Lyn01]. Tapscott [Pon01]. target [TW03b, TR05]. Task [LCB^{+}07, AYN13, BBI12, BT05, DW04, DWA13, HTT07, JJ10, KdHWA04, KA02, KAB05, KJ12, LM09, Li09b, LB10, MPT09, MB13, PK11b, Tan07a, TTT12, VH12, XTY06, Yi08, SSM^{+}08a]. Task-based [LCB^{+}07, KAB05, KJ12, Yi08]. task-related [AYN13]. task-technology [DW04, DWA13]. task-urgency [MB13]. tasks [Bil01, Bil02, Bys02, BH05, HC02, KWS07, Lan06, MW12, Sun12, WRM^{+}02, WHTS12]. Tawney [Duf11]. taxonomic [CH07]. taxonomies [Pom05b, Sac06]. taxonomy [HLC03, LBTN09, PK05, RPWM06]. Taylor [SEDC11]. TEACH [Lip03]. teachers [Chu09]. teaching [BSM^{+}05, Chu05, Col05, Coo07, Gaz07, KT08, Kra02a, MQH13, ST08, Opp10]. team [SC11, Wor07]. teams [AT02, BT08, LBNB04, MG07, SHS^{+}11]. tech [Fou01a, Lai09]. Technical [ZL07a, ALT09, KMK03]. Technicon [Plo10]. technique [LWH11, SPR09]. Techniques [Van03, ASAK04, ABJ^{+}06, ADD^{+}06, BBB^{+}08, Bur06b, CC01a, DBP^{+}07, DARK09, GS03, KASGAM09, MO04, OYM09, Oya08, TL05, WW13, YB10b, ZLCH06, vEWDvdB10]. techno [PDJCRBBM12]. techno-scientific [PDJCRBBM12]. technoculture [Nay07b]. Technological [Gal12b, WL06b, Has03a, Lip03, PGL13, RPWM06, TWW^{+}06]. technologically [Bre05]. technologie [MDM06]. Technologies [Cho10b, Fau09, Gup09, Hal09, Lar09, LMH09, Lun09, MR09, Mar09b, McQ09, Rin09, Sau09, SHS09, Ehi02, Fou02, HPGG^{+}07, Ley04, LR11, PGFMP13, SSM^{+}08a, SSM^{+}08b, VC09, VC08, YZLC13, Dum01a, Bri10, Ivo03, Gaz07]. Technology [Ano01e, Ano06b, Ano08b, Ano08c, Ano10, Ano11, Ano12, Ano13b, Ano13c, Ano14, AFR07, Cho09b, Ekb12, Has09, Jam10, JP08b, Lar10a, Mos09, PN06, Pal08, Rob10b, Saw10, SH05, Tan07a, VC06, WCC10b, Ago07b, AKS11, BWD02, BBB^{+}08, DW04, DWA13, Day02a, Fle06b, Gu02, Har07a, Hol01, HLC05, Jac01, KRK^{+}10, KTR10, KO05, Lap07, Lim07, Lip13, MA12c, Nah07, OT08, PM11, SHO11, SH07a, SBT^{+}07, Su12, SZ08, Sun12, SK13, VC09, Vis05, VC08, vRD09, vRC11, Ano03f, McQ09, SML12, Fou12, Enn01b, Hol01, Enn08a, Gal06, Ma07, MC03b, Sch05, Tom05b, Vit08]. TED [ST13]. TEDS [SEDC11]. teenagers [AHH06a, AHH06b]. teens [Ago08, AAN12]. teledensity [MKMM09]. Telemedicine [KMD06]. teleophthalmology [KMD06]. telepistemology [Hal01]. Telerobotics [Hal01]. televised [DC09]. television [Fat07]. tell [ESF^{+}08, Jap07]. Temporal [BJC^{+}07, ACH^{+}04, DH13, JSP05, LWY01, ZQI^{+}10]. tentative [MBG03]. Term [Los01, NRD11, BR01, BKRN03, CL01, CMH10, DLA^{+}10, DML02, DGKS04, Efr10, HGGL05, HCO03, JLR04, LAG08, Liu10, LB12c, LB03, Los13, LT02, McD11, NSQI03, VBMM11, WL06a, WZ02, YLZ10, YHWL13, ZW01, ZWW^{+}08, Wan06]. term-term [BKRN03]. terminological [Bod09]. terminologies [MK13]. terminology [Gre01a]. terms [AMFiA11, GW12, KW01, LB05, Ley06, LP09, LS11b, LB11a, LZB13, PCY02, Ror10, SM13, SS13, ZL07a]. terror [ESL^{+}10]. terror-related [ESL^{+}10]. terrorism [HKLG05]. test [LK06, OJ04, Ror07, HWD03, Wal04b]. testbed [DBP^{+}07]. tested [Tol09]. Testing [ADC10, CCC10, Kos05, MLO02, BL11, Jon02, JHW^{+}08, CL08, FN11, GSB08]. tests [BL13]. Text [HFC10, Sri04, The05, AN09, Ano11, AWC^{+}07, AMFiA11, BLH^{+}10, BR01, BSK13, CLH13, CMH10, CSA^{+}02, DW13, DRM^{+}04, DSG^{+}12, Duw06, DARK09, FS10, GS03, GRJS01, HZ03a, HSSV07, HB11, HC07, JLB08, KASGAM09, KDL02, KdRH^{+}01, KB05, KH12, LZ08b, Liu09, Liu10, LYJ^{+}10, LZG13, MH11, MC12, MG09a, MG10, Mor10, MK13, MAR04, Mus05, OW04, Rig02, RPYZ07, SLN03, TBP^{+}10, Tse01, WY06, WH10, WRM^{+}02, WHTS12, YL05, YLL06, AYX^{+}13, War10d, And02]. Text-based [HFC10]. text-classification [KASGAM09]. textbooks [Mel04]. textiles [McQ09]. texts [CKB^{+}08, DLVÁ12, LC05a, LH11, WS06]. textual [CZT06, MLO02, WP05]. texture [LP03]. their [Ben01b, BL09, CA07, CK08, CLL^{+}07, DAB^{+}07, Day11, Egg10c, GRSFV^{+}12b, GRSFV12a, GLDM02, HHWD05, HW05, Ju07, KGL12, KW02, KY04, LSB13, LV06, LS11b, SWJS01, SZ08, TW10, Tra08, VC08, Wil10b]. Thelwall [Ano11]. them [Sch12, Sch13d, Kal05]. thematic [FPZAGBdMA05, HM13, KP03]. theme [CL11, Hjø01b]. theme-based [CL11]. theorem [Duf11, Egg10a]. theoretic [ASW05, DK07, ESK06, War08, ZI01]. theoretical [BLH12, CCC10, GPG^{+}11, JR10, KO05, rP08a, rP08b, The06, VW12, Ye11a, YR08, ZZ03, AHH06a, SWF^{+}02a]. Theories [RI08, Ahn11, Hjø11d, Hjø11c, Ma12a, VG06, Van11, Ago07a]. Theorizing [Jap08]. Theory [Cox09, Dav11, Egg13a, Fan10, For13, Hal10, Hil11, Koe11, Luk09, San10, Ack02a, AKS11, BN08, Bea06, Buc08, Bur06a, Bus07, Cas03, Col11, Dai06, Des08b, Egg05c, Egg06, Egg08b, GD07, Har07b, HPGG^{+}07, Hav02, Hjø01b, Hjø09a, Hjø10b, HBSC13, Jör07, Kar07, KTL^{+}13, KJ03, Ley05, Ley10, LCW06, McK04, MvL12, PM01, Plo10, RK01, RI08, Sag06, SR07, Su06, Szo10, Tad06, War07a, War07d, Whi07a, Whi07b, Wor08, Xu07a, XL07, ZIB01, Su03a, Egg07a]. theory-grounded [Wor08]. theory-practice [Har07b]. therapy [Gen06]. there [BMD08, SM13]. thermodynamic [Pra11a]. thermodynamics [De 12, Pra12b]. thermodynamics-bibliometrics [De 12, Pra12b]. thesauri [CLL^{+}07, Gre01a, TK06]. thesaurus [Ano13a, BCT06, LY05, LVdA^{+}04, MH11, SR06, Tse02, Wan06, WL13, YL03b, SL10]. thesaurus-based [BCT06]. thesaurus-enhanced [SR06]. Thierry [VT13]. thing [BR13]. Things [Els09, KH07]. think [Nay05a]. Thinking [Cro13g, Has03b, BM12, Sar07a, Sar07b]. Thinks [Les11]. third [Sag06]. Thomas [Joh01]. thought [Cro13d]. Thoughts [EGR11, Bur12]. thousand [Rob01]. thousands [CSL03]. thread [VW12]. threats [SH05, Wor08]. Three [TW03b, CH12, CBDH08, Chu09, DBM02, HZ12, KDL02, Ma12a, PDU02, STS02, TW03a, VY12, WIY01, WWZ09, Ye11b]. three-character [KDL02]. three-descriptor [WIY01]. thumbnails [WRM^{+}02]. Thuraisingham [Cul01]. Ticoll [Pon01]. ties [MG07]. Time [BBB04, MYWA03, All11, AWBB09, BI02, Bur04a, CMFZ01, Che03d, DO01, Efr10, Egg07b, HB08, KP03, Ler13, LBMO11, LT13, MBC10, McC08, MB13, Pet02, Slo07, TDC13, Van12, Whi01, WT10, ZZL12, dBvSW10, Åst07, Kar06]. time-constrained [All11]. Time-sliced [Åst07]. Time-tracking [BBB04]. Timelines [CM07, Rou06]. times [GRZ12, War02]. tissue [KR08]. tissue-centric [KR08]. title [Kil01, Kil04, MSYD11, TSW12, Wan06]. titles [HB07, PCR13]. tobacco [Hen13a]. Together [Ekb12, CLL^{+}05, Cro04]. Token [Bur03c, Egg03]. Token-Taken [Bur03c, Egg03]. Tony [Fos01]. too [Ami01]. tool [ACHVH10, CLHHVH12, Kwo05, MBBS02, MTS^{+}13, NJN08, NBYA03, RPL10, RMC12, Tan07b, Wij12, Win08a]. toolbox [Ley09]. Tools [Coo07, Nel06, BBB04, CWZ^{+}10, CLHHVH11, ERJ07, GLDM02, HH01, Mar07c, SB07a, SF08, Sre02, SH05, Ahm07]. top [Bor13b, CvLB12, PBN02, PD08, SRS^{+}13, WJR05, ERR07]. top- [PD08, SRS^{+}13]. top-cited [Bor13b]. TOP-curves [ERR07]. top-ranking [WJR05]. Topic [Ano01g, Arg03, Che03c, Din11b, LMNS13, MH04, TB13, ACK12, AH05b, AM11, Ano01h, Ano01f, Ano01i, Ano02c, Ano02e, Ano02d, Ano02f, Ano03e, Ano03g, Ano04c, Ano05, Arg06, BIP04, BIP09, CKB^{+}10, Che05, ER08b, GO03, HZ07, HVP06, LYM07, LS10b, LDS^{+}11, LW12, MC12, Mos04, NB06, OCO08, PK05, QCL^{+}12, WLC09, WHP10, WHTS12, XBM13, YL06, YL09, YHL11, Yi08]. Topic-based [Din11b, TB13, LW12]. topic-focused [PK05]. topic-hierarchy [YL09]. Topical [SL13, Cri11, Egg13a, Hua13, KK11, KBP10, YD12]. topicality [GO03, Hjø01b, XC06, XY08]. topically [BJC^{+}07]. topics [Alb10b, Ano01a, CS04, LC13, WT12]. topographic [BC09]. Topological [SKM07, ZZ10, LK10, Meg01]. toponyms [OTP13]. total [Ben07, BD07b, ZZL12, Bla10]. Totality [Day01]. traces [KTL^{+}13]. Tracing [Gun13, TH13]. Track [Mea10, Dan11]. Tracking [Enn03c, SH07c, ACK12, BR12a, BBB04, CKB^{+}10, CCMM02, LHB^{+}08, Haw01, HLC03, WV13]. Trade [HPGG^{+}07, ADC10, CM08b, SL12]. trade-off [ADC10]. trade-offs [SL12]. tradeoffs [Wag03]. trader [SC08]. Trading [CF06]. Tradition [MQH13]. Traditional [Esc10b, HB09, KJ03, MBG^{+}07]. traffic [HM03].

Training [GCR09, LT04, CBS10, HC07, KC12, MW08, SF08, Sti05]. transaction [Dav04a, HHvdHdR10]. transactions [Ago04b]. transactive [HBSC13]. Transfer [Col07b, Dav04b, FFMPM13, KK10a, McQ09, YD13]. transformation [KSF04, LB06, SRDZ06]. transformations [Egg08a]. Transformative [PM11]. transformed [War02]. Transforming [Ago04b, Det11]. transforms [Joh01]. transient [Che06b]. transition [Ben05, BHDK13, Chr03, CBL^{+}13, Mac06, Men03, Chr04, Buc10]. Transitions [XJ10, LAL08]. transitive [LKJ08]. translated [DS10b]. Translating [CLYL06]. translation [FCjM12, LKJ08, LR13, Nie03, WTLC06]. transliterating [CLYL06]. transliteration [FR08, Kar03, KLY08]. transliteration-pairs [FR08]. transmuting [LX13]. transparency [Pel11, Nay05b]. Transposition [PCBL04]. traps [Ben07]. travel [DW04]. traversal [ZATH09]. TREC [KSK01, VG07b]. TRECVid [TJSL11]. tree [DML02, JTLG12, JTDG13, WS06, YLW10]. trees [Yag06, YWC08]. Trend [ACH^{+}04]. Trending [WT12]. Trends [Ami01, BR08, LJLlW08, ZC04, Che06b, DO01, KP11, NBG11, RFV10, SSS13, WBY03]. trendy [NBG11]. triangle [CvL12, Tij10]. triangulation [Wij12]. tricitation [McC09]. Trigg [Fou01b]. triggers [SU11]. triple [KP11, LS09, SK13, YYL13]. true [JTLG12]. Trust [KFW08, LC05b, ADD^{+}06, CL06, HFW09, IA12, KH09, LS11d, Lyn01]. Trust-based [LC05b]. Trustworthiness [ADR11]. truth [Cox08]. TSDW [LSD13]. TT&T [PPH06]. TTF [DW04]. TTR [LDS^{+}11]. tune [PSB04]. Tuning [Gua11]. tunneling [HHHR13]. Turkish [AŞT06, CKB^{+}08, CKB^{+}10]. Turkle [Ekb12]. turn [HLR09]. turn-taking [HLR09]. Turning [FW09, LBMO11, RHWB11]. TV [EM07]. TV-Anytime [EM07]. Tversky [Ror07]. Tweets [JZSC09]. twelve [AJ10]. Twenty [Har04, Mar07a, Mar07b, Ago08, Zin07a]. Twenty-eight [Zin07a]. Twenty-First [Har04, Ago08]. Twenty-five [Mar07a, Mar07b]. Twitter [BL12, GGR10, JZSC09, MSC13, NBG11, TBP11, WT12]. twitterers [BL12]. two [BGY^{+}13, CJA^{+}04, Chu07, ER02, HSSV07, Jär07, KDL02, KW02, KHC12, LC05c, MH11, MS09b, PGSH06, RH12, WLLH09, WCC08, WCC10b, vEWDvdB10, Bur05, CCMM02, Hei06, LBNB04, LSD13]. two- [KDL02]. two-dimensional [KHC12]. two-factor [WCC08]. two-level [WLLH09]. two-stage [MH11, LSD13]. two-step [CJA^{+}04]. Type [Bur09b, Ano06b, BDBA13, De 12, HTT07, HRK^{+}06, LCB^{+}07, PK11b, Pra12b, RJ08, RY08, ZRG11, Bur03c, Egg03]. Type/ [Bur03c, Egg03]. typed [CSA^{+}02]. types [Chu09, NV09, PNHG07, PK11b, SV05, SKTM09, ZS09]. typology [Hua13, MBG03].

U.S. [DC09, GKL09, GGR10, TH13]. UK [FMMF13, San08]. ultimate [SSPA10]. Uncertain [Wil10a]. Uncertainties [Sch13d]. Uncertainty [Wil10a, DW04, FS11, Ley12, LAL09, Ano01h, WFE^{+}02, Egg07a]. uncited [Hen13b, MM10]. uncitedness [Ame12, Bur12, EGR11]. Uncovering [CCQ^{+}08, VW12]. underdeveloped [CC09]. Undergraduate [Mac04b, MB13, Zha08, DC01b, KP03, LCB^{+}07, Mac04a, PGSH06, WRB^{+}09, Whi02d]. undergraduates [CLB^{+}05, CLL^{+}07, Su03b]. underlife [Tra08]. underrepresentation [Ago07b]. Understand [Wil10a, LMV^{+}12, PSO06]. Understanding [Bur09a, CL06, CLL13, DBT11, Gaz09, KKGB11, KTW05, LCC07, LK10, LCW06, MC12, McD07, RU07, TYS13, UR07, Xie06, XC09, YR08, YC11, ZB07, BD10, BPT07, DSG^{+}12, Hep04, Kim06, MW10, Vei09, Ago04b, Des08b, Gaz07, Kal05, Vei06a]. uneven [WM06]. Unfounded [SK02b]. unification [Ye11a, Ye11b]. Unified [LKL10, BEK^{+}01, BR04]. unimportant [RN09]. Unintended [Ago04a, vDH12]. union [CH05]. unique [KT07c]. unit [LR10, SL06]. unit-based [SL06]. Unitary [Hoe03]. unite [GALC11, Kar06]. United [War02, CTS04, HLB05, KBP10, NHCW13, NHL^{+}10, Shi08]. units [Lim07, OM13, Zha12]. Universal [War09b, Mar07c]. Universality [BD09, WvEvR12, GBF^{+}12]. universitie [HO09]. Universities [BdMAM13, GRSFV^{+}12b, GTHZ05, LS11b, RN09, vR08a]. University [Cox09, Ano03d, BMD13, CKR13, KP11, Kim10b, Ley04, OM13, SLS06, The02, TH03, TW03b, The05, YYL13, Bai04, LS09]. university-industry-government [KP11]. unknown [WTLC06]. unlayered [FL01]. Unlocking [Jör04]. unsupervised [AK13, CFN^{+}10, WLL10]. Untangling [Egg07c]. updates [Ley02]. upon [TH03]. Upper [DJF^{+}10]. Urban [Mos09, AHH06a, AHH06b, Bur02a]. urgency [MB13]. URL [KT07b, LS10a, TSW12]. URLs [Dav04a]. Urquhart [Ben05, Ben01b]. USA [Ano01e, Pel11, Ano03h]. usability [BLA07, Bol07, CGY09, GSB08, JG05, JHW^{+}08, Ma02a, MBC08, XW02, ZL07d, Jon02, Vil10b, Wal02]. Usage [BV08, Coo01, ALT09, BHC06, Bro03a, Bro03b, CC01a, CC02, DS03, DP06, HLLS07, HW05, HNJT06, JdMAA12, LCW06, LT02, McD07, SG11]. usage-based [BV08]. usages [MTS^{+}13]. Use [DW04, HZ12, TS01, AXP11, Awa06, Beg05, BMK03, Bil01, Bil02, BPJR07, Bur06b, CKR13, CH05, CWZ^{+}10, CBDH08, Cho09a, CN07, DJWR02, DAB^{+}07, FDH04, Fle06b, GGR10, Har02, Har08b, HML02, HL05, HDG^{+}07, Hol01, JJ10, Jar11, JP02, KEA^{+}05a, KO05, LKK03, Ley08a, Lim09, LHHVH09, MW10, MQH13, MPT09, McD07, MBBS02, NY08, NS12, O’C13, Pal08, PM11, PM01, RK01, RLvR02, Sav06, Sav09, SB07c, Sha06a, SD03, Slo07, SC06, TVFW07, TL05, Tom06, TSRGJC^{+}13, Tre12, Tse01, Vak08, Vau03, VT03, VBMM11, Whi02c, Wil02, WT08, WRM^{+}02, YH04, YZLC13, Zha01, ZS09, Col05, Kap09]. used [Bor13c, DBT11]. User [BIKYL07, Bea06, BR12b, CL09, CS06a, Dav06, DP07, Efr12, GD01, Gre06, Hen06, HLC05, JBCJ06, JG05, Kal06, KO09, KSM06, MS02, Moc06, NY08, OTP13, RV06, Ros06, Ros08, Rot06, Sac06, VR06c, Vil10a, WL06a, Wil06a, YBC^{+}06, ZZZR04, vRFBvW10, Abb05, APLG13, AMS10, Alb10a, Alb10b, BIKLY09, Ben01a, Ben04a, BBMN03, BBMN08, BCT06, BMTV02, CTZ05, CHH09, CR03, CL03, CLL^{+}05, Cop11, Cot02, DS03, Day11, Erc06, GBK05, GO03, HML02, Hep04, HBSZ13, HTT07, hHyW04, HSCL07, HJ13, JZS09, Ju07, JLB08, KKGB11, KLP09, Kos05, KdRH^{+}01, LM01, LCC07, Lee10, LWH11, LFC01, LKK07, LT13, LR13, MBG^{+}07, Mar02, Mar07a, Mar07b, NSQI03, OT08, OT10, OL13, OKG07, POC09, Pat11, PBS^{+}04, PLBS06, Pet02, PCY02, RLvR03, SKM01, SDWG01, SR06, SG01, SBBF09, Ste08]. user [Su03a, Su03b, SZ08, Sun12, SL13, SS13, Tan07b, TD02, TSA11, VŽ08, WEA^{+}12, Whi02c, WIY01, WL11b, YRW07, YSD^{+}08, Yoo10, Yoo12, YR02, ZC01, Zha02, ZL08, ZATH09, ZL07d, Ben04a, EWF^{+}02]. User-based [Ros08]. user-centered [Erc06, HML02, Pat11, PBS^{+}04, SR06, WEA^{+}12, ZL08, ZL07d]. user-defined [Abb05]. user-generated [SS13]. User-process [JG05]. User-producer [vRFBvW10]. user-side [KLP09]. user-subjective [BBMN03, BBMN08, GBK05]. user-system [BCT06]. Users [KAB05, BAB07, CG06, FMCL07, HTT07, LBR02, LDL^{+}08, Lip03, NCRJ13, NS07, PBS^{+}04, Rad06, SK12, TRJ05, XC06]. Uses [AdG12, SHO11, CK08, PGSH06]. Using [APSA12, Ano13a, CC01a, CL03, CM06, CJ04, HS01, HWD03, KJ02, KWR^{+}07, KDL02, KB11, LMV^{+}12, MBC08, MCK08, McC09, MAR04, NS07, Oak09, Opp07, Oya08, PSO06, SCC^{+}12, SPH08, SKM11, TVY12, WKdJvdBV01, WJR05, WOE^{+}13, WHTS12, dMFRN^{+}10, AJ10, ADR11, Ale03, ACHVH10, AWC^{+}07, AR13, AMFiA11, BWW03, BBB04, BMD08, BD09, BSMD11, BL11, BCT02, BWD02, BSK13, CH03, Che08, Dan11, DLA^{+}10, DBP^{+}07, DDO11, ESL^{+}10, GCR09, GdMA01, GLLL11, HKBY^{+}10, HFW09, Hen12, HV01, HASAT^{+}10, HDB07, JBLOHM10, KK04, Kil01, Kil04, KR08, Koh13, KTR10, KHC09, KC09, LM01, LHLL08, LKJ08, LCB^{+}07, Ley06, LZ08a, LRR12, LB12b, LO13, LK10, LSD13, LLXS13, LZC^{+}07, LR10, LVdA^{+}04, LGM^{+}09, Ma12b, MH11, MPS12, MSWK11, MC12, McC08, MG09a, MG10, MC03a, Muk09]. using [MPR^{+}03, MMMM^{+}01, Mus05, NZY13, NSLD13, OW04, PCR13, PLN09, PDU02, PAL12, RTM04, RPYZ07, RHCP12, ST08, SU11, SRDZ06, SLS06, SY07, SKTM09, SB12b, SJTB13, Sta11, SZC11, SLR^{+}11, Suz09, TYS13, Tan04, TCW11, Vau06, WS11, WmcsW09, WWZ09, WOB09, WIZ01, WL08, WCC08, WCC10b, YWC08, Yi08, Yi12, ZWW09, ZJS09, ZA13, Cha09, CH12, GLDM02, Nic03a, VC06, WL13, WT08]. USPTO [LB12b]. utility [Wal06]. utilization [Tij10]. Utilizing [Gre06, YHWL13, Yoo10, Laz07, Zha12].

Vagueness [War10b]. valid [Moe11]. Validation [Kim09b, jLB05, BD09, CCCJ10, KK08, MM04, OCO08, SH07c, SSPA10]. validity [Ben01b, BSMD11, VvdHH^{+}13]. valuable [Tan07c]. value [Col07a, Day02b, DGKS04, Enn04, GPG^{+}11, HRC12b, KdHWA04, LFC01, LR13, MSWK11, PN06, Pal08, RR06, Rab09, Tom04, FS11, SEDC11]. Value-Added [SEDC11]. value-adding [LFC01]. value-based [DGKS04]. values [CFW^{+}12, Egg02b, KH12]. variable [MEH07]. variables [BW06]. Varian [Tom01a]. variants [BMD08, BMD09, Bor11]. variation [Ale03, Che12, WLC09]. Variations [LTO11, LAG08]. vary [SM05b]. varying [Bys02]. vector [BBM02, DRM^{+}04, Egg06, Egg09a, GBK05, KS01, Lak03, LSZ04, WOB09]. vectors [Egg09a]. Vegas [Ano01e, Ano03h]. vehicular [Cro09f]. Venue [NSJ13]. Venue-author-coupling [NSJ13]. Vernacular [Cro09f]. version [Bur06a, Erc06]. versioned [HZ03b]. versus [Ben08, BL09, BGBIBM13, CD06, FT12, GOR06, HKKP13, HO09, HNB^{+}09, Jam08b, KSS03, Ley08b, LS09, LO10, LS10b, MY07, NDD^{+}09, SS04, SG11, Shi08, YH04, AL08]. very [ACHVH10, BJC^{+}07, Kis10, MH04]. via [CL12, Gre01c]. Vickery [Cro10b]. victims [Wes07b]. Video [MWG06, Ago08, Alb10a, Alb10b, AM11, FS11, Goo01, HFC10, IL07, KK10b, Kim11a, LWY09, SSM^{+}08a, SSM^{+}08b, TLM06]. video-sharing [HFC10]. Videogame [Win11]. videogames [McD11]. videos [KTA12, ST13, TSV12]. view [BAB02, Bur01, Gar01b, HSKS03, KK10a, PCBL04, WL09, YNF09]. viewed [RFV10]. viewership [FS11]. ViewFinder [Alb10a]. viewpoint [Mar11, Sav06]. viewpoints [FCM04]. views [CA07, GL12, LBTN09, MM01]. villain [LY11]. violence [DRW08, HW13, Wes07b, Wes09]. violent [HZ12]. Virtual [GSB08, AT02, BBC01, FG01, HBSZ13, KSF04, Mon04, O’C13, Rad06, SOH08, SK12, Tad06, TTT12, Wor07, Kom11, Koe04a, Koe04b, Sti05]. Virtue [Bla10]. viruses [SSB01]. visibility [DLM09, ECP08, MCRdMA11]. Visible [CSL04, Enn04, Hjø02d]. vision [Enn06a, GS11, Lun09, Ste03a]. Visioning [SS10]. Visions [Bea06, Dav06, Gre06, Hen06, JBCJ06, Kal06, KSM06, Moc06, RV06, Ros06, Rot06, Sac06, VR06c, Wil06a, Sri07]. visiting [HC09]. Visscher [De 13, Egg12]. Visual [ASQR09, HT11, Iye01, TCW11, ZATH09, Alb10b, BLH^{+}10, BJ09, CK03, CHMS13, GRZ12, HY09, JJBC01, LBTN09, MLO02, RHS10, RCD08, TL01]. visualisation [MCDA03]. Visualization [Kos09b, Ley07c, Su12, ZW01, ZWW^{+}08, BWD02, CC11, CPO01, DSG^{+}12, FMM10, HH01, JLB08, Kos05, Ley08b, MYWA03, ZWW09, Ano01i, Sch04a]. visualization-based [Kos05]. visualizations [ASR^{+}10, HS01, LCB^{+}07, vdHR11]. Visualizing [CCMM02, Che03b, JBLOHM10, Nea11, SCLC06, SYL^{+}13, Vau06, dMAVQCR^{+}07, Che06b, BB03]. visually [STS12]. Vivisimo [KSJ06]. vocabularies [Abb05]. vocabulary [LMRB08, MM01, QH06, TS08, vHVD^{+}13]. vocabulary-based [LMRB08]. Vol [Vit08]. volume [Sch05]. volution [Bro01b]. VOS [vEWDvdB10]. Vox [WSJS01]. VR [MC03a]. vs [RKO10]. vu [CS06b]. VxInsight(R) [BWD02].

W [Ano06b]. Waddington [McC08, McC09]. Waikato [Bai04]. wait [RR12]. walk [Ano13a, WL13]. Walter [Ham11]. want [KAB05]. war [RB08]. warning [Cho09a]. Watch [Mea10]. water [TVF06]. wave [Ye11a]. wave-heat [Ye11a]. wavelet [LP03]. Way [Det11, BM12, Min10]. weak [MG07]. weakening [LLG12]. weaknesses [HB09, Sti01a]. wealth [CFW^{+}12]. weapons [SSB01]. Weaving [Reu05]. Web [Ack02b, Ano03d, Bro02, Che06a, Che04, Cho10b, CC09, Enn01a, Fou01b, Jon02, Kap09, Ost08, Reu05, Sti02, The09a, Wal02, Wil06b, Ahm07, Ano13c, BT08, BR12b, Bur11, Cha11, CMFZ01, CC02, Che03c, Cho13b, Coo01, DRW08, DG12, DS11, Gun13, Har02, HH01, HB11, HC11, HCO03, KP11, LDvM12, LX13, LDM12, Lyn01, MW12, Meg01, OM13, OCJ12, PBN02, RFV12, SWJS01, Sre02, TH03, TW03b, TS11, TBP12, TB13, WBY03, WSJS01, WHC12, YR08, ZS08b, Ago02, ACHVH10, Ano01a, Ano04c, Ano05, ACGL09, AN06, BYHM07, BR12a, BIP04, BI06, BIB07, BIP09, BLT07, BJC^{+}07, BP10a, Bil01, Bil02, BW05, BHC06, Bro07, CCG^{+}06, CS04, CFS05, CSCC07, CFY07, CFCZ01, CWZ^{+}01, CC01a, CMD05, CCQ^{+}08, CGY09, CL08, CHH09, Cho10a, CCR09, CC09, Col05, Cot02]. Web [Cot04, CG06, CJJ^{+}02, CTS04, DW04, DC01b, DLM09, DBM02, DPV11, DP07, ESL^{+}10, EHD05, ECP08, FFPW04, FG03, FMM03, FMM05b, FMM05a, FAC08, FAC10, Fu10, GOR06, GP10, GRSFV12a, Gor02, GS11, Gwi10, HZ07, HLLS07, HVPLHP06, HVP06, HY09, Hon06, HTT07, HCH^{+}05, HPAS04, HSCL07, HC09, HC07, HD06, Iye01, JP01, JSP05, JR06, JSBK07, JSK07, JBS09, JSI^{+}04, Käk06, Kal06, KS08, KOS09a, KWS07, KK04, KA02, Kim05, Kim06, Kim09a, Kim10b, KBP10, Koe02, KSJ06, KT07b, KT07c, KC07, KLY08, Lak03, LYM07, LBR02, LBC02, LBNB04, LC06, Laz07, LM09, LV06, LO13, LY05, LY06, LfBW08, LZC09, Li09a, LHA07, LZC^{+}07, LPSS06, LT02, LT04, Ma02b, MPS12, MH10, MY07, MR08, MS09b, Men04, Met07, Moc06, MV08, NLC06, Nic03a, NSE^{+}06, OW04, OAU11, ÖAU^{+}04]. Web [PK05, PZ05, PMBdS10, PRNZ^{+}11, PR07, PCBL04, PCY02, QZCC06, RLF02, RFQ^{+}05, RTM04, Rie02, Rie04, RHS10, RFV10, Ros06, Ros08, SWT04, Sel05, SEMK05, SLS06, SY07, ST09, Slo02, Slo07, Spi02, Su03a, Su03b, SL06, TS09, The01, The02, TH04, The05, TP06, TS06, TVF06, TSJ09, TRJ05, VS03, VT03, VS05, Vau06, VTD10, VY12, WTLC06, WY07, WB02, WCLZ01, Whi02b, WWZ09, WRM^{+}02, WdV04, WIZ01, XJ10, YC02, YL09, YR02, ZZZR04, Zha08, ZJS09, Zha10b, ZS07, Zuc06a, dBvSW10, dMFRN^{+}10, vRFBvW10, Fos01, Che10b, Des07, Koc06, Par10, ZL07d, ZL07b]. Web-accessible [Enn01a]. web-based [BR12b, CC02, Coo01, Ago02, Bro07, CC01a, CL08, HC07, KWS07, Kim06, Nic03a, PRNZ^{+}11, TS09, XJ10, Che10b]. Web-crawling [Cot04]. Web-located [Sel05]. web-prevalent [Ano13c, WHC12]. Web/URL [KT07b]. webcasts [DBT11]. weblib [Sre02]. weblogs [Luz09]. Webometric [FPZAGBdMA05]. webometrics [BI04, Ano02c, TV04, TKV^{+}10]. webpage [TCW11]. Webs [PT05, TW03a, Pon01]. website [Cha11, LMV^{+}12]. websites [BIA12, Rab09, Reu05]. weight [DLA^{+}10, WvE10, WL08]. Weighted [LYJ^{+}10, YD10, Din11a, ER03, RTM04]. weighting [Efr10, NRD11, WZ08, ZS08c]. we’ll [WK13, vEW09]. well-known [vEW09]. Weller [Sti01a]. wellness [WP04]. were [Ano08c, BD08]. West [War09a]. Western [LDM12]. Wheeler [Pet02]. Where [Kos12, PZ12, Dil02, HO09]. whether [WvE07]. Which [BLT07, BL11, PLBS06, LTO11, BK10]. while [JW09, PV03, Sau09]. Whitney [HWD03, Wal04b]. Who [DC01a, EHD05, RKBS11, Tan07a, XTY06, CA07, Wal11a]. Wide [Cot02, DW04, Lak03, Ma02b, YR02]. widely [Wu10]. Widén [BI13]. wiki [SP11]. Wikipedia [AYN13, CH11, CCC10, Fal08, GS11, HSH10, LS11a, LSD13, Lim09, LATH08, LT10, Luy11, Luy12, Luy13, MSWK11, STSG08, Tsi13, YHHL12]. wikis [ASR^{+}10, YWM10]. wikisphere [SP11]. Will [Bla10, Bur02b, Jar11, Tan07a, Tre12, XTY06, LT04]. William [Hjø01a, Hol01, Pet02, Sch01]. willingness [RR12]. win [RKBS11]. window [Che03a]. windows [The08a, ZSR09, Nay05b]. winners [DC09]. winning [Har07a]. WIPO [Lip03]. wired [Plo10, Enn08b, Saw10]. Wisecrackers [Wor08]. Witheford [Day02a]. within [Cha09, HWT^{+}13, Hen12, KH12, LT09b, MPT09, Moe05, SR06, WJD^{+}13, WW13, XC09]. without [CW04, CBS10, EW10, Hav02, SD01]. Wolfgang [Pee12]. women [Gen06, LKG^{+}10, PK11a, YR08, WP04, Ago07b]. Wooster [Wil06c]. Word [Ano06b, HRK^{+}06, Mus05, AR13, CW04, GOR06, JZSC09, Kil04, LZ08a, LSD13, LW12, TP06, ZLB11, SD01]. word-based [LW12]. word-matching [GOR06]. Word-oriented [Mus05]. words [HVPLHP06, HASAT^{+}10, KDL02, Kil01, MSYD11, PT05, Rob01, TB13]. Work [CKR13, Cox10, HC02, Lan06, ALT09, Ano08a, Bat08a, Bat11, CH13, CSL04, Dan11, GLDM02, HSJB12, Huv13, Jay12, Li09b, LB10, Mar07d, MPT09, Nel06, Pal05, rP08a, rP08b, RMC12, SV03, STSG08, Vak08, VV12, Wil00, Xie06, Yak05, Gal12b, McI10, Wal02]. workbook [Whi02c]. Working [AKS11, hChC05, FW09, Mos09]. Working-Class [Mos09]. workplace [ALT09, FFMPM13, Fou01b]. Works [Day08, Les11, Nic03a, Cul01]. Workspaces [Hen06]. World [Luy13, Tra09, Wil10a, Ano13b, AYSC06, CZH^{+}04, FC01, GSD12, KJ03, Ler10, MHR13, SCLC06, SSS13, TTT12, XMCM13, Cot02, DW04, HPGG^{+}07, Lak03, Ma02b, YR02, Bar01a, War10d, Wil06b]. Worlds [Swa11, BBC01, GSB08]. Worldwide [KEA^{+}05a, vDH12]. worth [HB08, Rob01]. WoS [ACHVH10]. wrapper [HCH^{+}05]. Writing [Bia10b, CTZ05, Cro04, Har06b, PM11, ZLCH06]. Writing-style [ZLCH06]. writings [KW02]. Written [CD06, Ano14, SMAB09, Tra08]. Wu [Ano13c, WCC10b]. Wurster [Joh01]. WWW [HS01, Mar02]. www.hotelodisej.hr [Ano03d]. www.hr [Ano03d]. www.pedos.hr [Ano03d].

XML [Hol04, KLFG04, Nel06, AFMM02, BLH12, CK02, CKK^{+}02, CSA^{+}02, Kle08, LLD^{+}02, NJN08, PMK11, PK11b, Rig02, SM02, Kal01, Hol04, Wad01]. XML-based [PK11b, Kle08]. XQL [BYN02]. Xu [Tan07a].

Y. [Tan07a]. Yahoo [Bil12]. Yahooligans [Bil01, Bil02]. Yale [Buc13]. Yang [Tan07a]. Yangarber [VT13]. Ye [Bur09b]. year [Cam11, Coo02a, DC09, GL12, KLFG04, Koe02, Wal11b]. years [Ard13, DO01, Jay12, LSC12, Mar07a, Mar07b]. Yee [Spe01]. Yi [Ano13c]. yield [Cam10a]. York [FDH04]. young [Ago02, Coo02b, WH09, PSWL13]. youngster [SD03]. yours [WK13]. yourself [Fle06b]. youth [DAB^{+}07]. YouTube [Hov13, KTA12, LLD^{+}12, TSV12].

Zipf [Egg10a, Los01]. Zipfian [Egg05c]. Ziv [ANdlF07]. zones [BBRdMA12].

References

Agosto:2012:RSR

[AAN12] Denise E. Agosto, June Abbas, and Robin Naughton. Relationships and social rules: Teens’ social network and other ICT selection practices. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1108– 1124, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Almeida:2009:OKM

[AB09] Mauricio Barcellos Almeida and Ricardo Rodrigues Barbosa. Ontologies in knowledge management support: a case study. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2032–2047, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Abbas:2005:MMS

[Abb05] June Abbas. Out of the mouths of middle school children: I. Developing user-defined controlled vocabularies for subject access in a digital library. Journal of the American Society for Information Science and Technology: JASIST, 56(14): 1512–1524, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Abdelali:2004:LMS

[Abd04] Ahmed Abdelali. Localization in Modern Standard Arabic. Journal of the American Society for Information Science and Technology: JASIST, 55(1):23–28, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aqeel:2006:DNS

[ABJ^{+}06] Syed Uzair Aqeel, Steve Beitzel, Eric Jensen, David Grossman, and Ophir Frieder. On the development of name search techniques for Arabic. Journal of the American Society for Information Science and Technology: JASIST, 57(6):728– 739, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Archambault:2009:CBS

[ACGL09] Éric Archambault, David Campbell, Yves Gingras, and Vincent Larivière. Comparing bibliometric statistics obtained from the Web of Science and Scopus. Journal of the American Society for Information Science and Technology: JA-SIST, 60(7):1320–1326, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Amitay:2004:TDT

[ACH^{+}04] Einat Amitay, David Carmel, Michael Herscovici, Ronny Lempel, and Aya Soffer. Trend detection through temporal link analysis. Journal of the American Society for Information Science and Technology: JASIST, 55(14):1270–1281, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Alonso:2010:WQP

[ACHVH10] Sergio Alonso, Francisco Javier Cabrerizo, Enrique Herrera-Viedma, and Francisco Herrera. WoS query partitioner: a tool to retrieve very large numbers of items from the Web of Science using different source-based partitioning approaches. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1582–1597, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ackermann:2002:BRBa

[Ack02a] Eric G. Ackermann. Book review: A sociological theory of communication: the self-organization of the knowledge-based society, by Loet Leydesdorff. Journal of the American Society for Information Science and Technology: JASIST, 53 (1):61–62, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ackermann:2002:BRBb

[Ack02b] Eric G. Ackermann. Book review: The laws of the Web: Patterns in the ecology of information. Journal of the American Society for Information Science and Technology: JASIST, 53(11):969–970, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ackermann:2003:BR

[Ack03] Eric G. Ackermann. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 54(9):906–908, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aksoy:2012:NDT

[ACK12] Cem Aksoy, Fazli Can, and Seyit Kocberber. Novelty detection for topic tracking. Journal of the American Society for Information Science and Technology: JASIST, 63(4):777– 795, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Azagra-Caro:2011:SLD

[ACMP11] Joaquín M. Azagra-Caro, Pauline Mattsson, and François Perruchas. Smoothing the lies: the distinctive effects of patent characteristics on examiner and applicant citations. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1727–1740, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Allison:2005:DIM

[ACMS05] Arthur Allison, James Currall, Michael Moss, and Susan Stuart. Digital identity matters. Journal of the American Society for Information Science and Technology: JASIST, 56(4):364–372, February 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Awazu:2004:BCO

[AD04] Yukika Awazu and Kevin C. Desouza. Brief communication: Open knowledge management: Lessons from the open source revolution. Journal of the American Society for Information Science and Technology: JASIST, 55(11):1016–1019, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Abramo:2009:DSS

[AD09] Giovanni Abramo and Ciriaco Andrea D’Angelo. A decision support system for public research organizations participating in national research assessment exercises. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2095–2106, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Abramo:2010:TTB

[ADC10] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Testing the trade-off between productivity and quality in research activities. Journal of the American Society for Information Science and Technology: JASIST, 61(1): 132–140, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Abramo:2012:IIT

[ADC12] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Identifying interdisciplinarity through the disciplinary classification of coauthors of scientific publications. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2206–2222, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aringhieri:2006:FTT

[ADD^{+}06] Roberto Aringhieri, Ernesto Damiani, Sabine De Capitani Di Vimercati, Stefano Paraboschi, and Pierangelo Samarati. Fuzzy techniques for trust and reputation management in anonymous peer-to-peer systems. Journal of the American Society for Information Science and Technology: JASIST, 57(4):528–537, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Arolas:2012:UEI

[AdG12] Enrique Estellés Arolas and Fernando González Ladrón de Guevar. Uses of explicit and implicit tags in social bookmarking. Journal of the American Society for Information Science and Technology: JASIST, 63(2):313–322, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Akter:2011:TMI

[ADR11] Shahriar Akter, John D’Ambra, and Pradeep Ray. Trustworthiness in mHealth information services: an assessment of a hierarchical model with mediating and moderating effects using partial least squares (PLS). Journal of the American Society for Information Science and Technology: JASIST, 62(1):100–116, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Abramo:2010:RBI

[ADV10] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Fulvio Viel. A robust benchmark for the h- and g-indexes. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1275–1280, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Abramo:2013:AAI

[ADV13] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Fulvio Viel. Assessing the accuracy of the h- and g-indexes for measuring researchers’ productivity. Journal of the American Society for Information Science and Technology: JA-SIST, 64(6):1224–1234, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Antin:2010:LHM

[AE10] Judd Antin and Matthew Earp. With a little help from my friends: Self-interested and prosocial behavior on MySpace Music. Journal of the American Society for Information Science and Technology: JASIST, 61(5):952–963, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Alberts:2012:EPA

[AF12] Inge Alberts and Dominic Forest. Email pragmatics and automatic classification: a study in the organizational context. Journal of the American Society for Information Science and Technology: JASIST, 63(5):904–922, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aljlayl:2002:BEA

[AFG02] M. Aljlayl, O. Frieder, and D. Grossman. On bidirectional English–Arabic search. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1139– 1151, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Afifi:2003:BRB

[Afi03] Marianne Afifi. Book review: Electronic collection development: a practical guide. Journal of the American Society for Information Science and Technology: JASIST, 54(6):588– 589, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Azagury:2002:NNP

[AFMM02] Alain Azagury, Michael E. Factor, Yoelle S. Maarek, and Benny Mandler. A novel navigation paradigm for XML repositories. Journal of the American Society for Information Science and Technology: JASIST, 53(6):515–525, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Asonuma:2006:RAD

[AFR06] Akihiro Asonuma, Yong Fang, and Ronald Rousseau. Reflections on the age distribution of Japanese scientists. Journal of the American Society for Information Science and Technology: JASIST, 57(3):342–346, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [AFR07].

Asonuma:2007:EAF

[AFR07] Akihiro Asonuma, Yong Fang, and Ronald Rousseau. Erratum: A. Asonuma, Y. Fang and R. Rousseau, ‘Reflections on the age distribution of Japanese scientists’. Journal of the American Society for Information Science and Technology 57 (3) 2006, 342–346. Journal of the American Society for Information Science and Technology: JASIST, 58 (5):762, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [AFR06].

Abt:2002:RBN

[AG02] Helmut A. Abt and Eugene Garfield. Is the relationship between numbers of references and paper lengths the same for all sciences? Journal of the American Society for Information Science and Technology: JASIST, 53(13):1106–1112, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2001:BR

[Ago01] Denise E. Agosto. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 52(1):81–83, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2002:BRS

[Ago02] Denise E. Agosto. Bounded rationality and satisficing in young people’s Web-based decision making. Journal of the American Society for Information Science and Technology: JASIST, 53(1):16–27, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2004:BRBb

[Ago04a] Denise E. Agosto. Book review: Net Effects: How Librarians Can Manage the Unintended Consequences of the Internet. Journal of the American Society for Information Science and Technology: JASIST, 55(11):1025–1026, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2004:BRBa

[Ago04b] Denise E. Agosto. Book review: Understanding reference transactions: Transforming an art into a science. Journal of the American Society for Information Science and Technology: JASIST, 55(4):365–366, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2005:BRB

[Ago05] Denise E. Agosto. Book review: Information literacy: Essential skills for the information age. Journal of the American Society for Information Science and Technology: JA-SIST, 56(9):1008–1009, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2006:BRBa

[Ago06a] Denise E. Agosto. Book review: Creating the customerdriven library: Building on the bookstore model. Journal of the American Society for Information Science and Technology: JASIST, 57(4):585–586, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2006:BRBb

[Ago06b] Denise E. Agosto. Book review: Qualitative research for the information professional: a practical handbook. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1708–1709, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2007:BRBa

[Ago07a] Denise E. Agosto. Book review: Theories of information behavior. Journal of the American Society for Information Science and Technology: JASIST, 58(2):303, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2007:BRBb

[Ago07b] Denise E. Agosto. Book review: Women and information technology: Research on underrepresentation. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1704, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2008:BRB

[Ago08] Denise E. Agosto. Book review: Sex, brains, and video games: a librarian’s guide to teens in the twenty-first century. Journal of the American Society for Information Science and Technology: JASIST, 59(12):2033–2034, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2013:BRB

[Ago13] Denise E. Agosto. Book review: Human Information Interaction: An Ecological Approach to Information Behavior — By Raya Fidel. Journal of the American Society for Information Science and Technology: JASIST, 64(1):213–214, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aguillo:2006:SRA

[AGOP06] Isidro F. Aguillo, Begoña Granadino, José L. Ortega, and José A. Prieto. Scientific research activity and communication measured with cybermetrics indicators. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1296–1302, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Al-Hawamdeh:2005:DIG

[AH05a] Suliman Al-Hawamdeh. Designing an interdisciplinary graduate program in knowledge management. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1200–1206, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Al-Hawamdeh:2005:IST

[AH05b] Suliman Al-Hawamdeh. Introduction to the special topic section: Knowledge management in Asia. Journal of the American Society for Information Science and Technology: JA-SIST, 56(11):1154–1155, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2006:TMEa

[AHH06a] Denise E. Agosto and Sandra Hughes-Hassell. Toward a model of the everyday life information needs of urban teenagers, part 1: Theoretical model. Journal of the American Society for Information Science and Technology: JA-SIST, 57(10):1394–1403, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agosto:2006:TMEb

[AHH06b] Denise E. Agosto and Sandra Hughes-Hassell. Toward a model of the everyday life information needs of urban teenagers, Part 2: Empirical model. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1418–1426, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Azadeh:2008:DII

[AHK08] A. Azadeh, M. Haghnevis, and Y. Khodadadegan. Design of the integrated information system, business, and production process by simulation. Journal of the American Society for Information Science and Technology: JASIST, 59(2):216– 234, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ahmad:2007:BRB

[Ahm07] Ashraf M. A. Ahmad. Book review: Multimedia content and the semantic web: Methods, standards and tools. Journal of the American Society for Information Science and Technology: JASIST, 58(3):457–458, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ahn:2011:AIS

[Ahn11] June Ahn. Advances in information science: The effect of social network sites on adolescents’ social and academic development: Current theories and controversies. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1435–1445, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ahlgren:2010:MIT

[AJ10] Per Ahlgren and Kalervo Järvelin. Measuring impact of twelve information scientists using the DCI index. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1424–1439, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Arencibia-Jorge:2008:BCA

[AJBAFHCE08] Ricardo Arencibia-Jorge, Ismaray Barrios-Almaguer, Sandra Fernández-Hernández, and Rachel Carvajal-Espino. Brief communications: Applying successive H indices in the institutional evaluation: a case study. Journal of the American Society for Information Science and Technology: JA-SIST, 59(1):155–157, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ahlgren:2003:RCS

[AJR03] Per Ahlgren, Bo Jarneving, and Ronald Rousseau. Requirements for a cocitation similarity measure, with special reference to Pearson’s correlation coefficient. Journal of the American Society for Information Science and Technology: JASIST, 54(6):550–560, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ahlgren:2004:LEA

[AJR04a] Per Ahlgren, Bo Jarneving, and Ronald Rousseau. Letters to the Editor: Author cocitation analysis and Pearson’s r. Journal of the American Society for Information Science and Technology: JASIST, 55(9):843, July 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ahlgren:2004:LER

[AJR04b] Per Ahlgren, Bo Jarneving, and Ronald Rousseau. Letters to the Editor: Rejoinder: In defense of formal methods. Journal of the American Society for Information Science and Technology: JASIST, 55(10):936, August 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Arazy:2011:MIQ

[AK11] Ofer Arazy and Rick Kopak. On the measurability of information quality. Journal of the American Society for Information Science and Technology: JASIST, 62(1):89–99, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Akiva:2013:GUM

[AK13] Navot Akiva and Moshe Koppel. A generic unsupervised method for decomposing multi-author documents. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2256–2264, November 2013. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aksnes:2006:CRP

[Aks06] Dag W. Aksnes. Citation rates and perceptions of scientific contribution. Journal of the American Society for Information Science and Technology: JASIST, 57(2):169–185, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aksnes:2008:BCW

[Aks08] Dag W. Aksnes. Brief communications: When different persons have an identical author name. How frequent are homonyms? Journal of the American Society for Information Science and Technology: JASIST, 59(5):838–841, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Allen:2011:WAT

[AKS11] David Allen, Stan Karanasios, and Mira Slavova. Working with activity theory: Context, technology, and information behavior. Journal of the American Society for Information Science and Technology: JASIST, 62(4):776–788, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agnew:2007:PMM

[AKW07] Grace Agnew, Dan Kniesner, and Mary Beth Weber. Perspectives on MPEG-7: Metadata for multimedia: Integrating MPEG-7 into the Moving Image Collections portal. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1357–1363, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Antonakis:2008:QSI

[AL08] John Antonakis and Rafael Lalive. Quantifying Scholarly Impact: IQp Versus the Hirsch h. Journal of the American Society for Information Science and Technology: JASIST, 59(6):956–969, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Albertson:2010:AUI

[Alb10a] Dan Albertson. Analyzing user interaction with the ViewFinder video retrieval system. Journal of the American Society for Information Science and Technology: JASIST, 61(2):238–252, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Albertson:2010:IUF

[Alb10b] Dan Albertson. Influences of users’ familiarity with visual search topics on interactive video digital libraries. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2490–2502, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Alemayehu:2003:APV

[Ale03] Nega Alemayehu. Analysis of performance variation using query expansion. Journal of the American Society for Information Science and Technology: JASIST, 54(5):379–391, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Allen:2011:IBD

[All11] David Allen. Information behavior and decision making in time-constrained practice: a dual-processing perspective. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2165–2181, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Almeida:2013:RON

[Alm13] Mauricio Barcellos Almeida. Revisiting ontologies: a necessary clarification. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1682–1693, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Allard:2009:DET

[ALT09] Suzie Allard, Kenneth J. Levine, and Carol Tenopir. Design engineers and technical professionals at work: Observing information usage in the workplace. Journal of the American Society for Information Science and Technology: JASIST, 60(3):443–454, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ajiferuke:2010:CCC

[ALW10] Isola Ajiferuke, Kun Lu, and Dietmar Wolfram. A comparison of citer and citation-based measure outcomes for multiple disciplines. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2086–2096, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Albertson:2011:STC

[AM11] Dan Albertson and Charles Meadows III. Situated topic complexity in interactive video retrieval. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1676–1695, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Amez:2012:CMM

[Ame12] Lucy Amez. Citation measures at the micro level: Influence of publication age, field, and uncitedness. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1459–1465, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Atlam:2011:NAA

[AMFiA11] El-Sayed Atlam, Kazuhiro Morita, Masao Fuketa, and Jun ichi Aoe. A new approach for Arabic text classification using Arabic field-association terms. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2266–2276, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Amitay:2001:TFP

[Ami01] Einat Amitay. Trends, fashions, patterns, norms, conventions … and hypertext too. Journal of the American Society for Information Science and Technology: JASIST, 52 (1):36–43, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Al-Maskari:2010:RFI

[AMS10] Azzah Al-Maskari and Mark Sanderson. A review of factors influencing user satisfaction in information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 61(5):859–868, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aula:2006:MSP

[AN06] Anne Aula and Klaus Nordhausen. Modeling successful performance in Web searching. Journal of the American Society for Information Science and Technology: JASIST, 57(12): 1678–1693, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ahmed:2009:EGC

[AN09] Farag Ahmed and Andreas Nürnberger. Evaluation of ngram conflation approaches for Arabic text retrieval. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1448–1465, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Andersen:2002:BRB

[And02] Jack Andersen. Book review: Information, knowledge, text. Journal of the American Society for Information Science and Technology: JASIST, 53(8):690–691, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Adiego:2007:LZC

[ANdlF07] Joaquín Adiego, Gonzalo Navarro, and Pablo de la Fuente. Lempel–Ziv compression of highly structured documents. Journal of the American Society for Information Science and Technology: JASIST, 58(4):461–478, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2001:CPCa

[Ano01a] Anonymous. Call for papers: Call for papers: Special topics issue of JASIST: Web retrieval and mining. Journal of the American Society for Information Science and Technology: JASIST, 52(7):603, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2001:CPCb

[Ano01b] Anonymous. Call for papers: CoLIS4: Fourth international conference on conceptions of library science and information science: Emerging frameworks and methods. Journal of the American Society for Information Science and Technology: JASIST, 52(11):986, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2001:EAI

[Ano01c] Anonymous. Editorials: Announcement: International summer school on the digital library. Journal of the American Society for Information Science and Technology: JASIST, 52(10):876, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2001:ECPc

[Ano01d] Anonymous. Editorials: Call for papers: 12th International Conference on Database and Expert Systems Applications (DEXA 2001). Journal of the American Society for Information Science and Technology: JASIST, 52(4):367–368, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2001:ECPd

[Ano01e] Anonymous. Editorials: Call for papers: Special session on Information Retrieval: International Conference on Information Technology: Coding and Computing (ITCC 2002) The Orleans, Las Vegas, Nevada, USA April 8–10, 2002 sponsored by the IEEE Computer Society. Journal of the American Society for Information Science and Technology: JASIST, 52(14):1280, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2001:ECPe

[Ano01f] Anonymous. Editorials: Call for papers: Special topic issue of JASIST: Document search interface design for large-scale collections and intelligent access. Journal of the American Society for Information Science and Technology: JASIST, 52(14):1281, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2001:ECPa

[Ano01g] Anonymous. Editorials: Call for papers: Special topic issue of JASIST on information architecture. Journal of the American Society for Information Science and Technology: JASIST, 52(2):184, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2001:ECPb

[Ano01h] Anonymous. Editorials: Call for papers: Special topic issue of JASIST: Uncertainty and logic in information access. Journal of the American Society for Information Science and Technology: JASIST, 52(3):280, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2001:EST

[Ano01i] Anonymous. Editorials: Special topic issue of JASIST: Visualization of scientific paradigms. Journal of the American Society for Information Science and Technology: JASIST, 52(6):514, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2002:ASI

[Ano02a] Anonymous. Announcement: Seventh International Summer School on the Digital Library. Journal of the American Society for Information Science and Technology: JASIST, 53 (9):779–780, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2002:CPP

[Ano02b] Anonymous. Call for papers: A perspectives issue on knowledge management in Asia. Journal of the American Society for Information Science and Technology: JASIST, 53(11): 971, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2002:CPC

[Ano02c] Anonymous. Call for papers: Call for papers: Special topic issue: Webometrics. Journal of the American Society for Information Science and Technology: JASIST, 53(9):778, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2002:CPSb

[Ano02d] Anonymous. Call for papers: Special topic issue of JASIST: Research on Information Seeking. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1189, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2002:CPSa

[Ano02e] Anonymous. Call for papers: Special topic issue of JASIST: Soft power: Informational ambiguities and asymmetries in the network age. Journal of the American Society for Information Science and Technology: JASIST, 53(3):254, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2002:EST

[Ano02f] Anonymous. Editorial: Special topic issue of JASIST: Bioinformatics. Journal of the American Society for Information Science and Technology: JASIST, 53(1):63, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2002:E

[Ano02g] Anonymous. Erratum. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1277, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2003:BR

[Ano03a] Anonymous. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 54(2): 182–184, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2003:CPP

[Ano03b] Anonymous. Call for papers: A perspectives on knowledge management in Asia. Journal of the American Society for Information Science and Technology: JASIST, 54(3):277, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2003:CPAb

[Ano03c] Anonymous. Call for papers: ACM CIKM 2003: Preliminary call for papers. Journal of the American Society for Information Science and Technology: JASIST, 54(5):469, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2003:CPAa

[Ano03d] Anonymous. Call for papers: Annual course and conference: Libraries in the digital age (Lida) 2003 Dubrovnik and Mljet, Croatia 26–30 May, 2003 inter-university centre (http://www.hr/iuc) Don Ivana Bulica 4, 20000 Dubrovnik, Croatia hotel Odisej, island Mljet, Pomena, Croatia (http://www.hotelodisej.hr) course Web site: http://www.pedos.hr/lida Course e-mail: lida@pedos.hr. Journal of the American Society for Information Science and Technology: JASIST, 54(3):278, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2003:CPC

[Ano03e] Anonymous. Call for papers: Call for papers: Special topic issue of JASIST: Multilingual information systems. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1171, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2003:CPIb

[Ano03f] Anonymous. Call for papers: Information Resources Management Association (IRMA) 15th Annual International Conference Innovations through Information Technology. Journal of the American Society for Information Science and Technology: JASIST, 54(13):1265, November 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2003:CPS

[Ano03g] Anonymous. Call for papers: Special topic issue of JASIST: Intelligence and security informatics. Journal of the American Society for Information Science and Technology: JA-SIST, 54(7):698, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2003:CPIa

[Ano03h] Anonymous. Call for papers: The 2003 International Conference on Information and Knowledge Engineering (IKE’03): June 23 26, 2003 Monte Carlo Resort, Las Vegas, Nevada, USA. Journal of the American Society for Information Science and Technology: JASIST, 54(5):470, March 2003. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2004:CPJ

[Ano04a] Anonymous. Call for papers: Joint conference on digital libraries (JCDL) 2004: Global reach and diverse impact. Journal of the American Society for Information Science and Technology: JASIST, 55(4):373–374, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2004:CPSb

[Ano04b] Anonymous. Call for papers: Second Symposium on Intelligence and Security Informatics (ISI-2004). Journal of the American Society for Information Science and Technology: JASIST, 55(11):1028–1029, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2004:CPSa

[Ano04c] Anonymous. Call for papers: Special topic issue of JASIST: Soft approaches to information retrieval and information access on the Web. Journal of the American Society for Information Science and Technology: JASIST, 55(4):372, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2005:CPS

[Ano05] Anonymous. Call for papers: Special topic issue of JASIST: Mining Web resources for enhancing information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 56(9):1011, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2006:A

[Ano06a] Anonymous. Announcement. Journal of the American Society for Information Science and Technology: JASIST, 57(8): 1133, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2006:EMH

[Ano06b] Anonymous. Erratum: S. M. Humphrey, W. J. Rogers, H. Kilicoglu, D. Demner-Fushman, and T. C. Rindflesch, ‘Word sense disambiguation by selecting the best semantic type based on Journal Descriptor Indexing: Preliminary experiment’. Journal of the American Society for Information Science and Technology 57 (1) 2006, 96–113. Journal of the American Society for Information Science and Technology: JASIST, 57(5):726, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [HRK^{+}06].

Anonymous:2008:ERD

[Ano08a] Anonymous. Erratum re: “Digital library resources as a basis for collaborative work” JASIST 51(3) 2000, 228–245.

Journal of the American Society for Information Science and Technology: JASIST, 59(9):1535, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:2008:ERP

[Ano08b] Anonymous. Erratum Re: Perspectives on Global Information Ethics, Journal of the American Society for Information Science and Technology 59(7) 2008, 1171–1183. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1698, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Bri08].

Anonymous:2008:ERS

[Ano08c] Anonymous. Erratum Re: ‘Selecting manuscripts for a highimpact journal through peer review: a citation analysis of communications that were accepted by Agewandte Chemie International Edition, or rejected but published elsewhere’. Journal of the American Society for Information Science and Technology 59(11) 2008, 1841–1852. Journal of the American Society for Information Science and Technology: JA-SIST, 59(12):2037–2038, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [BD08].

Anonymous:2010:ECS

[Ano10] Anonymous. Erratum: Correction to Shah, C. and Marchionini, G. (2010). Awareness in collaborative information seeking. Journal of the American Society for Information Science and Technology, 61(10), 1970–1986. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2377, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [SM10].

Anonymous:2011:ECT

[Ano11] Anonymous. Erratum: Correction to Thelwall, M., Buckley, K., Paltoglou, G., Cai, D., and Kappas, A., ‘Sentiment in short strength detection informal text’. Journal of the American Society for Information Science and Technology 61(12) 2010, 2544–2558. Journal of the American Society for Information Science and Technology: JASIST, 62(2):419, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Ano11].

Anonymous:2012:EKJ

[Ano12] Anonymous. Erratum: K. Jonkers and G. E. Derrick. The bibliometric bandwagon: Characteristics of bibliometric articles outside the field literature. Journal of the American Society for Information Science and Technology 63(4) 2012, 829–836. DOI: 10.1002/asi.22620. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1280, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [JD12].

Anonymous:2013:ERW

[Ano13a] Anonymous. Erratum: A random walk on an ontology: Using thesaurus structure for automatic subject indexing 64(7) 2013, 1330–1344. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1757, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [WL13].

Anonymous:2013:EGS

[Ano13b] Anonymous. Erratum: Gazni, A., Sugimoto, C. R., and Didegah, F. (2012). Mapping world scientific collaboration: Authors, institutions, and countries. Journal of the American Society for Information Science and Technology, 63(2), 323–335. DOI: 10.1002/asi.21688. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2600, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [GSD12].

Anonymous:2013:ELL

[Ano13c] Anonymous. Erratum: Ling-Ling Wu, Mu-Hsuan Huang, and Ching-Yi Chen. Citation patterns of the pre-web and web-prevalent environments: the moderating effects of domain knowledge Journal of the American Society for Information Science and Technology 63(11) 2012, 2182–2194. DOI: 10.1002/asi.22710. Journal of the American Society for Information Science and Technology: JASIST, 64(1):215, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [WHC12].

Anonymous:2013:EPF

[Ano13d] Anonymous. Erratum: Publication fees in open access publishing: Sources of funding and factors influencing choice of journal 63(1) 2012, 98–107. Journal of the American Society for Information Science and Technology: JASIST, 64(5): 1089, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [SB12a].

Anonymous:2014:EBS

[Ano14] Anonymous. Erratum: “Backstage” solidarity in Spanishand English-written medical research papers: Publication context and the acknowledgment paratext. Journal of the American Society for Information Science and Technology 60(2) 2009, 307–317. DOI: 10.1002/asi.20981. Journal of the Association for Information Science and Technology, 65 (10):2164, October 2014. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic). See [SMAB09].

Aker:2013:DHC

[APLG13] Ahmet Aker, Laura Plaza, Elena Lloret, and Robert Gaizauskas. Do humans have conceptual models about geographic objects? A user study. Journal of the American Society for Information Science and Technology: JASIST, 64(4):689–700, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Alzahrani:2012:USI

[APSA12] Salha Alzahrani, Vasile Palade, Naomie Salim, and Ajith Abraham. Using structural information and citation evidence to detect significant plagiarism cases in scientific publications. Journal of the American Society for Information Science and Technology: JASIST, 63(2):286–312, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Assefa:2013:BMS

[AR13] Shimelis G. Assefa and Abebe Rorissa. A bibliometric mapping of the structure of STEM education using co-word analysis. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2513–2536, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Albarran:2011:RMC

[ARC11] Pedro Albarrán and Javier Ruiz-Castillo. References made and citations received by scientific articles. Journal of the American Society for Information Science and Technology: JASIST, 62(1):40–49, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ardanuy:2013:BCS

[Ard13] Jordi Ardanuy. Brief communication: Sixty years of citation analysis studies in the humanities (1951–2010). Journal of the American Society for Information Science and Technology: JASIST, 64(8):1751–1755, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Argamon:2003:CPS

[Arg03] Shlomo Argamon. Call for papers: Special topic issue of JA-SIST: Computational methods for style analysis and synthesis. Journal of the American Society for Information Science and Technology: JASIST, 54(7):699, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Argamon:2006:IST

[Arg06] Shlomo Argamon. Introduction to the special topic section on the computational analysis of style. Journal of the American Society for Information Science and Technology: JA-SIST, 57(11):1503–1505, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aksnes:2011:FRL

[ARPS11] Dag W. Aksnes, Kristoffer Rorstad, Fredrik Piro, and Gunnar Sivertsen. Are female researchers less cited? A largescale study of Norwegian scientists. Journal of the American Society for Information Science and Technology: JA-SIST, 62(4):628–636, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Al-Shawakfa:2010:CSS

[ASABS^{+}10] Emad Al-Shawakfa, Amer Al-Badarneh, Safwan Shatnawi, Khaleel Al-Rabab’ah, and Basel Bani-Ismail. A comparison study of some Arabic root finding algorithms. Journal of the American Society for Information Science and Technology: JASIST, 61(5):1015–1024, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Al-Sughaiyer:2004:AMA

[ASAK04] Imad A. Al-Sughaiyer and Ibrahim A. Al-Kharashi. Arabic morphological analysis techniques: a comprehensive survey. Journal of the American Society for Information Science and Technology: JASIST, 55(3):189–213, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aris:2009:VOD

[ASQR09] Aleks Aris, Ben Shneiderman, Vahed Qazvinian, and Dragomir Radev. Visual overviews for discovering key papers and influences across research fronts. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2219–2228, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Arazy:2010:RCW

[ASR^{+}10] Ofer Arazy, Eleni Stroulia, Stan Ruecker, Cristina Arias, Carlos Fiorentino, Veselin Ganev, and Timothy Yau. Recognizing contributions in wikis: Authorship categories, algorithms, and visualizations. Journal of the American Society for Information Science and Technology: JASIST, 61(6): 1166–1179, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Al:2006:AHL

[AŞT06] Umut Al, Mustafa Şahiner, and Yaşar Tonta. Arts and humanities literature: Bibliometric characteristics of contributions by Turkish authors. Journal of the American Society for Information Science and Technology: JASIST, 57(8): 1011–1022, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aastrom:2007:CLR

[Åst07] Fredrik Åström. Changes in the LIS research front: Timesliced cocitation analyses of LIS journal articles, 1990–2004. Journal of the American Society for Information Science and Technology: JASIST, 58(7):947–957, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Artymiuk:2005:GTM

[ASW05] Peter J. Artymiuk, Ruth V. Spriggs, and Peter Willett. Graph theoretic methods for the analysis of structural relationships in biological macromolecules. Journal of the American Society for Information Science and Technology: JA-SIST, 56(5):518–528, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Alavi:2002:KIV

[AT02] Maryam Alavi and Amrit Tiwana. Knowledge integration in virtual teams: the potential role of KMS. Journal of the American Society for Information Science and Technology: JASIST, 53(12):1029–1037, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Atkinson:2001:CCP

[Atk01] Ross Atkinson. Contingency and contradiction: the place(s) of the library at the dawn of the new millennium. Journal of the American Society for Information Science and Technology: JASIST, 52(1):3–11, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Austin:2001:RML

[Aus01] Brice Austin. Research: Mooers’ law: In and out of context. Journal of the American Society for Information Science and Technology: JASIST, 52(8):607–609, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Aversa:2013:BRB

[Ave13] Elizabeth Aversa. Book review: Introduction to information science — By David Bawden and Lyn Robinson. Journal of the American Society for Information Science and Technology: JASIST, 64(5):1081–1083, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Allen:2005:MSC

[AW05] Robert B. Allen and Yejun Wu. Metrics for the scope of a collection. Journal of the American Society for Information Science and Technology: JASIST, 56(12):1243–1249, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Awazu:2006:BRB

[Awa06] Yukika Awazu. Book review: Creation, use, and deployment of digital information. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1709– 1710, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Althouse:2009:DIF

[AWBB09] Benjamin M. Althouse, Jevin D. West, Carl T. Bergstrom, and Theodore Bergstrom. Differences in impact factor across fields and over time. Journal of the American Society for Information Science and Technology: JASIST, 60(1):27–34, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Argamon:2007:STC

[AWC^{+}07] Shlomo Argamon, Casey Whitelaw, Paul Chase, Sobhan Raj Hota, Navendu Garg, and Shlomo Levitan. Stylistic text classification using functional lexical features. Journal of the American Society for Information Science and Technology: JASIST, 58(6):802–822, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Agarwal:2011:CBI

[AXP11] Naresh Kumar Agarwal, Yunjie (Calvin) Xu, and Danny C. C. Poo. A context-based investigation into source use by information seekers. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1087– 1104, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Arazy:2013:SWT

[AYN13] Ofer Arazy, Lisa Yeo, and Oded Nov. Stay on the Wikipedia task: When task-related disagreements slip into personal and procedural conflicts. Journal of the American Society for Information Science and Technology: JASIST, 64(8): 1634–1648, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Avrahami:2006:FPF

[AYSC06] Thi Truong Avrahami, Lawrence Yau, Luo Si, and Jamie Callan. The FedLemur project: Federated search in the real world. Journal of the American Society for Information Science and Technology: JASIST, 57(3):347–358, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Apostolova:2013:IRS

[AYX^{+}13] Emilia Apostolova, Daekeun You, Zhiyun Xue, Sameer Antani, Dina Demner-Fushman, and George R. Thoma. Image retrieval from scientific publications: Text and image content processing to separate multipanel figures. Journal of the American Society for Information Science and Technology: JASIST, 64(5):893–908, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Baruchson-Arbib:2002:VFL

[BAB02] Shifra Baruchson-Arbib and Jenny Bronstein. A view to the future of the library and information science profession: A Delphi study. Journal of the American Society for Information Science and Technology: JASIST, 53(5):397–408, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Baruchson-Arbib:2007:HIU

[BAB07] Shifra Baruchson-Arbib and Jenny Bronstein. Humanists as information users in the digital age: the case of Jewish studies scholars in Israel. Journal of the American Society for Information Science and Technology: JASIST, 58(14): 2269–2279, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Badurek:2003:BR

[Bad03] Christopher A. Badurek. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 54(9):905–906, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bainbridge:2004:MIR

[Bai04] David Bainbridge. Music information retrieval research and its context at the University of Waikato. Journal of the American Society for Information Science and Technology: JASIST, 55(12):1092–1099, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barr:2001:BRB

[Bar01a] Charles Jeffrey Barr. Book review: The digital enterprise: How to reshape your business for a connected world, edited by Nicholas G. Carr. Journal of the American Society for Information Science and Technology: JASIST, 52(12):1084– 1086, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barreau:2001:BR

[Bar01b] Deborah Barreau. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 52(1):80, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2005:IIa

[Bar05a] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1129, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2005:IIb

[Bar05b] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 56(12):1241–1242, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2005:IIc

[Bar05c] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1361–1362, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2005:IId

[Bar05d] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 56(14):1467–1468, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barjak:2006:RII

[Bar06a] Franz Barjak. The role of the Internet in informal scholarly communication. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1350–1367, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIg

[Bar06b] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(7):847–850, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIh

[Bar06c] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(8):993–996, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIi

[Bar06d] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1135–1137, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIj

[Bar06e] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1281–1283, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIk

[Bar06f] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1415–1417, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIl

[Bar06g] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1563–1565, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIm

[Bar06h] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1713–1714, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIn

[Bar06i] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1859–1861, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIa

[Bar06j] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(1):1–3, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIb

[Bar06k] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(2):149–151, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIc

[Bar06l] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(3):297–298, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IId

[Bar06m] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(4):447, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIe

[Bar06n] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 57(5):587, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2006:IIf

[Bar06o] Carol L. Barry. Introduction: In this issue. Journal of the American Society for Information Science and Technology:

JASIST, 57(6):727, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIa

[Bar07a] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(1):1–2, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIb

[Bar07b] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(2):155–156, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIc

[Bar07c] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(3):307–308, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IId

[Bar07d] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(4):459–460, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIe

[Bar07e] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(5):611–612, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIf

[Bar07f] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(6):763–765, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIg

[Bar07g] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology:

JASIST, 58(7):913–914, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIh

[Bar07h] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1069–1070, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIi

[Bar07i] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1225–1226, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIj

[Bar07j] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1389–1391, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIk

[Bar07k] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1557–1558, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIl

[Bar07l] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1707–1708, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIm

[Bar07m] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 58(13):1911–1914, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2007:IIn ogy: JASIST, 58(14):2163–2166, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barreau:2008:PBF

[Bar08a] Deborah Barreau. The persistence of behavior and form in the organization of personal information. Journal of the American Society for Information Science and Technology: JASIST, 59(2):307–317, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIa

[Bar08b] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(1):1–2, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIb

[Bar08c] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(2):163–165, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIc

[Bar08d] Carol L. Barry. In this issue: In This Issue. Journal of the American Society for Information Science and Technology: JASIST, 59(3):337–338, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IId

[Bar08e] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(4):507–509, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIe

[Bar08f] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(5):681–683, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIf

[Bar08g] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology:

JASIST, 59(6):857–858, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIg

[Bar08h] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(7):1025, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIh

[Bar08i] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1193–1194, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIi

[Bar08j] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1363–1365, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIj

[Bar08k] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1537–1538, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIk

[Bar08l] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1699–1701, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIl

[Bar08m] Carol L. Barry. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 59(12):1867–1869, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIm ogy: JASIST, 59(13):2039–2041, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barry:2008:IIn

[Bar08o] Carol L. Barry. In this issue: In This Issue. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2195–2197, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bartolucci:2012:LEP

[Bar12] Francesco Bartolucci. Letter to the Editor: On a possible decomposition of the h-index. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2126–2127, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bates:2006:FFI

[Bat06] Marcia J. Bates. Fundamental forms of information. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1033–1045, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bates:2008:BCH

[Bat08a] Marcia J. Bates. Brief communications: Hjørland’s critique of Bates’ work on defining information. Journal of the American Society for Information Science and Technology: JA-SIST, 59(5):842–844, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bellaachia:2008:PNE

[BAT08b] Abdelghani Bellaachia and Ghita Amor-Tijani. Proper nouns in English–Arabic cross language information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 59(12):1925–1932, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bates:2011:BHM

[Bat11] Marcia J. Bates. Birger Hjørland’s Manichean misconstruction of Marcia Bates’ work. Journal of the American Society for Information Science and Technology: JASIST, 62(10): 2038–2044, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bawden:2010:BRB

[Baw10] David Bawden. Book review: Perspectives on Knowledge Management. Journal of the American Society for Information Science and Technology: JASIST, 61(3):629, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyack:2003:IAE

[BB03] Kevin W. Boyack and Katy Börner. Indicator-assisted evaluation and funding of research: Visualizing the influence of grants on the number and citation counts of research papers. Journal of the American Society for Information Science and Technology: JASIST, 54(5):447–461, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bertoli-Barsotti:2013:LEI

[BB13] Lucio Bertoli-Barsotti. Letter to the Editor: Improving a decomposition of the h-index. Journal of the American Society for Information Science and Technology: JASIST, 64(7): 1522, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bordons:2004:TTR

[BBB04] María Bordons, Carmen Bravo, and Santos Barrigón. Timetracking of the research profile of a drug using bibliometric tools. Journal of the American Society for Information Science and Technology: JASIST, 55(5):445–461, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Britt:2008:BCD

[BBB^{+}08] Barry L. Britt, Michael W. Berry, Murray Browne, Mary Ann Merrell, and James Kolpack. Brief communications: Document classification techniques for automated technology readiness level analysis. Journal of the American Society for Information Science and Technology: JASIST, 59(4):675–680, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burnett:2001:RSW

[BBC01] Gary Burnett, Michele Besant, and Elfreda A. Chatman. Research: Small worlds: Normative behavior in virtual communities and feminist bookselling. Journal of the American Society for Information Science and Technology: JASIST, 52(7):536–547, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bache:2009:BCM

[BBC09] Richard Bache, Mark Baillie, and Fabio Crestani. Brief communications: Measuring the likelihood property of scoring functions in general retrieval models. Journal of the American Society for Information Science and Technology: JA-SIST, 60(6):1294–1297, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barsky:2012:ITP

[BBI12] Elena Barsky and Judit Bar-Ilan. The impact of task phrasing on the choice of search keywords and on the search process and success. Journal of the American Society for Information Science and Technology: JASIST, 63(10):1987–2005, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bhavnani:2006:SHD

[BBJ^{+}06] Suresh K. Bhavnani, Christopher K. Bichakjian, Timothy M. Johnson, Roderick J. Little, Frederick A. Peck, Jennifer L. Schwartz, and Victor J. Strecher. Strategy hubs: Domain portals to help find comprehensive information. Journal of the American Society for Information Science and Technology: JASIST, 57(1):4–24, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Billhardt:2002:CVM

[BBM02] Holger Billhardt, Daniel Borrajo, and Victor Maojo. A context vector model for information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 53(3):236–249, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bergman:2003:USA

[BBMN03] Ofer Bergman, Ruth Beyth-Marom, and Rafi Nachmias. The user-subjective approach to personal information management systems. Journal of the American Society for Information Science and Technology: JASIST, 54(9):872–878, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bergman:2008:USA

[BBMN08] Ofer Bergman, Ruth Beyth-Marom, and Rafi Nachmias. The user-subjective approach to personal information management systems design: Evidence and implementations. Journal of the American Society for Information Science and Technology: JASIST, 59(2):235–246, January 15, 2008. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barrantes:2012:CFZ

[BBRdMA12] Bárbara S. Lancho Barrantes, Vicente P. Guerrero Bote, Zaida Chinchilla Rodríguez, and Félix de Moya Anegón. Citation flows in the zones of influence of scientific collaborations. Journal of the American Society for Information Science and Technology: JASIST, 63(3):481–489, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bean:2005:EFO

[BC05] Carol A. Bean and Milton Corn. Extramural funding opportunities in bioinformatics from the National Library of Medicine. Journal of the American Society for Information Science and Technology: JASIST, 56(5):551–556, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bose:2009:MEG

[BC09] Indranil Bose and Xi Chen. A method for extension of generative topographic mapping for fuzzy clustering. Journal of the American Society for Information Science and Technology: JASIST, 60(2):363–371, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boughanem:2002:UGA

[BCT02] M. Boughanem, C. Chrisment, and L. Tamine. On using genetic algorithms for multimodal relevance optimization in information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 53(11):934– 942, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Blocks:2006:RMU

[BCT06] Dorothee Blocks, Daniel Cunliffe, and Douglas Tudhope. A reference model for user-system interaction in thesaurusbased searching. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1655–1665, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Braun:2005:GIE

[BD05] Tibor Braun and Ildikó Dióspatonyi. Gatekeeping indicators exemplified by the main players in the international gatekeeping orchestration of analytical chemistry journals. Journal of the American Society for Information Science and Technology: JASIST, 56(8):854–860, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2007:BCW

[BD07a] Lutz Bornmann and Hans-Dieter Daniel. Brief communication: What do we know about the h index? Journal of the American Society for Information Science and Technology: JASIST, 58(9):1381–1385, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2007:MPS

[BD07b] Lutz Bornmann and Hans-Dieter Daniel. Multiple publication on a single research study: Does it pay? The influence of number of research articles on total citation counts in biomedicine. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1100–1107, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2008:SMH

[BD08]  Lutz Bornmann and Hans-Dieter Daniel.  Selecting

manuscripts for a high-impact journal through peer review: a citation analysis of communications that were accepted by Angewandte Chemie International Edition, or rejected but published elsewhere. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1841– 1852, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Ano08c].

Bornmann:2009:UCD

[BD09] Lutz Bornmann and Hans-Dieter Daniel. Universality of citation distributions —- a validation of Radicchi et al.’s relative indicator c_{f} = c/c_{0} at the micro level using data from chemistry. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1664–1670, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Balatsoukas:2010:EGS

[BD10] Panos Balatsoukas and Peter Demian. Effects of granularity of search results on the relevance judgment behavior of engineers: Building systems for retrieval and understanding of context. Journal of the American Society for Information Science and Technology: JASIST, 61(3):453–467, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brown:2013:KSK

[BDBA13] Susan A. Brown, Alan R. Dennis, Diana Burley, and Priscilla Arling. Knowledge sharing and knowledge management system avoidance: the role of knowledge type and the social network in bypassing an organizational knowledge management system. Journal of the American Society for Information Science and Technology: JASIST, 64(10):2013–2023, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Baloh:2012:COI

[BDH12] Peter Baloh, Kevin C. Desouza, and Ray Hackney. Contextualizing organizational interventions of knowledge management systems: a design science perspective. Journal of the American Society for Information Science and Technology: JASIST, 63(5):948–966, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2013:DUR

[BdMAM13] Lutz Bornmann, Felix de Moya Anegón, and Rüdiger Mutz. Do universities or research institutions with a specific subject profile have an advantage or a disadvantage in institutional rankings? A latent class analysis with data from the SCImago Ranking. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2310– 2316, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Berendsen:2013:AEP

[BdRB^{+}13] Richard Berendsen, Maarten de Rijke, Krisztian Balog, Toine Bogers, and Antal van den Bosch. On the assessment of expertise profiles. Journal of the American Society for Information Science and Technology: JASIST, 64(10): 2024–2044, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Booker:2012:FAA

[BDS12] Lorne D. Booker, Brian Detlor, and Alexander Serenko. Factors affecting the adoption of online library resources by business students. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2503–2520, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Beale:2006:PSU

[Bea06] Russell Beale. Perspectives on search user interfaces: Best practices and future visions: Improving Internet interaction: From theory to practice. Journal of the American Society for Information Science and Technology: JASIST, 57(6):829– 833, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bounhas:2010:TCS

[BEES10] Ibrahim Bounhas, Bilel Elayeb, Fabrice Evrard, and Yahya Slimani. Toward a computer study of the reliability of Arabic stories. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1686–1705, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Beghtol:2005:EDM

[Beg05] Clare Beghtol. Ethical decision-making for knowledge representation and organization systems for global use. Journal of the American Society for Information Science and Technology: JASIST, 56(9):903–912, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bodoff:2001:RUM

[BEK^{+}01] David Bodoff, Daniel Enache, Ajit Kambil, Gary Simon, and Alex Yukhimets. Research: A unified maximum likelihood approach to document retrieval. Journal of the American Society for Information Science and Technology: JASIST, 52(10):785–796, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Benoit:2001:BRB

[Ben01a] G. Benoît. Book review: Snap to grid: a user’s guide to digital arts, media, and cultures, by Peter Lunenfeld. Journal of the American Society for Information Science and Technology: JASIST, 52(6):509–510, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bensman:2001:UGL

[Ben01b] Stephen J. Bensman. Urquhart’s and Garfield’s Laws: The British controversy over their validity. Journal of the American Society for Information Science and Technology: JA-SIST, 52(9):714–724, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Benoit:2002:DDN

[Ben02] G. Benoît. Data discretization for novel relationship discovery in information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 53 (9):736–746, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Benoit:2004:PBR

[Ben04a] Gerald Benoît. Properties-based retrieval and user decision states: User control and behavior modeling. Journal of the American Society for Information Science and Technology: JASIST, 55(6):488–497, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bensman:2004:LEP

[Ben04b] Stephen J. Bensman. Letters to the Editor: Pearson’s r and author cocitation analysis: a commentary on the controversy. Journal of the American Society for Information Science and Technology: JASIST, 55(10):935, August 2004.

CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bensman:2005:UPT

[Ben05] Stephen J. Bensman. Urquhart and probability: the transition from librarianship to library and information science. Journal of the American Society for Information Science and Technology: JASIST, 56(2):189–214, January 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Benoit:2006:BRB

[Ben06] Gerald Benoît. Book review: Link analysis: an information science approach. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1855–1858, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bensman:2007:BCI

[Ben07] Stephen J. Bensman. Brief communications: The impact factor, total citations, and better citation mouse traps: a commentary. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1904–1908, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bensman:2008:DDI

[Ben08] Stephen J. Bensman. Distributional differences of the impact factor in the sciences versus the social sciences: an analysis of the probabilistic structure of the 2005 journal citation reports. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1366–1382, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bensman:2011:BRB

[Ben11] Stephen J. Bensman. Book review: The evaluation of research by scientometric indicators. Journal of the American Society for Information Science and Technology: JASIST, 62(1):208–210, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barnett:2008:IIS

[BF08] George A. Barnett and Edward L. Fink. Impact of the Internet and scholar age distribution on academic citation age. Journal of the American Society for Information Science and Technology: JASIST, 59(4):526–534, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bergman:2013:FVT

[BGBIBM13] Ofer Bergman, Noa Gradovitch, Judit Bar-Ilan, and Ruth Beyth-Marom. Folder versus tag preference in personal information management. Journal of the American Society for Information Science and Technology: JASIST, 64(10): 1995–2012, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Birnholtz:2013:CCC

[BGY^{+}13] Jeremy Birnholtz, Shion Guha, Y. Connie Yuan, Geri Gay, and Caren Heller. Cross-campus collaboration: a scientometric and network case study of publication activity across two campuses of a single institution. Journal of the American Society for Information Science and Technology: JA-SIST, 64(1):162–172, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bystrom:2005:CFT

[BH05] Katriina Byström and Preben Hansen. Conceptual framework for tasks in information studies. Journal of the American Society for Information Science and Technology: JA-SIST, 56(10):1050–1061, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Benoit:2011:RDO

[BH11] G. Benoît and Lisa Hussey. Repurposing digital objects: Case studies across the publishing industry. Journal of the American Society for Information Science and Technology: JASIST, 62(2):363–374, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bhavnani:2005:WID

[Bha05] Suresh K. Bhavnani. Why is it difficult to find comprehensive information? Implications of information scatter for search and design. Journal of the American Society for Information Science and Technology: JASIST, 56(9):989–1003, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brody:2006:EWU

[BHC06] Tim Brody, Stevan Harnad, and Leslie Carr. Earlier Web usage statistics as predictors of later citation impact. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1060–1072, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bernius:2013:EET

[BHDK13] Steffen Bernius, Matthias Hanauske, Berndt Dugall, and Wolfgang König. Exploring the effects of a transition to open access: Insights from a simulation study. Journal of the American Society for Information Science and Technology: JASIST, 64(4):701–726, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2002:MMS

[BI02] Judit Bar-Ilan. Methods for measuring search engine performance over time. Journal of the American Society for Information Science and Technology: JASIST, 53(4):308– 319, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bjorneborn:2004:TBF

[BI04] Lennart Björneborn and Peter Ingwersen. Toward a basic framework for webometrics. Journal of the American Society for Information Science and Technology: JASIST, 55(14): 1216–1227, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2006:WLS

[BI06] Judit Bar-Ilan. Web links and search engine ranking: the case of Google and the query “jew”. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1581–1589, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2013:BRB

[BI13] Judit Bar-Ilan. Book review: Social Information Research — By Gunilla Widén and Kim Holmberg. Journal of the American Society for Information Science and Technology: JA-SIST, 64(11):2400–2402, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bias:2010:BRBa

[Bia10a] Randolph G. Bias. Book review: Research Methods for Human–Computer Interaction. Journal of the American Society for Information Science and Technology: JASIST, 61 (1):204–205, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bias:2010:BRBb

[Bia10b] Randolph G. Bias. Book reviews: The Case for Books: Past, Present, and Future; Reading and Writing the Electronic Book. Synthesis Lectures on Information Concepts, Retrieval, and Services. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2588– 2591, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2012:MNO

[BIA12] Judit Bar-Ilan and Rina Azoulay. Map of nonprofit organization websites in Israel. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1142– 1167, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2007:CAW

[BIB07] Judit Bar-Ilan and Yifat Belous. Children as architects of Web directories: an exploratory study. Journal of the American Society for Information Science and Technology: JA-SIST, 58(6):895–907, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2009:PBS

[BIKLY09] Judit Bar-Ilan, Kevin Keenoy, Mark Levene, and Eti Yaari. Presentation bias is significant in determining user preference for search results — a user study. Journal of the American Society for Information Science and Technology: JA-SIST, 60(1):135–149, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2007:URS

[BIKYL07] Judit Bar-Ilan, Kevin Keenoy, Eti Yaari, and Mark Levene. User rankings of search engine results. Journal of the American Society for Information Science and Technology: JA-SIST, 58(9):1254–1266, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bilal:2001:CUY

[Bil01] Dania Bilal. Children’s use of the Yahooligans! Web search engine: II. Cognitive and physical behaviors on research tasks. Journal of the American Society for Information Science and Technology: JASIST, 52(2):118–136, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bilal:2002:CUY

[Bil02] Dania Bilal. Children’s use of the Yahooligans! Web search engine. III. Cognitive and physical behaviors on fully selfgenerated search tasks. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1170– 1183, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bilal:2012:RRJ

[Bil12] Dania Bilal. Ranking, relevance judgment, and precision of information retrieval on children’s queries: Evaluation of Google, Yahoo!, Bing, Yahoo! Kids, and Ask Kids. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1879–1896, September 2012. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2004:ECD

[BIP04] Judit Bar-Ilan and Bluma C. Peritz. Evolution, continuity, and disappearance of documents on a specific topic on the Web: a longitudinal study of “informetrics”. Journal of the American Society for Information Science and Technology: JASIST, 55(11):980–990, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brener:2005:CMR

[BIP05] Nathan E. Brener, S. S. Iyengar, and O. S. Pianykh. A conclusive methodology for rating OCR performance. Journal of the American Society for Information Science and Technology: JASIST, 56(12):1274–1287, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2009:MME

[BIP09] Judit Bar-Ilan and Bluma C. Peritz. A method for measuring the evolution of a topic on the Web: the case of “informetrics”. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1730–1740, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Birnholtz:2006:WDI

[Bir06] Jeremy P. Birnholtz. What does it mean to be an author? The intersection of credit, contribution, and collaboration in science. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1758–1770, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Birnholtz:2007:WDR

[Bir07] Jeremy P. Birnholtz. When do researchers collaborate? Toward a model of collaboration propensity. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2226–2239, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bishop:2011:LBQ

[Bis11] Bradley Wade Bishop. Location-based questions and local knowledge. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1594–1603, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bar-Ilan:2010:EBI

[BIZGMS10] Judit Bar-Ilan, Maayan Zhitomirsky-Geffet, Yitzchak Miller, and Snunith Shoham. The effects of background information and social interaction on image tagging. Journal of the American Society for Information Science and Technology: JASIST, 61(5):940–951, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bondarenko:2009:CVC

[BJ09] Olha Bondarenko and Ruud Janssen. Connecting visual cues to semantic judgments in the context of the office environment. Journal of the American Society for Information Science and Technology: JASIST, 60(5):933–952, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Beitzel:2004:FER

[BJC^{+}04] Steven M. Beitzel, Eric C. Jensen, Abdur Chowdhury, David Grossman, Ophir Frieder, and Nazli Goharian. Fusion of effective retrieval strategies in the same information retrieval system. Journal of the American Society for Information Science and Technology: JASIST, 55(10):859–868, August 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Beitzel:2007:TAV

[BJC^{+}07] Steven M. Beitzel, Eric C. Jensen, Abdur Chowdhury, Ophir Frieder, and David Grossman. Temporal analysis of a very large topically categorized Web query log. Journal of the American Society for Information Science and Technology: JASIST, 58(2):166–178, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bondarenko:2010:RDP

[BJD10] Olha Bondarenko, Ruud Janssen, and Samuël Driessen. Requirements for the design of a personal documentmanagement system. Journal of the American Society for Information Science and Technology: JASIST, 61(3):468– 482, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bjork:2012:HMO

[Bjö12] Bo-Christer Björk. The hybrid model for open access publication of scholarly articles: a failed experiment? Journal of the American Society for Information Science and Technology: JASIST, 63(8):1496–1504, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyack:2010:CCA

[BK10] Kevin W. Boyack and Richard Klavans. Co-citation analysis, bibliographic coupling, and direct citation: Which citation approach represents the research front most accurately? Journal of the American Society for Information Science and Technology: JASIST, 61(12):2389–2404, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bookstein:2003:AMC

[BKRN03] Abraham Bookstein, Vladimir Kulyukin, Timo Raita, and John Nicholson. Adapting measures of clumping strength to assess term-term similarity. Journal of the American Society for Information Science and Technology: JASIST, 54 (7):611–620, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bensman:2009:DIJ

[BL09] Stephen J. Bensman and Loet Leydesdorff. Definition and identification of journals as bibliographic and subject entities: Librarianship versus ISI Journal Citation Reports methods and their effect on citation measures. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1097–1117, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2011:WCP

[BL11] Lutz Bornmann and Loet Leydesdorff. Which cities produce more excellent papers than can be expected? A new mapping approach, using Google Maps, based on statistical significance testing. Journal of the American Society for Information Science and Technology: JASIST, 62(10):1954–1962, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bae:2012:SAT

[BL12] Younggue Bae and Hongchul Lee. Sentiment analysis of Twitter audiences: Measuring the positive or negative influence of popular twitterers. Journal of the American Society for Information Science and Technology: JASIST, 63 (12):2521–2535, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2013:LES

[BL13] Lutz Bornmann and Loet Leydesdorff. Letter to the Editor: Statistical tests and research assessments: a comment on Schneider (2012). Journal of the American Society for Information Science and Technology: JASIST, 64(6):1306– 1308, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Blair:2002:KMH

[Bla02] David C. Blair. Knowledge management: Hype, hope, or help? Journal of the American Society for Information Science and Technology: JASIST, 53(12):1019–1028, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Blazic:2007:AUI

[BLA07] Borka Jerman Blažič, Effie Lai-Chong Law, and Tanja Arh. An assessment of the usability of an Internet-based education system in a cross-cultural environment: the case of the Interreg crossborder program in Central Europe. Journal of the American Society for Information Science and Technology: JASIST, 58(1):66–75, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Blanchette:2010:BRB

[Bla10] Jean-François Blanchette. Book reviews: Total Recall: How the E-memory Revolution Will Change Everything; DELETE: The Virtue of Forgetting in the Digital Age. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1727–1731, August 2010. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Blanchette:2011:MHB

[Bla11] Jean-François Blanchette. A material history of bits. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1042–1057, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Black:2013:BRB

[Bla13] Alistair Black. Book review: Paper Machines: About Cards & Catalogs, 1548–1929 — By Markus Krajewski. Journal of the American Society for Information Science and Technology: JASIST, 64(2):431–432, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bias:2010:ESV

[BLH^{+}10] Randolph G. Bias, Kevin Larson, Sheng-Cheng Huang, Paul R. Aumer-Ryan, and Chris Montesclaros. An exploratory study of visual and psychological correlates of preference for onscreen subpixel-rendered text. Journal of the American Society for Information Science and Technology: JASIST, 61(4):745–757, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Blanke:2012:FTE

[BLH12] Tobias Blanke, Mounia Lalmas, and Theo Huibers. A framework for the theoretical evaluation of XML retrieval. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2463–2473, December 2012. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barjak:2007:WFE

[BLT07] Franz Barjak, Xuemei Li, and Mike Thelwall. Which factors explain the Web impact of scientists’ personal homepages? Journal of the American Society for Information Science and Technology: JASIST, 58(2):200–211, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bouhnik:2006:IDL

[BM06] Dan Bouhnik and Tali Marcus. Interaction in distancelearning courses. Journal of the American Society for Information Science and Technology: JASIST, 57(3):299–305, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bouyssou:2011:RSD

[BM11] Denis Bouyssou and Thierry Marchant. Ranking scientists and departments in a consistent manner. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1761–1769, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2012:AKP

[BM12] Lutz Bornmann and Werner Marx. The Anna Karenina principle: a way of thinking about success in science. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2037–2051, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2008:TBI

[BMD08] Lutz Bornmann, Rüdiger Mutz, and Hans-Dieter Daniel. Are there better indices for evaluation purposes than the h index? A comparison of nine different variants of the h index using data from biomedicine. Journal of the American Society for Information Science and Technology: JASIST, 59(5):830–837, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2009:BCD

[BMD09] Lutz Bornmann, Rüdiger Mutz, and Hans-Dieter Daniel. Brief communications: Do we need the h index and its variants in addition to standard bibliometric measures? Journal of the American Society for Information Science and Technology: JASIST, 60(6):1286–1289, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2013:MSR

[BMD13] Lutz Bornmann, Rüdiger Mutz, and Hans-Dieter Daniel. Multilevel-statistical reformulation of citation-based university rankings: The Leiden ranking 2011/2012. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1649–1658, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bailon-Moreno:2006:SNS

[BMJARB06] Rafael Bailón-Moreno, Encarnación Jurado-Alameda, and Rosario Ruiz-Baños. The scientific network of surfactants: Structural analysis. Journal of the American Society for Information Science and Technology: JASIST, 57(7):949– 960, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Belefant-Miller:2003:BCP

[BMK03] Helen Belefant-Miller and Donald W. King. Brief communication: A profile of faculty reading and information-use behaviors on the cusp of the electronic age. Journal of the American Society for Information Science and Technology: JASIST, 54(2):179–181, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brajnik:2002:SHU

[BMTV02] Giorgio Brajnik, Stefano Mizzaro, Carlo Tasso, and Fabio Venuti. Strategic help in user interfaces for information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 53(5):343–358, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barzilai-Nahon:2008:TTN

[BN08] Karine Barzilai-Nahon. Toward a theory of network gatekeeping: a framework for exploring information control. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1493–1512, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Batlle:2004:RAA

[BNUA04] Eloi Batlle, Helmut Neuschmied, Peter Uray, and Gerd Ackermann. Recognition and analysis of audio for copyright protection: The RAA project. Journal of the American Society for Information Science and Technology: JASIST, 55(12): 1084–1091, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bodoff:2006:RBR

[Bod06] David Bodoff. Relevance for browsing, relevance for searching. Journal of the American Society for Information Science and Technology: JASIST, 57(1):69–86, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bodoff:2009:ETC

[Bod09] David Bodoff. Emergence of terminological conventions as a searcher–indexer coordination game. Journal of the American Society for Information Science and Technology: JA-SIST, 60(12):2509–2529, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bote:2013:QBI

[BOGdMA13] Vicente P. Guerrero Bote, Carlos Olmeda-Gómez, and Félix de Moya-Anegón. Quantifying the benefits of international scientific collaboration. Journal of the American Society for Information Science and Technology: JASIST, 64(2):392– 404, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bolstad:2007:BRB

[Bol07] Judy P. Bolstad. Book review: Design and usability of digital libraries: Case studies in the Asia Pacific. Journal of the American Society for Information Science and Technology: JASIST, 58(1):152–153, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bolstad:2008:BRB

[Bol08] Judy P. Bolstad. Book review: Personal Information Management. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2192–2193, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bookstein:2001:IAS

[Boo01] A. Bookstein. Implications of ambiguity for scientometric measurement. Journal of the American Society for Information Science and Technology: JASIST, 52(1):74–79, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bookstein:2013:M

[Boo13] Abraham Bookstein. In memoriam. Journal of the American Society for Information Science and Technology: JASIST, 64(5):1086–1088, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Borlund:2003:CRI

[Bor03] Pia Borlund. The concept of relevance in IR. Journal of the American Society for Information Science and Technology: JASIST, 54(10):913–925, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Borlund:2009:BRB

[Bor09] Pia Borlund. Book review: Interactive Information Retrieval in Digital Environments. Journal of the American Society for Information Science and Technology: JASIST, 60 (9):1944–1945, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2011:LED

[Bor11] Lutz Bornmann. Letters to the Editor: Do we need the Eindex in addition to the h-index and its variants? Journal of the American Society for Information Science and Technology: JASIST, 62(7):1433–1434, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Borgman:2012:AIS

[Bor12] Christine L. Borgman. Advances in information science: The conundrum of sharing research data. Journal of the American Society for Information Science and Technology: JA-SIST, 63(6):1059–1078, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2013:AIS

[Bor13a] Lutz Bornmann. Advances in information science: What is societal impact of research and how can it be assessed? a literature survey. Journal of the American Society for Information Science and Technology: JASIST, 64(2):217–233, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2013:HAP

[Bor13b] Lutz Bornmann. How to analyze percentile citation impact data meaningfully in bibliometrics: the statistical analysis of distributions, percentile rank classes, and top-cited papers. Journal of the American Society for Information Science and Technology: JASIST, 64(3):587–595, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2013:LEP

[Bor13c] Lutz Bornmann. Letter to the Editor: The problem of percentile rank scores used with small reference sets. Journal of the American Society for Information Science and Technology: JASIST, 64(3):650, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bowler:2010:SRC

[Bow10] Leanne Bowler. The self-regulation of curiosity and interest during the information search process of adolescent students. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1332–1344, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EId

[Boy01a] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 52(5):369–370, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIf

[Boy01b] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 52(7):515–516, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIg

[Boy01c] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 52(8):605–606, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIh

[Boy01d] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 52(9):699–700, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIk

[Boy01e] Bert R. Boyce. Editorial: In this Issue. Journal of the American Society for Information Science and Technology: JA-SIST, 52(12):987–988, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIl

[Boy01f] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 52(13):1087–1088, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIa

[Boy01g] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 52(2):85–86, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIb

[Boy01h] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 52(3):185–186, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIc

[Boy01i] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 52(4):281–282, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIe

[Boy01j] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 52(6):443–444, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIi

[Boy01k] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 52(10):783, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIj

[Boy01l] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 52(11):877, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2001:EIm

[Boy01m] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology:

JASIST, 52(14):1195–1196, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIa

[Boy02a] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 53(1):1–2, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIb

[Boy02b] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 53(3):197, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIc

[Boy02c] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 53(4):255–256, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EId

[Boy02d] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 53(5):325–326, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIe

[Boy02e] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 53(7):529–530, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIg

[Boy02f] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 53(9):693–694, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIh

[Boy02g] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology:

JASIST, 53(10):781, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIk

[Boy02h] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 53(13):1083–1084, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIl

[Boy02i] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 53(14):1191, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIf

[Boy02j] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 53(8):615–616, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIi

[Boy02k] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 53(11):877–878, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2002:EIj

[Boy02l] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 53(12):973, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:IIa

[Boy03a] Bert Boyce. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 54(1):1–2, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIa

[Boy03b] Bert R. Boyce. Editorial: In this Issue. Journal of the American Society for Information Science and Technology:

JASIST, 54(2):95–96, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIb

[Boy03c] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 54(3):191–192, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIc

[Boy03d] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 54(5):361, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EId

[Boy03e] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 54(6):471–472, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIe

[Boy03f] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 54(8):704–705, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIg

[Boy03g] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 54(10):911–912, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIh

[Boy03h] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 54(11):987–988, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIi

[Boy03i] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 54(12):1079–1080, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIj

[Boy03j] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 54(13):1173–1174, November 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIk

[Boy03k] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 54(14):1267–1268, December 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:EIf

[Boy03l] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 54(9):808–809, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2003:IIb

[Boy03m] Bert R. Boyce. In this issue: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 54(7):593, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2004:EIa

[Boy04a] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 55(1):1–2, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2004:EIb

[Boy04b] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 55(2):95–96, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2004:EIc

[Boy04c] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 55(3):187–188, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2004:EId

[Boy04d] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 55(4):281–282, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See comments [Mac04b].

Boyce:2004:EIe

[Boy04e] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 55(5):375–376, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2004:EIf

[Boy04f] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 55(6):469–470, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2004:EIg

[Boy04g] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 55(7):563–564, May 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2004:EIh

[Boy04h] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 55(10):845, August 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2004:EIi

[Boy04i] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 55(11):937–938, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2004:EIj

[Boy04j] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 55(13):1117–1118, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2005:EIa

[Boy05a] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 56(1):1–2, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2005:EIb

[Boy05b] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 56(2):109–110, January 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2005:EIc

[Boy05c] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science and Technology: JA-SIST, 56(4):325–326, February 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2005:IIa

[Boy05d] Bert R. Boyce. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 56(6):557–558, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2005:IIb

[Boy05e] Bert R. Boyce. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 56(7):661–663, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2005:IIc

[Boy05f] Bert R. Boyce. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 56(8):777–778, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2005:IId

[Boy05g] Bert R. Boyce. Introduction: In this issue. Journal of the American Society for Information Science and Technology:

JASIST, 56(9):891–892, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyce:2005:IIe

[Boy05h] Bert R. Boyce. Introduction: In this issue. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1013–1014, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Blake:2006:CISa

[BP06a] Catherine Blake and Wanda Pratt. Collaborative information synthesis I: a model of information behaviors of scientists in medicine and public health. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1740–1749, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Blake:2006:CISb

[BP06b] Catherine Blake and Wanda Pratt. Collaborative information synthesis II: Recommendations for information systems to support synthesis activities. Journal of the American Society for Information Science and Technology: JASIST, 57 (14):1888–1895, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bhavnani:2010:SMR

[BP10a] Suresh K. Bhavnani and Frederick A. Peck. Scatter matters: Regularities and implications for the scatter of healthcare information on the Web. Journal of the American Society for Information Science and Technology: JASIST, 61(4):659– 676, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bordogna:2010:FCB

[BP10b] Gloria Bordogna and Marco Pagani. A flexible content-based image retrieval model and a customizable system for the retrieval of shapes. Journal of the American Society for Information Science and Technology: JASIST, 61(5):907–926, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Buchanan:2007:DMO

[BPJR07] Tom Buchanan, Carina Paine, Adam N. Joinson, and Ulf-Dietrich Reips. Development of measures of online privacy concern and protection for use on the Internet. Journal of the American Society for Information Science and Technology: JASIST, 58(2):157–165, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bhandar:2007:TUR

[BPT07] Mamata Bhandar, Shan-Ling Pan, and Bernard C. Y. Tan. Towards understanding the roles of social capital in knowledge integration: a case study of a collaborative information systems project. Journal of the American Society for Information Science and Technology: JASIST, 58(2):263–274, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bookstein:2001:RDT

[BR01] Abraham Bookstein and T. Raita. Research: Discovering term occurrence structure in text. Journal of the American Society for Information Science and Technology: JASIST, 52(6):476–486, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bodoff:2004:NUP

[BR04] David Bodoff and Stephen Robertson. A new unified probabilistic model. Journal of the American Society for Information Science and Technology: JASIST, 55(6):471–487, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Buente:2008:TII

[BR08] Wayne Buente and Alice Robbin. Trends in Internet information behavior, 2000–2004. Journal of the American Society for Information Science and Technology: JASIST, 59 (11):1743–1760, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bashir:2011:RBQ

[BR11a] Shariq Bashir and Andreas Rauber. On the relationship between query characteristics and IR functions retrieval bias. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1515–1532, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Becker:2011:DCD

[BR11b] Christoph Becker and Andreas Rauber. Decision criteria in digital preservation: What to measure and how. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1009–1028, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Balatsoukas:2012:ETA

[BR12a] Panos Balatsoukas and Ian Ruthven. An eye-tracking approach to the analysis of relevance judgments on the Web: the case of Google search engine. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1728–1746, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bodoff:2012:UMR

[BR12b] David Bodoff and Daphne Raban. User models as revealed in web-based research services. Journal of the American Society for Information Science and Technology: JASIST, 63(3):584–599, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bawden:2013:ONS

[BR13] David Bawden and Lyn Robinson. Opinion: No such thing as society? On the individuality of information behavior. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2587–2590, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Braman:2009:BRB

[Bra09] Sandra Braman. Book review: Governing Global Electronic Networks: International Perspectives on Policy and Power. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1720–1721, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Braman:2011:BRB

[Bra11] Sandra Braman. Book review: Information Policies and Strategies. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1213–1214, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Breitzman:2005:AIT

[Bre05] Anthony Breitzman. Automated identification of technologically similar organizations. Journal of the American Society for Information Science and Technology: JASIST, 56(10): 1015–1023, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Britz:2008:PGI

[Bri08] Johannes J. Britz. Perspectives on global information ethics: Making the global information society good: a social justice perspective on the ethical dimensions of the global information society. Journal of the American Society for Information Science and Technology: JASIST, 59(7):1171–1183, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Ano08b].

Britz:2010:BRB

[Bri10] Johannes Britz. Book review: The Oxford Handbook of Information and Communication Technologies. Journal of the American Society for Information Science and Technology: JASIST, 61(1):209–210, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brooks:2001:ESF

[Bro01a] Terrence A. Brooks. Editorials: Still the frontier: Information science at the Millennium. Journal of the American Society for Information Science and Technology: JASIST, 52(1):1–2, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brown:2001:VPS

[Bro01b] Cecelia Brown. The E-volution of preprints in the scholarly communication of physicists and astronomers. Journal of the American Society for Information Science and Technology:

JASIST, 52(3):187–200, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brooks:2002:BRB

[Bro02] Terrence A. Brooks. Book review: Designing with Javascript: Creating dynamic Web pages. Journal of the American Society for Information Science and Technology: JASIST, 53(11):967–968, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brown:2003:CFS

[Bro03a] Cecelia Brown. The changing face of scientific discourse: Analysis of genomic and proteomic database usage and acceptance. Journal of the American Society for Information Science and Technology: JASIST, 54(10):926–938, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brown:2003:REP

[Bro03b] Cecelia Brown. The role of electronic preprints in chemical communication: Analysis of citation, usage, and acceptance in the journal literature. Journal of the American Society for Information Science and Technology: JASIST, 54(5):362– 371, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brown:2007:RWB

[Bro07] Cecelia Brown. The role of Web-based information in the scholarly communication of chemists: Citation and content analyses of American Chemical Society Journals. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2055–2065, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brody:2008:PGI

[Bro08] Roberta Brody. Perspectives on global information ethics: The problem of information naïveté. Journal of the American Society for Information Science and Technology: JA-SIST, 59(7):1124–1127, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bletsas:2009:HIR

[BS09] Aggelos Bletsas and John N. Sahalos. Hirsch index rankings require scaling and higher moment. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2577–2586, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyack:2013:IAC

[BSK13] Kevin W. Boyack, Henry Small, and Richard Klavans. Improving the accuracy of co-citation clustering using full text. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1759–1767, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Borgman:2005:CFI

[BSM^{+}05] Christine L. Borgman, Laura J. Smart, Kelli A. Millwood, Jason R. Finley, Leslie Champeny, Anne J. Gilliland, and Gregory H. Leazer. Comparing faculty information seeking in teaching and research: Implications for the design of digital libraries. Journal of the American Society for Information Science and Technology: JASIST, 56(6):636–657, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bornmann:2011:IOA

[BSMD11] Lutz Bornmann, Hermann Schier, Werner Marx, and Hans-Dieter Daniel. Is interactive open access publishing able to identify high-impact submissions? A study on the predictive validity of atmospheric chemistry and physics by using percentile rank classes. Journal of the American Society for Information Science and Technology: JASIST, 62(1):61–71, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bensman:2010:MCR

[BSP10] Stephen J. Bensman, Lawrence J. Smolinsky, and Alexander I. Pudovkin. Mean citation rate per article in mathematics journals: Differences from the scientific model. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1440–1463, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bartlett:2005:DPB

[BT05] Joan C. Bartlett and Elaine G. Toms. Developing a protocol for bioinformatics analysis: an integrated information behavior and task analysis approach. Journal of the American Society for Information Science and Technology: JA-SIST, 56(5):469–482, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Barjak:2008:SAW

[BT08] Franz Barjak and Mike Thelwall. A statistical analysis of the web presences of European life sciences research teams. Journal of the American Society for Information Science and Technology: JASIST, 59(4):628–643, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Buckland:2008:LED

[Buc08] Michael K. Buckland. Letter to the Editor: Democratic theory in library information science. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1534, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Buckland:2010:BRB

[Buc10] Michael Buckland. Book review: Information Science in Transition. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1505, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Buchanan:2011:RIS

[Buc11] Sarah A. Buchanan. Regional information science: Education and disciplinary precept in the Los Angeles Chapter of ASIST, 1961–2011. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2277– 2289, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Buckland:2012:WKS

[Buc12] Michael Buckland. What kind of science can information science be? Journal of the American Society for Information Science and Technology: JASIST, 63(1):1–7, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Buckland:2013:MYM

[Buc13] Michael K. Buckland. In memoriam: Yale Mitchell Braunstein, 1945–2012. Journal of the American Society for Information Science and Technology: JASIST, 64(2):433, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Budd:2007:IAI

[Bud07] John M. Budd. Information, analysis, and ideology: a case study of science and the public interest. Journal of the American Society for Information Science and Technology: JA-SIST, 58(14):2366–2371, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2001:BCA

[Bur01] Quentin L. Burrell. Brief communication: “ambiguity” and scientometric measurement: a dissenting view. Journal of the American Society for Information Science and Technology: JASIST, 52(12):1075–1080, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burke:2002:DNU

[Bur02a] Lauren Burke. Designing a new urban Internet. Journal of the American Society for Information Science and Technology: JASIST, 53(10):863–865, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2002:WPE

[Bur02b] Quentin L. Burrell. Will this paper ever be cited? Journal of the American Society for Information Science and Technology: JASIST, 53(3):232–235, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2003:LES

[Bur03a] Quentin L. Burrell. Letters to the Editor: The sample size dependency of statistical measures in informetrics? Some comments. Journal of the American Society for Information Science and Technology: JASIST, 54(11):1076–1077, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See comments [YKT03a].

Burrell:2003:PFC

[Bur03b] Quentin L. Burrell. Predicting future citation behavior. Journal of the American Society for Information Science and Technology: JASIST, 54(5):372–378, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2003:TTT

[Bur03c] Quentin L. Burrell. “Type/ token-taken” informetrics: Some comments and further examples. Journal of the American Society for Information Science and Technology: JASIST, 54(13):1260–1263, November 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2004:LES

[Bur04a] Quentin L. Burrell. Letter to the Editor: Sample-size dependence or time dependence of statistical measures in informetrics? Journal of the American Society for Information Science and Technology: JASIST, 55(2):183–184, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2004:LEF

[Bur04b] Quentin L. Burrell. Letters to the Editor: Fitting Lotka’s Law: Some cautionary observations on a recent paper by Newby et al. (2003). Journal of the American Society for Information Science and Technology: JASIST, 55(13):1209– 1210, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [NGJ03].

Burrell:2005:MSC

[Bur05] Quentin L. Burrell. Measuring similarity of concentration between different informetric distributions: Two new approaches. Journal of the American Society for Information Science and Technology: JASIST, 56(7):704–714, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2006:BCE

[Bur06a] Quentin L. Burrell. Brief communication: On Egghe’s version of continuous concentration theory. Journal of the American Society for Information Science and Technology:

JASIST, 57(10):1406–1411, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2006:LES

[Bur06b] Quentin L. Burrell. Letter to the Editor: Some comments on the use of reliability techniques. Journal of the American Society for Information Science and Technology: JASIST, 57(7):991–992, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2007:BCE

[Bur07] Quentin L. Burrell. Brief communication: Egghe’s construction of Lorenz curves resolved. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2157–2159, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burke:2009:BRB

[Bur09a] Colin Burke. Book review: European Modernism and the Information Society: Informing the Present, Understanding the Past. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1298–1299, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2009:BCS

[Bur09b] Quentin L. Burrell. Brief communications: Some comments on “A Proposal for a Dynamic h-Type Index” by Rousseau and Ye. Journal of the American Society for Information Science and Technology: JASIST, 60(2):418–419, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [RY08].

Burford:2011:CPW

[Bur11] Sally Burford. Complexity and the practice of web information architecture. Journal of the American Society for Information Science and Technology: JASIST, 62(10):2024–2037, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2012:BCA

[Bur12] Quentin L. Burrell. Brief communication: Alternative thoughts on uncitedness. Journal of the American Society for Information Science and Technology: JASIST, 63 (7):1466–1470, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Burrell:2013:FIL

[Bur13] Quentin L. Burrell. Formulae for the h-index: a lack of robustness in Lotkaian informetrics? Journal of the American Society for Information Science and Technology: JASIST, 64(7):1504–1514, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Buschman:2007:DTL

[Bus07] John Buschman. Democratic theory in library information science: Toward an emendation. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1483–1496, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bollen:2008:UIF

[BV08] Johan Bollen and Herbert Van de Sompel. Usage impact factor: the effects of sample characteristics on usage-based impact metrics. Journal of the American Society for Information Science and Technology: JASIST, 59(1):136–149, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bilal:2005:CCS

[BW05] Dania Bilal and Peiling Wang. Children’s conceptual structures of science categories and the design of Web directories. Journal of the American Society for Information Science and Technology: JASIST, 56(12):1303–1313, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bodoff:2006:DQR

[BW06] David Bodoff and Samuel Po-Shing Wong. Documents and queries as random variables: History and implications. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1138–1154, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Boyack:2002:DVU

[BWD02] Kevin W. Boyack, Brian N. Wylie, and George S. Davidson. Domain visualization using VxInsight(R) for science and technology management. Journal of the American Society for Information Science and Technology: JASIST, 53 (9):764–774, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bergman:2010:EFS

[BWS^{+}10] Ofer Bergman, Steve Whittaker, Mark Sanderson, Rafi Nachmias, and Anand Ramamoorthy. The effect of folder structure on personal file navigation. Journal of the American Society for Information Science and Technology: JA-SIST, 61(12):2426–2441, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bodoff:2003:RDL

[BWW03] David Bodoff, Bin Wu, and K. Y. Michael Wong. Relevance data for language models using maximum likelihood. Journal of the American Society for Information Science and Technology: JASIST, 54(11):1050–1061, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Baeza-Yates:2002:P

[BYCMS02] Ricardo Baeza-Yates, David Carmel, Yoelle Maarek, and Aya Soffer. Preface. Journal of the American Society for Information Science and Technology: JASIST, 53(6):413– 414, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Baeza-Yates:2007:MWR

[BYHM07] Ricardo Baeza-Yates, Carlos Hurtado, and Marcelo Mendoza. Mining Web resources for enhancing information retrieval: Improving search engines by query clustering. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1793–1804, October 2007. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Baeza-Yates:2002:XPN

[BYN02] Ricardo Baeza-Yates and Gonzalo Navarro. XQL and proximal nodes. Journal of the American Society for Information Science and Technology: JASIST, 53(6):504–514, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Bystrom:2002:IIS

[Bys02] Katriina Byström. Information and information sources in tasks of varying complexity. Journal of the American Society for Information Science and Technology: JASIST, 53 (7):581–591, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Benitez:2007:PMM

[BZC07] Ana B. Benitez, Di Zhong, and Shih-Fu Chang. Perspectives on MPEG-7: Metadata for multimedia: Enabling MPEG-7 structural and semantic descriptions in retrieval applications. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1377–1380, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Campanario:2007:RHC

[CA07] Juan Miguel Campanario and Erika Acedo. Rejecting highly cited papers: the views of scientists who encounter resistance to their discoveries from other scientists. Journal of the American Society for Information Science and Technology: JASIST, 58(5):734–743, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cabanac:2012:SLR

[Cab12] Guillaume Cabanac. Shaping the landscape of research in information systems from the perspective of editorial boards: a scientometric study of 77 leading journals. Journal of the American Society for Information Science and Technology: JASIST, 63(5):977–996, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cheshire:2010:BAE

[CAC10] Coye Cheshire, Judd Antin, and Elizabeth Churchill. Behaviors, adverse events, and dispositions: an empirical study of online discretion and information control. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1487–1501, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Caldeira:2008:BRB

[Cal08] Carlos Caldeira. Book review: Group cognition: Computer support for building collaborative knowledge. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1531, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Campanario:2010:BCS

[Cam10a] Juan Miguel Campanario. Brief communication: Selfcitations that contribute to the journal impact factor: an investment-benefit-yield analysis. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2575–2580, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Campanario:2010:BCD

[Cam10b] Juan Miguel Campanario. Brief communications: Distribution of ranks of articles and citations in journals. Journal of the American Society for Information Science and Technology: JASIST, 61(2):419–423, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Campanario:2011:LID

[Cam11] Juan Miguel Campanario. Large increases and decreases in journal impact factors in only one year: the effect of journal self-citations. Journal of the American Society for Information Science and Technology: JASIST, 62(2):230–235, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Capurro:2008:PGI

[Cap08] Rafael Capurro. Perspectives on global information ethics: Information ethics for and from Africa. Journal of the American Society for Information Science and Technology: JA-SIST, 59(7):1162–1170, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Case:2003:BRB

[Cas03] Donald O. Case. Book review: Current theory in library and information science. Journal of the American Society for Information Science and Technology: JASIST, 54(4):358– 359, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Case:2008:CFH

[Cas08] Donald O. Case. Collection of family health histories: the link between genealogy and public health. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2312–2319, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chua:2013:FGA

[CB13] Alton Y. K. Chua and Snehasish Banerjee. So fast so good: an analysis of answer quality and answer speed in community Question-answering sites. Journal of the American Society for Information Science and Technology: JASIST, 64(10): 2058–2068, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Choo:2008:ICI

[CBDH08] Chun Wei Choo, Pierrette Bergeron, Brian Detlor, and Lorna Heaton. Information culture and information use: an exploratory study of three organizations. Journal of the American Society for Information Science and Technology: JASIST, 59(5):792–804, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cole:2013:SIM

[CBL^{+}13] Charles Cole, Jamshid Behesthi, Andrew Large, Isabelle Lamoureux, Dhary Abuhimed, and Mohammed AlGhamdi. Seeking information for a middle school history project: the concept of implicit knowledge in the students’ transition from Kuhlthau’s Stage 3 to Stage 4. Journal of the American Society for Information Science and Technology: JA-SIST, 64(3):558–573, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cui:2010:SAB

[CBS10] Hong Cui, David Boufford, and Paul Selden. Semantic annotation of biosystematics literature without training examples. Journal of the American Society for Information Science and Technology: JASIST, 61(3):522–542, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Costas:2009:SRS

[CBvLvR09] Rodrigo Costas, Maria Bordons, Thed N. van Leeuwen, and Anthony F. J. van Raan. Scaling rules in the science system: Influence of field-specific citation characteristics on the impact of individual researchers. Journal of the American Society for Information Science and Technology: JASIST, 60(4):740–753, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2001:UCT

[CC01a] Hui-Min Chen and Michael D. Cooper. Using clustering techniques to detect usage patterns in a Web-based information system. Journal of the American Society for Information Science and Technology: JASIST, 52(11):888–904, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cooper:2001:RPR

[CC01b] Michael D. Cooper and Hui-Min Chen. Research: Predicting the relevance of a library catalog search. Journal of the American Society for Information Science and Technology: JASIST, 52(10):813–827, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2002:SMU

[CC02] Hui-Min Chen and Michael D. Cooper. Stochastic modeling of usage patterns in a web-based information system. Journal of the American Society for Information Science and Technology: JASIST, 53(7):536–548, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Covi:2004:RCL

[CC04] Lisa M. Covi and Melissa H. Cragin. Reconfiguring control in library collection development: a conceptual framework for assessing the shift toward electronic collections. Journal of the American Society for Information Science and Technology: JASIST, 55(4):312–325, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chung:2009:BUW

[CC09] Wingyan Chung and Hsinchun Chen. Browsing the underdeveloped Web: an experiment on the Arabic Medical Web Directory. Journal of the American Society for Information Science and Technology: JASIST, 60(3):595–607, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chang:2011:CVS

[CC11] Y. F. Chang and C.-M. Chen. Classification and visualization of the social science network by the minimum span clustering method. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2404–2413, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cox:2013:AIS

[CC13] Andrew M. Cox and Sheila Corrall. Advances in information science: Evolving academic library specialties. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1526–1542, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cho:2010:TIT

[CCC10] Hichang Cho, MeiHui Chen, and Siyoung Chung. Testing an integrative theoretical model of knowledge-sharing behavior in the context of Wikipedia. Journal of the American Society for Information Science and Technology: JASIST, 61(6): 1198–1212, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cabanac:2010:SVC

[CCCJ10] Guillaume Cabanac, Max Chevalier, Claude Chrisment, and Christine Julien. Social validation of collective annotations: Definition and experiment. Journal of the American Society for Information Science and Technology: JASIST, 61(2): 271–287, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Calado:2006:LBS

[CCG^{+}06] Pável Calado, Marco Cristo, Marcos André Gonçalves, Edleno S. de Moura, Berthier Ribeiro-Neto, and Nivio Ziviani. Link-based similarity measures for the classification of Web documents. Journal of the American Society for Information Science and Technology: JASIST, 57(2):208–221, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2002:VTG

[CCMM02] Chaomei Chen, Timothy Cribbin, Robert Macredie, and Sonali Morar. Visualizing and tracking the growth of competing paradigms: Two case studies. Journal of the American Society for Information Science and Technology: JA-SIST, 53(8):678–689, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2008:UDW

[CCQ^{+}08] Hsinchun Chen, Wingyan Chung, Jialun Qin, Edna Reid, Marc Sageman, and Gabriel Weimann. Uncovering the dark Web: a case study of Jihad on the Web. Journal of the American Society for Information Science and Technology: JA-SIST, 59(8):1347–1359, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chung:2009:BSA

[CCR09] Wingyan Chung, Hsinchun Chen, and Edna Reid. Business stakeholder analyzer: an experiment of classifying stakeholders on the Web. Journal of the American Society for Information Science and Technology: JASIST, 60(1):59–74, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cox:2011:DMC

[CCS11] Andrew Cox, Paul Clough, and Stefan Siersdorfer. Developing metrics to characterize Flickr groups. Journal of the American Society for Information Science and Technology: JASIST, 62(3):493–506, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Crestani:2006:WVS

[CD06] Fabio Crestani and Heather Du. Written versus spoken queries: a qualitative and quantitative comparative analysis. Journal of the American Society for Information Science and Technology: JASIST, 57(7):881–890, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Crestani:2003:MLF

[CDLvR03] Fabio Crestani, Sandor Dominich, Mounia Lalmas, and Cornelis Joost (Keith) van Rijsbergen. Mathematical, logical, and formal methods in information retrieval: an introduction to the special issue. Journal of the American Society for Information Science and Technology: JASIST, 54(4):281– 284, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cao:2011:RRQ

[CDLY11] Yunbo Cao, Huizhong Duan, Chin-Yew Lin, and Yong Yu. Re-ranking question search results by clustering questions. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1177–1187, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cortez:2009:FAE

[CdSG^{+}09] Eli Cortez, Altigran S. da Silva, Marcos André Gonçalves, Filipe Mesquita, and Edleno S. de Moura. A flexible approach for extracting metadata from bibliographic citations. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1144–1158, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ceruzzi:2010:BRB

[Cer10] Paul Ceruzzi. Book review: Punched-Card Systems and the Early Information Explosion, 1880–1945. Journal of the American Society for Information Science and Technology: JASIST, 61(2):428, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2006:TCR

[CF06] Blaise Cronin and Sara Franks. Trading cultures: Resource mobilization and service rendering in the life sciences as revealed in the journal article’s paratext. Journal of the American Society for Information Science and Technology: JA-SIST, 57(14):1909–1918, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Camara:2007:IPO

[CF07a] Gilberto Câmara and Frederico Fonseca. Information policies and open source software in developing countries. Journal of the American Society for Information Science and Technology: JASIST, 58(1):121–132, January 1, 2007. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2007:E

[CF07b] Zhixiang Chen and Bin Fu. Erratum. Journal of the American Society for Information Science and Technology: JA-SIST, 58(14):2414, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2007:CRS

[CF07c] Zhixiang Chen and Bin Fu. On the complexity of Rocchio’s similarity-based relevance feedback algorithm. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1392–1400, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Colazo:2009:ILC

[CF09] Jorge Colazo and Yulin Fang. Impact of license choice on Open Source Software development activity. Journal of the American Society for Information Science and Technology: JASIST, 60(5):997–1011, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2001:MMS

[CFCZ01] Hsinchun Chen, Haiyan Fan, Michael Chau, and Daniel Zeng. MetaSpider: Meta-searching and categorization on the Web. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1134–1147, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cota:2010:UHB

[CFN^{+}10] Ricardo G. Cota, Anderson A. Ferreira, Cristiano Nascimento, Marcos André Gonçalves, and Alberto H. F. Laender. An unsupervised heuristic-based hierarchical method for name disambiguation in bibliographic citations. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1853–1870, September 2010. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chau:2005:AQL

[CFS05] Michael Chau, Xiao Fang, and Olivia R. Liu Sheng. Analysis of the query logs of a Web site search engine. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1363–1376, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cheng:2012:RIW

[CFW^{+}12] An-Shou Cheng, Kenneth R. Fleischmann, Ping Wang, Emi Ishita, and Douglas W. Oard. The role of innovation and wealth in the net neutrality debate: a content analysis of human values in Congressional and FCC hearings. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1360–1373, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chau:2007:WSC

[CFY07] Michael Chau, Xiao Fang, and Christopher C. Yang. Web searching in Chinese: a study of a search engine in Hong Kong. Journal of the American Society for Information Science and Technology: JASIST, 58(7):1044–1054, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Crystal:2006:RCI

[CG06] Abe Crystal and Jane Greenberg. Relevance criteria identified by health information users during Web searches. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1368–1382, August 2006. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chalmeta:2008:MIK

[CG08] Ricardo Chalmeta and Reyes Grangel. Methodology for the implementation of knowledge management systems. Journal of the American Society for Information Science and Technology: JASIST, 59(5):742–755, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2009:EPL

[CGY09] Yu-Hui Chen, Carol Anne Germain, and Huahai Yang. An exploration into the practices of library Web usability in ARL academic libraries. Journal of the American Society for Information Science and Technology: JASIST, 60(5):953– 968, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2003:NMD

[CH03] Ruey-Shun Chen and Yi-Chung Hu. A novel method for discovering fuzzy sequential patterns using the simple fuzzy partition method. Journal of the American Society for Information Science and Technology: JASIST, 54(7):660–670, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chaplan:2005:RRL

[CH05] Margaret A. Chaplan and Edward J. Hertenstein. Rolerelated library use by local union officials. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1062–1074, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cui:2007:RIK

[CH07] Hong Cui and P. Bryan Heidorn. The reusability of induced knowledge for the automatic semantic markup of taxonomic descriptions. Journal of the American Society for Information Science and Technology: JASIST, 58(1):133–149, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Callahan:2011:CBW

[CH11] Ewa S. Callahan and Susan C. Herring. Cultural bias in Wikipedia content on famous persons. Journal of the American Society for Information Science and Technology: JA-SIST, 62(10):1899–1915, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chang:2012:SEI

[CH12] Yu-Wei Chang and Mu-Hsuan Huang. A study of the evolution of interdisciplinarity in library and information science: Using three bibliometric methods. Journal of the American Society for Information Science and Technology: JASIST, 63(1):22–33, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cabanac:2013:BCI

[CH13] Guillaume Cabanac and James Hartley. Brief communication: Issues of work–life balance among JASIST authors and editors. Journal of the American Society for Information Science and Technology: JASIST, 64(10):2182–2186, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chawner:2003:BR

[Cha03] Brenda Chawner. Book review. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1166–1167, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chang:2009:OSI

[Cha09] Hui-Jung Chang. Online supportive interactions: Using a network approach to examine communication patterns within a psychosis social support group in Taiwan. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1504–1517, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chang:2011:MWC

[Cha11] Hui-Jung Chang. Multinationals on the web: Cultural similarities and differences in English-language and Chineselanguage website designs. Journal of the American Society for Information Science and Technology: JASIST, 62(6): 1105–1117, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chawner:2013:BRB

[Cha13] Brenda Chawner. Book review: Opening standards: the global politics of interoperability — Edited by Laura DeNardis. Journal of the American Society for Information Science and Technology: JASIST, 64(4):868–870, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cortez:2011:LML

[CHdS^{+}11] Eli Cortez, Mauro Rojas Herrera, Altigran S. da Silva, Edleno S. de Moura, and Marden Neubert. Lightweight methods for large-scale product categorization. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1839–1848, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2003:BRB

[Che03a] Chaomei Chen. Book review: Patents, citations & innovations: a window on the knowledge economy. Journal of the American Society for Information Science and Technology: JASIST, 54(8):802–803, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2003:EVS

[Che03b] Chaomei Chen. Editorial: Visualizing scientific paradigms: an introduction. Journal of the American Society for Information Science and Technology: JASIST, 54(5):392–393, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2003:IBS

[Che03c] Hsinchun Chen. Introduction to the JASIST special topic issue on web retrieval and mining: a machine learning perspective. Journal of the American Society for Information Science and Technology: JASIST, 54(7):621–624, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chersky:2003:LEW

[Che03d] Joseph Chersky. Letters to the Editor: What is time? Journal of the American Society for Information Science and Technology: JASIST, 54(11):1078, September 2003. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2004:BRB

[Che04] Chaomei Chen. Book review: Mining the Web: Discovering knowledge from hypertext data. Journal of the American Society for Information Science and Technology: JASIST, 55(3):275–276, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2005:IIS

[Che05] Hsinchun Chen. Introduction: Introduction to the special topic issue: Intelligence and security informatics. Journal of the American Society for Information Science and Technology: JASIST, 56(3):217–220, February 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Che:2006:BRB

[Che06a] Haoyang Che. Book review: A Semantic Web primer. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1132, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2006:CID

[Che06b] Chaomei Chen. CiteSpace II: Detecting and visualizing emerging trends and transient patterns in scientific literature. Journal of the American Society for Information Science and Technology: JASIST, 57(3):359–377, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2006:LKB

[Che06c] Jiangping Chen. A lexical knowledge base approach for English–Chinese cross-language information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 57(2):233–243, January 15, 2006. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2008:CSN

[Che08] C.-M. Chen. Classification of scientific networks using aggregated journal–journal citation relations in the Journal Citation Reports. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2296–2304, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2009:BRB

[Che09] Chaomei Chen. Book review: The New Invisible College: Science for Development. Journal of the American Society for Information Science and Technology: JASIST, 60(8): 1719–1720, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Callahan:2010:CCA

[CHE10a] Alison Callahan, Stephen Hockema, and Gunther Eysenbach. Contextual cocitation: Augmenting cocitation analysis and its applications. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1130– 1143, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2010:BRB

[Che10b] Sherry Chen. Book review: Web-Based Learning Through Educational Informatics: Information Science Meets Educational Computing. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2164– 2165, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2012:PES

[Che12] Chaomei Chen. Predictive effects of structural variation on citation counts. Journal of the American Society for Information Science and Technology: JASIST, 63(3):431–449, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chiang:2009:CWU

[CHH09] I-Ping Chiang, Chun-Yao Huang, and Chien-Wen Huang. Characterizing Web users’ degree of Web 2.0-ness. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1349–1357, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2013:VAS

[CHMS13] Chaomei Chen, Zhigang Hu, Jared Milbank, and Timothy Schultz. A visual analytic study of retracted articles in scientific literature. Journal of the American Society for Information Science and Technology: JASIST, 64(2):234–253, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Choo:2009:IUE

[Cho09a] Chun Wei Choo. Information use and early warning effectiveness: Perspectives and prospects. Journal of the American Society for Information Science and Technology: JA-SIST, 60(5):1071–1082, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chopra:2009:BRB

[Cho09b] Samir Chopra. Book review: Open Source: Technology and Policy. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2382–2383, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Choi:2010:ECF

[Cho10a] Youngok Choi. Effects of contextual factors on image searching on the Web. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2011–2028, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chowdhury:2010:BRB

[Cho10b] Gobinda Chowdhury. Book review: Social Information Retrieval Systems: Emerging Technologies and Applications for Searching the Web Effectively. Journal of the American Society for Information Science and Technology: JASIST, 61 (12):2587–2588, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chowdhury:2012:AIS

[Cho12] Gobinda Chowdhury. Advances in information science: Building environmentally sustainable information services: a green is research agenda. Journal of the American Society for Information Science and Technology: JASIST, 63 (4):633–647, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Choi:2013:ISA

[Cho13a] Namjoo Choi. Information systems attachment: an empirical exploration of its antecedents and its impact on community participation intention. Journal of the American Society for Information Science and Technology: JASIST, 64 (11):2354–2365, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Choi:2013:AIS

[Cho13b] Youngok Choi. Analysis of image search queries on the web: Query modification patterns and semantic attributes. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1423–1441, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chrzastowski:2003:MTP

[Chr03] Tina E. Chrzastowski. Making the transition from print to electronic serial collections: a new model for academic chemistry libraries? Journal of the American Society for Information Science and Technology: JASIST, 54(12):1141–1148, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chrzastowski:2004:LER

[Chr04] Tina Chrzastowski. Letters to the Editor: Rejoinder: Transition from print to electronic serials. Journal of the American Society for Information Science and Technology: JA-SIST, 55(4):370–371, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chu:2001:RII

[Chu01] Heting Chu. Research in image indexing and retrieval as reflected in the literature. Journal of the American Society for Information Science and Technology: JASIST, 52(12): 1011–1018, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chua:2005:DIS

[Chu05] Alton Y. K. Chua. The design and implementation of a simulation game for teaching knowledge management. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1207–1216, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chua:2007:TTH

[Chu07] Alton Y. K. Chua. A tale of two hurricanes: Comparing Katrina and Rita through a knowledge management perspective. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1518–1528, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chu:2009:IPB

[Chu09] Kai Wah Samuel Chu. Inquiry project-based learning with a partnership of three types of teachers and the school librarian. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1671–1686, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2010:SDC

[CISH10] Chaomei Chen, Fidelia Ibekwe-SanJuan, and Jianhua Hou. The structure and dynamics of cocitation clusters: a multiple-perspective cocitation analysis. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1386–1409, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Crudge:2004:UIS

[CJ04] Sarah E. Crudge and Frances C. Johnson. Using the information seeker to elicit construct models for search engine evaluation. Journal of the American Society for Information Science and Technology: JASIST, 55(9):794–806, July 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Case:2004:TSF

[CJA^{+}04] Donald O. Case, J. David Johnson, James E. Andrews, Suzanne L. Allard, and Kimberly M. Kelly. From two-step flow to the Internet: the changing array of sources for genetics information seeking. Journal of the American Society for Information Science and Technology: JASIST, 55(8):660– 669, June 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cathey:2007:EPS

[CJB^{+}07] Rebecca J. Cathey, Eric C. Jensen, Steven M. Beitzel, Ophir Frieder, and David Grossman. Exploiting parallelism to support scalable hierarchical clustering. Journal of the American Society for Information Science and Technology: JA-SIST, 58(8):1207–1221, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cunliffe:2002:IAB

[CJJ^{+}02] Daniel Cunliffe, Helen Jones, Melanie Jarvis, Kevin Egan, Rhian Huws, and Siân Munro. Information architecture for bilingual Web sites. Journal of the American Society for Information Science and Technology: JASIST, 53(10):866– 873, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chinenyanga:2002:EEL

[CK02] Taurai Tapiwa Chinenyanga and Nicholas Kushmerick. An expressive and efficient language for XML information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 53(6):438–453, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2003:RLV

[CK03] Chaomei Chen and Jasna Kuljis. The rising landscape: a visual exploration of superstring revolutions in physics. Journal of the American Society for Information Science and Technology: JASIST, 54(5):435–446, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chung:2008:BAA

[CK08] Deborah S. Chung and Sujin Kim. Blogging activity among cancer patients and their companions: Uses, gratifications, and predictors of outcomes. Journal of the American Society for Information Science and Technology: JASIST, 59(2): 297–306, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Can:2008:IRT

[CKB^{+}08] Fazli Can, Seyit Kocberber, Erman Balcik, Cihan Kaynak, H. Cagdas Ocalan, and Onur M. Vursavas. Information retrieval on Turkish texts. Journal of the American Society for Information Science and Technology: JASIST, 59(3):407– 421, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Can:2010:NED

[CKB^{+}10] Fazli Can, Seyit Kocberber, Ozgur Baglioglu, Suleyman Kardas, H. Cagdas Ocalan, and Erkan Uyar. New event detection and topic tracking in Turkish. Journal of the American Society for Information Science and Technology: JA-SIST, 61(4):802–819, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chua:2007:ADR

[CKF07] Alton Y. K. Chua, Selcan Kaynak, and Schubert S. B. Foo. An analysis of the delayed response to Hurricane Katrina through the lens of knowledge management. Journal of the American Society for Information Science and Technology: JASIST, 58(3):391–403, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2003:SSC

[CKH^{+}03] Hsin-Hsi Chen, June-Jei Kuo, Sheng-Jie Huang, Chuan-Jie Lin, and Hung-Chia Wung. A summarization system for Chinese news from multiple sources. Journal of the American Society for Information Science and Technology: JASIST, 54(13):1224–1236, November 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cohen:2002:ESQ

[CKK^{+}02] Sara Cohen, Yaron Kanza, Yakov Kogan, Yehoshua Sagiv, Werner Nutt, and Alexander Serebrenik. EquiX — a search and query language for XML. Journal of the American Society for Information Science and Technology: JASIST, 53 (6):454–466, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Capra:2013:WPM

[CKR13] Robert Capra, Julia Khanova, and Sarah Ramdeen. Work and personal e-mail use by university employees: PIM practices across domain boundaries. Journal of the American Society for Information Science and Technology: JASIST, 64(5):1029–1044, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chan:2006:ISC

[CKT06] Hock Chuan Chan, Hee-Woong Kim, and Weai Chee Tan. Information systems citation patterns from International Conference on Information Systems articles. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1263–1274, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chung:2001:CBA

[CL01] Young Mee Chung and Jae Yun Lee. A corpus-based approach to comparative evaluation of statistical term association measures. Journal of the American Society for Information Science and Technology: JASIST, 52(4):283–296, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cole:2003:UUM

[CL03] Charles Cole and John E. Leide. Using the user’s mental model to guide the integration of information space into information need. Journal of the American Society for Information Science and Technology: JASIST, 54(1):39–46, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cheung:2006:UCT

[CL06] Christy M. K. Cheung and Matthew K. O. Lee. Understanding consumer trust in Internet shopping: a multidisciplinary approach. Journal of the American Society for Information Science and Technology: JASIST, 57(4):479–492, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cheung:2008:SWB

[CL08] Christy M. K. Cheung and Matthew K. O. Lee. The structure of Web-based information systems satisfaction: Testing of competing models. Journal of the American Society for Information Science and Technology: JASIST, 59(10): 1617–1630, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cheung:2009:USI

[CL09] Christy M. K. Cheung and Matthew K. O. Lee. User satisfaction with an Internet-based portal: an asymmetric and nonlinear approach. Journal of the American Society for Information Science and Technology: JASIST, 60(1):111–122, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cai:2011:ESL

[CL11] Xiaoyan Cai and Wenjie Li. Enhancing sentence-level clustering with integrated and interactive frameworks for themebased summarization. Journal of the American Society for Information Science and Technology: JASIST, 62(10):2067– 2082, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cortese:2012:CTI

[CL12] Juliann Cortese and Mia Liza A. Lustria. Can tailoring increase elaboration of health messages delivered via an adaptive educational site on adolescent sexual health and decision making? Journal of the American Society for Information Science and Technology: JASIST, 63(8):1567–1580, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cole:2005:IAS

[CLB^{+}05] Charles Cole, John Leide, Jamshid Beheshti, Andrew Large, and Martin Brooks. Investigating the Anomalous States of Knowledge hypothesis in a real-life problem situation: a study of history and psychology undergraduates seeking information for a course essay. Journal of the American Society for Information Science and Technology: JASIST, 56 (14):1544–1554, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cao:2013:RBB

[CLD^{+}13] Qilin Cao, Yong Lu, Dayong Dong, Zongming Tang, and Yongqiang Li. The roles of bridging and bonding in social media communities. Journal of the American Society for Information Science and Technology: JASIST, 64(8): 1671–1681, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2013:TMA

[CLH13] Yen-Liang Chen, Yi-Hung Liu, and Wu-Liang Ho. A text mining approach to assist the general public in the retrieval of legal documents. Journal of the American Society for Information Science and Technology: JASIST, 64(2):280–290, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cobo:2011:SMS

[CLHHVH11] M. J. Cobo, A. G. López-Herrera, E. Herrera-Viedma, and F. Herrera. Science mapping software tools: Review, analysis, and cooperative study among tools. Journal of the American Society for Information Science and Technology: JA-SIST, 62(7):1382–1402, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cobo:2012:SNS

[CLHHVH12] M. J. Cobo, A. G. López-Herrera, E. Herrera-Viedma, and F. Herrera. SciMAT: a new science mapping analysis software tool. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1609–1630, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cole:2005:PIT

[CLL^{+}05] Charles Cole, John E. Leide, Andrew Large, Jamshid Beheshti, and Martin Brooks. Putting it together online: Information need identification for the domain novice user. Journal of the American Society for Information Science and Technology: JASIST, 56(7):684–694, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cole:2007:CMM

[CLL^{+}07] Charles Cole, Yang Lin, John Leide, Andrew Large, and Jamshid Beheshti. A classification of mental models of undergraduates seeking information for a course essay in history and psychology: Preliminary investigations into aligning their mental models with online thesauri. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2092–2104, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cheung:2013:UCI

[CLL13] Christy M. K. Cheung, Matthew K. O. Lee, and Zach W. Y. Lee. Understanding the continuance intention of knowledge sharing in online communities of practice through the post-knowledge-sharing evaluation processes. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1357–1374, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cecchini:2010:MEA

[CLMB10] Rocío L. Cecchini, Carlos M. Lorenzetti, Ana G. Maguitman, and Nélida B. Brignole. Multiobjective evolutionary algorithms for context-based search. Journal of the American Society for Information Science and Technology: JA-SIST, 61(6):1258–1274, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Clyde:2004:EQR

[Cly04] Laurel A. Clyde. Evaluating the quality of research publications: a pilot study of school librarianship. Journal of the American Society for Information Science and Technology: JASIST, 55(13):1119–1130, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2006:TTN

[CLYL06] Hsin-Hsi Chen, Wen-Cheng Lin, Changhua Yang, and Wei-Hao Lin. Translating–transliterating named entities for multilingual information access. Journal of the American Society for Information Science and Technology: JASIST, 57(5): 645–659, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2003:HIS

[CLZC03] Hsinchun Chen, Ann M. Lally, Bin Zhu, and Michael Chau. HelpfulMed: Intelligent searching for medical information over the Internet. Journal of the American Society for Information Science and Technology: JASIST, 54(7):683–694, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2002:CSH

[CM02] Sherry Y. Chen and Robert D. Macredie. Cognitive styles and hypermedia navigation: Development of a learning model. Journal of the American Society for Information Science and Technology: JASIST, 53(1):3–15, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2006:BCU

[CM06] Blaise Cronin and Lokman Meho. Brief communication: Using the h-index to rank influential information scientistss. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1275–1278, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2007:TCS

[CM07] Blaise Cronin and Lokman I. Meho. Timelines of creativity: a study of intellectual innovators in information science. Journal of the American Society for Information Science and Technology: JASIST, 58(13):1948–1959, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2008:BCA

[CM08a] Blaise Cronin and Lokman I. Meho. Brief communications: Applying the author affiliation index to library and information science journals. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1861– 1865, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2008:SBI

[CM08b] Blaise Cronin and Lokman I. Meho. The shifting balance of intellectual trade in information studies. Journal of the American Society for Information Science and Technology: JASIST, 59(4):551–564, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Case:2011:DBC

[CM11] Donald O. Case and Joseph B. Miller. Do bibliometricians cite differently from other scholars? Journal of the American Society for Information Science and Technology: JASIST, 62(3):421–432, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2005:FID

[CMD05] Sherry Y. Chen, George D. Magoulas, and Dionisios Dimakopoulos. A flexible interface design for Web directories to accommodate different cognitive styles. Journal of the American Society for Information Science and Technology: JASIST, 56(1):70–83, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2001:RFR

[CMFZ01] Zhixiang Chen, Xiannong Meng, Richard H. Fowler, and Binhai Zhu. Research: Features: Real-time adaptive feature and document learning for web search. Journal of the American Society for Information Science and Technology: JASIST, 52(8):655–665, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chung:2010:FAS

[CMH10] EunKyung Chung, Shawne Miksa, and Samantha K. Hastings. A framework of automatic subject term assignment for text categorization: an indexing conception-based approach. Journal of the American Society for Information Science and Technology: JASIST, 61(4):688–699, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Carpineto:2009:MIR

[CMRS09] Claudio Carpineto, Stefano Mizzaro, Giovanni Romano, and Matteo Snidero. Mobile information retrieval with search results clustering: Prototypes and evaluations. Journal of the American Society for Information Science and Technology: JASIST, 60(5):877–895, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chung:2007:HSS

[CN07] Jin Soo Chung and Delia Neuman. High school students’ Information seeking and use for class projects. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1503–1517, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Coleman:2005:ICU

[Col05] Anita Coleman. Instruments of cognition: Use of citations and Web links in online teaching materials. Journal of the American Society for Information Science and Technology: JASIST, 56(4):382–392, February 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Coleman:2007:AVJ

[Col07a] Anita Coleman. Assessing the value of a journal beyond the impact factor. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1148–1161, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Coleman:2007:SAC

[Col07b] Anita Coleman. Self-archiving and the Copyright Transfer Agreements of ISI-ranked library and information science journals. Journal of the American Society for Information Science and Technology: JASIST, 58(2):286–296, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cole:2011:TIN

[Col11] Charles Cole. A theory of information need for information retrieval that connects information to knowledge. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1216–1231, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cooper:2001:RUP

[Coo01] Michael D. Cooper. Research: Usage patterns of a webbased library catalog. Journal of the American Society for Information Science and Technology: JASIST, 52(2):137– 148, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cooper:2002:CSI

[Coo02a] Linda Z. Cooper. A case study of information-seeking behavior in 7-year-old children in a semistructured situation. Journal of the American Society for Information Science and Technology: JASIST, 53(11):904–922, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cooper:2002:MPE

[Coo02b] Linda Z. Cooper. Methodology for a project examining cognitive categories for library information in young children. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1223–1231, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cook:2007:BRB

[Coo07] Darrell Cook. Book review: Stimulated recall and mental models: Tools for teaching and learning computer information literacy. Journal of the American Society for Information Science and Technology: JASIST, 58(3):456–457, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Copeland:2011:APL

[Cop11] Andrea J. Copeland. Analysis of public library users’ digital preservation practices. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1288– 1300, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Copeland:2012:BRB

[Cop12] Andrea J. Copeland. Book reviews: From Fear to Flow: Personality and information interaction. Journal of the American Society for Information Science and Technology: JASIST, 63(2):424–425, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cothey:2001:LE

[Cot01] Viv Cothey. Letter to the Editor. Journal of the American Society for Information Science and Technology: JASIST, 52(7):601–602, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cothey:2002:LSW

[Cot02] Vivian Cothey. A longitudinal study of World Wide Web users’ information-searching behavior. Journal of the American Society for Information Science and Technology: JA-SIST, 53(2):67–78, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cothey:2004:WCR

[Cot04] Viv Cothey. Web-crawling reliability. Journal of the American Society for Information Science and Technology: JA-SIST, 55(14):1228–1238, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cox:2008:PGI

[Cox08] Richard J. Cox. Perspectives on global information ethics: Archival ethics: the truth of the matter. Journal of the American Society for Information Science and Technology: JASIST, 59(7):1128–1133, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cox:2009:BRB

[Cox09] Richard J. Cox. Book review: College and University Archives: Readings in Theory and Practice. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2592–2593, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cox:2010:BRB

[Cox10] Andrew M. Cox. Book review: Communities of Practice: Fostering Peer-to-Peer Learning and Informal Knowledge Sharing in the Work Place. Journal of the American Society for Information Science and Technology: JASIST, 61 (3):629–631, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cabanac:2013:COE

[CP13] Guillaume Cabanac and Thomas Preuss. Capitalizing on order effects in the bids of peer-reviewed conferences to secure reviews by expert referees. Journal of the American Society for Information Science and Technology: JASIST, 64(2): 405–415, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2001:RFJ

[CPO01] Chaomei Chen, Ray J. Paul, and Bob O’Keefe. Research: Fitting the jigsaw of citation: Information visualization in domain analysis. Journal of the American Society for Information Science and Technology: JASIST, 52(4):315–330, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Child:2009:BCP

[CPP09] Jeffrey T. Child, Judy C. Pearson, and Sandra Petronio. Blogging, communication, and privacy management: Development of the Blogging Privacy Management Measure. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2079–2094, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Choi:2003:SIA

[CR03] Youngok Choi and Edie M. Rasmussen. Searching for images: the analysis of users’ queries for image retrieval in American history. Journal of the American Society for Information Science and Technology: JASIST, 54(6):498–511, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chinchilla-Rodriguez:2012:ICM

[CRBPdMAM12] Zaida Chinchilla-Rodríguez, Maria Benavent-Pérez, Félix de Moya-Anegón, and Sandra Miguel. International collaboration in Medical Research in Latin America and the Caribbean (2003–2007). Journal of the American Society for Information Science and Technology: JASIST, 63(11): 2223–2238, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cribbin:2011:DLT

[Cri11] Timothy Cribbin. Discovering latent topical structure by second-order similarity analysis. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1188–1207, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2001:RHP

[Cro01] Blaise Cronin. Research: Hyperauthorship: a postmodern perversion or evidence of a structural shift in scholarly communication practices? Journal of the American Society for Information Science and Technology: JASIST, 52(7):558– 569, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2004:BCB

[Cro04] Blaise Cronin. Brief communications: Bowling alone together: Academic writing as distributed cognition. Journal of the American Society for Information Science and Technology: JASIST, 55(6):557–560, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2008:BCE

[Cro08] Blaise Cronin. Brief communications: On the epistemic significance of place. Journal of the American Society for Information Science and Technology: JASIST, 59(6):1002–1006, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2009:EST

[Cro09a] Blaise Cronin. Editorial: A seat at the table. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2387, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2009:ECG

[Cro09b] Blaise Cronin. Editorial: Changing of the guard. Journal of the American Society for Information Science and Technology: JASIST, 60(1):1–2, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2009:EEE

[Cro09c] Blaise Cronin. Editorial: End of an era. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1729, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2009:ERS

[Cro09d] Blaise Cronin. Editorial: Rigor and Speed. Journal of the American Society for Information Science and Technology: JASIST, 60(4):645, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2009:ECP

[Cro09e] Blaise Cronin. Editorial: The changing profile of JASIST authors. Journal of the American Society for Information Science and Technology: JASIST, 60(10):1949, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2009:EVV

[Cro09f] Blaise Cronin. Editorial: Vernacular and vehicular language. Journal of the American Society for Information Science and Technology: JASIST, 60(3):433, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2010:EAI

[Cro10a] Blaise Cronin. Editorial: Advances in information science. Journal of the American Society for Information Science and Technology: JASIST, 61(4):639, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2010:MBV

[Cro10b] Blaise Cronin. In memoriam: Brian Vickery: an appreciation. Journal of the American Society for Information Science and Technology: JASIST, 61(4):850–851, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2011:EDC

[Cro11a] Blaise Cronin. Editorial: Don’t confuse accreditation with reputation. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2299–2300, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2011:EPR

[Cro11b] Blaise Cronin. Editorial: Peer review. Journal of the American Society for Information Science and Technology: JA-SIST, 62(7):1215, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2011:EID

[Cro11c] Blaise Cronin. Editorial: The intelligence disconnect. Journal of the American Society for Information Science and Technology: JASIST, 62(10):1867–1868, October 2011. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2012:EDM

[Cro12a] Blaise Cronin. Editorial: Do me a favor. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1281, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2012:ELM

[Cro12b] Blaise Cronin. Editorial: Language matters. Journal of the American Society for Information Science and Technology: JASIST, 63(2):217, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2012:ERR

[Cro12c] Blaise Cronin. Editorial: The resilience of rejected manuscripts. Journal of the American Society for Information Science and Technology: JASIST, 63(10):1903–1904, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2013:EC

[Cro13a] Blaise Cronin. Editorial: Canonicity. Journal of the American Society for Information Science and Technology: JA-SIST, 64(11):2189–2190, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2013:EMM

[Cro13b] Blaise Cronin. Editorial: Metrics à la mode. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1091, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2013:ESP

[Cro13c] Blaise Cronin. Editorial: Self-plagiarism: an odious oxymoron. Journal of the American Society for Information Science and Technology: JASIST, 64(5):873, May 2013. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2013:ESF

[Cro13d] Blaise Cronin. Editorial: Slow Food for thought. Journal of the American Society for Information Science and Technology: JASIST, 64(1):1, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2013:ESC

[Cro13e] Blaise Cronin. Editorial: Standing on ceremony. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1309–1310, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2013:EEI

[Cro13f] Blaise Cronin. Editorial: The evolving indicator space (iSpace). Journal of the American Society for Information Science and Technology: JASIST, 64(8):1523–1525, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2013:ETA

[Cro13g] Blaise Cronin. Editorial: Thinking about data. Journal of the American Society for Information Science and Technology: JASIST, 64(3):435–436, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2002:BCB

[CS02] Blaise Cronin and Debora Shaw. Brief communications: Banking (on) different forms of symbolic capital. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1267–1270, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ceaparu:2004:FGS

[CS04] Irina Ceaparu and Ben Shneiderman. Finding governmental statistical data on the Web: a study of categorically organized links for the FedStats topics page. Journal of the American Society for Information Science and Technology: JASIST, 55(11):1008–1015, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Clough:2006:UEE

[CS06a] Paul Clough and Mark Sanderson. User experiments with the Eurovision cross-language image retrieval system. Journal of the American Society for Information Science and Technology: JASIST, 57(5):697–708, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Conrad:2006:MDV

[CS06b] Jack G. Conrad and Cindy P. Schriber. Managing déjà vu: Collection building for the identification of nonidentical duplicate documents. Journal of the American Society for Information Science and Technology: JASIST, 57(7):921–932, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Calegari:2008:OFC

[CS08a] Silvia Calegari and Elie Sanchez. Object-fuzzy concept network: an enrichment of ontologies in semantic information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2171–2185, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Carbo:2008:PGI

[CS08b] Toni Carbo and Martha M. Smith. Perspectives on global information ethics: Global information ethics: Intercultural perspectives on past and future research. Journal of the American Society for Information Science and Technology: JASIST, 59(7):1111–1123, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cetintas:2012:EQG

[CS12] Suleyman Cetintas and Luo Si. Effective query generation and postprocessing strategies for prior art patent search. Journal of the American Society for Information Science and Technology: JASIST, 63(3):512–527, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Colazzo:2002:TTR

[CSA^{+}02] Dario Colazzo, Carlo Sartiani, Antonio Albano, Paolo Manghi, Giorgio Ghelli, Luca Lini, and Michele Paoli. A typed text retrieval query language for XML documents. Journal of the American Society for Information Science and Technology: JASIST, 53(6):467–488, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chau:2007:RBS

[CSCC07] Michael Chau, Boby Shiu, Ivy Chan, and Hsinchun Chen. Redips: Backlink search and analysis on the Web for business intelligence analysis. Journal of the American Society for Information Science and Technology: JASIST, 58(3): 351–365, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2003:CTC

[CSL03] Blaise Cronin, Debora Shaw, and Kathryn La Barre. A cast of thousands: Coauthorship and subauthorship collaboration in the 20th century as manifested in the scholarly journal literature of psychology and philosophy. Journal of the American Society for Information Science and Technology: JASIST, 54(9):855–871, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cronin:2004:VLV

[CSL04] Blaise Cronin, Debora Shaw, and Kathryn La Barre. Visible, less visible, and invisible work: Patterns of collaboration in 20th century chemistry. Journal of the American Society for Information Science and Technology: JASIST, 55(2):160– 168, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Collins-Thompson:2005:PRD

[CTC05] Kevyn Collins-Thompson and Jamie Callan. Predicting reading difficulty with statistical language models. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1448–1462, November 2005. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cyr:2004:LWD

[CTS04] Dianne Cyr and Haizley Trevor-Smith. Localization of Web design: an empirical comparison of German, Japanese, and United States Web site characteristics. Journal of the American Society for Information Science and Technology: JA-SIST, 55(13):1199–1208, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chan:2005:ENE

[CTZ05] H. C. Chan, H. H. Teo, and X. H. Zeng. An evaluation of novice end-user computing performance: Data modeling, query writing, and comprehension. Journal of the American Society for Information Science and Technology: JASIST, 56(8):843–853, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cui:2010:CEP

[Cui10] Hong Cui. Competency evaluation of plant character ontologies against domain literature. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1144–1165, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cui:2012:CFG

[Cui12] Hong Cui. CharaParser for fine-grained semantic annotation of organism morphological descriptions. Journal of the American Society for Information Science and Technology: JASIST, 63(4):738–754, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cullen:2001:BRB

[Cul01] John Cullen. Book review: Knowledge management: Classic and contemporary works, edited by Daryl Morey, Mark Maybury, and Bhavani Thuraisingham. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1190–1191, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cullen:2003:BR

[Cul03] John Cullen. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 54 (6):591–592, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cullen:2004:BR

[Cul04] John Cullen. Book review. Journal of the American Society for Information Science and Technology: JASIST, 55(1):88– 89, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cullen:2006:BRB

[Cul06] John Cullen. Book review: Democratizing innovation. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1711–1712, October 2006. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cushing:2013:SSM

[Cus13] Amber L. Cushing. “it’s stuff that speaks to me”: Exploring the characteristics of digital possessions. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1723–1734, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Costas:2012:ART

[CvL12] Rodrigo Costas and Thed N. van Leeuwen. Approaching the “reward triangle”: General analysis of the presence of funding acknowledgments and “peer interactive communication” in scientific publications. Journal of the American Society for Information Science and Technology: JASIST, 63 (8):1647–1661, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Costas:2010:BCA

[CvLB10] Rodrigo Costas, Thed N. van Leeuwen, and María Bordons. A bibliometric classificatory approach for the study and assessment of research performance at the individual level: the effects of age on productivity and impact. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1564–1581, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Costas:2012:RPI

[CvLB12] Rodrigo Costas, Thed N. van Leeuwen, and María Bordons. Referencing patterns of individual researchers: Do top scientists rely on more extensive information sources? Journal of the American Society for Information Science and Technology: JASIST, 63(12):2433–2450, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Costas:2010:SLS

[CvLvR10] Rodrigo Costas, Thed N. van Leeuwen, and Anthony F. J. van Raan. Is scientific literature subject to a ‘Sell-By-Date’? A general methodology to analyze the ‘durability’ of scientific documents. Journal of the American Society for Information Science and Technology: JASIST, 61(2):329–339, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cannane:2001:GPC

[CW01] Adam Cannane and Hugh E. Williams. General-purpose compression for efficient retrieval. Journal of the American Society for Information Science and Technology: JASIST, 52(5):430–437, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Comeau:2004:NWI

[CW04] Donald C. Comeau and W. John Wilbur. Non-word identification or spell checking without a dictionary. Journal of the American Society for Information Science and Technology: JASIST, 55(2):169–177, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2001:RWM

[CWZ^{+}01] Zheng Chen, Liu Wenyin, Feng Zhang, Mingjing Li, and Hongjiang Zhang. Research: Web mining for Web image retrieval. Journal of the American Society for Information Science and Technology: JASIST, 52(10):831–839, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chau:2010:EUS

[CWZ^{+}10] Michael Chau, Cho Hung Wong, Yilu Zhou, Jialun Qin, and Hsinchun Chen. Evaluating the use of search engine development tools in IT education. Journal of the American Society for Information Science and Technology: JASIST, 61(2): 288–299, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chua:2008:STM

[CY08] Alton Y. K. Chua and Christopher C. Yang. The shift towards multi-disciplinarity in information science. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2156–2170, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chung:2004:ISB

[CZH^{+}04] Wingyan Chung, Yiwen Zhang, Zan Huang, Gang Wang, Thian-Huat Ong, and Hsinchun Chen. Internet searching and browsing in a multilingual world: an experiment on the Chinese Business Intelligence Portal (CBizPort). Journal of the American Society for Information Science and Technology: JASIST, 55(9):818–831, July 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2006:SDR

[CZT06] Liang Chen, Jia Zeng, and Naoyuki Tokuda. A “stereo” document representation for textual information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 57(6):768–774, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Desouza:2005:BCM

[DA05] Kevin C. Desouza and Yukika Awazu. Brief communication: Maintaining knowledge management systems: a strategic imperative. Journal of the American Society for Information Science and Technology: JASIST, 56(7):765–768, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DElia:2007:IYU

[DAB^{+}07] George D’Elia, June Abbas, Kay Bishop, Donald Jacobs, and Eleanor Jo Rodger. The impact of youth’s use of the Internet on their use of the public library. Journal of the American Society for Information Science and Technology: JA-SIST, 58(14):2180–2196, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Daizadeh:2006:EIM

[Dai06] Iraj Daizadeh. An example of information management in biology: Qualitative data economizing theory applied to the Human Genome Project databases. Journal of the American Society for Information Science and Technology: JASIST, 57(2):244–250, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Danell:2011:CQS

[Dan11] Rickard Danell. Can the quality of scientific work be predicted using information on the author’s track record? Journal of the American Society for Information Science and Technology: JASIST, 62(1):50–60, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Carrillo-de-Albornoz:2013:EBM

[dAP13] Jorge Carrillo de Albornoz and Laura Plaza. An emotionbased model of negation, intensifiers, and modality for polarity and intensity classification. Journal of the American Society for Information Science and Technology: JASIST, 64 (8):1618–1633, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Duwairi:2009:FRT

[DARK09] Rehab Duwairi, Mohammad Nayef Al-Refai, and Natheer Khasawneh. Feature reduction techniques for Arabic text categorization. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2347–2352, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davenport:2002:MKM

[Dav02] Elisabeth Davenport. Mundane knowledge management and microlevel organizational learning: an ethological approach. Journal of the American Society for Information Science and Technology: JASIST, 53(12):1038–1046, ???? 2002. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davis:2004:ISB

[Dav04a] Philip M. Davis. Information-seeking behavior of chemists: a transaction log analysis of referral URLs. Journal of the American Society for Information Science and Technology: JASIST, 55(4):326–332, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davis:2004:LET

[Dav04b] Philip M. Davis. Letters to the Editor: Transfer from print to electronic serials. Journal of the American Society for Information Science and Technology: JASIST, 55(4):369– 370, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davis:2006:PSU

[Dav06] Lynne Davis. Perspectives on search user interfaces: Best practices and future visions: Designing a search user interface for a digital library. Journal of the American Society for Information Science and Technology: JASIST, 57(6):788– 791, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davis:2008:BCE

[Dav08] Philip M. Davis. Brief communications: Eigenfactor: Does the principle of repeated improvement result in better estimates than raw citation counts? Journal of the American Society for Information Science and Technology: JASIST, 59(13):2186–2188, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davenport:2009:BRB

[Dav09a] Elisabeth Davenport. Book review: Everyday Information Practices: A Social Phenomenological Perspective. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2590–2591, December 2009. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davis:2009:ACO

[Dav09b] Philip M. Davis. Author-choice open-access publishing in the biological and medical literature: a citation analysis. Journal of the American Society for Information Science and Technology: JASIST, 60(1):3–8, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davis:2011:BRB

[Dav11] Charles H. Davis. Book review: The Information: A History, a Theory, a Flood. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2543– 2545, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Desouza:2006:FGC

[DAW06] Kevin C. Desouza, Yukika Awazu, and Yun Wan. Factors governing the consumption of explicit knowledge. Journal of the American Society for Information Science and Technology: JASIST, 57(1):36–43, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Day:2001:RTR

[Day01] Ronald E. Day. Research: Totality and representation: a history of knowledge management through European documentation, critical modernity, and post-Fordism. Journal of the American Society for Information Science and Technology: JASIST, 52(9):725–735, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Day:2002:BRB

[Day02a] Ron Day. Book review: Cyber-Marx: Cycles and circuits of struggle in high-technology capitalism, by Nick Dyer-Witheford. Journal of the American Society for Information Science and Technology: JASIST, 53(1):60–61, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Day:2002:SCV

[Day02b] Ronald E. Day. Social capital, value, and measure: Antonio Negri’s challenge to capitalism. Journal of the American Society for Information Science and Technology: JASIST, 53(12):1074–1082, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Day:2003:LER

[Day03] Ron Day. Letters to the Editor: Reaction to a book review. Journal of the American Society for Information Science and Technology: JASIST, 54(1):93, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Day:2005:CIK

[Day05] Ronald E. Day. Clearing up “implicit knowledge”: Implications for Knowledge Management, information science, psychology, and social epistemology. Journal of the American Society for Information Science and Technology: JASIST, 56(6):630–635, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Day:2007:KCS

[Day07] Ronald E. Day. Kling and the “critical”: Social informatics and critical informatics. Journal of the American Society for Information Science and Technology: JASIST, 58(4):575– 582, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Day:2008:WR

[Day08] Ronald E. Day. Works and representation. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1644–1652, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Day:2010:LE

[Day10] Ronald E. Day. Letter to the Editor. Journal of the American Society for Information Science and Technology: JA-SIST, 61(1):213, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Day:2011:DUR

[Day11] Ronald E. Day. Death of the user: Reconceptualizing subjects, objects, and their relations. Journal of the American Society for Information Science and Technology: JASIST, 62(1):78–88, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dennis:2002:WSP

[DBM02] Simon Dennis, Peter Bruza, and Robert McArthur. Web searching: a process-oriented experimental study of three interactive search paradigms. Journal of the American Society for Information Science and Technology: JASIST, 53 (2):120–133, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dannenberg:2007:CES

[DBP^{+}07] Roger B. Dannenberg, William P. Birmingham, Bryan Pardo, Ning Hu, Colin Meek, and George Tzanetakis. A comparative evaluation of search techniques for query-by-humming using the MUSART testbed. Journal of the American Society for Information Science and Technology: JA-SIST, 58(5):687–701, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dufour:2011:UHW

[DBT11] Christine Dufour, Joan C. Bartlett, and Elaine G. Toms. Understanding how webcasts are used as sources of information. Journal of the American Society for Information Science and Technology: JASIST, 62(2):343–362, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

deBoer:2010:EHT

[dBvSW10] Viktor de Boer, Maarten van Someren, and Bob J. Wielinga. Extracting historical time periods from the Web. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1888–1908, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davenport:2001:BCW

[DC01a] Elisabeth Davenport and Blaise Cronin. Brief communication: Who dunnit? Metatags and hyperauthorship. Journal of the American Society for Information Science and Technology: JASIST, 52(9):770–773, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davis:2001:REW

[DC01b] Philip M. Davis and Suzanne A. Cohen. Research: The effect of the Web on undergraduate citation behavior 1996–1999. Journal of the American Society for Information Science and Technology: JASIST, 52(4):309–314, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Doerfel:2009:SNC

[DC09] Marya L. Doerfel and Stacey L. Connaughton. Semantic networks and competition: Election year winners and losers in U.S. televised presidential debates, 1960–2004. Journal of the American Society for Information Science and Technology: JASIST, 60(1):201–218, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Carvalho:2012:LLP

[dCCRdM^{+}12] André L. da Costa Carvalho, Cristian Rossi, Edleno S. de Moura, Altigran S. da Silva, and David Fernandes. LePrEF: Learn to precompute evidence fusion for efficient query evaluation. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1383–1397, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

deCampos:2003:IRF

[dCFLH03] Luis M. de Campos, Juan M. Fernández-Luna, and Juan F. Huete. Implementing relevance feedback in the Bayesian Network Retrieval model. Journal of the American Society for Information Science and Technology: JASIST, 54 (4):302–313, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Devaul:2011:CAA

[DDO11] Holly Devaul, Anne R. Diekema, and Jonathan Ostwald. Computer-assisted assignment of educational standards using natural language processing. Journal of the American Society for Information Science and Technology: JASIST, 62(2):395–405, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DeVisscher:2010:IMS

[De 10] Alex De Visscher. An index to measure a scientist’s specific impact. Journal of the American Society for Information Science and Technology: JASIST, 61(2):319–328, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DeBellis:2011:BRB

[De 11a] Nicola De Bellis. Book review: The Publish or Perish Book. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1426–1429, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DeVisscher:2011:BCW

[De 11b] Alex De Visscher. Brief communications: What does the g-index really measure? Journal of the American Society for Information Science and Technology: JASIST, 62(11): 2290–2293, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See response [De 13] and reply [Egg12].

DeVisscher:2012:LET

[De 12] Alex De Visscher. Letter to the Editor: The thermodynamicsbibliometrics consilience and the meaning of h-type indices — reply. Journal of the American Society for Information Science and Technology: JASIST, 63(3):630–631, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DeVisscher:2013:LER

[De 13] Alex De Visscher. Letter to the Editor: Response to “Remarks on the paper by A. De Visscher, ‘What does the gindex really measure?’ ”. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1960– 1962, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DeNicola:2009:BRB

[DeN09] Lane DeNicola. Book review: A Disciplinary Blueprint for the Assessment of Information Literacy. Journal of the American Society for Information Science and Technology: JASIST, 60(3):639–641, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Desouza:2007:BRB

[Des07] Kevin C. Desouza. Book review: Information politics on the Web. Journal of the American Society for Information Science and Technology: JASIST, 58(4):608–609, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Desouza:2008:BRBa

[Des08a] Kevin C. Desouza. Book review: Advancing knowledge and the knowledge economy. Journal of the American Society for Information Science and Technology: JASIST, 59(2):331– 333, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Desouza:2008:BRBb

[Des08b] Kevin C. Desouza. Book review: Understanding knowledge as a commons: From theory to practice. Journal of the American Society for Information Science and Technology: JASIST, 59(2):333–334, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Detlor:2011:BRB

[Det11] Brian Detlor. Book review: Dancing With Digital Natives: Staying in Step with the Generation That’s Transforming the Way Business Is Done. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2297– 2298, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Doran:2012:BCC

[DG12] Derek Doran and Swapna S. Gokhale. Brief communication: A classification framework for web robots. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2549–2554, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DAngelo:2011:HAA

[DGA11] Ciriaco Andrea D’Angelo, Cristiano Giuffrida, and Giovanni Abramo. A heuristic approach to author name disambiguation in bibliometrics databases for large-scale research assessments. Journal of the American Society for Information Science and Technology: JASIST, 62(2):257–269, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dominich:2004:EBI

[DGKS04] Sándor Dominich, Júlia Góth, Tamás Kiezer, and Zoltán Szlávik. An entropy-based interpretation of retrieval status value-based retrieval, and its application to the computation of term and query discrimination value. Journal of the American Society for Information Science and Technology: JASIST, 55(7):613–627, May 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Daoud:2013:MGT

[DH13] Mariam Daoud and Jimmy Xiangji Huang. Modeling geographic, temporal, and proximity contexts for improving geotemporal search. Journal of the American Society for Information Science and Technology: JASIST, 64(1):190–212, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Du:2010:CLM

[DHKW10] Helen Du, Jin-Xing Hao, Ron Kwok, and Christian Wagner. Can a lean medium enhance large-group communication? Examining the impact of interactive mobile learning. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2122–2137, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dow:2013:DAO

[DHW13] Kevin E. Dow, Gary Hackbarth, and Jeffrey Wong. Data architectures for an organizational memory information system. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1345–1356, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dianu:2002:BRB

[Dia02] An-Chi Hoh Dianu. Book review: The social life of information. Journal of the American Society for Information Science and Technology: JASIST, 53(4):320, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dillon:2002:EIA

[Dil02] Andrew Dillon. Editorial: Information architecture in JA-SIST: Just where did we come from? Journal of the American Society for Information Science and Technology: JA-SIST, 53(10):821–823, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dimitrova:2009:BRB

[Dim09] Dimitrina Dimitrova. Book review: Scientific Collaboration on the Internet. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1721–1723, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ding:2005:PML

[Din05] Chris H. Q. Ding. A probabilistic model for Latent Semantic Indexing. Journal of the American Society for Information Science and Technology: JASIST, 56(6):597–608, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dincer:2007:SPC

[Din07] Bekir Taner Dinçer. Statistical principal components analysis for retrieval experiments. Journal of the American Society for Information Science and Technology: JASIST, 58 (4):560–574, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ding:2011:AWP

[Din11a] Ying Ding. Applying weighted PageRank to author citation networks. Journal of the American Society for Information Science and Technology: JASIST, 62(2):236–245, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ding:2011:TBP

[Din11b] Ying Ding. Topic-based PageRank on author cocitation networks. Journal of the American Society for Information Science and Technology: JASIST, 62(3):449–466, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ding:2012:BRI

[Din12] Ying Ding. Book reviews: Information Concepts, Retrieval, and Services [Series]. Journal of the American Society for Information Science and Technology: JASIST, 63(2):421–424, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ding:2010:UTO

[DJF^{+}10] Ying Ding, Elin K. Jacob, Michael Fried, Ioan Toma, Erjia Yan, Schubert Foo, and Staša Milojević. Upper tag ontology for integrating social tagging data. Journal of the American Society for Information Science and Technology: JASIST, 61(3):505–521, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Detlor:2011:LOI

[DJW^{+}11] Brian Detlor, Heidi Julien, Rebekah Willson, Alexander Serenko, and Maegen Lavallee. Learning outcomes of information literacy instruction at business schools. Journal of the American Society for Information Science and Technology: JASIST, 62(3):572–585, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DElia:2002:IIP

[DJWR02] George D’Elia, Corinne Jörgensen, Joseph Woelfel, and Eleanor Jo Rodger. The impact of the Internet on public library use: an analysis of the current consumer market for library and Internet services. Journal of the American Society for Information Science and Technology: JASIST, 53 (10):802–820, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ding:2009:PST

[DJZ^{+}09] Ying Ding, Elin K. Jacob, Zhixiong Zhang, Schubert Foo, Erjia Yan, Nicolas L. George, and Lijiang Guo. Perspectives on social tagging. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2388–2401, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dominich:2007:MTA

[DK07] Sándor Dominich and Tamás Kiezer. A measure theoretic approach to information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1108–1122, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dang:2010:NCD

[DLA^{+}10] E. K. F. Dang, R. W. P. Luk, J. Allan, K. S. Ho, S. C. F. Chan, K. F. L. Chung, and D. L. Lee. A new contextdependent term weight computed by boost and discount using relevance information. Journal of the American Society for Information Science and Technology: JASIST, 61(12):

2514–2530, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dang:2008:NMC

[DLH^{+}08] E. K. F. Dang, R. W. P. Luk, K. S. Ho, S. C. F. Chan, and D. L. Lee. A new measure of clustering effectiveness: Algorithms and experimental studies. Journal of the American Society for Information Science and Technology: JASIST, 59(3):390–406, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DeAndres:2009:EFF

[DLM09] Javier De Andrés, Pedro Lorca, and Ana B. Martínez. Economic and financial factors for the adoption and visibility effects of Web accessibility: the case of European banks. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1769–1780, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Diaz:2012:MLA

[DLVÁ12] Noa P. Cruz Díaz, Manuel J. Maña López, Jacinto Mata Vázquez, and Victoria Pachón Álvarez. A machine-learning approach to negation and speculation detection in clinical texts. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1398–1410, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

deMoya-Anegon:2007:VMS

[dMAVQCR^{+}07] Scimago Group: Félix de Moya-Anegón, Benjamín Vargas-Quesada, Zaida Chinchilla-Rodríguez, Elena Corera-Álvarez, Francisco J. Munoz-Fernández, and Victor Herrero-Solana. Visualizing the marrow of science. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2167–2179, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Camacho-Minano:2009:MNR

[dMCMNN09] María del Mar Camacho-Miñano and Manuel Núñez-Nickel. The multilayered nature of reference selection. Journal of the American Society for Information Science and Technology:

JASIST, 60(4):754–777, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

deMoura:2010:USI

[dMFRN^{+}10] Edleno S. de Moura, David Fernandes, Berthier Ribeiro-Neto, Altigran S. da Silva, and Marcos André Gonçalves. Using structural information to improve search in Web collections. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2503–2513, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Diaz:2002:ATC

[DML02] Irene Diaz, Jorge Morato, and Juan Lloréns. An algorithm for term conflation based on tree structures. Journal of the American Society for Information Science and Technology: JASIST, 53(3):199–208, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dubnov:2006:SAA

[DMR06] Shlomo Dubnov, Stephen McAdams, and Roger Reynolds. Structural and affective aspects of music from statistical audio signal analysis. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1526– 1536, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dore:2001:HAP

[DO01] Jean-Christophe Doré and Tiiu Ojasoo. How to analyze publication time trends by correspondence factor analysis: Analysis of publications by 48 countries in 18 disciplines over 12 years. Journal of the American Society for Information Science and Technology: JASIST, 52(9):763–769, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Doty:2013:BRB

[Dot13] Philip Doty. Book review: How information matters: Networks and public policy innovation — By Kathleen Hale. Journal of the American Society for Information Science and Technology: JASIST, 64(3):647–649, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dovey:2004:OOP

[Dov04] Matthew J. Dovey. Overview of the OMRAS project: Online music retrieval and searching. Journal of the American Society for Information Science and Technology: JASIST, 55(12):1100–1107, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Downie:2004:IOS

[Dow04] J. Stephen Downie. Introduction and overview: A sample of music information retrieval approaches. Journal of the American Society for Information Science and Technology: JASIST, 55(12):1033–1036, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davis:2006:EIC

[DP06] Philip M. Davis and Jason S. Price. eJournal interface can influence usage statistics: Implications for libraries, publishers, and Project COUNTER. Journal of the American Society for Information Science and Technology: JASIST, 57 (9):1243–1248, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ding:2007:UMP

[DP07] Chen Ding and Jagdish C. Patra. User modeling for personalized Web search with self-organizing map. Journal of the American Society for Information Science and Technology: JASIST, 58(4):494–507, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DiIorio:2011:SWA

[DPV11] Angelo Di Iorio, Silvio Peroni, and Fabio Vitali. A Semantic Web approach to everyday overlapping markup. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1696–1716, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dragulanescu:2002:EIS

[Dra02] Nicolae Dragulanescu. Emerging information society and history of information science in Romania. Journal of the American Society for Information Science and Technology: JASIST, 53(1):41–46, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Drabenstott:2003:DNE

[Dra03] Karen M. Drabenstott. Do nondomain experts enlist the strategies of domain experts? Journal of the American Society for Information Science and Technology: JASIST, 54 (9):836–854, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Diaz:2004:IPT

[DRM^{+}04] Irene Díaz, José Ranilla, Elena Montañes, Javier Fernández, and Elías F. Combarro. Improving performance of text categorization by combining filtering and support vector machines. Journal of the American Society for Information Science and Technology: JASIST, 55(7):579–592, May 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davenport:2008:GAS

[DRW08] Dee Dee Davenport, Jennifer Richey, and Lynn Westbrook. E-government access to social service information: State web resources for domestic violence survivors. Journal of the American Society for Information Science and Technology: JASIST, 59(6):903–915, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Davis:2003:ILA

[DS03] Philip M. Davis and Leah R. Solla. An IP-level analysis of usage statistics for electronic journals in chemistry: Making inferences about user behavior. Journal of the American Society for Information Science and Technology: JASIST, 54(11):1062–1068, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Debole:2005:ARH

[DS05a] Franca Debole and Fabrizio Sebastiani. An analysis of the relative hardness of Reuters-21578 subsets. Journal of the American Society for Information Science and Technology: JASIST, 56(6):584–596, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dominich:2005:PII

[DS05b] Sándor Dominich and Adrienn Skrop. PageRank and Interaction Information Retrieval. Journal of the American Society for Information Science and Technology: JASIST, 56(1): 63–69, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dolamic:2009:ISS

[DS09] Ljiljana Dolamic and Jacques Savoy. Indexing and searching strategies for the Russian language. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2540–2547, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dolamic:2010:BCW

[DS10a] Ljiljana Dolamic and Jacques Savoy. Brief communication: When stopword lists make the difference. Journal of the American Society for Information Science and Technology: JASIST, 61(1):200–203, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dolamic:2010:REM

[DS10b] Ljiljana Dolamic and Jacques Savoy. Retrieval effectiveness of machine translated queries. Journal of the American Society for Information Science and Technology: JASIST, 61 (11):2266–2273, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Du:2011:TWS

[DS11] Jia Tina Du and Amanda Spink. Toward a web search model: Integrating multitasking, cognitive coordination, and cognitive shifts. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1446–1472, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dunne:2012:RUS

[DSG^{+}12] Cody Dunne, Ben Shneiderman, Robert Gove, Judith Klavans, and Bonnie Dorr. Rapid understanding of scientific paper collections: Integrating statistics, text analytics, and visualization. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2351–2369, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Didegah:2013:DRC

[DT13] Fereshteh Didegah and Mike Thelwall. Determinants of research citation impact in nanoscience and nanotechnology. Journal of the American Society for Information Science and Technology: JASIST, 64(5):1055–1064, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Duff:2011:RTT

[Duf11] Alistair S. Duff. The Rawls-Tawney theorem and the digital divide in postindustrial society. Journal of the American Society for Information Science and Technology: JASIST, 62(3):604–612, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dumova:2001:BRB

[Dum01a] Tatyana Dumova. Book review: Community informatics: Enabling communities with information and communication technologies, Edited by Michael Gurstein. Journal of the American Society for Information Science and Technology: JASIST, 52(11):981–983, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dumova:2001:BR

[Dum01b] Tatyana Dumova. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 52(1):83–84, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Duwairi:2006:MLA

[Duw06] Rehab M. Duwairi. Machine learning for Arabic text categorization. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1005–1010, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

deVries:2001:CIM

[dV01] Arjen P. de Vries. Content independence in multimedia databases. Journal of the American Society for Information Science and Technology: JASIST, 52(11):954–960, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DAmbra:2004:UWW

[DW04] John D’Ambra and Concepción S. Wilson. Use of the World Wide Web for international travel: Integrating the construct of uncertainty in information seeking and the tasktechnology fit (TTF) model. Journal of the American Society for Information Science and Technology: JASIST, 55 (8):731–742, June 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Daranyi:2013:DCD

[DW13] Sándor Darányi and Peter Wittek. Demonstrating conceptual dynamics in an evolving text collection. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2564–2572, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

DAmbra:2013:ATT

[DWA13] John D’Ambra, Concepción S. Wilson, and Shahriar Akter. Application of the task-technology fit model to structure and evaluate the adoption of E-books by Academics. Journal of the American Society for Information Science and Technology: JASIST, 64(1):48–64, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ding:2009:PRA

[DYFC09] Ying Ding, Erjia Yan, Arthur Frazho, and James Caverlee. PageRank for ranking authors in co-citation networks. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2229–2243, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Dang:2009:ALM

[DZC^{+}09] Yan Dang, Yulei Zhang, Hsinchun Chen, Paul Jen-Hwa Hu, Susan A. Brown, and Cathy Larson. Arizona Literature Mapper: an integrated approach to monitor and analyze global bioterrorism research literature. Journal of the American Society for Information Science and Technology: JA-SIST, 60(7):1466–1485, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Edwards:2005:CIM

[EAPL05] Ron Edwards, Ajith Abraham, and Sonja Petrovic-Lazarevic. Computational intelligence to model the export behavior of multinational corporation subsidiaries in Malaysia. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1177–1186, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Eastman:2002:HMB

[Eas02] Caroline M. Eastman. 30,000 hits may be better than 300: Precision anomalies in Internet searches. Journal of the American Society for Information Science and Technology: JASIST, 53(11):879–882, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Espadas:2008:WSV

[ECP08] Javier Espadas, Coral Calero, and Mario Piattini. Web site visibility evaluation. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1727– 1742, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Edwards:2007:BRB

[Edw07] Phillip M. Edwards. Book review: Blogging and RSS: a librarian’s guide. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2160–2161, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Engqvist:2010:LEN

[EF10] Leif Engqvist and Joachim G. Frommen. Letter to the Editor: New insights into the relationship between the h-index and self-citations? Journal of the American Society for Information Science and Technology: JASIST, 61(7):1514– 1516, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Efron:2005:EBM

[Efr05] Miles Efron. Eigenvalue-based model selection during latent semantic indexing. Journal of the American Society for Information Science and Technology: JASIST, 56(9):969–988, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Efron:2008:SMS

[Efr08] Miles Efron. Shannon meets Shortz: A probabilistic model of crossword puzzle difficulty. Journal of the American Society for Information Science and Technology: JASIST, 59 (6):875–886, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Efron:2010:LTS

[Efr10] Miles Efron. Linear time series models for term weighting in information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1299– 1312, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Efron:2011:AIS

[Efr11] Miles Efron. Advances in information science: Information search and retrieval in microblogs. Journal of the American Society for Information Science and Technology: JASIST, 62(6):996–1008, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Efron:2012:BRB

[Efr12] Miles Efron. Book reviews: Search User Interfaces — By Marti A. Hearst. Interactive Information Seeking, Behaviour and Retrieval — Edited by Ian Ruthven and Diane Kelly. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2555–2556, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2012:BCA

[EG12] Leo Egghe and Raf Guns. Brief communication: Applications of the generalized law of Benford to informetric data. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1662–1665, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2001:NAR

[Egg01] L. Egghe. A noninformetric analysis of the relationship between citation age and journal productivity. Journal of the American Society for Information Science and Technology: JASIST, 52(5):371–377, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2002:BRB

[Egg02a] L. Egghe. Book review: Mathematical foundations of information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 53(5):409–410, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2002:SCV

[Egg02b] L. Egghe. Sampling and concentration values of incomplete bibliographies. Journal of the American Society for Information Science and Technology: JASIST, 53(4):271–281, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2003:TTT

[Egg03] Leo Egghe. Type/ token-taken informetrics. Journal of the American Society for Information Science and Technology: JASIST, 54(7):603–610, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2005:PPL

[Egg05a] L. Egghe. The power of power laws and an interpretation of Lotkaian informetric systems as self-similar fractals. Journal of the American Society for Information Science and Technology: JASIST, 56(7):669–675, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2005:RBC

[Egg05b] L Egghe. Relations between the continuous and the discrete Lotka power function. Journal of the American Society for Information Science and Technology: JASIST, 56(7):664– 668, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2005:ZLC

[Egg05c] L. Egghe. Zipfian and Lotkaian continuous concentration theory. Journal of the American Society for Information Science and Technology: JASIST, 56(9):935–945, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2006:POV

[Egg06] L. Egghe. Properties of the n-overlap vector and n-overlap similarity theory. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1165–1177, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2007:BRB

[Egg07a] Leo Egghe. Book review: Uncertainty and information: Foundations of generalized information theory. Journal of the American Society for Information Science and Technology: JASIST, 58(5):756–758, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2007:BCD

[Egg07b] Leo Egghe. Brief communication: Dynamic h-index: The Hirsch index in function of time. Journal of the American Society for Information Science and Technology: JASIST, 58(3):452–454, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2007:UHL

[Egg07c] Leo Egghe. Untangling Herdan’s Law and Heaps’ Law: Mathematical and informetric arguments. Journal of the American Society for Information Science and Technology: JASIST, 58(5):702–709, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2008:ITI

[Egg08a] L. Egghe. The influence of transformations on the h-index and the g-index. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1304–1312, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2008:MTI

[Egg08b] Leo Egghe. Mathematical theory of the h- and g-index in case of fractional counting of authorship. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1608–1616, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2008:MSF

[Egg08c] Leo Egghe. A model for the size-frequency function of coauthor pairs. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2133–2137, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2009:NRB

[Egg09a] Leo Egghe. New relations between similarity measures for vectors based on vector norms. Journal of the American Society for Information Science and Technology: JASIST, 60(2):232–239, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2009:RHI

[Egg09b] Leo Egghe. A rationale for the Hirsch-index rank-order distribution and a comparison with the impact factor rankorder distribution. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2142– 2144, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2010:BCN

[Egg10a] Leo Egghe. Brief communication: A new short proof of Naranan’s theorem, explaining Lotka’s Law and Zipf’s Law. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2581–2583, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2010:BCR

[Egg10b] Leo Egghe. Brief communication: On the relation between the association strength and other similarity measures. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1502–1504, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2010:GPS

[Egg10c] Leo Egghe. Good properties of similarity measures and their complementarity. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2151–2160, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2010:IAD

[Egg10d] Leo Egghe. Influence of adding or deleting items and sources on the h-index. Journal of the American Society for Information Science and Technology: JASIST, 61(2):370–373, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2012:BCR

[Egg12] L. Egghe. Brief communication: Remarks on the paper by A. De Visscher, “What does the g-index really measure?”. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2118–2121, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [De 11b] and reply [De 13].

Egghe:2013:TTC

[Egg13a] L. Egghe. Theory of the topical coverage of multiple databases. Journal of the American Society for Information Science and Technology: JASIST, 64(1):126–131, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2013:LEN

[Egg13b] Leo Egghe. Letter to the Editor: Note on a possible decomposition of the h-index. Journal of the American Society for Information Science and Technology: JASIST, 64(4): 871, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2011:TUN

[EGR11] Leo Egghe, Raf Guns, and Ronald Rousseau. Thoughts on uncitedness: Nobel Laureates and Fields Medalists as case studies. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1637–1644, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2012:EE

[EGRL12] Leo Egghe, Raf Guns, Ronald Rousseau, and K. U. Leuven. Erratum. Journal of the American Society for Information Science and Technology: JASIST, 63(2):429, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Eschenfelder:2005:WPD

[EHD05] Kristin R. Eschenfelder, Robert Glenn Howard, and Anuj C. Desai. Who posts DeCSS and why?: a content analysis of Web sites posting DVD circumvention software. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1405–1418, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ehikhamenor:2002:SEF

[Ehi02] Fabian A. Ehikhamenor. Socio-economic factors in the application of information and communication technologies in Nigerian print media. Journal of the American Society for Information Science and Technology: JASIST, 53(7):602– 611, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ekbia:2009:DAQ

[Ekb09] Hamid R. Ekbia. Digital artifacts as quasi-objects: Qualification, mediation, and materiality. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2554–2566, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ekbia:2012:BRB

[Ekb12] Hamid R. Ekbia. Book reviews: Alone Together: Why We Expect More from Technology and Less from Each Other. By Sherry Turkle. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1897–1898, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2009:RBP

[EL09] Leo Egghe and Loet Leydesdorff. The relation between Pearson’s correlation coefficient r and Salton’s cosine measure. Journal of the American Society for Information Science and Technology: JASIST, 60(5):1027–1036, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2008:FPR

[ELR08] Leo Egghe, Liming Liang, and Ronald Rousseau. Fundamental properties of rhythm sequences. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1469–1478, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2009:RBI

[ELR09] Leo Egghe, Liming Liang, and Ronald Rousseau. A relation between h-index and impact factor in the power-law model. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2362–2365, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Elsweiler:2009:BRB

[Els09] David Elsweiler. Book review: Keeping Found Things Found: The Study and Practice of Personal Information Management. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1725–1727, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Evain:2007:PMM

[EM07] Jean-Pierre Evain and José M. Martínez. Perspectives on MPEG-7: Metadata for multimedia: TV-Anytime Phase 1 and MPEG-7. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1367–1373, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2001:BRBa

[Enn01a] Lisa A. Ennis. Book review: Creating Web-accessible databases: Case studies for libraries, museums, and other nonprofits, by Julie M. Still. Journal of the American Society for Information Science and Technology: JASIST, 52 (11):980–981, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2001:BRBb

[Enn01b] Lisa A. Ennis. Book review: Peer-to-peer: Harnessing the benefits of a disruptive technology, edited by Andy Oram. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1191–1192, ???? 2001. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2002:BRBb

[Enn02a] Lisa A. Ennis. Book review: The digital divide: Facing a crisis or creating a myth? Journal of the American Society for Information Science and Technology: JASIST, 53 (7):612–613, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2002:BRBa

[Enn02b] Lisa A. Ennis. Book review: The map library in the new millennium. Journal of the American Society for Information Science and Technology: JASIST, 53(3):250–251, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2003:BRBa

[Enn03a] Lisa Ennis. Book review: Digital futures: Strategies for the information age. Journal of the American Society for Information Science and Technology: JASIST, 54(9):908–909, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2003:BRa

[Enn03b] Lisa A. Ennis. Book review. Journal of the American Society for Information Science and Technology: JASIST, 54(1):91– 92, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2003:BRBb

[Enn03c] Lisa A. Ennis. Book review: Dark Fiber: Tracking Critical Internet Culture. Journal of the American Society for Information Science and Technology: JASIST, 54(10):981–982, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2003:BRb

[Enn03d] Lisa A. Ennis. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1165–1166, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2004:BRB

[Enn04] Lisa A. Ennis. Book review: The visible librarian: Asserting your value with marketing and advocacy. Journal of the American Society for Information Science and Technology: JASIST, 55(14):1304, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2006:BRBb

[Enn06a] Lisa A. Ennis. Book review: E-government: From vision to implementation. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1280, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2006:BRBa

[Enn06b] Lisa A. Ennis. Book review: The digital age and local studies. Journal of the American Society for Information Science and Technology: JASIST, 57(3):443, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2007:BRBa

[Enn07a] Lisa A. Ennis. Book review: Information ethics: Privacy, property, and power. Journal of the American Society for Information Science and Technology: JASIST, 58(2):302, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2007:BRBb

[Enn07b] Lisa A. Ennis. Book review: The access principle: the case for open access to research and scholarship. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1386, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2008:BRBa

[Enn08a] Lisa A. Ennis. Book review: Technology made simple: an improvement guide for small and medium libraries. Journal of the American Society for Information Science and Technology: JASIST, 59(2):334–335, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ennis:2008:BRBb

[Enn08b] Lisa A. Ennis. Book review: Wired shut: Copyright and the shape of digital culture. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2191– 2192, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2002:DRC

[ER02] L. Egghe and I. K. Ravichandra Rao. Duality revisited: Construction of fractional frequency distributions based on two dual Lotka laws. Journal of the American Society for Information Science and Technology: JASIST, 53(10):789– 801, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2003:MCW

[ER03] Leo Egghe and Ronald Rousseau. A measure for the cohesion of weighted networks. Journal of the American Society for Information Science and Technology: JASIST, 54(3):193– 202, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2008:SDI

[ER08a] L. Egghe and I. K. Ravichandra Rao. Study of different h-indices for groups of authors. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1276–1281, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2008:BCI

[ER08b] Leo Egghe and I. K. Ravichandra Rao. Brief communication: The influence of the broadness of a query of a topic on its h-index: Models and examples of the h-index of n-grams. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1688–1693, August 2008.

CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2012:HIS

[ER12] Leo Egghe and Ronald Rousseau. The Hirsch index of a shifted Lotka function and its relation with the impact factor. Journal of the American Society for Information Science and Technology: JASIST, 63(5):1048–1053, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ercegovac:2006:MVR

[Erc06] Zorana Ercegovac. Multiple-version resources in digital libraries: Towards user-centered displays. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1023–1032, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Elsweiler:2007:TMS

[ERJ07] David Elsweiler, Ian Ruthven, and Christopher Jones. Towards memory supporting personal information management tools. Journal of the American Society for Information Science and Technology: JASIST, 58(7):924–946, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Egghe:2007:TC

[ERR07] Leo Egghe, Ronald Rousseau, and Sandra Rousseau. TOPcurves. Journal of the American Society for Information Science and Technology: JASIST, 58(6):777–785, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Eschenfelder:2010:BRBb

[Esc10a] Kristin R. Eschenfelder. Book review: Memory Practices in the Sciences. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1945–1946, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Eschenfelder:2010:BRBa

[Esc10b] Kristin R. Eschenfelder. Book review: Review of Intellectual Property and Traditional Cultural Expressions in a Digital Environment. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1295–1297, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Elkiss:2008:BME

[ESF^{+}08] Aaron Elkiss, Siwei Shen, Anthony Fader, Güneş Erkan, David States, and Dragomir Radev. Blind men and elephants: What do citation summaries tell us about a research article? Journal of the American Society for Information Science and Technology: JASIST, 59(1):51–62, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Elovici:2006:DTA

[ESK06] Yuval Elovici, Bracha Shapira, and Paul B. Kantor. A decision theoretic approach to combining information filters: an analytical and empirical evaluation. Journal of the American Society for Information Science and Technology: JA-SIST, 57(3):306–320, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Elovici:2010:DAT

[ESL^{+}10] Yuval Elovici, Bracha Shapira, Mark Last, Omer Zaafrany, Menahem Friedman, Moti Schneider, and Abraham Kandel. Detection of access to terror-related Web sites using an Advanced Terror Detection System (ATDS). Journal of the American Society for Information Science and Technology: JASIST, 61(2):405–418, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ess:2010:BRB

[Ess10] Charles Ess. Book review: Handbook of Research on Computer Mediated Communication. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1725–1727, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Efron:2010:QPR

[EW10] Miles Efron and Megan Winget. Query polyrepresentation for ranking retrieval systems without relevance judgments. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1081–1091, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ellis:2002:ISM

[EWF^{+}02] David Ellis, T. D. Wilson, Nigel Ford, Allen Foster, H. M. Lam, R. Burton, and Amanda Spink. Information seeking and mediated searching. Part 5. User–intermediary interaction. Journal of the American Society for Information Science and Technology: JASIST, 53(11):883–893, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Exner:2002:BRB

[Exn02] Frank Exner. Book review: The modern invention of information: Discourse, history, and power. Journal of the American Society for Information Science and Technology: JASIST, 53(11):966, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Epperson:2008:RRR

[EZ08] Terrence W. Epperson and Alan Zemel. Reports, requests, and recipient design: the management of patron queries in online reference chats. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2268– 2283, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fu:2008:HAW

[FAC08] Tianjun Fu, Ahmed Abbasi, and Hsinchun Chen. A hybrid approach to Web forum interactional coherence analysis. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1195–1209, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fu:2010:FCD

[FAC10] Tianjun Fu, Ahmed Abbasi, and Hsinchun Chen. A focused crawler for Dark Web forums. Journal of the American Society for Information Science and Technology: JASIST, 61(6): 1213–1231, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Faherty:2002:BRB

[Fah02] Roisin Faherty. Book review: Strategies for electronic commerce and the Internet. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1187– 1188, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fallis:2008:TEW

[Fal08] Don Fallis. Toward an epistemology of Wikipedia. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1662–1674, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fang:2010:BRB

[Fan10] Jin-Qing Fang. Book review: Network Science — Theory and Application. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1297–1298, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fernandez-Alles:2009:ISH

[FARR09] Mariluz Fernandez-Alles and Antonio Ramos-Rodríguez. Intellectual structure of human resources management research: a bibliometric analysis of the journal Human Resource Management, 1985–2005. Journal of the American Society for Information Science and Technology: JASIST, 60(1):161–175, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fatemi:2007:PMM

[Fat07] Nastaran Fatemi. Perspectives on MPEG-7: Metadata for multimedia: MPEG-7 in practice: Analysis of a television news retrieval application. Journal of the American Society for Information Science and Technology: JASIST, 58 (9):1364–1366, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Faust:2009:PDI

[Fau09] Jurgen Faust. Perspectives on design: Information technologies and creative practices: Positive design. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1887–1894, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2006:FND

[fBWLBC06] Yi fang Brook Wu, Quanzhi Li, Razvan Stefan Bot, and Xin Chen. Finding nuggets in documents: a machine learning approach. Journal of the American Society for Information Science and Technology: JASIST, 57(6):740–752, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fritch:2001:EIR

[FC01] John W. Fritch and Robert L. Cromwell. Evaluating Internet resources: Identity, affiliation, and cognitive authority in a networked world. Journal of the American Society for Information Science and Technology: JASIST, 52(6):499– 507, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Farreus:2012:SCA

[FCjM12] Mireia Farreús, Marta R. Costa-jussà, and Maja Popović Morse. Study and correlation analysis of linguistic, perceptual, and automatic machine translation evaluations. Journal of the American Society for Information Science and Technology: JASIST, 63(1):174–184, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

French:2004:MVA

[FCM04] James C. French, A. C. Chapin, and Worthy N. Martin. Multiple viewpoints as an approach to digital library interfaces. Journal of the American Society for Information Science and Technology: JASIST, 55(10):911–922, August 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Foss:2012:CSR

[FDB^{+}12] Elizabeth Foss, Allison Druin, Robin Brewer, Phillip Lo, Luis Sanchez, Evan Golub, and Hilary Hutchinson. Children’s search roles at home: Implications for designers, researchers, educators, and parents. Journal of the American Society for Information Science and Technology: JASIST, 63(3):558–573, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fisher:2004:IGU

[FDH04] Karen E. Fisher, Joan C. Durrance, and Marian Bouch Hinton. Information grounds and the use of need-based services by immigrants in Queens, New York: a context-based, outcome evaluation approach. Journal of the American Society for Information Science and Technology: JASIST, 55 (8):754–766, June 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Foss:2013:ASR

[FDY^{+}13] Elizabeth Foss, Allison Druin, Jason Yip, Whitney Ford, Evan Golub, and Hilary Hutchinson. Adolescent search roles. Journal of the American Society for Information Science and Technology: JASIST, 64(1):173–189, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Feinberg:2011:CAP

[Fei11] Melanie Feinberg. Compiler to author: a process for designing rhetorically aware document collections. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1784–1796, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fenstermacher:2005:IB

[Fen05] David Fenstermacher. Introduction to bioinformatics. Journal of the American Society for Information Science and Technology: JASIST, 56(5):440–446, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ferguson:2004:BRB

[Fer04] Stephen Ferguson. Book review: JSTOR: A History. Journal of the American Society for Information Science and Technology: JASIST, 55(11):1026–1027, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fricke:2004:IAA

[FF04] Martin Frické and Don Fallis. Indicators of accuracy for answers to ready reference questions on the Internet. Journal of the American Society for Information Science and Technology: JASIST, 55(3):238–245, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ferran-Ferrer:2013:KFT

[FFMPM13] Núria Ferran-Ferrer, Julià Minguillón, and Mario Pérez-Montoro. Key factors in the transfer of information-related competencies between academic, workplace, and daily life contexts. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1112–1121, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fan:2004:EFF

[FFPW04] Weiguo Fan, Edward A. Fox, Praveen Pathak, and Harris Wu. The effects of fitness functions on genetic programmingbased ranking discovery for Web search. Journal of the American Society for Information Science and Technology: JASIST, 55(7):628–636, May 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fitzgerald:2001:RJE

[FG01] Mary Ann Fitzgerald and Chad Galloway. Relevance judging, evaluation, and decision making in virtual libraries: a descriptive study. Journal of the American Society for Information Science and Technology: JASIST, 52(12):989–1010, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fenstermacher:2003:CSM

[FG03] Kurt D. Fenstermacher and Mark Ginsburg. Client-side monitoring for Web mining. Journal of the American Society for Information Science and Technology: JASIST, 54 (7):625–637, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fidel:2013:LEF

[Fid13] Raya Fidel. Letters to the Editor: Factual errors in the review of human information interaction. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2597, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Figueredo:2006:LEN

[Fig06] Eduardo Figueredo. Letter to the Editor: The numerical equivalence between the impact factor of journals and the quality of the articles. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1561, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Freitas-Junior:2006:CDC

[FJRNV^{+}06] Hermes R. Freitas-Junior, Berthier Ribeiro-Neto, Rodrigo F. Vale, Alberto H. F. Laender, and Luciano R. S. Lima. Categorization-driven cross-language retrieval of medical information. Journal of the American Society for Information Science and Technology: JASIST, 57(4):501–510, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Finn:2006:LCD

[FK06] Aidan Finn and Nicholas Kushmerick. Learning to classify documents according to genre. Journal of the American Society for Information Science and Technology: JASIST, 57 (11):1506–1518, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fraser:2001:ELA

[FL01] Landon Fraser and Craig Locatis. Effects of link annotations on search performance in layered and unlayered hierarchically organized information spaces. Journal of the American Society for Information Science and Technology: JASIST, 52(14):1255–1261, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fleischmann:2006:BRB

[Fle06a] Kenneth R. Fleischmann. Book review: The digital sublime: Myth, power, and cyberspace. Journal of the American Society for Information Science and Technology: JASIST, 57 (7):989–990, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fleischmann:2006:DIY

[Fle06b] Kenneth R. Fleischmann. Do-it-yourself information technology: Role hybridization and the design–use interface. Journal of the American Society for Information Science and Technology: JASIST, 57(1):87–95, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fonseca:2005:TAN

[FM05] Frederico T. Fonseca and James E. Martin. Toward an alternative notion of information systems ontologies: Information engineering as a hermeneutic enterprise. Journal of the American Society for Information Science and Technology: JASIST, 56(1):46–57, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frias-Martinez:2007:ACS

[FMCL07] Enrique Frias-Martinez, Sherry Y. Chen, and Xiaohui Liu. Automatic cognitive style identification of digital library users for personalization. Journal of the American Society for Information Science and Technology: JASIST, 58 (2):237–251, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ford:2001:RID

[FMM01] Nigel Ford, David Miller, and Nicola Moss. The role of individual differences in Internet searching: an empirical study. Journal of the American Society for Information Science and Technology: JASIST, 52(12):1049–1066, ???? 2001. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ford:2003:WSS

[FMM03] Nigel Ford, David Miller, and Nicola Moss. Web search strategies and approaches to studying. Journal of the American Society for Information Science and Technology: JA-SIST, 54(6):473–489, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ford:2005:WSSb

[FMM05a] Nigel Ford, David Miller, and Nicola Moss. Web search strategies and human individual differences: a combined analysis. Journal of the American Society for Information Science and Technology: JASIST, 56(7):757–764, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ford:2005:WSSa

[FMM05b] Nigel Ford, David Miller, and Nicola Moss. Web search strategies and human individual differences: Cognitive and demographic factors, Internet attitudes, and approaches. Journal of the American Society for Information Science and Technology: JASIST, 56(7):741–756, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fabrikant:2010:NLM

[FMM10] Sara Irina Fabrikant, Daniel R. Montello, and David M. Mark. The natural landscape metaphor in information visualization: the role of commonsense geomorphology. Journal of the American Society for Information Science and Technology: JASIST, 61(2):253–270, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Franceschini:2013:NAE

[FMM13] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. A novel approach for estimating the omittedcitation rate of bibliometric databases with an application to the field of bibliometrics. Journal of the American Society for Information Science and Technology: JASIST, 64(10): 2149–2156, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fullwood:2013:SBB

[FMMF13] Chris Fullwood, Karen Melrose, Neil Morris, and Sarah Floyd. Sex, blogs, and baring your soul: Factors influencing UK blogging strategies. Journal of the American Society for Information Science and Technology: JASIST, 64(2): 345–355, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fernandez-Molina:2001:MRA

[FMP01] J. Carlos Fernández-Molina and Eduardo Peis. The moral rights of authors in the age of digital information. Journal of the American Society for Information Science and Technology: JASIST, 52(2):109–117, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frandsen:2008:IDD

[FN08] Tove Faber Frandsen and Jeppe Nicolaisen. Intradisciplinary differences in database coverage and the consequences for bibliometric research. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1570– 1581, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frandsen:2011:PBC

[FN11] Tove Faber Frandsen and Jeppe Nicolaisen. Praise the bridge that carries you over: Testing the flattery citation hypothesis. Journal of the American Society for Information Science and Technology: JASIST, 62(5):807–818, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frandsen:2012:EAE

[FN12] Tove Faber Frandsen and Jeppe Nicolaisen. Effects of academic experience and prestige on researchers’ citing behavior. Journal of the American Society for Information Science and Technology: JASIST, 63(1):64–71, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frandsen:2013:REC

[FN13] Tove Faber Frandsen and Jeppe Nicolaisen. The ripple effect: Citation chain reactions of a Nobel Prize. Journal of the American Society for Information Science and Technology: JASIST, 64(3):437–447, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fonseca:2007:DRO

[Fon07] Frederico Fonseca. The double role of ontologies in information science research. Journal of the American Society for Information Science and Technology: JASIST, 58(6):786– 793, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ford:2004:CCI

[For04a] Nigel Ford. Creativity and convergence in information science research: the roles of objectivity and subjectivity, constraint, and control. Journal of the American Society for Information Science and Technology: JASIST, 55(13):1169– 1182, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ford:2004:MCP

[For04b] Nigel Ford. Modeling cognitive processes in information seeking: From Popper to Pask. Journal of the American Society for Information Science and Technology: JASIST, 55 (9):769–782, July 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ford:2009:BRB

[For09] Kevin M. Ford. Book review: Ambivalence Towards Convergence: Digitalization and Media Change. Journal of the American Society for Information Science and Technology: JASIST, 60(2):427–428, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ford:2013:BRB

[For13] Nigel Ford. Book review: Information Need: A Theory Connecting Information Search to Knowledge Formation — By Charles Cole. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2595–2596, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fosmire:2001:BRB

[Fos01] Michael Fosmire. Book review: Electronic expectations: Science journals on the Web, by Tony Stankus. Journal of the American Society for Information Science and Technology: JASIST, 52(6):508–509, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Foster:2004:NMI

[Fos04] Allen Foster. A nonlinear model of information-seeking behavior. Journal of the American Society for Information Science and Technology: JASIST, 55(3):228–237, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fourie:2001:BRBb

[Fou01a] Ina Fourie. Book review: Academic libraries as high-tech gateways: a guide to design & space decisions (2nd Edition), by Richard J. Bazillion & Connie L. Braun. Journal of the American Society for Information Science and Technology:

JASIST, 52(11):983, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fourie:2001:BRBa

[Fou01b] Ina Fourie. Book review: From Web to workplace: Designing open hypermedia systems, by Kaj Grønbaek and Randall H. Trigg. Journal of the American Society for Information Science and Technology: JASIST, 52(6):510–512, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fourie:2002:BRB

[Fou02] Ina Fourie. Book review: Intelligent technologies in library and information service applications. Journal of the American Society for Information Science and Technology: JA-SIST, 53(4):321, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fourie:2003:BR

[Fou03] Ina Fourie. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 54 (2):184–185, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fourie:2012:BRB

[Fou12] Ina Fourie. Book review: Introduction to Information Science and Technology. Edited by Charles H. Davis & Deborah Shaw. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1673–1674, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frank:2004:PLC

[FP04] Eibe Frank and Gordon W. Paynter. Predicting Library of Congress classifications from Library of Congress subject headings. Journal of the American Society for Information Science and Technology: JASIST, 55(3):214–227, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fidel:2004:MAS

[FPCB04] Raya Fidel, Annelise Mark Pejtersen, Bryan Cleal, and Harry Bruce. A multidimensional approach to the study of human-information interaction: a case study of collaborative information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 55(11):939– 953, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Faba-Perez:2005:CAW

[FPZAGBdMA05] Cristina Faba-Pérez, Felipe Zapico-Alonso, Vicente P. Guerrero-Bote, and Félix de Moya-Anegón. Comparative analysis of Webometric measurements in thematic environments. Journal of the American Society for Information Science and Technology: JASIST, 56(8):779–785, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frohlich:2001:APC

[FR01] Cliff Frohlich and Lynn Resler. Analysis of publications and citations from a geophysics research institute. Journal of the American Society for Information Science and Technology: JASIST, 52(9):701–713, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frandsen:2005:AIC

[FR05] Tove Faber Frandsen and Ronald Rousseau. Article impact calculated over arbitrary periods. Journal of the American Society for Information Science and Technology: JASIST, 56(1):58–62, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fattah:2008:EAP

[FR08] Mohamed Abdel Fattah and Fuji Ren. English-Arabic proper-noun transliteration-pairs creation. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1675–1687, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ferrara:2013:SIE

[FR13] Emilio Ferrara and Alfonso E. Romero. Scientific impact evaluation and the effect of self-citations: Mitigating the bias by discounting the h-index. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2332–2339, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frandsen:2003:LES

[Fra03] Tove Faber Frandsen. Letters to the Editor: Simplifying the reversed duplicate removal procedure. Journal of the American Society for Information Science and Technology: JASIST, 54(3):275–276, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Franceschet:2009:CAS

[Fra09] Massimo Franceschet. A cluster analysis of scholar and journal bibliometric indicators. Journal of the American Society for Information Science and Technology: JASIST, 60(10): 1950–1964, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Franceschet:2011:CCS

[Fra11] Massimo Franceschet. Collaboration in computer science: a network science approach. Journal of the American Society for Information Science and Technology: JASIST, 62(10): 1992–2012, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Franceschet:2012:LSS

[Fra12] Massimo Franceschet. The large-scale structure of journal citation networks. Journal of the American Society for Information Science and Technology: JASIST, 63(4):837–842, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frohmann:2008:SIE

[Fro08] Bernd Frohmann. Subjectivity and information ethics. Journal of the American Society for Information Science and Technology: JASIST, 59(2):267–277, January 15, 2008. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fry:2010:BRB

[Fry10] Jenny Fry. Book review: Scholarship in the Digital Age: Information, Infrastructure, and the Internet. Journal of the American Society for Information Science and Technology: JASIST, 61(3):636–637, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Flammia:2007:BCL

[FS07] Madelyn Flammia and Carol Saunders. Brief communications: Language as power on the Internet. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1899–1903, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fautsch:2009:ASM

[FS09] Claire Fautsch and Jacques Savoy. Algorithmic stemmers or morphological analysis? An evaluation. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1616–1624, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fagni:2010:SNE

[FS10] Tiziano Fagni and Fabrizio Sebastiani. Selecting negative examples for hierarchical text classification: an experimental comparison. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2256–2265, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fu:2011:ABE

[FS11] W. Wayne Fu and Clarice C. Sim. Aggregate bandwagon effect on online videos’ viewership: Value uncertainty, popularity cues, and heuristics. Journal of the American Society for Information Science and Technology: JASIST, 62(12): 2382–2395, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fegley:2012:RPV

[FT12] Brent D. Fegley and Vetle I. Torvik. On the role of poetic versus nonpoetic features in “kindred” and diachronic poetry attribution. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2165–2181, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Fu:2010:TMI

[Fu10] Xin Fu. Towards a model of implicit feedback for Web search. Journal of the American Society for Information Science and Technology: JASIST, 61(1):30–49, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Furner:2002:R

[Fur02] Jonathan Furner. On recommending. Journal of the American Society for Information Science and Technology: JA-SIST, 53(9):747–763, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Furner:2010:EAI

[Fur10] Jonathan Furner. Editorial: Advances in Information Science. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1733, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Furner:2011:EAI

[Fur11] Jonathan Furner. Editorial: Advances in Information Science. Journal of the American Society for Information Science and Technology: JASIST, 62(6):995, June 2011. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Frandsen:2009:TWP

[FW09] Tove Faber Frandsen and Paul Wouters. Turning working papers into journal articles: an exercise in microbibliometrics. Journal of the American Society for Information Science and Technology: JASIST, 60(4):728–739, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ford:2002:ISM

[FWF^{+}02] Nigel Ford, T. D. Wilson, Allen Foster, David Ellis, and Amanda Spink. Information seeking and mediated searching. Part 4. Cognitive styles in information seeking. Journal of the American Society for Information Science and Technology: JASIST, 53(9):728–735, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gonzalez-Alcaide:2008:BCL

[GACCNM^{+}08] Gregorio González-Alcaide, Lourdes Castelló-Cogollos, Carolina Navarro-Molina, Rafael Aleixandre-Benavent, and Juan Carlos Valderrama-Zurián. Brief communications: Library and information science research areas: Analysis of journal articles in lisa. Journal of the American Society for Information Science and Technology: JASIST, 59(1):150– 154, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gaffney:2007:BRB

[Gaf07] Sean Gaffney. Book review: The Internet imaginaire. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2405–2407, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Galloway:2006:BRB

[Gal06] Patricia Galloway. Book review: Social capital and information technology. Journal of the American Society for Information Science and Technology: JASIST, 57(5):723–724, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Galloway:2012:BRBa

[Gal12a] Patricia Galloway. Book review: I, Digital: Personal Collections in the Digital Era. Edited by Christopher A. Lee. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1278–1279, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Galloway:2012:BRBb

[Gal12b] Patricia Galloway. Book reviews: Designing Culture: The Technological Imagination at Work. By Anne Balsamo. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1899–1900, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Goh:2011:FUI

[GALC11] Dion Hoe-Lian Goh, Rebecca P. Ang, Chei Sian Lee, and Alton Y. K. Chua. Fight or unite: Investigating game genres for image tagging. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1311– 1324, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garfield:2001:RIH

[Gar01a] Eugene Garfield. Recollections of Irving H. Sher 1924–1996: Polymath/ information scientist extraordinaire. Journal of the American Society for Information Science and Technology: JASIST, 52(14):1197–1202, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garfield:2001:RPV

[Gar01b] Eugene Garfield. A retrospective and prospective view of information retrieval and artificial intelligence in the 21st Century. Journal of the American Society for Information Science and Technology: JASIST, 52(1):18–21, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gaule:2009:ASL

[Gau09] Patrick Gaulé. Access to scientific literature in India. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2548–2553, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gazan:2007:BRB

[Gaz07] Rich Gazan. Book review: Understanding and communicating social informatics: a framework for studying and teaching the human contexts of information and communication technologies. Journal of the American Society for Information Science and Technology: JASIST, 58(1):151–152, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gazan:2009:BRB

[Gaz09] Rich Gazan. Book review: Fundamentals of Information Studies: Understanding Information and Its Environment, Second Edition. Journal of the American Society for Information Science and Technology: JASIST, 60(2):429, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gazan:2011:AIS

[Gaz11] Rich Gazan. Advances in information science: Social Q&A. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2301–2312, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Grauwin:2012:CSS

[GBF^{+}12] Sebastian Grauwin, Guillaume Beslon, Éric Fleury, Sara Franceschelli, Celine Robardet, Jean-Baptiste Rouquier, and Pablo Jensen. Complex systems science: Dreams of universality, interdisciplinarity reality. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1327–1338, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Goren-Bar:2005:SUS

[GBK05] Dina Goren-Bar and Tsvi Kuflik. Supporting user-subjective categorization with self-organizing maps and learning vector quantization. Journal of the American Society for Information Science and Technology: JASIST, 56(4):345–355, February 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gentil-Beccot:2009:IRH

[GBMH^{+}09] Anne Gentil-Beccot, Salvatore Mele, Annette Holtkamp, Heath B. O’Connell, and Travis C. Brooks. Information resources in High-Energy Physics: Surveying the present landscape and charting the future course. Journal of the American Society for Information Science and Technology: JASIST, 60(1):150–160, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gonzalez:2007:SIF

[GC07] Lidia González and Juan Miguel Campanario. Structure of the impact factor of journals included in the Social Sciences Citation Index: Citations from documents labeled “editorial material”. Journal of the American Society for Information Science and Technology: JASIST, 58(2):252–262, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gauch:2009:THC

[GCR09] Susan Gauch, Aravind Chandramouli, and Shankar Ranganathan. Training a hierarchical classifier using inter document relationships. Journal of the American Society for Information Science and Technology: JASIST, 60(1):47–58, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gottlieb:2001:RUP

[GD01] Lisa Gottlieb and Juris Dilevko. Research: User preferences in the classification of electronic bookmarks: Implications for a shared system. Journal of the American Society for Information Science and Technology: JASIST, 52(7):517– 535, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gottlieb:2003:IHI

[GD03] Lisa Gottlieb and Juris Dilevko. Investigating how individuals conceptually and physically structure file folders for electronic bookmarks: the example of the financial services industry. Journal of the American Society for Information Science and Technology: JASIST, 54(2):124–139, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Graham:2007:KST

[GD07] Paul J. Graham and Harley D. Dickinson. Knowledgesystem theory in society: Charting the growth of knowledgesystem models over a decade, 1994–2003. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2372–2381, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Guerrero:2001:RDD

[GdMA01] Vicente P. Guerrero and Félix de Moya Anegón. Reduction of the dimension of a document space using the fuzzified output of a Kohonen network. Journal of the American Society for Information Science and Technology: JASIST, 52(14): 1234–1241, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gelernter:2009:IIA

[Gel09] Judith Gelernter. Image indexing in article component databases. Journal of the American Society for Information Science and Technology: JASIST, 60(10):1965–1976, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Genuis:2006:ERM

[Gen06] Shelagh K. Genuis. Exploring the role of medical and consumer literature in the diffusion of information related to hormone therapy for menopausal women. Journal of the American Society for Information Science and Technology: JASIST, 57(7):974–988, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Genuis:2012:CSE

[Gen12] Shelagh K. Genuis. Constructing “sense” from evolving health information: a qualitative investigation of information seeking and sense making across sources. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1553–1566, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Guan:2009:EIP

[GG09] Jian Cheng Guan and Xia Gao. Exploring the h-index at patent level. Journal of the American Society for Information Science and Technology: JASIST, 60(1):35–40, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Golbeck:2010:TUU

[GGR10] Jennifer Golbeck, Justin M. Grimes, and Anthony Rogers. Twitter use by the U.S. Congress. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1612–1621, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Golbeck:2011:ESS

[GKE11] Jennifer Golbeck, Jes Koepfler, and Beth Emmerling. An experimental study of social tagging behavior and image content. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1750–1760, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gerlach:2009:SSR

[GKL09] James H. Gerlach, Feng-Yang Bob Kuo, and Cathy S. Lin. Self sanction and regulative sanction against copyright infringement: a comparison between U.S. and China college students. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1687–1701, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gwizdka:2009:RSF

[GL09] Jacek Gwizdka and Irene Lopatovska. The role of subjective factors in the information search process. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2452–2464, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gross:2012:WSG

[GL12] Melissa Gross and Don Latham. What’s skill got to do with it?: Information literacy skills and self-views of ability among first-year college students. Journal of the American Society for Information Science and Technology: JASIST, 63(3):574–583, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gil-Leiva:2007:KGA

[GLAA07] Isidoro Gil-Leiva and Adolfo Alonso-Arroyo. Keywords given by authors of scientific articles in database descriptors. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1175–1187, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gorman:2002:FEW

[GLDM02] Paul Gorman, Mary Lavelle, Lois Delcambre, and David Maier. Following experts at work in their own information spaces: Using observational methods to develop tools for the digital library. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1245–1250, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Guo:2011:FSR

[GLLL11] Zixiu Guo, Xiongwen Lu, Yuan Li, and Yifan Li. A framework of students’ reasons for using CMC media in learning contexts: a structural approach. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2182–2200, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gluck:2001:RME

[Glu01] Myke Gluck. Research: Multimedia exploratory data analysis for geospatial data mining: the case for augmented seriation. Journal of the American Society for Information Science and Technology: JASIST, 52(8):686–696, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Glusker:2008:BRB

[Glu08] Ann Glusker. Book review: Government documents librarianship: a guide for the neo-depository era. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1695–1696, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Galvez:2007:APN

[GMA07] Carmen Galvez and Félix Moya-Anegón. Approximate personal name-matching through finite-state graphs. Journal of the American Society for Information Science and Technology: JASIST, 58(13):1960–1976, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gianoli:2009:BCI

[GMM09] Ernesto Gianoli and Marco A. Molina-Montenegro. Brief communications: Insights into the relationship between the h-index and self-citations. Journal of the American Society for Information Science and Technology: JASIST, 60(6): 1283–1285, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Goggins:2013:GIM

[GMV13] Sean P. Goggins, Christopher Mascaro, and Giuseppe Valetto. Group informatics: a methodological approach and ontology for sociotechnical group research. Journal of the American Society for Information Science and Technology: JASIST, 64(3):516–539, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Goh:2007:LDL

[GN07] Dion Hoe-Lian Goh and Peng Kin Ng. Link decay in leading information science journals. Journal of the American Society for Information Science and Technology: JASIST, 58(1):

15–24, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Greisdorf:2003:NTM

[GO03] Howard Greisdorf and Brian O’Connor. Nodes of topicality: Modeling user notions of on topic documents. Journal of the American Society for Information Science and Technology: JASIST, 54(14):1296–1304, December 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Godejord:2007:FCP

[God07] Per Arne Godejord. Fighting child pornography: Exploring didactics and student engagement in social informatics. Journal of the American Society for Information Science and Technology: JASIST, 58(3):446–451, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gonzalez:2008:BRB

[Gon08] Allegra Gonzalez. Book review: Digital citizenship: The Internet, society, and participation. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2189–2190, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Goodrum:2001:RMS

[Goo01] Abby A. Goodrum. Research: Multidimensional scaling of video surrogates. Journal of the American Society for Information Science and Technology: JASIST, 52(2):174–182, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gordon:2001:BIC

[Gor01] Andrew S. Gordon. Browsing image collections with representations of common-sense activities. Journal of the American Society for Information Science and Technology: JA-SIST, 52(11):925–929, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gorbunov:2002:RWD

[Gor02] Andrey L. Gorbunov. Relevance of Web documents: Ghosts consensus method. Journal of the American Society for Information Science and Technology: JASIST, 53(10):783– 788, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garces:2006:CMI

[GOR06] Pablo J. Garcés, José A. Olivas, and Francisco P. Romero. Concept-matching IR systems versus word-matching information retrieval systems: Considering fuzzy interrelations for indexing Web pages. Journal of the American Society for Information Science and Technology: JASIST, 57(4):564– 576, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gordon:2012:ISP

[Gor12] Avishag Gordon. The invisibility of science publications in hebrew: a comparative database study. Journal of the American Society for Information Science and Technology: JA-SIST, 63(3):607–615, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garcia-Perez:2010:ACP

[GP10] Miguel A. García-Pérez. Accuracy and completeness of publication and citation records in the Web of Science, PsycINFO, and Google Scholar: a case study for the computation of h indices in Psychology. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2070–2085, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Grad:2011:PAV

[GPG^{+}11] Roland Grad, Pierre Pluye, Vera Granikov, Janique Johnson-Lafleur, Michael Shulha, Soumya Bindiganavile Sridhar, Jonathan L. Moscovici, Gillian Bartlett, Alain C. Vandal, Bernard Marlow, and Lorie Kloda. Physicians’ assessment of the value of clinical information: Operationalization of a theoretical model. Journal of the American Society for Information Science and Technology: JASIST, 62(10):1884–1891, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gorraiz:2013:OLB

[GPG13] Juan Gorraiz, Philip J. Purnell, and Wolfgang Glänzel. Opportunities for and limitations of the Book Citation Index.

Journal of the American Society for Information Science and Technology: JASIST, 64(7):1388–1398, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garfield:2003:WDW

[GPI03] Eugene Garfield, A. I. Pudovkin, and V. S. Istomin. Why do we need algorithmic historiography? Journal of the American Society for Information Science and Technology: JA-SIST, 54(5):400–412, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Guns:2009:SGI

[GR09] Raf Guns and Ronald Rousseau. Simulating growth of the h-index. Journal of the American Society for Information Science and Technology: JASIST, 60(2):410–417, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Graef:2005:BRB

[Gra05] Jean Graef. Book review: Information strategy in practice. Journal of the American Society for Information Science and Technology: JASIST, 56(14):1556–1557, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Greenberg:2001:OQE

[Gre01a] Jane Greenberg. Optimal query expansion (QE) processing methods with semantically encoded structured thesauri terminology. Journal of the American Society for Information Science and Technology: JASIST, 52(6):487–498, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Greenberg:2001:QCA

[Gre01b] Jane Greenberg. A quantitative categorical analysis of metadata elements in image-applicable metadata schemas. Journal of the American Society for Information Science and Technology: JASIST, 52(11):917–924, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Greenberg:2001:RAQ

[Gre01c] Jane Greenberg. Research: Automatic query expansion via lexical–semantic relationships. Journal of the American Society for Information Science and Technology: JASIST, 52 (5):402–415, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gremett:2006:PSU

[Gre06] Peter Gremett. Perspectives on search user interfaces: Best practices and future visions: Utilizing a user’s context to improve search results. Journal of the American Society for Information Science and Technology: JASIST, 57(6):808– 812, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Goodrum:2001:OSA

[GRJS01] Abby A. Goodrum, Mark E. Rorvig, Ki-Tai Jeong, and Chitturi Suresh. An open source agenda for research linking text and image content features. Journal of the American Society for Information Science and Technology: JASIST, 52 (11):948–953, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Given:2007:IID

[GRS^{+}07] Lisa M. Given, Stan Ruecker, Heather Simpson, Elizabeth (Bess) Sadler, and Andrea Ruskin. Inclusive interface design for seniors: Image-browsing for a health information context. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1610–1617, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garcia:2011:RSA

[GRSFV11] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Ranking of the subject areas of Scopus. Journal of the American Society for Information Science and Technology: JA-SIST, 62(10):2013–2023, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garcia:2012:SSC

[GRSFV12a] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Scientific subject categories of Web of Knowledge ranked according to their multidimensional prestige of influential journals. Journal of the American Society for Information Science and Technology: JASIST, 63(5):1017–1029, May 2012.

CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garcia:2012:MAI

[GRSFV^{+}12b] J. A. García, Rosa Rodríguez-Sánchez, J. Fdez-Valdivia, N. Robinson-García, and D. Torres-Salinas. Mapping academic institutions according to their journal publication profile: Spanish universities as a case study. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2328–2340, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gruzd:2007:BRB

[Gru07] Anatoliy Gruzd. Book review: New directions in cognitive information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 58(5):758– 760, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Glanzel:2012:VRR

[GRZ12] Wolfgang Glänzel, Ronald Rousseau, and Lin Zhang. A visual representation of relative first-citation times. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1420–1425, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Giorgetti:2003:ASC

[GS03] Daniela Giorgetti and Fabrizio Sebastiani. Automating survey coding by multiclass text categorization techniques. Journal of the American Society for Information Science and Technology: JASIST, 54(14):1269–1277, December 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gorgeon:2011:WAW

[GS11] Arnaud Gorgeon and E. Burton Swanson. Web 2.0 according to Wikipedia: Capturing an organizing vision. Journal of the American Society for Information Science and Technology: JASIST, 62(10):1916–1932, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garoufallou:2008:VMV

[GSB08] Emmanouel Garoufallou, Rania Siatri, and Panagiotis Balatsoukas. Virtual maps–virtual worlds: Testing the usability of a Greek virtual cultural map. Journal of the American Society for Information Science and Technology: JASIST, 59(4):591–601, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gazni:2012:MWS

[GSD12] Ali Gazni, Cassidy R. Sugimoto, and Fereshteh Didegah. Mapping world scientific collaboration: Authors, institutions, and countries. Journal of the American Society for Information Science and Technology: JASIST, 63(2):323–335, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Ano13b].

Geng:2005:CKM

[GTHZ05] Qian Geng, Charles Townley, Kun Huang, and Jing Zhang. Comparative knowledge management: a pilot study of Chinese and American universities. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1031–1044, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gu:2002:ESM

[Gu02] Yinian Gu. An exploratory study of Malaysian publication productivity in computer science and information technology. Journal of the American Society for Information Science and Technology: JASIST, 53(12):974–986, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Guastavino:2011:BRB

[Gua11] Catherine Guastavino. Book review: The Tuning of Place: Sociable Spaces and Pervasive Digital Media. Journal of the American Society for Information Science and Technology: JASIST, 62(4):805, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Guns:2013:TOS

[Gun13] Raf Guns. Tracing the origins of the semantic web. Journal of the American Society for Information Science and Technology: JASIST, 64(10):2173–2181, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gupta:2009:PDI

[Gup09] Sanjay Gupta. Perspectives on design: Information technologies and creative practices: Communication clothing design. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1915–1919, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Garnatje:2013:LED

[GV13] Teresa Garnatje and Joan Vallès. Letter to the Editor: Does aggressiveness in evaluation improve the quality of scientific research? Journal of the American Society for Information Science and Technology: JASIST, 64(8):1756, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Guo:2012:ESS

[GW12] Liqiang Guo and Xiaojun Wan. Exploiting syntactic and semantic relationships between terms for opinion retrieval. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2269–2282, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gwizdka:2010:DCL

[Gwi10] Jacek Gwizdka. Distribution of cognitive load in Web search. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2167–2187, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Gracy:2013:EMI

[GZS13] Karen F. Gracy, Marcia Lei Zeng, and Laurence Skirvin. Exploring methods to improve access to Music resources by aligning library Data with Linked Data: a report of methodologies and preliminary findings. Journal of the American Society for Information Science and Technology: JASIST, 64(10):2078–2099, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Haas:2003:ISE

[Haa03] Stephanie W. Haas. Improving the search environment: Informed decision making in the search for statistical information. Journal of the American Society for Information Science and Technology: JASIST, 54(8):782–797, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Haigh:2012:BRB

[Hai12] Thomas Haigh. Book reviews: Emanuel Goldberg and his knowledge machine. Journal of the American Society for Information Science and Technology: JASIST, 63(2):427–428, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Halavais:2001:BRB

[Hal01] Alexander Halavais. Book review: The robot in the garden: Telerobotics and telepistemology in the age of the Internet, edited by Ken Goldberg. Journal of the American Society for Information Science and Technology: JASIST, 52(7):598– 599, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hall:2004:LEM

[Hal04a] Homer J. Hall. Letters to the Editor: Meet to agree or meet to argue? How to run a committee meeting for action. Journal of the American Society for Information Science and Technology: JASIST, 55(13):1210–1211, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hall:2004:LEQ

[Hal04b] Homer J. Hall. Letters to the Editor: Quality of information. Journal of the American Society for Information Science and Technology: JASIST, 55(1):92, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hall:2009:PDI

[Hal09] Peter Hall. Perspectives on design: Information technologies and creative practices: Disorderly reasoning in information design. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1877–1882, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hall:2010:BRB

[Hal10] Hazel Hall. Book review: Principles of Knowledge Management: Theory, Practice, and Cases. Journal of the American Society for Information Science and Technology: JASIST, 61(2):430–432, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hamilton:2007:ISD

[Ham07] Elizabeth C. Hamilton. The impact of survey data: Measuring success. Journal of the American Society for Information Science and Technology: JASIST, 58(2):190–199, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hammarfelt:2011:CAM

[Ham11] Björn Hammarfelt. Citation analysis on the micro level: the example of Walter Benjamin’s Illuminations. Journal of the American Society for Information Science and Technology: JASIST, 62(5):819–830, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hanson:2001:BRB

[Han01] Charles Hanson. Book review: Information appliances and beyond: Interaction design for consumer products, Eric Bergman, editor. Journal of the American Society for Information Science and Technology: JASIST, 52(4):365–366, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hargittai:2002:BLS

[Har02] Eszter Hargittai. Beyond logs and surveys: In-depth measures of people’s web use skills. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1239–1244, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Harris:2004:BRB

[Har04] Lydia Eato Harris. Book review: Research Questions for the Twenty-First Century. Journal of the American Society for Information Science and Technology: JASIST, 55(7):652– 655, May 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Harris:2005:BRB

[Har05] Lydia Eato Harris. Book review: Benchmarks in distance education: The LIS experience. Journal of the American Society for Information Science and Technology: JASIST, 56(6):658–659, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Harris:2006:BRB

[Har06a] Lydia Eato Harris. Book review: Foundations of library and information science, 2nd edition. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1279–1280, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hartley:2006:RWB

[Har06b] James Hartley. Reading and writing book reviews across the disciplines. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1194–1207, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hara:2007:ITS

[Har07a] Noriko Hara. Information technology support for communities of practice: How public defenders learn about winning and losing in court. Journal of the American Society for Information Science and Technology: JASIST, 58(1):76–87, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Harris:2007:BRB

[Har07b] Lydia Eato Harris. Book review: Spanning the theorypractice divide in library and information science. Journal of the American Society for Information Science and Technology: JASIST, 58(4):606–607, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Harris:2008:BRB

[Har08a] Lydia E. Harris. Book review: Social software in libraries: Building collaborations, communication, and community online. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1866, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Harwood:2008:BCC

[Har08b] Nigel Harwood. Brief communications: Citers’ use of citees’ names: Findings from a qualitative interview-based study. Journal of the American Society for Information Science and Technology: JASIST, 59(6):1007–1011, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Harmon:2009:BRB

[Har09] Glynn Harmon. Book review: Connections: Patterns of Discovery. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2596–2598, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hassan:2003:SMI

[Has03a] E. Hassan. Simultaneous mapping of interactions between scientific and technological knowledge bases: the case of space communications. Journal of the American Society for Information Science and Technology: JASIST, 54(5):462– 468, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hassell:2003:BRB

[Has03b] Lewis Hassell. Book review: Social Thinking — Software Practice. Journal of the American Society for Information Science and Technology: JASIST, 54(10):979–981, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hassanien:2004:RSA

[Has04] Aboul-Ella Hassanien. Rough set approach for attribute reduction and rule generation: a case of patients with suspected breast cancer. Journal of the American Society for Information Science and Technology: JASIST, 55(11):954– 962, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hassell:2009:BRB

[Has09] Lew Hassell. Book review: Computerization Movements and Technology Diffusion. Journal of the American Society for Information Science and Technology: JASIST, 60(3):638– 639, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hmeidi:2010:NAE

[HASAT^{+}10] Ismail I. Hmeidi, Riyad F. Al-Shalabi, Ahmad T. Al-Taani, Hassan Najadat, and Shaker A. Al-Hazaimeh. A novel approach to the extraction of roots from Arabic words using bigrams. Journal of the American Society for Information Science and Technology: JASIST, 61(3):583–591, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2012:CCI

[HAT12] Hong Huang, James Andrews, and Jiang Tang. Citation characterization and impact normalization in bioinformatics journals. Journal of the American Society for Information Science and Technology: JASIST, 63(3):490–497, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hauptman:2005:CPP

[Hau05] Robert Hauptman. Call for papers: Papers on ethics for publication. Journal of the American Society for Information Science and Technology: JASIST, 56(7):775, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hauptman:2008:LE

[Hau08] Robert Hauptman. Letter to the Editor. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1697, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Haverty:2002:IAI

[Hav02] Marsha Haverty. Information architecture without internal theory: an inductive design process. Journal of the American Society for Information Science and Technology: JASIST, 53(10):839–845, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hawkins:2001:ISA

[Haw01] Donald T. Hawkins. Information science abstracts: Tracking the literature of information science. Part 1: Definition and map. Journal of the American Society for Information Science and Technology: JASIST, 52(1):44–53, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Haythornthwaite:2006:LKN

[Hay06] Caroline Haythornthwaite. Learning and knowledge networks in interdisciplinary collaborations. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1079–1092, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hayden:2007:BRB

[Hay07] Lance Hayden. Book review: Covert and overt: Recollecting and connecting intelligence service and information science. Journal of the American Society for Information Science and Technology: JASIST, 58(2):303–305, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hartley:2007:EST

[HB07] James Hartley and Lucy Betts. The effects of spacing and titles on judgments of the effectiveness of structured abstracts. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2335–2340, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hartley:2008:RPS

[HB08] James Hartley and Lucy Betts. Revising and polishing a structured abstract: Is it worth the time and effort? Journal of the American Society for Information Science and Technology: JASIST, 59(12):1870–1877, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hartley:2009:CWT

[HB09] James Hartley and Lucy Betts. Common weaknesses in traditional abstracts in the social sciences. Journal of the American Society for Information Science and Technology: JA-SIST, 60(10):2010–2018, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Herdagdelen:2011:SGA

[HB11] Amaç Herdağdelen and Marco Baroni. Stereotypical gender actions can be extracted from web text. Journal of the American Society for Information Science and Technology: JA-SIST, 62(9):1741–1749, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hofmann:2010:CFF

[HBBdR10] Katja Hofmann, Krisztian Balog, Toine Bogers, and Maarten de Rijke. Contextual factors for finding similar experts. Journal of the American Society for Information Science and Technology: JASIST, 61(5):994–1014, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2013:WDG

[HBSC13] Meikuan Huang, Joshua Barbour, Chunke Su, and Noshir Contractor. Why do group members provide information to digital knowledge repositories? a multilevel application of transactive memory theory. Journal of the American Society for Information Science and Technology: JASIST, 64 (3):540–557, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ho:2013:SBI

[HBSZ13] Shuyuan Mary Ho, Michael Bieber, Min Song, and Xiangmin Zhang. Seeking beyond with IntegraL: a user study of sense-making enabled by anchor-based virtual integration of library systems. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1927–1945, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hu:2009:DSQ

[HBT^{+}09] Paul Jen-Hwa Hu, Susan A. Brown, James Y. L. Thong, Frank K. Y. Chan, and Kar Yan Tam. Determinants of service quality and continuance intention of online services: the case of eTax. Journal of the American Society for Information Science and Technology: JASIST, 60(2):292–306, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2002:BCW

[HC02] Birger Hjørland and Frank Sejer Christensen. Brief communications: Work tasks and socio-cognitive relevance: a specific example. Journal of the American Society for Information Science and Technology: JASIST, 53(11):960–965, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hung:2007:WBT

[HC07] Chen-Ming Hung and Lee-Feng Chien. Web-based text classification in the absence of manually labeled training documents. Journal of the American Society for Information Science and Technology: JASIST, 58(1):88–96, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2009:CWS

[HC09] Chun-Yao Huang and Shin-Shin Chang. Commonality of Web site visiting among countries. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1168–1177, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hsu:2011:EEP

[HC11] Ming-Hung Hsu and Hsin-Hsi Chen. Efficient and effective prediction of social tags to enhance web search. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1473–1487, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2004:GMC

[HCC04] Zan Huang, Wingyan Chung, and Hsinchun Chen. A graph model for E-commerce recommender systems. Journal of the American Society for Information Science and Technology: JASIST, 55(3):259–274, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hsu:2005:RWW

[HCH^{+}05] Chun-Nan Hsu, Chia-Hui Chang, Chang-Huain Hsieh, Jiann-Jyh Lu, and Chien-Chi Chang. Reconfigurable Web wrapper agents for biological information integration. Journal of the American Society for Information Science and Technology: JASIST, 56(5):505–517, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2005:WEC

[hChC05] Hsueh hua Chen and Tzu heng Chiu. The working environment and changing role of corporate librarians in Taiwan. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1227–1236, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2003:RTS

[HCO03] Chien-Kang Huang, Lee-Feng Chien, and Yen-Jen Oyang. Relevant term suggestion in interactive web search based on contextual information in query session logs. Journal of the American Society for Information Science and Technology: JASIST, 54(7):638–649, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hirsh:2004:SIO

[HD04] Sandra Hirsh and Jamie Dinkelacker. Seeking information in order to produce information: an empirical study at Hewlett Packard Labs. Journal of the American Society for Information Science and Technology: JASIST, 55(9):807–817, July 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hupfer:2006:GWI

[HD06] Maureen E. Hupfer and Brian Detlor. Gender and Web information seeking: a self-concept orientation model. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1105–1115, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hutchinson:2007:SEA

[HDB07] Hilary Browne Hutchinson, Allison Druin, and Benjamin B. Bederson. Supporting elementary-age children’s searching and browsing: Design and evaluation using the international children’s digital library. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1618– 1630, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hert:2007:IMM

[HDG^{+}07] Carol A. Hert, Sheila O. Denn, Daniel W. Gillman, Jung Sun Oh, Maria Cristina Pattuelli, and Naybell Hernández. Investigating and modeling metadata use to support information architecture development in the statistical knowledge network. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1267–1284, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

He:2011:MEC

[HDN11] Bing He, Ying Ding, and Chaoqun Ni. Mining enriched contextual information of scientific collaboration: a meso perspective. Journal of the American Society for Information Science and Technology: JASIST, 62(5):831–845, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

He:2009:ICD

[He09] Zi-Lin He. International collaboration does not have greater epistemic authority. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2151– 2164, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heath:2005:BRB

[Hea05] Angela Heath. Book review: Information representation and retrieval in the digital age. Journal of the American Society for Information Science and Technology: JASIST, 56(2): 215–216, January 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heinstrom:2006:BEP

[Hei06] Jannica Heinström. Broad exploration or precise specificity: Two basic information seeking patterns among students. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1440–1450, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hellqvist:2010:RHI

[Hel10] Björn Hellqvist. Referencing in the humanities and its implications for citation analysis. Journal of the American Society for Information Science and Technology: JASIST, 61(2): 310–318, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hemminger:2005:IIS

[Hem05] Bradley M. Hemminger. Introduction: Introduction to the special issue on bioinformatics. Journal of the American Society for Information Science and Technology: JASIST, 56(5):437–439, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hemphill:2007:BRB

[Hem07] Libby Hemphill. Book review: Human–machine reconfigurations: Plans and situated actions, 2nd ed. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2404–2405, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hendry:2006:PSU

[Hen06] David G. Hendry. Perspectives on search user interfaces: Best practices and future visions: Workspaces for search. Journal of the American Society for Information Science and Technology: JASIST, 57(6):800–802, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hennemann:2012:EPG

[Hen12] Stefan Hennemann. Evaluating the performance of geographical locations within scientific networks using an aggregation — randomization — re-sampling approach (ARR). Journal of the American Society for Information Science and Technology: JASIST, 63(12):2393–2404, December 2012.

CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heneberg:2013:LFS

[Hen13a] Petr Heneberg. Lifting the fog of scientometric research artifacts: On the scientometric analysis of environmental tobacco smoke research. Journal of the American Society for Information Science and Technology: JASIST, 64(2):334– 344, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heneberg:2013:SUA

[Hen13b] Petr Heneberg. Supposedly uncited articles of Nobel laureates and Fields medalists can be prevalently attributed to the errors of omission and commission. Journal of the American Society for Information Science and Technology: JA-SIST, 64(3):448–454, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hepworth:2004:FUU

[Hep04] Mark Hepworth. A framework for understanding user requirements for an information service: Defining the needs of informal carers. Journal of the American Society for Information Science and Technology: JASIST, 55(8):695–708, June 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2010:TBV

[HFC10] Chunneng Huang, Tianjun Fu, and Hsinchun Chen. Textbased video content classification for online video-sharing sites. Journal of the American Society for Information Science and Technology: JASIST, 61(5):891–906, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

He:2009:RTP

[HFW09] Wei He, Yulin Fang, and Kwok-Kee Wei. The role of trust in promoting organizational knowledge seeking using knowledge management systems: an empirical investigation. Journal of the American Society for Information Science and Technology: JASIST, 60(3):526–537, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hembrooke:2005:EEF

[HGGL05] Helene A. Hembrooke, Laura A. Granka, Geraldine K. Gay, and Elizabeth D. Liddy. The effects of expertise and feedback on search term selection and subsequent learning. Journal of the American Society for Information Science and Technology: JASIST, 56(8):861–871, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heo:2001:REC

[HH01] Misook Heo and Stephen C. Hirtle. Research: An empirical comparison of visualization tools to assist information retrieval on the web. Journal of the American Society for Information Science and Technology: JASIST, 52(8):666– 675, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hew:2007:KSO

[HH07] Khe Foon Hew and Noriko Hara. Knowledge sharing in online environments: a qualitative case study. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2310–2324, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hallonsten:2013:ASS

[HH13] Olof Hallonsten and Daniel Holmberg. Analyzing structural stratification in the Swedish higher education system: Data contextualization with policy-history analysis. Journal of the American Society for Information Science and Technology: JASIST, 64(3):574–586, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hu:2010:ASE

[HHfHC10] Paul Jen-Hwa Hu, Fang-Ming Hsu, Han fen Hu, and Hsunchun Chen. Agency satisfaction with electronic record management systems: a large-scale survey. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2559–2574, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heinze:2013:NPS

[HHHR13] Thomas Heinze, Richard Heidler, Raphael Heiko Heiberger, and Jan Riebling. New patterns of scientific growth: How research expanded after the invention of scanning tunneling microscopy and the discovery of Buckminsterfullerenes. Journal of the American Society for Information Science and Technology: JASIST, 64(4):829–843, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hirvonen:2012:IBS

[HHNK12] Noora Hirvonen, Maija-Leena Huotari, Raimo Niemelä, and Raija Korpelainen. Information behavior in stages of exercise behavior change. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1804– 1819, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huurnink:2010:SBM

[HHvdHdR10] Bouke Huurnink, Laura Hollink, Wietske van den Heuvel, and Maarten de Rijke. Search behavior of media professionals at an audiovisual archive: a transaction log analysis. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1180–1197, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2008:CRO

[hHwC08] Mu hsuan Huang and Yu wei Chang. Characteristics of research output in social sciences and humanities: From a research evaluation perspective. Journal of the American Society for Information Science and Technology: JASIST, 59 (11):1819–1828, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Havemann:2005:FLB

[HHWD05] Frank Havemann, Michael Heinz, and Roland Wagner-Döbler. Firm-like behavior of journals? Scaling properties of their output and impact growth dynamics. Journal of the American Society for Information Science and Technology: JASIST, 56(1):3–12, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2004:IDP

[hHyW04] Mu hsuan Huang and Hui yu Wang. The influence of document presentation order and number of documents judged on users’ judgments of relevance. Journal of the American Society for Information Science and Technology: JASIST, 55(11):970–979, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hill:2011:BRB

[Hil11] Heather L. Hill. Book review: Critical Theory for Library and Information Science Exploring the Social from Across the Disciplines. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1425–1426, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Himma:2003:BRB

[Him03] Kenneth Einar Himma. Book review: Beyond our control? Confronting the limits of our legal system in the age of cyberspace. Journal of the American Society for Information Science and Technology: JASIST, 54(6):589–591, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Himma:2004:BRB

[Him04] Kenneth Einar Himma. Book review: The library’s legal answer book. Journal of the American Society for Information Science and Technology: JASIST, 55(3):276–277, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Himma:2008:PGI

[Him08] Kenneth Einar Himma. Perspectives on global information ethics: The justification of intellectual property: Contemporary philosophical disputes. Journal of the American Society for Information Science and Technology: JASIST, 59(7): 1143–1161, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Holsapple:2004:FKM

[HJ04] C. W. Holsapple and K. D. Joshi. A formal knowledge management ontology: Conduct, activities, resources, and influences. Journal of the American Society for Information Science and Technology: JASIST, 55(7):593–612, May 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2013:CUT

[HJ13] Hong Huang and Corinne Jörgensen. Characterizing user tagging and co-occurring metadata in general and specialized metadata collections. Journal of the American Society for Information Science and Technology: JASIST, 64 (9):1878–1889, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2001:BRB

[Hjø01a] Birger Hjørland. Book review: Digital libraries, edited by William Y. Arms. Journal of the American Society for Information Science and Technology: JASIST, 52(2):183, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2001:BCT

[Hjø01b] Birger Hjørland. Brief communication: Towards a theory of aboutness, subject, topicality, theme, domain, field, content … and relevance. Journal of the American Society for Information Science and Technology: JASIST, 52(9):774–778, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2001:LEW

[Hjø01c] Birger Hjørland. Letters to the Editor: Why is meta analysis neglected by information scientists? Journal of the American Society for Information Science and Technology: JA-SIST, 52(13):1193–1194, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2001:RLL

[Hjø01d] Birger Hjørland. Research letter: Letter to the Editor. Journal of the American Society for Information Science and Technology: JASIST, 52(8):697, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2002:BRB

[Hjø02a] Birger Hjørland. Book review: The future of classification, edited by Rita Marcella and Arthur Maltby. Journal of the American Society for Information Science and Technology: JASIST, 53(1):57, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2002:ESC

[Hjø02b] Birger Hjørland. Epistemology and the socio-cognitive perspective in information science. Journal of the American Society for Information Science and Technology: JASIST, 53 (4):257–270, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2002:LES

[Hjø02c] Birger Hjørland. Letter to the Editor: The special competency of information specialists. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1275–1276, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2002:LEM

[Hjø02d] Birger Hjørland. Letter to the Editors: Meta-analysis should also be visible inside information science. Journal of the American Society for Information Science and Technology: JASIST, 53(4):324, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2003:LEA

[Hjø03] Birger Hjørland. Letters to the Editor: Arguments for epistemology in information science. Journal of the American Society for Information Science and Technology: JASIST, 54(8):805–806, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2007:IOS

[Hjø07] Birger Hjørland. Information: Objective or subjective/ situational? Journal of the American Society for Information Science and Technology: JASIST, 58(10):1448–1456, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2009:CT

[Hjø09a] Birger Hjørland. Concept theory. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1519–1536, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2009:LED

[Hjø09b] Birger Hjørland. Letter to the Editor: Does JASIST have an archive? Journal of the American Society for Information Science and Technology: JASIST, 60(2):432, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2009:LEC

[Hjø09c] Birger Hjørland. Letter to the Editor: The controversy over the concept of “information”: a rejoinder to Professor Bates. Journal of the American Society for Information Science and Technology: JASIST, 60(3):643, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2010:FCR

[Hjø10a] Birger Hjørland. The foundation of the concept of relevance. Journal of the American Society for Information Science and Technology: JASIST, 61(2):217–237, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2010:LEA

[Hjø10b] Birger Hjørland. Letters to the Editor: Answer to Professor Szostak (concept theory). Journal of the American Society for Information Science and Technology: JASIST, 61(5): 1078–1080, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2011:EIS

[Hjø11a] Birger Hjørland. Evaluation of an information source illustrated by a case study: Effect of screening for breast cancer. Journal of the American Society for Information Science and Technology: JASIST, 62(10):1892–1898, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2011:EBP

[Hjø11b] Birger Hjørland. Evidence-based practice: an analysis based on the philosophy of science. Journal of the American Society for Information Science and Technology: JASIST, 62 (7):1301–1310, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2011:ITKb

[Hjø11c] Birger Hjørland. The importance of theories of knowledge: Browsing as an example. Journal of the American Society for Information Science and Technology: JASIST, 62(3):594– 603, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hjorland:2011:ITKa

[Hjø11d] Birger Hjørland. The importance of theories of knowledge: Indexing and information retrieval as an example. Journal of the American Society for Information Science and Technology: JASIST, 62(1):72–77, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

HaCohen-Kerner:2010:CCU

[HKBY^{+}10] Yaakov HaCohen-Kerner, Hananya Beck, Elchai Yehudai, Mordechay Rosenstein, and Dror Mughaz. Cuisine: Classification using stylistic feature sets and/or name-based feature sets. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1644–1657, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hu:2009:ISF

[HKC09] Daning Hu, Siddharth Kaza, and Hsinchun Chen. Identifying significant facilitators of dark network evolution. Journal of the American Society for Information Science and Technology: JASIST, 60(4):655–665, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

HaCohen-Kerner:2010:HHA

[HKKP10] Yaakov HaCohen-Kerner, Ariel Kass, and Ariel Peretz. HAADS: A Hebrew Aramaic abbreviation disambiguation system. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1923–1932, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

HaCohen-Kerner:2013:IDM

[HKKP13] Yaakov HaCohen-Kerner, Ariel Kass, and Ariel Peretz. Initialism disambiguation: Man versus machine. Journal of the American Society for Information Science and Technology: JASIST, 64(10):2133–2148, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Haynes:2005:OAT

[HKLG05] Steven R. Haynes, Thomas George Kannampallil, Lawrence L. Larson, and Nitesh Garg. Optimizing anti-terrorism resource allocation. Journal of the American Society for Information Science and Technology: JASIST, 56(3):299–309, February 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heinrichs:2005:MOK

[HL05] John H. Heinrichs and Jeen-Su Lim. Model for organizational knowledge creation and strategic use of information. Journal of the American Society for Information Science and Technology: JASIST, 56(6):620–629, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hemminger:2005:SBP

[HLB05] Bradley M. Hemminger, Trish Losi, and Anne Bauers. Survey of bioinformatics programs in the United States. Journal of the American Society for Information Science and Technology: JASIST, 56(5):529–537, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hawkins:2003:ISA

[HLC03] Donald T. Hawkins, Signe E. Larson, and Bari Q. Caton. Information science abstracts: Tracking the literature of information science. Part 2: a new taxonomy for information science. Journal of the American Society for Information Science and Technology: JASIST, 54(8):771–781, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hu:2005:UAI

[HLC05] Paul Jen-Hwa Hu, Chienting Lin, and Hsinchun Chen. User acceptance of Intelligence and Security Informatics technology: a study of COPLINK. Journal of the American Society for Information Science and Technology: JASIST, 56(3): 235–244, February 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2011:CMC

[HLC11] Mu-Hsuan Huang, Chi-Shiou Lin, and Dar-Zen Chen. Counting methods, country rank changes, and counting inflation in the assessment of national research productivity and impact. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2427–2436, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heinrichs:2007:DFA

[HLLS07] John H. Heinrichs, Kee-Sook Lim, Jeen-Su Lim, and Melissa Allen Spangenberg. Determining factors of academic library Web site usage. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2325– 2334, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Haus:2004:SDA

[HLP04] Goffredo Haus, Maurizio Longari, and Emanuele Pollastri. A score-driven approach to music information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 55(12):1045–1052, October 2004. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hu:2013:CEG

[HLP13] Guangwei Hu, Hui Lin, and Wenwen Pan. Conceptualizing and examining E-government service capability: A Review and empirical study. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2379– 2395, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hawes:2009:ECM

[HLR09] Timothy Hawes, Jimmy Lin, and Philip Resnik. Elements of a computational model for multi-party discourse: the turntaking behavior of Supreme Court justices. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1607–1615, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hemminger:2007:ISB

[HLVA07] Bradley M. Hemminger, Dihui Lu, K. T. L. Vaughan, and Stephanie J. Adams. Information seeking behavior of academic scientists. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2205–2225, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hinich:2003:PIF

[HM03] Melvin J. Hinich and Robert E. Molyneux. Predicting information flows in network traffic. Journal of the American Society for Information Science and Technology: JASIST, 54(2):161–168, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hubert:2009:ASE

[HM09] Gilles Hubert and Josiane Mothe. An adaptable search engine for multimodal information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1625–1634, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Halevi:2013:TCF

[HM13] Gali Halevi and Henk F. Moed. The thematic and conceptual flow of disciplinary research: a citation context analysis of the journal of informetrics, 2007. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1903–1913, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

He:2011:RDB

[HMdR11] Jiyin He, Edgar Meij, and Maarten de Rijke. Result diversification based on query-specific cluster ranking. Journal of the American Society for Information Science and Technology: JASIST, 62(3):550–571, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2012:LCB

[HMFW12] Lan Huang, David Milne, Eibe Frank, and Ian H. Witten. Learning a concept-based document similarity measure. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1593–1608, August 2012.

CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heidorn:2002:CUC

[HML02] P. Bryan Heidorn, Bharat Mehra, and Mary F. Lokhaiser. Complementary user-centered methodologies for information seeking and use: System’s design in the Biological Information Browsing Environment (BIBE). Journal of the American Society for Information Science and Technology: JASIST, 53(14):1251–1258, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Humphrey:2009:CRB

[HNB^{+}09] Susanne M. Humphrey, Aurélie Névéol, Allen Browne, Julien Gobeil, Patrick Ruch, and Stéfan J. Darmoni. Comparing a rule-based versus statistical system for automatic categorization of MEDLINE documents according to biomedical specialty. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2530–2539, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huntington:2006:ADD

[HNJT06] Paul Huntington, David Nicholas, Hamid R. Jamali, and Carol Tenopir. Article decay in the digital environment: an analysis of usage of OhioLINK by date of publication, employing deep log methods. Journal of the American Society for Information Science and Technology: JASIST, 57(13): 1840–1851, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ho:2008:BRB

[Ho08] Shuk Ying Ho. Book review: Human-computer interaction and management information systems: Foundations. Journal of the American Society for Information Science and Technology: JASIST, 59(12):2034–2035, October 2008. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Holsapple:2009:HMW

[HO09] Clyde W. Holsapple and Daniel O’Leary. How much and where? Private versus public universities’ publication patterns in the information systems discipline. Journal of the American Society for Information Science and Technology: JASIST, 60(2):318–331, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hoenkamp:2003:UOD

[Hoe03] Eduard Hoenkamp. Unitary operators on the document space. Journal of the American Society for Information Science and Technology: JASIST, 54(4):314–320, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Holbrooks:2001:BRB

[Hol01] M. Zoe Holbrooks. Book review: High technology and lowincome communities. Prospects for the positive use of advanced information technology, edited by Donald A. Schön, Bish Sanyal, and William J. Mitchell. Journal of the American Society for Information Science and Technology: JA-SIST, 52(3):275–277, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Holbrooks:2004:BRB

[Hol04] Zoe Holbrooks. Book review: The ABCs of XML: the librarian’s guide to the eXtensible Markup Language; XML in libraries; and Learning XML. Journal of the American Society for Information Science and Technology: JASIST, 55 (14):1304–1305, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Holsapple:2008:PPA

[Hol08] Clyde W. Holsapple. A publication power approach for identifying premier information systems journals. Journal of the American Society for Information Science and Technology: JASIST, 59(2):166–185, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hong:2006:ISM

[Hon06] Traci Hong. The influence of structural and message features on Web site credibility. Journal of the American Society for Information Science and Technology: JASIST, 57(1):114– 127, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hong:2012:IHS

[Hon12] Traci Hong. Internet health search: When process complements goals. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2283–2293, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hovden:2013:BIM

[Hov13] Robert Hovden. Bibliometrics for Internet media: Applying the h-index to YouTube. Journal of the American Society for Information Science and Technology: JASIST, 64(11): 2326–2331, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2004:DWL

[HPAS04] Xiangji Huang, Fuchun Peng, Aijun An, and Dale Schuurmans. Dynamic Web log session identification with statistical language models. Journal of the American Society for Information Science and Technology: JASIST, 55(14):1290– 1303, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Harrison:2007:GIT

[HPGG^{+}07] Teresa M. Harrison, Theresa Pardo, J. Ramon Gil-Garcia, Fiona Thompson, and Dubravka Juraga. Geographic information technologies, structuration theory, and the World Trade Center crisis. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2240– 2254, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Herranz:2012:MFS

[HRC12a] Neus Herranz and Javier Ruiz-Castillo. Multiplicative and fractional strategies when journals are assigned to several subfields. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2195–2205, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hu:2012:NAM

[HRC12b] Xiaojun Hu, Ronald Rousseau, and Jin Chen. A new approach for measuring the value of patents based on structural indicators for ego patent citation networks. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1834–1842, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Humphrey:2006:WSD

[HRK^{+}06] Susanne M. Humphrey, Willie J. Rogers, Halil Kilicoglu, Dina Demner-Fushman, and Thomas C. Rindflesch. Word sense disambiguation by selecting the best semantic type based on Journal Descriptor Indexing: Preliminary experiment. Journal of the American Society for Information Science and Technology: JASIST, 57(1):96–113, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Ano06b].

Hochheiser:2001:RUI

[HS01] Harry Hochheiser and Ben Shneiderman. Research: Using interactive visualizations of WWW log data to characterize access patterns and inform site design. Journal of the American Society for Information Science and Technology: JASIST, 52(4):331–343, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

He:2002:CFA

[HS02] Shaoyi He and Amanda Spink. A comparison of foreign authorship distribution in JASIST and the Journal of Documentation. Journal of the American Society for Information Science and Technology: JASIST, 53(11):953–959, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hogan:2013:OPS

[HS13a] Niamh M. Hogan and Karl J. Sweeney. Opinion paper: Social networking and scientific communication: a paradoxical return to Mertonian roots? Journal of the American Society for Information Science and Technology: JASIST, 64 (3):644–646, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2013:RIF

[HS13b] Xiaoli Huang and Dagobert Soergel. Relevance: an improved framework for explicating the notion. Journal of the American Society for Information Science and Technology: JA-SIST, 64(1):18–35, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2007:CWU

[HSCL07] Chun-Yao Huang, Yung-Cheng Shen, I-Ping Chiang, and Chen-Shun Lin. Characterizing Web users’ online information behavior. Journal of the American Society for Information Science and Technology: JASIST, 58(13):1988–1997, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hara:2010:CCA

[HSH10] Noriko Hara, Pnina Shachaf, and Khe Foon Hew. Crosscultural analysis of the Wikipedia community. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2097–2108, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2012:PDQ

[HSJB12] Hong Huang, Besiki Stvilia, Corinne Jörgensen, and Hank W. Bass. Prioritization of data quality dimensions and skills requirements in genome annotation work. Journal of the American Society for Information Science and Technology: JASIST, 63(1):195–207, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hara:2003:EVS

[HSKS03] Noriko Hara, Paul Solomon, Seung-Lye Kim, and Diane H. Sonnenwald. An emerging view of scientific collaboration: Scientists’ perspectives on collaboration and factors that impact collaboration. Journal of the American Society for Information Science and Technology: JASIST, 54(10):952– 965, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hemminger:2007:CFT

[HSSV07] Bradley M. Hemminger, Billy Saelim, Patrick F. Sullivan, and Todd J. Vision. Comparison of full-text searching to metadata searching for genes in two biomedical literature cohorts. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2341–2352, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hartel:2011:VAP

[HT11] Jenna Hartel and Leslie Thomson. Visual approaches and photography for the study of immediate information space. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2214–2224, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2011:IPP

[HTC11] Mu-Hsuan Huang, Muh-Chyun Tang, and Dar-Zen Chen. Inequality of publishing performance and international collaboration in physics. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1156– 1165, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hollink:2011:SSL

[HTdV11] Vera Hollink, Theodora Tsikrika, and Arjen P. de Vries. Semantic search log analysis: a method and a study on professional image search. Journal of the American Society for Information Science and Technology: JASIST, 62(4):691– 713, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hong:2007:HDW

[HTT07] Weiyin Hong, James Y. L. Thong, and Kar Yan Tam. How do Web users respond to non-banner-ads animation? The effects of task type and user experience. Journal of the American Society for Information Science and Technology: JA-SIST, 58(10):1467–1482, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hu:2009:BCL

[Hu09] Xiaojun Hu. Brief communication: Loads of special authorship functions: Linear growth in the percentage of “equal first authors” and corresponding authors. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2378–2381, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huang:2013:AGF

[Hua13] Xiaoli Huang. Applying a generic function-based topical relevance typology to structure clinical questions and answers. Journal of the American Society for Information Science and Technology: JASIST, 64(1):65–85, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huber:2001:NMA

[Hub01] John C. Huber. A new method for analyzing scientific productivity. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1089–1099, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huber:2002:NMG

[Hub02] John C. Huber. A new model that generates Lotka’s Law. Journal of the American Society for Information Science and Technology: JASIST, 53(3):209–219, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hutchinson:2008:BRBa

[Hut08a] Alvin Hutchinson. Book review: Evaluations of individual scientists and research institutions. Journal of the American Society for Information Science and Technology: JASIST, 59(1):160–161, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hutchinson:2008:BRBb

[Hut08b] Alvin Hutchinson. Book review: Laughing at the CIO: a parable and prescription for IT leadership. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1694–1695, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huvila:2010:ISP

[Huv10] Isto Huvila. Information sources and perceived success in corporate finance. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2219– 2229, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huvila:2011:PBO

[Huv11] Isto Huvila. The politics of boundary objects: Hegemonic interventions and the making of a document. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2528–2539, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huvila:2013:HMK

[Huv13] Isto Huvila. How a museum knows? Structures, work roles, and infrastructures of information work. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1375–1387, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Herrera-Viedma:2001:RMR

[HV01] E. Herrera-Viedma. Research: Modeling the retrieval process for an information retrieval system using an ordinal fuzzy linguistic approach. Journal of the American Society for Information Science and Technology: JASIST, 52 (6):460–475, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Harzing:2009:GSI

[HvdW09] Anne-Wil Harzing and Ron van der Wal. A Google Scholar h-index for journals: an alternative metric to measure journal impact in economics and business. Journal of the American Society for Information Science and Technology: JA-SIST, 60(1):41–46, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Herrera-Viedma:2006:SAI

[HVP06] Enrique Herrera-Viedma and Gabriella Pasi. Soft approaches to information retrieval and information access on the Web: an introduction to the special topic section. Journal of the American Society for Information Science and Technology: JASIST, 57(4):511–514, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Herrera-Viedma:2006:EIQ

[HVPLHP06] Enrique Herrera-Viedma, Gabriella Pasi, Antonio G. Lopez-Herrera, and Carlos Porcel. Evaluating the information quality of Web sites: a methodology based on fuzzy computing with words. Journal of the American Society for Information Science and Technology: JASIST, 57(4):538–549, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hood:2001:SDD

[HW01] William W. Hood and Concepción S. Wilson. The scatter of documents over databases in different subject domains: How many databases are needed? Journal of the American Society for Information Science and Technology: JASIST, 52(14):1242–1254, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hauck:2002:WBI

[HW02] Roslin V. Hauck and Suzanne Weisband. When a better interface and easy navigation aren’t enough: Examining the information architecture in a law enforcement agency. Journal of the American Society for Information Science and Technology: JASIST, 53(10):846–854, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hood:2003:OBD

[HW03] William W. Hood and Concepción S. Wilson. Overlap in bibliographic databases. Journal of the American Society for Information Science and Technology: JASIST, 54(12): 1091–1103, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hood:2005:BCR

[HW05] William Wellesley Hood and Concepción Shimizu Wilson. Brief communication: The relationship of records in multiple databases to their usage or citedness. Journal of the American Society for Information Science and Technology: JA-SIST, 56(9):1004–1007, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hicks:2011:CON

[HW11] Diana Hicks and Jian Wang. Coverage and overlap of the new social sciences and humanities journal lists. Journal of the American Society for Information Science and Technology: JASIST, 62(2):284–294, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Houston:2013:IBM

[HW13] Ron Houston and Lynn Westbrook. Information-based mitigation of intimate partner violence. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1694–1706, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Huber:2003:BCU

[HWD03] John C. Huber and Roland Wagner-Döbler. Brief communications: Using the Mann–Whitney Test on informetric data. Journal of the American Society for Information Science and Technology: JASIST, 54(8):798–801, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hughes:2010:DOI

[HWJ10] Benjamin Hughes, Jonathan Wareham, and Indra Joshi. Doctors’ online information needs, cognitive search strategies, and judgments of information quality and cognitive authority: How predictive judgments introduce bias into cognitive search models. Journal of the American Society for Information Science and Technology: JASIST, 61(3):433– 452, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

He:2013:SAP

[HWT^{+}13] Ru He, Jiong Wang, Jin Tian, Cheng-Tao Chu, Bradley Mauney, and Igor Perisic. Session analysis of people search within a professional social network. Journal of the American Society for Information Science and Technology: JA-SIST, 64(5):929–950, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hoeber:2009:HSV

[HY09] Orland Hoeber and Xue Dong Yang. HotMap: Supporting visual exploration of Web search results. Journal of the American Society for Information Science and Technology: JASIST, 60(1):90–110, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hyland:2003:SCS

[Hyl03] Ken Hyland. Self-citation and self-reference: Credibility and promotion in academic publication. Journal of the American Society for Information Science and Technology: JASIST, 54(3):251–259, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heinz:2003:ESP

[HZ03a] Steffen Heinz and Justin Zobel. Efficient single-pass index construction for text databases. Journal of the American Society for Information Science and Technology: JASIST, 54(8):713–729, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hoad:2003:MIV

[HZ03b] Timothy C. Hoad and Justin Zobel. Methods for identifying versioned and plagiarized documents. Journal of the American Society for Information Science and Technology: JA-SIST, 54(3):203–215, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Hawking:2007:DTM

[HZ07] David Hawking and Justin Zobel. Does topic metadata help with Web search? Journal of the American Society for Information Science and Technology: JASIST, 58(5):613–628, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Heverin:2012:UMC

[HZ12] Thomas Heverin and Lisl Zach. Use of microblogging for collective sense-making during violent crises: a study of three campus shootings. Journal of the American Society for Information Science and Technology: JASIST, 63(1):34–47, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ibrahim:2012:IST

[IA12] Nurain Hassan Ibrahim and David Allen. Information sharing and trust during major incidents: Findings from the oil industry. Journal of the American Society for Information Science and Technology: JASIST, 63(10):1916–1928, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ferrer-i-Cancho:2009:FSF

[iCG09] Ramon Ferrer i Cancho and Ricard Gavaldà. The frequency spectrum of finite samples from the intermittent silence process. Journal of the American Society for Information Science and Technology: JASIST, 60(4):837–843, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Iyer:2007:PSV

[IL07] Hemalata Iyer and Caitlain Devereaux Lewis. Prioritization strategies for video storyboard keyframes. Journal of the American Society for Information Science and Technology: JASIST, 58(5):629–644, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ingwersen:2012:BRB

[Ing12] Peter Ingwersen. Book review: Interactive Information Seeking, Behaviour and Retrieval. Edited by Ian Ruthven and Diane Kelly. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2122–2125, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ivanisevic:2006:MAS

[IS06] Ranka Ivanišević and Damir Sapunar. Multiple authorship in a small medical journal: a case study of the Croatian Medical Journal. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1073–1078, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ibekwe-SanJuan:2012:AIS

[IS12] Fidelia Ibekwe-SanJuan. Advances in information science: The French conception of information science: “Une exception française”? Journal of the American Society for Information Science and Technology: JASIST, 63(9):1693–1709, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ivancheva:2001:NGN

[Iva01] Ludmila E. Ivancheva. The non-Gaussian nature of bibliometric and scientometric distributions: a new approach to interpretation. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1100–1105, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ivory:2003:BRB

[Ivo03] Melody Y. Ivory. Book review: Leonardo’s Laptop: Human Needs and the New Computing Technologies. Journal of the American Society for Information Science and Technology: JASIST, 54(10):977–979, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Iyengar:2001:EVB

[Iye01] S. S. Iyengar. Editorials: Visual based retrieval systems and Web mining — Introduction. Journal of the American Society for Information Science and Technology: JASIST, 52(10):829–830, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jacobs:2001:ITI

[Jac01] Neil Jacobs. Information technology and interests in scholarly communication: a discourse analysis. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1122–1133, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

James:2008:BCR

[Jam08a] Jeffrey James. Brief communication: Re-estimating the difficulty of closing the digital divide. Journal of the American Society for Information Science and Technology: JASIST, 59(12):2024–2032, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

James:2008:DPV

[Jam08b] Jeffrey James. Digital preparedness versus the digital divide: a confusion of means and ends. Journal of the American Society for Information Science and Technology: JASIST, 59(5):785–791, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

James:2010:BRB

[Jam10] Jeffrey James. Book review: Ict4d: Information and Communication Technology for Development. Journal of the American Society for Information Science and Technology: JASIST, 61(1):208–209, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Janes:2002:DRR

[Jan02] Joseph Janes. Digital reference: Reference librarians’ experiences and attitudes. Journal of the American Society for Information Science and Technology: JASIST, 53(7):549– 566, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Japzon:2007:BRB

[Jap07] Andrea Japzon. Book review: What they didn’t tell you about knowledge management. Journal of the American Society for Information Science and Technology: JASIST, 58(12): 1909–1910, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Japzon:2008:BRB

[Jap08] Andrea Japzon. Book review: Theorizing digital cultural heritage: a critical discourse. Journal of the American Society for Information Science and Technology: JASIST, 59(8): 1360–1361, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jarvelin:2007:ATA

[Jär07] Kalervo Järvelin. An analysis of two approaches in information retrieval: From frameworks to study designs. Journal of the American Society for Information Science and Technology: JASIST, 58(7):971–986, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jaric:2011:LEU

[Jar11] Ivan Jarić. Letter to the Editor: The use of h-index for the assessment of journals’ performance will lead to shifts in editorial policies. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2546, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jayroe:2012:HSW

[Jay12] Tina J. Jayroe. A humble servant: the work of Helen L. Brownson and the early years of information science research. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2052–2061, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jones:2006:PSU

[JBCJ06] Matt Jones, George Buchanan, Tzu-Chiang Cheng, and Preeti Jain. Perspectives on search user interfaces: Best practices and future visions: Changing the pace of search: Supporting “background” information seeking. Journal of the American Society for Information Science and Technology: JASIST, 57(6):838–842, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jorge-Botana:2010:VPU

[JBLOHM10] Guillermo Jorge-Botana, José A. León, Ricardo Olmos, and Yusef Hassan-Montero. Visualizing polysemy using LSA and the predication algorithm. Journal of the American Society for Information Science and Technology: JASIST, 61(8): 1706–1724, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jaeger:2004:ECI

[JBM04] Paul T. Jaeger, John Carlo Bertot, and Charles R. McClure. The effects of the Children’s Internet Protection Act (CIPA) in public libraries and its implications for research: a statistical, policy, and legal analysis. Journal of the American Society for Information Science and Technology: JASIST, 55(13):1131–1139, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2009:PQR

[JBS09] Bernard J. Jansen, Danielle L. Booth, and Amanda Spink. Patterns of query reformulation during Web searching. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1358–1371, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jimenez-Contreras:2006:BMJ

[JCDR^{+}06] Evaristo Jiménez-Contreras, Mercedes De La Moneda, Elvira Ruiz de Osma, Rafael Bailón-Moreno, and Rosario Ruiz-Baños. A bibliometric model for journal discarding policy at academic libraries. Journal of the American Society for Information Science and Technology: JASIST, 57(2): 198–207, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jonkers:2012:BBC

[JD12] K. Jonkers and G. E. Derrick. The bibliometric bandwagon: Characteristics of bibliometric articles outside the field literature. Journal of the American Society for Information Science and Technology: JASIST, 63(4):829–836, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Ano12].

Jonkers:2012:MUR

[JdMAA12] Koen Jonkers, Felix de Moya Anegon, and Isidro F. Aguillo. Measuring the usage of e-research infrastructure as an indicator of research activity. Journal of the American Society for Information Science and Technology: JASIST, 63 (7):1374–1382, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jeong:2003:MGI

[JG03] Wooseob Jeong and Myke Gluck. Multimodal geographic information systems: Adding haptic and auditory display. Journal of the American Society for Information Science and Technology: JASIST, 54(3):229–242, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ju:2005:UPM

[JG05] Boryung Ju and Myke Gluck. User-process model approach to improve user interface usability. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1098–1112, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jupp:2006:QFB

[JG06] Julie R. Jupp and John S. Gero. A qualitative featurebased characterization of 2D architectural style. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1537–1550, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Janes:2001:AES

[JHR01] Joseph Janes, Chrystie Hill, and Alex Rolfe. Ask-an-expert services analysis. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1106–1121, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jung:2008:LSD

[JHW^{+}08] Seikyung Jung, Jonathan L. Herlocker, Janet Webster, Margaret Mellinger, and Jeremy Frumkin. LibraryFind: System design and usability testing of academic metasearch system. Journal of the American Society for Information Science and Technology: JASIST, 59(3):375–389, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jorgensen:2005:IQI

[JJ05] Corinne Jörgensen and Peter Jörgensen. Image querying by image professionals. Journal of the American Society for Information Science and Technology: JASIST, 56(12): 1346–1359, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jagtap:2010:RUS

[JJ10] Santosh Jagtap and Aylmer Johnson. Requirements and use of in-service information in an engineering redesign task: Case studies from the aerospace industry. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2442–2460, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jorgensen:2001:CFE

[JJBC01] Corinne Jörgensen, Alejandro Jaimes, Ana B. Benitez, and Shih-Fu Chang. A conceptual framework and empirical research for classifying visual descriptors. Journal of the American Society for Information Science and Technology: JA-SIST, 52(11):938–947, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lin:2005:IFA

[jL05] Shin jeng Lin. Internetworking of factors affecting successive searches over multiple episodes. Journal of the American Society for Information Science and Technology: JASIST, 56(4):416–436, February 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lin:2005:VMI

[jLB05] Shin jeng Lin and Nick Belkin. Validation of a model of information seeking over multiple search sessions. Journal of the American Society for Information Science and Technology: JASIST, 56(4):393–415, February 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Julien:2008:DRC

[JLB08] Charles-Antoine Julien, John E. Leide, and France Bouthillier. In-depth review: Controlled user evaluations of information visualization interfaces for text retrieval: Literature review and meta-analysis. Journal of the American Society for Information Science and Technology: JASIST, 59(6):1012– 1024, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jang:2004:RDM

[JLCL04] Jyh-Shing Roger Jang, Hong-Ru Lee, Jiang-Chuen Chen, and Cheng-Yuan Lin. Research and developments of a multimodal MIR engine for commercial applications in East Asia. Journal of the American Society for Information Science and Technology: JASIST, 55(12):1067–1076, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jeong:2009:YIS

[JLK09] Senator Jeong, Sungin Lee, and Hong-Gee Kim. Are you an invited speaker? A bibliometric analysis of elite groups for scholarly events in bioinformatics. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1118–1131, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jin:2004:LTI

[JLR04] Bihui Jin, Ling Li, and Ronald Rousseau. Long-term influences of interventions in the normal development of science: China and the Cultural Revolution. Journal of the American Society for Information Science and Technology: JA-SIST, 55(6):544–550, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2013:EAR

[JLS13] Bernard J. Jansen, Zhe Liu, and Zach Simon. The effect of ad rank on the performance of keyword advertising campaigns. Journal of the American Society for Information Science and Technology: JASIST, 64(10):2115–2132, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2005:EEP

[JM05] Bernard J. Jansen and Michael D. McNeese. Evaluating the effectiveness of and patterns of interactions with automated searching assistance. Journal of the American Society for Information Science and Technology: JASIST, 56(14):1480– 1503, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Johnson:2001:BRB

[Joh01] Lonnie L. Johnson. Book review: Blown to bits: How the new economics of information transforms strategy, by Philip Evans and Thomas S. Wurster. Journal of the American Society for Information Science and Technology: JASIST, 52(9):780–781, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Johnson:2007:SCS

[Joh07] Catherine A. Johnson. Social capital and the search for information: Examining the role of social capital in information seeking behavior in Mongolia. Journal of the American Society for Information Science and Technology: JASIST, 58 (6):883–894, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jones:2002:BRB

[Jon02] Matt Jones. Book review: Usability testing for library Web sites: a hands-on guide. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1184, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jorgensen:2001:IO

[Jör01] Corinne Jörgensen. Introduction and overview. Journal of the American Society for Information Science and Technology: JASIST, 52(11):906–910, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jorgensen:2004:BCU

[Jör04] Corinne Jörgensen. Brief communications: Unlocking the museum: a manifesto. Journal of the American Society for Information Science and Technology: JASIST, 55(5):462– 464, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jorgensen:2007:PMM

[Jör07] Corinne Jörgensen. Perspectives on MPEG-7: Metadata for multimedia: The MPEG-7 standard: Multimedia description in theory and application. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1323–1328, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2001:RWS

[JP01] Bernard J. Jansen and Udo Pooch. A review of Web searching studies and a framework for future research. Journal of the American Society for Information Science and Technology: JASIST, 52(3):235–246, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jones:2002:AED

[JP02] Steve Jones and Gordon W. Paynter. Automatic extraction of document keyphrases for use in digital libraries: Evaluation and applications. Journal of the American Society for Information Science and Technology: JASIST, 53(8):653– 677, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jarvelin:2008:DID

[JP08a] Kalervo Järvelin and Olle Persson. The DCI index: Discounted cumulated impact-based research evaluation. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1433–1440, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [JP08b].

Jarvelin:2008:ERD

[JP08b] Kalervo Järvelin and Olle Persson. Erratum re: “The DCI-index: Discounted cumulated impact-based research evaluation”, Journal of the American Society for Information Science and Technology, 59(9), 1433–1440. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2350–2352, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [JP08a].

Jansen:2006:ESP

[JR06] Bernard J. Jansen and Marc Resnick. An examination of searcher’s perceptions of nonsponsored and sponsored links during ecommerce Web searching. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1949–1961, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2010:STC

[JR10] Bernard J. Jansen and Soo Young Rieh. The seventeen theoretical constructs of information searching and information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1517–1534, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2007:DSW

[JSBK07] Bernard J. Jansen, Amanda Spink, Chris Blakely, and Sherry Koshman. Defining a session on Web search engines. Journal of the American Society for Information Science and Technology: JASIST, 58(6):862–871, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jepsen:2004:CSW

[JSI^{+}04] Erik Thorlund Jepsen, Piet Seiden, Peter Ingwersen, Lennart Björneborn, and Pia Borlund. Characteristics of scientific Web publications: Preliminary data gathering and analysis. Journal of the American Society for Information Science and Technology: JASIST, 55(14):1239–1249, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2007:WSI

[JSK07] Bernard J. Jansen, Amanda Spink, and Sherry Koshman. Web searcher interaction with the Dogpile.com metasearch engine. Journal of the American Society for Information Science and Technology: JASIST, 58(5):744–755, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2005:TCA

[JSP05] Bernard J. Jansen, Amanda Spink, and Jan Pedersen. A temporal comparison of AltaVista Web searching. Journal of the American Society for Information Science and Technology: JASIST, 56(6):559–570, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jiang:2010:CSB

[JT10] Xing Jiang and Ah-Hwee Tan. CRCTOL: a semantic-based domain ontology learning system. Journal of the American Society for Information Science and Technology: JASIST, 61(1):150–168, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Julien:2013:RST

[JTDG13] Charles-Antoine Julien, Pierre Tirilly, Jesse David Dinneen, and Catherine Guastavino. Reducing subject tree browsing complexity. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2201–2223, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Julien:2012:CTL

[JTLG12] Charles-Antoine Julien, Pierre Tirilly, John E. Leide, and Catherine Guastavino. Constructing a true LCSH tree of a science and engineering collection. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2405–2418, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ju:2007:DDK

[Ju07] Boryung Ju. Does domain knowledge matter: Mapping users’ expertise to their information interactions. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2007–2020, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Julien:2012:BRB

[Jul12] Heidi Julien. Book reviews: Looking for Information: A Survey of Research on Information Seeking, Needs, and Behavior (3rd ed.) — Edited by Donald O. Case. New Directions in Information Behaviour — Edited by Amanda Spink and Jannica Heinström. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2557– 2558, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jordan:2009:AGK

[JW09] Chris Jordan and Carolyn Watters. Addressing gaps in knowledge while reading. Journal of the American Society for Information Science and Technology: JASIST, 60(11): 2255–2268, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2009:BEU

[JZS09] Bernard J. Jansen, Mimi Zhang, and Carsten D. Schultz. Brand and its effect on user perception of search engine performance. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1572–1595, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jansen:2009:TPT

[JZSC09] Bernard J. Jansen, Mimi Zhang, Kate Sobel, and Abdur Chowdury. Twitter power: Tweets as electronic word of mouth. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2169–2188, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Jiang:2011:ADC

[JZW^{+}11] Yong Jiang, Hai-Tao Zheng, Xinmin Wang, Binggan Lu, and Kaihua Wu. Affiliation disambiguation for constructing semantic digital libraries. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1029– 1041, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2002:CTI

[KA02] Kyung-Sun Kim and Bryce Allen. Cognitive and task influences on Web searching behavior. Journal of the American Society for Information Science and Technology: JASIST, 53(2):109–119, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kruschwitz:2005:UWM

[KAB05] Udo Kruschwitz and Hala Al-Bakour. Users want more sophisticated search assistants: Results of a task-based evaluation. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1377–1393, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Katsaros:2009:FIQ

[KAB09] Dimitrios Katsaros, Leonidas Akritidis, and Panayiotis Bozanis. The f index: Quantifying the impact of coterminal citations on scientists’ ranking. Journal of the American Society for Information Science and Technology: JASIST, 60 (5):1051–1056, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koppel:2006:FIC

[KAD06] Moshe Koppel, Navot Akiva, and Ido Dagan. Feature instability as a criterion for selecting potential style markers. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1519–1525, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kaki:2006:FBK

[Käk06] Mika Käki. fKWIC: Frequency-based keyword-in-context index for filtering Web search results. Journal of the American Society for Information Science and Technology: JA-SIST, 57(12):1606–1615, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kalbach:2001:BRB

[Kal01] James Kalbach. Book review: Learning XML, by Erik T. Ray. Journal of the American Society for Information Science and Technology: JASIST, 52(12):1082–1084, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kalbach:2005:BRB

[Kal05] James Kalbach. Book review: Understanding information systems: What they do and why we need them. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1463–1464, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kalbach:2006:PSU

[Kal06] James Kalbach. Perspectives on search user interfaces: Best practices and future visions: “i’m feeling lucky”: the role of emotions in seeking information on the Web. Journal of the American Society for Information Science and Technology: JASIST, 57(6):813–818, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kan:2004:BR

[Kan04] Min-Yen Kan. Book review. Journal of the American Society for Information Science and Technology: JASIST, 55(2): 178–179, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kapucu:2009:BRB

[Kap09] Aysegul Kapucu. Book review: American libraries and the Internet: The Social Construction of Web Appropriation and Use. Journal of the American Society for Information Science and Technology: JASIST, 60(2):430–431, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Karakos:2003:GEI

[Kar03] Alexandros Karakos. Greeklish: an experimental interface for automatic transliteration. Journal of the American Society for Information Science and Technology: JASIST, 54 (11):1069–1074, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Karamuftuoglu:2006:IAI

[Kar06] Murat Karamuftuoglu. Information arts and information science: Time to unite? Journal of the American Society for Information Science and Technology: JASIST, 57(13): 1780–1793, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Karamuftuoglu:2007:NST

[Kar07] Murat Karamuftuoglu. Need for a systemic theory of classification in information science. Journal of the American Society for Information Science and Technology: JASIST, 58(13):1977–1987, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Karamuftuoglu:2009:SLI

[Kar09] Murat Karamuftuoglu. Situating logic and information in information science. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2019–2031, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kanaan:2009:CTC

[KASGAM09] Ghassan Kanaan, Riyad Al-Shalabi, Sameh Ghwanmeh, and Hamda Al-Ma’adeed. A comparison of text-classification techniques applied to Arabic text. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1836–1844, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kazmer:2010:DDR

[Kaz10] Michelle M. Kazmer. Disengaging from a distributed research project: Refining a model of group departures. Journal of the American Society for Information Science and Technology: JASIST, 61(4):758–771, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kazmer:2011:BRB

[Kaz11] Michelle M. Kazmer. Book review: Digital Dead End: Fighting for Social Justice in the Information Age. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2296–2297, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kostoff:2005:FMT

[KB05] Ronald N. Kostoff and Joel A. Block. Factor matrix text filtering and clustering. Journal of the American Society for Information Science and Technology: JASIST, 56(9):946– 968, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Klavans:2006:IBM

[KB06] Richard Klavans and Kevin W. Boyack. Identifying a better measure of relatedness for mapping science. Journal of the American Society for Information Science and Technology: JASIST, 57(2):251–263, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Klavans:2009:TCM

[KB09] Richard Klavans and Kevin W. Boyack. Toward a consensus map of science. Journal of the American Society for Information Science and Technology: JASIST, 60(3):455–476, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Klavans:2011:UGM

[KB11] Richard Klavans and Kevin W. Boyack. Using global mapping to create more accurate document-level maps of research fields. Journal of the American Society for Information Science and Technology: JASIST, 62(1):1–18, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kantor:2012:PDM

[KBG12] Paul B. Kantor, Jonathan M. Bullinger, and Cecilia S. Gal. Patient decision-making modes and causes: a preliminary investigation. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1339–1349, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2010:HIN

[KBP10] Jang Hyun Kim, George A. Barnett, and Han Woo Park. A hyperlink and issue network analysis of the United States Senate: a rediscovery of the Web as a relational and topical medium. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1598–1611, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ku:2007:MWR

[KC07] Lun-Wei Ku and Hsin-Hsi Chen. Mining Web resources for enhancing information retrieval: Mining opinions from the Web: Beyond relevance retrieval. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1838–1850, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kwakkel:2009:MPS

[KC09] Jan H. Kwakkel and Scott W. Cunningham. Managing polysemy and synonymy in science mapping using the mixtures of factor analyzers model. Journal of the American Society for Information Science and Technology: JASIST, 60(10): 2064–2078, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kules:2012:ITS

[KC12] Bill Kules and Robert Capra. Influence of training and stage of search on gaze behavior in a library catalog faceted search interface. Journal of the American Society for Information Science and Technology: JASIST, 63(1):114–138, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2013:CKP

[KC13] Sungwon Kim and Seongyun Cho. Characteristics of Korean personal names. Journal of the American Society for Information Science and Technology: JASIST, 64(1):86–95, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Keikha:2012:EDD

[KCC12] Mostafa Keikha, Fabio Crestani, and Mark James Carman. Employing document dependency in blog search. Journal of the American Society for Information Science and Technology: JASIST, 63(2):354–365, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kabel:2004:AVT

[KdHWA04] Suzanne Kabel, Robert de Hoog, Bob J. Wielinga, and Anjo Anjewierden. The added value of task and ontology-based markup for information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 55(4):348–362, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Khoo:2002:USC

[KDL02] Christopher S. G. Khoo, Yubin Dai, and Teck Ee Loh. Using statistical and contextual information to identify twoand three-character words in Chinese text. Journal of the American Society for Information Science and Technology: JASIST, 53(5):365–377, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kostoff:2001:CMI

[KdRH^{+}01] Ronald N. Kostoff, J. Antonio del Río, James A. Humenik, Esther Ofilia García, and Ana María Ramírez. Citation mining: Integrating text mining and bibliometrics for research user profiling. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1148–1156, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kraft:2008:EFB

[KE08] Donald H. Kraft and Editor. Editorial: Final JASIST editorial. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2198, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kurtz:2005:WUI

[KEA^{+}05a] Michael J. Kurtz, Guenther Eichhorn, Alberto Accomazzi, Carolyn Grant, Markus Demleitner, and Stephen S. Murray.

Worldwide use and impact of the NASA Astrophysics Data System digital library. Journal of the American Society for Information Science and Technology: JASIST, 56(1):36–45, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kurtz:2005:BPA

[KEA^{+}05b] Michael J. Kurtz, Guenther Eichhorn, Alberto Accomazzi, Carolyn Grant, Markus Demleitner, Stephen S. Murray, Nathalie Martimbeau, and Barbara Elwell. The bibliometric properties of article readership information. Journal of the American Society for Information Science and Technology: JASIST, 56(2):111–128, January 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kelly:2006:MOIa

[Kel06a] Diane Kelly. Measuring online information seeking context, Part 1: Background and method. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1729–1739, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kelly:2006:MOIb

[Kel06b] Diane Kelly. Measuring online information seeking context, Part 2: Findings and discussion. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1862–1874, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kelton:2008:TDI

[KFW08] Kari Kelton, Kenneth R. Fleischmann, and William A. Wallace. Trust in digital information. Journal of the American Society for Information Science and Technology: JASIST, 59(3):363–374, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kirchik:2012:CPL

[KGL12] Olessia Kirchik, Yves Gingras, and Vincent Larivière. Changes in publication languages and citation practices and their effect on the scientific impact of Russian science (1993– 2010). Journal of the American Society for Information Science and Technology: JASIST, 63(7):1411–1419, July 2012.

CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kari:2007:IHT

[KH07] Jarkko Kari and Jenna Hartel. Information and higher things in life: Addressing the pleasurable and the profound in information science. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1131– 1147, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2009:RTB

[KH09] Byoungsoo Kim and Ingoo Han. The role of trust belief and its antecedents in a community-driven knowledge environment. Journal of the American Society for Information Science and Technology: JASIST, 60(5):1012–1026, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kozak:2012:PNV

[KH12] Marcin Kozak and James Hartley. Presenting numerical values within sentences and text tables. Journal of the American Society for Information Science and Technology: JA-SIST, 63(1):108–113, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kozak:2013:BCP

[KH13] Marcin Kozak and James Hartley. Brief communication: Publication fees for open access journals: Different disciplines — different methods. Journal of the American Society for Information Science and Technology: JASIST, 64 (12):2591–2594, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kayed:2008:BOR

[KHAN08] Ahmad Kayed, Nael Hirzallah, Luai A. Al Shalabi, and Mohannad Najjar. Building ontological relationships: a new approach. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1801–1809, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ku:2009:OMR

[KHC09] Lun-Wei Ku, Hsiu-Wei Ho, and Hsin-Hsi Chen. Opinion mining and relationship discovery using CopeOpi opinion analysis system. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1486–1503, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kuan:2012:TDA

[KHC12] Chung-Huei Kuan, Mu-Hsuan Huang, and Dar-Zen Chen. A two-dimensional approach to performance evaluation for a large number of research institutions. Journal of the American Society for Information Science and Technology: JA-SIST, 63(4):817–828, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kilgour:2001:KIO

[Kil01] Frederick G. Kilgour. Known-item online searches employed by scholars using surname plus first, or last, or first and last title words. Journal of the American Society for Information Science and Technology: JASIST, 52(14):1203–1209, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kilgour:2004:EUC

[Kil04] Frederick G. Kilgour. An experiment using coordinate title word searches. Journal of the American Society for Information Science and Technology: JASIST, 55(1):74–80, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2005:OIO

[Kim05] Hyun Hee Kim. ONTOWEB: Implementing an ontologybased Web retrieval system. Journal of the American Society for Information Science and Technology: JASIST, 56(11): 1167–1176, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2006:TUW

[Kim06] Jong-Ae Kim. Toward an understanding of Web-based subscription database acceptance. Journal of the American Society for Information Science and Technology: JASIST, 57 (13):1715–1728, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2009:DPI

[Kim09a] Jeonghyun Kim. Describing and predicting informationseeking behavior on the Web. Journal of the American Society for Information Science and Technology: JASIST, 60 (4):679–693, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2009:VPR

[Kim09b] Yong-Mi Kim. Validation of psychometric research instruments: the case of information science. Journal of the American Society for Information Science and Technology: JA-SIST, 60(6):1178–1191, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2010:FSA

[Kim10a] Jihyun Kim. Faculty self-archiving: Motivations and barriers. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1909–1922, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2010:AUL

[Kim10b] Yong-Mi Kim. The adoption of university library Web site resources: a multigroup analysis. Journal of the American Society for Information Science and Technology: JASIST, 61(5):978–993, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2011:TVS

[Kim11a] Hyun Hee Kim. Toward video semantic search based on a structured folksonomy. Journal of the American Society for Information Science and Technology: JASIST, 62(3):478– 492, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kimport:2011:BRB

[Kim11b] Katrina Kimport. Book review: Alternative and Activist New Media. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1649–1651, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kishida:2010:HSR

[Kis10] Kazuaki Kishida. High-speed rough clustering for very large document collections. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1092– 1104, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kekalainen:2002:UGR

[KJ02] Jaana Kekäläinen and Kalervo Järvelin. Using graded relevance assessments in IR evaluation. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1120–1129, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kaye:2003:HOM

[KJ03] Barbara K. Kaye and Thomas J. Johnson. From here to obscurity?: Media substitution theory and traditional media in an on-line world. Journal of the American Society for Information Science and Technology: JASIST, 54(3):260– 273, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kumpulainen:2012:BTB

[KJ12] Sanna Kumpulainen and Kalervo Järvelin. Barriers to taskbased information access in molecular medicine. Journal of the American Society for Information Science and Technology: JASIST, 63(1):86–97, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Khan:2004:EWD

[KK04] M. Shamim Khan and Sebastian Khor. Enhanced Web document retrieval using automatic query expansion. Journal of the American Society for Information Science and Technology: JASIST, 55(1):29–40, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2008:DVE

[KK08] Yong Ho Kim and Hyun Hee Kim. Development and validation of evaluation indicators for a consortium of institutional repositories: a case study of dCollection. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1282–1294, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kang:2010:MVI

[KK10a] Minhyung Kang and Young-Gul Kim. A multilevel view on interpersonal knowledge transfer. Journal of the American Society for Information Science and Technology: JASIST, 61(3):483–494, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2010:TCF

[KK10b] Hyun Hee Kim and Yong Ho Kim. Toward a conceptual framework of key-frame extraction and storyboard display for video summarization. Journal of the American Society for Information Science and Technology: JASIST, 61 (5):927–939, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kaptein:2011:EET

[KK11] Rianne Kaptein and Jaap Kamps. Explicit extraction of topical context. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1548–1563, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kas:2012:ASN

[KKF^{+}12] Miray Kas, Alla G. Khadka, William Frankenstein, Ahmed Y. Abdulla, Frank Kunkel, L. Richard Carley, and Kathleen M. Carley. Analyzing scientific networks for nuclear capabilities assessment. Journal of the American Society for Information Science and Technology: JASIST, 63(7): 1294–1312, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kang:2011:UES

[KKGB11] Minhyung Kang, Byoungsoo Kim, Peter Gloor, and Gee-Woo Bock. Understanding the effect of social networks on user behaviors in community-driven knowledge services. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1066–1074, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ku:2011:CRA

[KL11] Chih-Hao Ku and Gondy Leroy. A crime reports analysis system to identify related crimes. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1533–1547, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kruschwitz:2013:DQS

[KLAS13] Udo Kruschwitz, Deirdre Lungley, M-Dyaa Albakour, and Dawei Song. Deriving query suggestions for site search. Journal of the American Society for Information Science and Technology: JASIST, 64(10):1975–1994, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Klein:2008:PQM

[Kle08] Shmuel T. Klein. Processing queries with metrical constraints in XML-based IR systems. Journal of the American Society for Information Science and Technology: JASIST, 59(1):86–97, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kazai:2004:ERL

[KLFG04] Gabriella Kazai, Mounia Lalmas, Norbert Fuhr, and Norbert Gövert. European research letter: A report on the first year of the INitiative for the Evaluation of XML retrieval (INEX’02). Journal of the American Society for Information Science and Technology: JASIST, 55(6):551–556, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2009:DNC

[KLP09] Pan Jun Kim, Jae Yun Lee, and Ji-Hong Park. Developing a new collection-evaluation method: Mapping and the user-side h-index. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2366–2377, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kuo:2008:ALC

[KLY08] Jin-Shea Kuo, Haizhou Li, and Ying-Kuei Yang. Active learning for constructing transliteration lexicons from the Web. Journal of the American Society for Information Science and Technology: JASIST, 59(1):126–135, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kifle:2006:TSS

[KMD06] Mengistu Kifle, Victor W. A. Mbarika, and Pratim Datta. Telemedicine in sub-Saharan Africa: the case of teleophthalmology and eye care in Ethiopia. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1383–1393, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kling:2003:BMI

[KMK03] Rob Kling, Geoffrey McKim, and Adam King. A bit more to it: Scholarly Communication Forums as Socio-Technical Interaction Networks. Journal of the American Society for Information Science and Technology: JASIST, 54(1):47–67, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Knothe:2006:CCA

[Kno06] Gerhard Knothe. Comparative citation analysis of duplicate or highly related publications. Journal of the American Society for Information Science and Technology: JASIST, 57 (13):1830–1839, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kwon:2005:MFA

[KO05] Nahyun Kwon and Anthony J. Onwuegbuzie. Modeling the factors affecting individuals’ use of community networks: a theoretical explanation of community-based information and communication technology use. Journal of the American Society for Information Science and Technology: JASIST, 56(14):1525–1543, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2009:URC

[KO09] Soojung Kim and Sanghee Oh. Users’ relevance criteria for evaluating answers in a social Q&A site. Journal of the American Society for Information Science and Technology: JASIST, 60(4):716–727, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koshik:2012:CAS

[KO12] Irene Koshik and Hiromi Okazawa. A conversation analytic study of actual and potential problems in communication in library chat reference interactions. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2006–2019, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koc:2006:BRB

[Koc06] Mustafa Koc. Book review: Web site design with the patron in mind: a step-by-step guide for libraries. Journal of the American Society for Information Science and Technology: JASIST, 57(3):444–445, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koehler:2002:WPC

[Koe02] Wallace Koehler. Web page change and persistence — a fouryear longitudinal study. Journal of the American Society for Information Science and Technology: JASIST, 53(2):162– 171, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koehler:2003:BR

[Koe03] Wallace Koehler. Book review. Journal of the American Society for Information Science and Technology: JASIST, 54(13):1264, November 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koehler:2004:BRBa

[Koe04a] Wallace Koehler. Book review: Virtual inequality: Beyond the digital divide. Journal of the American Society for Information Science and Technology: JASIST, 55(5):467–468, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koehler:2004:BRBb

[Koe04b] Wallace Koehler. Book review: Virtual inequality: Beyond the digital divide. Journal of the American Society for Information Science and Technology: JASIST, 55(11):1024–1025, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koenig:2011:BRB

[Koe11] Michael E. D. Koenig. Book review: Knowledge Management in Theory and Practice (2nd ed.). Journal of the American Society for Information Science and Technology: JA-SIST, 62(10):2083, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koh:2013:AIC

[Koh13] Kyungwon Koh. Adolescents’ information-creating behavior embedded in digital Media practice using scratch. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1826–1841, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Komito:2011:SMM

[Kom11] Lee Komito. Social media and migration: Virtual community 2.0. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1075–1086, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kostoff:2003:LEP

[Kos03] Ronald N. Kostoff. Letter to the Editor: The practice and malpractice of stemming. Journal of the American Society for Information Science and Technology: JASIST, 54(10): 984–985, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kostek:2004:ASC

[Kos04] Bozena Kostek. Application of soft computing to automatic music information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 55(12): 1108–1116, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koshman:2005:TUI

[Kos05] Sherry Koshman. Testing user interaction with a prototype visualization-based information retrieval system. Journal of the American Society for Information Science and Technology: JASIST, 56(8):824–833, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kawamura:2009:SHN

[KOS09a] Shuntaro Kawamura, Yo-Hey Otake, and Takafumi Suzuki. The structure of the hyperlink network formed by the Web pages of Japanese public libraries. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1159–1167, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koshman:2009:BRB

[Kos09b] Sherry Koshman. Book review: Information Visualization: Human-Centered Issues and Perspectives. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2588–2590, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kostoff:2012:LEW

[Kos12] Ronald N. Kostoff. Letter to the Editor: Where is the research in the research literature? Journal of the American Society for Information Science and Technology: JASIST, 63(8):1675–1676, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kracker:2003:ELA

[KP03] Jacqueline Kracker and Howard R. Pollio. The experience of libraries across time: Thematic analysis of undergraduate recollections of library experiences. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1104–1116, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Khan:2011:MTH

[KP11] Gohar Feroz Khan and Han Woo Park. Measuring the triple helix on the web: Longitudinal trends in the universityindustry-government relationship in Korea. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2443–2455, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kang:2010:MSA

[KPEO10] Hyunmo Kang, Catherine Plaisant, Tamer Elsayed, and Douglas W. Oard. Making sense of archived e-mail: Exploring the Enron collection with NetLens. Journal of the American Society for Information Science and Technology: JASIST, 61(4):723–744, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kretschmer:2001:RAI

[KR01] Hildrun Kretschmer and Ronald Rousseau. Research: Author inflation leads to a breakdown of Lotka’s Law. Journal of the American Society for Information Science and Technology: JASIST, 52(8):610–614, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2008:CTC

[KR08] Sujin Kim and Edie Rasmussen. Characteristics of tissuecentric biomedical researchers using a survey and cluster analysis. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1210–1223, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kracker:2002:RASa

[Kra02a] Jacqueline Kracker. Research anxiety and students’ perceptions of research: an experiment. Part I. Effect of teaching Kuhlthau’s ISP model. Journal of the American Society for Information Science and Technology: JASIST, 53(4):282– 294, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kraft:2002:EWN

[Kra02b] Donald H. Kraft. Editorials: What’s new in JASIST. Journal of the American Society for Information Science and Technology: JASIST, 53(7):614, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kraft:2003:EHB

[Kra03a] Donald H. Kraft. Editorials: Honoring the JASIST referees. Journal of the American Society for Information Science and Technology: JASIST, 54(8):701–703, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kraft:2003:ERB

[Kra03b] Donald H. Kraft. Editorials: Refereeing for JASIST. Journal of the American Society for Information Science and Technology: JASIST, 54(9):807, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kraft:2004:CPC

[Kra04] Donald H. Kraft. Call for papers: A call for JASIST referees. Journal of the American Society for Information Science and Technology: JASIST, 55(2):185, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kraft:2008:EBO

[Kra08] Donald H. Kraft. Editorial: JASIST open access. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1539, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Krauskopf:2013:LED

[Kra13] Erwin Krauskopf. Letter to the Editor: Deceiving the research community through manipulation of the impact factor. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2403, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Keselman:2010:ASN

[KRK^{+}10] Alla Keselman, Graciela Rosemblat, Halil Kilicoglu, Marcelo Fiszman, Honglan Jin, Dongwook Shin, and Thomas C. Rindflesch. Adapting semantic natural language processing technology to address information overload in influenza epidemic management. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2531– 2543, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Karunakaran:2013:TMC

[KRS13] Arvind Karunakaran, Madhu C. Reddy, and Patricia Ruma Spence. Toward a model of collaborative information behavior in organizations. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2437– 2451, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kulyukin:2001:RRS

[KS01] Vladimir A. Kulyukin and Amber Settle. Ranked retrieval with semantic networks and vector spaces. Journal of the American Society for Information Science and Technology: JASIST, 52(14):1224–1233, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Katerattanakul:2008:FAI

[KS08] Pairin Katerattanakul and Keng Siau. Factors affecting the information quality of personal Web portfolios. Journal of the American Society for Information Science and Technology: JASIST, 59(1):63–76, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Khalifa:2010:ASN

[KS10] Mohamed Khalifa and Kathy Ning Shen. Applying semantic networks to hypertext design: Effects on knowledge structure acquisition and problem solving. Journal of the American Society for Information Science and Technology: JA-SIST, 61(8):1673–1685, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kelly:2013:SRI

[KS13] Diane Kelly and Cassidy R. Sugimoto. A systematic review of interactive information retrieval evaluation studies, 1967–2006. Journal of the American Society for Information Science and Technology: JASIST, 64(4):745–770, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koppel:2009:CMA

[KSA09] Moshe Koppel, Jonathan Schler, and Shlomo Argamon. Computational methods in authorship attribution. Journal of the American Society for Information Science and Technology: JASIST, 60(1):9–26, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kling:2004:RSV

[KSF04] Rob Kling, Lisa B. Spector, and Joanna Fortuna. The real stakes of virtual publishing: the transformation of E-Biomed into PubMed central. Journal of the American Society for Information Science and Technology: JASIST, 55(2):127– 148, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koshman:2006:WSV

[KSJ06] Sherry Koshman, Amanda Spink, and Bernard J. Jansen. Web searching on the Vivisimo search engine. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1875–1887, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Keenan:2001:REP

[KSK01] Sabrina Keenan, Alan F. Smeaton, and Gary Keogh. Research: The effect of pool depth on system evaluation in TREC. Journal of the American Society for Information Science and Technology: JASIST, 52(7):570–574, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Komlodi:2006:PSU

[KSM06] Anita Komlodi, Dagobert Soergel, and Gary Marchionini. Perspectives on search user interfaces: Best practices and future visions: Search histories for user support in user interfaces. Journal of the American Society for Information Science and Technology: JASIST, 57(6):803–807, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kuflik:2003:SBV

[KSS03] Tsvi Kuflik, Bracha Shapira, and Peretz Shoval. Stereotypebased versus personal-based filtering rules in information filtering systems. Journal of the American Society for Information Science and Technology: JASIST, 54(3):243–250, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Koike:2007:KDB

[KT07a] Asako Koike and Toshihisa Takagi. Knowledge discovery based on an implicit and explicit conceptual network. Journal of the American Society for Information Science and Technology: JASIST, 58(1):51–65, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kousha:2007:GSC

[KT07b] Kayvan Kousha and Mike Thelwall. Google Scholar citations and Google Web/URL citations: a multi-discipline exploratory analysis. Journal of the American Society for Information Science and Technology: JASIST, 58(7):1055– 1065, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kousha:2007:HSC

[KT07c] Kayvan Kousha and Mike Thelwall. How is science cited on the Web? A classification of Google unique Web citations. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1631–1644, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kousha:2008:AID

[KT08] Kayvan Kousha and Mike Thelwall. Assessing the impact of disciplinary research on teaching: an automatic analysis of online syllabuses. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2060– 2069, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kousha:2009:GBS

[KT09] Kayvan Kousha and Mike Thelwall. Google book search: Citation analysis for social science and the humanities. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1537–1549, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Krishnan:2012:MEG

[KT12] Satish Krishnan and Thompson S. H. Teo. Moderating effects of governance on information infrastructure and egovernment development. Journal of the American Society for Information Science and Technology: JASIST, 63(10): 1929–1946, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kousha:2012:ROV

[KTA12] Kayvan Kousha, Mike Thelwall, and Mahshid Abdoli. The role of online videos in research communication: a content analysis of YouTube videos cited in academic publications. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1710–1727, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Karanasios:2013:MSD

[KTL^{+}13] Stan Karanasios, Dhavalkumar Thakker, Lydia Lau, David Allen, Vania Dimitrova, and Alistair Norman. Making sense of digital traces: an activity theory driven ontological approach. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2452–2467, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kousha:2010:CIS

[KTR10] Kayvan Kousha, Mike Thelwall, and Somayeh Rezaie. Can the impact of scholarly images be assessed online? An exploratory study using image identification technology. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1734–1744, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kousha:2011:ACI

[KTR11] Kayvan Kousha, Mike Thelwall, and Somayeh Rezaie. Assessing the citation impact of books: the role of Google Books, Google Scholar, and Scopus. Journal of the American Society for Information Science and Technology: JA-SIST, 62(11):2147–2164, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kankanhalli:2005:USE

[KTW05] Atreyi Kankanhalli, Bernard C. Y. Tan, and Kwok-Kee Wei. Understanding seeking from electronic knowledge repositories: an empirical study. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1156– 1166, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kuperman:2006:PIM

[Kup06] Victor Kuperman. Productivity in the Internet mailing lists: a bibliometric analysis. Journal of the American Society for Information Science and Technology: JASIST, 57(1):51–59, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kvasny:2010:BRB

[Kva10] Lynette Kvasny. Book review: Google and the Digital Divide: The Bias of Online Knowledge. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2586–2587, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kim:2001:RCB

[KW01] Won Kim and W. John Wilbur. Research: Corpus-based statistical screening for content-bearing terms. Journal of the American Society for Information Science and Technology: JASIST, 52(3):247–259, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kracker:2002:RASb

[KW02] Jacqueline Kracker and Peiling Wang. Research anxiety and students’ perceptions of research: an experiment. Part II. Content analysis of their writings on two experiences. Journal of the American Society for Information Science and Technology: JASIST, 53(4):295–307, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kwon:2005:CNC

[Kwo05] Nahyun Kwon. Community networks: Community capital or merely an affordable Internet access tool? Journal of the American Society for Information Science and Technology: JASIST, 56(8):812–823, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kelly:2007:UID

[KWR^{+}07] Diane Kelly, Nina Wacholder, Robert Rittman, Ying Sun, Paul Kantor, Sharon Small, and Tomek Strzalkowski. Using interview data to identify evaluation criteria for interactive, analytical question-answering systems. Journal of the American Society for Information Science and Technology: JA-SIST, 58(7):1032–1043, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kellar:2007:FSC

[KWS07] Melanie Kellar, Carolyn Watters, and Michael Shepherd. A field study characterizing Web-based information-seeking tasks. Journal of the American Society for Information Science and Technology: JASIST, 58(7):999–1018, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kwok:2004:SPP

[KY04] Sai Ho Kwok and Christopher C. Yang. Searching the peerto-peer networks: the community and their queries. Journal of the American Society for Information Science and Technology: JASIST, 55(9):783–793, July 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kuo:2008:SIA

[KY08] Feng-Yang Kuo and Mei-Lien Young. A study of the intention–action gap in knowledge sharing practices. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1224–1237, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Kaszkiel:2001:RER

[KZ01] Marcin Kaszkiel and Justin Zobel. Research: Effective ranking with arbitrary passages. Journal of the American Society for Information Science and Technology: JASIST, 52 (4):344–364, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lariviere:2008:LTV

[LAG08] Vincent Larivière, Éric Archambault, and Yves Gingras. Long-term variations in the aging of scientific literature: From exponential growth to steady-state science (1900– 2004). Journal of the American Society for Information Science and Technology: JASIST, 59(2):288–296, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lariviere:2006:PSR

[LAGVG06] Vincent Larivière, Éric Archambault, Yves Gingras, and Étienne Vignola-Gagné. The place of serials in referencing practices: Comparing natural sciences and engineering with social sciences and humanities. Journal of the American Society for Information Science and Technology: JASIST, 57 (8):997–1004, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lai:2009:HRC

[Lai09] Jung-Yu Lai. How reward, computer self-efficacy, and perceived power security affect knowledge management systems success: an empirical investigation in high-tech companies. Journal of the American Society for Information Science and Technology: JASIST, 60(2):332–347, February 2009. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lakshminarayana:2003:LED

[Lak03] S. Lakshminarayana. Letters to the Editor: Dynamic ranking with n + 1 dimensional vector space models: an alternative search mechanism for World Wide Web. Journal of the American Society for Information Science and Technology: JASIST, 54(3):274–275, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lakshminarayana:2007:LEQ

[Lak07] S. Lakshminarayana. Letter to the Editor: Quality search content: a reality with next generation browsers. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1223, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lakshminarayana:2008:LEA

[Lak08] Sadasivuni Lakshminarayana. Letter to the Editor: Antiplagiarism certification be an academic mandate. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2194, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lucio-Arias:2008:MPA

[LAL08] Diana Lucio-Arias and Loet Leydesdorff. Main-path analysis and path-dependent transitions in HistCite^{T M}-based historiograms. Journal of the American Society for Information Science and Technology: JASIST, 59(12):1948–1962, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lucio-Arias:2009:IRF

[LAL09] Diana Lucio-Arias and Loet Leydesdorff. An indicator of research front activity: Measuring intellectual organization as uncertainty reduction in document sets. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2488–2498, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lankes:2004:DRR

[Lan04] R. David Lankes. The digital reference research agenda. Journal of the American Society for Information Science and Technology: JASIST, 55(4):301–311, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Landry:2006:WRT

[Lan06] Carol Fay Landry. Work roles, tasks, and the information behavior of dentists. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1896– 1908, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lapinski:2007:BRB

[Lap07] P. Scott Lapinski. Book review: Managing information technology: a handbook for systems librarians. Journal of the American Society for Information Science and Technology: JASIST, 58(6):908–909, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Larsen:2009:PDI

[Lar09] Jack Lenor Larsen. Perspectives on design: Information technologies and creative practices: The LongHouse proposal for objects classified by mediums. Journal of the American Society for Information Science and Technology: JA-SIST, 60(9):1907–1914, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Large:2010:BRB

[Lar10a] Andrew Large. Book review: Mobile Technology for Children: Designing for Interaction and Learning. Journal of the American Society for Information Science and Technology: JASIST, 61(4):855, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Larson:2010:BRBb

[Lar10b] Ray R. Larson. Book review: Information Retrieval: Searching in the 21st Century; Human Information Retrieval. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2370–2372, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Larson:2010:BRBa

[Lar10c] Ray R. Larson. Book review: Introduction to Information Retrieval. Journal of the American Society for Information Science and Technology: JASIST, 61(4):852–853, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Latham:2002:IAN

[Lat02] Don Latham. Information architecture: Notes toward a new curriculum. Journal of the American Society for Information Science and Technology: JASIST, 53(10):824–830, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luyt:2008:IWA

[LATH08] Brendan Luyt, Tay Chee Hsien Aaron, Lim Hai Thian, and Cheng Kian Hong. Improving Wikipedia’s accuracy: Is edit age a solution? Journal of the American Society for Information Science and Technology: JASIST, 59(2):318–330, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Law:2002:BR

[Law02] Derek G. Law. Book review. Journal of the American Society for Information Science and Technology: JASIST, 53(14): 1271–1272, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lazarinis:2007:EUS

[Laz07] Fotis Lazarinis. Engineering and utilizing a stopword list in Greek Web retrieval. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1645– 1652, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Losada:2003:ETS

[LB03] David E. Losada and Alvaro Barreiro. Embedding term similarity and inverse document frequency into a logical model of information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 54(4):285– 301, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2005:MCS

[LB05] Loet Leydesdorff and Jin Bihui. Mapping the Chinese Science Citation Database in terms of aggregated journal– journal citation relations. Journal of the American Society for Information Science and Technology: JASIST, 56(14): 1469–1479, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2006:CPL

[LB06] Loet Leydesdorff and Stephen Bensman. Classification and powerlaws: the logarithmic transformation. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1470–1486, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2008:ICA

[LB08] Jiunde Lee and Elizabeth Boling. Information-conveying approaches and cognitive styles of mental modeling in a hypermedia-based learning environment. Journal of the American Society for Information Science and Technology: JASIST, 59(4):644–661, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2010:ERB

[LB10] Yuelin Li and Nicholas J. Belkin. An exploration of the relationships between work task and interactive information search behavior. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1771–1789, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2011:HFC

[LB11a] Loet Leydesdorff and Lutz Bornmann. How fractional counting of citations affects the impact factor: Normalization in terms of differences in citation potentials among fields of science. Journal of the American Society for Information Science and Technology: JASIST, 62(2):217–229, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2011:III

[LB11b] Loet Leydesdorff and Lutz Bornmann. Integrated impact indicators compared with impact factors: an alternative research design with policy implications. Journal of the American Society for Information Science and Technology: JA-SIST, 62(11):2133–2146, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2012:LEP

[LB12a] Loet Leydesdorff and Lutz Bornmann. Letter to the Editor: Percentile ranks and the integrated impact indicator ( I3 ). Journal of the American Society for Information Science and Technology: JASIST, 63(9):1901–1902, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2012:MUP

[LB12b] Loet Leydesdorff and Lutz Bornmann. Mapping (USPTO) patent data using overlays to Google Maps. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1442–1458, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lor:2012:EPP

[LB12c] Peter Johan Lor and J. J. Britz. An ethical perspective on political-economic issues in the long-term preservation of digital heritage. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2153–2164, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Laakso:2013:DOA

[LB13a] Mikael Laakso and Bo-Christer Björk. Delayed open access: an overlooked high-impact category of openly available scientific literature. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1323–1329, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lingel:2013:KIS

[LB13b] Jessa Lingel and Danah Boyd. “keep it secret, keep it safe”: Information poverty, information norms, and stigma. Journal of the American Society for Information Science and Technology: JASIST, 64(5):981–991, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Large:2002:IAW

[LBC02] Andrew Large, Jamshid Beheshti, and Charles Cole. Information architecture for the Web: The IA matrix approach to designing children’s portals. Journal of the American Society for Information Science and Technology: JASIST, 53 (10):831–838, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2011:TTC

[LBMO11] Loet Leydesdorff, Lutz Bornmann, Rüdiger Mutz, and Tobias Opthof. Turning the tables on citation analysis one more time: Principles for comparing sets of documents. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1370–1381, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Large:2004:DWP

[LBNB04] Andrew Large, Jamshid Beheshti, Valerie Nesset, and Leanne Bowler. Designing Web portals in intergenerational teams: Two prototype portals for elementary school students. Journal of the American Society for Information Science and Technology: JASIST, 55(13):1140–1154, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Large:2002:DCC

[LBR02] Andrew Large, Jamshid Beheshti, and Tarjin Rahman. Design criteria for children’s Web portals: the users speak out. Journal of the American Society for Information Science and Technology: JASIST, 53(2):79–94, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2011:AAD

[LBS11] Hui Li, Sourav S. Bhowmick, and Aixin Sun. AffRank: Affinity-driven ranking of products in online social rating networks. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1345–1359, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Large:2009:DVT

[LBTN09] Andrew Large, Jamshid Beheshti, Nahid Tabatabaei, and Valerie Nesset. Developing a visual taxonomy: Children’s views on aesthetics. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1808– 1822, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2001:RAI

[lC01] Hsin liang Chen. Research: An analysis of image queries in the field of art history. Journal of the American Society for Information Science and Technology: JASIST, 52(3):260– 273, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leroy:2005:GOE

[LC05a] Gondy Leroy and Hsinchun Chen. Genescene: an ontologyenhanced integration of linguistic and co-occurrence based relations in biomedical texts. Journal of the American Society for Information Science and Technology: JASIST, 56 (5):457–468, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2005:TBS

[LC05b] Peng Liu and Amit Chetal. Trust-based secure information sharing between Federal Government agencies. Journal of the American Society for Information Science and Technology: JASIST, 56(3):283–298, February 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Losee:2005:BCT

[LC05c] Robert M. Losee and Lewis Church, Jr. Brief communication: Are two document clusters better than one? The Cluster Performance Question for information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 56(1):106–108, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lau:2006:BMP

[LC06] Annie Y. S. Lau and Enrico W. Coiera. A Bayesian model that predicts the impact of Web searching on decision making. Journal of the American Society for Information Science and Technology: JASIST, 57(7):873–880, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2012:MSI

[LC12] Lung-Hao Lee and Hsin-Hsi Chen. Mining search intents for collaborative cyberporn filtering. Journal of the American Society for Information Science and Technology: JASIST, 63(2):366–376, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2013:DBL

[LC13] Shengbo Liu and Chaomei Chen. The differences between latent topics in abstracts and citation contexts of citing papers. Journal of the American Society for Information Science and Technology: JASIST, 64(3):627–639, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leide:2007:TBI

[LCB^{+}07] John E. Leide, Charles Cole, Jamshid Beheshti, Andrew Large, and Yang Lin. Task-based information retrieval: Structuring undergraduate history essays for better course evaluation using essay-type visualizations. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1227–1241, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2007:UUA

[LCC07] Matthew K. O. Lee, Christy M. K. Cheung, and Zhaohui Chen. Understanding user acceptance of multimedia messaging services: an empirical study. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2066–2077, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lam:2005:CBG

[LCR^{+}05] Wai Lam, Ki Chan, Dragomir Radev, Horacio Saggion, and Simone Teufel. Context-based generic cross-lingual retrieval of documents and automated summaries. Journal of the American Society for Information Science and Technology: JASIST, 56(2):129–139, January 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lin:2006:UCA

[LCW06] Julian Lin, Hock Chuan Chan, and Kwok Kee Wei. Understanding competing application usage with the theory of planned behavior. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1338–1349, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lesaffre:2008:HPU

[LDL^{+}08] Micheline Lesaffre, Liesbeth De Voogdt, Marc Leman, Bernard De Baets, Hans De Meyer, and Jean-Pierre Martens. How potential users of music search and retrieval systems describe the semantic quality of music. Journal of the American Society for Information Science and Technology: JASIST, 59(5):695–707, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lorca:2012:SCD

[LDM12] Pedro Lorca, Javier De Andrées, and Ana B. Martínez. Size and culture as determinants of the web policy of listed firms: the case of web accessibility in Western European countries. Journal of the American Society for Information Science and Technology: JASIST, 63(2):392–405, February 2012. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2010:JMB

[LdMAGB10] Loet Leydesdorff, Félix de Moya-Anegón, and Vicente P. Guerrero-Bote. Journal maps on the basis of Scopus data: a comparison with the Journal Citation Reports of the ISI. Journal of the American Society for Information Science and Technology: JASIST, 61(2):352–369, February 2010. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2011:MTC

[LDS^{+}11] Daifeng Li, Ying Ding, Cassidy Sugimoto, Bing He, Jie Tang, Erjia Yan, Nan Lin, Zheng Qin, and Tianxi Dong. Modeling topic and community structure in social tagging: The TTR–LDA–Community model. Journal of the American Society for Information Science and Technology: JASIST, 62 (9):1849–1866, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lewandowski:2012:DQI

[LDvM12] Dirk Lewandowski, Jessica Drechsler, and Sonja von Mach. Deriving query intents from web search engine queries. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1773–1788, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2010:AUN

[Lee10] Jin Ha Lee. Analysis of user needs and information features in natural language queries seeking music information. Journal of the American Society for Information Science and Technology: JASIST, 61(5):1025–1045, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lercher:2010:ESC

[Ler10] Aaron Lercher. Efficiency of scientific communication: a survey of world science. Journal of the American Society for Information Science and Technology: JASIST, 61(10): 2049–2060, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lercher:2013:CTC

[Ler13] Aaron Lercher. Correlation over time for citations to mathematics articles. Journal of the American Society for Information Science and Technology: JASIST, 64(3):455–463, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leslie:2011:BRB

[Les11] Christopher Leslie. Book reviews: The Master Switch: The Rise and Fall of Information Empires; In the Plex: How Google Thinks, Works, and Shapes Our Lives; The Net Delusion: The Dark Side of Internet Freedom. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2540–2543, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2002:DEU

[Ley02] Loet Leydesdorff. Dynamic and evolutionary updates of classificatory schemes in scientific journal structures. Journal of the American Society for Information Science and Technology: JASIST, 53(12):987–994, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2003:LEE

[Ley03a] Loet Leydesdorff. Letters to the Editor: Empirical evidence of self-organization? A rejoinder. Journal of the American Society for Information Science and Technology: JASIST, 54(8):804, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [vdB03a].

Leydesdorff:2003:LER

[Ley03b] Loet Leydesdorff. Letters to the Editor: Rejoinder to Van den Besselaar’s letter entitled “Descriptive statistics, inferential statistics, rhetorical statistics”. Journal of the American Society for Information Science and Technology: JASIST, 54(11):1077–1078, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [vdB03b].

Leydesdorff:2004:UIK

[Ley04] Loet Leydesdorff. The university–industry knowledge relationship: Analyzing patents and the science base of technologies. Journal of the American Society for Information Science and Technology: JASIST, 55(11):991–1001, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2005:BCS

[Ley05] Loet Leydesdorff. Brief communication: Similarity measures, author cocitation analysis, and information theory. Journal of the American Society for Information Science and Technology: JASIST, 56(7):769–772, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2006:CSJ

[Ley06] Loet Leydesdorff. Can scientific journals be classified in terms of aggregated journal–journal citation relations using the Journal Citation Reports? Journal of the American Society for Information Science and Technology: JASIST, 57 (5):601–613, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2007:BCI

[Ley07a] Loet Leydesdorff. Betweenness centrality as an indicator of the interdisciplinarity of scientific journals. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1303–1319, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2007:LES

[Ley07b] Loet Leydesdorff. Letter to Editor: Should co-occurrence data be normalized? A rejoinder. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2411–2413, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2007:VCI

[Ley07c] Loet Leydesdorff. Visualization of the citation impact environments of scientific journals: an online mapping exercise. Journal of the American Society for Information Science and Technology: JASIST, 58(1):25–38, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2008:CUC

[Ley08a] Loet Leydesdorff. Caveats for the use of citation indicators in research and journal evaluations. Journal of the American Society for Information Science and Technology: JASIST, 59(2):278–287, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2008:NVA

[Ley08b] Loet Leydesdorff. On the normalization and visualization of author co-citation data: Salton’s Cosine versus the Jaccard index. Journal of the American Society for Information Science and Technology: JASIST, 59(1):77–85, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2008:PCI

[Ley08c] Loet Leydesdorff. Patent classifications as indicators of intellectual organization. Journal of the American Society for Information Science and Technology: JASIST, 59(10): 1582–1597, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2009:HNC

[Ley09] Loet Leydesdorff. How are new citation-based journal indicators adding to the bibliometric toolbox? Journal of the American Society for Information Science and Technology: JASIST, 60(7):1327–1336, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2010:CMS

[Ley10] Loet Leydesdorff. The communication of meaning and the structuration of expectations: Giddens’ “structuration theory” and Luhmann’s “self-organization”. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2138–2150, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2012:LEA

[Ley12] Loet Leydesdorff. Letter to the Editor: Accounting for the uncertainty in the evaluation of percentile ranks. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2349–2350, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2003:PKB

[LF03] M. C. Lee and Chun-Kan Fung. A public-key based authentication and key establishment protocol coupled with a client puzzle. Journal of the American Society for Information Science and Technology: JASIST, 54(9):810–823, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2008:PSS

[LfBW08] Quanzhi Li and Yi fang Brook Wu. People search: Searching people sharing similar interests from the Web. Journal of the American Society for Information Science and Technology: JASIST, 59(1):111–125, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liew:2001:PIE

[LFC01] Chern Li Liew, Schubert Foo, and K. R. Chennupati. A proposed integrated environment for enhanced user interaction and value-adding of electronic documents: an empirical evaluation. Journal of the American Society for Information Science and Technology: JASIST, 52(1):22–35, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lou:2013:CHQ

[LFLP13] Jie Lou, Yulin Fang, Kai H. Lim, and Jerry Zeyu Peng. Contributing high quantity and quality knowledge to online Q&A communities. Journal of the American Society for Information Science and Technology: JASIST, 64(2):356–371, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lariviere:2010:BCI

[LG10a] Vincent Larivière and Yves Gingras. Brief communications: The impact factor’s Matthew Effect: a natural experiment in bibliometrics. Journal of the American Society for Information Science and Technology: JASIST, 61(2):424–427, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lariviere:2010:RBI

[LG10b] Vincent Larivière and Yves Gingras. On the relationship between interdisciplinarity and scientific impact. Journal of the American Society for Information Science and Technology: JASIST, 61(1):126–131, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lariviere:2009:BCD

[LGA09] Vincent Larivière, Yves Gingras, and Éric Archambault. Brief communication: The decline in the concentration of citations, 1900–2007. Journal of the American Society for Information Science and Technology: JASIST, 60(4):858– 862, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2010:IIP

[LGCA10] Chei Sian Lee, Dion Hoe-Lian Goh, Alton Y. K. Chua, and Rebecca P. Ang. Indagator: Investigating perceived gratifications of an application that blends mobile content sharing with gameplay. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1244–1257, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lykourentzou:2009:EDS

[LGM^{+}09] Ioanna Lykourentzou, Ioannis Giannoukos, George Mpardis, Vassilis Nikolopoulos, and Vassili Loumos. Early and dynamic student achievement prediction in e-learning courses using neural networks. Journal of the American Society for Information Science and Technology: JASIST, 60(2):372– 380, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2001:SOE

[LH01] Loet Leydesdorff and Gaston Heimeriks. The selforganization of the European Information Society: the case of “biotechnology”. Journal of the American Society for Information Science and Technology: JASIST, 52(14):1262– 1274, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2011:REP

[LH11] Rey-Long Liu and Yi-Chih Huang. Ranker enhancement for proximity-based ranking of biomedical texts. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2479–2495, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2007:MWRa

[LHA07] Yang Liu, Xiangji Huang, and Aijun An. Mining Web resources for enhancing information retrieval: Personalized recommendation with adaptive mixture of Markov models. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1851–1870, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lorigo:2008:ETO

[LHB^{+}08] Lori Lorigo, Maya Haridasan, Hrönn Brynjarsdóttir, Ling Xia, Thorsten Joachims, Geri Gay, Laura Granka, Fabio Pellacini, and Bing Pan. Eye tracking and online search: Lessons learned and challenges ahead. Journal of the American Society for Information Science and Technology: JA-SIST, 59(7):1041–1052, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lopez-Herrera:2009:SUM

[LHHVH09] A. G. López-Herrera, E. Herrera-Viedma, and F. Herrera. A study of the use of multi-objective evolutionary algorithms to learn Boolean queries: a comparative study. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1192–1207, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lan:2008:DAR

[LHLL08] Kwok Cheung Lan, Kei Shiu Ho, Robert Wing Pong Luk, and Hong Va Leong. Dialogue act recognition using maximum entropy. Journal of the American Society for Information Science and Technology: JASIST, 59(6):859–874, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lokker:2011:DIF

[LHMW11] Cynthia Lokker, R. Brian Haynes, K. Ann McKibbon, and Nancy L. Wilczynski. Determining the impact factors of secondary journals: a retrospective cohort study. Journal of the American Society for Information Science and Technology: JASIST, 62(4):637–642, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2011:SAH

[LHS11] Loet Leydesdorff, Björn Hammarfelt, and Almila Salah. The structure of the Arts & Humanities Citation Index: a mapping on the basis of aggregated citations among 1,157 journals. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2414–2426, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2009:RNP

[Li09a] Rowena Li. The representation of national political freedom on Web interface design: the indicators. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1222–1248, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2009:ERB

[Li09b] Yuelin Li. Exploring the relationships between work task and search task in information search. Journal of the American Society for Information Science and Technology: JASIST, 60(2):275–291, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liang:2005:RSR

[Lia05] Liming Liang. R-sequences: Relative indicators for the rhythm of science. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1045–1049, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Larsen:2009:DFA

[LIL09] Birger Larsen, Peter Ingwersen, and Berit Lund. Data fusion according to the principle of polyrepresentation. Journal of the American Society for Information Science and Technology: JASIST, 60(4):646–654, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lim:2007:DIT

[Lim07] Sook Lim. Do information technology units have more power than other units in academic libraries? Journal of the American Society for Information Science and Technology: JA-SIST, 58(9):1242–1253, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lim:2009:HWD

[Lim09] Sook Lim. How and why do college students use Wikipedia? Journal of the American Society for Information Science and Technology: JASIST, 60(11):2189–2202, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lincicum:2002:BRB

[Lin02] Shirley J. Lincicum. Book review: Automatic summarization. Journal of the American Society for Information Science and Technology: JASIST, 53(5):410–411, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lincicum:2003:BR

[Lin03] Shirley J. Lincicum. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 54(4):356–357, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lipinski:2003:MTN

[Lip03] Tomas A. Lipinski. The myth of technological neutrality in copyright and the rights of institutional users: Recent legal challenges to the information organization as mediator and the impact of the DMCA, WIPO, and TEACH. Journal of the American Society for Information Science and Technology: JASIST, 54(9):824–835, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lipinski:2013:BRB

[Lip13] Tomas A. Lipinski. Book review: Digital rights movement: the role of technology in subverting digital copyright — by Hector Postigo. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1303–1305, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2009:CRH

[Liu09] Rey-Long Liu. Context recognition for hierarchical text classification. Journal of the American Society for Information Science and Technology: JASIST, 60(4):803–813, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2010:CBT

[Liu10] Rey-Long Liu. Context-based term frequency assessment for text classification. Journal of the American Society for Information Science and Technology: JASIST, 61(2):300–309, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2013:PEC

[Liu13a] Rey-Long Liu. A passage extractor for classification of disease aspect information. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2265– 2277, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2013:GMC

[Liu13b] Xiaozhong Liu. Generating metadata for cyberlearning resources through information retrieval and meta-search. Journal of the American Society for Information Science and Technology: JASIST, 64(4):771–786, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2013:AAQ

[LJ13] Xiaozhong Liu and Han Jia. Answering academic questions for education by recommending cyberlearning resources. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1707–1722, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luor:2008:TLF

[LJLlW08] Ted (Tainyi) Luor, Robert E. Johanson, Hsi-Peng Lu, and Ling ling Wu. Trends and lacunae for future computer assisted learning (CAL) research: an assessment of the literature in SSCI journals from 1998–2006. Journal of the American Society for Information Science and Technology: JA-SIST, 59(8):1313–1320, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lin:2006:BRT

[LK06] Jimmy Lin and Boris Katz. Building a reusable test collection for question answering. Journal of the American Society for Information Science and Technology: JASIST, 57 (7):851–861, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2010:ULSb

[LK10] Dandan Li and Chung-Ping Kwong. Understanding latent semantic indexing: a topological structure analysis using Qanalysis. Journal of the American Society for Information Science and Technology: JASIST, 61(3):592–608, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Levin:2012:CBB

[LKBJ12] Michael Levin, Stefan Krawczyk, Steven Bethard, and Dan Jurafsky. Citation-based bootstrapping for large-scale author disambiguation. Journal of the American Society for Information Science and Technology: JASIST, 63(5):1030– 1047, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lustria:2010:PDH

[LKG^{+}10] Mia Liza A. Lustria, Michelle M. Kazmer, Robert L. Glueckauf, Robert P. Hawkins, Ebrahim Randeree, Ivee B. Rosario, Casey McLaughlin, and Sarah Redmond. Participatory design of a health informatics system for rural health practitioners and disadvantaged women. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2243–2255, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lehtokangas:2008:ETD

[LKJ08] Raija Lehtokangas, Heikki Keskustalo, and Kalervo Järvelin. Experiments with transitive dictionary translation and pseudo-relevance feedback using graded relevance assessments. Journal of the American Society for Information Science and Technology: JASIST, 59(3):476–488, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lamb:2003:IEO

[LKK03] Roberta Lamb, John Leslie King, and Rob Kling. Informational environments: Organizational contexts of online information use. Journal of the American Society for Information Science and Technology: JASIST, 54(2):97–114, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lim:2007:PMM

[LKK07] Jeongyeon Lim, Sanggil Kang, and Munchurl Kim. Perspectives on MPEG-7: Metadata for multimedia: Automatic user preference learning for personalized electronic program guide applications. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1346– 1356, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2010:ULSa

[LKL10] Dandan Li, Chung-Ping Kwong, and Dik Lun Lee. Unified linear subspace approach to semantic analysis. Journal of the American Society for Information Science and Technology: JASIST, 61(1):175–189, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2011:DIS

[LKZC11] Xuan Liu, Siddharth Kaza, Pengzhu Zhang, and Hsinchun Chen. Determining inventor status and its effect on knowledge diffusion: a study on nanotechnology literature from China, Russia, and India. Journal of the American Society for Information Science and Technology: JASIST, 62(6): 1166–1176, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2012:IAM

[LL12] John S. Liu and Louis Y. Y. Lu. An integrated approach for main path analysis: Development of the Hirsch index as an example. Journal of the American Society for Information Science and Technology: JASIST, 63(3):528–542, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2005:EIF

[LLA05] Shan Li, Moon-Chuen Lee, and Donald Adjeroh. Effective invariant features for shape-based image retrieval. Journal of the American Society for Information Science and Technology: JASIST, 56(7):729–740, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lisee:2008:CPS

[LLA08] Cynthia Lisée, Vincent Larivière, and Éric Archambault. Conference proceedings as a source of scientific information: a bibliometric analysis. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1776– 1784, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2012:DRB

[LLC12] Duen-Ren Liu, Chin-Hui Lai, and Ya-Ting Chen. Document recommendations based on knowledge flows: a hybrid of personalized and group-based approaches. Journal of the American Society for Information Science and Technology: JA-SIST, 63(10):2100–2117, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luk:2002:SIS

[LLD^{+}02] Robert W. P. Luk, H. V. Leong, Tharam S. Dillon, Alvin T. S. Chan, W. Bruce Croft, and James Allan. A survey in indexing and searching XML documents. Journal of the American Society for Information Science and Technology: JASIST, 53(6):415–437, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lin:2012:DFD

[LLD^{+}12] Nan Lin, Daifeng Li, Ying Ding, Bing He, Zheng Qin, Jie Tang, Juanzi Li, and Tianxi Dong. The dynamic features of Delicious, Flickr, and YouTube. Journal of the American Society for Information Science and Technology: JASIST, 63(1):139–162, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lozano:2012:WRB

[LLG12] George A. Lozano, Vincent Larivière, and Yves Gingras. The weakening relationship between the impact factor and papers’ citations in the digital age. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2140–2145, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2009:PSO

[LLL09] Kun Chang Lee, Namho Lee, and Honglei Li. A particle swarm optimization-driven cognitive map approach to analyzing information systems project risk. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1208–1221, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lin:2013:LRU

[LLXS13] Yuan Lin, Hongfei Lin, Kan Xu, and Xiaoling Sun. Learning to rank using smoothing methods for language modeling. Journal of the American Society for Information Science and Technology: JASIST, 64(4):818–828, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lam:2001:MUI

[LM01] Wai Lam and Javed Mostafa. Modeling user interest shift using a Bayesian approach. Journal of the American Society for Information Science and Technology: JASIST, 52 (5):416–429, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2009:MWI

[LM09] Hyuk-Jin Lee and Gheorghe Muresan. Mediated Web information retrieval for a complex searching task. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1372–1391, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lariviere:2011:ICS

[LM11] Vincent Larivière and Benoit Macaluso. Improving the coverage of social science and humanities researchers’ output: the case of the Érudit journal platform. Journal of the American Society for Information Science and Technology: JA-SIST, 62(12):2437–2442, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luna-Morales:2009:EPS

[LMCRRPA09] Ma. Elena Luna-Morales, Francisco Collazo-Reyes, Jane M. Russell, and Miguel Ángel Pérez-Angón. Early patterns of scientific production by Mexican researchers in mainstream journals, 1900–1950. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1337– 1348, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lunin:2009:PDIa

[LMH09] Lois F. Lunin, Kathi Martin, and Samantha K. Hastings. Perspectives on design: Information technologies and creative practices: Design: Information technologies and creative practices. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1874–1876, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lucassen:2013:TFI

[LMNS13] Teun Lucassen, Rienco Muilwijk, Matthijs L. Noordzij, and Jan Maarten Schraagen. Topic familiarity and information skills in online credibility evaluation. Journal of the American Society for Information Science and Technology: JASIST, 64(2):254–264, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leroy:2008:BAH

[LMRB08] Gondy Leroy, Trudi Miller, Graciela Rosemblat, and Allen Browne. A balanced approach to health information evaluation: a vocabulary-based naïve Bayes classifier and readability formulas. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1409–1419, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lowry:2012:UEL

[LMV^{+}12] Paul Benjamin Lowry, Greg Moody, Anthony Vance, Matthew Jensen, Jeff Jenkins, and Taylor Wells. Using an elaboration likelihood approach to better understand the persuasiveness of website privacy assurance cues for online consumers. Journal of the American Society for Information Science and Technology: JASIST, 63(4):755–776, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liben-Nowell:2007:LPP

[LNK07] David Liben-Nowell and Jon Kleinberg. The link-prediction problem for social networks. Journal of the American Society for Information Science and Technology: JASIST, 58(7): 1019–1031, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2010:BCS

[LO10] Loet Leydesdorff and Tobias Opthof. Brief communication: Scopus’s source normalized impact per paper (SNIP) versus a journal impact factor based on fractional counting of citations. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2365–2369, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2011:LES

[LO11] Loet Leydesdorff and Tobias Opthof. Letters to the Editor: Scopus’ SNIP indicator: Reply to Moed. Journal of the American Society for Information Science and Technology: JASIST, 62(1):214–215, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2013:BCC

[LO13] Loet Leydesdorff and Tobias Opthof. Brief communication: Citation analysis with medical subject Headings (MeSH) using the Web of Knowledge: a new routine. Journal of the American Society for Information Science and Technology: JASIST, 64(5):1076–1080, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Losee:2001:TDB

[Los01] Robert M. Losee. Term dependence: a basis for Luhn and Zipf models. Journal of the American Society for Information Science and Technology: JASIST, 52(12):1019–1025, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Losee:2013:EAM

[Los13] Robert M. Losee. The effect of assigning a metadata or indexing term on document ordering. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2191–2200, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2003:RSI

[LP03] Moon-Chuen Lee and Chi-Man Pun. Rotation and scale invariant wavelet feature for content-based texture image retrieval. Journal of the American Society for Information Science and Technology: JASIST, 54(1):68–80, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lorigo:2007:FSL

[LP07] Lori Lorigo and Fabio Pellacini. Frequency and structure of long distance scholarly collaborations in a physics community. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1497–1502, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2009:BCD

[LP09] Loet Leydesdorff and Carole Probst. Brief communications: The delineation of an interdisciplinary specialty in terms of a journal set: the case of communication studies. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1709–1718, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2010:MGS

[LP10] Loet Leydesdorff and Olle Persson. Mapping the geography of science: Distribution patterns and networks of relations among cities and institutes. Journal of the American Society for Information Science and Technology: JASIST, 61(8): 1622–1634, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lopez-Pujalte:2003:OBF

[LPGBdMA03] Cristina López-Pujalte, Vicente P. Guerrero-Bote, and Félix de Moya-Anegón. Order-based fitness functions for genetic algorithms applied to relevance feedback. Journal of the American Society for Information Science and Technology: JASIST, 54(2):152–160, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Loia:2006:WNS

[LPSS06] Vincenzo Loia, Witold Pedrycz, Sabrina Senatore, and Maria I. Sessa. Web navigation support by means of proximity-driven assistant agents. Journal of the American Society for Information Science and Technology: JASIST, 57(4):515–527, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lehr:2002:OMF

[LR02] Jennifer K. Lehr and Ronald E. Rice. Organizational measures as a form of knowledge management: a multitheoretic, communication-based exploration. Journal of the American Society for Information Science and Technology: JASIST, 53(12):1060–1073, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Loyland:2008:DBD

[LR08] Knut Løyland and Vidar Ringstad. Determinants of borrowing demand from Norwegian local public libraries. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1295–1303, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2009:GMS

[LR09] Loet Leydesdorff and Ismael Rafols. A global map of science based on the ISI subject categories. Journal of the American Society for Information Science and Technology: JASIST, 60(2):348–362, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2010:KDT

[LR10] Yuxian Liu and Ronald Rousseau. Knowledge diffusion through publications and citations: a case study using ESIfields as unit of diffusion. Journal of the American Society for Information Science and Technology: JASIST, 61(2): 340–351, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2011:LEG

[LR11] Loet Leydesdorff and Ismael Rafols. Local emergence and global diffusion of research technologies: an exploration of patterns of network formation. Journal of the American Society for Information Science and Technology: JASIST, 62 (5):846–860, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lopes:2013:MVH

[LR13] Carla Teixeira Lopes and Cristina Ribeiro. Measuring the value of health query translation: an analysis by user language proficiency. Journal of the American Society for Information Science and Technology: JASIST, 64(5):951–963, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2013:FNI

[LRB^{+}13] Loet Leydesdorff, Filippo Radicchi, Lutz Bornmann, Claudio Castellano, and Wouter de Nooy. Field-normalized impact factors (IFs): a comparison of rescaling and fractionally counted IFs. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2299–2309, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2013:IOJ

[LRC13] Loet Leydesdorff, Ismael Rafols, and Chaomei Chen. Interactive overlays of journals and the measurement of interdisciplinarity on the basis of aggregated journal–journal citations. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2573–2586, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2012:BPM

[LRR12] Loet Leydesdorff, Daniele Rotolo, and Ismael Rafols. Bibliometric perspectives on medical innovation using the medical subject Headings of PubMed. Journal of the American Society for Information Science and Technology: JASIST, 63 (11):2239–2253, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Loureiro:2005:IAC

[LS05] Oscar Loureiro and Hava Siegelmann. Introducing an active cluster-based information retrieval paradigm. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1024–1030, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2008:DAJ

[LS08] Loet Leydesdorff and Thomas Schank. Dynamic animations of journal maps: Indicators of structural changes and interdisciplinary developments. Journal of the American Society for Information Science and Technology: JASIST, 59(11): 1810–1818, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2009:NID

[LS09] Loet Leydesdorff and Yuan Sun. National and international dimensions of the Triple Helix in Japan: University– industry–government versus international coauthorship relations. Journal of the American Society for Information Science and Technology: JASIST, 60(4):778–788, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2010:AUQ

[LS10a] Wei Meng Lee and Mark Sanderson. Analyzing URL queries. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2300–2310, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2010:MBA

[LS10b] Loet Leydesdorff and Alkim Almila Akdag Salah. Maps on the basis of the Arts & Humanities Citation Index: the journals Leonardo and Art Journal versus “digital humanities” as a topic. Journal of the American Society for Information Science and Technology: JASIST, 61(4):787–801, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lewandowski:2011:RWA

[LS11a] Dirk Lewandowski and Ulrike Spree. Ranking of Wikipedia articles in search engines revisited: Fair ranking for reasonable quality? Journal of the American Society for Information Science and Technology: JASIST, 62(1):117–132, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2011:HEU

[LS11b] Loet Leydesdorff and Jung C. Shin. How to evaluate universities in terms of their relative citation impacts: Fractional counting of citations and the normalization of differences among disciplines. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1146– 1155, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2011:HPC

[LS11c] Duen-Ren Liu and Meng-Jung Shih. Hybrid-patent classification based on patent-network analysis. Journal of the American Society for Information Science and Technology: JASIST, 62(2):246–256, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lucassen:2011:FAT

[LS11d] Teun Lucassen and Jan Maarten Schraagen. Factual accuracy and trust in information: the role of expertise. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1232–1242, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2012:STN

[LS12] Danielle H. Lee and Titus Schleyer. Social tagging is no substitute for controlled indexing: a comparison of Medical Subject Headings and CiteULike tags assigned to 231,388 papers. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1747–1757, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2013:SSI

[LS13] Loet Leydesdorff and Øivind Strand. The Swedish system of innovation: Regional synergies in a knowledge-based economy. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1890–1902, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lariviere:2013:TOI

[LSB13] Vincent Larivière, Cassidy R. Sugimoto, and Pierrette Bergeron. In their own image? a comparison of doctoral students’ and faculty members’ referencing behavior. Journal of the American Society for Information Science and Technology: JASIST, 64(5):1045–1054, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lariviere:2012:BCL

[LSC12] Vincent Larivière, Cassidy R. Sugimoto, and Blaise Cronin. A bibliometric chronicling of library and information science’s first hundred years. Journal of the American Society for Information Science and Technology: JASIST, 63 (5):997–1016, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2013:TTS

[LSD13] Chenliang Li, Aixin Sun, and Anwitaman Datta. TSDW: Two-stage word sense disambiguation using Wikipedia. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1203–1223, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2011:SBR

[LSJR11] Jung-Tae Lee, Jangwon Seo, Jiwoon Jeon, and Hae-Chang Rim. Sentence-based relevance flow analysis for high accuracy retrieval. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1666–1675, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2004:BCI

[LSZ04] Li-Juan Liu, Xu-Bang Shen, and Xue-Cheng Zou. Brief communication: An improved fast encoding algorithm for vector quantization. Journal of the American Society for Information Science and Technology: JASIST, 55(1):81–87, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2013:AIS

[LSZC13] Carole J. Lee, Cassidy R. Sugimoto, Guo Zhang, and Blaise Cronin. Advances in information science: Bias in peer review. Journal of the American Society for Information Science and Technology: JASIST, 64(1):2–17, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lucas:2002:FFI

[LT02] Wendy Lucas and Heikki Topi. Form and function: the impact of query term and operator usage on Web search results. Journal of the American Society for Information Science and Technology: JASIST, 53(2):95–108, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lucas:2004:TWS

[LT04] Wendy Lucas and Heikki Topi. Training for Web search: Will it get you in shape? Journal of the American Society for Information Science and Technology: JASIST, 55(13): 1183–1198, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Levitt:2008:MRM

[LT08] Jonathan M. Levitt and Mike Thelwall. Is multidisciplinary research more highly cited? A macrolevel study. Journal of the American Society for Information Science and Technology: JASIST, 59(12):1973–1984, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2009:RIC

[LT09a] Charlotte P. Lee and Ciaran B. Trace. The role of information in a community of hobbyist collectors. Journal of the American Society for Information Science and Technology: JASIST, 60(3):621–637, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Levitt:2009:CLC

[LT09b] Jonathan M. Levitt and Mike Thelwall. Citation levels and collaboration within library and information science. Journal of the American Society for Information Science and Technology: JASIST, 60(3):434–442, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luyt:2010:IWC

[LT10] Brendan Luyt and Daniel Tan. Improving Wikipedia’s credibility: References and citations in a sample of history articles. Journal of the American Society for Information Science and Technology: JASIST, 61(4):715–722, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

LaBarre:2012:ETL

[LT12] Kathryn A. La Barre and Carol L. Tilley. The elusive tale: leveraging the study of information seeking and knowledge organization to improve access to and discovery of folktales. Journal of the American Society for Information Science and Technology: JASIST, 63(4):687–701, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2013:RTU

[LT13] Xiaozhong Liu and Howard Turtle. Real-time user interest modeling for real-time ranking. Journal of the American Society for Information Science and Technology: JASIST, 64 (8):1557–1576, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Levitt:2011:VBS

[LTO11] Jonathan M. Levitt, Mike Thelwall, and Charles Oppenheim. Variations between subjects in the extent to which the social sciences have become more interdisciplinary. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1118–1129, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luk:2009:BRB

[Luk09] Robert Luk. Book review: A Generative Theory of Relevance. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2587–2588, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lunin:2009:PDIb

[Lun09] Lois F. Lunin. Perspectives on design: Information technologies and creative practices: Design: the vision and the plans: Additional resources. Journal of the American Society for Information Science and Technology: JASIST, 60(9): 1942, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lustria:2007:CIM

[Lus07] Mia Liza A. Lustria. Can interactivity make a difference? Effects of interactivity on the comprehension of and attitudes toward online health content. Journal of the American Society for Information Science and Technology: JASIST, 58 (6):766–776, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luyt:2011:NHR

[Luy11] Brendan Luyt. The nature of historical representation on Wikipedia: Dominant or alterative historiography? Journal of the American Society for Information Science and Technology: JASIST, 62(6):1058–1065, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luyt:2012:IWD

[Luy12] Brendan Luyt. The inclusivity of Wikipedia and the drawing of expert boundaries: an examination of talk pages and reference lists. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1868–1878, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luyt:2013:HWN

[Luy13] Brendan Luyt. History on Wikipedia: In need of a NWICO (New World Information and Communication Order)? The case of Cambodia. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1193– 1202, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Luzon:2009:SHF

[Luz09] María José Luzón. Scholarly hyperwriting: the function of links in academic weblogs. Journal of the American Society for Information Science and Technology: JASIST, 60(1):75– 89, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2006:COM

[LV06] Loet Leydesdorff and Liwen Vaughan. Co-occurrence matrices and their applications in information science: Extending ACA to the Web environment. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1616–1628, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Llorens:2004:AGD

[LVdA^{+}04] Juan Lloréns, Manuel Velasco, Antonio de Amescua, José A. Moreiro, and Vicente Martínez. Automatic generation of domain representations using thesaurus structures. Journal of the American Society for Information Science and Technology: JASIST, 55(10):846–858, August 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lu:2012:MAR

[LW12] Kun Lu and Dietmar Wolfram. Measuring author research relatedness: a comparison of word-based, topic-based, and author cocitation approaches. Journal of the American Society for Information Science and Technology: JASIST, 63 (10):1973–1986, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2011:OBT

[LWH11] Yen-Hsien Lee, Chih-Ping Wei, and Paul Jen-Hwa Hu. An ontology-based technique for preserving user preferences in document-category evolutions. Journal of the American Society for Information Science and Technology: JASIST, 62 (3):507–520, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lam:2001:RCD

[LWW01] Wai Lam, Kam-Fai Wong, and Chi-Yin Wong. Research: Chinese document indexing based on a new partitioned signature file: Model and evaluation. Journal of the American Society for Information Science and Technology: JASIST, 52(7):584–597, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2001:TAC

[LWY01] Wenjie Li, Kam-Fai Wong, and Chunfa Yuan. Toward automatic Chinese temporal information extraction. Journal of the American Society for Information Science and Technology: JASIST, 52(9):748–762, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lee:2009:BOE

[LWY09] Yue-Shi Lee, Yu-Chieh Wu, and Jie-Chi Yang. BVideoQA: Online English/ Chinese bilingual video question answering. Journal of the American Society for Information Science and Technology: JASIST, 60(3):509–525, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lin:2013:BCT

[LX13] Shinjeng Lin and Iris Xie. Behavioral changes in transmuting multisession successive searches over the web. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1259–1283, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2005:ACT

[LY05] Kar Wing Li and Christopher C. Yang. Automatic crosslingual thesaurus generated from the Hong Kong SAR Police Department Web corpus for crime analysis. Journal of the American Society for Information Science and Technology: JASIST, 56(3):272–282, February 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2006:CAP

[LY06] Kar Wing Li and Christopher C. Yang. Conceptual analysis of parallel corpus collected from the Web. Journal of the American Society for Information Science and Technology: JASIST, 57(5):632–644, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lu:2011:SGI

[LY11] Li Lu and Y. Connie Yuan. Shall I Google it or ask the competent villain down the hall? The moderating role of information need in information source selection. Journal of the American Society for Information Science and Technology: JASIST, 62(1):133–145, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2010:WHC

[LYJ^{+}10] Xinhai Liu, Shi Yu, Frizo Janssens, Wolfgang Glänzel, Yves Moreau, and Bart De Moor. Weighted hybrid clustering by combining text mining and bibliometrics on a large-scale journal database. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1105–1119, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lam:2007:MWR

[LYM07] Wai Lam, Christopher C. Yang, and Filippo Menczer. Mining Web resources for enhancing information retrieval: Introduction to the special topic section on mining Web resources for enhancing information retrieval. Journal of the American Society for Information Science and Technology: JA-SIST, 58(12):1791–1792, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lynch:2001:WDD

[Lyn01] Clifford A. Lynch. When documents deceive: Trust and provenance as new factors for information retrieval in a tangled web. Journal of the American Society for Information Science and Technology: JASIST, 52(1):12–17, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2008:BCC

[LZ08a] Loet Leydesdorff and Ping Zhou. Brief communications: Coword analysis using the Chinese character set. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1528–1530, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Lin:2008:QDS

[LZ08b] Ming Lin and Zhu Zhang. Question-driven segmentation of lecture speech text: Towards intelligent e-learning systems. Journal of the American Society for Information Science and Technology: JASIST, 59(2):186–200, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Leydesdorff:2013:HCJ

[LZB13] Loet Leydesdorff, Ping Zhou, and Lutz Bornmann. How can journal impact factors be normalized across fields of science? An assessment in terms of percentile ranks and fractional counts. Journal of the American Society for Information Science and Technology: JASIST, 64(1):96–107, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2007:MWRb

[LZC^{+}07] Yiqun Liu, Min Zhang, Rongwei Cen, Liyun Ru, and Shaoping Ma. Mining Web resources for enhancing information retrieval: Data cleansing for Web information retrieval using query independent features. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1884–1898, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2009:ECS

[LZC09] Jia Li, Pengzhu Zhang, and Jinwei Cao. External concept support for group support systems through Web mining. Journal of the American Society for Information Science and Technology: JASIST, 60(5):1057–1070, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Liu:2013:FTC

[LZG13] Xiaozhong Liu, Jinsong Zhang, and Chun Guo. Full-text citation analysis: a new method to enhance scholarly networks. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1852–1863, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Li:2008:KBL

[LZLC08] Jiexun Li, Zhu Zhang, Xin Li, and Hsinchun Chen. Kernelbased learning for biomedical relation extraction. Journal of the American Society for Information Science and Technology: JASIST, 59(5):756–769, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ma:2002:DSE

[Ma02a] Wei Ma. A database selection expert system based on reference librarian’s database selection strategy: a usability and empirical evaluation. Journal of the American Society for Information Science and Technology: JASIST, 53(7):567– 580, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ma:2002:DAM

[Ma02b] Yan Ma. A design analysis model for developing World Wide Web sites. Journal of the American Society for Information Science and Technology: JASIST, 53(7):531–535, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ma:2007:BRB

[Ma07] Hongyan Ma. Book review: Process-aware information systems: Bridging people and software through process technology. Journal of the American Society for Information Science and Technology: JASIST, 58(3):455–456, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ma:2012:MIA

[Ma12a] Lai Ma. Meanings of information: the assumptions and research consequences of three foundational LIS theories. Journal of the American Society for Information Science and Technology: JASIST, 63(4):716–723, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ma:2012:SPC

[Ma12b] Lai Ma. Some philosophical considerations in using mixed methods in library and information science research. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1859–1867, September 2012. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mervyn:2012:SCI

[MA12c] Kieran Mervyn and David K. Allen. Sociospatial context and information behavior: Social exclusion and the influence of mobile information technology. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1125–1141, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mutawa:2008:HAA

[MAA08] A. M. Mutawa, Salah Alnajem, and Fadi Alzhouri. An HPSG approach to Arabic nominal sentences. Journal of the American Society for Information Science and Technology: JASIST, 59(3):422–434, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Macpherson:2004:IPM

[Mac04a] Karen Macpherson. An information processing model of undergraduate electronic database information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 55(4):333–347, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Macpherson:2004:LES

[Mac04b] Karen Macpherson. Letter to the Editor: Summary of article appearing in “In This Issue,” JASIST, 55(4), 281–282 (2004): An Information Processing Model of Undergraduate Electronic Database Information Retrieval. Journal of the American Society for Information Science and Technology: JASIST, 55(14):1306, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Boy04d].

Mackenzie:2006:AAS

[Mac06] Maureen L. Mackenzie. ASIS to ASIS&T: a society in transition? Journal of the American Society for Information Science and Technology: JASIST, 57(13):1750–1757, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mai:2013:QQI

[Mai13] Jens-Erik Mai. The quality and qualities of information. Journal of the American Society for Information Science and Technology: JASIST, 64(4):675–688, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Matia:2005:SPG

[MAL^{+}05] Kaushik Matia, Luis A. Nunes Amaral, Marc Luwel, Henk F. Moed, and H. Eugene Stanley. Scaling phenomena in the growth dynamics of scientific output. Journal of the American Society for Information Science and Technology: JA-SIST, 56(9):893–902, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marchionini:2002:CEU

[Mar02] Gary Marchionini. Co-evolution of user and organizational interfaces: a longitudinal case study of WWW dissemination of national statistics. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1192– 1209, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mustafa:2004:UGA

[MAR04] Suleiman H. Mustafa and Qasem A. Al-Radaideh. Using Ngrams for Arabic text searching. Journal of the American Society for Information Science and Technology: JASIST, 55(11):1002–1007, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Markey:2007:TFYa

[Mar07a] Karen Markey. Twenty-five years of end-user searching, Part 1: Research findings. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1071– 1081, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Markey:2007:TFYb

[Mar07b] Karen Markey. Twenty-five years of end-user searching, Part 2: Future research directions. Journal of the American Society for Information Science and Technology: JASIST, 58(8): 1123–1130, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Martinez:2007:PMM

[Mar07c] José M. Martínez. Perspectives on MPEG-7: Metadata for multimedia: MPEG-7 tools for Universal Multimedia Access. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1374–1376, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marty:2007:CNI

[Mar07d] Paul F. Marty. The changing nature of information work in museums. Journal of the American Society for Information Science and Technology: JASIST, 58(1):97–107, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marchant:2009:SBB

[Mar09a] Thierry Marchant. Score-based bibliometric rankings of authors. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1132–1137, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marcus:2009:PDI

[Mar09b] Susan Marcus. Perspectives on design: Information technologies and creative practices: New basics for new literacies. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1933–1938, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marx:2011:SFH

[Mar11] Werner Marx. Special features of historical papers from the viewpoint of bibliometrics. Journal of the American Society for Information Science and Technology: JASIST, 62 (3):433–439, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Matzat:2009:QIA

[Mat09] Uwe Matzat. Quality of information in academic e-mailing lists. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1859–1870, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Matusiak:2013:IMR

[Mat13] Krystyna K. Matusiak. Image and multimedia resources in an academic environment: a qualitative study of students’ experiences and literacy practices. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1577–1589, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Maxwell:2004:MIP

[Max04] Terrence A. Maxwell. Mapping information policy frames: the politics of the Digital Millennium Copyright Act. Journal of the American Society for Information Science and Technology: JASIST, 55(1):3–12, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Maxwell:2005:PPD

[Max05] Terrence A. Maxwell. Parsing the public domain. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1130–1139, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mizrachi:2013:UPA

[MB13] Diane Mizrachi and Marcia J. Bates. Undergraduates’ personal academic information management and the consideration of time and task-urgency. Journal of the American Society for Information Science and Technology: JASIST, 64(8): 1590–1607, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Miksa:2007:DFA

[MBBK07] Shawne D. Miksa, Kathleen Burnett, Laurie J. Bonnici, and Joonmin Kim. The development of a facet analysis system to identify and measure the dimensions of interaction in online learning. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1569–1577, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mehra:2002:SAP

[MBBS02] Bharat Mehra, Ann Peterson Bishop, Imani Bazzell, and Cynthia Smith. Scenarios in the Afya project as a participatory action research (PAR) tool for studying information seeking and use across the “digital divide”. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1259–1266, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Makri:2008:UIB

[MBC08] Stephann Makri, Ann Blandford, and Anna L. Cox. Using information behaviors to evaluate the functionality and usability of electronic resources: From Ellis’s model to evaluation. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2244–2267, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marx:2010:RSR

[MBC10] Werner Marx, Lutz Bornmann, and Manuel Cardona. Reference standards and reference multipliers for the comparison of the citation impact of papers published in different time periods. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2061–2069, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Morillo:2003:IST

[MBG03] Fernanda Morillo, María Bordons, and Isabel Gómez. Interdisciplinarity in science: a tentative typology of disciplines and research areas. Journal of the American Society for Information Science and Technology: JASIST, 54(13):1237– 1249, November 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Makri:2007:LJA

[MBG^{+}07] Stephann Makri, Ann Blandford, Jeremy Gow, Jon Rimmer, Claire Warwick, and George Buchanan. A library or just another information resource? A case study of users’ mental models of traditional and digital libraries. Journal of the American Society for Information Science and Technology: JASIST, 58(3):433–445, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Modjeska:2003:IDE

[MC03a] David Modjeska and Mark Chignell. Individual differences in exploration using desktop VR. Journal of the American Society for Information Science and Technology: JASIST, 54(3):216–228, February 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Morgan:2003:BRB

[MC03b] Allison Morgan and Cong Chen. Book review: Integrity, internal control and security in information systems: Connecting governance and technology. Journal of the American Society for Information Science and Technology: JASIST, 54(10):976–977, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McArthur:2012:UPA

[MC12] David McArthur and Helen Crompton. Understanding public-access cyberlearning projects using text mining and topic analysis. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2146–2152, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McCain:2008:AAI

[McC08] Katherine W. McCain. Assessing an author’s influence using time series historiographic mapping: the oeuvre of Conrad Hal Waddington (1905–1975). Journal of the American Society for Information Science and Technology: JASIST, 59 (4):510–525, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McCain:2009:UTD

[McC09] Katherine W. McCain. Using tricitation to dissect the citation image: Conrad Hal Waddington and the rise of evolutionary developmental biology. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1301–1319, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McCain:2011:EOI

[McC11] Katherine W. McCain. Eponymy and obliteration by incorporation: the case of the “Nash Equilibrium”. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1412–1424, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McCain:2012:AOI

[McC12] Katherine W. McCain. Assessing obliteration by incorporation: Issues and caveats. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2129– 2139, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McDonald:2007:UJU

[McD07] John D. McDonald. Understanding journal usage: a statistical analysis of citation and use. Journal of the American Society for Information Science and Technology: JASIST, 58(1):39–50, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McDonough:2011:PVL

[McD11] Jerome P. McDonough. Packaging videogames for long-term preservation: Integrating FRBR and the OAIS reference model. Journal of the American Society for Information Science and Technology: JASIST, 62(1):171–184, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mothe:2003:DMD

[MCDA03] Josiane Mothe, Claude Chrisment, Bernard Dousset, and Joel Alaux. DocCube: Multi-dimensional visualisation and exploration of large document sets. Journal of the American Society for Information Science and Technology: JASIST, 54(7):650–659, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McInerney:2002:KMD

[McI02] Claire McInerney. Knowledge management and the dynamic nature of knowledge. Journal of the American Society for Information Science and Technology: JASIST, 53(12):1009– 1018, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McInerney:2010:BRB

[McI10] Claire R. McInerney. Book review: Managing Knowledge Networks; Mastering Organizational Knowledge Flow: How to Make Knowledge Sharing Work. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2372–2374, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McKenzie:2004:PTN

[McK04] Pamela J. McKenzie. Positioning theory and the negotiation of information needs in a clinical midwifery setting. Journal of the American Society for Information Science and Technology: JASIST, 55(8):685–694, June 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marshall:2008:UIF

[MCK08] Byron Marshall, Hsinchun Chen, and Siddharth Kaza. Using importance flooding to identify interesting networks of criminal activity. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2099–2114, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McLeod:2010:BRBa

[McL10a] Julie McLeod. Book review: Leading and Managing Archives and Records Programs: Strategies for Success. Journal of the American Society for Information Science and Technology: JASIST, 61(4):855–858, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McLeod:2010:BRBb

[McL10b] Julie McLeod. Book review: Managing Electronic Records, 4th Ed. Journal of the American Society for Information Science and Technology: JASIST, 61(5):1075, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McQuaid:2009:PDI

[McQ09] Matilda McQuaid. Perspectives on design: Information technologies and creative practices: From craft to production: Technology transfer in extreme textiles. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1920–1922, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Miguel:2011:OAS

[MCRdMA11] Sandra Miguel, Zaida Chinchilla-Rodriguez, and Félix de Moya-Anegón. Open access and Scopus: a new approach to scientific visibility from the standpoint of access. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1130–1145, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McInerney:2002:EIB

[MD02] Claire McInerney and Ronald Day. Editorial: Introduction to the JASIST special issue on Knowledge Management. Journal of the American Society for Information Science and Technology: JASIST, 53(12):1008, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

MacMullen:2005:IPM

[MD05] W. John MacMullen and Sheila O. Denn. Information problems in molecular biology and bioinformatics. Journal of the American Society for Information Science and Technology: JASIST, 56(5):447–456, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meso:2006:MIC

[MDM06] Peter Meso, Pratim Datta, and Victor Mbarika. Moderating information and communication technologies’ influences on socioeconomic development with good governance: a study of the developing countries. Journal of the American Society for Information Science and Technology: JASIST, 57(2): 186–197, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Malone:2003:ICE

[ME03] Cheryl Knott Malone and Fernando Elichirigoity. Information as commodity and economic sector: Its emergence in the discourse of industrial classification. Journal of the American Society for Information Science and Technology: JA-SIST, 54(6):512–520, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meadows:2010:BRB

[Mea10] Charles Meadows III. Book review: Spychips: How Major Corporations and Government Plan to Track Your Every Purchase and Watch Your Every Move. Journal of the American Society for Information Science and Technology: JASIST, 61(3):635–636, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meghabghab:2001:RGW

[Meg01] George Meghabghab. Research: Google’s web page ranking applied to different topological web graph structures. Journal of the American Society for Information Science and Technology: JASIST, 52(9):736–747, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meho:2006:MIQ

[Meh06] Lokman I. Meho. E-mail interviewing in qualitative research: a methodological discussion. Journal of the American Society for Information Science and Technology: JASIST, 57 (10):1284–1295, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Moore:2007:SEV

[MEH07] Joi L. Moore, Sanda Erdelez, and Wu He. The search experience variable in information behavior research. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1529–1546, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Melucci:2004:MDL

[Mel04] Massimo Melucci. Making digital libraries effective: Automatic generation of links for similarity search across hypertextbooks. Journal of the American Society for Information Science and Technology: JASIST, 55(5):414–430, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mendelsohn:2003:IOC

[Men03] Loren D. Mendelsohn. Introduction and overview: Chemistry journals: the transition from paper to electronic, with lessons for other disciplines. Journal of the American Society for Information Science and Technology: JASIST, 54(12):

1136–1137, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Menczer:2004:LSC

[Men04] Filippo Menczer. Lexical and semantic clustering by Web links. Journal of the American Society for Information Science and Technology: JASIST, 55(14):1261–1269, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Merriman:2009:BRB

[Mer09] Kevin W. Merriman. Book review: What’s the Alternative: Career Options for Librarians and Info Pros. Journal of the American Society for Information Science and Technology: JASIST, 60(3):641–642, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Metzger:2007:MSC

[Met07] Miriam J. Metzger. Making sense of credibility on the Web: Models for evaluating online information and recommendations for future research. Journal of the American Society for Information Science and Technology: JASIST, 58(13): 2078–2091, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Morris:2007:MRT

[MG07] Steven A. Morris and Michel L. Goldstein. Manifestation of research teams in journal literature: a growth model of papers, authors, collaboration, coauthorship, weak ties, and Lotka’s Law. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1764–1782, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mengle:2009:PDU

[MG09a] Saket Mengle and Nazli Goharian. Passage detection using text classification. Journal of the American Society for Information Science and Technology: JASIST, 60(4):814–825, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mengle:2009:AMF

[MG09b] Saket S. R. Mengle and Nazli Goharian. Ambiguity measure feature-selection algorithm. Journal of the American Society for Information Science and Technology: JASIST, 60(5): 1037–1050, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mengle:2010:DRA

[MG10] Saket S. R. Mengle and Nazli Goharian. Detecting relationships among categories using text classification. Journal of the American Society for Information Science and Technology: JASIST, 61(5):1046–1061, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Muresan:2004:TMM

[MH04] Gheorghe Muresan and David J. Harper. Topic modeling for mediated access to very large document collections. Journal of the American Society for Information Science and Technology: JASIST, 55(10):892–910, August 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marcial:2010:SDR

[MH10] Laura Haak Marcial and Bradley M. Hemminger. Scientific data repositories on the Web: an initial survey. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2029–2048, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Maghsoodi:2011:IFM

[MH11] Nooshin Maghsoodi and Mohammad Mehdi Homayounpour. Improving Farsi multiclass text classification using a thesaurus and two-stage feature selection. Journal of the American Society for Information Science and Technology: JA-SIST, 62(10):2055–2066, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mat-Hassan:2005:ASN

[MHL05] Mazlita Mat-Hassan and Mark Levene. Associating search and navigation behavior through log analysis. Journal of the American Society for Information Science and Technology: JASIST, 56(9):913–934, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mardis:2012:TBI

[MHM12] Marcia A. Mardis, Ellen S. Hoffman, and Flora P. McMartin. Toward broader impacts: Making sense of NSF’s merit review criteria in the context of the National Science Digital Library. Journal of the American Society for Information Science and Technology: JASIST, 63(9):1758–1772, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mulligan:2013:PRC

[MHR13] Adrian Mulligan, Louise Hall, and Ellen Raphael. Peer review in a changing world: an international study measuring the attitudes of researchers. Journal of the American Society for Information Science and Technology: JASIST, 64(1): 132–161, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Milojevic:2010:MCM

[Mil10a] Staša Milojević. Modes of collaboration in modern science: Beyond power laws and preferential attachment. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1410–1423, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Milojevic:2010:PLD

[Mil10b] Staša Milojević. Power law distributions in information science: Making the case for logarithmic binning. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2417–2425, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Miller:2012:BRB

[Mil12] Emily Miller. Book reviews: Information and living systems: Philosophical and scientific perspectives. Journal of the American Society for Information Science and Technology: JASIST, 63(2):425–427, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Minguillo:2010:TNW

[Min10] David Minguillo. Toward a new way of mapping scientific fields: Authors’ competence for publishing in scholarly journals. Journal of the American Society for Information Science and Technology: JASIST, 61(4):772–786, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Misa:2009:BRB

[Mis09] Thomas J. Misa. Book review: The Internet and American Business. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2165–2167, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mitchell:2008:BRB

[Mit08] Nicole Mitchell. Book review: Library 2.0: a guide to participatory library service. Journal of the American Society for Information Science and Technology: JASIST, 59(5):855, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mizzaro:2003:QCS

[Miz03] Stefano Mizzaro. Quality control in scholarly publishing: a new proposal. Journal of the American Society for Information Science and Technology: JASIST, 54(11):989–1005, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mizzaro:2012:OPR

[Miz12] Stefano Mizzaro. Opinion paper: Readersourcing — a manifesto. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1666–1672, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Muresan:2013:ITT

[MK13] Smaranda Muresan and Judith L. Klavans. Inducing terminologies from text: a case study for the consumer health domain. Journal of the American Society for Information Science and Technology: JASIST, 64(4):727–744, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meso:2009:RDS

[MKMM09] Peter Meso, Bangaly Kaba, Victor Mbarika, and Richard McCline. Regional differences in stakeholders’ perspectives on strategies to overcome teledensity limitations: an exploratory comparison of selected Latin American and Sub-Saharan African countries. Journal of the American Society for Information Science and Technology: JASIST, 60(1): 184–200, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Morse:2002:TVI

[MLO02] Emile Morse, Michael Lewis, and Kai A. Olsen. Testing visual information retrieval methodologies case study: Comparative analysis of textual, icon, graphical, and “spring” displays. Journal of the American Society for Information Science and Technology: JASIST, 53(1):28–40, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Miao:2010:FGO

[MLZ10] Qingliang Miao, Qiudan Li, and Daniel Zeng. Fine-grained opinion mining by integrating multiple review sources. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2288–2299, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marion:2001:CVS

[MM01] Linda S. Marion and Katherine W. McCain. Contrasting views of software engineering journals: Author cocitation choices and indexer vocabulary assignments. Journal of the American Society for Information Science and Technology: JASIST, 52(4):297–308, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mea:2004:MRE

[MM04] Vincenzo Della Mea and Stefano Mizzaro. Measuring retrieval effectiveness: a new proposal and a first experimental validation. Journal of the American Society for Information Science and Technology: JASIST, 55(6):530–543, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

MacRoberts:2010:PCA

[MM10] M. H. MacRoberts and B. R. MacRoberts. Problems of citation analysis: a study of uncited and seldom-cited influences.

Journal of the American Society for Information Science and Technology: JASIST, 61(1):1–12, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McKemmish:2009:CET

[MMBF09] Sue McKemmish, Rosetta Manaszewicz, Frada Burstein, and Julie Fisher. Consumer empowerment through metadata-based information quality reporting: The Breast Cancer Knowledge Online Portal. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1792–1807, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

MacIntosh-Murray:2005:IBC

[MMC05] Anu MacIntosh-Murray and Chun Wei Choo. Information behavior in the context of improving patient safety. Journal of the American Society for Information Science and Technology: JASIST, 56(12):1332–1345, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marshall:2004:ECA

[MMCC04] Byron Marshall, Daniel McDonald, Hsinchun Chen, and Wingyan Chung. EBizPort: Collecting and analyzing business intelligence information. Journal of the American Society for Information Science and Technology: JASIST, 55 (10):873–891, August 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Muller:2001:EIB

[MMMM^{+}01] Wolfgang Müller, Stéphane Marchand-Maillet, Henning Müller, David McG. Squire, and Thierry Pun. Evaluating image browsers using structured annotation. Journal of the American Society for Information Science and Technology: JASIST, 52(11):961–968, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Margaritopoulos:2012:QMM

[MMMM12] Merkourios Margaritopoulos, Thomas Margaritopoulos, Ioannis Mavridis, and Athanasios Manitsaris. Quantifying and measuring metadata completeness. Journal of the American Society for Information Science and Technology: JASIST, 63(4):724–737, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Melucci:2004:CMP

[MO04] Massimo Melucci and Nicola Orio. Combining melody processing and information retrieval techniques: Methodology, evaluation, and system implementation. Journal of the American Society for Information Science and Technology: JASIST, 55(12):1058–1066, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Melucci:2007:DIE

[MO07] Massimo Melucci and Nicola Orio. Design, implementation, and evaluation of a methodology for automatic stemmer generation. Journal of the American Society for Information Science and Technology: JASIST, 58(5):673–686, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Montesi:2008:CJP

[MO08] Michela Montesi and John Mackenzie Owen. From conference to journal publication: How conference papers in software engineering are extended for publication in journals. Journal of the American Society for Information Science and Technology: JASIST, 59(5):816–829, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mochel:2006:PSU

[Moc06] Karl Mochel. Perspectives on search user interfaces: Best practices and future visions: Search in the Web shopping environment. Journal of the American Society for Information Science and Technology: JASIST, 57(6):819–824, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Moed:2005:SRB

[Moe05] Henk F. Moed. Statistical relationships between downloads and citations at the level of individual documents within a single journal. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1088–1097, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Moed:2007:EOA

[Moe07] Henk F. Moed. The effect of “open access” on citation impact: an analysis of ArXiv’s condensed matter section. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2047–2054, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Moed:2011:LES

[Moe11] Henk F. Moed. Letters to the Editor: The source normalized impact per paper is a valid and sophisticated indicator of journal citation impact. Journal of the American Society for Information Science and Technology: JASIST, 62(1): 211–213, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mon:2004:BRB

[Mon04] Lorri Mon. Book review: Chat reference: a guide to live virtual reference services. Journal of the American Society for Information Science and Technology: JASIST, 55(4):366– 368, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Morris:2005:MES

[Mor05] Steven A. Morris. Manifestation of emerging specialties in journal literature: a growth model of papers, references, exemplars, bibliographic coupling, cocitation, and clustering coefficient distribution. Journal of the American Society for Information Science and Technology: JASIST, 56(12): 1250–1273, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Morris:2010:IDI

[Mor10] Jane Morris. Individual differences in the interpretation of text: Implications for information science. Journal of the American Society for Information Science and Technology: JASIST, 61(1):141–149, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mostafa:2004:EDS

[Mos04] Javed Mostafa. Editorial: Document search interface design: Background and introduction to special topic section. Journal of the American Society for Information Science and Technology: JASIST, 55(10):869–872, August 2004. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mosco:2009:BRB

[Mos09] Vincent Mosco. Book review: Working-Class Network Society: Communication Technology and the Information Have-Less in Urban China. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2593– 2595, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mukhopadhyay:2003:MAI

[MPR^{+}03] Snehasis Mukhopadhyay, Shengquan Peng, Rajeev Raje, Mathew Palakal, and Javed Mostafa. Multi-agent information classification using dynamic acquaintance lists. Journal of the American Society for Information Science and Technology: JASIST, 54(10):966–975, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mukhopadhyay:2005:DMA

[MPR^{+}05] S. Mukhopadhyay, S. Peng, R. Raje, J. Mostafa, and M. Palakal. Distributed multi-agent information filtering — a comparative study. Journal of the American Society for Information Science and Technology: JASIST, 56(8):834– 842, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Makris:2012:WQD

[MPS12] Christos Makris, Yannis Plegas, and Sofia Stamou. Web query disambiguation using PageRank. Journal of the American Society for Information Science and Technology: JA-SIST, 63(8):1581–1592, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McCay-Peet:2009:IUW

[MPT09] Lori McCay-Peet and Elaine Toms. Image use within the work task model: Images as information and illustration. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2416–2429, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Martin:2013:BRP

[MQH13] Kim Martin and Anabel Quan-Haase. Are e-books replacing print books? Tradition, serendipity, and opportunity in the adoption and use of e-books for historical research and teaching. Journal of the American Society for Information Science and Technology: JASIST, 64(5):1016–1028, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meho:2008:CCC

[MR08] Lokman I. Meho and Yvonne Rogers. Citation counting, citation ranking, and h-index of human-computer interaction researchers: a comparison of Scopus and Web of Science. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1711–1726, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marcoux:2009:PDI

[MR09] Yves Marcoux and Élias Rizkallah. Perspectives on design: Information technologies and creative practices: Intertextual semantics: a semantics for information design. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1895–1906, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Maglaughlin:2002:UPR

[MS02] Kelly L. Maglaughlin and Diane H. Sonnenwald. User perspectives on relevance criteria: a comparison among relevant, partially relevant, and not-relevant judgments. Journal of the American Society for Information Science and Technology: JASIST, 53(5):327–342, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meho:2005:RRP

[MS05] Lokman I. Meho and Kristina M. Spurgin. Ranking the research productivity of library and information science faculty and schools: an evaluation of data sources and research methods. Journal of the American Society for Information Science and Technology: JASIST, 56(12):1314–1331, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

McCool:2009:FDC

[MS09a] Matthew McCool and Kirk St. Amant. Field dependence and classification: Implications for global information systems. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1258–1266, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meho:2009:ASI

[MS09b] Lokman I. Meho and Cassidy R. Sugimoto. Assessing the scholarly impact of information studies: a tale of two citation databases — Scopus and Web of Science. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2499–2508, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ma:2013:PPN

[MSC13] Zongyang Ma, Aixin Sun, and Gao Cong. On predicting the popularity of newly emerging hashtags in Twitter. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1399–1410, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Malo:2011:CBD

[MSWK11] Pekka Malo, Ankur Sinha, Jyrki Wallenius, and Pekka Korhonen. Concept-based document classification using Wikipedia and value function. Journal of the American Society for Information Science and Technology: JASIST, 62 (12):2496–2511, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Milojevic:2011:CSL

[MSYD11] Staša Milojević, Cassidy R. Sugimoto, Erjia Yan, and Ying Ding. The cognitive structure of Library and Information Science: Analysis of article title words. Journal of the American Society for Information Science and Technology: JA-SIST, 62(10):1933–1953, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meho:2003:MIS

[MT03] Lokman I. Meho and Helen R. Tibbo. Modeling the information-seeking behavior of social scientists: Ellis’s study revisited. Journal of the American Society for Information Science and Technology: JASIST, 54(6):570–587, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mirel:2013:SPU

[MTS^{+}13] Barbara Mirel, Jennifer Steiner Tonks, Jean Song, Fan Meng, Weijian Xuan, and Rafiqa Ameziane. Studying PubMed usages in the field for complex problem solving: Implications for tool design. Journal of the American Society for Information Science and Technology: JASIST, 64 (5):874–892, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mueller:2010:BRB

[Mue10] Milton Mueller. Book review: Internet Governance: Infrastructure and Institutions. Journal of the American Society for Information Science and Technology: JASIST, 61 (7):1511–1512, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mukherjee:2009:DOA

[Muk09] Bhaskar Mukherjee. Do open-access journals in library and information science have any scholarly impact? A bibliometric study of selected open-access journals using Google Scholar. Journal of the American Society for Information Science and Technology: JASIST, 60(3):581–594, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Mustafa:2005:WOA

[Mus05] Suleiman H. Mustafa. Word-oriented approximate string matching using occurrence heuristic tables: a heuristic for searching Arabic text. Journal of the American Society for Information Science and Technology: JASIST, 56(14):1504– 1511, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Moreno:2008:TIW

[MV08] Nathalie Moreno and Antonio Vallecillo. Towards interoperable Web engineering methods. Journal of the American Society for Information Science and Technology: JASIST, 59(7):1073–1092, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Medina:2012:SJC

[MvL12] C. M. Calero Medina and T. N. van Leeuwen. Seed journal citation network maps: a method based on network theory. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1226–1234, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Medelyan:2008:DIA

[MW08] Olena Medelyan and Ian H. Witten. Domain-independent automatic keyphrase indexing with small training sets. Journal of the American Society for Information Science and Technology: JASIST, 59(7):1026–1040, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Makri:2010:IIU

[MW10] Stephann Makri and Claire Warwick. Information for inspiration: Understanding architects’ information seeking and use behaviors to inform design. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1745–1770, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

MacKay:2012:EMW

[MW12] Bonnie MacKay and Carolyn Watters. An examination of multisession web tasks. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1183– 1197, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Marchionini:2006:OVD

[MWG06] Gary Marchionini, Barbara M. Wildemuth, and Gary Geisler. The Open Video Digital Library: A Möbius strip of research and practice. Journal of the American Society for Information Science and Technology: JASIST, 57(12): 1629–1643, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Meho:2007:IDS

[MY07] Lokman I. Meho and Kiduk Yang. Impact of data sources on citation counts and rankings of LIS faculty: Web of Science versus Scopus and Google Scholar. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2105–2125, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Morris:2003:TLV

[MYWA03] Steven A. Morris, G. Yen, Zheng Wu, and Benyam Asnake. Time line visualization of research fronts. Journal of the American Society for Information Science and Technology: JASIST, 54(5):413–422, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nahl:2007:SBI

[Nah07] Diane Nahl. Social–biological information technology: an integrated conceptual framework. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2021–2046, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nathan:2012:SIP

[Nat12] Lisa P. Nathan. Sustainable information practice: an ethnographic investigation. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2254– 2268, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nayar:2004:BRB

[Nay04] Pramod K. Nayar. Book review: The Internet in everyday life. Journal of the American Society for Information Science and Technology: JASIST, 55(3):278–279, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nayar:2005:BRBb

[Nay05a] Pramod K. Nayar. Book review: How images think. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1126–1127, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nayar:2005:BRBa

[Nay05b] Pramod K. Nayar. Book review: Windows and mirrors: Interaction design, digital art, and the myth of transparency.

Journal of the American Society for Information Science and Technology: JASIST, 56(7):773–774, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nayar:2006:BRB

[Nay06] Pramod K. Nayar. Book review: Shaping the network society: the new role of civil society in cyberspace. Journal of the American Society for Information Science and Technology: JASIST, 57(5):724–725, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nayar:2007:BRBb

[Nay07a] Pramod K. Nayar. Book review: Human rights in the global information society. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2407– 2408, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nayar:2007:BRBa

[Nay07b] Pramod K. Nayar. Book review: Media ecologies: Materialist energies in art and technoculture. Journal of the American Society for Information Science and Technology: JA-SIST, 58(8):1222, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Newman:2006:PTD

[NB06] David J. Newman and Sharon Block. Probabilistic topic decomposition of an eighteenth-century American newspaper. Journal of the American Society for Information Science and Technology: JASIST, 57(6):753–767, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Naaman:2011:HTC

[NBG11] Mor Naaman, Hila Becker, and Luis Gravano. Hip and trendy: Characterizing emerging trends on Twitter. Journal of the American Society for Information Science and Technology: JASIST, 62(5):902–918, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Navarro:2003:MFA

[NBYA03] Gonzalo Navarro, Ricardo Baeza-Yates, and João Marcelo Azevedo Arcoverde. Matchsimile: a flexible approximate matching tool for searching proper names. Journal of the American Society for Information Science and Technology: JASIST, 54(1):3–15, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nguyen:2013:IKM

[NC13] Son Hoang Nguyen and Gobinda Chowdhury. Interpreting the knowledge map of digital library research (1990–2010). Journal of the American Society for Information Science and Technology: JASIST, 64(6):1235–1258, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nicholas:2013:IGC

[NCRJ13] David Nicholas, David Clark, Ian Rowlands, and Hamid R. Jamali. Information on the go: a case study of Europeana mobile users. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1311–1322, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Neveol:2009:NLP

[NDD^{+}09] Aurélie Névéol, Thomas M. Deserno, Stéfan J. Darmoni, Mark Oliver Güld, and Alan R. Aronson. Natural language processing versus content-based image analysis for medical document retrieval. Journal of the American Society for Information Science and Technology: JASIST, 60(1):123–134, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Neal:2008:BRB

[Nea08] Diane Neal. Book review: Facebook: the missing manual. Journal of the American Society for Information Science and Technology: JASIST, 59(12):2035–2036, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Neal:2011:BRB

[Nea11] Diane Rasmussen Neal. Book review: Atlas of Science: Visualizing What We Know. Journal of the American Society for Information Science and Technology: JASIST, 62(6): 1212–1213, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nelson:2006:BRB

[Nel06] Jeremy Nelson. Book review: Putting XML to work in the library: Tools for improving access and management. Journal of the American Society for Information Science and Technology: JASIST, 57(2):294–295, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Newby:2001:CSI

[New01] Gregory B. Newby. Cognitive space and information space. Journal of the American Society for Information Science and Technology: JASIST, 52(12):1026–1048, ???? 2001. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Newby:2003:OSS

[NGJ03] Gregory B. Newby, Jane Greenberg, and Paul Jones. Open source software development and Lotka’s Law: Bibliometric patterns in programming. Journal of the American Society for Information Science and Technology: JASIST, 54(2): 169–178, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See comments [Bur04b].

Niu:2012:SFA

[NH12] Xi Niu and Bradley M. Hemminger. A study of factors that affect the information-seeking behavior of academic scientists. Journal of the American Society for Information Science and Technology: JASIST, 63(2):336–353, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Naidoo:2013:MRB

[NHCW13] Jeff Naidoo, Jeffrey T. Huber, Pamela Cupp, and Qishan Wu. Modeling the relationship between an emerging infectious disease epidemic and the body of scientific literature associated with it: the case of HIV/ AIDS in the United States. Journal of the American Society for Information Science and Technology: JASIST, 64(2):380–391, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Niemi:2003:MDM

[NHJ03] Timo Niemi, Lasse Hirvonen, and Kalervo Järvelin. Multidimensional Data Model and Query Language for Informetrics. Journal of the American Society for Information Science and Technology: JASIST, 54(10):939–951, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Niu:2010:NSI

[NHL^{+}10] Xi Niu, Bradley M. Hemminger, Cory Lown, Stephanie Adams, Cecelia Brown, Allison Level, Merinda McLure, Audrey Powers, Michele R. Tennant, and Tara Cataldo. National study of information seeking behavior of academic researchers in the United States. Journal of the American Society for Information Science and Technology: JASIST, 61 (5):869–890, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nicholson:2003:BAC

[Nic03a] Scott Nicholson. Bibliomining for automated collection development in a digital library setting: Using data mining to discover Web-based scholarly research works. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1081–1090, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nicholson:2003:BRB

[Nic03b] Scott Nicholson. Book review: Models for library management, decision-making, and planning. Journal of the American Society for Information Science and Technology: JA-SIST, 54(4):359–360, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nicolaisen:2009:CNL

[Nic09] Jeppe Nicolaisen. Compromised need and the label effect: an examination of claims and evidence. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2004–2009, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nicolaisen:2010:BRB

[Nic10] Jeppe Nicolaisen. Book review: Bibliometrics and Citation Analysis: From the Science Citation Index to Cybermetrics. Journal of the American Society for Information Science and Technology: JASIST, 61(1):205–207, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nie:2003:QEQ

[Nie03] Jian-Yun Nie. Query expansion and query translation as logical inference. Journal of the American Society for Information Science and Technology: JASIST, 54(4):335–346, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Niemi:2007:QLDa

[NJ07a] Timo Niemi and Janne Jämsen. A query language for discovering semantic associations, Part I: Approach and formal definition of query primitives. Journal of the American Society for Information Science and Technology: JASIST, 58 (11):1559–1568, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Niemi:2007:QLDb

[NJ07b] Timo Niemi and Janne Jämsen. A query language for discovering semantic associations, Part II: sample queries and query evaluation. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1686–1700, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nappila:2008:TDC

[NJN08] Turkka Näppilä, Kalervo Järvelin, and Timo Niemi. A tool for data cube construction from structurally heterogeneous XML documents. Journal of the American Society for Information Science and Technology: JASIST, 59(3):435–449, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nah:2009:BER

[NKLC09] In Wook Nah, Dai-Shin Kang, Dae-Hee Lee, and Yun-Chul Chung. A bibliometric evaluation of research performance in different subject categories. Journal of the American Society for Information Science and Technology: JASIST, 60(6): 1138–1143, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ng:2006:AJD

[NKS^{+}06] Kwong Bor Ng, Paul Kantor, Tomek Strzalkowski, Nina Wacholder, Rong Tang, Bing Bai, Robert Rittman, Peng Song, and Ying Sun. Automated judgment of document qualities. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1155–1164, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Naing:2006:ELC

[NLC06] Myo-Myo Naing, Ee-Peng Lim, and Roger H. L. Chiang. Extracting link chains of relationship instances from a Web site. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1590–1605, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Neuhaus:2009:PCI

[NMD09] Christoph Neuhaus, Werner Marx, and Hans-Dieter Daniel. The publication and citation impact profiles of Angewandte Chemie and the Journal of the American Chemical Society based on the sections of Chemical Abstracts: a case study on the limitations of the Journal Impact Factor. Journal of the American Society for Information Science and Technology: JASIST, 60(1):176–183, January 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nehmadi:2011:PSA

[NMPBA11] Liron Nehmadi, Joachim Meyer, Yisrael Parmet, and Noam Ben-Asher. Predicting a screen area’s perceived importance from spatial and physical attributes. Journal of the American Society for Information Science and Technology: JA-SIST, 62(9):1829–1838, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nov:2010:APO

[NNY10] Oded Nov, Mor Naaman, and Chen Ye. Analysis of participation in an online photo-sharing community: a multidimensional perspective. Journal of the American Society for Information Science and Technology: JASIST, 61(3):555– 566, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Norris:2008:CAO

[NOR08] Michael Norris, Charles Oppenheim, and Fytton Rowland. The citation advantage of open-access articles. Journal of the American Society for Information Science and Technology: JASIST, 59(12):1963–1972, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nunes:2011:TWB

[NRD11] Sérgio Nunes, Cristina Ribeiro, and Gabriel David. Term weighting based on document revision history. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2471–2478, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nicholson:2007:ULH

[NS07] Scott Nicholson and Catherine Arnott Smith. Using lessons from health care to protect the privacy of library users: Guidelines for the de-identification of library data based on HIPAA. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1198–1206, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nov:2012:DRC

[NS12] Oded Nov and William Schecter. Dispositional resistance to change and hospital physicians’ use of electronic medical records: a multidimensional perspective. Journal of the American Society for Information Science and Technology: JASIST, 63(4):648–656, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nicholson:2006:HMI

[NSE^{+}06] Scott Nicholson, Tito Sierra, U. Yeliz Eseryel, Ji-Hong Park, Philip Barkow, Erika J. Pozo, and Jane Ward. How much of it is real? Analysis of paid placement in Web search engine results. Journal of the American Society for Information Science and Technology: JASIST, 57(4):448–461, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ni:2013:VAC

[NSJ13] Chaoqun Ni, Cassidy R. Sugimoto, and Jiepu Jiang. Venue-author-coupling: a measure for identifying disciplines through author communities. Journal of the American Society for Information Science and Technology: JASIST, 64(2): 265–279, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ni:2013:JIP

[NSLD13] Chaoqun Ni, Debora Shaw, Sean M. Lind, and Ying Ding. Journal impact and proximity: an assessment using bibliographic features. Journal of the American Society for Information Science and Technology: JASIST, 64(4):802–817, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nakashima:2003:BBC

[NSQI03] Makoto Nakashima, Keizo Sato, Yanhua Qu, and Tetsuro Ito. Browsing-based conceptual information retrieval incorporating dictionary term relations, keyword association, and a user’s interest. Journal of the American Society for Information Science and Technology: JASIST, 54(1):16–28, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Numminen:2009:QTP

[NV09] Piritta Numminen and Pertti Vakkari. Question types in public libraries’ digital reference service in Finland: Comparing 1999 and 2006. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1249– 1257, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nov:2008:BCU

[NY08] Oded Nov and Chen Ye. Brief communications: Users’ personality and perceived ease of use of digital libraries: the case for resistance to change. Journal of the American Society for Information Science and Technology: JASIST, 59 (5):845–851, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Nov:2009:BCR

[NY09] Oded Nov and Chen Ye. Brief communications: Resistance to change and the adoption of digital libraries: an integrative model. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1702–1708, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Narock:2013:SSO

[NZY13] Tom Narock, Lina Zhou, and Victoria Yoon. Semantic similarity of ontology instances using polarity mining. Journal of the American Society for Information Science and Technology: JASIST, 64(2):416–427, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ortega:2012:SAE

[OA12] José Luis Ortega and Isidro F. Aguillo. Science is all in the eye of the beholder: Keyword maps in Google Scholar citations. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2370–2377, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Oakleaf:2009:URA

[Oak09] Megan Oakleaf. Using rubrics to assess information literacy: an examination of methodology and interrater reliability. Journal of the American Society for Information Science and Technology: JASIST, 60(5):969–983, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ozel:2004:MBM

[ÖAU^{+}04] S. Ayse Özel, I. Sengör Altingövde, Özgür Ulusoy, Gültekin Özsoyoğlu, and Z. Meral Özsoyoğlu. Metadata-based modeling of information resources on the Web. Journal of the American Society for Information Science and Technology: JASIST, 55(2):97–110, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ozcan:2011:ENQ

[OAU11] Rifat Ozcan, Ismail Sengor Altingovde, and Ozgür Ulusoy. Exploiting navigational queries for result presentation and caching in Web search engines. Journal of the American Society for Information Science and Technology: JASIST, 62(4):714–726, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

OBrien:2007:BRB

[O’B07] Heather L. O’Brien. Book review: Exploring information systems research approaches: Readings and reflections. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2402–2403, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

OBrien:2012:BRB

[O’B12] Heather O’Brien. Book review: Inventing the Medium: Principles of Interaction Design as a Cultural Practice. By Janet H. Murray. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2345–2348, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

OConnor:2013:IIS

[O’C13] Lisa G. O’Connor. Investors’ information sharing and use in virtual communities. Journal of the American Society for Information Science and Technology: JASIST, 64(1):36–47, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

ONeill:2008:EAL

[OCD08] Edward T. O’Neill, Lynn Silipigni Connaway, and Timothy J. Dickey. Estimating the audience level for library resources. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2042–2050, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ortiz-Cordova:2012:CWS

[OCJ12] Adan Ortiz-Cordova and Bernard J. Jansen. Classifying web search queries to identify high revenue generating customers. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1426–1441, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ozmutlu:2008:CVN

[OCO08] H. Cenk Ozmutlu, Fatih Cavdur, and Seda Ozmutlu. Crossvalidation of neural network applications for automatic new topic identification. Journal of the American Society for Information Science and Technology: JASIST, 59(3):339– 362, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Oh:2012:CMH

[Oh12] Sanghee Oh. The characteristics and motivations of health answerers for sharing information, knowledge, and experiences in online environments. Journal of the American Society for Information Science and Technology: JASIST, 63 (3):543–557, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Onodera:2011:MEA

[OIM^{+}11] Natsuo Onodera, Mariko Iwasawa, Nobuyuki Midorikawa, Fuyuki Yoshikane, Kou Amano, Yutaka Ootani, Tadashi Kodama, Yasuhiko Kiyama, Hiroyuki Tsunoda, and Shizuka Yamazaki. A method for eliminating articles by homonymous authors from the large number of articles retrieved by author search. Journal of the American Society for Information Science and Technology: JASIST, 62(4):677–690, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Onwuegbuzie:2004:ISP

[OJ04] Anthony J. Onwuegbuzie and Qun G. Jiao. Information search performance and research achievement: an empirical test of the Anxiety-Expectation Mediation model of library anxiety. Journal of the American Society for Information Science and Technology: JASIST, 55(1):41–54, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ou:2007:AMS

[OKG07] Shiyan Ou, Christopher S. G. Khoo, and Dion H. Goh. Automatic multidocument summarization of research abstracts: Design and user evaluation. Journal of the American Society for Information Science and Technology: JASIST, 58 (10):1419–1435, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ozkar:2006:MRS

[ÖL06] Mine Özkar and Nyssim Lefford. Modal relationships as stylistic features: Examples from Seljuk and Celtic patterns. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1551–1560, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

OBrien:2013:MMA

[OL13] Heather L. O’Brien and Mahria Lebow. Mixed-methods approach to measuring user experience in online news interactions. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1543–1556, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ouyang:2010:IIM

[OLLL10] You Ouyang, Wenjie Li, Sujian Li, and Qin Lu. Intertopic information mining for query-based summarization. Journal of the American Society for Information Science and Technology: JASIST, 61(5):1062–1072, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Osareh:2008:SIC

[OM08] Farideh Osareh and Katherine W. McCain. The structure of Iranian chemistry research, 1990–2006: an author cocitation analysis. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2146–2155, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Orduna-Malea:2013:AWP

[OM13] Enrique Orduña-Malea. Aggregation of the web performance of internal university units as a method of quantitative analysis of a university system: the case of Spain. Journal of the American Society for Information Science and Technology: JASIST, 64(10):2100–2114, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Orzano:2008:KMM

[OMS^{+}08] A. John Orzano, Claire R. McInerney, Davida Scharf, Alfred F. Tallia, and Benjamin F. Crabtree. A knowledge management model: Implications for enhancing quality in health care. Journal of the American Society for Information Science and Technology: JASIST, 59(3):489–505, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Oppenheim:2007:BCU

[Opp07] Charles Oppenheim. Brief communication: Using the hindex to rank influential British researchers in information science and librarianship. Journal of the American Society for Information Science and Technology: JASIST, 58 (2):297–301, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Oppenheim:2009:BRB

[Opp09] Charles Oppenheim. Book review: The Public Domain: Enclosing the Commons of the Mind. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2593, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Oppenheim:2010:BRB

[Opp10] Charles Oppenheim. Book review: Teaching of Intellectual Property: Principles and Methods. Journal of the American Society for Information Science and Technology: JASIST, 61(1):210–211, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Orme:2001:BRB

[Orm01] Marianne Orme. Book review: A place at the table: Participating in community building, by Kathleen de la Pena McCook. Journal of the American Society for Information Science and Technology: JASIST, 52(5):439–441, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Orme:2007:BRB

[Orm07] Marianne Orme. Book review: Radio frequency identification handbook for librarians. Journal of the American Society for Information Science and Technology: JASIST, 58 (14):2408–2409, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ostergren:2008:BRB

[Ost08] Marilyn Ostergren. Book review: The principles of beautiful Web design. Journal of the American Society for Information Science and Technology: JASIST, 59(1):159, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

OBrien:2008:WUE

[OT08] Heather L. O’Brien and Elaine G. Toms. What is user engagement? A conceptual framework for defining user engagement with technology. Journal of the American Society for Information Science and Technology: JASIST, 59 (6):938–955, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

OBrien:2010:DES

[OT10] Heather L. O’Brien and Elaine G. Toms. The development and evaluation of a survey to measure user engagement. Journal of the American Society for Information Science and Technology: JASIST, 61(1):50–69, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ostermann:2013:UEA

[OTP13] Frank O. Ostermann, Martin Tomko, and Ross Purves. User evaluation of automatically generated keywords and toponyms for geo-referenced images. Journal of the American Society for Information Science and Technology: JASIST, 64(3):480–499, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Oluic-Vukovic:2001:IKS

[OV01] Vesna Oluić-Vuković. From information to knowledge: Some reflections on the origin of the current shifting towards knowledge processing and further perspective. Journal of the American Society for Information Science and Technology: JASIST, 52(1):54–61, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Olsen:2004:BCS

[OW04] Kai A. Olsen and James G. Williams. Brief communication: Spelling and grammar checking using the Web as a text repository. Journal of the American Society for Information Science and Technology: JASIST, 55(11):1020–1023, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Oyarce:2008:USR

[Oya08] Guillermo Oyarce. Using the shape recovery method to evaluate indexing techniques. Journal of the American Society for Information Science and Technology: JASIST, 59(9): 1479–1492, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ofoghi:2009:IFS

[OYM09] Bahadorreza Ofoghi, John Yearwood, and Liping Ma. The impact of frame semantic annotation levels, frame-alignment techniques, and fusion methods on factoid answer processing. Journal of the American Society for Information Science and Technology: JASIST, 60(2):247–263, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Palmer:2005:SWS

[Pal05] Carole L. Palmer. Scholarly work and the shaping of digital access. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1140–1153, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Paling:2008:TGV

[Pal08] Stephen Paling. Technology, genres, and value change: Literary authors and artistic use of information technology. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1238–1251, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Piwowarski:2012:UQP

[PAL12] B. Piwowarski, M. R. Amini, and M. Lalmas. On using a quantum physics formalism for multidocument summarization. Journal of the American Society for Information Science and Technology: JASIST, 63(5):865–888, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Parmanto:2010:BRB

[Par10] Bambang Parmanto. Book review: Web Accessibility: A Foundation for Research. Journal of the American Society for Information Science and Technology: JASIST, 61(3): 634–635, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Piro:2013:MAP

[PAR13] Fredrik Niclas Piro, Dag W. Aksnes, and Kristoffer Rørstad. A macro analysis of productivity differences across fields: Challenges in the measurement of scientific publishing. Journal of the American Society for Information Science and Technology: JASIST, 64(2):307–320, February 2013. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pattuelli:2011:MDO

[Pat11] M. Cristina Pattuelli. Modeling a domain ontology for cultural heritage resources: a user-centered approach. Journal of the American Society for Information Science and Technology: JASIST, 62(2):314–342, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Peters:2001:ERL

[PB01] Carol Peters and Martin Braschler. European research letter: Cross-language system evaluation: The CLEF campaigns. Journal of the American Society for Information Science and Technology: JASIST, 52(12):1067–1072, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Park:2002:HAN

[PBN02] Han Woo Park, George A. Barnett, and In-Yong Nam. Hyperlink–affiliation network structure of top web sites: Examining affiliates with hyperlink in Korea. Journal of the American Society for Information Science and Technology: JASIST, 53(7):592–601, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Petrelli:2004:OUD

[PBS^{+}04] Daniela Petrelli, Micheline Beaulieu, Mark Sanderson, George Demetriou, Patrick Herring, and Preben Hansen.

Observing users, designing clarity: a case study on the usercentered design of a cross-language information retrieval system. Journal of the American Society for Information Science and Technology: JASIST, 55(10):923–934, August 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Poulos:2008:TSE

[PBV08] Marios Poulos, George Bokos, and Fotios Vaioulis. Towards the semantic extraction of digital signatures for librarian image-identification purposes. Journal of the American Society for Information Science and Technology: JASIST, 59 (5):708–718, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Prime-Claverie:2004:TCM

[PCBL04] Camille Prime-Claverie, Michael Beigbeder, and Thierry Lafouge. Transposition of the cocitation method with a view to classifying Web pages. Journal of the American Society for Information Science and Technology: JASIST, 55(14): 1282–1289, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Panicheva:2013:ISS

[PCR13] Polina Panicheva, John Cardiff, and Paolo Rosso. Identifying subjective statements in news titles using a personal sense annotation framework. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1411– 1422, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pu:2002:SCQ

[PCY02] Hsiao-Tieh Pu, Shui-Lung Chuang, and Chyan Yang. Subject categorization of query terms for exploring Web users’ search interests. Journal of the American Society for Information Science and Technology: JASIST, 53(8):617–630, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Polat:2008:PPT

[PD08] Huseyin Polat and Wenliang Du. Privacy-preserving top- N recommendation on distributed data. Journal of the American Society for Information Science and Technology: JA-SIST, 59(7):1093–1108, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pino-Diaz:2012:SKM

[PDJCRBBM12] José Pino-Díaz, Evaristo Jiménez-Contreras, Rosario Ruíz-Baños, and Rafael Bailón-Moreno. Strategic knowledge maps of the techno-scientific network (SK maps). Journal of the American Society for Information Science and Technology: JASIST, 63(4):796–804, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pettigrew:2002:FCI

[PDU02] Karen E. Pettigrew, Joan C. Durrance, and Kenton T. Unruh. Facilitating community information seeking using the Internet: Findings from three public library–community network systems. Journal of the American Society for Information Science and Technology: JASIST, 53(11):894–903, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Peekhaus:2012:BRB

[Pee12] Wilhelm Peekhaus. Book review: Information Markets: A Strategic Guideline for the I-Commerce. By Frank Linde and Wolfgang G. Stock. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1477– 1479, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Peled:2011:WTC

[Pel11] Alon Peled. When transparency and collaboration collide: The USA Open Data program. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2085–2094, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pepe:2011:RBA

[Pep11] Alberto Pepe. The relationship between acquaintanceship and coauthorship in scientific collaboration networks. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2121–2132, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Perry:2003:NIS

[Per03] Claudia A. Perry. Network influences on scholarly communication in developmental dyslexia: a longitudinal followup. Journal of the American Society for Information Science and Technology: JASIST, 54(14):1278–1295, December 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Petrou:2002:BRB

[Pet02] Anastasis (Tassos) D. Petrou. Book review: Saving the time of the library user through subject access innovation: Papers in honor of Pauline Atherton Cochrane, edited by William J. Wheeler. Journal of the American Society for Information Science and Technology: JASIST, 53(1):58–60, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Petrou:2003:BR

[Pet03] Anastasis D. Petrou. Book review. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1168–1170, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Petrou:2007:BRBa

[Pet07a] Anastasis D. Petrou. Book review: Introducing information management: an information research reader. Journal of the American Society for Information Science and Technology: JASIST, 58(4):607–608, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Petrou:2007:BRBb

[Pet07b] Anastasis D. Petrou. Book review: Metadata and its impact on libraries. Journal of the American Society for Information Science and Technology: JASIST, 58(6):909–910, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Peterson:2012:BRB

[Pet12] Gabriel M. Peterson. Book review: The Global Flow of Information — Legal, Social and Cultural Perspectives. By Ramesh Subramanian and Eddan Katz. Journal of the American Society for Information Science and Technology:

JASIST, 63(4):843–844, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Peterson:2013:CRO

[Pet13] Gabriel M. Peterson. Characteristics of retracted open access biomedical literature: a bibliographic analysis. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2428–2436, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pudovkin:2002:APF

[PG02] Alexander I. Pudovkin and Eugene Garfield. Algorithmic procedure for finding semantically related journals. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1113–1119, November 2002. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pabon:2013:LOD

[PGFMP13] Gustavo Pabón, Claudio Gutiérrez, Javier D. Fernández, and Miguel A. Martínez-Prieto. Linked Open Data technologies for publication of census microdata. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1802–1814, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pras:2013:ITA

[PGL13] Amandine Pras, Catherine Guastavino, and Maryse Lavoie. The impact of technological advances on recording studio practices. Journal of the American Society for Information Science and Technology: JASIST, 64(3):612–626, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pluye:2013:FLO

[PGR^{+}13] Pierre Pluye, Roland Grad, Carol Repchinsky, Barbara Jovaisas, Janique Johnson-Lafleur, Marie-Eve Carrier, Vera Granikov, Barbara Farrell, Charo Rodriguez, Gillian Bartlett, Carmen Loiselle, and France Légaré. Four levels of outcomes of information-seeking: a mixed methods study in primary health care. Journal of the American Society for Information Science and Technology: JASIST, 64(1):108–125, January 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pan:2006:ODL

[PGSH06] Bing Pan, Geri Gay, John Saylor, and Helene Hembrooke. One digital library, two undergraduate classes, and four learning modules: Uses of a digital library in classrooms. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1315–1325, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pillay:2013:LEA

[Pil13] Avin Pillay. Letters to the Editor: Academic promotion and the h-index. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2598–2599, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pirkola:2001:RER

[PJ01] Ari Pirkola and Kalervo Järvelin. Research: Employing the resolution power of search keys. Journal of the American Society for Information Science and Technology: JASIST, 52(7):575–583, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pharo:2006:ISI

[PJ06] Nils Pharo and Kalervo Järvelin. “irrational” searchers and IR-rational researchers. Journal of the American Society for Information Science and Technology: JASIST, 57(2):222– 232, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pahlevi:2005:CTC

[PK05] Said Mirza Pahlevi and Hiroyuki Kitagawa. Conveying taxonomy context for topic-focused Web search. Journal of the American Society for Information Science and Technology: JASIST, 56(2):173–188, January 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pellack:2011:REW

[PK11a] Lorraine J. Pellack and Lori Osmus Kappmeyer. The ripple effect of women’s name changes in indexing, citation, and authority control. Journal of the American Society for Information Science and Technology: JASIST, 62(3):440–448, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pharo:2011:ETT

[PK11b] Nils Pharo and Astrid Krahn. The effect of task type on preferred element types in an XML-based retrieval system. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1717–1726, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Petrelli:2006:WUI

[PLBS06] Daniela Petrelli, Stephen Levin, Micheline Beaulieu, and Mark Sanderson. Which user interaction for cross-language information retrieval? Design issues and reflections. Journal of the American Society for Information Science and Technology: JASIST, 57(5):709–722, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pera:2009:SLS

[PLN09] Maria Soledad Pera, William Lund, and Yiu-Kai Ng. A sophisticated library search strategy using folksonomies and similarity matching. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1392– 1406, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Plotnick:2010:CST

[Plo10] Rachel Plotnick. Computers, systems theory, and the making of a wired hospital: a history of Technicon Medical Information System, 1964–1987. Journal of the American Society for Information Science and Technology: JASIST, 61(6): 1281–1294, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pachet:2004:PMA

[PLZA04] François Pachet, Amaury La Burthe, Aymeric Zils, and Jean-Julien Aucouturier. Popular music access: The Sony music browser. Journal of the American Society for Information Science and Technology: JASIST, 55(12):1037–1044, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pettigrew:2001:UTI

[PM01] Karen E. Pettigrew and Lynne (E. F.) McKechnie. The use of theory in information science research. Journal of the American Society for Information Science and Technology: JASIST, 52(1):62–73, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Paling:2011:TUI

[PM11] Stephen Paling and Crystle Martin. Transformative use of information technology in American literary writing: a pilot survey of literary community members. Journal of the American Society for Information Science and Technology: JASIST, 62(5):947–962, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pepe:2010:AAM

[PMBdS10] Alberto Pepe, Matthew Mayernik, Christine L. Borgman, and Herbert Van de Sompel. From artifacts to aggregations: Modeling scientific life cycles on the semantic Web. Journal of the American Society for Information Science and Technology: JASIST, 61(3):567–582, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pal:2011:EER

[PMK11] Sukomal Pal, Mandar Mitra, and Jaap Kamps. Evaluation effort, reliability and reusability in XML retrieval. Journal of the American Society for Information Science and Technology: JASIST, 62(2):375–394, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Paling:2006:TGV

[PN06] Stephen Paling and Michael Nilan. Technology, genres, and value change: the case of little magazines. Journal of the American Society for Information Science and Technology: JASIST, 57(7):862–872, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pera:2009:SSM

[PN09] Maria Soledad Pera and Yiu-Kai Ng. SpamED: a spam Email detection approach based on phrase similarity. Journal of the American Society for Information Science and Technology: JASIST, 60(2):393–409, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pan:2007:OKM

[PNHG07] Shan L. Pan, Sue Newell, Jimmy Huang, and Robert D. Galliers. Overcoming knowledge management challenges during ERP implementation: the need to integrate and share different types of knowledge. Journal of the American Society for Information Science and Technology: JASIST, 58(3):404– 419, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Panzarasa:2009:PDU

[POC09] Pietro Panzarasa, Tore Opsahl, and Kathleen M. Carley. Patterns and dynamics of users’ behavior and interaction: Network analysis of an online community. Journal of the American Society for Information Science and Technology: JASIST, 60(5):911–932, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Poder:2010:LEL

[Põd10] Endel Põder. Letter to the Editor: Let’s correct that small mistake. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2593–2594, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pomerantz:2005:CFO

[Pom05a] Jeffrey Pomerantz. A conceptual framework and open research questions for chat-based reference service. Journal of the American Society for Information Science and Technology: JASIST, 56(12):1288–1302, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pomerantz:2005:LAQ

[Pom05b] Jeffrey Pomerantz. A linguistic analysis of question taxonomies. Journal of the American Society for Information Science and Technology: JASIST, 56(7):715–728, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Perea-Ortega:2013:IPC

[POMVULMC13] José M. Perea-Ortega, M. Teresa Martín-Valdivia, L. Alfonso Ureña-López, and Eugenio Martínez-Cámara. Improving polarity classification of bilingual parallel corpora combining machine learning and semantic orientation approaches. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1864–1877, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ponelis:2001:BRB

[Pon01] Shana R. Ponelis. Book review: Digital capital: Harnessing the power of business webs, by Don Tapscott, David Ticoll, & Alex Lowy. Journal of the American Society for Information Science and Technology: JASIST, 52(5):438–439, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Poworoznek:2003:LEC

[Pow03] Emily L. Poworoznek. Linking of errata: Current practices in online physical sciences journals. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1153–1159, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pan:2006:DLA

[PPH06] Shan Pan, Gary Pan, and Ming H. Hsieh. A dual-level analysis of the capability development process: a case study of TT&T. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1814–1829, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pan:2007:ISP

[PPN07] Gary Pan, Shan L. Pan, and Michael Newman. Information systems project post-mortems: Insights from an attribution perspective. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2255– 2268, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Perugini:2007:MWR

[PR07] Saverio Perugini and Naren Ramakrishnan. Mining Web resources for enhancing information retrieval: Mining Web functional dependencies for flexible information access. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1805–1819, October 2007. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pouratashi:2010:AFI

[PR10] Mahtab Pouratashi and Ahmad Rezvanfar. Analysis of factors influencing application of ICT by agricultural graduate students. Journal of the American Society for Information Science and Technology: JASIST, 61(1):81–87, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Prathap:2010:LEB

[Pra10] Gangan Prathap. Letter to the Editor: Big Macs and Eigenfactor scores: the correlation conundrum. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2378–2379, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Prathap:2011:BCT

[Pra11a] Gangan Prathap. Brief communications: A thermodynamic explanation for the Glänzel–Schubert model for the h-index. Journal of the American Society for Information Science and Technology: JASIST, 62(5):992–994, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Prathap:2011:BCF

[Pra11b] Gangan Prathap. Brief communications: Fractionalized exergy for evaluating research performance. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2294–2295, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Prathap:2012:LEI

[Pra12a] Gangan Prathap. Letter to the Editor: The inconsistency of the H-index. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1480–1481, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Prathap:2012:LET

[Pra12b] Gangan Prathap. Letter to the Editor: The thermodynamicsbibliometrics consilience and the meaning of h-type indices. Journal of the American Society for Information Science and Technology: JASIST, 63(2):430, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Prentice:2009:BRB

[Pre09] Ann E. Prentice. Book review: Knowledge Management in Practice: Connections and Context. Journal of the American Society for Information Science and Technology: JASIST, 60(3):642, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pereira:2011:GWB

[PRNZ^{+}11] Denilson Alves Pereira, Berthier Ribeiro-Neto, Nivio Ziviani, Alberto H. F. Laender, and Marcos André Gonçalves. A generic Web-based entity resolution framework. Journal of the American Society for Information Science and Technology: JASIST, 62(5):919–932, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Picard:2003:ERH

[PS03] Justin Picard and Jacques Savoy. Enhancing retrieval with hyperlinks: a general model based on propositional argumentation systems. Journal of the American Society for Information Science and Technology: JASIST, 54(4):347– 355, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pardo:2004:NTP

[PSB04] Bryan Pardo, Jonah Shifrin, and William Birmingham. Name that tune: a pilot study in finding a melody from a sung query. Journal of the American Society for Information Science and Technology: JASIST, 55(4):283–300, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pahor:2008:ENP

[PŠD08] Marko Pahor, Miha Škerlavaj, and Vlado Dimovski. Evidence for the network perspective on organizational learning. Journal of the American Society for Information Science and Technology: JASIST, 59(12):1985–1994, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Perer:2006:URR

[PSO06] Adam Perer, Ben Shneiderman, and Douglas W. Oard. Using rhythms of relationships to understand e-mail archives. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1936–1948, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Prichard:2013:YPC

[PSWL13] Jeremy Prichard, Caroline Spiranovic, Paul Watters, and Christopher Lueg. Young people, child pornography, and subcultural norms on the Internet. Journal of the American Society for Information Science and Technology: JASIST, 64(5):992–1000, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Price:2005:BCC

[PT05] Liz Price and Mike Thelwall. Brief communication: The clustering power of low frequency words in academic Webs. Journal of the American Society for Information Science and Technology: JASIST, 56(8):883–888, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Pennanen:2003:SCS

[PV03] Mikko Pennanen and Pertti Vakkari. Students’ conceptual structure, search process, and outcome while preparing a research proposal: a longitudinal case study. Journal of the American Society for Information Science and Technology: JASIST, 54(8):759–770, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Parmanto:2005:MWA

[PZ05] Bambang Parmanto and Xiaoming Zeng. Metric for Web accessibility evaluation. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1394– 1404, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Peng:2012:WYP

[PZ12] Tai-Quan Peng and Jonathan J. H. Zhu. Where you publish matters most: a multilevel analysis of factors affecting citations of Internet studies. Journal of the American Society for Information Science and Technology: JASIST, 63 (9):1789–1803, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Quirin:2008:QMB

[QCGB^{+}08] Arnaud Quirin, Oscar Cordón, Vicente P. Guerrero-Bote, Benjamín Vargas-Quesada, and Felix Moya-Anegón. A quick MST-based algorithm to obtain Pathfinder networks (∞, n−1). Journal of the American Society for Information Science and Technology: JASIST, 59(12):1912–1924, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Qu:2012:ECM

[QCL^{+}12] Bo Qu, Gao Cong, Cuiping Li, Aixin Sun, and Hong Chen. An evaluation of classification models for question topic categorization. Journal of the American Society for Information Science and Technology: JASIST, 63(5):889–903, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Qin:2006:BIV

[QH06] Jian Qin and Naybell Hernández. Building interoperable vocabulary and structures for learning objects. Journal of the American Society for Information Science and Technology: JASIST, 57(2):280–292, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Qin:2006:MWR

[QZCC06] Jialun Qin, Yilu Zhou, Michael Chau, and Hsinchun Chen. Multilingual Web retrieval: an experiment in English– Chinese business intelligence. Journal of the American Society for Information Science and Technology: JASIST, 57 (5):671–683, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Raban:2008:ISO

[Rab08] Daphne Ruth Raban. The incentive structure in an online information market. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2284– 2295, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Raban:2009:SPV

[Rab09] Daphne Ruth Raban. Self-presentation and the value of information in Q&A websites. Journal of the American Society for Information Science and Technology: JASIST, 60 (12):2465–2473, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Radford:2006:EVU

[Rad06] Marie L. Radford. Encountering virtual users: a qualitative investigation of interpersonal communication in chat reference. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1046–1059, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rawlins:2010:BRB

[Raw10] Gregory J. E. Rawlins. Book review: Glut: Mastering Information Through the Ages. Journal of the American Society for Information Science and Technology: JASIST, 61 (1):207, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Robbin:2008:IIC

[RB08] Alice Robbin and Wayne Buente. Internet information and communication behavior during a political moment: The Iraq war, March 2003. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2210– 2231, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rodriguez-Barcenas:2013:SAH

[RBLH13] Gustavo Rodríguez-Bárcenas and María J. López-Huertas. Saaty’s analytic hierarchies method for knowledge organization in decision making. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1454– 1467, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rorissa:2008:ERB

[RCD08] Abebe Rorissa, Paul Clough, and Thomas Deselaers. Exploring the relationship between feature and perceptual visual spaces. Journal of the American Society for Information Science and Technology: JASIST, 59(5):770–784, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rodriguez:2005:MAR

[RCT05] Andrés Rodríguez, José María Carazo, and Oswaldo Trelles. Mining association rules from biological databases. Journal of the American Society for Information Science and Technology: JASIST, 56(5):493–504, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Renaud:2004:BRB

[Ren04] John P. Renaud. Book review: Democracy and New Media. Journal of the American Society for Information Science and Technology: JASIST, 55(6):561–562, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Reuter:2005:BRB

[Reu05] Kara Reuter. Book review: Weaving a library Web: a guide to developing children’s Websites. Journal of the American Society for Information Science and Technology: JASIST, 56(14):1555–1556, December 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Reuter:2007:AAR

[Reu07] Kara Reuter. Assessing aesthetic relevance: Children’s book selection in a digital library. Journal of the American Society for Information Science and Technology: JASIST, 58(12): 1745–1763, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Radev:2005:PQA

[RFQ^{+}05] Dragomir Radev, Weiguo Fan, Hong Qi, Harris Wu, and Amardeep Grewal. Probabilistic question answering on the Web. Journal of the American Society for Information Science and Technology: JASIST, 56(6):571–583, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Romero-Frias:2010:EPT

[RFV10] Esteban Romero-Frías and Liwen Vaughan. European political trends viewed through patterns of Web linking. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2109–2121, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Romero-Frias:2012:ERB

[RFV12] Esteban Romero-Frías and Liwen Vaughan. Exploring the relationships between media and political parties through web hyperlink analysis: the case of Spain. Journal of the American Society for Information Science and Technology: JASIST, 63(5):967–976, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Robinson-Garcia:2013:LEA

[RGLCTSC13] Nicolás Robinson-García, Emilio Delgado López-Cózar, Daniel Torres-Salinas, and Juan Miguel Campanario. Letter to the Editor: Against the resilience of rejected manuscripts. Journal of the American Society for Information Science and Technology: JASIST, 64(10):2187–2188, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ramdeen:2012:TTI

[RH12] Sarah Ramdeen and Bradley M. Hemminger. A tale of two interfaces: How facets affect the library catalog search. Journal of the American Society for Information Science and Technology: JASIST, 63(4):702–715, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Roman:2012:EDU

[RHCP12] Jorge H. Román, Kevin J. Hulin, Linn M. Collins, and James E. Powell. Entity disambiguation using semantic networks. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2087–2099, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rhodes:2004:BR

[Rho04] Nicholas Rhodes. Book review. Journal of the American Society for Information Science and Technology: JASIST, 55(1):90–91, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Robins:2010:CHI

[RHS10] David Robins, Jason Holmes, and Mary Stansbury. Consumer health information on the Web: the relationship of visual design and perceptions of credibility. Journal of the American Society for Information Science and Technology: JASIST, 61(1):13–29, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Reimer:2011:TDI

[RHWB11] Yolanda Jacobs Reimer, Matthew Hagedal, Peter Wolf, and Bradley Bahls. Turning the desktop inside-out: Evaluating information access and management through a single interface. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2327–2346, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rorissa:2008:TCI

[RI08] Abebe Rorissa and Hemalata Iyer. Theories of cognition and image categorization: What category labels reveal about basic level theory. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1383–1392, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rieh:2002:JIQ

[Rie02] Soo Young Rieh. Judgment of information quality and cognitive authority in the Web. Journal of the American Society for Information Science and Technology: JASIST, 53 (2):145–161, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rieh:2004:WHI

[Rie04] Soo Young Rieh. On the Web at home: Information seeking and Web searching in the home environment. Journal of the American Society for Information Science and Technology:

JASIST, 55(8):743–753, June 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Riggs:2002:XFT

[Rig02] Ken Roger Riggs. XML and free text. Journal of the American Society for Information Science and Technology: JA-SIST, 53(6):526–528, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rinehart:2009:PDI

[Rin09] Richard Rinehart. Perspectives on design: Information technologies and creative practices: Artists and subversive metadata. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1883–1886, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Riviera:2013:SCA

[Riv13] Emanuela Riviera. Scientific communities as autopoietic systems: the reproductive function of citations. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1442–1453, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rising:2007:PMM

[RJ07] Hawley K. Rising III and Corinne Jörgensen. Perspectives on MPEG-7: Metadata for multimedia: Semantic description in MPEG-7: the rich recursion of ripeness. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1338–1345, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rousseau:2008:ADT

[RJ08] Ronald Rousseau and Bihui Jin. The age-dependent h-type AR^{2}-index: Basic properties and a case study. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2305–2311, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Robbin:2001:SET

[RK01] Alice Robbin and Heather Koball. Seeking explanation in theory: Reflections on the social practices of organizations that distribute public use microdata files for research purposes. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1169–1189, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rokach:2011:WGW

[RKBS11] Lior Rokach, Meir Kalech, Ido Blank, and Rami Stern. Who is going to win the next Association for the Advancement of Artificial Intelligence Fellowship Award? Evaluating researchers by mining bibliographic data. Journal of the American Society for Information Science and Technology: JA-SIST, 62(12):2456–2470, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Roitblat:2010:DCL

[RKO10] Herbert L. Roitblat, Anne Kershaw, and Patrick Oot. Document categorization in legal electronic discovery: computer classification vs. manual review. Journal of the American Society for Information Science and Technology: JASIST, 61 (1):70–80, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rafols:2009:CBA

[RL09] Ismael Rafols and Loet Leydesdorff. Content-based and algorithmic classifications of journals: Perspectives on the dynamics of scientific communication and indexer effects. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1823–1835, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Radev:2002:GAN

[RLF02] Dragomir R. Radev, Kelsey Libner, and Weiguo Fan. Getting answers to natural language questions on the Web. Journal of the American Society for Information Science and Technology: JASIST, 53(5):359–364, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ruthven:2002:CSC

[RLvR02] Ian Ruthven, Mounia Lalmas, and Keith van Rijsbergen. Combining and selecting characteristics of information use. Journal of the American Society for Information Science and Technology: JASIST, 53(5):378–396, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ruthven:2003:IUS

[RLvR03] Ian Ruthven, Mounia Lalmas, and Keith van Rijsbergen. Incorporating user search behavior into relevance feedback. Journal of the American Society for Information Science and Technology: JASIST, 54(6):529–549, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rokach:2013:BCP

[RM13] Lior Rokach and Prasenjit Mitra. Brief communication: Parsimonious citer-based measures: the artificial intelligence domain as a case study. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1951– 1959, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Reitsma:2012:CIB

[RMC12] René Reitsma, Byron Marshall, and Trevor Chart. Can intermediary-based science standards crosswalking work? Some evidence from mining the standard alignment tool (SAT). Journal of the American Society for Information Science and Technology: JASIST, 63(9):1843–1858, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rodriguez-Navarro:2009:SRU

[RN09] Alonso Rodríguez-Navarro. Sound research, unimportant discoveries: Research, universities, and formal evaluation of research in Spain. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1845–1858, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ribeiro-Neto:2001:ESA

[RNLdL01] Berthier Ribeiro-Neto, Alberto H. F. Laender, and Luciano R. S. de Lima. An experimental study in automatically categorizing medical documents. Journal of the American Society for Information Science and Technology: JASIST, 52 (5):391–401, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Roberts:2001:PWT

[Rob01] Helene E. Roberts. A picture is worth a thousand words: Art indexing in electronic databases. Journal of the American Society for Information Science and Technology: JASIST, 52(11):911–916, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Robinson:2010:EAE

[Rob10a] Mark A. Robinson. An empirical analysis of engineers’ information behaviors. Journal of the American Society for Information Science and Technology: JASIST, 61(4):640– 658, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Rob10b].

Robinson:2010:ECR

[Rob10b] Mark A. Robinson. Erratum: Correction to Robinson, M. A. (2010). An empirical analysis of engineers’ information behaviors. Journal of the American Society for Information Science and Technology, 61(4), 640–658. Journal of the American Society for Information Science and Technology: JA-SIST, 61(9):1947, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Rob10a].

Rokach:2012:BCA

[Rok12] Lior Rokach. Brief communication: Applying the publication power approach to artificial intelligence journals. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1270–1277, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rorissa:2007:RBP

[Ror07] Abebe Rorissa. Relationships between perceived features and similarity of images: a test of Tversky’s contrast model. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1401–1418, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rorissa:2010:CSF

[Ror10] Abebe Rorissa. A comparative study of Flickr tags and index terms in a general image collection. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2230–2242, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rosenfeld:2002:IAL

[Ros02] Louis Rosenfeld. Information architecture: Looking ahead. Journal of the American Society for Information Science and Technology: JASIST, 53(10):874–876, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rose:2006:PSU

[Ros06] Daniel E. Rose. Perspectives on search user interfaces: Best practices and future visions: Reconciling informationseeking behavior with search user interfaces for the Web. Journal of the American Society for Information Science and Technology: JASIST, 57(6):797–799, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rosso:2008:UBI

[Ros08] Mark A. Rosso. User-based identification of Web genres. Journal of the American Society for Information Science and Technology: JASIST, 59(7):1053–1072, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Roto:2006:PSU

[Rot06] Virpi Roto. Perspectives on search user interfaces: Best practices and future visions: Search on mobile phones. Journal of the American Society for Information Science and Technology: JASIST, 57(6):834–837, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rousseau:2006:BCT

[Rou06] Ronald Rousseau. Brief communication: Timelines in citation research. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1404–1405, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rousseau:2007:BCE

[Rou07] Ronald Rousseau. Brief communication: On Egghe’s construction of Lorenz curves. Journal of the American Society for Information Science and Technology: JASIST, 58(10): 1551–1552, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rousseau:2012:BCB

[Rou12] Ronald Rousseau. Brief communication: Basic properties of both percentile rank scores and the I3 indicator. Journal of the American Society for Information Science and Technology: JASIST, 63(2):416–420, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rowley:2009:BRB

[Row09] Jennifer Rowley. Book review: Digital Consumers: Reshaping the Information Professions. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2383–2385, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Park:2008:LPFa

[rP08a] Jung ran Park. Linguistic politeness and face-work in computer-mediated communication, Part 1: a theoretical framework. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2051–2059, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Park:2008:LPFb

[rP08b] Jung ran Park. Linguistic politeness and face-work in computer mediated communication, Part 2: an application of the theoretical framework. Journal of the American Society for Information Science and Technology: JASIST, 59 (14):2199–2209, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rooney:2008:ISC

[RPG^{+}08] Niall Rooney, David Patterson, Mykola Galushka, Vladimir Dobrynin, and Elena Smirnova. An investigation into the stability of contextual document clustering. Journal of the American Society for Information Science and Technology: JASIST, 59(2):256–266, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rotman:2012:SCC

[RPH^{+}12] Dana Rotman, Kezia Procita, Derek Hansen, Cynthia Sims Parr, and Jennifer Preece. Supporting content curation communities: the case of the Encyclopedia of Life. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1092–1107, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rafols:2010:SOM

[RPL10] Ismael Rafols, Alan L. Porter, and Loet Leydesdorff. Science overlay maps: a new tool for research policy and library management. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1871–1887, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Park:2009:CPD

[rPLM09] Jung ran Park, Caimei Lu, and Linda Marion. Cataloging professionals in the digital environment: a content analysis of job descriptions. Journal of the American Society for Information Science and Technology: JASIST, 60(4):844– 857, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rantanen:2006:FDT

[RPWM06] Esa M. Rantanen, Brent O. Palmer, Douglas A. Wiegmann, and Kevin M. Musiorski. Five-dimensional taxonomy to relate human errors and technological interventions in a human factors literature database. Journal of the American Society for Information Science and Technology: JASIST, 57(9):1221–1232, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rishel:2007:DCT

[RPYZ07] Tom Rishel, Louise A. Perkins, Sumanth Yenduri, and Farnaz Zand. Determining the context of text using augmented latent semantic indexing. Journal of the American Society for Information Science and Technology: JASIST, 58(14): 2197–2204, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Raban:2006:ESN

[RR06] Daphne R. Raban and Sheizaf Rafaeli. The effect of source nature and status on the subjective value of information. Journal of the American Society for Information Science and Technology: JASIST, 57(3):321–329, February 1, 2006.

CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rousseau:2012:IBJ

[RR12] Sandra Rousseau and Ronald Rousseau. Interactions between journal attributes and authors’ willingness to wait for editorial decisions. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1213– 1225, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rosemblat:2013:ESP

[RRA^{+}13] Graciela Rosemblat, Melissa P. Resnick, Ione Auston, Dongwook Shin, Charles Sneiderman, Marcelo Fiszman, and Thomas C. Rindflesch. Extending SemRep to the public health domain. Journal of the American Society for Information Science and Technology: JASIST, 64(10):1963–1974, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Raban:2011:ARP

[RRG11] Daphne Ruth Raban, Inbal Ronen, and Ido Guy. Acting or reacting? Preferential attachment in a people-tagging system. Journal of the American Society for Information Science and Technology: JASIST, 62(4):738–747, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Raghu:2005:AHS

[RRW05] T. S. Raghu, R. Ramesh, and Andrew B. Whinston. Addressing the homeland security problem: a collaborative decision-making framework. Journal of the American Society for Information Science and Technology: JASIST, 56(3): 310–324, February 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Reed:2008:PGI

[RS08] G. M. Reed and J. W. Sanders. Perspectives on global information ethics: The principle of distribution. Journal of the American Society for Information Science and Technology: JASIST, 59(7):1134–1142, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rosenbaum:2010:SAO

[RS10] Howard Rosenbaum and Pnina Shachaf. A structuration approach to online communities of practice: the case of Q&A communities. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1933–1944, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rushdi-Saleh:2011:OOC

[RSMVULPO11] Mohammed Rushdi-Saleh, M. Teresa Martín-Valdivia, L. Alfonso Ureña-López, and José M. Perea-Ortega. OCA: Opinion corpus for Arabic. Journal of the American Society for Information Science and Technology: JASIST, 62(10):2045– 2054, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Robertson:2008:MKS

[RT08] Stephen Robertson and John Tait. In memoriam: Karen Spärck Jones. Journal of the American Society for Information Science and Technology: JASIST, 59(5):852–854, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Reitsma:2004:SWS

[RTM04] René F. Reitsma, Lehana Thabane, and J. Michael B. MacLeod. Spatialization of Web sites using a weighted frequency model of navigation data. Journal of the American Society for Information Science and Technology: JASIST, 55(1):13–22, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rowley:2007:USI

[RU07] Jennifer Rowley and Christine Urquhart. Understanding student information behavior in relation to electronic information services: Lessons from longitudinal monitoring and evaluation, Part 1. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1162– 1174, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ruch:2010:BRB

[Ruc10] Patrick Ruch. Book review: Literature-based Discovery. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1506–1508, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Resnick:2006:PSU

[RV06] Marc L. Resnick and Misha W. Vaughan. Perspectives on search user interfaces: Best practices and future visions: Best practices and future visions for search user interfaces. Journal of the American Society for Information Science and Technology: JASIST, 57(6):781–787, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rapp:2005:BRN

[RW05] Barbara A. Rapp and David L. Wheeler. Bioinformatics resources from the National Center for Biotechnology Information: an integrated foundation for discovery. Journal of the American Society for Information Science and Technology: JASIST, 56(5):538–550, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rousseau:2008:BCP

[RY08] Ronald Rousseau and Fred Y. Ye. Brief communications: A proposal for a dynamic h-type index. Journal of the American Society for Information Science and Technology: JA-SIST, 59(11):1853–1855, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See comments [Bur09b].

Ryan:2009:BRB

[Rya09] Michael Ryan. Book review: Reading Matters: Five Centuries of Discovering Books. Journal of the American Society for Information Science and Technology: JASIST, 60(8): 1727–1728, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Rousseau:2004:CAC

[RZ04] Ronald Rousseau and Alesia Zuccala. A classification of author co-citations: Definitions and search strategies. Journal of the American Society for Information Science and Technology: JASIST, 55(6):513–529, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sacco:2006:PSU

[Sac06] Giovanni M. Sacco. Perspectives on search user interfaces: Best practices and future visions: Dynamic taxonomies and guided searches. Journal of the American Society for Information Science and Technology: JASIST, 57(6):792–796, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Saggion:2006:BRB

[Sag06] Horacio Saggion. Book review: Indexing and abstracting in theory and practice, third edition. Journal of the American Society for Information Science and Technology: JASIST, 57(1):144–146, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sandusky:2002:BRB

[San02] Robert J. Sandusky. Book review: Accessing and browsing information and communication. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1185–1187, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sanderson:2008:BCR

[San08] Mark Sanderson. Brief communication: Revisiting h measured on UK LIS and IR academics. Journal of the American Society for Information Science and Technology: JA-SIST, 59(7):1184–1190, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sandstrom:2010:BRB

[San10] Pamela Effrein Sandstrom. Book review: Information Foraging Theory: Adaptive Interaction with Information. Journal of the American Society for Information Science and Technology: JASIST, 61(10):2161–2164, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Saracevic:2007:RRLa

[Sar07a] Tefko Saracevic. Relevance: a review of the literature and a framework for thinking on the notion in information science.

Part II: nature and manifestations of relevance. Journal of the American Society for Information Science and Technology: JASIST, 58(13):1915–1933, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Saracevic:2007:RRLb

[Sar07b] Tefko Saracevic. Relevance: a review of the literature and a framework for thinking on the notion in information science. Part III: Behavior and effects of relevance. Journal of the American Society for Information Science and Technology: JASIST, 58(13):2126–2144, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Satyanesan:2008:BRB

[Sat08] Jessie Satyanesan. Book review: Libraries and Google. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1531–1533, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sauperl:2004:CCG

[Šau04] Alenka Šauperl. Catalogers’ common ground and shared knowledge. Journal of the American Society for Information Science and Technology: JASIST, 55(1):55–63, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sauro:2009:PDI

[Sau09] Clare Sauro. Perspectives on design: Information technologies and creative practices: Digitized historic costume collections: Inspiring the future while preserving the past. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1939–1941, September 2009. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Savolainen:2003:BR

[Sav03] Reijo Savolainen. Book reviews. Journal of the American Society for Information Science and Technology: JASIST, 54(7):695–697, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Savolainen:2006:IUG

[Sav06] Reijo Savolainen. Information use as gap-bridging: the viewpoint of sense-making methodology. Journal of the American Society for Information Science and Technology: JA-SIST, 57(8):1116–1125, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Savolainen:2007:ISH

[Sav07] Reijo Savolainen. Information source horizons and source preferences of environmental activists: a social phenomenological approach. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1709–1719, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Savolainen:2009:IIC

[Sav09] Reijo Savolainen. Interpreting informational cues: an explorative study on information use among prospective homebuyers. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2244–2254, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Savolainen:2011:JQC

[Sav11] Reijo Savolainen. Judging the quality and credibility of information in Internet discussion forums. Journal of the American Society for Information Science and Technology: JA-SIST, 62(7):1243–1256, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Savolainen:2012:SAP

[Sav12] Reijo Savolainen. The structure of argument patterns on a social Q&A site. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2536–2548, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sawyer:2010:BRB

[Saw10] Steve Sawyer. Book review: Wired for Innovation: How Information Technology is Reshaping the Economy. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2375–2376, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Song:2003:TCS

[SB03] D. Song and P. D. Bruza. Towards context sensitive information inference. Journal of the American Society for Information Science and Technology: JASIST, 54(4):321– 334, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Salembier:2007:PMM

[SB07a] Philippe Salembier and Ana B. Benitez. Perspectives on MPEG-7: Metadata for multimedia: Structure description tools. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1329–1337, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schneider:2007:MCPa

[SB07b] Jesper W. Schneider and Pia Borlund. Matrix comparison, Part 1: Motivation and important issues for measuring the resemblance between proximity measures or ordination results. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1586–1595, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schneider:2007:MCPb

[SB07c] Jesper W. Schneider and Pia Borlund. Matrix comparison, Part 2: Measuring the resemblance between proximity measures or ordination results by use of the Mantel and Procrustes statistics. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1596–1609, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Solomon:2012:PFO

[SB12a] David J. Solomon and Bo-Christer Björk. Publication fees in open access publishing: Sources of funding and factors influencing choice of journal. Journal of the American Society for Information Science and Technology: JASIST, 63(1): 98–107, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Ano13d].

Solomon:2012:SOA

[SB12b] David J. Solomon and Bo-Christer Björk. A study of open access journals using article processing charges. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1485–1495, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Smith:2012:EIJ

[SBA12] Gavin Smith, Chris Brien, and Helen Ashman. Evaluating implicit judgments from image search clickthrough data. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2451–2462, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Srinivasan:2009:BDM

[SBBF09] Ramesh Srinivasan, Robin Boast, Katherine M. Becvar, and Jonathan Furner. Blobgects: Digital museum catalogs and diverse user communities. Journal of the American Society for Information Science and Technology: JASIST, 60 (4):666–678, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schaalje:2013:OSS

[SBF13] G. Bruce Schaalje, Natalie J. Blades, and Tomohiko Funai. An open-set size-adjusted Bayesian classifier for authorship attribution. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1815–1825, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sharma:2011:ESF

[SBIS11] Nikhil Sharma, Brian S. Butler, Jeannie Irwin, and Heiko Spallek. Emphasizing social features in information portals: Effects on new member engagement. Journal of the American Society for Information Science and Technology: JA-SIST, 62(11):2106–2120, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sun:2011:TBS

[SBNB11] Aixin Sun, Sourav S. Bhowmick, Khanh Tran Nam Nguyen, and Ge Bai. Tag-based social image retrieval: an empirical evaluation. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2364–2381, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Scalise:2007:ATL

[SBT^{+}07] Kathleen Scalise, Diana J. Bernbaum, Mike Timms, S. Veeragoudar Harrell, Kristen Burmester, Cathleen A. Kennedy, and Mark Wilson. Adaptive technology for e-learning: principles and case studies of an emerging field. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2295–2309, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spink:2004:Ia

[SC04a] Amanda Spink and Charles Cole. Introduction. Journal of the American Society for Information Science and Technology: JASIST, 55(8):657–659, June 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spink:2004:Ib

[SC04b] Amanda Spink and Charles Cole. Introduction. Journal of the American Society for Information Science and Technology: JASIST, 55(9):767–768, July 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spink:2006:HIB

[SC06] Amanda Spink and Charles Cole. Human information behavior: Integrating diverse approaches and information use. Journal of the American Society for Information Science and Technology: JASIST, 57(1):25–35, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schumaker:2008:ENA

[SC08] Robert P. Schumaker and Hsinchun Chen. Evaluating a news-aware quantitative trader: the effect of momentum and contrarian stock selection strategies. Journal of the American Society for Information Science and Technology: JA-SIST, 59(2):247–255, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Su:2011:MNA

[SC11] Chunke Su and Noshir Contractor. A multidimensional network approach to studying team members’ information seeking from human and digital knowledge sources in consulting firms. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1257–1275, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sugimoto:2012:BPE

[SC12] Cassidy R. Sugimoto and Blaise Cronin. Biobibliometric profiling: an examination of multifaceted approaches to scholarship. Journal of the American Society for Information Science and Technology: JASIST, 63(3):450–468, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sood:2012:AIP

[SCA12] Sara Owsley Sood, Elizabeth F. Churchill, and Judd Antin. Automatic identification of personal insults on social news sites. Journal of the American Society for Information Science and Technology: JASIST, 63(2):270–285, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Souza:2012:USL

[SCC^{+}12] Jucimar Souza, André Carvalho, Marco Cristo, Edleno Moura, Pavel Calado, Paul-Alexandru Chirita, and Wolfgang Nejdl. Using site-level connections to estimate link confidence. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2294–2312, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schroeder:2001:BRB

[Sch01] Alan T. Schroeder, Jr. Book review: The information resources policy handbook: Research for the information age. Benjamin M. Compaine and William H. Read, editors. Journal of the American Society for Information Science and Technology: JASIST, 52(3):278–279, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schroeder:2002:BRB

[Sch02] Alan T. Schroeder, Jr. Book review: Data privacy in the information age. Journal of the American Society for Information Science and Technology: JASIST, 53(3):251–253, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schamber:2003:BRB

[Sch03] Linda Schamber. Book review: Research Methods for Students, Academics and Professionals: Information Management and Systems. Journal of the American Society for Information Science and Technology: JASIST, 54(10):982– 983, August 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schneider:2004:BRB

[Sch04a] Jesper W. Schneider. Book review: Mapping scientific frontiers: the quest for knowledge visualization. Journal of the American Society for Information Science and Technology: JASIST, 55(4):363–365, February 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Scholl:2004:BR

[Sch04b] Jochen Scholl. Book review. Journal of the American Society for Information Science and Technology: JASIST, 55(2): 181–182, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schwartz:2005:BRB

[Sch05] Candy Schwartz. Book review: Annual review of information science and technology, volume 38, 2004. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1464–1465, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schreiber:2008:EII

[Sch08] Michael Schreiber. An empirical investigation of the g-index for 26 physicists in comparison with the h-index, the Aindex, and the R-index. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1513– 1522, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schreiber:2009:CSM

[Sch09a] Michael Schreiber. A case study of the modified Hirsch index h_{m} accounting for multiple coauthors. Journal of the American Society for Information Science and Technology: JA-SIST, 60(6):1274–1282, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schreiber:2009:FCP

[Sch09b] Michael Schreiber. Fractionalized counting of publications for the g-index. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2145–2150, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schreiber:2010:RIA

[Sch10] Michael Schreiber. Revisiting the g-index: the average number of citations in the g-core. Journal of the American Society for Information Science and Technology: JASIST, 61(1): 169–174, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schreiber:2012:IRP

[Sch12] Michael Schreiber. Inconsistencies of recently proposed citation impact indicators and how to avoid them. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2062–2073, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schreiber:2013:BCD

[Sch13a] Michael Schreiber. Brief communication: Do we need the gindex? Journal of the American Society for Information Science and Technology: JASIST, 64(11):2396–2399, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schreiber:2013:BCE

[Sch13b] Michael Schreiber. Brief communication: Empirical evidence for the relevance of fractional scoring in the calculation of percentile rank scores. Journal of the American Society for Information Science and Technology: JASIST, 64(4):861– 867, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schreiber:2013:BCI

[Sch13c] Michael Schreiber. Brief communication: Inconsistencies in the highly cited publications indicator. Journal of the American Society for Information Science and Technology: JA-SIST, 64(6):1298–1302, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schreiber:2013:BCU

[Sch13d] Michael Schreiber. Brief communication: Uncertainties and ambiguities in percentiles and how to avoid them. Journal of the American Society for Information Science and Technology: JASIST, 64(3):640–643, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Samoylenko:2006:VSW

[SCLC06] I. Samoylenko, T.-C. Chao, W.-C. Liu, and C.-M. Chen. Visualizing the scientific world and its evolution. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1461–1469, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sha:2013:KPH

[SCZZ13] Xiqing Sha, Klarissa Ting-Ting Chang, Cheng Zhang, and Chenghong Zhang. Knowledge popularity in a heterogeneous network: Exploiting the contextual effects of document popularity in knowledge management systems. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1842–1851, September 2013. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Saeed:2001:BCN

[SD01] Khalid Saeed and Agnieszka Dardzińska. Brief communication: Natural language processing: Word recognition without segmentation. Journal of the American Society for Information Science and Technology: JASIST, 52(14):1275– 1279, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shenton:2003:CYU

[SD03] Andrew K. Shenton and Pat Dixon. A comparison of youngsters’ use of CD-ROM and the Internet as information resources. Journal of the American Society for Information Science and Technology: JASIST, 54(11):1029–1049, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Serenko:2012:MSL

[SDJB12] Alexander Serenko, Brian Detlor, Heidi Julien, and Lorne D. Booker. A model of student learning outcomes of information literacy instruction in a business school. Journal of the American Society for Information Science and Technology: JASIST, 63(4):671–686, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shepherd:2001:RUP

[SDWG01] Michael Shepherd, John F. (Jack) Duffy, Carolyn Watters, and Nitin Gugle. The role of user profiles for news filtering. Journal of the American Society for Information Science and Technology: JASIST, 52(2):149–160, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Scholl:2011:TFA

[SEDC11] Hans J. Scholl, Michael B. Eisenberg, Lee Dirks, and Timothy S. Carlson. The TEDS framework for assessing information systems from a human actors’ perspective: Extending and repurposing Taylor’s Value-Added Model. Journal of the American Society for Information Science and Technology: JASIST, 62(4):789–804, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sellitto:2005:IIW

[Sel05] Carmine Sellitto. The impact of impermanent Web-located citations: a study of 123 scholarly conference publications. Journal of the American Society for Information Science and Technology: JASIST, 56(7):695–703, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shapira:2005:PPM

[SEMK05] Bracha Shapira, Yuval Elovici, Adlay Meshiach, and Tsvi Kuflik. PRAW — A PRivAcy model for the Web. Journal of the American Society for Information Science and Technology: JASIST, 56(2):159–172, January 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schwartz:2007:CDA

[SF07] Frank Schwartz and Y. C. Fang. Citation data analysis on hydrogeology. Journal of the American Society for Information Science and Technology: JASIST, 58(4):518–525, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shachak:2008:ETB

[SF08] Aviv Shachak and Sara Fine. The Effect of training on biologists acceptance of bioinformatics tools: a field experiment. Journal of the American Society for Information Science and Technology: JASIST, 59(5):719–730, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spink:2001:RRL

[SG01] Amanda Spink and Howard Greisdorf. Research: Regions and levels: Measuring and mapping users’ relevance judgments. Journal of the American Society for Information Science and Technology: JASIST, 52(2):161–173, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schloegl:2011:GUV

[SG11] Christian Schloegl and Juan Gorraiz. Global usage versus global citation metrics: the case of pharmacology journals. Journal of the American Society for Information Science and Technology: JASIST, 62(1):161–170, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Serpa:2013:SCA

[SGJ13] Francisco G. Serpa, Adam M. Graves, and Artjay Javier. Statistical common author networks. Journal of the American Society for Information Science and Technology: JA-SIST, 64(12):2507–2512, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stvilia:2007:FIQ

[SGTS07] Besiki Stvilia, Les Gasser, Michael B. Twidale, and Linda C. Smith. A framework for information quality assessment. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1720–1733, October 2007.

CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Strickland:2005:TSI

[SH05] Lee S. Strickland and Laura E. Hunt. Technology, security, and individual privacy: New tools, new threats, and new public perceptions. Journal of the American Society for Information Science and Technology: JASIST, 56(3):221–234, February 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sawyer:2007:CIT

[SH07a] Steve Sawyer and Haiyan Huang. Conceptualizing information, technology, and people: Comparing information science and information systems literatures. Journal of the American Society for Information Science and Technology: JA-SIST, 58(10):1436–1447, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sotudeh:2007:CPO

[SH07b] Hajar Sotudeh and Abbas Horri. The citation performance of open access journals: a disciplinary investigation of citation distribution models. Journal of the American Society for Information Science and Technology: JASIST, 58(13): 2145–2156, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sotudeh:2007:TOA

[SH07c] Hajar Sotudeh and Abbas Horri. Tracking open access journals evolution: Some considerations in open access data collection validation. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1578–1585, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shachak:2006:DPU

[Sha06a] Aviv Shachak. Diffusion pattern of the use of genomic databases and analysis of biological sequences from 1970– 2003: Bibliographic record analysis of 12 journals. Journal of the American Society for Information Science and Technology: JASIST, 57(1):44–50, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shantz:2006:BRB

[Sha06b] Jeff Shantz. Book review: Code: Collaborative ownership and the digital economy. Journal of the American Society for Information Science and Technology: JASIST, 57(13): 1854–1855, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shankar:2007:OCP

[Sha07] Kalpana Shankar. Order from chaos: the poetics and pragmatics of scientific recordkeeping. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1457–1466, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shankar:2008:BRB

[Sha08] Kalpana Shankar. Book review: Research methods in information. Journal of the American Society for Information Science and Technology: JASIST, 59(1):158–159, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shah:2013:EAC

[Sha13a] Chirag Shah. Effects of awareness on coordination in collaborative information seeking. Journal of the American Society for Information Science and Technology: JASIST, 64(6): 1122–1143, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shaw:2013:AIS

[Sha13b] Ryan Shaw. Advances in information science: Information organization and the philosophy of history. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1092–1103, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shepherd:2010:BRB

[She10] Elizabeth Shepherd. Book review: Electronic Records in the Manuscript Repository. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1512– 1513, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shin:2008:NGI

[Shi08] Dong-Hee Shin. Next generation of information infrastructure: a comparative case study of Korea versus the United States of America. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1785– 1800, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shilton:2012:AIS

[Shi12] Katie Shilton. Advances in information science: Participatory personal data: an emerging research challenge for the information sciences. Journal of the American Society for Information Science and Technology: JASIST, 63(10): 1905–1915, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Salovaara:2011:EAI

[SHO11] Antti Salovaara, Sacha Helfenstein, and Antti Oulasvirta. Everyday appropriations of information technology: a study of creative uses of digital cameras. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2347–2363, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Snyder:2009:PDI

[SHS09] Jaime Snyder, Robert Heckman, and Michael J. Scialdone. Perspectives on design: Information technologies and creative practices: Information studios: Integrating arts-based learning into the education of information professionals. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1923–1932, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stvilia:2011:CST

[SHS^{+}11] Besiki Stvilia, Charles C. Hinnant, Katy Schindler, Adam Worrall, Gary Burnett, Kathleen Burnett, Michelle M. Kazmer, and Paul F. Marty. Composition of scientific teams and publication productivity at a national science lab. Journal of the American Society for Information Science and Technology: JASIST, 62(2):270–283, February 2011. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Simon:2007:BRB

[Sim07] Scott J. Simon. Book review: Computer models of musical creativity. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1553–1555, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sin:2011:ICC

[Sin11] Sei-Ching Joanna Sin. International coauthorship and citation impact: a bibliometric study of six LIS journals, 1980– 2008. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1770–1783, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sinn:2012:IDA

[Sin12] Donghee Sinn. Impact of digital archival collections on historical research. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1521–1537, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stvilia:2010:MAQ

[SJ10] Besiki Stvilia and Corinne Jörgensen. Member activities and quality of tags in a collection of historical photographs in Flickr. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2477–2489, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Soulier:2013:RRE

[SJTB13] Laure Soulier, Lamjed Ben Jabeur, Lynda Tamine, and Wahiba Bahsoun. On ranking relevant entities in heterogeneous networks using a language-based model. Journal of the American Society for Information Science and Technology: JASIST, 64(3):500–515, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Szava-Kovats:2001:RIC

[SK01] Endre Száva-Kováts. Research: Indirect-collective referencing (ICR) in the elite journal literature of physics. I. A literature science study on the journal level. Journal of the American Society for Information Science and Technology: JASIST, 52(3):201–211, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Szava-Kovats:2002:ICR

[SK02a] Endre Száva-Kováts. Indirect-collective referencing (ICR) in the elite journal literature of physics. II. A literature science study on the level of communications. Journal of the American Society for Information Science and Technology: JASIST, 53(1):47–56, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Szava-Kovats:2002:UAH

[SK02b] Endre Száva-Kováts. Unfounded attribution of the “halflife” index-number of literature obsolescence to Burton and Kebler: a literature science study. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1098–1105, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sun:2006:CEN

[SK06] Ying Sun and Paul B. Kantor. Cross-evaluation: a new model for information system evaluation. Journal of the American Society for Information Science and Technology: JASIST, 57(5):614–628, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shah:2012:SQV

[SK12] Chirag Shah and Vanessa Kitzie. Social Q&A and virtual reference — comparing apples and oranges with the help of experts and users. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2020–2036, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Swar:2013:AIT

[SK13] Bobby Swar and Gohar Feroz Khan. An analysis of the information technology outsourcing domain: a social network and Triple helix approach. Journal of the American Society for Information Science and Technology: JASIST, 64 (11):2366–2378, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sauperl:2008:CAL

[ŠKL08] Alenka Šauperl, Janko Klasinc, and Simona Lužar. Components of abstracts: Logical structure of scholarly abstracts in pharmacology, sociology, and linguistics and literature. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1420–1432, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shapira:2001:EEM

[SKM01] Bracha Shapira, Paul B. Kantor, and Benjamin Melamed. The effect of extrinsic motivation on user behavior in a collaborative information finding system. Journal of the American Society for Information Science and Technology: JA-SIST, 52(11):879–887, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shibata:2007:TAC

[SKM07] Naoki Shibata, Yuya Kajikawa, and Katsumori Matsushima. Topological analysis of citation networks to discover the future core articles. Journal of the American Society for Information Science and Technology: JASIST, 58(6):872–882, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sun:2011:UCE

[SKM11] Ying Sun, Paul B. Kantor, and Emile L. Morse. Using crossevaluation to evaluate interactive QA systems. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1653–1665, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shibata:2011:MRB

[SKS11] Naoki Shibata, Yuya Kajikawa, and Ichiro Sakata. Measuring relatedness between communities in a citation network. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1360–1369, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shibata:2012:LPC

[SKS12] Naoki Shibata, Yuya Kajikawa, and Ichiro Sakata. Link prediction in citation networks. Journal of the American Society for Information Science and Technology: JASIST, 63 (1):78–85, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shibata:2009:CSM

[SKTM09] Naoki Shibata, Yuya Kajikawa, Yoshiyuki Takeda, and Katsumori Matsushima. Comparative study on methods of detecting research fronts using different types of citation. Journal of the American Society for Information Science and Technology: JASIST, 60(3):571–580, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sundar:2007:NCI

[SKWH07] S. Shyam Sundar, Silvia Knobloch-Westerwick, and Matthias R. Hastall. News cues: Information scent and cognitive heuristics. Journal of the American Society for Information Science and Technology: JASIST, 58(3):366–378, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sun:2006:WUB

[SL06] Aixin Sun and Ee-Peng Lim. Web unit-based mining of homepage relationships. Journal of the American Society for Information Science and Technology: JASIST, 57(3):394– 407, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shaw:2010:EAT

[SL10] Debora Shaw and Sean M. Lind. Editorial: ASIS&T Thesaurus descriptors chosen by JASIST submitting authors and peer reviewers. Journal of the American Society for Information Science and Technology: JASIST, 61(10):1949– 1950, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shelton:2012:PPB

[SL12] R. D. Shelton and Loet Leydesdorff. Publish or patent: Bibliometric evidence for empirical trade-offs in national funding strategies. Journal of the American Society for Information Science and Technology: JASIST, 63(3):498–511, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sun:2013:TCD

[SL13] Xiaoling Sun and Hongfei Lin. Topical community detection from mining user tagging behavior and interest. Journal of the American Society for Information Science and Technology: JASIST, 64(2):321–333, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sun:2003:PMF

[SLN03] Aixin Sun, Ee-Peng Lim, and Wee-Keong Ng. Performance measurement framework for hierarchical text classification. Journal of the American Society for Information Science and Technology: JASIST, 54(11):1014–1028, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Slone:2002:IMM

[Slo02] Debra J. Slone. The influence of mental models and goals on search patterns during Web interaction. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1152–1169, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Slone:2007:ITC

[Slo07] Debra J. Slone. The impact of time constraints on Internet and Web use. Journal of the American Society for Information Science and Technology: JASIST, 58(4):508– 517, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sugimoto:2011:SSD

[SLR^{+}11] Cassidy R. Sugimoto, Daifeng Li, Terrell G. Russell, S. Craig Finlay, and Ying Ding. The shifting sands of disciplinary development: Analyzing North American Library and Information Science dissertations using latent Dirichlet allocation. Journal of the American Society for Information Science and Technology: JASIST, 62(1):185–204, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shen:2006:ECU

[SLS06] Xiangxing Shen, Dongmin Li, and Chong Shen. Evaluating China’s university library Web sites using correspondence analysis. Journal of the American Society for Information Science and Technology: JASIST, 57(4):493–500, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Su:2009:PPN

[SLSK09] Fang Pei Su, Kuei Kuei Lai, R. R. K. Sharma, and Tsung Hsien Kuo. Patent priority network: Linking patent portfolio to strategic goals. Journal of the American Society for Information Science and Technology: JASIST, 60 (11):2353–2361, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schlieder:2002:QRX

[SM02] Torsten Schlieder and Holger Meuss. Querying and ranking XML documents. Journal of the American Society for Information Science and Technology: JASIST, 53(6):489–503, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stefl-Mabry:2003:SJA

[SM03] Joette Stefl-Mabry. A social judgment analysis of information source preference profiles: an exploratory study to empirically represent media selection patterns. Journal of the American Society for Information Science and Technology: JASIST, 54(9):879–904, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sombatsompop:2005:MEI

[SM05a] Narongrit Sombatsompop and Teerasak Markpin. Making an equality of ISI impact factors for different subject fields. Journal of the American Society for Information Science and Technology: JASIST, 56(7):676–683, May 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stefl-Mabry:2005:RMP

[SM05b] Joette Stefl-Mabry. The reality of media preferences: Do professional groups vary in awareness? Journal of the American Society for Information Science and Technology: JA-SIST, 56(13):1419–1426, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shah:2010:ACI

[SM10] Chirag Shah and Gary Marchionini. Awareness in collaborative information seeking. Journal of the American Society for Information Science and Technology: JASIST, 61(10): 1970–1986, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Ano10].

Schubert:2013:PAL

[SM13] Torben Schubert and Carolin Michels. Placing articles in the large publisher nations: Is there a “free lunch” in terms of higher impact? Journal of the American Society for Information Science and Technology: JASIST, 64(3):596–611, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Small:2003:PCM

[Sma03] Henry Small. Paradigms, citations, and maps of science: a personal history. Journal of the American Society for Information Science and Technology: JASIST, 54(5):394–399, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Smalheiser:2005:LEE

[Sma05] Neil R. Smalheiser. Letter to the Editor: Erratum to publisher. Journal of the American Society for Information Science and Technology: JASIST, 56(9):1010, July 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Smalheiser:2012:AIS

[Sma12] Neil R. Smalheiser. Advances in information science: Literature-based discovery: Beyond the ABCs. Journal of the American Society for Information Science and Technology: JASIST, 63(2):218–224, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Salager-Meyer:2009:BSS

[SMAB09] Françoise Salager-Meyer, María Ángeles Alcaraz Ariza, and Maryelis Pabón Berbesí. “Backstage solidarity” in Spanishand English-written medical research papers: Publication context and the acknowledgment paratext. Journal of the American Society for Information Science and Technology:

JASIST, 60(2):307–317, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Ano14].

Smith:2001:QAI

[Smi01] John R. Smith. Quantitative assessment of image retrieval effectiveness. Journal of the American Society for Information Science and Technology: JASIST, 52(11):969–979, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Smith:2004:BRB

[Smi04] Derek G. Smith. Book review: A history of online information services, 1963–1976. Journal of the American Society for Information Science and Technology: JASIST, 55 (7):651–652, May 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Steele:2012:AIS

[SML12] Robert Steele, Kyongho Min, and Amanda Lo. Advances in information science: Personal health record architectures: Technology infrastructure implications and dependencies. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1079–1091, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stvilia:2009:MOC

[SMY09] Besiki Stvilia, Lorri Mon, and Yong Jeong Yi. A model for online consumer health information quality. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1781–1791, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sugimoto:2011:AGI

[SNRB11] Cassidy R. Sugimoto, Chaoqun Ni, Terrell G. Russell, and Brenna Bychowski. Academic genealogy as an indicator of interdisciplinarity: an examination of dissertation networks in Library and Information Science. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1808–1828, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Soares:2006:BRB

[Soa06] Andrey Soares. Book review: Legitimate applications of peer-to-peer networks. Journal of the American Society for Information Science and Technology: JASIST, 57(2):293– 294, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shachaf:2008:SEV

[SOH08] Pnina Shachaf, Shannon M. Oltmann, and Sarah M. Horowitz. Service equality in virtual reference. Journal of the American Society for Information Science and Technology: JASIST, 59(4):535–550, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sonnenwald:2010:BRB

[Son10] Diane H. Sonnenwald. Book review: E-Collaboration in Modern Organizations: Initiating and Managing Distributed Projects. Journal of the American Society for Information Science and Technology: JASIST, 61(3):633–634, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spink:2002:MIS

[SOO02] Amanda Spink, H. Cenk Ozmutlu, and Seda Ozmutlu. Multitasking information seeking and searching processes. Journal of the American Society for Information Science and Technology: JASIST, 53(8):639–652, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Srinivasan:2007:DIE

[SP07] Ramesh Srinivasan and Ajit Pyati. Diasporic information environments: Reframing immigrant-focused information research. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1734–1744, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stuckman:2011:AWM

[SP11] Jeffrey Stuckman and James Purtilo. Analyzing the wikisphere: Methodology and data to support quantitative wiki research. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1564–1576, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spern:2001:BRB

[Spe01] Karen Spern. Book review: Organizing audiovisual and electronic resources for access: a cataloging guide, by Ingrid Hsieh-Yee. Journal of the American Society for Information Science and Technology: JASIST, 52(6):512–513, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sugimoto:2008:UFC

[SPH08] Cassidy R. Sugimoto, Jean A. Pratt, and Karina Hauser. Using field cocitation analysis to assess reciprocal and shared impact of LIS/ MIS fields. Journal of the American Society for Information Science and Technology: JASIST, 59 (9):1441–1453, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spink:2002:EIS

[Spi02] Amanda Spink. Editorial: Introduction to the special issue on Web research. Journal of the American Society for Information Science and Technology: JASIST, 53(2):65–66, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spoerri:2008:ARE

[Spo08] Anselm Spoerri. Authority and ranking effects in data fusion. Journal of the American Society for Information Science and Technology: JASIST, 59(3):450–460, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Srinivasan:2009:CBS

[SPR09] Ramesh Srinivasan, Alberto Pepe, and Marko A. Rodriguez. A clustering-based semi-automated technique to build cultural ontologies. Journal of the American Society for Information Science and Technology: JASIST, 60(3):608–620, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shiri:2006:QEB

[SR06] Ali Shiri and Crawford Revie. Query expansion behavior within a thesaurus-enhanced search environment: a usercentered evaluation. Journal of the American Society for Information Science and Technology: JASIST, 57(4):462– 478, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Simkin:2007:MTC

[SR07] Mikhail V. Simkin and Vwani P. Roychowdhury. A mathematical theory of citing. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1661– 1673, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shaalan:2009:NNE

[SR09] Khaled Shaalan and Hafsa Raza. NERA: Named Entity Recognition for Arabic. Journal of the American Society for Information Science and Technology: JASIST, 60(8): 1652–1663, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Saab:2011:IO

[SR11] David J. Saab and Uwe V. Riss. Information as ontologization. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2236–2246, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shah:2006:CBA

[SRDZ06] Biren Shah, Vijay Raghavan, Praveen Dhatric, and Xiaoquan Zhao. A cluster-based approach for efficient contentbased image retrieval using a similarity-preserving space transformation method. Journal of the American Society for Information Science and Technology: JASIST, 57(12): 1694–1707, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sreenivasulu:2002:LEE

[Sre02] V. Sreenivasulu. Letter to the Editor: Engineering a search engine (weblib) and browser (knowledge navigator) for digital libraries: global knowledge discovery tools exclusively for librarians and libraries on the web. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1273–1275, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Srinivasan:2004:TMG

[Sri04] Padmini Srinivasan. Text mining: Generating hypotheses from MEDLINE. Journal of the American Society for Information Science and Technology: JASIST, 55(5):396–413, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Srinivasan:2007:EAI

[Sri07] Ramesh Srinivasan. Ethnomethodological architectures: Information systems driven by cultural and community visions. Journal of the American Society for Information Science and Technology: JASIST, 58(5):723–733, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shani:2013:ICD

[SRS^{+}13] Guy Shani, Lior Rokach, Bracha Shapira, Sarit Hadash, and Moran Tangi. Investigating confidence displays for top- N recommendations. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2548– 2563, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stieger:2007:FRO

[SRV07] Stefan Stieger, Ulf-Dietrich Reips, and Martin Voracek. Forced-response in online surveys: Bias from reactance and an increase in sex-specific dropout. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1653–1660, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schloegl:2004:IRL

[SS04] Christian Schloegl and Wolfgang G. Stock. Impact and relevance of LIS journals: a scientometric analysis of international and German-language LIS journals — Citation analysis versus reader survey. Journal of the American Society for Information Science and Technology: JASIST, 55(13):1155– 1168, November 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stock:2006:IPI

[SS06] Mechtild Stock and Wolfgang G. Stock. Intellectual property information: a comparative analysis of main information providers. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1794–1803, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schmidt:2009:CIE

[SS09] Stefanie Schmidt and Wolfgang G. Stock. Collective indexing of emotions in images. A study in emotional information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 60(5):863–876, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Soderholm:2010:VFE

[SS10] Hanna M. Söderholm and Diane H. Sonnenwald. Visioning future emergency healthcare collaboration: Perspectives from large and small medical centers. Journal of the American Society for Information Science and Technology: JA-SIST, 61(9):1808–1823, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Syn:2013:FST

[SS13] Sue Yeon Syn and Michael B. Spring. Finding subject terms for classificatory metadata from user-generated social tags. Journal of the American Society for Information Science and Technology: JASIST, 64(5):964–980, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Swanson:2001:RID

[SSB01] Don R. Swanson, Neil R. Smalheiser, and A. Bookstein. Research: Information discovery from complementary literatures: Categorizing viruses as potential weapons. Journal of the American Society for Information Science and Technology: JASIST, 52(10):797–812, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Soderholm:2008:EPV

[SSM^{+}08a] Hanna M. Söderholm, Diane H. Sonnenwald, James E. Manning, Bruce Cairns, Greg Welch, and Henry Fuchs. Exploring the potential of video technologies for collaboration in emergency medical care: Part II. Task performance. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2335–2349, December 2008. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sonnenwald:2008:EPV

[SSM^{+}08b] Diane H. Sonnenwald, Hanna M. Söderholm, James E. Manning, Bruce Cairns, Greg Welch, and Henry Fuchs. Exploring the potential of video technologies for collaboration in emergency medical care: Part I. Information sharing. Journal of the American Society for Information Science and Technology: JASIST, 59(14):2320–2334, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stringer:2010:SVG

[SSPA10] Michael J. Stringer, Marta Sales-Pardo, and Luís A. Nunes Amaral. Statistical validation of a global model for the distribution of the ultimate number of citations accrued by papers published in a scientific journal. Journal of the American Society for Information Science and Technology: JASIST, 61 (7):1377–1385, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Segev:2013:WGF

[SSS13] Elad Segev, Tamir Sheafer, and Shaul R. Shenhav. Is the world getting flatter? A new method for examining structural trends in the news. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2537– 2547, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Swanson:2006:RIC

[SST06] Don R. Swanson, Neil R. Smalheiser, and Vetle I. Torvik. Ranking indirect connections in literature-based discovery: the role of medical subject headings. Journal of the American Society for Information Science and Technology: JA-SIST, 57(11):1427–1439, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shapira:2009:EPM

[SSTM09] Bracha Shapira, Peretz Shoval, Noam Tractinsky, and Joachim Meyer. ePaper: a personalized mobile newspaper. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2333–2346, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sandusky:2008:FUJ

[ST08] Robert J. Sandusky and Carol Tenopir. Finding and using journal-article components: Impacts of disaggregation on teaching and research practice. Journal of the American Society for Information Science and Technology: JASIST, 59(6):970–982, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shifman:2009:AGD

[ST09] Limor Shifman and Mike Thelwall. Assessing global diffusion with Web memetics: the spread and evolution of a popular joke. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2567–2576, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sugimoto:2013:SSB

[ST13] Cassidy R. Sugimoto and Mike Thelwall. Scholars on soap boxes: Science communication and dissemination in TED videos. Journal of the American Society for Information Science and Technology: JASIST, 64(4):663–674, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stamatatos:2009:SMA

[Sta09] Efstathios Stamatatos. A survey of modern authorship attribution methods. Journal of the American Society for Information Science and Technology: JASIST, 60(3):538–556, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stanton:2010:BRB

[Sta10] Jeffrey M. Stanton. Book review: Security of Information and Communication Networks. Journal of the American Society for Information Science and Technology: JASIST, 61(5):1074, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stamatatos:2011:PDU

[Sta11] Efstathios Stamatatos. Plagiarism detection using stopword n-grams. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2512–2527, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Steinerova:2003:ISR

[Ste03a] Jela Steinerová. Information science research agenda in Slovakia: History and emerging vision. Journal of the American Society for Information Science and Technology: JASIST, 54(1):81–86, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stern:2003:NKM

[Ste03b] David Stern. New knowledge management systems: the implications for data discovery, collection development, and the changing role of the librarian. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1138–1140, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stenmark:2008:ICU

[Ste08] Dick Stenmark. Identifying clusters of user behavior in intranet search engine log files. Journal of the American Society for Information Science and Technology: JASIST, 59 (14):2232–2243, December 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stirling:2001:BRBb

[Sti01a] Dale A. Stirling. Book review: Editorial peer review: Its strengths and weaknesses, by Ann C. Weller. Journal of the American Society for Information Science and Technology: JASIST, 52(11):984–985, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stirling:2001:BRBa

[Sti01b] Dale A. Stirling. Book review: Knowledge management for the information professional. T. Kanti Srikantaiah and Michael E. D. Koenig, editors. Journal of the American Society for Information Science and Technology: JASIST, 52 (3):277–278, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stirling:2002:BRB

[Sti02] Dale A. Stirling. Book review: Principles of Web design. Journal of the American Society for Information Science and Technology: JASIST, 53(11):968–969, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stirling:2003:BRB

[Sti03] Dale A. Stirling. Book review: Encoded archival description on the Internet. Journal of the American Society for Information Science and Technology: JASIST, 54(9):909– 910, July 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stirling:2005:BRB

[Sti05] Dale A. Stirling. Book review: Virtual reference training: the complete guide to providing anytime anywhere answers. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1238–1239, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stock:2006:BCR

[Sto06] Wolfgang G. Stock. Brief communication: On relevance distributions. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1126–1129, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stock:2010:CSR

[Sto10] Wolfgang G. Stock. Concepts and semantic relations in information science. Journal of the American Society for Information Science and Technology: JASIST, 61(10):1951–1969, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stock:2011:ICA

[Sto11] Wolfgang G. Stock. Informational cities: Analysis and construction of cities in the knowledge society. Journal of the American Society for Information Science and Technology: JASIST, 62(5):963–986, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Southon:2002:KMT

[STS02] F. C. Gray Southon, Ross J. Todd, and Megan Seneque. Knowledge management in three organizations: an exploratory study. Journal of the American Society for Information Science and Technology: JASIST, 53(12):1047– 1059, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sahib:2012:CAI

[STS12] Nuzhah Gooda Sahib, Anastasios Tombros, and Tony Stockman. A comparative analysis of the information-seeking behavior of visually impaired and sighted searchers. Journal of the American Society for Information Science and Technology: JASIST, 63(2):377–391, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Stvilia:2008:IQW

[STSG08] Besiki Stvilia, Michael B. Twidale, Linda C. Smith, and Les Gasser. Information quality work organization in Wikipedia. Journal of the American Society for Information Science and Technology: JASIST, 59(6):983–1001, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Su:2003:CSMa

[Su03a] Louise T. Su. A comprehensive and systematic model of user evaluation of Web search engines: I. Theory and background. Journal of the American Society for Information Science and Technology: JASIST, 54(13):1175–1192, November 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Su:2003:CSMb

[Su03b] Louise T. Su. A comprehensive and systematic model of user evaluation of Web search engines: II. An evaluation by undergraduates. Journal of the American Society for Information Science and Technology: JASIST, 54(13):1193–1223, November 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Su:2006:BRB

[Su06] Yu Su. Book review: Classifying science: Phenomena, data, theory, method, practice. Journal of the American Society for Information Science and Technology: JASIST, 57(14): 1977–1978, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Seki:2011:ODR

[SU11] Kazuhiro Seki and Kuniaki Uehara. Opinionated document retrieval using subjective triggers. Journal of the American Society for Information Science and Technology: JASIST, 62(5):861–876, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Su:2012:VGS

[Su12] Hsin-Ning Su. Visualization of global science and technology policy research structure. Journal of the American Society for Information Science and Technology: JASIST, 63(2): 242–255, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sundin:2006:BRB

[Sun06] Olof Sundin. Book review: Deflating information: From science studies to documentation. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1852–1853, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sun:2012:WDP

[Sun12] Jun Sun. Why different people prefer different systems for different tasks: an activity perspective on technology adoption in a dynamic user environment. Journal of the American Society for Information Science and Technology: JA-SIST, 63(1):48–63, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Suzuki:2009:ESS

[Suz09] Takafumi Suzuki. Extracting speaker-specific functional expressions from political speeches using random forests in order to investigate speakers’ political styles. Journal of the American Society for Information Science and Technology:

JASIST, 60(8):1596–1606, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spring:2003:PSA

[SV03] Michael B. Spring and Vichita Vathanophas. Peripheral social awareness information in collaborative work. Journal of the American Society for Information Science and Technology: JASIST, 54(11):1006–1013, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Serola:2005:AAC

[SV05] Sami Serola and Pertti Vakkari. The anticipated and assessed contribution of information types in references retrieved for preparing a research proposal. Journal of the American Society for Information Science and Technology: JASIST, 56(4):373–381, February 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schroeder:2012:GOC

[SW12] Andreas Schroeder and Christian Wagner. Governance of open content creation: a conceptualization and analysis of control and guiding mechanisms in the open content domain. Journal of the American Society for Information Science and Technology: JASIST, 63(10):1947–1959, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Swain:2011:BRB

[Swa11] Deborah E. Swain. Book review: Complex Worlds: Digital Culture, Rhetoric and Professional Communication. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1645–1647, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spink:2002:ISMa

[SWF^{+}02a] Amanda Spink, T. D. Wilson, Nigel Ford, Allen Foster, and David Ellis. Information-seeking and mediated searching. Part 1. Theoretical framework and research design. Journal of the American Society for Information Science and Technology: JASIST, 53(9):695–703, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spink:2002:ISMb

[SWF^{+}02b] Amanda Spink, T. D. Wilson, Nigel Ford, Allen Foster, and David Ellis. Information seeking and mediated searching study. Part 3. Successive searching. Journal of the American Society for Information Science and Technology: JASIST, 53(9):716–727, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Spink:2001:RSW

[SWJS01] Amanda Spink, Dietmar Wolfram, Major B. J. Jansen, and Tefko Saracevic. Research: Searching the web: the public and their queries. Journal of the American Society for Information Science and Technology: JASIST, 52(3):226–234, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Scholer:2004:QAS

[SWT04] Falk Scholer, Hugh E. Williams, and Andrew Turpin. Query association surrogates for Web search. Journal of the American Society for Information Science and Technology: JA-SIST, 55(7):637–650, May 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Schroeder:2007:ACL

[SXCC07] Jennifer Schroeder, Jennifer Xu, Hsinchun Chen, and Michael Chau. Automated criminal link analysis based on domain knowledge. Journal of the American Society for Information Science and Technology: JASIST, 58(6):842–855, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Squicciarini:2011:CEC

[SXZ11] Anna C. Squicciarini, Heng Xu, and Xiaolong (Luke) Zhang. CoPE: Enabling collaborative privacy management in online social networks. Journal of the American Society for Information Science and Technology: JASIST, 62(3):521–534, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shi:2007:MWR

[SY07] Xiaodong Shi and Christopher C. Yang. Mining Web resources for enhancing information retrieval: Mining related queries from Web search engine query logs using an improved association rule mining model. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1871–1883, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shen:2013:VHE

[SYL^{+}13] Jiantong Shen, Leye Yao, Youping Li, Mike Clarke, Li Wang, and Dan Li. Visualizing the history of evidence-based medicine: a bibliometric analysis. Journal of the American Society for Information Science and Technology: JASIST, 64(10):2157–2172, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Simeoni:2008:MHC

[SYNC08] Fabio Simeoni, Murat Yakici, Steve Neely, and Fabio Crestani. Metadata harvesting for content-based distributed information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 59(1):12–24, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Sun:2008:EAF

[SZ08] Heshan Sun and Ping Zhang. An exploration of affect factors and their role in user technology acceptance: Mediation and causality. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1252–1263, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Shapira:2011:PSI

[SZ11] Bracha Shapira and Boaz Zabar. Personalized search: Integrating collaboration and social networks. Journal of the American Society for Information Science and Technology: JASIST, 62(1):146–160, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Strotmann:2012:AND

[SZ12] Andreas Strotmann and Dangzhi Zhao. Author name disambiguation: What difference does it make in author-based citation analysis? Journal of the American Society for Information Science and Technology: JASIST, 63(9):1820–1833, September 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Suakkaphong:2011:DNE

[SZC11] Nichalin Suakkaphong, Zhu Zhang, and Hsinchun Chen. Disease named entity recognition using semisupervised learning and conditional random fields. Journal of the American Society for Information Science and Technology: JASIST, 62 (4):727–737, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Szostak:2010:LEC

[Szo10] Rick Szostak. Letters to the Editor: Comment on Hjørland’s concept theory. Journal of the American Society for Information Science and Technology: JASIST, 61(5):1076–1077, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Szostak:2011:CCB

[Szo11] Rick Szostak. Complex concepts into basic concepts. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2247–2265, November 2011. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Taddeo:2006:BRB

[Tad06] Laura Taddeo. Book review: The virtual reference experience: Integrating theory into practice. Journal of the American Society for Information Science and Technology: JA-SIST, 57(6):843–844, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Talja:2005:SDC

[Tal05] Sanna Talja. The social and discursive construction of computing skills. Journal of the American Society for Information Science and Technology: JASIST, 56(1):13–22, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Taniguchi:2004:DCR

[Tan04] Shoichi Taniguchi. Design of cataloging rules using conceptual modeling of cataloging process. Journal of the American Society for Information Science and Technology: JA-SIST, 55(6):498–512, April 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Taniguchi:2005:REB

[Tan05] Shoichi Taniguchi. Recording evidence in bibliographic records and descriptive metadata. Journal of the American Society for Information Science and Technology: JA-SIST, 56(8):872–882, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Taniguchi:2006:SSE

[Tan06] Shoichi Taniguchi. A system for supporting evidence recording in bibliographic records. Journal of the American Society for Information Science and Technology: JASIST, 57(9): 1249–1262, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tan:2007:EXC

[Tan07a] Bernard C. Y. Tan. Erratum: Y. Xu, C. Y. Tan and L. Yang, ‘Who will you ask? An empirical study of interpersonal task information seeking’. Journal of the American Society for Information Science and Technology 57 (12) 2006, 1666–1677. Journal of the American Society for Information Science and Technology: JASIST, 58(9):1387, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [XTY06].

Tang:2007:BSF

[Tan07b] Muh-Chyun Tang. Browsing and searching in a faceted information space: a naturalistic study of PubMed users’ interaction with a display tool. Journal of the American Society for Information Science and Technology: JASIST, 58 (13):1998–2006, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Taniguchi:2007:SSE

[Tan07c] Shoichi Taniguchi. A system for supporting evidence recording in bibliographic records, Part II: What is valuable evidence for catalogers? Journal of the American Society for Information Science and Technology: JASIST, 58(6):823– 841, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tarulli:2008:BRB

[Tar08] Laurel Tarulli. Book review: The law of libraries and archives. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2190–2191, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2013:TBS

[TB13] Mike Thelwall and Kevan Buckley. Topic-based sentiment analysis for the social web: the role of mood and issuerelated words. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1608–1617, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2010:SSD

[TBP^{+}10] Mike Thelwall, Kevan Buckley, Georgios Paltoglou, Di Cai, and Arvid Kappas. Sentiment strength detection in short informal text. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2544–2558, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Ano11].

Thelwall:2011:STE

[TBP11] Mike Thelwall, Kevan Buckley, and Georgios Paltoglou. Sentiment in Twitter events. Journal of the American Society for Information Science and Technology: JASIST, 62(2): 406–418, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2012:SSD

[TBP12] Mike Thelwall, Kevan Buckley, and Georgios Paltoglou. Sentiment strength detection for the social web. Journal of the American Society for Information Science and Technology: JASIST, 63(1):163–173, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tzanetakis:2004:MAR

[TC04] George Tzanetakis and Perry Cook. Music analysis and retrieval systems for audio signals. Journal of the American Society for Information Science and Technology: JASIST, 55(12):1077–1083, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tsai:2011:VWB

[TCW11] Richard Tzong-Han Tsai, Borong Chiu, and Chi-En Wu. Visual webpage block importance prediction using conditional random fields. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2225–2235, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Toms:2002:SHS

[TD02] Elaine G. Toms and Wendy Duff. “I spent 1\frac{1}{2} hours sifting through one large box. … ”: Diaries as information behavior of the archives user: Lessons learned. Journal of the American Society for Information Science and Technology: JA-SIST, 53(14):1232–1238, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Taylor:2013:SNS

[TDC13] Nolan J. Taylor, Alan R. Dennis, and Jeff W. Cummings. Situation normality and the shape of search: the effects of time delays and information presentation on search behavior. Journal of the American Society for Information Science and Technology: JASIST, 64(5):909–928, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tennis:2012:SCE

[Ten12] Joseph T. Tennis. The strange case of eugenics: a subject’s ontogeny in a long-lived classification scheme and the question of collocative integrity. Journal of the American Society for Information Science and Technology: JASIST, 63 (7):1350–1359, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2003:CBR

[TH03] Mike Thelwall and Gareth Harries. The connection between the research of a university and counts of links to its web pages: an investigation based upon a classification of the relationships of pages to the research of the host university. Journal of the American Society for Information Science and Technology: JASIST, 54(7):594–602, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2004:DWS

[TH04] Mike Thelwall and Gareth Harries. Do the Web sites of higher rated scholars have significantly more online impact? Journal of the American Society for Information Science and Technology: JASIST, 55(2):149–159, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tang:2013:TFK

[TH13] Li Tang and Guangyuan Hu. Tracing the footprint of knowledge spillover: Evidence from U.S.–China collaboration in nanotechnology. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1791–1801, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2001:EMI

[The01] Mike Thelwall. Extracting macroscopic information from Web links. Journal of the American Society for Information Science and Technology: JASIST, 52(13):1157–1168, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2002:CDW

[The02] Mike Thelwall. Conceptualizing documentation on the Web: an evaluation of different heuristic-based models for counting links between university Web sites. Journal of the American Society for Information Science and Technology: JASIST, 53(12):995–1005, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2005:TCE

[The05] Mike Thelwall. Text characteristics of English language university Web sites. Journal of the American Society for Information Science and Technology: JASIST, 56(6):609–619, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2006:ISS

[The06] Mike Thelwall. Interpreting social science link analysis research: a theoretical framework. Journal of the American Society for Information Science and Technology: JASIST, 57(1):60–68, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2008:EAC

[The08a] Mike Thelwall. Extracting accurate and complete results from search engines: Case study windows live. Journal of the American Society for Information Science and Technology: JASIST, 59(1):38–50, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2008:QCS

[The08b] Mike Thelwall. Quantitative comparisons of search engine results. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1702–1710, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2008:SNG

[The08c] Mike Thelwall. Social networks, gender, and friending: an analysis of MySpace member profiles. Journal of the American Society for Information Science and Technology: JA-SIST, 59(8):1321–1330, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2009:BRB

[The09a] Mike Thelwall. Book review: Handbook of Research on Web Log Analysis. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1943, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2009:HM

[The09b] Mike Thelwall. Homophily in MySpace. Journal of the American Society for Information Science and Technology: JASIST, 60(2):219–231, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thornton:2009:BRB

[Tho09] Grace Thornton. Book review: Everything is Miscellaneous: The Power of the New Digital Disorder. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1299–1300, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thomas:2012:OWP

[Tho12] Paul Thomas. Opinion: To what problem is distributed information retrieval the solution? Journal of the American Society for Information Science and Technology: JASIST, 63(7):1471–1476, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tierney:2010:BRB

[Tie10] William G. Tierney. Book review: For the Common Good: Principles of American Academic Freedom. Journal of the American Society for Information Science and Technology: JASIST, 61(1):211–212, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tijssen:2010:DBS

[Tij10] Robert J. W. Tijssen. Discarding the ‘basic science/applied science’ dichotomy: a knowledge utilization triangle classification system of research journals. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1842–1852, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Talvensaari:2007:CBC

[TJLJ07] Tuomas Talvensaari, Martti Juhola, Jorma Laurikkala, and Kalervo Järvelin. Corpus-based cross-language information retrieval in retrieval of highly relevant documents. Journal of the American Society for Information Science and Technology: JASIST, 58(3):322–334, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thornley:2011:SIT

[TJSL11] Clare V. Thornley, Andrea C. Johnson, Alan F. Smeaton, and Hyowon Lee. The scholarly impact of TRECVid (2003– 2009). Journal of the American Society for Information Science and Technology: JASIST, 62(4):613–627, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tsuji:2006:AGJ

[TK06] Keita Tsuji and Kyo Kageura. Automatic generation of Japanese–English bilingual thesauri based on bilingual corpora. Journal of the American Society for Information Science and Technology: JASIST, 57(7):891–906, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2008:OPS

[TK08] Mike Thelwall and Kayvan Kousha. Online presentations as a source of scientific impact? An analysis of PowerPoint files citing academic journals. Journal of the American Society for Information Science and Technology: JASIST, 59 (5):805–815, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tenopir:2005:REJ

[TKB^{+}05a] Carol Tenopir, Donald W. King, Peter Boyce, Matt Grayson, and Keri-Lynn Paulson. Relying on electronic journals: Reading patterns of astronomers. Journal of the American Society for Information Science and Technology: JASIST, 56(8):786–802, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thompson:2005:SIR

[TKB05b] E. Dale Thompson and Michelle L. Kaarst-Brown. Sensitive information: a review and research agenda. Journal of the American Society for Information Science and Technology: JASIST, 56(3):245–257, February 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2010:PRW

[TKV^{+}10] Mike Thelwall, Antje Klitkou, Arnold Verbeek, David Stuart, and Celine Vincent. Policy-relevant Webometrics for individual scientific fields. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1464– 1475, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tam:2001:SNL

[TL01] A. M. Tam and C. H. C. Leung. Structured natural-language descriptions for semantic content retrieval of visual materials. Journal of the American Society for Information Science and Technology: JASIST, 52(11):930–937, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thornton:2005:URT

[TL05] K. E. B. Thornton and A. Lazzarini. The use of reliability techniques to predict the recovery rate of recovering alcohol addicts. Journal of the American Society for Information Science and Technology: JASIST, 56(4):356–363, February 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tsagkias:2010:PPP

[TLdR10] Manos Tsagkias, Martha Larson, and Maarten de Rijke. Predicting podcast preference: an analysis framework and its application. Journal of the American Society for Information Science and Technology: JASIST, 61(2):374–391, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Torra:2006:ICE

[TLM06] Vicenç Torra, Sergi Lanau, and Sadaaki Miyamoto. Image clustering for the exploration of video sequences. Journal of the American Society for Information Science and Technology: JASIST, 57(4):577–584, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tashiro:2012:MNL

[TLM^{+}12] Hisato Tashiro, Antonio Lau, Junichiro Mori, Nobuzumi Fujii, and Yuya Kajikawa. E-mail networks and leadership performance. Journal of the American Society for Information Science and Technology: JASIST, 63(3):600–606, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tsakonas:2013:EDL

[TMPP13] Giannis Tsakonas, Angelos Mitrelis, Leonidas Papachristopoulos, and Christos Papatheodorou. An exploration of the digital library evaluation literature based on an ontological representation. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1914–1926, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Toldo:2001:LEM

[Tol01a] Luca Toldo. Letter to the Editor: Measurement of cluster quality. Journal of the American Society for Information Science and Technology: JASIST, 52(7):602, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Toldo:2001:BRB

[Tol01b] Luca I. G. Toldo. Book review: After the Internet: Alien intelligence, by James Martin. Journal of the American Society for Information Science and Technology: JASIST, 52 (7):599–600, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tol:2009:BCM

[Tol09] Richard S. J. Tol. Brief communications: The Matthew effect defined and tested for the 100 most prolific economists. Journal of the American Society for Information Science and Technology: JASIST, 60(2):420–426, February 2009. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tomer:2001:BRB

[Tom01a] Christinger Tomer. Book review: Internet publishing and beyond: the economics of digital information and intellectual property, edited by Brian Kahin and Hal R. Varian. Journal of the American Society for Information Science and Technology: JASIST, 52(12):1081–1082, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tompson:2001:BRB

[Tom01b] Sara R. Tompson. Book review: Change management in information services, by Lyndon Pugh. Journal of the American Society for Information Science and Technology: JA-SIST, 52(9):779–780, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Toms:2002:IIP

[Tom02] Elaine G. Toms. Information interaction: Providing a framework for information architecture. Journal of the American Society for Information Science and Technology: JASIST, 53(10):855–862, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tompson:2004:BRB

[Tom04] Sara R. Tompson. Book review: The bottom line: Determining and communicating the value of the special library. Journal of the American Society for Information Science and Technology: JASIST, 55(5):465–467, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tomer:2005:BRBa

[Tom05a] Christinger Tomer. Book review: Connected, or what it means to live in the network society. Journal of the American Society for Information Science and Technology: JA-SIST, 56(10):1127–1128, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tomer:2005:BRBb

[Tom05b] Christinger Tomer. Book review: Humanizing information technology. Journal of the American Society for Information Science and Technology: JASIST, 56(12):1360, October 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tompson:2006:BRB

[Tom06] Sara R. Tompson. Book review: Digital library use: Social practice in design and evaluation. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1130–1132, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2006:LES

[TP06] Mike Thelwall and Liz Price. Language evolution and the spread of ideas on the Web: a procedure for identifying emergent hybrid word family members. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1326–1337, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2007:ICP

[TP07] Mike Thelwall and Rudy Prabowo. Identifying and characterizing public science-related fears from RSS feeds. Journal of the American Society for Information Science and Technology: JASIST, 58(3):379–390, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tsakonas:2011:ORD

[TP11] Giannis Tsakonas and Christos Papatheodorou. An ontological representation of the digital library evaluation domain. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1577–1593, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2006:RRF

[TPF06] Mike Thelwall, Rudy Prabowo, and Ruth Fairclough. Are raw RSS feeds suitable for broad issue scanning? A science concern case study. Journal of the American Society for Information Science and Technology: JASIST, 57(12): 1644–1654, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tan:2005:MKC

[TPLC05] Chee Wee Tan, Shan Ling Pan, Eric Tze Kuan Lim, and Calvin Meng Lai Chan. Managing knowledge conflicts in an interorganizational project: a case study of the Infocomm Development Authority of Singapore. Journal of the American Society for Information Science and Technology: JA-SIST, 56(11):1187–1199, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Toldo:2005:IBA

[TR05] Luca Toldo and Friedrich Rippmann. Integrated bioinformatics application for automated target discovery. Journal of the American Society for Information Science and Technology: JASIST, 56(5):483–492, March 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Trace:2008:RUI

[Tra08] Ciaran B. Trace. Resistance and the underlife: Informal written literacies and their relationship to human information behavior. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1540–1554, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Trauth:2009:BRB

[Tra09] Eileen M. Trauth. Book review: The Information Revolution and World Politics. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2598– 2600, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Trevorrow:2012:LEU

[Tre12] Paul Trevorrow. Letter to the Editor: The use of H-index for the assessment of journals’ performance will lead to shifts in editorial policies — a response. Journal of the American Society for Information Science and Technology: JASIST, 63(4):845–846, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tombros:2005:HUA

[TRJ05] Anastasios Tombros, Ian Ruthven, and Joemon M. Jose. How users assess Web pages for information seeking. Journal of the American Society for Information Science and Technology: JASIST, 56(4):327–344, February 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tosaka:2013:AIS

[TrP13] Yuji Tosaka and Jung ran Park. Advances in information science: RDA: Resource description & access — a survey of the current state of the art. Journal of the American Society for Information Science and Technology: JASIST, 64(4):651–662, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tang:2001:RUR

[TS01] Rong Tang and Paul Solomon. Research: Use of relevance criteria across stages of document evaluation: On the complementarity of experimental and naturalistic studies. Journal of the American Society for Information Science and Technology: JASIST, 52(8):676–685, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2006:WCE

[TS06] Mike Thelwall and David Stuart. Web crawling ethics revisited: Cost, privacy, and denial of service. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1771–1779, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tennis:2008:ESK

[TS08] Joseph T. Tennis and Stuart A. Sutton. Extending the simple knowledge organization system for concept management in vocabulary development applications. Journal of the American Society for Information Science and Technology: JASIST, 59(1):25–37, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tann:2009:BCW

[TS09] Chadwyn Tann and Mark Sanderson. Brief communications: Are Web-based informational queries changing? Journal of the American Society for Information Science and Technology: JASIST, 60(6):1290–1293, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2011:CMC

[TS11] Mike Thelwall and Pardeep Sud. A comparison of methods for collecting web citation data for academic organizations. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1488–1497, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Truran:2011:EUI

[TSA11] Mark Truran, Jan-Felix Schmakeit, and Helen Ashman. The effect of user intent on the stability of search engine results. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1276–1287, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tseng:2001:ACS

[Tse01] Yuen-Hsien Tseng. Automatic cataloguing and searching for retrospective data by use of OCR text. Journal of the American Society for Information Science and Technology: JASIST, 52(5):378–390, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tseng:2002:ATG

[Tse02] Yuen-Hsien Tseng. Automatic thesaurus generation for Chinese documents. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1130–1138, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tsikerdekis:2013:EPA

[Tsi13] Michail Tsikerdekis. The effects of perceived anonymity and anonymity states on conformity and groupthink in online communities: A Wikipedia study. Journal of the American Society for Information Science and Technology: JASIST, 64(5):1001–1015, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tjondronegoro:2009:SCM

[TSJ09] Dian Tjondronegoro, Amanda Spink, and Bernard J. Jansen. A study and comparison of multimedia Web searching: 1997–2006. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1756–1768, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Torres-Salinas:2013:UBA

[TSRGJC^{+}13] Daniel Torres-Salinas, Nicolás Robinson-García, Evaristo Jiménez-Contreras, Francisco Herrera, and Emilio Delgado López-Cózar. On the use of biplot analysis for multivariate bibliometric and scientific indicators. Journal of the American Society for Information Science and Technology: JA-SIST, 64(7):1468–1479, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2012:CYV

[TSV12] Mike Thelwall, Pardeep Sud, and Farida Vis. Commenting on YouTube videos: From Guatemalan rock to El Big Bang. Journal of the American Society for Information Science and Technology: JASIST, 63(3):616–629, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2012:LCI

[TSW12] Mike Thelwall, Pardeep Sud, and David Wilkinson. Link and co-inlink network diagrams with URL citations or title mentions. Journal of the American Society for Information Science and Technology: JASIST, 63(4):805–816, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tuominen:2003:MDL

[TTS03] Kimmo Tuominen, Sanna Talja, and Reijo Savolainen. Multiperspective digital libraries: the implications of constructionism for the development of digital libraries. Journal of the American Society for Information Science and Technology: JASIST, 54(6):561–569, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thellefsen:2013:EIC

[TTS13] Torkild Thellefsen, Martin Thellefsen, and Bent Sørensen. Emotion, information, and cognition, and some possible consequences for library and information science. Journal of the American Society for Information Science and Technology: JASIST, 64(8):1735–1750, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tan:2012:CIE

[TTT12] Wee-Kek Tan, Chuan-Hoo Tan, and Hock-Hai Teo. Conveying information effectively in a virtual world: Insights from synthesized task closure and media richness. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1198–1212, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tonta:2005:SJL

[TÜ05] Yaşar Tonta and Yurdagül Ünal. Scatter of journals and literature obsolescence reflected in document delivery requests. Journal of the American Society for Information Science and Technology: JASIST, 56(1):84–94, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2004:IWI

[TV04] Mike Thelwall and Liwen Vaughan. Introduction: Webometrics: an introduction to the special issue. Journal of the American Society for Information Science and Technology: JASIST, 55(14):1213–1215, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2006:WIA

[TVF06] Mike Thelwall, Katie Vann, and Ruth Fairclough. Web issue analysis: an integrated water resource management case study. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1303–1314, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Talja:2007:IRC

[TVFW07] Sanna Talja, Pertti Vakkari, Jenny Fry, and Paul Wouters. Impact of research cultures on the use of digital library resources. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1674–1685, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Torney:2012:UPF

[TVY12] Rosemary Torney, Peter Vamplew, and John Yearwood. Using psycholinguistic features for profiling first language of authors. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1256–1269, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2003:GST

[TW03a] Mike Thelwall and David Wilkinson. Graph structure in three national academic Webs: Power laws with anomalies. Journal of the American Society for Information Science and Technology: JASIST, 54(8):706–712, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2003:TTD

[TW03b] Mike Thelwall and David Wilkinson. Three target document range metrics for university web sites. Journal of the American Society for Information Science and Technology: JASIST, 54(6):490–497, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2010:PDS

[TW10] Mike Thelwall and David Wilkinson. Public dialogs in social network sites: What is their purpose? Journal of the American Society for Information Science and Technology:

JASIST, 61(2):392–404, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2008:BCI

[TWF08] Mike Thelwall, Paul Wouters, and Jenny Fry. Brief communications: Information-centered research for large-scale analyses of new information sources. Journal of the American Society for Information Science and Technology: JA-SIST, 59(9):1523–1527, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tartanus:2013:BCG

[TWKH13] Małgorzata Tartanus, Agnieszka Wnuk, Marcin Kozak, and James Hartley. Brief communication: Graphs and prestige in agricultural journals. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1946– 1950, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Torvik:2005:PSM

[TWSS05] Vetle I. Torvik, Marc Weeber, Don R. Swanson, and Neil R. Smalheiser. A probabilistic similarity metric for Medline records: a model for author name disambiguation. Journal of the American Society for Information Science and Technology: JASIST, 56(2):140–158, January 15, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Thelwall:2010:DME

[TWU10] Mike Thelwall, David Wilkinson, and Sukhvinder Uppal. Data mining emotion in social network communication: Gender differences in MySpace. Journal of the American Society for Information Science and Technology: JASIST, 61(1):190–199, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Teo:2006:OLC

[TWW^{+}06] Hock-Hai Teo, Xinwei Wang, Kwok-Kee Wei, Choon-Ling Sia, and Matthew K. O. Lee. Organizational learning capacity and attitude toward complex technological innovations: an empirical study. Journal of the American Society for Information Science and Technology: JASIST, 57(2):264–279, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tang:2013:UEM

[TYS13] Xuning Tang, Christopher C. Yang, and Min Song. Understanding the evolution of multiple scientific research domains using a content and network approach. Journal of the American Society for Information Science and Technology: JA-SIST, 64(5):1065–1075, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Urbain:2008:PPM

[UGF08] Jay Urbain, Nazli Goharian, and Ophir Frieder. Probabilistic passage models for semantic search of genomics literature. Journal of the American Society for Information Science and Technology: JASIST, 59(12):2008–2023, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Haque:2009:PEC

[uHG09] Asif ul Haque and Paul Ginsparg. Positional effects on citation and readership in arXiv. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2203–2218, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Haque:2010:LLA

[uHG10] Asif ul Haque and Paul Ginsparg. Last but not least: Additional positional effects on citation and readership in arXiv. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2381–2388, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Urquhart:2007:USI

[UR07] Christine Urquhart and Jennifer Rowley. Understanding student information behavior in relation to electronic information services: Lessons from longitudinal monitoring and evaluation, Part 2. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1188–1197, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Urquhart:2009:BRB

[Urq09] Christine Urquhart. Book review: Paper to Digital: Documents in the Information Age. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1945–1947, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Uitdenbogerd:2004:AEM

[UZ04] Alexandra L. Uitdenbogerd and Justin Zobel. An architecture for effective music information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 55(12):1053–1057, October 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vakkari:2008:PIU

[Vak08] Pertti Vakkari. Perceived influence of the use of electronic information resources on scholarly work and publication productivity. Journal of the American Society for Information Science and Technology: JASIST, 59(4):602–612, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

VanRoekel:2003:BRB

[Van03] James L. Van Roekel. Book review: Digital creativity: Techniques for digital media and the Internet. Journal of the American Society for Information Science and Technology: JASIST, 54(4):357–358, February 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vanclay:2007:BCR

[Van07] Jerome K. Vanclay. Brief communication: On the robustness of the h-index. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1547–1550, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

VanderSanden:2010:BRB

[Van10] Maarten C. A. Van der Sanden. Book review: Investigating Science Communication in the Information Age. Practising Science Communication in the Information Age. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1508–1511, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

VanderVeerMartens:2011:PPT

[Van11] Betsy Van der Veer Martens. The production of practice theories. Journal of the American Society for Information Science and Technology: JASIST, 62(3):586–593, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

VanCalster:2012:BCI

[Van12] Ben Van Calster. Brief communication: It takes time: a remarkable example of delayed recognition. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2341–2344, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vaughan:2003:CUP

[Vau03] K. T. L. Vaughan. Changing use patterns of print journals in the digital age: Impacts of electronic equivalents on print chemistry journal use. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1149–1152, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vaughan:2006:VLC

[Vau06] Liwen Vaughan. Visualizing linguistic and cultural differences using Web co-link data. Journal of the American Society for Information Science and Technology: JASIST, 57 (9):1178–1193, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vivanco:2011:BCB

[VBMM11] Luis Vivanco, Blanca Bartolomé, Montserrat San Martín, and Alfredo Martínez. Brief communications: Bibliometric analysis of the use of the term preembryo in scientific literature. Journal of the American Society for Information Science and Technology: JASIST, 62(5):987–991, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vishwanath:2006:TCU

[VC06] Arun Vishwanath and Hao Chen. Technology clusters: Using multidimensional scaling to evaluate and structure technology clusters. Journal of the American Society for Information Science and Technology: JASIST, 57(11):1451–1460, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vishwanath:2008:PCT

[VC08] Arun Vishwanath and Hao Chen. Personal communication technologies as an extension of the self: a cross-cultural comparison of people’s associations with technology and their symbolic proximity with others. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1761–1775, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

VanRijnsoever:2009:PTC

[VC09] Frank J. Van Rijnsoever and Carolina Castaldi. Perceived technology clusters and ownership of related technologies: the case of consumer electronics. Journal of the American Society for Information Science and Technology: JASIST, 60(2):381–392, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanDijk:2010:BRB

[vD10] Jan A. G. M. van Dijk. Book review: The Myth of Digital Democracy; Digital Citizenship, the Internet, Society and Participation. Journal of the American Society for Information Science and Technology: JASIST, 61(3):631–633, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vandenBesselaar:2003:BCE

[vdB03a] Peter van den Besselaar. Brief communications: Empirical evidence of self-organization? Journal of the American Society for Information Science and Technology: JASIST, 54(1): 87–90, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See rejoinder [Ley03a].

vandenBesselaar:2003:LED

[vdB03b] Peter van den Besselaar. Letters to the Editor: Descriptive statistics, inferential statistics, rhetorical statistics. Journal of the American Society for Information Science and Technology: JASIST, 54(11):1077, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See rejoinder [Ley03b].

vanderEijk:2004:CAC

[vdEvMK^{+}04] C. Christiaan van der Eijk, Erik M. van Mulligen, Jan A. Kors, Barend Mons, and Jan van den Berg. Constructing an associative concept space for literature-based discovery. Journal of the American Society for Information Science and Technology: JASIST, 55(5):436–444, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanDalen:2012:IUC

[vDH12] Hendrik P. van Dalen and Kène Henkens. Intended and unintended consequences of a publish-or-perish culture: a worldwide survey. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1282–1293, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vandenHeuvel:2011:FIP

[vdHR11] Charles van den Heuvel and W. Boyd Rayward. Facing interfaces: Paul Otlet’s visualizations of data integration. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2313–2326, December 2011. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Verleysen:2013:BCL

[VE13] Frederik T. Verleysen and Tim C. E. Engels. Brief communication: A label for peer-reviewed books. Journal of the American Society for Information Science and Technology: JASIST, 64(2):428–430, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Veith:2006:BRB

[Vei06a] Richard H. Veith. Book review: Understanding digital libraries. Journal of the American Society for Information Science and Technology: JASIST, 57(6):844–846, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Veith:2006:MII

[Vei06b] Richard H. Veith. Memex at 60: Internet or iPod? Journal of the American Society for Information Science and Technology: JASIST, 57(9):1233–1242, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Veinot:2009:IAS

[Vei09] Tiffany C. Veinot. Interactive acquisition and sharing: Understanding the dynamics of HIV/ AIDS information networks. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2313–2332, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanEck:2008:ASM

[vEW08] Nees Jan van Eck and Ludo Waltman. Appropriate similarity measures for author co-citation analysis. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1653–1661, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanEck:2009:HNC

[vEW09] Nees Jan van Eck and Ludo Waltman. How to normalize cooccurrence data? An analysis of some well-known similarity measures. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1635–1651, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanEck:2010:CTT

[vEWDvdB10] Nees Jan van Eck, Ludo Waltman, Rommert Dekker, and Jan van den Berg. A comparison of two techniques for bibliometric mapping: Multidimensional scaling and VOS. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2405–2416, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

VanderVeerMartens:2006:DTF

[VG06] Betsy Van der Veer Martens and Abby A. Goodrum. The diffusion of theories: a functional approach. Journal of the American Society for Information Science and Technology: JASIST, 57(3):330–341, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vicedo:2007:BRBb

[VG07a] José L. Vicedo and Jaime Gómez. Book review: Georeferencing: the geographic associations of information. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2401–2402, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vicedo:2007:BRBa

[VG07b] José L. Vicedo and Jaime Gómez. Book review: Trec: Experiment and evaluation in information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 58(6):910–911, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vickers:2012:M

[VG12] Peter Vickers and Alan Gilchrist. In memoriam. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1482–1483, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vasalou:2011:PDN

[VGM^{+}11] Asimina Vasalou, Alastair J. Gill, Fadhila Mazanderani, Chrysanthi Papoutsi, and Adam Joinson. Privacy dictionary: a new resource for the automated content analysis of privacy. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2095–2105, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vaughan:2005:PNE

[VH05] Liwen Vaughan and Trudi Bellardo Hahn. Profile, needs, and expectations of information professionals: What we learned from the 2003 ASIST membership survey. Journal of the American Society for Information Science and Technology:

JASIST, 56(1):95–105, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vakkari:2012:SED

[VH12] Pertti Vakkari and Saila Huuskonen. Search effort degrades search output but improves task outcome. Journal of the American Society for Information Science and Technology: JASIST, 63(4):657–670, April 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanHooland:2013:ESV

[vHVD^{+}13] Seth van Hooland, Ruben Verborgh, Max De Wilde, Johannes Hercher, Erik Mannens, and Rik Van de Walle. Evaluating the success of vocabulary reconciliation for cultural heritage collections. Journal of the American Society for Information Science and Technology: JASIST, 64(3):464–479, March 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vilar:2010:BRBb

[Vil10a] Polona Vilar. Book review: Designing the User Interface: Strategies for Effective Human-Computer Interaction (5th edition). Journal of the American Society for Information Science and Technology: JASIST, 61(5):1073–1074, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vilar:2010:BRBa

[Vil10b] Polona Vilar. Book review: Usability Engineering: Process, Products and Examples. Journal of the American Society for Information Science and Technology: JASIST, 61(4):854– 855, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vinkler:2004:CIS

[Vin04] Peter Vinkler. Characterization of the impact of sets of scientific papers: The Garfield (impact) factor. Journal of the American Society for Information Science and Technology: JASIST, 55(5):431–435, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vinkler:2011:ADC

[Vin11a] Péter Vinkler. Application of the distribution of citations among publications in scientometric evaluations. Journal of the American Society for Information Science and Technology: JASIST, 62(10):1963–1978, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vinkler:2011:LES

[Vin11b] Peter Vinkler. Letters to the Editor: Scientometrics and scientometricians in 2011. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1430– 1432, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vinkler:2013:LEQ

[Vin13] Peter Vinkler. Letter to the Editor: Quantity and impact through a single indicator. Journal of the American Society for Information Science and Technology: JASIST, 64(5): 1084–1085, May 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vishwanath:2005:IPT

[Vis05] Arun Vishwanath. Impact of personality on technology adoption: an empirical model. Journal of the American Society for Information Science and Technology: JASIST, 56 (8):803–811, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vital:2008:BRB

[Vit08] Sarah M. Vital. Book review: Annual review of information science and technology, Vol. 41. Journal of the American Society for Information Science and Technology: JASIST, 59(1):161–162, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Velden:2013:ECS

[VL13] Theresa Velden and Carl Lagoze. The extraction of community structures from publication networks to support ethnographic observations of field differences in scientific communication. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2405–2427, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanRaan:2006:PRD

[vR06a] Anthony F. J. van Raan. Performance-related differences of bibliometric statistical properties of research groups: Cumulative advantages and hierarchically layered networks. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1919–1935, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanRaan:2006:SPB

[vR06b] Anthony F. J. van Raan. Statistical properties of bibliometric indicators: Research group indicator distributions and correlations. Journal of the American Society for Information Science and Technology: JASIST, 57(3):408–430, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vaughan:2006:PSU

[VR06c] Misha W. Vaughan and Marc L. Resnick. Perspectives on search user interfaces: Best practices and future visions: Search user interfaces: Best practices and future visions. Journal of the American Society for Information Science and Technology: JASIST, 57(6):777–780, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanRaan:2008:BSP

[vR08a] Anthony F. J. van Raan. Bibliometric statistical properties of the 100 largest European research universities: Prevalent scaling rules in the science system. Journal of the American Society for Information Science and Technology: JASIST, 59(3):461–475, February 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanRaan:2008:SRS

[vR08b] Anthony F. J. van Raan. Scaling rules in the science system: Influence of field-specific citation characteristics on the impact of research groups. Journal of the American Society for Information Science and Technology: JASIST, 59 (4):565–576, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanRaan:2008:SCI

[vR08c] Anthony F. J. van Raan. Self-citation as an impactreinforcing mechanism in the science system. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1631–1643, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanRijnsoever:2011:ECC

[vRC11] Frank J. van Rijnsoever and Carolina Castaldi. Extending consumer categorization based on innovativeness: Intentions and technology clusters in consumer electronics. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1604–1613, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanRijnsoever:2009:EID

[vRD09] Frank J. van Rijnsoever and A. Rogier T. Donders. The effect of innovativeness on different levels of technology adoption. Journal of the American Society for Information Science and Technology: JASIST, 60(5):984–996, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vanRijnsoever:2010:UPI

[vRFBvW10] Frank J. van Rijnsoever, Jan Faber, Marnix L. J. Brinkman, and Marijn A. van Weele. User-producer interaction in Web site development: Motives, modes, and misfits. Journal of the American Society for Information Science and Technology: JASIST, 61(3):495–504, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vibert:2009:EDK

[VRL^{+}09] Nicolas Vibert, Christine Ros, Ludovic Le Bigot, Mélanie Ramond, Jérôme Gatefin, and Jean-François Rouet. Effects of domain knowledge on reference search with the PubMed database: an experimental study. Journal of the American Society for Information Science and Technology: JASIST, 60(7):1423–1447, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vaughan:2003:BWC

[VS03] Liwen Vaughan and Debora Shaw. Bibliographic and Web citations: What is the difference? Journal of the American Society for Information Science and Technology: JASIST, 54(14):1313–1322, December 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vakkari:2004:IRL

[VS04] Pertti Vakkari and Eero Sormunen. The influence of relevance levels on the effectiveness of interactive information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 55(11):963–969, September 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vaughan:2005:WCD

[VS05] Liwen Vaughan and Debora Shaw. Web citation data for impact assessment: a comparison of four science disciplines. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1075–1087, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vaughan:2003:SUW

[VT03] Liwen Vaughan and Mike Thelwall. Scholarly use of the Web: What are the key inducers of links to journal Web sites? Journal of the American Society for Information Science and Technology: JASIST, 54(1):29–38, January 1, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

vonThaden:2008:DIB

[vT08] Terry L. von Thaden. Distributed information behavior: a study of dynamic practice in a safety critical environment. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1555–1569, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vicedo:2013:BRB

[VT13] José L. Vicedo and David Tomás. Book review: Multisource, multilingual information extraction and summarization — Edited by Thierry Poibeau, Horacio Saggion, Jakub Piskorski and Roman Yangarber. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1519–1521, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vaughan:2010:CBP

[VTD10] Liwen Vaughan, Juan Tang, and Jian Du. Constructing business profiles based on keyword patterns on Web sites. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1120–1129, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

VanderVeerMartens:2012:OBB

[VV12] Betsy Van der Veer Martens and Connie Van Fleet. Opening the black box of “relevance work”: a domain analysis. Journal of the American Society for Information Science and Technology: JASIST, 63(5):936–947, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Verberne:2013:RVQ

[VvdHH^{+}13] Suzan Verberne, Maarten van der Heijden, Max Hinne, Maya Sappelli, Saskia Koldijk, Eduard Hoenkamp, and Wessel Kraaij. Reliability and validity of query intent assessments. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2224–2237, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vanopstal:2012:PSD

[VVLB12] Klaar Vanopstal, Robert Vander Stichele, Godelieve Laureys, and Joost Buysschaert. PubMed searches by Dutchspeaking nursing students: the impact of language and system experience. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1538–1552, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Veinot:2012:AIS

[VW12] Tiffany C. Veinot and Kate Williams. Advances in information science: Following the “community” thread from sociology to information behavior and informatics: Uncovering theoretical continuities and research opportunities. Journal of the American Society for Information Science and Technology: JASIST, 63(5):847–864, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vaughan:2012:WDA

[VY12] Liwen Vaughan and Rongbin Yang. Web data as academic and business quality estimates: a comparison of three data sources. Journal of the American Society for Information Science and Technology: JASIST, 63(10):1960–1972, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Vilar:2008:PIU

[VŽ08] Polona Vilar and Maja Žumer. Perceptions and importance of user friendliness of IR systems according to users’ individual characteristics and academic discipline. Journal of the American Society for Information Science and Technology: JASIST, 59(12):1995–2007, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Watters:2003:GLB

[WA03] Carolyn Watters and Ghada Amoudi. GeoSearcher: Location-based ranking of search engine results. Journal of the American Society for Information Science and Technology: JASIST, 54(2):140–151, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wademan:2001:BRB

[Wad01] Mark R. Wademan. Book review: XML: a manager’s guide, by Kevin Dick. Journal of the American Society for Information Science and Technology: JASIST, 52(3):274–275, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wagner:2003:MTE

[Wag03] A. Ben Wagner. Managing tradeoffs in the electronic age. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1160–1164, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wallace:2002:BRB

[Wal02] Raven Wallace. Book review: Usability for the Web: Designing Web sites that work. Journal of the American Society for Information Science and Technology: JASIST, 53 (9):775–777, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Walker:2004:BR

[Wal04a] Jessie Walker. Book review. Journal of the American Society for Information Science and Technology: JASIST, 55(2): 180–181, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wallace:2004:LEM

[Wal04b] Danny P. Wallace. Letters to the Editor: The Mann– Whitney Test. Journal of the American Society for Information Science and Technology: JASIST, 55(1):93, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Walters:2006:MUJ

[Wal06] Glenn D. Walters. Measuring the utility of journals in the crime-psychology field: Beyond the impact factor. Journal of the American Society for Information Science and Technology: JASIST, 57(13):1804–1813, November 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Walters:2007:IJC

[Wal07] William H. Walters. Institutional journal costs in an open access environment. Journal of the American Society for Information Science and Technology: JASIST, 58(1):108– 120, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Waller:2011:JIW

[Wal11a] Vivienne Waller. Not just information: Who searches for what on the search engine Google? Journal of the American Society for Information Science and Technology: JASIST, 62(4):761–775, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Walters:2011:CLC

[Wal11b] Glenn D. Walters. The citation life cycle of articles published in 13 American Psychological Association journals: a 25-year longitudinal analysis. Journal of the American Society for Information Science and Technology: JASIST, 62(8):

1629–1636, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Walsh:2012:IGF

[Wal12] John A. Walsh. “images of God and friends of God”: the holy icon as document. Journal of the American Society for Information Science and Technology: JASIST, 63(1): 185–194, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2006:ATD

[Wan06] Jun Wang. Automatic thesaurus development: Term extraction from title metadata. Journal of the American Society for Information Science and Technology: JASIST, 57 (7):907–920, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2007:BRB

[Wan07] Chingning Wang. Book review: New directions in human information behavior. Journal of the American Society for Information Science and Technology: JASIST, 58(10):1553, August 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2009:ESA

[Wan09] Jun Wang. An extensive study on automated Dewey Decimal Classification. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2269– 2286, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2002:BRB

[War02] Julian Warner. Book review: A nation transformed by information: How information has shaped the United States from colonial times to the present. Journal of the American Society for Information Science and Technology: JASIST, 53(9):775, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2004:BR

[War04] Julian Warner. Book review. Journal of the American Society for Information Science and Technology: JASIST, 55(1):

89–90, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2005:BRB

[War05] Julian Warner. Book review: Information systems and the economics of innovation. Journal of the American Society for Information Science and Technology: JASIST, 56 (8):889–890, June 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2006:BRB

[War06] Julian Warner. Book review: Protocol: How control exists after decentalization. Journal of the American Society for Information Science and Technology: JASIST, 57(1):146– 147, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2007:ABL

[War07a] Julian Warner. Analogies between linguistics and information theory. Journal of the American Society for Information Science and Technology: JASIST, 58(3):309–321, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2007:BRB

[War07b] Julian Warner. Book review: The information revolution and Ireland: Prospects and challenges. Journal of the American Society for Information Science and Technology: JA-SIST, 58(5):760–761, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2007:DSL

[War07c] Julian Warner. Description and search labor for information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1783–1790, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2007:LIT

[War07d] Julian Warner. Linguistics and information theory: Analytic advantages. Journal of the American Society for Information Science and Technology: JASIST, 58(2):275–285, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2007:SPS

[War07e] Julian Warner. Selection power and selection labor for information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 58(7):915–923, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2008:LTA

[War08] Julian Warner. A labor theoretic approach to information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 59(5):731–741, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2009:BRB

[War09a] Julian Warner. Book review: Information Revolutions in the History of the West. Journal of the American Society for Information Science and Technology: JASIST, 60(12): 2591–2592, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warwick:2009:BRB

[War09b] Claire Warwick. Book review: From Papyrus to Hypertext: Towards the Universal Digital Library. Journal of the American Society for Information Science and Technology: JA-SIST, 60(9):1947–1948, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2010:ACF

[War10a] Julian Warner. The absence of creativity in Feist and the computational process. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2324– 2336, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2010:BRB

[War10b] Julian Warner. Book review: Not Exactly: In Praise of Vagueness. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2584–2585, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2010:MRO

[War10c] Julian Warner. So mechanical or routine: the not original in Feist. Journal of the American Society for Information Science and Technology: JASIST, 61(4):820–834, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warwick:2010:BRB

[War10d] Claire Warwick. Book review: Text Editing, Print and the Digital World. Journal of the American Society for Information Science and Technology: JASIST, 61(2):428–430, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2013:CF

[War13] Julian Warner. Creativity for Feist. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1173–1192, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wathen:2002:BIF

[WB02] C. Nadine Wathen and Jacquelyn Burkell. Believe it or not: Factors influencing credibility on the Web. Journal of the American Society for Information Science and Technology: JASIST, 53(2):134–144, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

West:2010:BME

[WBB10a] Jevin West, Theodore Bergstrom, and Carl T. Bergstrom. Big Macs and Eigenfactor scores: Don’t let correlation coefficients fool you. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1800–1807, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

West:2010:LER

[WBB10b] Jevin D. West, Theodore Bergstrom, and Carl T. Bergstrom. Letter to the Editor: Response to “Big Macs and Eigenfactor scores: the correlation conundrum”. Journal of the American Society for Information Science and Technology: JA-SIST, 61(12):2592, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Whitty:2012:ALS

[WBJM12] Monica T. Whitty, Tom Buchanan, Adam N. Joinson, and Alex Meredith. Not all lies are spontaneous: an examination of deception across different modes of communication. Journal of the American Society for Information Science and Technology: JASIST, 63(1):208–216, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2003:MLW

[WBY03] Peiling Wang, Michael W. Berry, and Yiheng Yang. Mining longitudinal web queries: Trends and patterns. Journal of the American Society for Information Science and Technology: JASIST, 54(8):743–758, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2009:LMC

[WBY^{+}09] Howard D. White, Sebastian K. Boell, Hairong Yu, Mari Davis, Concepción S. Wilson, and Fletcher T. H. Cole. Libcitations: a measure for comparative assessment of book publications in the humanities and social sciences. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1083–1096, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2008:MUS

[WCC08] Ling-Ling Wu, Ya-Lan Chuang, and Pin-Yuen Chen. Motivation for using search engines: a two-factor model. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1829–1840, September 2008. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [WCC10b].

Westerman:2010:HAD

[WCC10a] S. J. Westerman, T. Cribbin, and J. Collins. Human assessments of document similarity. Journal of the American Society for Information Science and Technology: JASIST, 61(8): 1535–1542, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2010:ECW

[WCC10b] Ling-Ling Wu, Amber Chuang, and Pin-Yuen Chen. Erratum: Correction to Wu, L., Chuang, A., & Chen, P. (2008). Motivation for using search engines: a two factor model. Journal of American Society for Information Science and Technology, 59(11), 1829–1840. Journal of the American Society for Information Science and Technology: JASIST, 61(1):214–216, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [WCC08].

Wan-Chik:2013:IRI

[WCCS13] Rita Wan-Chik, Paul Clough, and Mark Sanderson. Investigating religious information searching through analysis of a search engine log. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2492–2506, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wenyin:2001:RMA

[WCLZ01] Liu Wenyin, Zheng Chen, Mingjing Li, and Hongjiang Zhang. Research: A media agent for automatically building a personalized semantic index of Web media objects. Journal of the American Society for Information Science and Technology: JASIST, 52(10):853–855, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Waltman:2012:LRD

[WCMK^{+}12] Ludo Waltman, Clara Calero-Medina, Joost Kosten, Ed C. M. Noyons, Robert J. W. Tijssen, Nees Jan van Eck, Thed N. van Leeuwen, Anthony F. J. van Raan, Martijn S. Visser, and Paul Wouters. The Leiden ranking 2011/2012: Data collection, indicators, and interpretation. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2419–2432, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wouters:2004:FCW

[WdV04] Paul Wouters and Repke de Vries. Formally citing the Web. Journal of the American Society for Information Science and Technology: JASIST, 55(14):1250–1260, December 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2012:RDK

[WEA^{+}12] Xin Wang, Sanda Erdelez, Carla Allen, Blake Anderson, Hongfei Cao, and Chi-Ren Shyu. Role of domain knowledge in developing user-centered medical-image indexing. Journal of the American Society for Information Science and Technology: JASIST, 63(2):225–241, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Weessies:2007:PDC

[Wee07] Kathleen W. Weessies. The publishing dynamics of catastrophic events. Journal of the American Society for Information Science and Technology: JASIST, 58(6):856–861, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Whittle:2007:DMS

[WEF^{+}07] Martin Whittle, Barry Eaglestone, Nigel Ford, Valerie J. Gillet, and Andrew Madden. Data mining of search engine logs. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2382–2400, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Weinberg:2010:COP

[Wei10] Bella Hass Weinberg. Citation, obliteration, and plagiarism, as discussed in ancient Jewish sources. Journal of the American Society for Information Science and Technology: JA-SIST, 61(11):2337–2364, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Westbrook:2007:BRB

[Wes07a] Lynn Westbrook. Book review: Human perspectives in the Internet society: Culture, psychology and gender. Journal of the American Society for Information Science and Technology: JASIST, 58(1):150–151, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Westbrook:2007:DIS

[Wes07b] Lynn Westbrook. Digital information support for domestic violence victims. Journal of the American Society for Information Science and Technology: JASIST, 58(3):420–432, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Westbrook:2009:IMI

[Wes09] Lynn Westbrook. Information myths and intimate partner violence: Sources, contexts, and consequences. Journal of the American Society for Information Science and Technology: JASIST, 60(4):826–836, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wilson:2002:ISM

[WFE^{+}02] T. D. Wilson, N. J. Ford, D. Ellis, A. E. Foster, and A. Spink. Information seeking and mediated searching: Part 2. Uncertainty and its correlates. Journal of the American Society for Information Science and Technology: JASIST, 53(9):704– 715, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2005:AEK

[WG05] Junxia Wang and Jiancheng Guan. The analysis and evaluation of knowledge efficiency in research groups. Journal of the American Society for Information Science and Technology: JASIST, 56(11):1217–1226, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Workman:2007:PED

[WG07] Michael Workman and John Gathegi. Punishment and ethics deterrents: a study of insider security contravention. Journal of the American Society for Information Science and Technology: JASIST, 58(2):212–222, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Willson:2010:ESR

[WG10] Rebekah Willson and Lisa M. Given. The effect of spelling and retrieval system familiarity on search behavior in online public access catalogs: a mixed methods study. Journal of the American Society for Information Science and Technology: JASIST, 61(12):2461–2476, December 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wallace:2009:NAD

[WGD09] Matthew L. Wallace, Yves Gingras, and Russell Duhon. A new approach for detecting scientific specialties from raw cocitation networks. Journal of the American Society for Information Science and Technology: JASIST, 60(2):240–246, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Willis:2012:ASM

[WGW12] Craig Willis, Jane Greenberg, and Hollie White. Analysis and synthesis of metadata goals for scientific data. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1505–1520, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Woelfer:2009:SHY

[WH09] Jill Palzkill Woelfer and David G. Hendry. Stabilizing homeless young people with information and place. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2300–2312, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2010:AAA

[WH10] Wendy Wang and Drew Hwang. Abstraction Assistant: an automatic text abstraction system. Journal of the American Society for Information Science and Technology: JASIST, 61(9):1790–1799, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2012:CPP

[WHC12] Ling-Ling Wu, Mu-Hsuan Huang, and Ching-Yi Chen. Citation patterns of the pre-web and web-prevalent environments: the moderating effects of domain knowledge. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2182–2194, November 2012. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Ano13c].

Wheeler:2003:BR

[Whe03] William J. Wheeler. Book review. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1167–1168, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2001:RAC

[Whi01] Howard D. White. Research: Authors as citers over time. Journal of the American Society for Information Science and Technology: JASIST, 52(2):87–108, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2002:LEL

[Whi02a] Howard D. White. Letter to the Editor: Library and information science (LIS) in aid of meta-analysis. Journal of the American Society for Information Science and Technology: JASIST, 53(4):323, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Whitley:2002:ASS

[Whi02b] Katherine M. Whitley. Analysis of Scifinder Scholar and Web of Science citation searches. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1210–1215, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Whitmire:2002:BRB

[Whi02c] Ethelene Whitmire. Book review: Identifying and analyzing user needs: a complete handbook and ready-to-use assessment workbook with disk. Journal of the American Society for Information Science and Technology: JASIST, 53(11): 966–967, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Whitmire:2002:DDU

[Whi02d] Ethelene Whitmire. Disciplinary differences and undergraduates’ information-seeking behavior. Journal of the American Society for Information Science and Technology: JA-SIST, 53(8):631–638, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2003:ACA

[Whi03a] Howard D. White. Author cocitation analysis and Pearson’s r. Journal of the American Society for Information Science and Technology: JASIST, 54(13):1250–1259, November 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2003:PNA

[Whi03b] Howard D. White. Pathfinder networks and author cocitation analysis: a remapping of paradigmatic information scientists. Journal of the American Society for Information Science and Technology: JASIST, 54(5):423–434, March 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2004:LER

[Whi04] Howard D. White. Letters to the Editor: Replies and a correction. Journal of the American Society for Information Science and Technology: JASIST, 55(9):843–844, July 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2006:BRB

[Whi06] Ryen W. White. Book review: Information retrieval design: Principles and options for information description, organization, display, and access in information retrieval databases, digital libraries, catalogs, and indexes. Journal of the American Society for Information Science and Technology: JASIST, 57(10):1412–1413, August 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2007:CBIa

[Whi07a] Howard D. White. Combining bibliometrics, information retrieval, and relevance theory, Part 1: First examples of a synthesis. Journal of the American Society for Information Science and Technology: JASIST, 58(4):536–559, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2007:CBIb

[Whi07b] Howard D. White. Combining bibliometrics, information retrieval, and relevance theory, Part 2: Some implications for information science. Journal of the American Society for Information Science and Technology: JASIST, 58(4):583– 605, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2010:SIT

[WHP10] Hao Wu, Jun He, and Yijian Pei. Scientific impact at the topic level: a case study in computational linguistics. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2274–2287, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2012:UAT

[WHTS12] Mingfang Wu, David Hawking, Andrew Turpin, and Falk Scholer. Using anchor text for homepage and topic distillation search tasks. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1235– 1255, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wijnhoven:2012:HIS

[Wij12] Fons Wijnhoven. The Hegelian inquiring system and a critical triangulation tool for the Internet information slave: a design science study. Journal of the American Society for Information Science and Technology: JASIST, 63(6):1168– 1182, June 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wilensky:2000:RDL

[Wil00] Robert Wilensky. Research: Digital library resources as a basis for collaborative work. Journal of the American Society for Information Science, 51(3):228–245, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See erratum [Ano08a].

Wildemuth:2002:EMS

[Wil02] Barbara M. Wildemuth. Effective methods for studying information seeking and use. Journal of the American Society for Information Science and Technology: JASIST, 53(14): 1218–1222, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wildemuth:2004:EDK

[Wil04] Barbara M. Wildemuth. The effects of domain knowledge on search tactic formulation. Journal of the American Society for Information Science and Technology: JASIST, 55(3):

246–258, February 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wilkerson:2005:BRB

[Wil05] Laura Michelle Wilkerson. Book review: Jeremiad Jottings. Journal of the American Society for Information Science and Technology: JASIST, 56(6):659–660, April 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wildemuth:2006:PSU

[Wil06a] Barbara M. Wildemuth. Perspectives on search user interfaces: Best practices and future visions: Evidence-based practice in search interface design. Journal of the American Society for Information Science and Technology: JASIST, 57(6):825–828, April 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Williams:2006:BRB

[Wil06b] James G. Williams. Book review: The Internet, the Web, and eBusiness: Formalizing applications for the real world. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1979–1980, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Williams:2006:MHA

[Wil06c] Robert V. Williams. In memoriam: Harold Abbott Wooster: a memorial essay. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1974–1976, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wilkinson:2010:BRB

[Wil10a] Leland Wilkinson. Book review: Picturing the Uncertain World: How to Understand, Communicate, and Control Uncertainty Through Graphical Display. Journal of the American Society for Information Science and Technology: JA-SIST, 61(4):853–854, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Williams:2010:HPL

[Wil10b] Robert V. Williams. Hans Peter Luhn and Herbert M. Ohlman: Their roles in the origins of keyword-in-context/ permutation automatic indexing. Journal of the American Society for Information Science and Technology: JASIST, 61(4):835–849, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Windsor:2002:LEI

[Win02] Donald A. Windsor. Letters to the Editor: Integer partitions result in skewed rank-frequency distributions. Journal of the American Society for Information Science and Technology: JASIST, 53(14):1276, December 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Winget:2008:AMS

[Win08a] Megan A. Winget. Annotations on musical scores by performing musicians: Collaborative models, interactive methods, and music digital library tool development. Journal of the American Society for Information Science and Technology: JASIST, 59(12):1878–1897, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Winter:2008:QAI

[Win08b] Carlos E. Winter. Quantitative analysis of indexed publications on seventeen model organisms in nine countries, from 1974 to 2006. Journal of the American Society for Information Science and Technology: JASIST, 59(10):1598–1607, August 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Winston:2009:BRB

[Win09] Mark D. Winston. Book review: Library Ethics. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2595–2596, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Winget:2011:AIS

[Win11] Megan A. Winget. Advances in information science: Videogame preservation and massively multiplayer online role-playing games: a review of the literature. Journal of the American Society for Information Science and Technology: JASIST, 62(10):1869–1883, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Widyantoro:2001:RLU

[WIY01] Dwi H. Widyantoro, Thomas R. Ioerger, and John Yen. Research: Learning user interest dynamics with a threedescriptor representation. Journal of the American Society for Information Science and Technology: JASIST, 52 (3):212–225, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2001:RWI

[WIZ01] Qishi Wu, S. Sitharama Iyengar, and Mengxia Zhu. Research: Web image retrieval using self-organizing feature map. Journal of the American Society for Information Science and Technology: JASIST, 52(10):868–875, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

West:2013:ALE

[WJD^{+}13] Jevin D. West, Michael C. Jensen, Ralph J. Dandrea, Gregory J. Gordon, and Carl T. Bergstrom. Author-level Eigenfactor metrics: Evaluating the influence of authors, institutions, and countries within the social science research network community. Journal of the American Society for Information Science and Technology: JASIST, 64(4):787–801, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2005:UTR

[WJR05] Ryen W. White, Joemon M. Jose, and Ian Ruthven. Using top-ranking sentences to facilitate effective information access. Journal of the American Society for Information Science and Technology: JASIST, 56(10):1113–1125, August 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2013:YSS

[WK13] Philip Fei Wu and Nikolaos Korfiatis. You scratch someone’s back and we’ll scratch yours: Collective reciprocity in social Q&A communities. Journal of the American Society for Information Science and Technology: JASIST, 64(10): 2069–2077, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Weeber:2001:RUC

[WKdJvdBV01] Marc Weeber, Henny Klein, Lolkje T. W. de Jong-van den Berg, and Rein Vos. Research: Using concepts in literaturebased discovery: Simulating Swanson’s Raynaud–fish oil and migraine–magnesium discoveries. Journal of the American Society for Information Science and Technology: JASIST, 52(7):548–557, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wacholder:2007:MQE

[WKK^{+}07] Nina Wacholder, Diane Kelly, Paul Kantor, Robert Rittman, Ying Sun, Bing Bai, Sharon Small, Boris Yamrom, and Tomek Strzalkowski. A model for quantitative evaluation of an end-to-end question-answering system. Journal of the American Society for Information Science and Technology: JASIST, 58(8):1082–1099, June 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2003:IIS

[WL03] Mei-Mei Wu and Ying-Hsang Liu. Intermediary’s information seeking, inquiring minds, and elicitation styles. Journal of the American Society for Information Science and Technology: JASIST, 54(12):1117–1133, October 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wacholder:2006:UPM

[WL06a] Nina Wacholder and Lu Liu. User preference: a measure of query-term quality. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1566– 1580, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2006:TIA

[WL06b] Yen-Chun Jim Wu and Hsing-Piao Liu. Technological innovation assessment of business-to-business electronic marketplaces. Journal of the American Society for Information Science and Technology: JASIST, 57(8):1093–1104, June 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2008:CPA

[WL08] Dian-Song Wu and Tyne Liang. Chinese pronominal anaphora resolution using lexical knowledge and entropybased weight. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2138–2145, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2009:SBP

[WL09] Ing-Long Wu and Han-Chang Lin. A strategy-based process for implementing knowledge management: an integrative view and empirical study. Journal of the American Society for Information Science and Technology: JASIST, 60 (4):789–802, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Walters:2011:BIC

[WL11a] William H. Walters and Anne C. Linvill. Bibliographic index coverage of open-access journals in six subject areas. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1614–1628, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2011:IIM

[WL11b] Mei-Mei Wu and Ying-Hsang Liu. On intermediaries’ inquiring minds, elicitation styles, and user satisfaction. Journal of the American Society for Information Science and Technology: JASIST, 62(12):2396–2403, December 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Willis:2013:RWO

[WL13] Craig Willis and Robert M. Losee. A random walk on an ontology: Using thesaurus structure for automatic subject indexing. Journal of the American Society for Information Science and Technology: JASIST, 64(7):1330–1344, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Ano13a].

Wu:2009:LDI

[WLC09] I-Chin Wu, Duen-Ren Liu, and Pei-Cheng Chang. Learning dynamic information needs: a collaborative topic variation inspection approach. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2430– 2451, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Westman:2011:DEM

[WLHO11] Stina Westman, Mari Laine-Hernandez, and Pirkko Oittinen. Development and evaluation of a multifaceted magazine image categorization model. Journal of the American Society for Information Science and Technology: JASIST, 62(2): 295–313, February 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wei:2010:IRL

[WLL10] Furu Wei, Wenjie Li, and Shixia Liu. iRANK: a ranklearn-combine framework for unsupervised ensemble ranking. Journal of the American Society for Information Science and Technology: JASIST, 61(6):1232–1243, June 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wei:2009:ATL

[WLLH09] Furu Wei, Wenjie Li, Qin Lu, and Yanxiang He. Applying two-level reinforcement ranking in query-oriented multidocument summarization. Journal of the American Society for Information Science and Technology: JASIST, 60(10): 2119–2131, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2006:IHA

[WM06] Shengli Wu and Sally McClean. Improving high accuracy retrieval by eliminating the uneven correlation effect in data fusion. Journal of the American Society for Information Science and Technology: JASIST, 57(14):1962–1973, December 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wan:2007:BML

[WM07] Raymond Wan and Alistair Moffat. Block merging for offline compression. Journal of the American Society for Information Science and Technology: JASIST, 58(1):3–14, January 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wilson:2009:EAS

[WmcsW09] Max L. Wilson, m. c. schraefel, and Ryen W. White. Evaluating advanced search interfaces using established information-seeking models. Journal of the American Society for Information Science and Technology: JASIST, 60 (7):1407–1422, July 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wolfram:2009:MCM

[WOB09] Dietmar Wolfram, Hope A. Olson, and Raina Bloom. Measuring consistency for multiple taggers using vector space modeling. Journal of the American Society for Information Science and Technology: JASIST, 60(10):1995–2003, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Woeginger:2009:GEI

[Woe09] Gerhard J. Woeginger. Generalizations of Egghe’s g-index. Journal of the American Society for Information Science and Technology: JASIST, 60(6):1267–1273, June 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Williams:2013:UBS

[WOE^{+}13] Christopher J. Williams, Michael O’Rourke, Sanford D. Eigenbrode, Ian O’Loughlin, and Stephen J. Crowley. Using bibliometrics to support the facilitation of cross-disciplinary communication. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1768–1779, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Workman:2007:PVT

[Wor07] Michael Workman. The proximal-virtual team continuum: a study of performance. Journal of the American Society for Information Science and Technology: JASIST, 58(6):794– 801, April 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Workman:2008:WTG

[Wor08] Michael Workman. Wisecrackers: a theory-grounded investigation of phishing and pretext social engineering threats to information security. Journal of the American Society for Information Science and Technology: JASIST, 59(4):662– 674, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Workman:2009:BRB

[Wor09] Michael Workman. Book review: Encyclopedia of Information and Ethics Security. Journal of the American Society for Information Science and Technology: JASIST, 60(8): 1723–1724, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warner:2004:TWW

[WP04] Dorothy Warner and J. Drew Procaccino. Toward wellness: Women seeking health information. Journal of the American Society for Information Science and Technology: JASIST, 55(8):709–730, June 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2005:SSA

[WP05] Tianhao Wu and William M. Pottenger. A semi-supervised active learning algorithm for information extraction from textual data. Journal of the American Society for Information Science and Technology: JASIST, 56(3):258–271, February 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2006:SIS

[WR06] Ryen W. White and Ian Ruthven. A study of interface support mechanisms for interactive information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 57(7):933–948, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Warwick:2009:CES

[WRB^{+}09] Claire Warwick, Jon Rimmer, Ann Blandford, Jeremy Gow, and George Buchanan. Cognitive economy and satisficing in information seeking: a longitudinal study of undergraduate information behavior. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2402– 2415, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wright:2010:BRB

[Wri10] Scott Wright. Book review: The Internet and Democratic Citizenship. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2374–2375, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Woodruff:2002:CUT

[WRM^{+}02] Allison Woodruff, Ruth Rosenholtz, Julie B. Morrison, Andrew Faulring, and Peter Pirolli. A comparison of the use of text summaries, plain thumbnails, and enhanced thumbnails for Web search tasks. Journal of the American Society for Information Science and Technology: JASIST, 53(2):172– 185, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2006:AIT

[WS06] Kieran J. White and Richard F. E. Sutcliffe. Applying incremental tree induction to retrieval from manuals and medical texts. Journal of the American Society for Information Science and Technology: JASIST, 57(5):588–600, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2011:BBC

[WS11] Kieran White and Richard F. E. Sutcliffe. Butcher, baker, or candlestick maker? Predicting occupations using predicate– argument relations. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1325– 1344, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wiggins:2012:IDF

[WS12] Andrea Wiggins and Steven Sawyer. Intellectual diversity and the faculty composition of iSchools. Journal of the American Society for Information Science and Technology: JASIST, 63(1):8–21, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Waltman:2013:CPB

[WS13] Ludo Waltman and Michael Schreiber. On the calculation of percentile-based bibliometric indicators. Journal of the American Society for Information Science and Technology: JASIST, 64(2):372–379, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wolfram:2001:BCV

[WSJS01] Dietmar Wolfram, Amanda Spink, Bernard J. Jansen, and Tefko Saracevic. Brief communication: Vox populi: the public searching of the web. Journal of the American Society for Information Science and Technology: JASIST, 52(12):1073– 1074, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wilson:2008:LCA

[WT08] Concepción S. Wilson and Carol Tenopir. Local citation analysis, publishing and reading patterns: Using multiple methods to evaluate faculty use of an academic library’s research collection. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1393–1408, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wilkinson:2010:SNS

[WT10] David Wilkinson and Mike Thelwall. Social network site changes over time: the case of MySpace. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2311–2323, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wilkinson:2012:TTT

[WT12] David Wilkinson and Mike Thelwall. Trending Twitter topics in English: an international comparison. Journal of the American Society for Information Science and Technology: JASIST, 63(8):1631–1646, August 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Weinberger:2008:OBE

[WTF08] Hadas Weinberger, Dov Te’eni, and Ariel J. Frank. Ontology-based evaluation of organizational memory. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1454–1468, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2006:EWM

[WTLC06] Jenq-Haur Wang, Jei-Wen Teng, Wen-Hsiang Lu, and Lee-Feng Chien. Exploiting the Web as the multilingual corpus for unknown query translation. Journal of the American Society for Information Science and Technology: JASIST, 57(5):660–670, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wong:2012:CQA

[WTP12] Wilson Wong, John Thangarajah, and Lin Padgham. Contextual question answering for the health domain. Journal of the American Society for Information Science and Technology: JASIST, 63(11):2313–2327, November 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wu:2010:IMA

[Wu10] Qiang Wu. The w-index: a measure to assess scientific impact by focusing on widely cited papers. Journal of the American Society for Information Science and Technology: JASIST, 61(3):609–614, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wainer:2013:WHC

[WV13] Jacques Wainer and Eduardo Valle. What happens to computer science research after it is published? Tracking CS research lines. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1104–1111, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wiesman:2004:IRM

[WvdHH04] F. Wiesman, H. J. van den Herik, and A. Hasman. Information retrieval by metabrowsing. Journal of the American Society for Information Science and Technology: JASIST, 55(7):565–578, May 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Waltman:2007:BCS

[WvE07] Ludo Waltman and Nees Jan van Eck. Brief communications: Some comments on the question whether cooccurrence data should be normalized. Journal of the American Society for Information Science and Technology: JA-SIST, 58(11):1701–1703, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Waltman:2010:RBE

[WvE10] Ludo Waltman and Nees Jan van Eck. The relation between Eigenfactor, audience factor, and influence weight. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1476–1486, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Waltman:2012:II

[WvE12a] Ludo Waltman and Nees Jan van Eck. The inconsistency of the h-index. Journal of the American Society for Information Science and Technology: JASIST, 63(2):406–415, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Waltman:2012:NMC

[WvE12b] Ludo Waltman and Nees Jan van Eck. A new methodology for constructing a publication-level classification system of science. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2378–2392, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Waltman:2012:UCD

[WvEvR12] Ludo Waltman, Nees Jan van Eck, and Anthony F. J. van Raan. Universality of citation distributions revisited. Journal of the American Society for Information Science and Technology: JASIST, 63(1):72–77, January 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Walters:2003:BIC

[WW03] William H. Walters and Esther I. Wilder. Bibliographic index coverage of a multidisciplinary field. Journal of the American Society for Information Science and Technology: JASIST, 54(14):1305–1312, December 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wilson:2013:CTM

[WW13] Mathew J. Wilson and Max L. Wilson. A comparison of techniques for measuring sensemaking and learning within participant-generated summaries. Journal of the American Society for Information Science and Technology: JASIST, 64(2):291–306, February 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

White:2004:DCR

[WWN04] Howard D. White, Barry Wellman, and Nancy Nazer. Does citation reflect social structure?: Longitudinal evidence from the “Globenet” interdisciplinary research group. Journal of the American Society for Information Science and Technology: JASIST, 55(2):111–126, January 15, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wolfram:2009:IWS

[WWZ09] Dietmar Wolfram, Peiling Wang, and Jin Zhang. Identifying Web search session patterns using cluster analysis: a comparison of three search environments. Journal of the American Society for Information Science and Technology: JASIST, 60(5):896–910, May 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2006:IAL

[WY06] Fu Lee Wang and Christopher C. Yang. The impact analysis of language differences on an automatic multilingual text summarization system. Journal of the American Society for Information Science and Technology: JASIST, 57(5):684– 696, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wang:2007:MWR

[WY07] Fu Lee Wang and Christopher C. Yang. Mining Web resources for enhancing information retrieval: Mining Web data for Chinese segmentation. Journal of the American Society for Information Science and Technology: JASIST, 58(12):1820–1837, October 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wolfram:2002:III

[WZ02] Dietmar Wolfram and Jin Zhang. An investigation of the influence of indexing exhaustivity and term distributions on a document space. Journal of the American Society for Information Science and Technology: JASIST, 53(11):943–952, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wolfram:2008:IIP

[WZ08] Dietmar Wolfram and Jin Zhang. The influence of indexing practices and weighting algorithms on document spaces. Journal of the American Society for Information Science and Technology: JASIST, 59(1):3–11, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wei:2009:EGM

[WZL09] Jiuchang Wei, Dingtao Zhao, and Liang Liang. Estimating the growth models of news stories on disasters. Journal of the American Society for Information Science and Technology: JASIST, 60(9):1741–1755, September 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xamena:2013:SRP

[XBM13] Eduardo Xamena, Nélida Beatriz Brignole, and Ana G. Maguitman. A study of relevance propagation in large topic ontologies. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2238–2255, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xu:2006:RJW

[XC06] Yunjie (Calvin) Xu and Zhiwei Chen. Relevance judgment: What do information users consider beyond topicality? Journal of the American Society for Information Science and Technology: JASIST, 57(7):961–973, May 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xie:2009:UHS

[XC09] Iris Xie and Colleen Cool. Understanding help seeking within the context of searching digital libraries. Journal of the American Society for Information Science and Technology: JASIST, 60(3):477–494, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xiao:2013:DCM

[XFCR13] Lu Xiao, Umer Farooq, John M. Carroll, and Mary Beth Rosson. The development of community members’ roles in partnership research projects: an empirical study. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2340–2353, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xia:2010:BCL

[Xia10] Jingfeng Xia. Brief communications: A longitudinal study of scholars attitudes and behaviors toward open-access journal publishing. Journal of the American Society for Information Science and Technology: JASIST, 61(3):615–624, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xie:2006:UHW

[Xie06] Hong (Iris) Xie. Understanding human–work domain interaction: Implications for the design of a corporate digital library. Journal of the American Society for Information Science and Technology: JASIST, 57(1):128–143, January 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xie:2011:EIL

[Xie11a] Bo Xie. Experimenting on the impact of learning methods and information presentation channels on older adults’ e-health literacy. Journal of the American Society for Information Science and Technology: JASIST, 62(9):1797–1807, September 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xie:2011:OAH

[Xie11b] Bo Xie. Older adults, e-health literacy, and collaborative learning: an experimental study. Journal of the American Society for Information Science and Technology: JASIST, 62(5):933–946, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xie:2010:TST

[XJ10] Iris Xie and Soohyung Joo. Transitions in search tactics during the Web-based search process. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2188–2205, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xu:2007:DIIb

[XL07] Yunjie Xu and Chengliang Liu. The dynamics of interactive information retrieval, Part II: an empirical study from the activity theory perspective. Journal of the American Society for Information Science and Technology: JASIST, 58 (7):987–998, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xu:2013:STS

[XMCM13] Chen Xu, Benjiang Ma, Xiaohong Chen, and Feicheng Ma. Social tagging in the scholarly world. Journal of the American Society for Information Science and Technology: JA-SIST, 64(10):2045–2057, October 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xie:2007:LMK

[XR07] Ying Xie and Vijay V. Raghavan. Language-modeling kernel based approach for information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2353–2365, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xu:2006:WWY

[XTY06] Yunjie (Calvin) Xu, Bernard C. Y. Tan, and Li Yang. Who will you ask? An empirical study of interpersonal task information seeking. Journal of the American Society for Information Science and Technology: JASIST, 57(12):1666–1677, October 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See erratum [Tan07a].

Xu:2007:DIIa

[Xu07a] Yunjie Xu. The dynamics of interactive information retrieval behavior, Part I: an activity theory perspective. Journal of the American Society for Information Science and Technology: JASIST, 58(7):958–970, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xu:2007:RJE

[Xu07b] Yunjie Xu. Relevance judgment in epistemic and hedonic information searches. Journal of the American Society for Information Science and Technology: JASIST, 58(2):179– 189, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xie:2002:SDL

[XW02] Hong (Iris) Xie and Dietmar Wolfram. State digital library usability: Contributing organizational factors. Journal of the American Society for Information Science and Technology: JASIST, 53(13):1085–1097, November 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xu:2008:OER

[XW08] Yunjie Xu and Dong Wang. Order effect in relevance judgment. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1264–1275, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xu:2008:NTI

[XY08] Yunjie Xu and Hainan Yin. Novelty and topicality in interactive information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 59(2): 201–215, January 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Xiao:2013:PAG

[XZW13] Chunjing Xiao, Fan Zhou, and Yue Wu. Predicting audience gender in online content-sharing social networks. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1284–1297, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yager:2006:KTP

[Yag06] Ronald R. Yager. Knowledge trees and protoforms in question-answering systems. Journal of the American Society for Information Science and Technology: JASIST, 57 (4):550–563, February 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yakel:2005:BRB

[Yak05] Elizabeth Yakel. Book review: Introduction to reference work in the digital age. Journal of the American Society for Information Science and Technology: JASIST, 56(11):

1237–1238, September 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2001:ISA

[Yan01] Shu Ching Yang. An interpretive and situated approach to an evaluation of Perseus digital libraries. Journal of the American Society for Information Science and Technology: JASIST, 52(14):1210–1223, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2010:BRB

[Yan10] Christopher C. Yang. Book review: Search Engines Information Retrieval in Practice. Journal of the American Society for Information Science and Technology: JASIST, 61 (2):430, February 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yuan:2010:EIS

[YB10a] Xiaojun Yuan and Nicholas J. Belkin. Evaluating an integrated system supporting multiple information-seeking strategies. Journal of the American Society for Information Science and Technology: JASIST, 61(10):1987–2010, October 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yuan:2010:IIR

[YB10b] Xiaojun Yuan and Nicholas J. Belkin. Investigating information retrieval support techniques for different informationseeking strategies. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1543– 1563, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yi:2006:USB

[YBC^{+}06] Kwan Yi, Jamshid Beheshti, Charles Cole, John E. Leide, and Andrew Large. User search behavior of domain-specific information retrieval systems: an analysis of the query logs from PsycINFO and ABC-Clio’s historical abstracts/ America: History and life. Journal of the American Society for Information Science and Technology: JASIST, 57 (9):1208–1220, July 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2002:PAC

[YC02] Christopher C. Yang and Alan Chung. A personal agent for Chinese financial news on the Web. Journal of the American Society for Information Science and Technology: JASIST, 53(2):186–196, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yi:2010:RSS

[YC10] Kwan Yi and Lois Mai Chan. Revisiting the syntactical and structural analysis of Library of Congress Subject Headings for the digital environment. Journal of the American Society for Information Science and Technology: JASIST, 61 (4):677–687, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yoon:2011:UIN

[YC11] JungWon Yoon and EunKyung Chung. Understanding image needs in daily life by analyzing questions in a social Q&A site. Journal of the American Society for Information Science and Technology: JASIST, 62(11):2201–2213, November 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yuan:2010:IAA

[YCE10] Y. Connie Yuan, Inga Carboni, and Kate Ehrlich. The impact of awareness and accessibility on expertise retrieval: a multilevel network perspective. Journal of the American Society for Information Science and Technology: JASIST, 61 (4):700–714, April 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yeganova:2009:HIP

[YCKW09] Lana Yeganova, Donald C. Comeau, Won Kim, and W. John Wilbur. How to interpret PubMed queries and why it matters. Journal of the American Society for Information Science and Technology: JASIST, 60(2):264–274, February 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yan:2009:ACM

[YD09] Erjia Yan and Ying Ding. Applying centrality measures to impact analysis: a coauthorship network analysis. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2107–2118, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yan:2010:WCI

[YD10] Erjia Yan and Ying Ding. Weighted citation: an indicator of an article’s prestige. Journal of the American Society for Information Science and Technology: JASIST, 61(8): 1635–1643, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yan:2012:SNS

[YD12] Erjia Yan and Ying Ding. Scholarly network similarities: How bibliographic coupling networks, citation networks, cocitation networks, topical networks, coauthorship networks, and coword networks relate to each other. Journal of the American Society for Information Science and Technology: JASIST, 63(7):1313–1326, July 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yan:2013:EBT

[YD13] Yalan Yan and Robert M. Davison. Exploring behavioral transfer from knowledge seeking to knowledge contributing: the mediating role of intrinsic motivation. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1144–1157, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yan:2011:RIM

[YDS11] Erjia Yan, Ying Ding, and Cassidy R. Sugimoto. P -Rank: an indicator measuring prestige in heterogeneous scholarly networks. Journal of the American Society for Information Science and Technology: JASIST, 62(3):467–477, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ye:2011:BCT

[Ye11a] Fred Y. Ye. Brief communication: A theoretical approach to the unification of informetric models by wave-heat equations. Journal of the American Society for Information Science and Technology: JASIST, 62(6):1208–1211, June 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ye:2011:BCU

[Ye11b] Fred Y. Ye. Brief communication: A unification of three models for the h-index. Journal of the American Society for Information Science and Technology: JASIST, 62(1): 205–207, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yeo:2011:BRB

[Yeo11] ShinJoung Yeo. Book review: Access Controlled: The Shaping of Power, Rights, and Rule in Cyberspace; Networks and States: The Global Politics of Internet Governance. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1647–1649, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yao:2008:ACR

[YEV08] Haining Yao, Letha H. Etzkorn, and Shamsnaz Virani. Automated classification and retrieval of reusable software components. Journal of the American Society for Information Science and Technology: JASIST, 59(4):613–627, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yitzhaki:2004:AUI

[YH04] Moshe Yitzhaki and Gloria Hammershlag. Accessibility and use of information sources among computer scientists and software engineers in Israel: Academy versus industry. Journal of the American Society for Information Science and Technology: JASIST, 55(9):832–842, July 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ye:2012:MMA

[YHHL12] Zheng Ye, Jimmy Xiangji Huang, Ben He, and Hongfei Lin. Mining a multilingual association dictionary from Wikipedia for cross-language information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2474–2487, December 2012. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ye:2011:FGQ

[YHL11] Zheng Ye, Jimmy Xiangji Huang, and Hongfei Lin. Finding a good query-related topic for boosting pseudo-relevance feedback. Journal of the American Society for Information Science and Technology: JASIST, 62(4):748–760, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ye:2013:UTP

[YHWL13] Zheng Ye, Ben He, Lifeng Wang, and Tiejian Luo. Utilizing term proximity for blog post retrieval. Journal of the American Society for Information Science and Technology: JA-SIST, 64(11):2278–2298, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yi:2008:IOR

[Yi08] Myongho Yi. Information organization and retrieval using a topic maps-based ontology: Results of a task-based evaluation. Journal of the American Society for Information Science and Technology: JASIST, 59(12):1898–1911, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yi:2010:SSA

[Yi10] Kwan Yi. A semantic similarity approach to predicting Library of Congress subject headings for social tags. Journal of the American Society for Information Science and Technology: JASIST, 61(8):1658–1672, August 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yi:2012:HCI

[Yi12] Kwan Yi. Harnessing collective intelligence in social tagging using Delicious. Journal of the American Society for Information Science and Technology: JASIST, 63(12):2488–2502, December 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yakel:2005:ADE

[YK05] Elizabeth Yakel and Jihyun Kim. Adoption and diffusion of Encoded Archival Description. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1427–1437, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2012:AIS

[YKCL12] Ming Yang, Melody Kiang, Hsinchun Chen, and Yijun Li. Artificial immune system for illicit content identification in social media. Journal of the American Society for Information Science and Technology: JASIST, 63(2):256–269, February 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2007:GDD

[YKL07] Haixuan Yang, Irwin King, and Michael R. Lyu. The generalized dependency degree between attributes. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2280–2294, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yoshikane:2003:LES

[YKT03a] Fuyuki Yoshikane, Kyo Kageura, and Keita Tsuji. Letters to the Editor: The sample size dependency of statistical measures and synchronic potentiality in informetrics. Some comments on some comments by Professor Burrell. Journal of the American Society for Information Science and Technology: JASIST, 54(11):1075–1076, September 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Bur03a].

Yoshikane:2003:MCA

[YKT03b] Fuyuki Yoshikane, Kyo Kageura, and Keita Tsuji. A method for the comparative analysis of concentration of author productivity, giving consideration to the effect of sample size dependency of statistical measures. Journal of the American Society for Information Science and Technology: JA-SIST, 54(6):521–528, April 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2003:ACE

[YL03a] Christopher C. Yang and Kar Wing Li. Automatic construction of English/ Chinese parallel corpora. Journal of the American Society for Information Science and Technology: JASIST, 54(8):730–742, June 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2003:AGE

[YL03b] Christopher C. Yang and Johnny Luk. Automatic generation of English/ Chinese thesaurus based on a parallel corpus in laws. Journal of the American Society for Information Science and Technology: JASIST, 54(7):671–682, May 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2005:HMB

[YL05] Christopher C. Yang and K. W. Li. A heuristic method based on a statistical approach for Chinese text segmentation. Journal of the American Society for Information Science and Technology: JASIST, 56(13):1438–1447, November 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2006:IST

[YL06] Christopher C. Yang and Wai Lam. Introduction to the special topic section on multilingual information systems. Journal of the American Society for Information Science and Technology: JASIST, 57(5):629–631, March 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2009:WST

[YL09] Christopher C. Yang and Nan Liu. Web site topic-hierarchy generation based on link structure. Journal of the American Society for Information Science and Technology: JASIST, 60(3):495–508, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yoon:2006:EPC

[YLL06] Yongwook Yoon, Changki Lee, and Gary Geunbae Lee. An effective procedure for constructing a hierarchical text classification system. Journal of the American Society for Information Science and Technology: JASIST, 57(3):431–442, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2010:RKD

[YLW10] Christopher C. Yang, Jianfeng Lin, and Chih-Ping Wei. Retaining knowledge for document management: Categorytree integration by exploiting category relationships and hierarchical structures. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1313– 1331, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2010:DBG

[YLZ10] Yuhang Yang, Qin Lu, and Tiejun Zhao. A delimiter-based general approach for Chinese term extraction. Journal of the American Society for Information Science and Technology: JASIST, 61(1):111–125, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yee:2009:VDP

[YNF09] Wai Gen Yee, Linh Thai Nguyen, and Ophir Frieder. A view of the data on P2P file-sharing systems. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2132–2141, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yoo:2008:E

[Yoo08] Eun-Young Yoo. Erratum. Journal of the American Society for Information Science and Technology: JASIST, 59(7): 1191, May 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yoon:2010:UQU

[Yoo10] JungWon Yoon. Utilizing quantitative users’ reactions to represent affective meanings of an image. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1345–1359, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yoon:2012:CSA

[Yoo12] Kyunghye Yoon. Conceptual syntagmatic associations in user tagging. Journal of the American Society for Information Science and Technology: JASIST, 63(5):923–935, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yu:2002:EMD

[YR02] Byeong-Min Yu and Seak-Zoon Roh. The effects of menu design on information-seeking performance and user’s attitude on the World Wide Web. Journal of the American Society for Information Science and Technology: JASIST, 53(11): 923–933, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yoo:2008:UMA

[YR08] Eun-Young Yoo and Louise S. Robbins. Understanding middle-aged women’s health information seeking on the web: a theoretical approach. Journal of the American Society for Information Science and Technology: JASIST, 59(4):577– 590, February 15, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yao:2007:PUC

[YRW07] Mike Z. Yao, Ronald E. Rice, and Kier Wallis. Predicting user concerns about online privacy. Journal of the American Society for Information Science and Technology: JASIST, 58(5):710–722, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yuan:2011:IBI

[YRXS11] Y. Connie Yuan, Laura N. Rickard, Ling Xia, and Clifford Scherer. The interplay between interpersonal and electronic resources in knowledge seeking among co-located and distributed employees. Journal of the American Society for Information Science and Technology: JASIST, 62(3):535–549, March 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yan:2011:IIE

[YS11] Erjia Yan and Cassidy R. Sugimoto. Institutional interactions: Exploring social, cognitive, and geographic relationships between institutions as demonstrated through citation networks. Journal of the American Society for Information Science and Technology: JASIST, 62(8):1498–1514, August 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ye:2008:IIS

[YSD^{+}08] Chen Ye, DongBack Seo, Kevin C. Desouza, Sridhar Papagari Sangareddy, and Sanjeev Jha. Influences of IT substitutes and user experience on post-adoption user switching: an empirical investigation. Journal of the American Society for Information Science and Technology: JASIST, 59 (13):2115–2132, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Tsay:2004:LGJ

[yT04] Ming yueh Tsay. Literature growth, journal characteristics, and author productivity in subject indexing, 1977 to 2000. Journal of the American Society for Information Science and Technology: JASIST, 55(1):64–73, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yang:2008:HSL

[YW08] Christopher C. Yang and Fu Lee Wang. Hierarchical summarization of large documents. Journal of the American Society for Information Science and Technology: JASIST, 59 (6):887–902, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yeh:2008:OBS

[YWC08] Jui-Feng Yeh, Chung-Hsien Wu, and Ming-Jun Chen. Ontology-based speech act identification in a bilingual dialog system using partial pattern trees. Journal of the American Society for Information Science and Technology: JASIST, 59(5):684–694, March 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yates:2010:FAS

[YWM10] Dave Yates, Christian Wagner, and Ann Majchrzak. Factors affecting shapers of organizational wikis. Journal of the American Society for Information Science and Technology: JASIST, 61(3):543–554, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Ye:2013:THU

[YYL13] Fred Y. Ye, Susan S. Yu, and Loet Leydesdorff. The Triple Helix of university–industry–government relations at the country level and its dynamic evolution under the pressures of globalization. Journal of the American Society for Information Science and Technology: JASIST, 64(11):2317–2325, November 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Yuan:2013:UDI

[YZLC13] Y. Connie Yuan, Xuan Zhao, Qinying Liao, and Changyan Chi. The use of different information and communication technologies to support knowledge sharing in organizations: From e-mail to micro-blogging. Journal of the American Society for Information Science and Technology: JASIST, 64 (8):1659–1670, August 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhu:2013:DCU

[ZA13] W. Z. Zhu and R. B. Allen. Document clustering using the LSI subspace signature model. Journal of the American Society for Information Science and Technology: JASIST, 64 (4):844–860, April 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zach:2005:WEE

[Zac05] Lisl Zach. When is “enough” enough? Modeling the information-seeking and stopping behavior of senior arts administrators. Journal of the American Society for Information Science and Technology: JASIST, 56(1):23–35, January 1, 2005. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2009:VHS

[ZATH09] Jin Zhang, Lu An, Tao Tang, and Yi Hong. Visual health subject directory analysis based on users’ traversal activities. Journal of the American Society for Information Science and Technology: JASIST, 60(10):1977–1994, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2007:UIR

[ZB07] Ping Zhang and Robert I. Benjamin. Understanding information related fields: a conceptual framework. Journal of the American Society for Information Science and Technology: JASIST, 58(13):1934–1947, November 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zheng:2009:ECR

[ZBK09] Hai-Tao Zheng, Charles Borchert, and Hong-Gee Kim. Exploiting corpus-related ontologies for conceptualizing document corpora. Journal of the American Society for Information Science and Technology: JASIST, 60(11):2287–2299, November 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2001:AEU

[ZC01] Xiangmin Zhang and Mark Chignell. Assessment of the effects of user characteristics on mental models of information retrieval systems. Journal of the American Society for Information Science and Technology: JASIST, 52(6):445–459, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zeng:2004:TIE

[ZC04] Marcia Lei Zeng and Lois Mai Chan. Trends and issues in establishing interoperability among knowledge organization systems. Journal of the American Society for Information Science and Technology: JASIST, 55(5):377–395, March 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhou:2008:OSP

[ZC08] Lina Zhou and Pimwadee Chaovalit. Ontology-supported polarity mining. Journal of the American Society for Information Science and Technology: JASIST, 59(1):98–110, January 1, 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2013:CCA

[ZDM13] Guo Zhang, Ying Ding, and Staša Milojević. Citation content analysis (CCA): a framework for syntactic and semantic analysis of citation content. Journal of the American Society for Information Science and Technology: JASIST, 64(7): 1490–1503, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2001:RSU

[Zha01] Yin Zhang. Research: Scholarly use of Internet-based electronic resources. Journal of the American Society for Information Science and Technology: JASIST, 52(8):628–654, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2002:CRJ

[Zha02] Xiangmin Zhang. Collaborative relevance judgment: a group consensus method for evaluating user search performance. Journal of the American Society for Information Science and Technology: JASIST, 53(3):220–231, ???? 2002. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2004:LER

[Zha04] Xiangmin Zhang. Letters to the Editors: Rejoinder: Quality of information. Journal of the American Society for Information Science and Technology: JASIST, 55(1):92, January 1, 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2008:USM

[Zha08] Yan Zhang. Undergraduate students’ mental models of the Web as an information retrieval system. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2087–2098, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2010:BCR

[Zha10a] Chun-Ting Zhang. Brief communications: Relationship of the h-index, g-index, and e-index. Journal of the American Society for Information Science and Technology: JASIST, 61(3):625–628, March 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2010:DEP

[Zha10b] Yan Zhang. Dimensions and elements of people’s mental models of an information-rich Web space. Journal of the American Society for Information Science and Technology: JASIST, 61(11):2206–2218, November 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2010:DHM

[Zha10c] Ying Zhang. Developing a holistic model for digital library evaluation. Journal of the American Society for Information Science and Technology: JASIST, 61(1):88–110, January 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2012:GSJ

[Zha12] Lei Zhang. Grasping the structure of journal articles: Utilizing the functions of information units. Journal of the American Society for Information Science and Technology: JA-SIST, 63(3):469–480, March 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2013:TLM

[Zha13] Yan Zhang. Toward a layered model of context for health information searching: an analysis of consumer-generated questions. Journal of the American Society for Information Science and Technology: JASIST, 64(6):1158–1172, June 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zachary:2001:RIT

[ZI01] John Zachary and S. S. Iyengar. Research: Information theoretic similarity measures for content based image retrieval. Journal of the American Society for Information Science and Technology: JASIST, 52(10):856–867, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zachary:2001:RCB

[ZIB01] John Zachary, S. S. Iyengar, and Jacob Barhen. Research: Content based image retrieval and information theory: a general approach. Journal of the American Society for Information Science and Technology: JASIST, 52(10):840–852, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zins:2007:CSI

[Zin07a] Chaim Zins. Classification schemes of Information Science: Twenty-eight scholars map the field. Journal of the American Society for Information Science and Technology: JA-SIST, 58(5):645–672, March 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zins:2007:CIS

[Zin07b] Chaim Zins. Conceptions of information science. Journal of the American Society for Information Science and Technology: JASIST, 58(3):335–350, February 1, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zins:2007:CAD

[Zin07c] Chaim Zins. Conceptual approaches for defining data, information, and knowledge. Journal of the American Society for Information Science and Technology: JASIST, 58(4):479– 493, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zins:2007:KMI

[Zin07d] Chaim Zins. Knowledge map of information science. Journal of the American Society for Information Science and Technology: JASIST, 58(4):526–535, February 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zins:2009:BRB

[Zin09] Chaim Zins. Book review: Information Science 101. Journal of the American Society for Information Science and Technology: JASIST, 60(12):2600–2602, December 2009. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2009:IFP

[ZJS09] Ying Zhang, Bernard J. Jansen, and Amanda Spink. Identification of factors predicting clickthrough in Web searching using neural network analysis. Journal of the American Society for Information Science and Technology: JASIST, 60 (3):557–570, March 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zilberman:2013:AGM

[ZKSE13] Polina Zilberman, Gilad Katz, Asaf Shabtai, and Yuval Elovici. Analyzing group E-mail exchange to detect data leakage. Journal of the American Society for Information Science and Technology: JASIST, 64(9):1780–1790, September 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhou:2007:CBC

[ZL07a] Ping Zhou and Loet Leydesdorff. A comparison between the China Scientific and Technical Papers and Citations Database and the Science Citation Index in terms of journal hierarchies and interjournal citation relations. Journal of the American Society for Information Science and Technology: JASIST, 58(2):223–236, January 15, 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhu:2007:BRBc

[ZL07b] Xiaomin Zhu and Jianxin Liao. Book review: Fuzzy logic and the semantic Web. Journal of the American Society for Information Science and Technology: JASIST, 58(14):2409– 2410, December 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhu:2007:BRBb

[ZL07c] Xiaomin Zhu and Jianxin Liao. Book review: The IMS: IP multimedia concepts and services in the mobile domain. Journal of the American Society for Information Science and Technology: JASIST, 58(11):1705–1706, September 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhu:2007:BRBa

[ZL07d] Xiaomin Zhu and Jianxin Liao. Book review: Web usability: a user-centered design approach. Journal of the American Society for Information Science and Technology: JASIST, 58(7):1066–1067, May 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2008:UCF

[ZL08] Ying Zhang and Yuelin Li. A user-centered functional metadata evaluation of moving image collections. Journal of the American Society for Information Science and Technology: JASIST, 59(8):1331–1346, June 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zitt:2011:HCW

[ZLB11] Michel Zitt, Alain Lelu, and Elise Bassecoulard. Hybrid citation-word representations in science mapping: Portolan charts of research fields? Journal of the American Society for Information Science and Technology: JASIST, 62(1):19– 39, January 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zheng:2006:FAI

[ZLCH06] Rong Zheng, Jiexun Li, Hsinchun Chen, and Zan Huang. A framework for authorship identification of online messages: Writing-style features and classification techniques. Journal of the American Society for Information Science and Technology: JASIST, 57(3):378–393, February 1, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2011:QSI

[ZLXW11] Cheng Zhang, Xu Liu, Yunjie (Calvin) Xu, and Youwei Wang. Quality-structure index: a new metric to measure scientific journal influence. Journal of the American Society for Information Science and Technology: JASIST, 62 (4):643–653, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2010:LAB

[ZQI^{+}10] Haizheng Zhang, Baojun Qiu, Kristinka Ivanova, C. Lee Giles, Henry C. Foley, and John Yen. Locality and attachedness-based temporal social network growth dynamics analysis: a case study of evolving nanotechnology scientific collaboration networks. Journal of the American Society for Information Science and Technology: JASIST, 61 (5):964–977, May 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2011:DTC

[ZRG11] Lin Zhang, Ronald Rousseau, and Wolfgang Glänzel. Document-type country profiles. Journal of the American Society for Information Science and Technology: JASIST, 62(7):1403–1411, July 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhao:2007:CCA

[ZS07] Dangzhi Zhao and Andreas Strotmann. Can citation analysis of Web publications better detect research fronts? Journal of the American Society for Information Science and Technology: JASIST, 58(9):1285–1302, July 2007. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhao:2008:ERA

[ZS08a] Dangzhi Zhao and Andreas Strotmann. Evolution of research activities and intellectual influences in information science 1996–2005: Introducing author bibliographiccoupling analysis. Journal of the American Society for Information Science and Technology: JASIST, 59(13):2070–2086, November 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhao:2008:ISD

[ZS08b] Dangzhi Zhao and Andreas Strotmann. Information science during the first decade of the web: an enriched author cocitation analysis. Journal of the American Society for Information Science and Technology: JASIST, 59(6):916–937, April 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zitt:2008:BCM

[ZS08c] Michel Zitt and Henry Small. Brief communications: Modifying the journal impact factor by fractional citation weighting: the audience factor. Journal of the American Society for Information Science and Technology: JASIST, 59(11):1856– 1860, September 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2009:CDT

[ZS09] Ping Zhang and Heshan Sun. The complexity of different types of attitudes in initial and continued ICT use. Journal of the American Society for Information Science and Technology: JASIST, 60(10):2048–2063, October 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhao:2011:CFL

[ZS11a] Dangzhi Zhao and Andreas Strotmann. Counting first, last, or all authors in citation analysis: a comprehensive comparison in the highly collaborative stem cell research field. Journal of the American Society for Information Science and Technology: JASIST, 62(4):654–676, April 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zins:2011:MKC

[ZS11b] Chaim Zins and Plácida L. V. A. C. Santos. Mapping the knowledge covered by library classification systems. Journal of the American Society for Information Science and Technology: JASIST, 62(5):877–901, May 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhou:2010:CSC

[ZSL10] Ping Zhou, Xinning Su, and Loet Leydesdorff. A comparative study on communication structures of Chinese journals in the social sciences. Journal of the American Society for Information Science and Technology: JASIST, 61(7):1360– 1376, July 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhu:2009:IMW

[ZSR09] Jianhan Zhu, Dawei Song, and Stefan Rüger. Integrating multiple windows and document features for expert finding. Journal of the American Society for Information Science and Technology: JASIST, 60(4):694–715, April 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2013:BCW

[ZTG13] Lin Zhang, Bart Thijs, and Wolfgang Glänzel. Brief communication: What does scientometrics share with other “metrics” sciences? Journal of the American Society for Information Science and Technology: JASIST, 64(7):1515–1518, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhao:2012:DIM

[ZTY12] Star X. Zhao, Alice M. Tan, and Fred Y. Ye. Distributive h-indices for measuring multilevel impact. Journal of the American Society for Information Science and Technology: JASIST, 63(10):2074–2086, October 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zuccala:2006:ACA

[Zuc06a] Alesia Zuccala. Author Cocitation Analysis is to intellectual structure as Web Colink Analysis is to … ? Journal of the American Society for Information Science and Technology: JASIST, 57(11):1487–1502, September 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zuccala:2006:MIC

[Zuc06b] Alesia Zuccala. Modeling the invisible college. Journal of the American Society for Information Science and Technology: JASIST, 57(2):152–168, January 15, 2006. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zuccala:2011:BRH

[ZvL11] Alesia Zuccala and Thed van Leeuwen. Book reviews in humanities research evaluations. Journal of the American Society for Information Science and Technology: JASIST, 62(10):1979–1991, October 2011. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2001:RVT

[ZW01] Jin Zhang and Dietmar Wolfram. Research: Visualization of term discrimination analysis. Journal of the American Society for Information Science and Technology: JASIST, 52(8):615–627, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2008:VHS

[ZWW^{+}08] Jin Zhang, Dietmar Wolfram, Peiling Wang, Yi Hong, and Rick Gillis. Visualization of health-subject analysis based on query term co-occurrences. Journal of the American Society for Information Science and Technology: JASIST, 59(12): 1933–1947, October 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2009:AQK

[ZWW09] Jin Zhang, Dietmar Wolfram, and Peiling Wang. Analysis of query keywords of sports-related queries using visualization and clustering. Journal of the American Society for Information Science and Technology: JASIST, 60(8):1550–1571, August 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhao:2013:PLL

[ZY13] Star X. Zhao and Fred Y. Ye. Power-law link strength distribution in paper cocitation networks. Journal of the American Society for Information Science and Technology: JA-SIST, 64(7):1480–1489, July 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2013:ICI

[ZYH13] Ping Zhang, Jasy Liew Suet Yan, and Katie DeVries Hassman. The intellectual characteristics of the information field: Heritage and substance. Journal of the American Society for Information Science and Technology: JASIST, 64(12):2468– 2491, December 2013. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhou:2003:NTF

[ZZ03] Lina Zhou and Dongsong Zhang. NLPIR: a theoretical framework for applying natural language processing to information retrieval. Journal of the American Society for Information Science and Technology: JASIST, 54(2):115–123, January 15, 2003. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhuge:2010:TCS

[ZZ10] Hai Zhuge and Junsheng Zhang. Topological centrality and its e-Science applications. Journal of the American Society for Information Science and Technology: JASIST, 61 (9):1824–1841, September 2010. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2009:SAC

[ZZL^{+}09] Changli Zhang, Daniel Zeng, Jiexun Li, Fei-Yue Wang, and Wanli Zuo. Sentiment analysis of Chinese documents: From sentence to document level. Journal of the American Society for Information Science and Technology: JASIST, 60(12): 2474–2487, December 2009. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2012:BCE

[ZZL12] Zili Zhang, Ziqiong Zhang, and Rob Law. Brief communication: Editorial responsiveness, journal quality, and total review time: an empirical analysis. Journal of the American Society for Information Science and Technology: JASIST, 63(5):1054–1058, May 2012. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Zhang:2004:UIS

[ZZZR04] Dongming Zhang, Caroline Zambrowicz, Hong Zhou, and Nancy K. Roderer. User information-seeking behavior in a medical Web portal environment: a preliminary study. Journal of the American Society for Information Science and Technology: JASIST, 55(8):670–684, June 2004. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).